A Bibliography of Publications in the Journals Computer Graphics and Image Processing (1972–1982), Computer Vision, Graphics, and Image Processing (1983–1990), Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing (1991–1994), Computer Vision, Graphics, and Image Processing. Image Understanding (1991–1994), Computer Vision

and Image Understanding: CVIU (1995–date) Graphical Models and Image Processing: GMIP (1995–1999) and Graphical Models (2000–date)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org, beebe@ieee.org (Internet)
                  WWW URL: http://www.math.utah.edu/~beebe/

10 February 2022 Version 1.72

Title word cross-reference

(18, 6) [MW00]. + [BCF06]. 1 [AOR94, BDL^{+}06, DLP13, Her98, TSK94]. 101 [FFFP07]. 113 [MBMC11]. 12 [PK99]. 16 [MMS97]. 2 [AKL93, Ano01s, AS08b, AM97, BBC00, BL94, Bd96, BRdBS99, BZ99, BCF06, BS04b, CYW04a, CYW04b, CL00b, CFM^{+}13, CC96, DB03, DAM12, DBB13, FPC^{+}08, FAB97, FKL^{+}98, FST94, GSPL10, GH90, GK04, HB98a, HB98b, IAP^{+}11, JDP97, JC98, JM92, KMB97, KS91b, KADS02, KK95, KS04a, KM03, KMN11, KNO^{+}09, Lau97, LP91, LPR93, LST13, LT90b, LS12, Luc01, MST85, Mil09, MBMC11, NT10, Neg12, NKPT13, NSEA13, Nur86, OJRT08, PPK93, RH85, Sal90, Ste01, TTIM96, TH04, VPAM12, WCZ02, YK95]. 2.5 [MCB13, SRHC13, ZP11]. 2008 [SCOG09]. 3 [AHRW87, ACF00, ACG^{+}09, ÁB13, AS08b, AM97, ARARCE11, ACDB12, BM99, BIP00, BI10, BI11, BCA98, Bar05, BLT05, BAM87, BT05, BD94a, BM95, BJ86, BR95, BY12, Bd96, BRdBS99, BZ99, BCF06, BGK95, BF05, BS00a, BSB87, BH95, COW98, CGH08, CM12, CCH91, CK11, CVB09, CS89, CJ93, CSDC96, CS98, CK00, CYNO11, CC11, CA86a, CLCO13, CL00b, CFM^{+}13, CP91, CCA92, CC96, CJ82, CG04, CS00, CPS10, DT96b, Dam08, DWB11, Dan97, DF01, DPR92, DN82, DSY10, EK98, ES04, EM96, FWH13, FBF08, FF09, FHMB84, FRL^{+}98, FDMA97, FAB97, FKL^{+}98, Fly92, FL96, GA91, GSPL10, GHMT09, Gol13, GR87a, GR85, GSV05, GW07, Gui98, Gui99, GPC^{+}10, GB93, GSK02, HFKN97, HASS10, HRS02, HR99, Hen98, HHI95, HGSM11, HG11]. 3 [HMF10, HMD93, HGB98, IAP^{+}11, JGR85, JH98, Jok98, dOSJVBS12, KL93, KWK94, Kd88, KMA^{+}00, KMGC84, KNO^{+}09, LCT09, LM96, Lat97, Lau97, LF83, LC85b, LR90, LK91, LP91, LPR93, LKC94, LPS^{+}11, LST13, LS92, Li92, LS08, LL13, LM95, LCC89, LSHT02, LT90b, LS12, LSTF12, LEA^{+}10, LK00, Ma94, Ma96, MS96a, MW00, MFJ95, MC09b, MMA06, MWTN04, MCT10, Mer88, MN95, Mil09, MBMC11, MKY01, MDR91, MB95, Mur87, NSK^{+}97, NG98b, NT10, NKP11, NFA04, NL96, NDO09, NSEA13, Nur86, OG98, OBS05, OMBH06, OJRT08, OCVV04, PK99, PSR08, PS07, PMW05, PCV94, PF87, PHK92, Pot87, PBN^{+}09, Pud98, QL96, RAH97, RG12, Rem04, RNDA13, Ros10a, SAA93, SC96, Sal90, SCD11, ST96, SI96, STV09, SM06, SN99, SGHM00, SCS91, Shi99, SKU^{+}09, ST10, SKVS13, SB00]. 3 [Ste01, SWS11, SRML09, SKBS13, SS11b, SB02, TB99, TLGS05, TCH07, TCMS04, Thi92, TG96, THO94, TN05, TN08, TML00, TF81, TCCK90, TA88, THL03, Udu81, Udu82, UK12b, UFF06, VV02, Ver81, VPAM12, VKP98, WSV91, WW94, WPS03, WWLV11, Wei88, WEY06, WWB84, XOF05, XL88, XP11, YC78a, YB07, YHR^{+}05, YT99, YH83, YC98, YYF89, YJC^{+}09, YL90, ZW97, ZZZL13, ZF94, ZXY^{+}12, ZZC^{+}13, ZZLZ13, ZZZY13, ZH04, Ziv10, tHV09]. 3 × 3 [SW04, GV84]. 4 [Gol13, HCLL89, RWWH00, Ron86]. 5 × 5 × 5 [SB02]. 6 [Ma96, SIT07]. 8 [CPC99, Lat93]. ^{+} [Abe84]. _{2} [CH11]. β [ABE98]. C^{1} [Klu78]. d [Pat13]. F [LMRMJ08]. G^{1} [BFRA12, FH12]. G^{2} [KP11]. H^{2} [TJ12]. k [JLD12]. kd [GK04]. k × k [O’G90]. l_{1} [DOSD11]. L_{2} [ZW03]. λ [CL05]. M [HBH11, Kle85]. μ [CW02]. N [BB83, Dou92a, Dou92b, YK95, DSdlH^{+}11, KCD00, RFLSA11, Thü03, UCB13]. O(N) [Ano92a, KT89, San90]. O(n log n) [WBR86]. P [Loh10]. q [MRW^{+}97]. S^{+} [DSS94]. X, Y [SPS81]. Z^{2} [Egg98].

-based [PLLL03]. -basis [CW02]. -best [YK95]. -Connected [Ron86]. -Connectedness [Lat93]. -D [LEA^{+}10, AOR94, BT05, BM95, BGK95, BSB87, CGH08, CCA92, CC96, DPR92, DN82, FHMB84, FRL^{+}98, FST94, FL96, GR87a, HHI95, JDP97, JH98, KK95, KMGC84, LCT09, LF83, LK91, LPS^{+}11, LSHT02, MKY01, NT10, Neg12, NL96, Rem04, TTIM96, TSK94, WSV91, WCZ02, WWB84, YC78a, YHR^{+}05, YH83].

-Dimensional [CJ82, GR85, Kle85, Nur86, UCB13, KCD00, Pat13, Thü03]. -dimensions [Gol13]. -DOF [SIT07]. -estimator [HBH11]. -Manifold [KADS02, GK04]. -means [JLD12]. -measure [LMRMJ08]. -medial [CL05]. -Neighborhood [MMS97]. -norm [ZW03]. -observation [Dou92a, Dou92b]. -Point [CPC99, HCLL89]. -Series [MRW^{+}97]. -simple [Loh10]. -Simply [Her98]. -Skeleton [ABE98]. -Space [BB83]. -state [Ros10a]. -subfields [GH90]. -Subiteration [Ma96, PK99]. -Tree [Abe84]. -trees [GK04, DSS94].

05 [KSM^{+}06].

1 [McD81b]. 1-form [HXS09]. 10th [LV03]. 186 [Ano92b]. 1978-1980 [Haw82]. 1994-1995 [Ano95a]. 1999 [Ros00b].

2.5D [LS09]. 2000 [BSW01]. 2002 [HCS03, Wyv03]. 2003 [Ano04q]. 2008 [Jam09]. 2010 [KB12]. 21 [Ano92a]. 214 [Oli01]. 2a [WR96]. 2D [BB04, CBK03].

329 [KK93]. 3D [CBK03, DR03, LZWP03, NFU02, PS03, She03].

4 [FSF07]. 4D [GK03].

62 [GW93a]. 6DOF [SE11].

74 [HQ12b].

8th [GHPW12].

’95 [Ano95b]. ’97 [Ano96g]. ’98 [WP00]. ’99 [KS00].

A. [Rab92]. AAM [ARARCE11]. AAMs [HDF12]. Abdominal [SP81]. abnormal [XG08a]. abnormalities [KS91c]. Absolute [DPB00, Kis96b, BK07, Dem05, LD95].

Absorption [Sbe00]. abstraction [MDFS11a]. Accelerated [AHDM10]. Acceptors [ITN84]. Access [DCCL99, KH83b, SGA12]. accidentalness [HBA93]. Accumulation [Ris89, BCM13]. Accuracy [ACB98, BKLO87, Ber84, CPK99, DH92, GK77, HS05, LHH^{+}98, Sha06, Tan95, BHMB10, GBF12, MN06, MM06]. Accuracy-Based [Tan95]. Accurate [AK10, AK11, AS08b, BGK98, CJC01, FS03, HBKN87, LK91, Lin02, MC09a, MG95a, SRTBS91, SLN95, TLCH05, VM06, VCVQ^{+}98, Coe12, CYE91, KS91c, LSKK10, PZX13, SB89]. Accurately [LMC09]. ACCV [Ano95b]. Achieve [CPK99]. Achieving [ZS09]. Acknowledgement [Ano02j, Ano03w, Ano04w, Ano05v, Ano06t, Ano07m]. Acknowledgment [Ano94a, Ano95a, Ano96a, Ano97h, Ano98f, Ano00f, Ano09s, Ano10q, Ano12r, Ano12s, Ano13t, Ano13u, Ano11q, Ano99g, Ano01p]. Acoustic [CFM02, NT10]. acquired [PS12]. Acquiring [CH06a, WWB84, ZB05]. Acquisition [AR77, LPF78, GCEC07, MHW89, PS03, TLGS05, WNH05, YAK^{+}08]. across [AVBK10, JSRS08]. Action [EK12, IB01, MU11, ZG10, AAASC11, ASCF13, CCFC13, JLD12, JLD13, KRK11, KH13, LYSS12, PC05, TCZ^{+}12, WRB06, WRB11]. action-recognition [PC05]. actions [PD11, UK12a, YS06, YS08]. Active [Alo92, BC92, Car96, CTCG95, DM01, DCTO97, IP98, KR99, LVW97, LSHT02, SI03, SO01, TM94, TS92, TS86, WM92, WCH98, YYL96, BH12, CUAT13, Coh91, CCD11, DBZ07, MFB11, MCB13, Mil09, MBMC11, PD05, TP05, UM05, WB12, WWJ13a, XAB07, YLA09, TRG^{+}13]. Active-Passive [TS86]. Active-Space [SO01]. Activities [YB99, DIMT12, VZP^{+}09]. activity [CCFC13, CPT07, HNB04, NN13, OGH04, PKK^{+}09, RR06, RS03, SOD10, SSdVL06]. actor [FR11]. AdaBoost [YCA^{+}10]. AdaBoost-based [YCA^{+}10]. adaptation [CSS^{+}13a, DD11a, HG11, YNCO11]. Adapted [Ku84, LCSL07, VMP03]. Adapting [QT10]. Adaption [EW91]. Adaptive [CT12, CS04, CHB86, CYC10, DD11b, FCG01, HH97, HGS08, IM06, JV97, JM92, KD86, KSC97, Mas85, MPC94, MC95, PMR92, PM89, PP95, PAA^{+}87, PEF92, PW86, RM02, STEK96, SF96, Tan95, WH00, WB82, WWJ13a, YCKA10, Zha97a, ZGLP12, BSM10, CRT90, EDB12, FLHK08, GS08, HYJ11, HBB^{+}12, HBG13, LRW08, LL04, LC88b, LYG07, MTAA11, MĆK09, OBS06b, WSSS13, XG08a, ZH04, PCC13]. adaptive-binning [LL04]. Adaptive-neighborhood [PMR92]. adaptive-resolution [ZH04]. Addition [DBB83]. Additive [BN84, Bie87, HNRR90, LJ90]. Address [JC94, DF91]. Adjacency [KCD00]. Adjustable [CSS13b]. adjustment [BS05, DSH04, GA09]. Adolescence [Zuc76a]. ADR [KŽ12]. Advanced [KK88a, ZS11]. Advances [HD07, GHMT09, dOSJVBS12, KHA^{+}05, MHK06]. Advantage [FL96]. advantages [KHK10]. Aerial [AMA79, BM99, CJC^{+}98, CJC01, FKL^{+}98, FTW81, FMR01, Gro82, GN98, HN88, May99, NMI79, PCJC98, SM93, UA77, WH01, ZC93, JRH03, Lap88, LSC08, LP90b, MHW89, TDWH07, YZ06, ZN13]. Affine [Ano01s, BH99, Che96, HJK02, Luc01, MB94, NG98a, SBZ97, ACAAC^{+}08, FB12, HY11, HN95]. affinities [CU10a, CU10b]. Affinity [CU10a, CU10b, Fre76, PDTE06]. affordances [KRK11]. Aftereffects [Sav87]. against [CCYC12, RH06]. Age [KdVL99]. agent [KK13]. agents [UM05]. Aggarwal [CV13]. Aggregation [FKL^{+}98, MYLP98, NP87]. aging [XFSC13]. Agnostic [FH12]. Aid [Lor83]. Aided [BBK78, CLR80, Sel81, FR80, GC80, SB13]. aiming [FLB06]. Air [Fog93, BKP10]. aircraft [WW80]. airport [HCN90]. Akio [KK93]. albedo [TS11]. Alexandrov [Ano94g]. Algebra [AS93a, AS93b, Gho91, HJS89, JC93, RWD90]. Algebraic [BGSdVL98, BD94a, CCF01, CP99, DC01, MNSK98, Sar83, UTB^{+}11, HR90, MD95, RH91]. algebras [EW91]. Algorithm [AS83, AB88, Ami90, ACB98, BGT94, BM98, BP84, BF87, Cap84, Cha81, CPC99, CMW^{+}97, CRC97, CP79, CD93, CC01, CWSI87, CCMW97, CCS95, CHRM96, Dan81b, DM82, DW87, DJG01, Ebe76, Ehr78, ER96, FDMA97, FS84, GW93a, GR81, GPP88, GSK02, HE81, HW83, JB92, JS87, KS91a, KPS76, KLK88, KU92, KU95, KSC97, LC79, LMR84, LM96, LHS01, LH88b, LD98, LHB87, LSZ83, Lum83, MS96a, MPJN87, MNHO00, Mer81, Min79, Mis84, NDBT95, OCON82, O’R85, OS95, OWW85, PK99, PPK93, PKP97, Pav80, Per76, Pet85, Peu79, PA83, Pud98, QL96, RW88, RMR85, Rob85, Sam82b, SSF94, Sch76, Sch78, SCS99, SLY89, Shi81, SP97d, SHKP98, Sin87, Spe92, SF96, Sur86, TV99, TG96, THN92, TF81, VV92b, Wec81, WW88, WM92, You86, ZHAH88]. Algorithm [dFP92, AFH81, BGPD09, BD94b, Bha91, Buz03, CBD^{+}03, CMBV04, CT12, Cha74, CL90, CCL04, CR03, CL91, CR90, Cre08, DBF04, Dam08, DV82, Fra81, GW93b, HDS08, HWW06, HZW^{+}10, KK95, DFP^{+}13, LS91, LY90, LZLP10, LPZ08, Lio91, Loh10, PCC13, QHXC12, RFLSA11, Rie75, SAS12, Sez90, SC93, SW94, TL05, VRKL13, WW80, WSSS13, Wil78, YB07, ZSCP08]. Algorithmic [WS91]. Algorithms [BB87, BM95, BN84, BDL92a, BW93, BS00b, Bur83, CKK^{+}12, CP81, DK79, DRCF95, DUC97, Dor79, Ede87, Enk88, FH84b, FSS94, Fie86, FHP01, Fon90, Fra79, Gla93, GH92, GO87, Kas94, KY86, KK81, KMP05, KK83, LPH01, LHH^{+}98, Ma96, MW00, MD95, Mil99, MWL99, MEDT96, MN94, MR92, ND92, O’L88, Oli00, Oli01, OW86, Pav78, Pav79, Pog85, QY02, ROH88, SUO00, SU01b, SWG02, Sam84, SS87, TH86, TSK94, THT^{+}98, US96, WF78, WH94, WWW95, Yan93b, Zhu89, dM92, vvv88, BDL92b, CX11, Chu77, DR03, DSdlH^{+}11, Fra89, GRGB^{+}13, GH90, GS12, HD07, HZLM11, Kaw78, KL11, LKC94, MUS06, MG95b, PDK96, PM82, PMW05, QKH^{+}12, SW05, SRS11, SKS11, TCAC90, VBN11, VCBC88, VD90, VPAM12, XO93]. Aliased [BBD^{+}94, Fie86]. Alignment [BU93, CVB09, KTNO97, BAP08, CPS10, FR11, HJ12, KA08, LH03, MCB13]. Allocation [WB82]. allowing [KDV12]. Almost [Bid86]. Aloimonos [Zha97b]. Alone [Mur87]. Alpha [McC82, SC99a]. Alpha-Root [McC82]. Alphanumerical [Tan81a]. alternate [ZZ10]. Alternating [PLS97]. Alternative [BP95, Mil99, SF96, SM13b]. Alternatives [Dub77]. ambiguities [Neg12]. Ambiguity [CM99a, Pel79, YK08]. Ambiguous [Pel79, San77]. American [VM01]. Amodal [BF05]. among [SU01b, UK12a]. Amount [KABP98]. Amplitude [BRW85, MY87b]. Analog [LSVD85, MM81, MM80]. Analogs [KR85a, LM00]. Analysis [ACLS98, AC99, AMA79, ABW97, Ano96g, AK96, BEPW00, BC88b, Bri98, CV92, CMW^{+}97, CRC97, CL97a, Che98, Che96, CR97, CN95, Coh85, CH78, DR93, EK98, EU85, FF79, För87, FS80, GSP01, GPK99, Gav99, GBR79, Gla93, GSU00, Har80a, Har80b, HGA86, Hsu79, IF99, JC94, JM92, Kan94b, KS95a, KS95b, Kis96a, Kov89, KH94, LZ97a, LBD92, Lee81b, LS01, LBS80, LW85, Lu78, MA78, MMN83, Muk97, NMI79, Nag78, NDN^{+}97, Nis97, Oka88, OP96, Pag99, PE92, PD79, Pav78, Pen92, Pen99, PS97, PA98, PSM80, Ram76, Ros88, Ros92, Ros93a, Ros94, Ros95, Ros96a, Ros97, Ros98a, Ros99a, Ros00a, Ros00b, Ros01, RLC^{+}11, SB96a, ST80, SI96, SB85, SP97b, SJ93b, Shu97, SHKP98, Sko86, Spi98, SA85, Tam83, TZ82]. Analysis [TS01, TSK94, UA77, VDO85, Wec79, WW97, WH00, WC92, YYL98, ZT80, AC07, Ang07, AZN11, BC10, BCM06, BRP04, CHP^{+}11, CAF09, CLC91, CPT07, CP09, CLCO13, CC03, CH06b, CKS^{+}05, DB03, DRK03, DIMT12, FL92, FLB06, GYTL09, Her90, Hu08, HW06, ITNP12, JTEA91, KZD^{+}11, KLL^{+}11, KB12, KSG^{+}13, LM89, LYKL12, LFMP13, LL04, LJ90, LLE^{+}09, LPVM13, LP10, LWH03, MPF07, MVP06, MP09a, MR05, Mat89, MHK06, MĆK09, OH05, PE09, PSE^{+}11, PKK^{+}09, Pop07, RM91, Ros89, Ros90, Ros91, SJST07, SYK96, SHC^{+}12, SSdVL06, SD90, SB89, SCCP05, SRML09, TJ12, TCZ^{+}12, TDT12, TCC90, UTB^{+}11, VMP03, WS89, WJG02, WY07, WS08, WLI08, WLMG08, XG08b, YLM11, YYF89, YSD03, ZMCA05, ZG10, ZZP12, NLW13]. Analytic [HCLL89, Tri90, WW90, XSD12]. Analytical [AAS97, TTA94, TS86, YSL11]. Analyze [Gro82, WN87]. Analyzer [CH80, MSN82]. Analyzers [Ley85]. Analyzing [AM00, Bic98, Bd96, CCR^{+}05, KW87, Pie79, TM86, TA88, CKS^{+}05, FS03, MB05, RSPD12]. Anatomical [HRS02, LSB^{+}00, LK00, MMA06, ZZC^{+}13]. Angiograms [SGHM00, NBDB04, YHS89]. angiography [BT05]. Angle [Kan88, BPBS13]. angles [KPE90]. Angular [ABMT87, APV99]. Angular-Based [APV99]. Anharmonic [BB83]. Animal [FRDC06]. Animate [BB92]. Animated [FM99, MKS^{+}08]. Animating [ATN83, GS01, BGA05, DM78]. Animation [Ano94f, FM96, Jam09, KB01, LK01, YKC^{+}86, DLP13, yKL11, OH06, PY08b, Ros10b, SJ12, UPBS08, ZCCD06]. animations [GVK06, TM07a]. Anisotropic [BS00a, JGR85, BI11, GR05, KGC05, PDA03, SGS^{+}10]. Annealed [RRR11]. Annealing [BCG95, KBZ96, PB99, JLL13].

Anno [VDO85]. Annotated [Ros01, EHG^{+}10]. Annotation [STEK96, XL98, ABC^{+}03, BSMK13, MR05, TLWT12, WHM^{+}09]. annotators [SYPK13]. Announcement [Ano94b, Ano95c, Ano97a, Ano01a, Ano01b, Ano01j, Ano01i, Ano02a, Ano02b, Ano03a, Ano03b, Ano03c, Ano96b, Ano96c]. anomalies [CHP^{+}11, RL13a]. Anomalous [JYTK11]. anomaly [BDS12, YGC13]. Anthropometry [BK01]. Anthropomorphic [TGB00]. Anthropomorpic [ST80]. Anti [Fie86]. Anti-Aliased [Fie86]. Antialiasing [KB91c, KON87]. ANTICS [DM78]. antipodal [LB10]. any [AVBK10, MG95b]. Anything [Pri86a]. Anytime [BAP08]. Aperture [FY85, Lee81b, SGA12, Zit88]. Apparent [Ish84, KMB97]. Appearance [BFY00, CW00, HF01, MKK02, SN99, TRG^{+}13, BF10, CD13, DZL07, DB03, ESS10, EL07, HFR06, JSRS08, LSD^{+}07, LHYK05, LPS^{+}11, MC09a, MCB13, MU11, SI03, ŠRDC09, TC11, YO11, YT13]. Appearance-Based [CW00, SN99, ESS10, MC09a, ŠRDC09, TC11]. applicability [KHK10]. Application [ACF00, AM01, ADRY94, BVL02, BKA84, GK98, HVD^{+}89, IKS86, JLD12, KABP98, Kri84, LSB^{+}00, LTS93, MCPB00, MAM97, OH81, OMLL98, RRS83, RAC^{+}13, RMFB02, Sch76, SP81, Sko86, SRHC13, TW98, TZ00, Ull81, VMP03, VY94, WSKH13, ZF94, ZBV93, vdWvO96, AO03, BvdHL^{+}13, BB13, CTCG95, ES81b, FWH13, GCFMT12, GWT09, KGK10, KGFP10, KS91c, KMBH09, Lea92, MUS06, Mar07, MDR91, OC90, PMC13, RC03, RCTV12, SA04, Sez90, TTF04, kWwZ13, YYF89, Ang07, BC10]. Applications [Ano98e, AS93a, AM78b, BY98, BS87, CS89, EPB05, För87, Gui99, Gui00, HT98, KJRA96, MS96a, MKK02, Pag92, Per76, SU01b, SWG02, SS79, Taj83, iTTF82, TPR^{+}00, US96, YKC^{+}86, YBDC93, AM93, Big90, CL90, CBT^{+}04, DB03, Elb05, HS05, KLBP11, KPPK09, LL04, MPVF11, MM05, RC13, SC96, Sah05, TMB12, WS08, WB12, Wei90, XSD12, YJC^{+}09, ZT09]. Applied [Ali77, EOS84, Kov89, KU95, SMB95, WF02, WF78, Her90, MJ11, Moo77]. Apply [WWW89b]. Applying [Bar84, JT86]. Approach [AK77, ABMT87, APV99, Ano94f, AD84, AMMV99, BKLO87, BC88a, BB91, BZ99, BSB87, BS88, CH96, CCP97, CT88, CT93, DGH98, De 93, DY98, Dod98, DC01, Dou81, FF79, Fre76, FM99, Goo92, HL84, HB86, HP78, HLF^{+}97, Hsu79, HP96, HW94, JW87, JN93, JJ83, KG94, KW00, KH98, KKO98, KH86, KKH96, LM00, LK91, LS01, Lem79a, LN85, LLL86, LS92, LAS94, LSHT02, LJ91, LH84, Lu78, MRW^{+}97, MZ96, MMP85, MYLP98, MSH86, NDN^{+}97, OMLL98, PG94, PLL00, Pha89, Pun81, RJ00, RH95, SGHM00, TC87, TS86, Tsa85, Tsa96, Wec79, WP93b, Woj84, YB95, ZXK02, Ano06m, BT05, BDS12, BCM06, BNG03, BPB11, CTM^{+}13, CDT11, CLHW94, DK13, DPR92, Dor89, FFFP07, FKV^{+}11, FSV07, GRGB^{+}13, GKK05, Gus07, HPR90, HBH10, HRC09, HW07, HC13c, KS89]. approach [KS91b, KK95, KL11, KS12, LJHH07, LBM04, LT90a, LC88a, LS12, MPST08, MCQ05, MHMO09, MMP09, NHSC09, Nic95, OBS06a, OS81, PTE12, PIK90, RCVA11, SM12, SMT04, Sha06, SCL13, SA81, SAC09, SG82, TMNM09, TH06, The83, TC89, TBN95, THL03, VCT09, WDB12, XSD12, YS08, dP10, Ano94g]. Approaches [Ahu83, LCZ^{+}01, RC97, YBDC93, BCF06, DCFM07, GH90, GJ10, HHWP03, KYM13, KMN11, SJST07]. Appropriate [Tan86]. Approximate [AKL93, BS04a, CT97, Che96, DBB13, SZKD99, ZCK09]. Approximating [ET94, MH98, O’L88, SC99a, TD83]. Approximation [AHD94, BM98, BB88, CLL^{+}99, CMPP99, DGH98, FHMB84, GK77, HFF93, JB99, JB89, KSG84, KP97, KŽ99, Kri84, KD82, LKE98, LM99b, LL97b, Pen94, PCR86, Sug93, TC86, WD84, Wil81, Wil84, Woj87, WW93, AK91, BG91, CCS05, Coe12, Gou91, GW90, KA08, KHK10, LM95, LRLB11, MK05, OBS05, Ram72, SC93, VCT09, VM06, Wil78, Fiu91b]. Approximations [Bra94, DG01, Gad91, II86, Pat13, WHHB12]. Arbitrarily [DFP85, Klu78]. Arbitrary [ANM98, BCL96, Bor84, DC88, Mor76, TC87, WA87, APB10, BBB11, Coe12, DMMP03, GW90, KK09, LBM04, PSF07, UCB13, WBOL07]. Arborescent [WR93]. Arc [CT95, Lan87, Rut81, TC95, Van84, WWW95, Alb74, Ber89, GW90, TC89, dMFU10, EL91]. arc-weight [dMFU10]. Arc/Chord [Rut81]. Architecture [CL83, CG87, GR87a, MS85, RS88, Tan86, WWW89a, CGL92, DRAB08, MFG10, RM91, SIT07]. Architecture-independent [WWW89a]. Architectures [HS83, Ree84b, TV99, WWW89b, Fon90]. Arcs [DGH98, Dor84, HB98b, Jos94, Lil97, Sau93, WV97, WJW94, BE11, KPMR91, SC93]. Area [Bri98, DBB83, Jok98, KSI98, Kul77, Mil99, MSW96, NS91, CKM11, DGZ12, EL91, ELA91, Fiu91a, GE08, KM03, ZN13]. Area-Based [Jok98]. Area-Invariants [NS91]. Areas [FMR01, Kul83, Ros80a]. ARG [PLLL03]. Arranged [MSN82]. Array [DW87, Mar80, Ree79, Sob78, Sub79]. Arrays [SS76, SSS82, THT^{+}98, CPT07, Sam80]. Art [EOW84, McL96, CH06b, CGW^{+}07, JM09b, KTP08, SCD11]. art-directed [CGW^{+}07]. Artefacts [PMV00]. article [Ano01s, GW93a, KK93]. Articulated [ACLS98, DF01, GESB95, Tay00, BCMCB09, DGC12, HW07, IAP^{+}11, MFB11, RRR11]. articulating [HER81, NHY10]. Artifact [SHS79]. artifacts [BTCH05]. artificial [FY06, HC13a]. Artistic [RL13b].

Ascender [CJC^{+}98]. Asian [Ano95b]. ASL [SLCP85]. ASM [CUAT13]. Aspect [Mun95, NWP97, SB90, ACDB12]. Aspect-Trees [Mun95]. Aspects [Ger85, SKOS95, TS92, VM01]. assembling [RG12]. ASSERT [SBK^{+}99]. Assessing [JOvW^{+}05, Ros96b, CCTCR09, YZY11]. Assessment [BS00a, YY84a, YY84b, SRP10]. Assessments [KOY86, KON87]. assignment [MEYD11]. assistance [HPvB^{+}10, WWH07]. Assisted [YKC^{+}86, ÁB13, PJW11]. Assumption [CM99a]. assumptions [WS06]. Asteroid [CHB86]. Astronomical [De 88]. asymmetry [LSCM03]. Asynchronous [Ede87, JDP97]. atlas [ZZC^{+}13]. atmospheric [LWGP08]. Atoms [Max84]. ATR [LCZ^{+}01]. Attending [TLMT^{+}05]. Attention [DCTO97, GFW13, HRC09, SKOS95, TW98, BBHF10, DL05, Ham05, IKST05, JOvW^{+}05, SFWG08, WRKP05, Ano05p, FRNS05, HH05]. Attention-from-motion [HRC09]. Attentional [MNE00, YYL96]. attraction [RM03]. Attribute [BJ96, GK95, ZRKZ^{+}11]. Attributed [CTF^{+}98, YF80, PLLL03, SRS11]. Attributes [Hen98, LSTF12, RFS03, TRS06]. Audiovisual [DGG08]. augmented [CKM11, MBH^{+}12, WS03]. Augmenting [Ros96c]. Aurora [GFL^{+}11]. authentication [DIMT12, PY08a, UBEP09]. Author [Ano93a, Ano93b, Ano93c, Ano94c, Ano94d, Ano94e, Ano95e, Ano95f, Ano95d, Ano96e, Ano96f, Ano96d, Ano97c, Ano97d, Ano97e, Ano97f, Ano97b, Ano98b, Ano98c, Ano98a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00a, Ano01d, Ano01e, Ano01f, Ano01g, Ano01c, Ano02c, Ano02d, Ano02e, Ano02f, Ano02k, Ano03r, Ano03t, Ano03u, Ano03v, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano06p, Ano06q, Ano06r, Ano06s, Ano07l, Ano09a, Ano03s]. Auto [AHZ96]. Automata [HH96, Ros10a]. Automated [BNL90, Chi88, CJC^{+}98, ES06, HVD^{+}89, HPvB^{+}10, Hsu79, LSB^{+}00, Lor83, NJ95, Oka88, PKD07, RCJ^{+}13, SZ03, SRP10, TZ82, CYP^{+}10, MO11, TDK10]. Automatic [Abu89, ARARCE11, BW76, BL98b, CM92, Cho79, CNC03, De 83a, EU85, FG89, FT84, GN98, HH77, HF80, HH98, IHTA90, KN04, KY06, KB12, LSVD85, Lhu08, LSHT02, MG95a, May99, Mey86, MEDT96, Oka84, Oli94, STEK96, Tan11, VV02, WH84, Wha91, Wil98, XYW^{+}08, YKC^{+}86, YJC^{+}09, ZZZ06, ABC^{+}03, DK13, FFY^{+}04, HDS08, IK89, IH91, MDdMG09, MCT10, MKA73, QKH^{+}12, RC13, SDR91, USKB10]. Automating [MHW89]. automation [CMH13]. Autonomous [GVK06, KR99, ST07, SK85, BKP10, JBC08]. Autonomously [KP00]. autoregressive [KS91b]. auxiliary [BW11, ZQ11]. avatar [LL06]. AVCD [DK13]. AVCD-FRA [DK13]. Average [FS85, GMT00]. Averaging [HC96, MMA06]. Avinash [KK93]. avoidance [CSS13b, JM09a]. avoiding [GB13]. Award [Ano12q, Ano13r, Ano07j, Ano07k, Ano08p]. aware [GWCO11, PL10]. Axes [Fri86, SB98c, Wec79, Nac82]. Axial [Ros86a, PA13]. Axiomatic [SU01a]. Axiomatized [KR85a]. Axis [BM95, CCMW97, CS01, Sam85, SB96b, SPW96, Wal88, WBR86, WBR88, CL05, LKC94]. Axis-Translation [Wal88].

B [Ano93d, Ano93e, Ano94f, GW93a, Hor79, Rab92, Abe84, BG79, BG80, CXY^{+}09, CL00b, CL91, CLR80, FSSL86, FSS94, GSS00, GW93b, GCB90, HB91, MY87a, Pha89, Rab92, RAH97, RG12, TC87, XWYY10, YKC^{+}86, YZZ^{+}10, ZK05]. B-DNA [RG12]. B-Solids [RAH97]. B-Spline [GW93a, Rab92, BG80, CL00b, FSSL86, Rab92, TC87, YKC^{+}86, BG79, CXY^{+}09, CL91, GW93b, GCB90, HB91, XWYY10, YZZ^{+}10]. B-Splines [CLR80, FSS94, GSS00, MY87a, Pha89, ZK05]. B. [Kul79a]. Back [Gul79, BK07]. back-off [BK07]. Back-Projection [Gul79]. Background [Ant98, DS07, KA94, LAS94, YCH07, JBR08, LRLB11, SZ07, TA11, VAWW10, ZCF13]. Background-subtraction [DS07]. background-weighted [JBR08]. Backgrounds [YK87, HK93, LC88a, LBNS09]. Backpack [HCHD01]. Backtracking [KW12]. Backviews [SK02]. Badler [Ano94f]. Ball [MSSS09, CG09, ROJX09, YJC^{+}09]. Balloon [CM95]. balloons [Coh91]. Band [BBD^{+}94, OS87, Mil09, MBMC11]. Band-Pass [OS87]. Band-to-Band [BBD^{+}94]. Bank [KC92, TKL^{+}09]. Bank-Based [KC92]. Bar [Maa94, dFCS93]. barium [KS91c]. barium-filled [KS91c]. barrier [Liu10, SCMS13]. Barsky [Rab92]. bas [ZZZY13]. bas-relief [ZZZY13]. Base [BW76, HMD93, KPH02, KZW12]. baseball [GHHX04]. Based [AS83, AMA79, AB88, AEM98, APV99, Ano93e, Ano01s, AA93, BIP00, BY01, BGSdVL98, BM98, BS99, BD94a, BL00, BL01, BRW85, Bra97, CFS98, CC97, CL98, CMVM86, Che00, CCS01, CT93, CH80, CPD93, CL97b, CW00, DWS83, DFP85, De 93, DRCF95, DCCL99, DUC97, DTG96, DLHT99, DY98, Egg98, FWL88, FDMA97, FL96, GY99, Gla93, Gle01, GM94, GY01, GPP88, GY88, HTEB11, HKD95, HR99, HSIW98, HF01, HLKF95, HLF^{+}97, HY98, IF95, JB99, Jok98, JB91, JEK98, KG94, KMI79, KW00, KKO98, KIF85, KR98, KABP98, Koh81, KK92, KK93, KMA^{+}00, KP00, KU95, KC92, KR99, LCS84, LSVD85, LL99, LHS01, LW85, Leu92, LN85, LHHC98, LLSV00, LK00, Luc01, MBKB02, MS97a, Mar82, MPJN87, MBK81, MHN84, MS97b, MN85, MZ96, MB94]. Based [MWL99, Min94, MG01, Mok97, Mok92, Muk97, MSH86, Mur87, NK00, NFJ93, Nis97, Nis98, OG98, OY92, PLL00, PBQ99, PR92a, PM97a, PMV00, PM97b, PR92b, Ree79, RWWH00, Roh94, SK02, SUO00, SYF99, SSF94, SB98a, SMK02, SB85, SLST99, SN99, She86, SLK86, SBK^{+}99, SPK^{+}02, SHKP98, SK88, SLL01, SL96, Spe92, SB94, SBxx, SK84, Tam83, TI01, Tan95, TY01, TB99, TM94, TS01, THCG84, VFV93, VKP98, WF02, WW97, WL88, WW93, WS91, YC98, YB01, ZM94, ZBV93, AAASC11, ALP06, AQ09, AS09, ACG^{+}09, ABEN09, AK10, AK11, AWK04, Ang07, AS08b, AZN11, AO04, ARARCE11, BI10, BBF^{+}11, BZS08, BY08, BL89, BL04, BL09, Big90, BMM^{+}07, BBB11, BPW91, BH12, BPB11, CBD^{+}03, CGU11, CPC08, CGAY13, CM12, CBC^{+}07, CTM^{+}13]. based [CL90, CK11, CSJ13, CLC91, CS10, CHZ^{+}13, CSS13b, CJL06, CP09, CD13, CU10a, CU10b, CR88, CG04, CZZS07, CGG91, DK13, DT10, DWB11, DS07, DN91, DD11a, DRK03, DQ05, DQ04, DLP13, ESS10, EDB12, EBN^{+}07, EyGS11, FPC^{+}08, FG89, FMGA^{+}12, FML12, FFY^{+}04, Far11, FB12, FKV^{+}11, FJJ91, FAB12, FSV07, FKS10, FK09, GRGB^{+}13, GB10, GSPL10, GBHS06, GRB13, GS12, GGMV08, GB13, GH08, GHHX04, GCPF08, GK90, GM90, GFW13, Gus07, Ham05, HDS08, HD09, HSH07, HN82, HXS09, HGR^{+}13, Hei04, HHWP03, HSKH07, HFR06, HNB04, HQN05, Hu08, HC13b, HMA10, HWW06, HDF12, Hub12, HGS08, ILRB04, ITNP12, JBC08, JBWK11, JXCZ13, JLD13, JM09a, JMPG11, KZD^{+}11, KK07, KK09, KLL^{+}11, KS12, KY06, KZ05, KDV12, KT07, KGU10, KL10, LBK10, LMRMJ08, LY05, LJHH07]. based [LFMP13, LZLP10, LPZ08, LL12, LFL08, LWGP08, LYCG08, LC09, LLC11, LEA^{+}10, LBCA10, LAL^{+}10, LN10, LWH03, ML13, MP09a, MC09a, MSG10, MTG07, MBH^{+}12, MR05, Mat89, MCT10, MCQ05, MHSP10, MGPP11, MW13, Mig12, Mil09, MBMC11, MHK06, MP09b, MTAA11, NHK08, NRJ11, NBPF11, OMBH06, PLLL03, PFV^{+}11, PL07, PSR08, Pec91, PD11, Pen03, PBM^{+}11, PKK^{+}09, PA10b, PFGG09, PR03, Pop07, PBN^{+}09, PZV13, PBG04, PMF90, PIK90, RM03, RSS07, RCVA11, RFS03, RNDA13, SGS^{+}10, SE11, SBB10, SM12, SB91, SIK92, SI03, ŠRDC09, SG11, SMT04, Sez90, SS06, SW94, SW05, SH08, SFWG08, SHS03, TAK09, TL05, TMT10, TA13, TM07a, TCMS04, TMN06, TC11, UBEP09, VAWW10, WJG02, WPS03, WZL^{+}03, WLZW04, WZ04, WLW06, WHHB12, WRB11, WS06, WLI08, WR08, WB11, XAB07, XYW^{+}08]. based [YB07, YHR^{+}05, YCA^{+}10, YGC13, YYF89, ZZZL13, ZZL13, ZZLZ13, ZLS^{+}13, ZCF13, ZUS06, ZCK09, dSdSF^{+}12, FRNS05]. Bases [MT84, Nis95, SK88, GS12]. Basic [ME98a, KP12, XO93]. Basis [ADRY94, Bar85, CL86, Hum79, BSM10, BH12, CW02, FSF07, HR90, LPR^{+}03, OBS06b, RH91, TFB80, WR08, ZK05]. basketball [CD10, PKK^{+}09]. Batch [CWC94]. Bayes [She86]. Bayesian [Car96, CC07, DLF06, FFFP07, KDV12, LWH03, MC09a, PP95, QC04, RH95, SC00a, SJ93a, SAC09, SS11b, TN07, YC98, ZCK09]. Be [Bog88]. bead [FLCdA06]. Beam [DB76, Gul79, PA82]. Beating [AHRW87]. Beck [Bri86]. Beckmann [RH06]. before [Mor90]. Behavior [GJH01, SC00a, GZJ05, KDV12, LL06, TDT12]. Behaviors [GMW12, SVS97, GVK06, WWH07]. Behaviour [CX11, CGH08, HFR06, SGH07, WMBY12, XG08a]. belief [BCMCB09, CS07, PL08]. benchmark [EHG^{+}10, THL13]. Benchmarking [MNCG01]. benchmarks [DFS08]. bending [TG13]. Bernstein [GS12]. best [AQ09, TCB^{+}08, YK95, ZW03]. Better [Ham77, Pri86a]. Between [Cav87, MAN84b, SK84, Åst97, BS96, CXY^{+}09, CU11, Col97, CDH99, HKM12, KWK94, KHB01, KŽ12, Lio91, PA98, PRW97a, TMT10, Ül01, WDN^{+}12, WS91]. Beyond [CM99a, HD07]. Bézier [GS12, Gos89, HWJ96, SZ96b, WX91, WW95, Zha99, ZW03]. Bhanu [Ano93d]. Bi [TC87, CLHW94]. bi-level [CLHW94]. Bi-Quadratic [TC87]. Bias [Che98, WH00, Ber89]. Bias-Reduced [Che98]. Bias-Variance [WH00]. Bibliography [Ros01, Sch80b, Sch81, Sch82]. Bicubic [Dub76]. Bidimensional [MSM81]. Bilateral [ZW97, yKL11]. Bilevel [Alg83, JJN76]. Bimodal [FRNS05]. Bimodality [LBD92]. bin [MGW10]. Binarization [O’G94, CMH13]. Binary [Bur80, Cai88, FK83, FS84, Hei99, HJS89, JEK98, KZ93, KPS76, KLK88, Kle80, KR85b, KD96, Kub84, Kul77, Liu93, LB87, MW00, MMP85, MKW94, ND97, OY92, Pav80, PS95, RM98, Sob78, SA85, Wah83, Wal88, YK87, YPVv81, ZSN96, vv92a, BPBS13, BDHM09, Dou92a, GRGB^{+}13, GK03, GU89, HQN05, KB91b, LWGP08, MR89, MSS90, MB11, NKP11, OEK08, SC96, Sam80, SW05, SM13b, TBN95, VBN11, ZG91]. Binary-Image [LB87]. Binary-Valued [FS84]. binning [LL04]. Binocular [Cav87, CPC99, Che91, De 93, Gri84, GY88, LHB87, MB85, TM94, WD96, LS08, PMF90]. Binomial [Olk95]. Bintree [Bie87, OY92]. bio [BC10, BCDH10, EK12]. bio-inspired [BC10, BCDH10, EK12]. Biological [Sko86, SGDP01, FPC^{+}08, Hil83, MSG10]. Biologically [BL98a, HL13, MFG10]. biologically-inspired [MFG10]. Biomedical [ABW97, AM78b, LBS80, KORC10]. biometric [DIMT12, HBF09, LFMP13, WF05]. biometrics [AZN11, BHF08, HBL^{+}11, HNC05, YB07]. Bisector [ALSR11, Pet00]. Bit [TV99, WB82, Kaw78, Sau91]. bit-rate [Sau91]. Bit-Serial [TV99]. Bitmapped [vv92a]. Bivariate [SK84]. Black [Mor76, CM94b]. Black-White [Mor76]. blend [BFRA12]. blended [SSS13]. Blending [CCF01, Dub77, ZQ11, Elb05, LJHH07]. Blind [ZSN96]. Blobs [FSS84, Kul77, FB12, SI03]. Block [BLd95, Gaa77, JC94, SJ93a, WWC82, HMA10]. Block-Chords [Gaa77]. Block-Wise [SJ93a]. Blocks [Kri84, SS76, Wal88, NHY10, RG12, RL13b, WS89]. Blood [HL84, LNY83, WHL84, TDK10]. blossoming [GS12]. Blue [CEP84, PS94]. Blue-Dyed [CEP84]. Blur [HKZ87, YK97, FM91, KM94]. Blurred [Bra85, KH86, YK97, CG09, FM91, LcTT91, LTT91, LX88, TLT91a]. Blurring [NI82, BTCH05]. BMVC96 [Ano96c]. Board [Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano10a, Ano10b, Ano12b, Ano12f, Ano12g, Ano12h, Ano12k, Ano12l, Ano13a, Ano13n, ME98a, Ano05f, Ano05l, Ano05m, Ano06h, Ano06l, Ano10p, GSPL10, Ano02g, Ano02h, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano05e, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05n, Ano05o, Ano06d, Ano06e, Ano06f, Ano06g, Ano06b, Ano06c, Ano06i, Ano06j, Ano06k, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano08f]. Board [Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano12a, Ano12c, Ano12d, Ano12e, Ano12i, Ano12m, Ano12n, Ano12o, Ano12j, Ano12p, Ano13c, Ano13e, Ano13g, Ano13h, Ano13o, Ano13p, Ano13b, Ano13d, Ano13f, Ano13i, Ano13q, Ano13j, Ano13k, Ano13l, Ano13m]. Boards [ME98b, SDR91]. Bodies [CP86, GK98]. Body [BK83, HZ86, KSS92, LC85b, LH88a, YH83, BCMCB09, CGH08, CFCP11, CPT07, DLF06, HUF05, HH91, HW07, KL13, NESP10, PA06, PT08, PYS03, RRR11, Rem04, SMT04, UPBS08, UFF06]. Bone [MDFS11a, MDFS11b]. Bonnie [Ano94f]. Books [Ano97g, Ano98d]. Bookstein [Sam82b]. Boolean [GPK99, HD97, OWW85, Ser80, WW88]. Boosting [CWO^{+}11]. Bootstrap [KN11, BRP04]. Border [AGW85, CCP97, SA85]. bordered [PSF07]. borders [GA91]. Both [LI00]. bottom [KMN11]. bottom-up [KMN11]. bottom-up/top-down [KMN11]. Bound [SZ96a, SHKP98, Zha97b, Bre03]. Boundaries [Cap84, Cha81, LM00, Nai87, Sel86, Udu94, Vel95, WSSD96, Lio91, Ros80a, ZYT10]. Boundary [AKL93, Bla85, BV99, CEC^{+}80, CP81, ES81a, GV78, GR81, GM87, GJP96, HKS06, HL78, Kaw79, KII98, LKE98, Lem79b, Liu77, LHHC98, LB87, Nis98, PH82, RB82, SPS81, SSP01b, TBN95, Udu81, Udu82, UA90, Van84, VTG95, WF78, AT89, BLH91, CR89, DCS05, ES81b, KA12, LS91, LK03, NRJ11, PDK96, PS07, RB89, RC03, SOD10, VPAM12, WPK09]. Boundary-Centered [RB82, RB89].

Bounded [AB88, Wil81, Yan93b, ZZ10]. Bounding [BE11]. Bounds [CT95, Han88, Kan98, Sub90, TC95, WZ97]. Box [McD81a]. Box-Filtering [McD81a]. Brain [CFYU12, Dav97, GMT00, WPS03, ASFP03, DCS05, LPR^{+}03, ZRL^{+}11, ZU09]. Branch [SHKP98, Bre03]. branch-and-bound [Bre03]. BRDF [AH08, GH03, YSL11]. breakdown [HBH11]. Breaking [TY01]. Breast [KHB01, CSY08, SRP10]. Breathe [ZCCD06]. Bright [BT88]. Bright-Spot [BT88]. brightness [TLCH05]. Brill [Bec86]. British [Ano96c]. Broadband [SM10]. broadcast [MSSS09, WHN08, YJC^{+}09]. broadcasts [DRK03]. bronchoscopy [HSKH07]. Brownian [KJRA96]. browsing [MĆK09]. brush [Pha91, XTLP04]. Bubbles [TK97, ZYP09]. Building [BM95, CJC01, DCH12, Dou81, FMR01, GN98, HB98a, Hen98, HHI95, JC81, LKC94, LN98, PCJC98, SGS01, SF95, VV02, Che08, HBH10, RG12]. Buildings [FKL^{+}98, HN88, May99, JRH03, KN04, LP90b]. bundle [BS05, GA09]. bundles [LAL^{+}10]. Burger [Ano93d]. Bus [THT^{+}98].

C [KK93, Zha99]. C-Bézier [Zha99]. C1 [Far82, GC80]. C2 [Ano06a, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano10a, Ano10b]. CAD [AA93, CFS98, EFF98, IF95, PHK92, ZZZ06]. CAD-Based [AA93, CFS98, IF95]. Cadastral [OMLL98]. calculating [MG95b]. Calculation [LS92]. Calculations [Bas81, MMS99, Shn81a]. Calibrated [WLD99]. Calibration [Buc88, CRC97, DC01, Gui00, HN91, KPE90, KC01, Man86, MBK81, PA13, PBSG12, Rob96a, TTG94, BHSD^{+}13, CXFS06, CF07, CDT11, CP04, CX11, DWW^{+}12, DMW10, FK09, GGO10, JF10, KK09, KGK10, KGFP10, LSKK10, LWLS12, LP10, MCT10, NNT11, QC04, RSL10, SW13, SP06, SCCP05, TM04, WCF10, YJC^{+}09, ZKRH04]. Call [Ano01h, Ano01q, Ano01r]. calligraphy [WLI08]. Camera [Buc88, CF07, CRC97, CYP^{+}10, CC00, DT96b, DC01, GY05, Gui00, HN91, iK87a, KS95b, KK09, KC01, KKH96, LT90b, Man86, MBK81, Rob96a, SW13, Sug88, Wam85, WC99, WCF10, XL98, BPS10, BBH^{+}12, CKM11, CA10, CDT11, DDLP10, ES06, GHA10, GB08, Gol05, GGO10, HC13c, JSRS08, JF10, KD10, KSR^{+}12, KGK10, LBK10, Lhu08, LDD09, LA05, LP10, MFB11, MCT10, NNT11, QC04, RP08, RCTV12, RLC^{+}11, SP06, SST06, SS11b, Tri90, UTB^{+}11, YCKA10, YS06, YJC^{+}09, Ziv10]. camera-captured [LDD09]. Cameras [WLD99, YC78b, AVBK10, BPS10, CVP10, CYP^{+}10, CS10, DWW^{+}12, DMW10, KHK10, KBJ^{+}10, LWLS12, MHSP10, MLH13, NFA04, PD11, PBSG12, RSL10, ROJX09, TM04, WZ08, YC78a, ZZ07]. Camouflage [TY01, WF02]. Can [KL77, Pri86a, KB91b]. Candidate [HFC96]. Canonical [DSNN08, LV96]. Can’t [Pri86a]. Capability [THN92]. captioned [JEF^{+}12]. Capture [CPK99, MG01, CFCP11, GODC07, MHK06, TCMS04]. captured [LDD09, PT08]. Capturing [PCP02, SKS97]. Car [ZBV93, DN82]. Car-Following [ZBV93]. Cardiac [RWWH00, GPDR13, TA13, WSKH13, WWJ13b]. carrier [SDC04]. Carrying [HCHD01]. cartilage [LPS^{+}11]. Cartographic [SP92]. Cartoon [LYKL12, CRH05]. cartoon-style [CRH05]. CartoonModes [LYKL12]. Cary [Ano94f]. Cascade [AVBK10]. Case [CGP85, MS96c, SU01a, Tan86, VF96, DBZ07, Got08, VD10]. Cases [GO87, Lin02, SCCP05]. Cast [Oka84, SCE04]. Casting [LZ97a, Rot82]. catadioptric [BDVK10, DWW^{+}12, GA09, Lhu08, PA13].

Catalog [AOR94]. categorical [SBM^{+}06]. Categories [SPK^{+}02, SB94, SBxx, FFFP07, FKS10]. Categorization [BKMSR98, MK01, GB10, MDFS11b, YZY11, ZG10, vGSV^{+}10]. Categorizing [BKMSR98]. category [GCPF08]. Catmull [SS11a]. Causal [CBB95, LA05]. caused [LX88]. CCD [Chi97]. Celebration [CV13]. Cell [CS82, LNY83, SRL82, Tam83, BEH^{+}81, KORC10, SH09, KL10, SM10]. Cells [HGA86, BEH^{+}81]. Cellular [Dav79, GV84, HH96, HJS89, Kim82, SC98, dM92, Ros10a]. Cellular-Logic [GV84]. Census [PCC13]. Center [CT95, Lan87, OD97, TC95, WWW95, Ami90, Ber89, Dem05, EK12, TC89]. center-surround [EK12]. Centered [GY99, PF87, RB82, SL85, RB89, SCL13]. Centerlines [NGC92]. Central [DPB00, Bar06, Dem05, DWW^{+}12, PA13, RSL10]. centre [DMW10]. centripetal [LYCG08]. Centroidal [AdVDI05]. centroids [KŽ12]. Cerebral [RSB93]. Certain [IKS86, SS79]. cervical [BEH^{+}81, BvdHL^{+}13]. CFA [LPVM13]. CG [CGW^{+}07]. CGI [WP00]. Chaiken [GW93a, GW93b]. Chaikin [Rie75]. Chain [AB88, Cai88, Ced79a, Cha81, KLK88, KD96, Mor76, PR79, SBT85, Tan99, vdWvO96, CL90, LF82, SW94]. Chain-Coding [KLK88]. Chain-Encoded [KD96]. Chain-Encoding [PR79]. Chain-Link [Ced79a, Mor76]. Chains [Cre99, MSM81]. Challenge [CDLD77, MST00, BGPD09]. Challenges [dOSJVBS12, BCF06]. Challenging [Ken86]. Chamfer [MMS99]. Change [Che00, HNR84, HKK08, Lai00, Ros02, SB98a, SJ89, XL98, CCYC12, IH91, MMP09, YCH07]. Changes [BFY00, Chi97, D’H86, JMA79, DD11b, Hil83, XFSC13, YNCO11]. changing [MTVM04]. channel [IJDAB13, NN13]. Channels [BSI87, OGH04, SGS^{+}10]. Character [CH87, Cou81a, KZ93, LSVD85, MLP97, Sin87, ALP06, CGW^{+}07, YT13]. Character/Graphics [KZ93]. Characteristic [BN84, Bie87, LSMS85]. Characteristics [BJ86, Fla89, HS87, Hod95, IE99, iK87a, CCR^{+}05, TG95c]. Characterization [DB94, Far86, Har94a, Har94b, HOPA91, KW99, NSK^{+}97, NS98, PA97, Ree84a, Shi94, SRT01, TTA94, VMUO95, Whi93, WM93, AQ09, ASFP03, BCM13, BB04, BPG05, TCB^{+}08, Žun03]. Characterizations [Eva11]. Characterizing [CZZF97, Kis96b, SC00b, Pon90]. Characters [Bra85, RD77, CBC^{+}07, Lan91]. Checking [DK79, KDK78]. Checks [KABP98]. Cheng [Rab92]. Chessboard [LH99]. Chest [De 83a, WF78]. Childhood [Zuc76a]. children [NKB11]. Chinese [FT84, WLI08]. chip [ZZ07]. Chips [HF80]. Choice [KCA81]. Choosing [Bar85]. Chord [BKW96, Ron86, Rut81, YJ84]. Chords [Gaa77]. Chromatic [JB91, GS95, LPVM13, VAWW10]. chrominance [dLAH07]. cine [WWJ13b]. Cinefilms [GBR79]. Circle [CT97, CL00a, Hor76, ALSR11, Chu77, Dan78a, Shi81, Kul79a]. Circles [CC01, Con88, Dor79, Fie86, Kul79b, KD81, MB79, NA84, WR87, Lea92]. Circuit [FWL88, HF80, ME98b, ME98a, SDR91]. Circular [BC85, CT95, CL00a, Dor84, HBA93, Jos94, KK83, Lan87, Lil97, Per81, Pla96, Sau93, TC95, Van84, Ami90, Ber89, HFF93, TC89]. Circular-Arc [Van84]. Circumscribing [DBB83, WL85]. Circumscription [O’R85]. Cited [Ano07j, Ano07k, Ano08p, Ano12q, Ano13r]. City [SJ01, IZKB12, JBWK11]. clamped [CXY^{+}09]. Clark [SS11a]. Class [Bid91, FL87, JLD12, KK79, MCPB99, PLS97, SS79, VLR84, CKLP09, CP09, Fon90, Pen03]. classes [ZYXZ13]. Classification [Bai88, BBC00, Chi81, DT09, DF02, Dav93, DWX83, FY85, GR92, HdVL99, HSD85, HB98c, HS83, JW94, KBZ96, KC92, KdVL99, LL97b, MCPB00, NFJ93, PZ92, Pel84, RW88, RFLSA11, STEK96, SHJB^{+}83, SL99, SW83b, SC98, TS00a, WS89, WLH85, Wei92, XL98, Yam80, DL10, FFM05, GHHX04, HL13, KORC10, KSS08, LLC11, PSR08, RRR11, RSS07, SB13, SYPK13, VMP03, ZZL13, dSdSF^{+}12]. Classified [SYF99]. Classifier [GK95, Hsu79, LLC11]. Classifiers [Lor83]. Classifying [AO04, Ros00a]. clay [DC04]. Clipping [Van84]. cliques [Eva11, PL08]. cloning [FSF07, Pan03]. Closed [ASS97, BB83, KPPK09, PCV94, Ree84a, RH85, SW86, Thü03, BGK95, Eva06, HA03, KG82, NRJ11]. Closed-Form [BB83]. Closed-world [KPPK09]. closely [DN91]. closeness [Mae90]. Closest [GSK02]. Closing [Boo79a]. closings [RH91]. closure [WWLV11]. cloth [UK12b]. clothing [CGW^{+}07]. Cloud [LH84, LM95, MS09, MPST08]. Clouds [Max86, CLK09, JXCZ13, MLF^{+}12, NBPF11, PLL12, ZN13]. clues [GSV05]. Cluster [Ris89, YXYW00, LZLP10]. Clustering [AW98, MH79, NC93, Pha01, Sto87, TB99, WF02, YYL98, AS09, CSY08, CFYU12, Cel90, CD13, FLHK08, HF11, JXC^{+}13, KBN12, MTG07, MMK04, RM03, TVC09, VAWW10, WSSS13]. clustering-based [VAWW10]. Clusters [ZF94, ZH79, SH09]. Clutter [WWHL88]. Cluttered [AA93, RKK^{+}00, AM04, Ano06m, GKK05, LBNS09, WRKP05]. Co [CP79, KP96, PA10b, THCG84, ZT80, GK90, LPVM13, Pec91]. Co-Occurrence [CP79, ZT80, KP96, PA10b, THCG84, GK90, LPVM13, Pec91]. Coalitional [DPT07]. Coarse [De 93, SY10, TB99, ML13]. Coarse-to-Fine [De 93, SY10, ML13]. cocycles [GDIIHK11]. Code [GV78, VS82, SGS^{+}10]. Codebook [JJM95, Mcl93]. codebooks [vGSV^{+}10]. coded [Sau91]. Codes [BBC00, Cai88, KLK88, Lam84, Maa94, SBT85, vdWvO96, BR90, CL90, LF82, SW94]. codeword [ATC^{+}13]. Coding [AB88, Bur80, CC97, Ced79a, EC88, HH97, KLK88, KG01, KU95, KCM85, KSC97, LC85a, Liu97, Mis84, Nad84, PR79, RP88, SP97c, Sob78, YB01, ŽA98, BRSSAL11, DF91, KYM13, LB04, OS81, San77, TD04]. Codon [RH85]. Codons [Ros93b]. Coefficients [CG94, SP97c, Yam79, Nad90, ZW03]. Cognitive [BBH^{+}12, Ham05, WWH07]. Coherence [LI00, SP92, MPF07]. coherent [KBD^{+}12, LYL10]. cohomology [GDIIHK11]. coincident [Tri90]. collaborative [PYS03]. collection [MSG10]. Collective [KS12]. Collective-reward [KS12]. Collinear [Cre99, DT96a, UTB^{+}11]. Collineation [CDH99]. Collision [VCVQ^{+}98, KG90, Sub90, TMT10, YR06]. Color [AMA79, AEM98, APV99, Ano94f, BFF97, BK07, BD02, CC97, Cav87, CYH94, Dre94, Ger85, GFS04, GB96, GB97, Gou84, HGA86, Hen98, IP98, JC94, LH93, LL97a, LPVM13, LPV07, MVP06, MTG07, MKK02, MM92, MV86, iOKS80, PS00, RY95, RPTB01, STEK96, Sap97, Sav87, SS95a, SG11, Shu97, SB79, SGK00, Taj83, UA77, VMP03, AQ09, ASVO12, BL04, BS04b, BH12, Cel90, CGG91, Dre96, HC13a, HO76, HWW06, HSJS10, HKK08, JWG04, JOvW^{+}05, KGU10, LLR10, LL04, LEB07, LMC09, LL08, LN10, MWF07, MN06, MGPJ11, MGPF08, NN04, PA10b, PBG04, PS12, QAB^{+}11, SCE04, SF07, SKU^{+}09, SAC09, TLEF06, VSP06, YZ06, YCL07, ZZ07, ZT09, ZCF13, PA10b]. Color-Based [AMA79, BL04, BH12, LN10]. Colored [Hor84, Oka81, DR04]. colors [HGS08]. Colour [BS87, Ang07, BG09, CT10, CT12, DCFM07, GE08, Hei04, PKD07, VBS^{+}04]. column [TH06]. column-space [TH06]. Combination [HC77, ZW93, KL11]. Combinations [AK78]. Combinatorial [ABE98, Fra95, KMT11, NKPT13, DSdlH^{+}11, FWWT13, WDN^{+}12, BK03]. Combinatorics [BACL97]. Combined [HYJ11, HGA86, LV11, NKP11, SKSR08, VRKL13]. Combining [GMA83, GCPF08, Hei04, QKH^{+}12, TID07, TCMS04, TLEF06, DUSL94, GFL^{+}11, GJ10, HDF12, KP12, MMK04, UKH88, XP11]. Command [Sab76]. Comment [BS89, Bri86, Hor79, Liu93, Nag86, Bor91]. Comments [Dan78a, Har94a, Kov86, Pie88, Tho86, Uhr86]. Commercial [Oka81, BLT05]. commercials [GS06]. common [KZW12, Lio91, SRS11]. communicating [UM05]. Commute [DDWZ12]. Comp [OBH04]. Compact [AGHN94, FWWT13, HB98c, Hub12, LL13, SGS^{+}10, vGSV^{+}10]. compactly [OBS05]. Comparability [Bre01]. Comparative [Che00, FRDC06, FS80, LCZ^{+}01, TT91, BFRA12, FML12, HS06, JM09b, LMRMJ08, LCP90, OH05, PSE^{+}11, SCD11, SYPK13]. Comparing [Gle01, GJ10, Sha11, vGSV^{+}10, CU11, OJRT08, TN05]. Comparison [Fra89, HSSB98, KYM13, OS87, RFC97, SGB01, Ste01, WS91, MSR07, PBSG12]. Comparisons [BM86, Sha79b]. Compass [PC89, Rob77, Nad90]. Compatibility [Cha83, KF86, Yam79]. competition [MMV06]. Complementary [LL97b, LL08]. Complete [BNG02, DG01, DY98, LT81, TG95b, ZN13, ACS03, KM03]. Completion [Rut79, WH96, WZWT99, BF05, LA11]. Complex [CM95, Jon97, KLL84, LM99b, MS97b, RBA94, SP97b, VKP98, BP09, ÇÖD08, CT10, FL09, HY11, Hu11, HCN90, KV06, KN04, LL12, MJ11, RSFdM04, SZ07, TN07, XYW11, YR06]. complex-cue [LL12]. complexes [CD11, Cou13]. complexities [XO93]. Complexity [EC88, Oka81, Pag99, PCP02, Ric84, Ris89, WL85, LT05]. Component [ERW93, FWL88, Jon99, AT89, BRSSAL11, CCL04, DB03, HHWP03, HQN05, MR89, Nic95, RCVA11, Ros08, Sub90, SHS03, WLMG08]. component-based [HHWP03]. component-labeling [CCL04]. Components [CCS01, LSZ83, Lum83, THN92, VK92, YD94, AHDM10, ACH^{+}13, DBB13]. Composed [LER95, Lat97, LL12, WB97]. Composite [HZLM11, SL99, OBS06a]. Compositing [KW99]. Composition [GT84, Mat89]. compositions [RL13a]. comprehensive [ASVO12, kWwZ13]. Compressed [Maa94, Spi98]. Compressing [IS02, IA03]. Compression [BIP00, CYH94, CHB86, FST94, GSK02, JEK98, KDRC98, Liu97, Mor76, NK00, PW86, SRK02b, WB82, SBS04, TVLS08, Vás11, WLZW04, YWMS08]. Compression-Based [BIP00]. Comput [AK11, Ano06m, MBMC11, PZ09]. Computable [LR90]. Computation [Bla85, BM00, BM02, BA92, CM99a, CCP97, CH99, DB88, Fuj97, Ger85, Kan94b, KM00, Kro86, LHKC97, MKY01, Neg96, OS87, OD99, SA96, SLB^{+}00, SKS97, SA92, VY94, Wat87, WBR86, WBR88, YAT97, BMR91, CW02, DRAB08, FKV^{+}11, Kle13, MSI10, MN06, OH05, San90, TLCH05, XSD12, Ano95h]. Computational [BB11, CM94b, DB79, EOS84, Gri83a, II86, Kan91a, Kan94a, KM84, LZ97a, MJS97, MZ96, Pog85, RA77, SMK02, TB94a, Ter83, FFY^{+}04, HM13, KTP08, LD95, MR90a, MR90b, Pec07, Pos77, SGA12, VBS^{+}04, XO93]. Computational-Geometric [II86]. computational-morphological [LD95]. Computationally [MN94]. compute [PHK92]. Computed [Art79, HL79, Her80, JGR85]. Computer [ATN83, AMA79, ABMT87, Ano94f, Ano94i, Ano95b, Ano98e, BBK78, BY98, BS87, BS88, CFS98, CLR80, Cou81a, DRDKE13, DB94, Ede87, EOS84, FMV93, FHP01, För87, GBR79, HTEB11, Har86, Har94b, HSKH07, Her72, Jam09, JH98, Jol94, Kan94a, LV03, LC79, Lee76, LHKC97, MP09a, Mee94, MST00, MG01, MK76, MT00, NH92, Per81, Ree84b, Ros88, Ros92, Ros93a, Ros94, Ros95, Ros96a, Ros97, Ros98a, Ros99a, Ros00a, Ros00b, Ros01, SI96, Sch80b, Sch81, Sch82, Sel81, SK86, Shi94, Sup79, Taj83, TTG94, UA77, WH94, Wil79, Yam78, YKC^{+}86, ZXK02, Ano05p, AO03, DLP13, FR80, GC80, HO76, Hil83, HBH11, JS07, JBS^{+}91, JTEA91, KPKH07, KMT11, LBK10, Mar93, NLM05, PZ08, PZ09, PYS03, Ros89, Ros90, Ros91, Sah05, SBB10]. computer [Sha75, SFWG08, TCB^{+}08, WKP13, ZCCD06, LLE^{+}09, STLH08]. Computer-Aided [CLR80, Sel81, GC80]. Computer-Assisted [YKC^{+}86]. Computer-based [HSKH07]. computer-generated [JTEA91]. Computerized [SP81, SHS79, XL88]. Computers [LSZ83, Ree82]. Computing [AD93, Ano98e, AM97, Beu91, Bie87, BY98, CXY^{+}09, CGL92, DT96a, DGZ12, Dav79, FK00, GK98, HKM12, KK88b, Kri92, LL86, LH99, MP03, NWP97, PA83, RS88, RS91b, SSN78, Tan99, TG95c, WJG02, WZWT99, AM93, CKK^{+}12, FYH11, JBK04, Pec91, QHXC12, SRS11]. Concentric [Con88]. Concept [GY88, Nag78, KYM13, KM03, THL13, USKB10]. concepts [LDC^{+}13]. Conciliating [IJDAB13]. Concurrent [CTE95, Lea92]. Condition [RM02, CCS05]. Conditional [SKM06, PV13]. Conditions [Bla85, CL00b, HA93, LTS93, OD01, OK04, ZJ05]. Cone [AAV96]. Cones [Sor81, Tan89]. Conference [Ano95b, Ano96g, LV03, SRK02a, Wyv03, Ano96c]. Confidence [Neg96, KN11, PMC13]. Configuration [KKO98, OD01]. Configurations [MRF96, Tan99, TZM98].

Confirmation [LF83]. confocal [KGK10]. Conformal [KG94]. Conforming [Spe97]. Congruence [LTS93]. Conic [Boo79b, Pha89, Sam82b, WJW94, Alb74, MG95b]. Conics [CD92, KL93, QV98, BA06, KGK10]. Connect [OBW87]. Connect-The-Dots [OBW87]. Connected [BBB11, Hei99, Her98, ITN84, Jon99, Kub84, LR90, LSZ83, Lum83, MR89, Mon84, Ron86, ROH88, SUO00, SU01a, THN92, AHDM10, HQN05, HQW^{+}12, Nic95, NFU02, RCVA11, SH09, SHS03, ZUS06]. connected-component [HQN05, SHS03]. connected-component-labeling-based [RCVA11]. Connectedness [Lat93, SU01b, US96, CUSZ07, CU10a, CU10b, CU11, MVP06]. connecting [GBL08]. Connectionist [Fel85]. connectivities [BNG05]. Connectivity [BDHM09, BNG02, Cha79, Ede87, KDK78, KADS02, KG01, LR90, Ma96, PA98, WB97, BNG03, GA91, O’G94, SDC04, ZS09]. Connectivity-preserving [BDHM09]. Connell [Ano94h]. Conquer [OS95]. Consecutive [Muk97]. Consecutivity [Cha79]. Consensus [CM97, LZ97b, Min94]. Conservative [YG07]. Consideration [SKOS95]. considerations [LTT91]. Considering [OD02]. Consistency [Gri83b, OMLL98, SF97, CBT^{+}04, CK09, FL92, MM06]. consistent [CPC08, JLD12, TY05, UK12b]. Conspicuousness [WV78]. Constancy [BFF97, BJ97, CT12, SAC09]. Constant [ACWK06, BR93, MS96c, Sch92]. Constant-volume [ACWK06]. Constrained [IP98, JW94, LBSP02, Ols99, Ull79, ZCL99, dFP92, LPR^{+}03, MFG10, Oli91, SMD^{+}08, WWJ13b, YZT^{+}13, ZW03]. Constraint [BZ99, Gle01, Jon97, Sch86, SM94, WY91, BHMB10, MZC^{+}05, PL08]. Constraint-Based [Gle01]. Constraint-Satisfaction [BZ99]. Constraints [Bri84, DM01, FL96, FB97, Gri83b, Hob97, JT86, Kan88, Pla92, RH85, TTA94, TP92, UN91, WN87, Zha97b, DQ05, FF09, FK09, IJDAB13, LB06, NNT11, NDO09, OCVV04, RC03, TR09, WDB12]. construct [HB91]. Constructing [BNG05, KZW12, Mil79b, TCH07, VTG95, Eva06, LH95]. Construction [CSDC96, SS87, SB90, Sze93, SRK02b, Td93, WW94, CH88, Far82, KS02, MK02, Sch06, ZZC^{+}13]. Constructions [SB87]. Constructive [PASS01, TC87, Vee97, MD82]. Contact [Bri98, Tan99, BHBF10, NLM05]. Containing [FT79]. Containment [Kal82, Gho90, HT89, VW80]. Content [BZS08, BS99, DCCL99, DRK03, GH08, GWCO11, JEK98, MBKB02, Mil80b, PBQ99, PA10b, SLST99, SBK^{+}99, SPK^{+}02, Ano13v, AO04, Hei04, ILRB04, KMBH09, LL12, MSG10, Pen03, WZ04, XG08b, YJC^{+}09]. Content-Based [BS99, DCCL99, JEK98, MBKB02, PBQ99, SLST99, SBK^{+}99, SPK^{+}02, DRK03, GH08, PA10b, Pen03]. Context [GB10, GDR04, RW76, RW79, CL08, HMF10, JYTK11, KK07, PSE^{+}11, PL10, WMBY12, YZY11]. Context-dependent [GDR04]. contexts [FYH11]. Contextual [ES81a, DFP^{+}13, SKM06]. Contingent [Sav87]. Continuity [Alg83, Bra94, Far82]. Continuous [AM97, BA92, Bra94, Cha83, GT84, GGR01, HQ12a, HQ12b, KR85a, LM00, ZZL13, CGR13, Eva06, JJN76, PV13, RCG^{+}09, TMN06]. continuous-discrete [PV13]. Contour [AM00, Arc81, AM78b, ASZ99a, BGLSS04, BM98, BR90, BMM^{+}07, Cai88, CM99a, CL98, CS98, DWS83, Dem96, DY98, Fog84, GPP88, KCM85, LL99, LF83, LAL^{+}10, MA83, NI82, Pet99, PR92a, Sob78, SM94, YTTT83, BB03, CCL04, CLHW94, Coh91, DT09, DS07, KG82, Lio91, Mig12, OS81, PDTE06, SD92, WO10, YLA09]. Contour-Based [CL98, DWS83, PR92a, BMM^{+}07, DS07]. Contours [CP99, DM01, DS87, FK99, GV78, JDP97, KMB97, Kd88, KSd88, KSS00, KD96, MPJN87, Pla96, Sau99, SGHM00, SC00b, SK84, VKP98, Wei88, WM92, Woj84, YLWY92, ZM96, Alb74, Beu91, KS91c, Mil09, MBMC11, SZ07, VRKL13, WS90, WW91, WWJ13a, XAB07]. contraction [JXC^{+}13]. Contrast [BL89, Gau92, Leu92, Mok92, MC95, Ney93, Ric84, SC97a, ZCL99]. contributed [IZKB12]. contribution [JOvW^{+}05]. Control [BG80, DCTO97, HER81, JH98, KKH96, LN85, Ley87a, MGMS01, PJ88, TM94, VFV93, WZ97, BG79, BBH^{+}12, Ham05, LM89, TKPR09, TM07b, VSR12, ZCCD06, Ano94f]. controllable [YL08]. Controlled [GS99, TTG94, BBB96, Reb89]. Controlling [WH00]. conventional [BPS10]. Convergence [Bra94, CRC97, GMT00, SK98, Wil89, YYL96, BE11]. Convergent [Mil79a, Bar05]. Conversion [Fog84, Sam84, LS91]. Conversions [Ül01]. Convex [DBB83, EKH01, GK98, Han88, KŽ99, OCON82, PW91, Rew84, Rob96b, Ron86, SR00, SB90, WL85, Bha91, DBB13, GS90, GBW89, HZLM11, Lan91, SM90]. Convexity [Cha83, Gaa77, Kis96b, KK81, LL99, MMS97, TY01, Mae90, RM06]. Convexity-Based [TY01]. Convolution [DB76, GV84, Hub12, PS83, WBB85, Fiu91a, KK95, MW91, SB89]. Convolutions [O’L88, Nad90]. Coons [For72]. Cooperating [CA97]. Cooperative [DC00a, LYA13, MLH13, UM05, ZKRH04]. Coordinate [Big97, LS01, Ül01]. Coordinate-Invariant [LS01]. coordinated [PKK^{+}09]. coordinates [JF10, ZHM11]. Coordinating [WWH07]. Coordination [BH86, YCKA10]. Coping [Wil98]. Coplanar [CRC97, QV98, Bar05, ODD96]. coplanarities [FK09]. copy [HO76]. Core [BC88a, KL10, TMT10]. Cores [MPPG98, PEFM98]. Corneal [GAD01, HSSH89, ZMCA05]. Corner [BY08, FT98, MY87a, Ros96c, Ros99c, LD90, MM06, RSv89, SS90b]. Corners [Dem96, SJ84]. coronary [YHS89]. Correct [LZ97b, DL10]. Correcting [YF80]. Correction [Gos89, McL96, PM97b, SKU^{+}09, Yan93a, Che08, MUS06, MKA73]. Corrections [JS87]. Correlation [HSD85, KC99, Yan93a, AS09, LRW08, LY06, MCF10]. Correlations [THO94]. correlogram [ZT09]. Correspondence [Chu02, CPD93, EW87, Jon97, JB91, Jur99, KG94, iK85, KHB01, Mit88, RS91a, SA96, Web83, LZLP10, LH03, MEYD11, PMW05, SAS12, XJK12]. Correspondences [CA97, CH99, KWK94, LH88a, LH88b, SBZ97, Tay00, CDT11, LH90, PW06, ZN08]. Corresponding [WB01, Fit88, Sha11]. Corrigendum [AK11, BM02, CYW04a, HQ12b, MBMC11, PZ09]. Cortex [CS82, GR92, Wat87]. Cosine [KSG84, Liu97, LL08, WLH85]. cosmetic [BHBF10]. Cost [FK00, CL91, KHH^{+}12, MSI10, MR90a, MR90b, MEYD11]. cost-optimal [CL91]. count [HBB^{+}12]. Counterexamples [O’R85]. Counterparts [FKW98]. Counting [AGW85, JP78, Mil99, MFA89, OR81, PS94]. counts [KRJ^{+}08]. Coupled [CBM01, YS09, GFW13, JXC^{+}13, SHC^{+}12, TRG^{+}13]. coupling [TMN06]. COV2 [Ano07e, Ano07f, Ano07g, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m]. Covariance [KWK84]. covariances [YO11]. Cover [LH84]. Coverage [Fiu91a, TG95b, ES06]. Covering [CM99a, HW94]. covers [Eva06]. crack [IO09]. Cramer [CT95, Kan98, TC95]. Cramer-Rao [CT95, Kan98, TC95]. Crease [PSK^{+}02, SLS01]. Creaseness [LLSV00]. created [SYPK13]. Creation [Ros10b]. Crest [MAM97]. Crested [Sch80a]. CRFs [YHN11]. Criteria [Bra94, IW97, Kim04]. Criterion [GCB92, GPP88, Kaw79, GBHS06, WX91]. critical [GB10]. Critique [Oli00, Oli01]. Crop [DWX83]. Cross [BCL96, CP99, EU85, Kan94a, KSd88, LCC89, LF98, NS91, THO94, Yan93a, Yau84, AWK04, Boi88, MCF10, WHN08, YC05, ZGK95]. Cross- [Yan93a]. cross-correlation [MCF10]. Cross-Correlations [THO94]. cross-lingual [WHN08]. cross-ratio [YC05]. Cross-Ratios [LF98]. cross-referencing [AWK04]. Cross-Sectional [CP99, KSd88]. Cross-Sections [EU85, LCC89]. cross-validation [ZGK95]. Crossing [MZ96, ZM94, DN91, SB91, WW91]. crossings [Yui89]. crowd [KB12, ZZP12]. crowds [CZZS07, GLOC10]. CRT [Yam78]. Crude [VV02]. Crust [ABE98]. CT [HRS02, Mar82, MDdMG09, SMD^{+}08]. CT-slice [MDdMG09]. Cube [CHC11]. Cubes [MS94]. Cubic [Buc88, CL00b, MY87a, PS83, PW86, SAG85, BE11, HN82, SB05, WJG02, ZK05]. cubical [Cou13]. Cubics [Boe82, SZ96b]. Cue [BL76, KR99, RJ00, RWWH00, EDB12, JC06, LL12]. Cue-Based [RWWH00]. Cues [LL97b, SLST99, SM93, CLZZ13, CRH05, CH06b, GW07, KN03, KSR^{+}12, KZW12, Mig12, NT10, ZTH^{+}11]. Culling [Med84, YG07]. cultural [dOSJVBS12]. Current [SK86]. Cursive [AHD98]. Curvature [DT97, FW97, FL92, Fog84, KG94, Kim97, Kis96b, Ley87b, LLSV00, MS96b, MKY01, Nac82, OD99, PSK^{+}02, SF97, Wil98, WM92, WS93, WW93, FB12, HSSH89, LR12, MSR07, Mil89, Yui89, YL90]. Curvature-Based [KG94, WW93, FB12]. Curve [ASS97, ABE98, Bar84, BP84, CW94, DB88, Elb01, GK77, GO94, II86, JXC^{+}13, Kan98, Kaw79, MY87a, Nis96b, Ols99, Pha89, SSF94, SB96b, SZ96b, SKS97, WW95, Yam78, YKC^{+}86, ŽA98, vdWvO96, Bot78, Cha74, CW02, CR90, LY90, SS78, SdB03, VCT09, WWWM12]. Curved [AKL93, EPRR79, For72, KOY86, KHB01, LT90b, ST96, Tan99, VKP98, WP93a, GW90, PHK92, Wal89]. Curves [Ano92a, Ano95h, BKD01, CV94, D’H86, DR93, EK97, EKH01, ET94, FAB97, GLR^{+}99, GW93a, IW97, iK84, LKE98, LM99a, LP79, MH98, Mok97, Rab92, Rei96, Rob85, Ros93b, SC99a, Sau93, SAG84, SAG85, SS95b, SZKD99, Sha78, Shl83, WD84, WV97, WZ97, Wil81, Zha99, BG91, Far02, GW93b, GS12, GCB90, HB91, HN95, KT89, LL13, MK02, MD95, OBH04, OH04, Ram72, SRTBS91, Sha75, SA81, Thü03, WX91, WJG02, Wil78, XWYY10, YZZ^{+}10, tHV09]. Curvilinear [HP96, SS84b, LCZ09]. Curvilinearly [Pra81]. cut [CUAT13, DK13, GPDR13, KT08, PSF07]. cut-graphs [PSF07]. cut/max [ZSCP08]. Cuts [LS94, CPP^{+}11, XAB07, ZSCP08]. cutting [BGTG04, FML12]. CVIU [SMHH04]. Cycle [RSB93]. Cycles [CM99a]. cyclic [TAK09]. cylinder [AO03]. Cylinders [Bid91, HC94, SB84]. Cytological [Mey86].

D [Ano94g, Ano01s, AS08b, BCF06, CYW04a, CFM^{+}13, FAB97, GSPL10, LEA^{+}10, MBMC11, AHRW87, ACF00, AKL93, ACG^{+}09, ÁB13, AS08b, AM97, AOR94, ARARCE11, ACDB12, BM99, BIP00, BBC00, BI10, BI11, BCA98, BL94, Bar05, BLT05, BAM87, BT05, BD94a, BM95, BJ86, BR95, BY12, Bd96, BRdBS99, BZ99, BCF06, BS04b, BGK95, BF05, BS00a, BSB87, BDL^{+}06, BH95, COW98, CGH08, CM12, CCH91, CK11, CVB09, CYW04b, CS89, CJ93, CSDC96, CS98, CK00, CYNO11, CC11, CA86a, CLCO13, CL00b, CFM^{+}13, CP91, CCA92, CC96, CG04, CS00, CPS10, DT96b, Dam08, DSdlH^{+}11, DWB11, Dan97, DB03, DF01, DPR92, DAM12, DN82, DLP13, DSY10, DBB13, EK98, ES04, EM96, FPC^{+}08, FWH13, FBF08, FF09, FHMB84, FRL^{+}98, FDMA97, FAB97, FKL^{+}98, Fly92, FST94, FL96]. D [GA91, GSPL10, GHMT09, GR87a, GSV05, GW07, Gui98, Gui99, GPC^{+}10, GB93, GSK02, HFKN97, HB98a, HASS10, HRS02, HR99, HB98b, Hen98, HHI95, HGSM11, HG11, HMF10, HMD93, HGB98, IAP^{+}11, JDP97, JGR85, JC98, JH98, Jok98, JM92, dOSJVBS12, KMB97, KL93, KWK94, KS91b, Kd88, KK95, KS04a, KM03, KMA^{+}00, KMGC84, KMN11, KNO^{+}09, LCT09, LM96, Lat97, Lau97, LF83, LC85b, LR90, LK91, LP91, LPR93, LKC94, LPS^{+}11, LST13, LS92, Li92, LS08, LL13, LM95, LCC89, LSHT02, LT90b, LS12, LSTF12, LK00, Luc01, Ma94, Ma96, MS96a, MW00, MFJ95, MC09b, MCB13, MMA06, MST85, MWTN04, MCT10, Mer88, MN95, Mil09, MKY01, MDR91, MB95, Mur87, NSK^{+}97, NG98b, NT10, Neg12, NKP11, NFA04, NKPT13, NL96, NDO09, NSEA13, OG98, OBS05, OMBH06, OJRT08]. D [OCVV04, PK99, PPK93, PSR08, PS07, PMW05, PCV94, PF87, PHK92, Pot87, PBN^{+}09, Pud98, QL96, RAH97, RG12, RWWH00, Rem04, RFLSA11, RH85, RNDA13, SAA93, SC96, Sal90, SCD11, ST96, SI96, STV09, SM06, SN99, SGHM00, SCS91, Shi99, SKU^{+}09, SP97c, ST10, SKVS13, SB00, Ste01, SWS11, SRHC13, SRML09, SKBS13, SS11b, SB02, TB99, TLGS05, TCH07, TCMS04, Thi92, TG96, THO94, TN05, TN08, TML00, TTIM96, TSK94, TH04, TF81, TCCK90, TA88, THL03, Udu81, Udu82, UK12b, UFF06, VV02, Ver81, VPAM12, VKP98, WSV91, WW94, WCZ02, WPS03, WWLV11, Wei88, WEY06, WWB84, XOF05, XL88, XP11, YC78a, YB07, YHR^{+}05, YT99, YH83, YK95, YC98, YYF89, YJC^{+}09, YL90, ZW97, ZP11, ZSCP08, ZZZL13, ZF94, ZXY^{+}12, ZZC^{+}13, ZZLZ13, ZZZY13, ZH04, Ziv10]. D [tHV09]. D- [FAB97]. D-based [GSPL10]. D-image [LS12]. D-Modeler [BD94a]. D-objects [PS07]. D-range [LS12]. D-rotations [RFLSA11]. D-Space [HR99]. D/ [CFM^{+}13]. Dagstuhl [GHPW12]. dandelion [LYG07]. Daniel [Ano92b]. Dashed [JvdBS99]. Data [AS83, AK78, AF81, AK96, BCA98, BL98a, BW76, Boo79b, BZ99, BS00a, Bri98, BS00b, BKA84, CKB96, CHB86, Chi81, CPK99, CN87b, Ein83, FS95, GS92, GKR02, Gro84, Gul79, GSK02, Har80a, HW81, Hob97, HMD93, Hsu79, Jac01, JC93, Kan91b, Kan94c, LR02, MBK81, Maz87, MM92, Min79, MAM97, ND92, Nis98, NWP97, OM84, OK07, PM89, Peu79, Peu83, PCV94, PW86, PS00, RAH97, RN93, Rob85, RF02, SAA93, Sam82b, Sch78, SHG^{+}88, Sha79a, SK88, SB00, SM97, Tan81a, TP75, VW80, VTG95, WW94, WLZW04, WALL00, WWB84, YD94, ZW93, ZOMK00, AM06, BBF^{+}11, BC10, BR12, BYN^{+}04, CH06a, CBT^{+}04, CD10, CP09, CC96, Cre08, FWWT13, FLHK08, GLOC10, GM90, HF11, JBC08, KK95, Kim04, Klu78, KS04a, KSS08, LL06, LB04, LY13]. data [LPR^{+}03, MSR07, MSS90, MK05, MC09b, OBS05, OBS06a, Pat13, PPT06, PMF90, QT10, RH06, RKG03, SY10, See89, Sez90, Sha11, SKVS13, SRHC13, TG11, TFL^{+}09, TBN95, TN05, TN08, Tri90, TZY08, VS08, VD90, WS08, WNH05, Wu93, YWMS08, YW07, YG07, YJKK91, ZZZ06, ZZ10]. Data- [CKB96, SM97]. data-driven [TZY08]. Data-Structures [AK78]. Database [BS99, SPK^{+}02, Tan81a, DR04, MTAA11, YAK^{+}08]. Databases [ADDK99, Gro84, KAES99, KR98, MK01, SBK^{+}99, GDR04, PA10b]. datasets [CCFC13, Kim13]. dead [Gre04]. Dealing [TO99]. Dean [Ano94g]. Deblurring [CT88, HKZ87, MRW^{+}97, Mar90, Mis84, Cha91]. Decade [Boo97]. Decentralization [Fau81]. decentralized [HW07].

Deciduous [HdVL99]. Decision [KCA81, RM98, HPvB^{+}10]. decomposable [CKK^{+}12]. Decomposing [DMMP03]. Decomposition [Ada93, Ahu83, AK78, AD84, BL92, Fau81, KD76, KC92, LL99, MW91, Maz87, MK01, O’R82, PCR86, PS95, RFLSA11, SW05, ZH86, BFR13, DAM12, JBK04, RB89, RDM^{+}11, SH09, SKS11, TMT10, WSC^{+}12, XYW^{+}08]. decomposition-like [DAM12]. Decompositions [Gad91]. Deconvolution [Lor83, KHS94]. decorative [Elb05]. decoupling [BDVK10]. Deep [SWYP00]. Default [VY94]. Defect [Mai76, CGAY13]. defect-laden [CGAY13]. Defects [Oka84]. Defined [DFP85, Gau92, LC79, AWC06, DR03, TWS06]. Defining [Bid92, CU10b]. Definition [ACF00, AM94, Cha83, MMP85, SU01a, US96, YK80, DBF04, KMBH09, Dam08]. Defocus [ZD01]. Defocused [RC97]. Deformable [BCA98, CCA92, CYES00, Dav97, DJG01, FB97, GSP02, Han88, KH98, LT05, LKE00, NFSK97, Pet99, RAH97, SY98, TI01, TC11, WRH97, BPB13, CBC^{+}07, CMD06, DQ04, HW06, KMBG09, ML13, MSF^{+}12, NÇ10, SI03, SBA13, SRHC13, TMT10, WB12, ZZC^{+}13]. Deformation [CSDC96, FK99, Gos89, KMB97, Nis96b, RW97, AWC06, FPC^{+}08, LB06, LPR^{+}03, Mar07, MWTN04, SY10, SKH08, WW90, WW91]. Deformations [FT98, LHH97, NMP97, RYN98, ASFP03, DC04, LF04]. Deformed [Nis97]. deforming [SOG09]. Degenerate [TZM98, MC09b]. Degradation [BHBF10, SK79]. Degraded [YK95, ZSN96, PS12]. degrades [HBF09]. Degree [Rab92, Sha11, ZW03]. degrees [LWLS12]. degrouping [ABD11]. Delaunay [DFP85, GYH13, dFP92]. Delaunay-Based [DFP85]. delay [NSEA13]. Deletable [Che98]. Deleting [TDMT85, Eva11]. Delineate [AM00]. delineated [Ano06m, GKK05]. Delineation [GR87b, SU01a, LCZ09]. Demand [Ken86]. dementia [HPvB^{+}10]. demodulation [WB11]. demonstration [KRK11]. demosaicing [dLAH07]. demosaicking [ZZ07]. denoising [HSJS10, MGPJ11, TH12]. Dense [FMR01, LSC08, XS98, HF11, IZKB12, WNH05, ZN13]. Density [BH99, PV97, Vel95, YKA01, Fit88, LCZ09, SRP10, WHM^{+}09, ZZP12]. density-image [Fit88]. Departure [Lee02, LY05]. Departures [SC00b]. Dependant [GKR02]. Dependence [SW83b]. dependencies [CHC11]. dependency [VBN11, XYW11]. Dependent [Chi97, CN87b, KWK84, OYTY98, Ree92, SY98, GDR04]. Depth [BL76, BK83, CH85, Che91, CP04, HZ86, KB91a, MNE00, Nev76, RC97, TM94, WL88, ZD01, JC06, PCR^{+}04, SB96a, SSL^{+}12, SCS91, SRML09, SKBS13, Wei90, WNH05]. depth-encoded [SKBS13]. Derivation [JW87, WN86]. Derivative [SSF94, DS90]. Derivatives [HS87, Lan84, WW95, MB95, Sub90]. Derive [MA85]. Derived [Nag83, DN82, LTT91]. Deriving [SYK96, Yam79]. Describing [BPYA85, RN93, KS91b]. Description [Ant98, Cai88, CM95, DG01, EF78, KW00, KD85, LN98, LL97b, MMP85, Mil79b, NP87, NB80, Pra83b, SRL82, SK85, ZH79, ASVO12, BGK95, BLH91, CL90, CGL92, CH09, CGG91, CNC03, FMGA^{+}12, Gho88, KCC89, KN04, MD82, SW94, TFB80, XHJF12, YJA96]. Descriptions [Haa82, KLL84, Ram76, WN86, YMA82, Nis96a]. Descriptor [DUC97, MWL86, HC13b, KŽ12, MPVF11, TG11, TWS06]. Descriptor-Based [DUC97]. Descriptors [ANM98, EK88, GAD01, Ros96c, AVBK10, GK90, HOH^{+}07, LL12, PZX13, PG13, PS12, WW80, ZZL13]. Design [BBK78, BW76, BS00a, BS87, CLR80, GB96, HJS89, Oka81, Oka88, PE92, SBB10, Sch93, Sel81, TST^{+}83, vdWvO96, AFH81, DQ05, FR80, GC80, SB89, TTF04]. Designing [DUC97]. Designs [CK84, CK87, LFMP13]. destinations [PHY^{+}11]. Destriping [HW79]. Detail [TKPR09, dBD98, RCG^{+}09, Rub82]. Detail-preserving [TKPR09]. Detailed [SJ12]. Detect [SM93, AVBK10, KS91c, ÜB05]. Detectability [TTA94]. detected [HBL^{+}11, UM90]. Detecting [Ano92a, CHP^{+}11, CC01, Con88, DT96a, GWT09, HS87, HN88, HCN90, IW97, iIK85, KT89, LW85, LB05, ST96, SJ84, SB87, SRHC13, SM99, WZ04, WW93, ZYT10, HRC09, Mar89, MG95b, RL13a]. Detection [BEH^{+}81, BB04, BP84, BT88, BCG95, BM86, BS00a, BO91, BP09, CK84, CK87, CL97a, Che98, CBM01, Che00, Chi97, CV94, CP81, CYES00, CSR83, DGH98, Dav76, DM80, DM81, De 83a, Dre94, Ebe76, ES81a, FD99, FTW81, FMR01, FC86, GS95, GPP88, GP85, GR87b, GJP96, HCHD01, HRS02, HJ83, HL78, HL01, HNR84, HT88, Hum79, HNR88, HNRR90, JMA79, JN93, JB99, JJM95, JB89, KA94, KMA^{+}00, Lan84, Lee02, LB98, LSMS85, LL97a, LN98, LJ89, Liu77, LD98, Loh10, Mai76, MML87, MGK00, Mcl93, MY87a, MZ96, Min94, MF77, NS98, Ols99, PSK^{+}02, PCJC98, Pat79, Per76, PH82, RY98, Rew84, Ris89, Rob77, RB92, Ros02, SP81, Sha78, SC92, She96, SJ89, Spi98, TW98, TZ82, TSK94, TZM98, THCG84, VMUO95, VCVQ^{+}98, WP93a, WF78]. Detection [XL98, YKA01, YBDC93, YW99, ZK81, ZBV93, de 83b, AZSVK05, ALK^{+}09, AHDM10, AFH81, BL89, BT05, BDS12, BBC^{+}07, BL09, BLH91, CSY08, CVP10, CWO^{+}11, CCYC12, dFCS93, CZZS07, Cum91, DLS^{+}09, DK13, DZL07, Dav75, DLF06, DD11b, FFM05, FLCdA06, GZP05, GS06, GSPL10, GG09, GHHX04, HK93, Hil83, HKK08, JWDF05, JYTK11, JXCZ13, KL07, KLL^{+}11, KS12, KYM13, KBD^{+}12, KL10, LMRMJ08, Lea92, LE09, LC88a, LAL^{+}10, MYC09, ML13, MJ88, MMP09, MDR91, MTAA11, NB10, OK04, PDK96, PZX13, PM82, PL10, PS05b, QKH^{+}12, RCTV12, SS78, Sau91, SJST07, Sez90, Sha75, SS09, SS90b, SOD10, SM13b, SKBS13, SMHH04, TMT10, TY05, TDK10, THL13, TCC90, VD90, VSP06, WJ07, WO10, WMBY12, WSKH13, Wu02, XG08a, YWZ11, YCA^{+}10, YGH11, YHN11, YGC13, YZ06, YO11, YJC^{+}09, YR06]. detection [ZS11, ZJ05]. Detector [BKD01, BS00a, CL00a, Har85, Mor81, SGB01, WM93, ZM94, vYB89, BA89, BPW91, FB12, KY06, RSv89, MAY^{+}10]. Detectors [Ber84, Bro78, HSSB98, KP00, Sha79b, Wil98, CHH09, DS90, DN91, MM06, SB91, USKB10]. Determination [BH95, BKA84, FT79, Lam84, LC85b, LF98, NÇ10, SBT85, TS86, Udu81, YK86, ZF94, VR95]. determine [HT89, LF82]. determines [ZGLG12]. Determining [BG79, BG80, Che91, D’H86, Hor74, HC13c, IH91, JB92, Kal82, Lev79, MA84, Pra81, Yac83, YH83, BPHB91, LH90, Tri90]. Deterministic [AO03, BGT94, ITN84, Mar93, GB13, KL11]. Developable [CLL^{+}99, LTS93]. development [ACS03, Cre08, Pos77]. Developmental [NR88b]. Device [LPF78, Bad77, NLM05]. Devices [Ced79a, Fra79, Hor76, Piz81, Dan78a, HSH07, SE11, Kul79a]. diagnosis [TDK10, YJKK91]. Diagram [BA92, Jar77, KSI98, LSMS85, KZW12]. Diagrams [FWL88, OS95, RM98, Sug93, AdVDI05]. Dialog [Kov86, Ros86b]. diameter [KŽ12, RNDA13]. diameter-based [RNDA13]. diamond [BFR13]. diary [RCJ^{+}13]. Dictionary [CWH^{+}13, GCPF08, ZZL13]. dictionary-based [ZZL13]. diffeomorphisms [Mar07]. Difference [JJ94, KD86, TMNM09, WJW94, YBDC93, ZXY^{+}12]. difference-of-Gaussian [ZXY^{+}12]. Differences [SJ93a]. Differencing [YMA82]. Different [Con88, KHB01, RWV95, SB85, Shi99, TS01, BKK11, CU11, FKS10, PBN^{+}09]. Differential [BCZ93, CCA92, GL95, KPH02, KMGC84, SPW96, TD04, VB98, WW97, KM89b, LT90a, RMD08, TG95c, YS08]. Differentiation [KS95a, GS12]. Diffrequency [CL86]. Diffusion [AG00, CBM01, KS96, Max86, SLS01, TÉSK11, Whi93, WP93b, BI11, KGC05, LYSS12, WWJ13a]. Digital [AK85, Ano92a, Ano94j, AM78a, BR93, BS89, BN84, Bie87, Bog88, Bor86, Bor96, Bou79, BM80, Bra85, Bre01, Bri98, Cag93, Cha79, CD92, CT88, CS89, CL97a, FH84a, For89, Gaa77, Gar76, HM84, Har80a, Har83b, Her93, Her98, Hey82, Hor77, JP78, JM79, KCD00, Kis96b, KK81, KSŽ96, KŽ99, KR85a, KR85b, KR89, KCM85, LM00, Lat93, Lee83b, LPR93, LSBG92, LP79, ML00, MML87, MM81, Mar82, Mas85, MT84, MMP85, MF77, MV86, NA84, NA85, NA90, NS96, OM84, Pha86, PA98, Pud98, Rab92, Ree84a, Rob96b, Ron86, RKW91, RB92, RYN98, SR00, Shl83, SBT85, SDC04, SB02, TC86, Udu82, Udu94, Vos88, Wal88, WVL83, WB97, WLH85, WS93, ŽA98, BRSSAL11, BT05, Bor91, Buz03, Coe12, dFCS93, Dou92a, Dou92b]. digital [EL03, Eva06, Eva11, FLCdA06, GCB90, HAGR91, KT89, LA11, LF82, LP91, Mar90, NKPT13, OC90, SC96, SS78, SC93, Shi81, SRP10, VRKL13, ZZ07]. Digitalization [ASS97]. DigitalSculpture [MCQ05]. Digitization [GL97, KOY86, KD81, Mae90, San78, SDPO81]. Digitizations [GL95]. Digitized [Bid86, DS87, EPRR79, FSS84, McL96, SA85, WD84, Wil81, Wil84, Alb74, CSY08]. digitizers [KB91b]. Digits [Por00]. dilation [CYE91, HBF09]. Dilations [HR90]. Dimension [Cav87, CD11, DL97, Ull83, And03, BS92, BBB11, CP09, Coe12]. Dimension-independent [CD11]. Dimensional [Art79, AT83, AD86, Bat84, CA86b, CJ82, Col77, Dou81, ES81a, EU85, FH84b, FS84, FSSL86, FL87, GY99, GR85, HP84, HD97, HGv87, HL79, ITN84, JT80, KK79, KK88b, Kle85, Kri84, LZ97a, LMKG85, LB97, LT81, LJ91, Liu77, Lum83, Lyn81, MA85, MG95a, MZ96, MNHO00, Mer81, Mit88, MN94, Muk92, NDC86, Nur86, PS97, PCR86, Pie88, RRS83, Ree84a, SSN78, SGS01, SF95, SCS99, Sko86, SS79, TK97, THN92, TQ97, Ull81, UCB13, WR93, WR96, WD96, WN86, YC78b, YAT97, ZM94, ZM96, Abu89, ASVO12, AFH81, AH08, BSMG05, Bha91, BEGB13, BKMV07, CR89, DBF04, DM12, GHZ^{+}13, Gar82, GU89, Got08, HQN05, KCD00, LB08, MS09, Nac82, PJW11, Pat13, SB05, TBN95, Thü03, UA90, UKH88, WW80]. Dimensionality [KAES99, RRR11, Zuc85, LLL13]. Dimensioning [DV98]. Dimensions [Bor84, Bor96, DP88, ET94, Jos99, MH98, Mul92, TML00, YAT97, CBT^{+}04, DMMP03, Dor89, Gol13, DC88]. Direct [Dre96, GMW83, GL98, HGv87, JK02, Neg96, SJ93b, SB90, BF07, HC13c, NH89, YJKK91]. Directed [Ano93d, RW88, Uhr86, BI11, CGW^{+}07, EKY08]. Direction [Cai88, GY99, PE09, Pra81, Tan89, VBN11, VY94, ACAAC^{+}08, CSS^{+}13a, Dre96, GWT09, HQW^{+}12, KM89b, YGH11]. Direction-Based [GY99]. Direction-dependency [VBN11]. Directional [AK78, BS00a, FD99, HC96, Pag97, Pag99, AS08a, LSPV04, SLKL11, TKL^{+}09, WS03]. Directionally [Kub84]. Directions [AT13]. disaster [KB12]. disc [QKH^{+}12]. Discontinuities [BGT94, Lan84]. Discontinuity [GP85, JB89, SP97d, Spe97, VB98].

Discontinuity-Preserving [SP97d, VB98]. Discontinuous [Nis96b, KS03, KMBG09]. discounting [BK07, SS11b]. Discovering [JEF^{+}12, FR11]. discovery [DHP08, LC09, MGPP11]. Discrete [AAS97, And03, BC85, Bog88, Bra94, BV99, BH83b, Cha83, CLR80, DRDKE13, Dor79, Dor84, Fra95, GGO10, HD97, Hen84, Her92, IE99, KII98, Kod76, KJRA96, Kul77, Kul79b, KK83, KC99, LV03, LL99, LL92, Liu97, MRW^{+}97, MMS97, Olk95, PZ08, PZ09, Pav80, Pel84, PW86, TG13, TL88, WLH85, YAT97, ZGLG12, ALSR11, CT12, HQ12a, HQ12b, PV13, Thü03, TMN06, VCT09, Žun03, LL08]. Discretization [CBK03, BTCH05, KM89b]. discriminant [ITNP12, SHC^{+}12, WJ07]. Discriminating [QV98]. Discrimination [AL99, DH00, Ekl79, HC77, KC87, LF79, Mok92, Mul88, SHJB^{+}83, LJ90]. Discriminative [GYTL09, LLC12, LSTARMB11]. Disk [Bad77]. Disks [Ada93, Dor79, Kul79b, ND97]. Disparity [BI11, Cav87, De 93, EW87, HH98, MGMS01, FJJ91, Gon09, JJT91, KN03, MSI10, PMF90]. Displaced [JK02]. Displacement [BB88, Nag83, PD83]. Display [Art79, AAV96, CS89, Fra79, GR87a, Gou84, HL79, Hor76, JGR85, JM79, LC79, Las92, NNT11, Piz81, PS00, RGC87, WC79, Yam78, Bad77, CD10, Dan78a, JJN76, Kul79a, Rub82, YJKK91]. Display-camera [NNT11]. Displaying [KK88b]. Displays [BC88a, BN85, GB96, KOY86, SGDP01, JJN76]. dissemination [Ros10b]. dissimilar [HVD^{+}89]. Dissimilarity [RPTB01]. Distance [ALK99, APV99, AS88, BS89, Bor84, Bor86, Bor96, BM00, BM02, CS01, Chu02, CM99b, DK79, Dan80, Egg98, ER96, KKK99, KSKB95, Kis96a, KŽ12, KTNO97, LHKC97, LH99, LL92, MMS99, Mas02, Mul92, NGC92, Pag92, Pag97, Por00, Pud98, Rag92, RW95, SWG02, SJ01, SWH84, SBA13, SB05, SB02, TV99, iTTF82, Wah83, WPR85, BMR91, BS92, BS04a, Bor91, BBB11, CCTCR09, CYW04a, CYW04b, CXY^{+}09, DT10, GH08, Gre04, HKM12, MGW10, MS10, NSEA13, PRR03, QHXC12, SS91, SW04, SCvW11, SCMS13, WDN^{+}12, dSdSF^{+}12, DC88, Vos88]. distance-based [BBB11]. Distance-Ordered [Pud98]. distances [Ang07, ITNP12, MS09, NSEA13]. distinctive [DDLP10, YK08]. Distinguishing [CHL05]. Distorted [YF80]. Distortion [Gos89, HJK02, KC87, PM97b, Tan76, CP04, KBJ^{+}10, TM04]. Distributed [HBA93, MM92, OMLL98, TM86, Ham05, IKST05, MCT10, SKS11]. Distribution [BRW85, Ell81, HB98c, TML00, YJ84, Coe12, FL09, FS03, JC90, Kim04, PKD07, PTE12, QAB^{+}11, QT10, TS11]. Distributions [Acu92, LH95, MS09, ZZAA92]. Disturbances [MPPG98]. Dither [Zha97a]. Dithered [Shi83]. Divergence [YAT97]. Diverse [WWW89b, DR04]. Divide [OS95]. Divide-and-Conquer [OS95]. Division [SRL82]. DNA [RG12]. Do [Hoc87, Pri86a]. Docking [SVS97]. Document [Ano96g, Doe98, KZ93, KB98, KH96, KDRC98, LPH01, Nad84, Nis98, Pav86, Spi98, Yan93a, CMH13, LDD09, O’G94]. Documents [BKMSR98, CB98, SHKP98, WWC82]. Does [Lau97]. DOF [SIT07]. Domain [Ano01s, BKMSR98, CCMW97, Luc01, ZD01, Hu11, KZW12, LBCA10, MR90b, NFSD13, TP05, WXRA07, YSD03]. Domains [DFP85, WBOL07]. Dominant [Spi98, WW93, KZ05, SS78, Wu02]. door [ESS10]. Dot [CCP97, SSN78, ZH79, AT89, BMR91]. Dots [OBW87]. Double [HA93, WR87]. Double-Step [WR87]. Doubly [Wam85]. Dougherty [Ano95g]. down [KMN11]. DP [SHKP98]. Drainage [OM84, See89]. Drawing [Fie86, FS80, JV97, Jar77, Peu79, SP97b, WJW94, NL90]. Drawings [BL92, CLD96, DL97, DP88, DV98, HQ82, KL77, LC85a, LCD97, MAN84b, PC99, TDMT85, Dor89, SDPO81, UN91, Wal89, ZXY^{+}12]. Drawn [BM79]. drift [RMD08]. Driven [CKB96, IW97, SM97, ABD11, BCM13, CYW04a, CYW04b, DM82, FAB12, RGA10, TZY08, Wor05]. driver [CPT07, TDT12]. driving [RCJ^{+}13]. Dual [Con88, Tau02a, WSV91, ÇÖD08, CT10, CS04, Hu11, KTP08, SKS11, WSKH13]. dual-point [CS04]. dual-tree [ÇÖD08, CT10, Hu11]. Duality [Ley87b, YL90]. Due [JMA79, BHBF10, BTCH05]. Duhamel [FS85]. duplicate [CHC11, JN09]. duplicated [ZTH^{+}11]. during [DLS^{+}09]. Durrant [Ano93d]. Durrant-Whyte [Ano93d]. Dyed [CEP84]. Dynamic [AD93, AT83, BPBS13, BB91, BBHF10, CSDC96, CS07, CC00, DQ05, FC86, GB13, GMW83, GSK02, HL78, JJ83, JT86, KAES99, KU95, LE09, LY90, LN85, LHB87, MA78, MS96c, Pie79, Pla92, RGC87, SGHM00, TW98, TS92, VR95, WPK09, XST04, YLM11, ZT98, ZKRH04, ALP06, Bar05, Bur81a, CLHW94, CH06b, DD11a, EL07, ES81b, GA13, GY05, HQW^{+}12, JBC08, KTP08, Lea92, LWH03, MSI10, MWTN04, MMP09, RM91, SCL13, SHK11, TN07, TMN06, VCT09, Vás11, WZL^{+}03, Wu02, XG08b, YR06]. Dynamical [FS95, OW84]. Dynamics [GM87, MJS97, TID07].

ear [AZN11, HNC05]. Early [HH97, Pog85, WBB85, Zuc85, AT89, SGS^{+}10]. easy [ZCCD06]. eccentricity [IAP^{+}11]. Echo [YTTT83]. Economical [OW86]. Ed [Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06b, Ano06c]. Ed.

[Ano07e, Ano07f, Ano07g, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m]. Edge [AGHN94, Ber84, BM86, BKD01, BS00a, Bro78, CN87a, CBM01, CY83b, CN87b, Cum91, Dav76, DM80, DM81, Ebe76, ES81a, For88, FC86, GMA83, Har80a, Har85, HSSB98, HS88, HLF^{+}97, HT88, JB99, JJM95, JB91, KA94, LL86, Lem79b, Leu92, LH92, LcTT91, Mcl93, MZ96, MH79, Min94, Mis84, MF77, MGPJ11, Mor81, NM79, NP87, PD79, PC89, Per76, PA10b, Pie88, PDTE06, Pra83b, RSB93, RM02, Rob77, RB92, SB93, SC97b, SSJ86, Sha79b, SC92, She96, SGB01, Shi86, Shn81b, SH84, VK92, Wil98, WS90, WM93, YH81, YBDC93, ZM94, ZW93, ZBV93, ZH79, de 83b, vYB89, BPW91, CGG91, Dav75, DUSL94, DN91, JM09a, KIK89, KM94, KM89b, KY06, LTT91, LMDB11, ML13, MJ88, MDR91, PM82, PMF90, SB91, Sau91, SS09, VR95, WO10, WPK09]. edge-avoidance [JM09a]. Edge-Based [HLF^{+}97, JB91, ZBV93, CGG91, PMF90]. Edge-Focused [RSB93]. Edge-Only [CN87a]. Edge-Preserving [CY83b, RM02, MGPJ11]. Edgel [NP87]. Edgels [HBA93]. Edges [HJ83, Ish84, LL97b, Med84, PR79, BL89, DN91, Eva11, Mor90, PE09, UM90]. edit [DT10]. Editing [Gle01, Jar77, TDMT85, Wal89, WXRA07, ZHM11]. Editor [Jar77, GSST03]. Editorial [Agg83, Ano01j, Ano01i, Ano01k, Ano05f, Ano05i, Ano05l, Ano05m, Ano06a, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06l, Ano07a, Ano07b, Ano07c, Ano07d, Ano07h, Ano07i, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano10a, Ano10b, Ano10p, Ano12b, Ano12f, Ano12g, Ano12h, Ano12k, Ano12l, Ano13a, Ano13n, BCM97, Kak95, Ano02g, Ano02h, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano05e, Ano05g, Ano05h, Ano05j, Ano05k, Ano05n, Ano05o, Ano06i, Ano06j, Ano06k, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano10n, Ano10o, Ano11a, Ano11b, Ano11c, Ano11d]. Editorial [Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano12a, Ano12c, Ano12d, Ano12e, Ano12i, Ano12m, Ano12n, Ano12o, Ano12j, Ano12p, Ano13c, Ano13e, Ano13g, Ano13h, Ano13o, Ano13p, Ano13b, Ano13d, Ano13f, Ano13i, Ano13q, Ano13j, Ano13k, Ano13l, Ano13m]. EDITORS [DCCL99, MT97]. edNLC [Fla89]. edNLC-graph [Fla89]. Effect [KC01, Lee91, THO94, YH81, KM89b]. Effective [dFCS93, LDGS^{+}13, CWO^{+}11, SSM06]. effectors [SRHC13]. Effects [CFA98, FT98, MPPG98, HC13a, KMP05, yKL11]. Efficiency [LHH^{+}98, PR79, WA87, Dye82, KTP08]. Efficient [Ahu86, BB87, BM00, BM02, CC01, CYES00, DOSD11, DG01, DMW10, FSSL86, FSS94, FKW98, GR81, GK04, Hob97, HP96, JB92, KK95, KB00, KG01, LC85a, LMR84, LSMS85, LP90a, LA05, MK01, MN94, OK04, Pag97, PZX13, PLS97, PG13, PA97, PL08, Ree80, RCTV12, RSS07, SKH08, SLB^{+}00, TGSH98, VD90, VCVQ^{+}98, VPAM12, Wal88, XOF05, XL98, Yan93b, ZK08, ZXY^{+}12, ZLH13, dM92, CBT^{+}04, CYNO11, DSS94, GRGB^{+}13, TLEF06, VAWW10, WW80, Wil78, XSD12]. ego [RN12]. ego-motion [RN12]. Egomotion [DT96a, DH00]. Eigenimages [LB00]. Eigenspace [CMW^{+}97]. eigenspaces [BWL04, EKY08].

Eigenvalues [SB98a]. Eigenvector [PLL00]. Eigenvectors [SB98a]. Elastic [ACLS98, AG00, Mos91, BK89, BL09, Far11, JKM07, NBDB04, RFS03, WR08, ZP11]. Elastically [Dav97]. elasticity [LV11, NÇ10, PDA03]. Electrical [Ble84]. Electron [Haw78, Haw82, Sel86]. Electronic [FWL88]. Electrophoretic [Sko86]. Element [SY98, TGSH98, ZH86, NÇ10]. Elementary [HH96, CKK^{+}12]. elements [BPG05, ITF06, SW05, SD90, TCZ^{+}12]. Elevation [Bri98, KK94, OM84, See89]. Eliminating [Kim04]. Elimination [BTCH05, CM99a, Goo92, GO87, Imm91]. Ellipse [Ros98b, Ros99b]. Ellipses [Fie86, MB79, Ros96b, Lea92]. Ellipsoids [KVW94]. Elliptic [KG82, LDGS^{+}13]. Elliptical [DGH98, GK77, WV97]. Elongated [HP84]. Embedded [EA95, AZSVK05, Bar05, CVP10, CKB10, HZW^{+}10, SBB10, VAWW10, YCA^{+}10]. Embedding [Ahu86, FKV^{+}11, GHZ^{+}13, XHW09, ZRKZ^{+}11]. Embeddings [SB84, KL07]. emergence [Ham05]. Emphasis [Mil80a, SH09]. Empirical [BKD01, FHP01, RPTB01, WSC^{+}12, DAM12]. enable [SSdVL06]. enables [TFL^{+}09, WRKP05]. Encapsulating [GSS00]. enclosure [VW80]. Encoded [Cou81b, KD96, NH92, PA82, Jea11, SKBS13, YLM11]. Encoder [MM81]. Encoding [BCZ93, KADS02, LMKG85, Mea82, OY92, PR79, RB82, Ram84, SLCP85, Tam84, TVLS08]. end [SRHC13]. end-effectors [SRHC13]. Endoscope [OD97]. endoscopic [HSKH07]. Endothelia [GAD01, ZMCA05]. Energy [Ano01s, HPB94, Luc01, MRF96, ACG^{+}09, BN90, EyGS11]. energy-based [ACG^{+}09, EyGS11]. Enforcement [AHD94]. engine [LEA^{+}10, SM10]. Engineering [DL97, DV98, EFF98, HQ82, NL90, PRW97b, SOJ^{+}95, Dor89].

Enhanced [BSMK13, GSP02, KS02, Shu97, ACDB12, HAKK91, KGC05, LSD^{+}07]. Enhancement [Bou79, CN87b, Fre77, FC86, HM84, HC77, HH96, Hum77, LBD92, Lem79b, Leu92, Lyn81, MBDB88, MY87b, MK76, Mok92, MC95, Ree92, Rew84, SC97a, Sau91, SWH84, Shi83, SLS01, SKB98, Stu76, TL78, Van77, WVL83, ZCL99, Ang07, BL89, Cho88, HWW06, HSJS10, Lan91, Ney93, PMR92, TKL^{+}09, YAK^{+}08]. Enhancing [Dem96, Hor77, LI00, Wam85]. ENO [SKS97]. enrollment [FBF08]. entirely [TN08]. Entities [HBA93]. Entropic [BLd95, Pun81]. Entropy [BGS83, KSW85, Min79, MR92, Shi86, Abu89, GHHX04, SE11]. Enumeration [RSB93]. Envelope [HGB98]. Environment [JW94, Nag78, SL85, CP09, LHM06, LY13, ST10]. Environment-Centered [SL85]. Environments [RRS83, AM04, Ano06m, CM12, CPS10, FPDK12, GKK05, GPC^{+}10, LS12, LA05, MP09a, NKB11]. Epicardial [YHS89]. Epiflow [ZN08]. Epipolar [KHB01, ACAAC^{+}08, CPC08, CKS^{+}05]. epipolar-based [CPC08]. epipolar-plane-image [CKS^{+}05]. epipole [LB10]. Epipoles [LF98]. Equal [CH78, Sch92]. Equalization [AF81, Gau92, O’G88, PAA^{+}87, SF96, ZCL99, BK07, PMR92]. Equation [CG94, KMI79, KS96, Sch86, CS10, MZC^{+}05]. Equations [CBM01, FH84b, SW86, VB98, VF96, WY91, CD95, CRT90]. Equivalence [CU10a]. equivalences [CU11]. erosion [CYE91]. erosion-dilation [CYE91]. erosions [HR90]. Erratum [Ano06m, KK93, OBH04, Ano99f]. erroneous [CX11]. Error [BVL02, BRP04, Jur99, KS95b, KOY86, KD81, OD02, Ros96b, SJ93b, SRT01, VS82, YF80, CPS05, LTT91, LD95, QAB^{+}11, SB96a, UTB^{+}11]. Error-Correcting [YF80]. Errors [CFA98, iK84, KW99, KB00, KON87, LZ97b, PR79, LcTT91, RFS03]. Establishing [KG90, RS91a]. Estimate [SC97b]. Estimates [Mil99, WALL00, YBDC93]. Estimating [BK01, BBD^{+}94, BFY00, Bur83, DGC12, GA09, Haa82, KC95, KRJ^{+}08, KH94, MC09b, PD83, QY02, RKH05, Shi99, TML00, TZM98, TZ00, WSV91, WSV05, Wec81, ZL01, LMC09, RN12, YSL11]. Estimation [AHRW87, Ano01s, ACB98, ACW96, BR93, BA96, BGK98, CT95, CD92, CSC96, CM94a, CP80, CL00a, CFA98, CEC^{+}80, CWC94, Dan97, DWS83, DC98, Enk88, FD99, FMV93, Fog91, För87, GM94, HD97, HH97, Imm96, Jos99, KS91a, KM94, KC87, Lan87, LK91, LB97, LB10, Lin02, LH88a, LH88b, Luc01, MS97a, MGMS01, Muk92, NDBT95, Ols93, PSK^{+}02, PP95, SA91, SP97a, SJ93b, SP97d, SA92, Sin87, Spe94, Spe97, SJB02, SM94, SK83b, Td92, TC95, WLD99, WS03, WM92, WS93, YH81, ZD01, AS08a, AS09, ACG^{+}09, AH08, BDVK10, Cha91, CSS^{+}13a, CS10, DM12, ES81b, EBN^{+}07, FL09, Gon09, HD09, HSH07, HBH11, HH12, JC06, JF10, KM89a, KHK10, KH90, KMN11, LSC08, LCZ09, LYA13, MSR07, MSSS09, MP09b, NT10, ODD96, PD05, PV06]. estimation [RDM^{+}11, RP08, RAC^{+}13, RNDA13, SRTBS91, Sau91, SM06, SO07, SRHC13, SM13b, TMNM09, TAK09, The83, TC89, TP05, TCCK90, UTB^{+}11, WHM^{+}09, WCF10, YCH07, YZT^{+}13, YA12, YC05, YHS89, ZBLS13, ZZAA92, ZZP12, ZDF10, dP10, dMFU10]. Estimation-Theoretic [SA92, The83]. Estimator [CT97, TZ00, CBT^{+}04, CYC10, Dre96, HBH11, KK90]. Estimators [DS87, MN94, Ber89]. Euclidean [BM02, AS88, BI10, BM00, BV99, Cou13, CM99b, Dan80, Egg98, ER96, KGK10, KTNO97, LHKC97, MMS99, SW04]. Euler [BN84, Bie87, IE99, LP91, Man84a]. Evaluating [BH12, Ste01]. Evaluation [BKD01, Che00, CY83b, CEP84, DL05, Elb01, FHP01, Fra79, GAD01, HRS02, Iiz87, LBD92, LCZ^{+}01, LPH01, Mas85, Ols93, PR03, RPTB01, VK92, WVL81, WY91, WWL92, YJ84, AFH81, BLH91, BG09, DL10, GE08, HYJ11, HMC10, HC13b, HWW06, LK03, LFL08, MO11, MM06, RN12, RLC^{+}11, SJST07, VD10, VR95, Wu93, YAK^{+}08, ZFG08]. Even [Rab92]. event [CGR13, HNB04, JYTK11, SM12, SMHH04, YLM11]. events [ABI^{+}04, DLS^{+}09, LCSL07, PSYZ13, RCJ^{+}13, TD04]. Evidence [ANM98, Mil79a, MYLP98, Ris89, Sav87]. Evidence-Gathering [ANM98]. Evolution [LL99, SC97a, SKS97, DCS05]. Evolutionary [CM94b, KBD^{+}12, RF02, TB94a, BPB11, SCD11]. Exact [Bid86, Far02, Hor84, DLP13, Fra81, KS89, Tri90]. Exaggeration [SKB98]. Example [CMB^{+}12, SMT04]. example-based [SMT04]. Example-guided [CMB^{+}12]. Examples [KLL84, PEF92, FFFP07, HXS09, XST04]. exemplar [OMBH06]. exemplar-based [OMBH06]. Exhaustive [Lin02, ALSR11]. Existence [CG94, Maa94, Fit88]. Expansion [RBA94, SB85, VF96, BKK11, KHS94, NH89]. Expansive [Jai94]. expected [MR90a, MR90b]. Experiment [DWX83, LFMP13]. Experimental [CK84, CK87, LCZ^{+}01, THO94, WY91, HF11]. Experiments [CH85, GMA83, KD76, PSWH84, Van77, Yam80, HMEB07, HKA13]. Expert [Ken86, Kov86, Mat89, Ros86b, She86, Tho86, Uhr86, Nag86]. experts [EKY08]. explicit [SOG09]. Explicitly [HFKN97]. Exploiting [CHC11, DDLP10, PMF90, WN87, Kui08]. exploration [OMW^{+}07]. Exploratory [BC92]. Exploring [Kui08, NF06]. Explosion [BY01]. Exponent [ZK01]. Exponential [AHZ96, Bar84, Bid92, BKA84, Wha91]. expression [CSG^{+}03, HOH^{+}07, LY06, LSCM03, MB11, SKVS13, SSS13, WY07]. Expressions [ADRY94, YB01, SHK11, SSS13]. Expressive [Pha91, NF06]. Extended [CTF^{+}98, KSS97, MC95, WB97]. Extending [GR05, Sta05, ZK05, ZK01]. Extension [BB83, FDMA97, GW93a, GW93b, MMV06, NS91, WB90]. exteriors [HBH10]. external [MLH13]. extract [MB95]. extracted [BY08]. Extracting [AB88, Cre99, CKS^{+}05, FKL^{+}98, KK94, RL93b, SC99b, SLY89, YC78b, FYH11, LP90b, YC78a]. Extraction [AK77, ANM98, AT89, AMMV99, AM78b, ADDK99, BC88a, CW94, Cap84, CCP97, DT96b, GR81, Gro82, GN98, HL84, HS83, KZ93, KII98, KS91c, KZ05, LSVD85, LAS94, LPH01, LHHC98, LB87, May99, MMP85, MNHO00, Mil79a, MK79, NMI79, NB80, Nis95, Nis98, OM84, PM89, RD93, Ree80, Rob96a, SCS99, Shi86, SS84b, SH84, SH77, TSP97, THN92, iTF78, Tou80, UZC97, Vee97, WWC82, WH01, YMA82, YK87, BB03, CM12, ÇÖD08, CR90, CNC03, DBF04, Dam08, DDWZ12, FWH13, FLCdA06, FS03, GHZ^{+}13, GK04, HXS09, HNC05, JXC^{+}13, KA12, LD90, LCZ09, LS09, MTG07, MZB^{+}10, MM90, MD82, PLL12, PQML11, PIK90, RC13, See89, Ste13, VPAM12, WWWM12, YT13, YR06]. Extractor [FWL88, Pav86]. Extrapolation [MR96, Kim04]. Extremely [SLCP85]. Extrinsic [LLSV00, PA13]. Eye [HP05, KMBH09, MM05, PE92, AZSVK05, HH07, JWDF05, NNT11, SFWG08, WSV05, WJ07, YC05, ZJ05]. eye-detection [AZSVK05]. eyebrow [LLC13].

F [Ano93d, Rab92]. FABRIK [AL11]. Face [Ano01q, BP94, CC03, HH77, HHWP03, HL01, KL07, LY06, MYLP98, RY98, SSN03, TTH07, YD94, YKA01, AM04, AC09a, AC09b, AKC11, ARARCE11, BC10, BCF06, BF10, CH06a, CFB05, DM12, EKY08, ESS10, FWH13, FBF08, GJ10, HASS10, Hu08, Hu11, HDF12, JLD12, KHA^{+}05, KMBH09, LRW08, LL08, MYK03, MCB13, PY08a, PZX13, PTE12, RM03, SHC^{+}12, SSM06, SKVS13, SM13b, TD04, WJ07, YCA^{+}10, YAK^{+}08, ZJ05, tHV09, BGPD09]. Faces [Ish84, RKK^{+}00, BL09, DBBB03, Kou03, ZKC03]. Facet [Coh85, HW81, PSWH84]. Facial [ADRY94, ÇÖD08, CSG^{+}03, FSF07, KdVL99, LSCM03, Pan03, TW98, YB01, DB03, GZJ05, HOH^{+}07, LB05, LY06, MB11, SHK11, SSS13, TLWT12, WY07, YLM11, tHV09]. Factor [Dan78b]. Factorization [SRT01, TI01, HRC09, LLL13, ZZ10]. Factors [BGPD09, CP09]. fairing [LBM04]. fall [ALK^{+}09]. False [NI82]. False-Contour [NI82]. Families [PSF07, ŽA98]. Family [AHZ96]. Fan [DB76, Gul79]. Fan-Beam [Gul79]. far [BBC^{+}07]. far-infrared [BBC^{+}07]. Farin [Ano95h]. Fast [BCMCB09, BDL92a, BM80, Bur81b, Bur83, CN87a, CH11, CR97, Coe12, CM99b, Egg98, FS84, GU89, GK95, GH92, HE81, HQN05, HH97, Imm96, IP98, KU92, KK92, KK93, KBJ^{+}10, Las92, LL92, LLL86, LCZ09, LK03, LB87, MWL86, MAP99, MPST08, MĆK09, NFSK97, NA79, OG98, OWW85, RM98, SW04, Sob78, Sup02, Sze91, TSR89, WD84, Wec81, WM92, WNH05, YAT97, YO11, YJKK91, vv92a, AL11, ARARCE11, BDL92b, BPB11, CL90, CBT^{+}04, CCYC12, CR90, CYE91, FL09, Fiu91a, Gar82, HDS08, HMA10, HZW^{+}10, LZLP10, MDdMG09, MU11, SW94, SB89, Tan11, TBN95, YB07]. Faster [Dan81a, BAP08, CDLD77]. fat [BE11]. FBA [FSF07]. FCC [Kim13]. Feasible [WSSD96]. Feature [Bai88, BC88a, BL98b, GHZ^{+}13, GYH13, GR81, HR99, Hum79, HS83, KSS97, KN99, LCD97, LSVD85, MFJ95, NB80, NFSD13, Nis95, Nis99, PLL00, PM89, Pav86, PBQ99, PM97a, RD93, Rob96a, RWV95, SB98a, Shn81b, SS84b, SH77, Tan76, TTA94, THN92, TS01, Tou80, TPR^{+}00, WF02, WC92, AK91, Big90, CBD^{+}03, CM12, ÇÖD08, CLC91, CWO^{+}11, CYNO11, CP09, CK09, DOSD11, DDWZ12, DG11, FWH13, Fly92, FYH11, HYJ11, HS89, HNC05, KGFP10, KS04a, KYM13, LK03, LFL08, LS09, MR90a, ODD96, PZX13, PLL12, PQML11, Pun03, QT10, SB13, TY05, TID07, UTB^{+}11, WHHB12, YO11, ZRL^{+}11, ZNG^{+}13]. Feature-Based [HR99, LFL08]. Feature-domain [NFSD13]. feature-oriented [FYH11]. Feature-preserving [GYH13]. featured [HH82]. Features [AK77, AM00, AOR94, BMZB02, COW98, CS98, Ehr78, Ekl79, HdVL99, Jon97, Kol83, KK94, NR88a, PA00, RL93a, RY98, SA95, Tsa96, Van77, BCM13, BDL92b, BEGB13, BDL^{+}06, CH09, DIOV06, DSNN08, EK12, FG89, FMGA^{+}12, FAB12, GS95, GBL08, KK11, KG82, LYSS12, MU11, MB95, Mor90, MD82, NHK08, SCE04, SKVS13, SM13b, TCC90, WJ07, ZYS09, dCCP12, AW09, BETV08, LL08]. Feedback [MBKB02, Mai81, KDV12, MW13, Pen03, RGA10, dSdSF^{+}12]. feedback-based [dSdSF^{+}12]. FEM [KMBG09]. femoral [KNO^{+}09]. Few [SGHM00, FFFP07]. Fewer [DBB83, OK07]. FFT [LF82]. fidelity [HAGR91, MWTN04]. fiducials [Ami90]. Field [DC98, Kas80, KCA81, KBZ96, MCPB00, Pag97, Pag99, PD83, PS97, SM94, TTA94, ZSN96, BD94b, CAF09, CYW04a, CYW04b, CMD06, GY05, Gou91, HE82, HC13b, HW06, HNC05, JC06, KS03, LB04, LL12, MHMO09, SM90, WB11, PV13, WKP13]. Field-of-View [TTA94]. Fields [BB88, BA96, CS82, DB88, DB79, DC86, Dod98, Enk88, Fog91, Mas02, MRF96, NC93, SP97a, TA88, WW97, WZWT99, WD92, WSSD96, Zuc85, BP05, BS04a, BK03, HS80, LPR^{+}03, QHXC12, RS91b, TCMS04].

Figural [MPPG98, PEFM98]. Figure [AL99, AD84, HER81]. figures [AK91]. File [KH83b, MHN84, Tam83]. filled [KS91c]. filleting [Elb05]. Filling [AGW86, Pav79, TL88, DV82, GA91, HKA13]. Film [ACW96, GMG92, Lee76, TDK10]. Filter [Bur81b, CGL98, CHB86, GMG92, Lee83b, LK91, SK83b, BP95, DD11a, HBB^{+}12, HSJS10, Imm91, KDV12, yKL11, LcTT91, LC88b, MHSP10, Ney93, TKL^{+}09, YNCO11, ZC89, RRR11]. filter-based [DD11a]. Filtering [Gul79, Jon99, Lee81a, Lyn81, McD81a, MF77, SO01, SMB95, YH81, Ang07, Ano06m, BL09, BKMV07, CNDS13, GKK05, KORC10, MWF07, MDR91, SA81]. Filters [CN87b, FS84, FS85, FSSL86, Hey82, KON87, Mas85, PLS97, Spe97, YBDC93, AS08a, AC09a, BW11, Dou92a, Dou92b, Fon90, GCB90, HDF12, Jea11, KHS94, LRW08, LST13, LY06, LD95, LSPV04, SBB10, SAC09, SD90, Rab92]. Find [Hob00]. Finder [PKP97]. Finding [AS88, CDH99, Fre76, Fri86, GS06, HE81, LH90, LF96, LP79, ND97, NA79, Sam82a, SB93, Sch92, SK83a, SBZ97, WWW95, ZF94, ZC93, ZT80, Ami90, CLHW94, Pon90, Sam89, VCBC88]. Fine [De 93, OD02, TB99, ML13, SY10]. Finger [WF05, ABEN09]. fingerprint [UBEP09]. Fingerprints [Pie88]. Finitary [Her98]. Finite [Kov89, Lee91, SY98, TGSH98, VS82, WBOL07, ITF06]. fir [SA81]. First [CV92, DPB00, Han93, Lan84, RM02, Tan79, VF96, DD11a, RCJ^{+}13, Sub90]. First-Order [CV92, Sub90]. first-person [RCJ^{+}13]. Fish [TML00]. Fit [BCA98, Ros96b, Ros99b, MB05]. Fits [KS95a, KSŽ96]. Fitted [Lil97]. Fitting [BA06, Boo79b, CYH94, HS91, Jac01, Jos94, Kan98, KB00, Man86, Pha89, Ros98b, Ros99b, Sam82b, SZ96b, Vee97, Yam78, YKC^{+}86, CL91, CC96, GO94, HSSH89, KPKPW90, LDGS^{+}13, LY90, MK02, WY11, Ano95g]. Fiume [Ano95h]. Five [Ros99b].

Five-Point [Ros99b]. Fixation [Dan97]. Fixed [GLR^{+}99, ROJX09]. fixing [IHTA90]. Flat [GLD93]. Flexible [BHSD^{+}13, BS99, KMBG09, NMP97, DLP13, LHJ^{+}09]. Flight [LSKK10, BHMB10, LBK10]. FLIR [LCZ^{+}01]. Floor [MCPB00, ES06]. Flow [BK83, BA96, CV92, CM94a, DC98, Enk88, FSA01, FST94, FS95, HZ86, JW87, iK86, iK87b, Kim97, LHH^{+}98, MNCG01, Mit88, NDBT95, PA83, Pra81, Sch86, SP97d, SA92, Spe97, SJB02, SW86, SS84a, TA88, VCVQ^{+}98, WALL00, XS98, ZHAH88, ZI87, BC91, BL09, CHZ^{+}13, CSS13b, DRAB08, FSV07, GYTL09, GPY^{+}07, Gon09, HMF10, JM09a, KN03, KN11, LS08, LB10, MN06, Mar07, MZC^{+}05, MEYD11, MKS^{+}08, MCF10, RDM^{+}11, RPG12, SM06, SG82, Sub90, TLCH05, TCMS04, WBOL07, WWJ13a, ZSCP08, ZLS^{+}13]. flow-based [BL09, CHZ^{+}13]. Flow-Motion [ZHAH88]. Flows [MS96b, Pra83a, SC97a, WD96, ACG^{+}09, HC13c, WW90]. Fluid [FS95, WALL00, LWGP08, TKPR09]. fluids [CD95]. fluoroscopic [KNO^{+}09]. fMRI [KGC05]. Focal [Che08, SCCP05]. Focus [SKOS95, TTA94, CXFS06, FM91, IKST05, NH89, DR04]. Focused [RSB93]. Focusing [BM99, May99]. FOE [Neg96]. Foldover [FM98]. Foldover-Free [FM98]. Folds [ATN83]. Following [LJ87, Sob78, SA85, ZBV93, CR89]. Font [KH96]. foot [TDT12]. footage [MR05]. Force [HNC05, IW97]. Force-Driven [IW97]. Forces [DF01, TG13]. foreground [AHDM10, CVP10, DD11b, YO11]. forest [CFYU12, dSdSF^{+}12]. Foresting [MSF^{+}12]. Fork [SHS79]. Form [BSF02, BB83, CF01, CS98, FAB97, HS06, Kol83, MST85, MKY01, Oka88, SW86, AGCA06, BP95, BvdHL^{+}13, HXS09, KP12, KMP05, Liu10, MFB11, TL05]. Form-Invariant [MST85]. formal [SB89]. Format [SHG^{+}88]. Formation [ES81a, MS97b, Nag78, Udu82]. Forms [Ül01, FR80]. Formulas [GS12]. Formulation [ACB98, SK83b, BLH91, Kaw78]. Fortran [Ree79]. Fortran-Based [Ree79]. forward [AT13]. Foundation [Kan94c]. Four [CJ82, Mid79, HN82, HF11, HQW^{+}12]. four-connected [HQW^{+}12]. Four-Point [Mid79, HN82]. Four-Source [CJ82]. Fourier [WEY06, ANM98, Bat84, DUC97, DG01, Ekl79, EU85, KG82, LEA^{+}10, MMN83, RRS83, TS00a, WW80, ZS11]. Fourier-interpolated [WEY06]. Fourier-Mellin [DG01]. Fourth [Ano96g]. Foveated [YYL96]. FPGA [MZB^{+}10, MAY^{+}10]. FPGAs [MZC^{+}05]. FRA [DK13]. Fractal [AK96, SLN95, YYF89, JTEA91, KCC89, LPZ08]. Fractal-based [YYF89, LPZ08]. Fractional [KJRA96]. Fragment [ASZ99a]. Fragmented [YLL12]. Fragments [EDB12, DT09]. Frame [ADDK99, DN82, HG11, LD90, PR03, SB89]. frame-based [PR03]. Frames [BH86, Far02]. Framework [ADDK99, CW94, Car96, GGR01, JM92, Kd88, LH95, PEF92, SA92, Td93, VM01, ASFP03, BGA05, BK03, CGR13, CMH13, CL08, CU11, DWB11, FFM05, FKV^{+}11, JLD13, KBN12, KM89a, KSR^{+}12, LC11, LV11, Li92, LLC13, LHJ^{+}09, LH03, PJW11, PL10, PMW05, RLS06, RS03, ŠRDC09, TÉSK11, TMB12, TH12, VCT09, WXRA07, YGC13, ZDF10, tHV09]. frameworks [CU11]. FReBIR [PFGG09]. Free [BvdHL^{+}13, BSF02, CF01, CS98, FAB97, FM98, KP12, Liu10, MKY01, TML00, WRB06, AGCA06, KMP05, RC03, TL05]. Free-Form [BSF02, CF01, CS98, FAB97, MKY01, BvdHL^{+}13, KP12, AGCA06, KMP05, TL05]. Free-Swimming [TML00]. freedom [LWLS12, Sha11]. Freeform [EK97]. Freeman [GV78, Kak97, KD81]. Freeman-Code [GV78]. French [KABP98]. frequencies [Fra89]. Frequency [Ano01s, BSI87, JW87, Luc01, SGS^{+}10]. friendly [CPP^{+}11]. Front [SK02, TMT10]. Front- [SK02]. front-based [TMT10]. FS [Neg12]. Full [BR95, UPBS08, LPR^{+}03, SB89]. Full-body [UPBS08]. full-frame [SB89]. Fully [ACB98, GH92, MS96a]. Function [ACW96, BRW85, GK98, GESB95, HD97, KSG84, KH96, O’R82, SB94, SBxx, BSM10, OC90, PFV^{+}11, PSR08, RSS07, RKH05, TTF04]. Function-Based [SB94, SBxx, PFV^{+}11, PSR08]. Functional [Hod95, RDR95, Elb05, KZD^{+}11, ZQ11]. Functionalities [RR95]. Functionality [BB95, Sta95]. Functionals [BN84, Bie87, EW87]. Functions [ADRY94, BGSdVL98, Bid92, CL00b, Dub77, FSSL86, Gro82, HFC96, HPB94, Hum79, KF86, MR96, Ree92, Ros96b, YKC^{+}86, ZK01, Big90, CGU11, CFG06, CU10a, CU10b, FSF07, HS91, OBS06b, PRR03, SDPO81, WR08]. Fundamental [BGK98, CZZF97, TZM98, ZL01, ASCF13]. Fundamentals [COK95, Pos77]. fundus [QKH^{+}12]. Further [Ros99b]. fuse [ZRL^{+}11]. Fusing [BC10, PS12, WW94, BKK11]. Fusion [ACW96, GY88, HSIW98, HSJS10, LMM95, LL08, RFL02, SM93, AM06, ABEN09, BF10, CA10, DS07, ES04, GLOC10, HD09, HGR^{+}13, JBC08, LB08, LFL08, LDC^{+}13, LBCA10, Mig12, TMB12, YW07, YR06, ZZZL13, ZZZP09]. fusion-based [HD09]. Futurist [CH06b]. Fuzzy [GSS00, KW00, KGU10, LSB^{+}00, MWF07, MCPB00, NC93, NFU02, Pha01, RMFB02, SUO00, SU01a, SU01b, SWG02, SP97a, SB13, TB99, US96, WDB12, ALK^{+}09, CUSZ07, CU10a, CU10b, CU11, DK13, ITNP12, LMDB11, PFGG09, WSSS13, ZUS06]. Fuzzy-connected [NFU02, ZUS06]. Fuzzy-rough [SB13]. fuzzy-rule-based [DK13].

G [Ano92a, Ano95h]. Gabor [CR97, Far11]. gain [YCH07]. gait [CNC03]. gaits [Boy04, FRDC06]. Galerkin [KMBG09]. Galleries [EOW84]. Game [YB95, PKK^{+}09]. Games [KBD^{+}12]. Gaps [Boo79a]. Gathering [ANM98, HAGR91]. Gauss [CRC97, JWG04]. Gaussian [AQ09, FL09, Har85, HKZ87, Jur99, KHS94, Kui08, KMN11, LB97, LBCA10, MSR07, MRW^{+}97, Ney93, OD99, OP96, RRR11, SB89, Ste13, TL05, UK12a, ZXY^{+}12]. Gaussian-Weighted [Har85]. Gaussians [GO94]. gaze [MM05, NKB11, NLM05, WSV05, YC05]. GC [CUAT13]. GC-ASM [CUAT13]. Gels [Sko86]. General [AKL93, BPHB91, Bro94, FSS94, Gra78, Kal82, KK88b, MWL99, MWLA99, Sob78, Wal87, ŽA98, CL08, DMW10, DSY10, KPMR91, RR06, RLC^{+}11, Sal90]. General-Purpose [Wal87]. Generalised [KCM85]. Generalization [HMESI13, Pla92, YAT97, SS90b]. Generalizations [KSŽ96]. Generalized [ALSR11, Bid91, CLCO13, Coh85, Elb05, GPY^{+}07, GL86, Hsu79, HC94, HWJ96, JP78, Lam84, LC79, LK97, LH93, MUS06, MP09b, SB84, SW83a, Sor81, Sug93, iTTF82, WLH85, AO03, FL09, Lea92, ZS11]. Generalizing [WO10]. Generate [BG80, BG79, CKLP09]. Generated [MWL99, MWLA99, MFV80, Pra81, TC87, JWG04, JTEA91, PHY^{+}11]. Generating [Cha81, IO09, LMDB11, NGC92, Pot87, Yau84, YB01, ZT98, Shi81]. Generation [Ali77, AM78a, BC85, Ced79b, Cho79, Cou81a, Dor79, Dor84, EK98, GLM78b, HMD93, Kod76, Las92, Liu93, LK00, MB79, MS85, Mey88, MR96, Mun95, NR88b, Nis99, NH92, OYTY98, SSF94, SI96, SA79, SPS81, WR87, YTTT83, ZG91, Cha74, CP09, Chu77, DM12, GLM78a, IK89, IH91, Moo77, SA81, SP06, SA90, ZZLZ13, ZZZY13]. Generative [Ley85, MCB13, PL07, FFM05, FFFP07, Pec07, XHW09, AW09]. generator [Dan78a, Kul79a]. Generators [Hor76, Bad77, GDIIHK11]. Generic [BKMSR98, GESB95, LD98, RJ94, RSL10, SB94, SBxx, CC03, DMW10, FKV^{+}11, OCVV04, RLS06, VK91]. Genetic [DUC97, QY02, SCS99, SC98, GRGB^{+}13, HDS08, SW05]. Genetically [HBL^{+}11]. Genus [LPR93, PSF07]. geo [WCF10]. geo-location [WCF10]. Geodesic [BDL92a, Kim97, PD05, RC13, MJ11, QHXC12]. geodesics [WPS03]. Geographic [MHN84]. Geographical [GS92]. Geometric [BBK78, BKLO87, BR95, COW98, CLR80, DUC97, EK97, GK98, GW93a, GHPW12, GBB98, GL95, HSIW98, II86, Kan91b, Kan94b, KS96, MAN84b, McL96, Mea82, MNSK98, MG95b, NA85, Pet00, RH95, RL93b, SSP01a, Sel81, Shn81a, SKS97, SLL01, Tam83, Tsa96, Vee97, Whi93, WBR86, WBR88, YAT97, Bar06, BPB13, Bre03, CHSV08, CK09, CPS05, Dor89, DQ05, FG89, FF09, FR80, Fra89, GW93b, GC80, GSV05, JBWK11, MS09, MMS^{+}07, PBN^{+}09, SRHC13, WB12, YS08]. Geometrical [BM80, Mil80a, van86, ABD11, Fit88, Nis96a, kWwZ13]. geometrical/statistical [Nis96a]. Geometrically [DLP13]. Geometries [LV96]. Geometry [Åst97, Ano95h, BVL02, BCZ93, BM98, Bog88, CFA98, Col97, DRDKE13, EOS84, FL96, GHMQ97, GSK02, Kan91a, LV03, LL13, MT84, PRW97a, Sch06, SA02, TZ00, Ver97, Whi93, WW97, Bar05, KCC89, KH90, LT90a, NNT11, PS05b, Pos77, SSM06, Tri90, VSP06, WPS03]. Geometry-Based [FL96]. Geometry-Limited [Whi93]. Geons [NL96]. Gestalt [AT89]. gesture [AAASC11, AWC06, HMF10, JM09b, TD04, TDT12, YS09]. Gestures [ZXK02, ZB05]. Getting [CDLD77, Dan81a]. Giant [MAY^{+}10]. Gibbs [Acu92, DC86, Gou91, LCH95, NC93, ZZAA92]. gist [HL13]. given [KS03]. Gleason [SM13a]. Glenn [Ano93e]. GLHS [LH93]. gliomas [RAC^{+}13]. Global [Ano01s, BVL02, GS92, HPB94, KA08, KB95a, KMGC84, Luc01, Ree80, SKB96, YZT^{+}13, YSL11, ZM96, DC04, GFW13, HHWP03, LCP90, PB11, VR95, WSZL13, RK11]. gloss [LMC09]. glossy [PK05]. GMOD [Ros10b]. GMSOLID [Sar83]. Goal [Mee94, Uhr86, DLS^{+}09, PSYZ13]. Goal-Directed [Uhr86]. good [BO05, ZS09]. Gorsky [Ano94g]. Goshtasby [Rab92]. GPA [CLCO13]. GPS [JF10]. GPU [CPP^{+}11, HKM12, Kim13]. GPU-friendly [CPP^{+}11]. grade [RAC^{+}13]. Gradient [BL76, Di 86, Ebe76, Lam84, Mis84, PD83, PC89, PA10b, Rob77, SKK83, WVL81, WSSD96, DS90, HC13b, KS03, LMDB11, SSL^{+}12, WXRA07, ZLH13, ZLS^{+}13, PE09]. Gradients [Ish84, BL04, BS04a]. grading [PKD07, SM13a]. Graeco [HK93]. Graeco-Latin [HK93]. Grammar [Ali77, DP88, LF79, Fla89, Moo77]. Grammars [Fu80, JRV82, JRV83, JR86, KD85, Sub79, YF80]. Grammatical [JvdBS99]. Grand [BGPD09]. Granular [VCVQ^{+}98]. granulometric [ZMCA05]. Granulometries [BJ96]. Graph [BPB11, DFP89, GPDR13, HTEB11, HP78, JO11, JRV82, JRV83, JBWK11, KCD00, LEB07, NRJ11, NBPF11, OTO06, RSB93, SLY89, SB90, YYL98, ZRKZ^{+}11, AS09, CHP^{+}11, CPP^{+}11, CK11, CUAT13, Far11, FKV^{+}11, Fla89, GDIIHK11, HA03, JXC^{+}13, Kaw78, KT08, MLF^{+}12, MMK04, PLLL03, RAHT11, SAS12, UK12a, XHW09, XYZH11, XAB07, ZP11]. Graph-Based [HTEB11, BPB11, JBWK11, NBPF11, AS09, CK11]. graph-cut [CUAT13]. graph-partitioning [MMK04]. Graph-Theoretic [HP78, RSB93]. Graphic [Bas81, BM79, KOY86, BS04b, JFS11]. Graphical [Ano95h, CYW04a, HQ12b, KB01, LM89, MS78, NN13, Ros10b, XG08b, KS00]. Graphics [ABMT87, Ano94f, Ano04q, BSW01, CLR80, EOS84, HB05, Hob00, HCS03, Hua80, KK88a, KZ93, KS00, Oka88, Pav79, PSM80, RGC87, SI96, Sch80b, Sch81, Sch82, SRK02a, Taj83, TDMT85, Wil79, DV82, Gon09, Kaw78, KLBP11, Mar93, RC06, SA81, Wal89]. Graphlike [Ram76]. Graphs [AHL96, BAM87, Ble84, Bre01, JC81, NWP97, NS96, CNDS13, MDFS11a, MDFS11b, PSF07, SRS11, ZS09, ZNG^{+}13, dMFU10]. Gray [DG01, Ham77, Iiz87, KSW85, KB91b, KD86, KU95, Nis98, PA00, PA97, PS95, Sha05, SW83b, iTTF82, THCG84, WP93a, WB97, WP88, Abu89, Dem05, Dou92b, EL91, ELA91, HSSH89, KL07, Pec91, YK95]. Gray-Level [DG01, KSW85, KU95, PA00, Abu89, Dem05, HSSH89, Pec91]. Gray-Levels [PA97]. Gray-Scale [Nis98, PS95, Dou92b, YK95]. Grayscale [Gou84, KZ93, Ste86, TSP97, WCZ02, BDL92b, YCL07]. Green [TL88]. Gregory [CR88, FH12]. Grey [GPK99, SS90b, iTF78, SRTBS91]. Grey-Level [SRTBS91]. Greyscale [Gad91]. Grid [Dor84, Fog84, HY94, Hob97, Kle85, NH92, San78, For89, IM06, MS10, Shi81, SLKL11]. Gridless [PB96]. Grids [RW76, RW79, WC79, WS91, SB05]. Grooves [LKK00]. Gross [Ano92b, Tan79]. Ground [AL99, Fog93, LB98, Cre08, LHM06, SYPK13]. Ground-to-Air [Fog93]. ground-truth [SYPK13]. Group [KC99, SC99b, EW91, MGPF08]. Grouping [ABD11, ASZ99a, BSI87, CH96, CA97, FMRV94, Hen98, HBA93, JDP97, KN99, LM99a, MRF96, PB99, SN99, YJA96, GZP05, LBNS09, YS09]. Grouping/degrouping [ABD11]. Groupings [CN95]. Groups [MFJ95, MJD^{+}00, SM97, KRJ^{+}08]. Grower [PSWH84]. Growing [BB88, RW88, Zuc76a]. Growth [YH78, RAC^{+}13]. Guaranteed [SK98]. Guards [EOW84]. Gudrun [Ano94f]. Guest [Agg83, Ano01k, GSST03, DCCL99, MT97]. guidance [BKP10, HSKH07, RGA10, Ano94g]. guide [TCB^{+}08]. Guided [EM96, JB89, KK92, KK93, AZN11, ASFP03, CMB^{+}12, DDWZ12, RS03, ZDL^{+}11]. guiding [OH05].

H [Ano92a, Ano93d, Ano94h]. Hadamard [Mis84]. Hair [PCP02, YXYW00]. Hairs [LKK00]. Hairstyle [LK01]. hairy [FJP06, XTLP04]. Halftoning [PB96]. Hand [ABEN09, CW00, PC99, Por00, SKOS95, ZXK02, EBN^{+}07, JM09b, OTO06, SGH07, dP10, DBZ07]. Hand-based [ABEN09]. Hand-Drawings [PC99]. hand-pose [dP10]. Hand-Printed [Por00]. handed [HOPA91]. handles [VZP^{+}09]. handling [CH11, LST13]. handoff [CYP^{+}10]. handwashing [HPvB^{+}10]. Handwriting [AHD98, Pel79]. Handwritten [AHD94, DLHT99, HY98]. Hanoch [Ano92b]. Haralick [CGL94]. Hard [FB97, HW94, HO76]. Hardware [BB87, GV84, MZC^{+}05, MNHO00, OR81, Pog85, AK10, AK11, AHDM10, Gon09, MSI10, PCC13]. hardware-based [AK10, AK11]. hardware-oriented [PCC13]. Harmonic [HXS09, LSBG92, HMF10, SGS^{+}10]. Harmonically [GK77]. Harmonics [BH95]. Hash [GK95]. Hashing [RH95, Tsa96, JBWK11]. Hausdorff [CS01, HKM12]. Haze [Max86]. Head [CSS^{+}13a, PE92, SGS01, YWZ11, YC05]. Head-Eye [PE92]. heading [RS03]. heading-guided [RS03]. Heads [FM99]. Healey [Ano93e]. Heart [AHRW87, LSB^{+}00]. Heat [KS96, YLL12, ZGLG12]. heavy [MSSS09].

HECOL [CPC08]. Height [Pag97, Pag99, CH06a, LSC08, Mas09]. hemispherical [GHA10]. Herb [Kak97]. heritages [dOSJVBS12]. hermeneutics [GMW12]. Hermite [KHS94]. Heterogeneous [WW94, GBL08, PZX13]. Heteroscedastic [KB00]. Heuristic [KVdG^{+}97, OBW87, O’R94]. Heuristics [VCBC88]. Hexagonal [SS76, Sub79, WS91]. Hexagonal-Based [WS91]. Hexagonally [Bur80]. hexahedral [ITF06, IA03]. Hidden [Che98, Goo92, GO87, KABP98, Med84, SP97a, BCM06, CLCO13, Fra81, NN13, YK95, ZYXZ13, ZZAA92]. Hidden-Line [Goo92]. Hidden-Surface [Goo92]. hiding [YCL07]. Hierarchic [MS78]. Hierarchical [Ahu83, CWH^{+}13, CN95, DFP89, DF91, FKL^{+}98, GS92, HUF05, HP96, KBZ96, KH98, KD96, Lee86, LLL86, LKE00, ML13, MJ88, NN13, PF87, Pri86b, PCR^{+}04, SP92, Shn81b, SL96, Tan95, TP75, TTIM96, VBH97, WHL84, YZ06, YNCO11, YW99, Cou13, DCL^{+}08, GL82, HBH10, JEF^{+}12, LC88a, PIK90]. hierarchies [JR09]. Hierarchy [Jon97, SN99, NFA04]. High [AHRW87, CJL06, CTH84, CJC01, DT96b, DH92, EA95, Fel85, Gud82, Kri84, Kro86, MCPB99, MN94, PCJC98, RL93a, RY95, SSF94, SLK86, Ver81, BC10, BEGB13, BKMV07, Cha74, CBT^{+}04, DRAB08, Fra89, HBH11, KA08, KZD^{+}11, KSS08, MWTN04, SP06, YAK^{+}08, ZYT10]. High-Accuracy [DH92]. High-Dimensional [MN94, BEGB13, BKMV07]. High-Level [Gud82, SLK86, ZYT10]. High-Order [Kri84, KA08]. high-performance [DRAB08]. high-precision [KSS08]. High-Resolution [CTH84, MCPB99, PCJC98, RY95, Ver81, KZD^{+}11, SP06]. High-Speed [AHRW87, DT96b]. Higher [WW95, YAT97, ZZP12, DS90, PL08]. Higher-order [ZZP12, PL08]. highlight [GHHX04]. Highlights [CTE95, MS00, ABC^{+}03]. Highly [SM10, DLP13]. Hildreth [For88]. hippocampus [XFSC13]. Hips [LCS84]. Histogram [Bid86, Fre77, Gau92, Gla93, HW79, Hum77, KSW85, KNJ84, Ku84, Lor83, MGW10, MAP99, MBDB88, O’G88, PAA^{+}87, RY95, SF96, WCZ02, ZCL99, BK07, KGU10, MHSP10, PMR92, Sez90, Wha91]. Histogram-Based [Gla93, KGU10, MHSP10, Sez90]. histogramming [LP90a]. Histograms [WPR85, Cho88, JWG04, LLR10, LL04, LL12, LDC^{+}13, NHY10, Zit88, PA10b]. histology [SM13a, Tan11]. history [WRB06]. Histosplines [TL79]. HMMs [WZL^{+}03]. hodograph [Far02]. HOG [HC13b]. Holes [AGW85, FHMB84, LPR93, CHSV08, QHXC12]. Holography [JM79]. Homeostatic [FY06]. Homogeneity [Wu93, KLL^{+}11, MVP06]. homogeneous [BFR13]. homographies [CPS05]. Homography [CPC08]. Homotopic [Pud98]. Hopfield [BBB96]. Horizon [MAL10]. Horn [Hor79, Kul79a]. Hough [BTNS90, CGR13, CS04, CL95, Con88, dFCS93, CSR83, DGH98, FMRV94, FS03, GLR^{+}99, GRB13, HBA93, HNR88, HNRR90, IK88, KB91c, KB00, KBD^{+}12, KC87, Lea92, Lea93, LY05, LLL86, MGK00, MNHO00, Ols99, PPK93, PKP97, PIK90, Ris89, ROH88, SYK96, Sha06, SK98, SKBS13, TD83, XO93]. Hough-based [GRB13]. houses [ÜB05]. HRCT [SBK^{+}99]. HtHT [KB00]. Hue [LH93]. Hull [EKH01, BL08, Wha91]. Human [AC99, Bie85, BP94, BL01, CFCP11, CMBP09, DLF06, Gav99, GMW12, GAD01, HH77, HL79, KB01, LC85b, LRD99, LLC13, LSTF12, MYLP98, MG01, PC05, RKK^{+}00, Roh94, Ros87b, SPK^{+}02, YD94, ZXK02, Ano06m, BCM13, CGH08, CCFC13, CYNO11, CNC03, DIMT12, FFY^{+}04, GKK05, HER81, HUF05, HWW06, ITNP12, JS07, KV06, KRK11, KPKH07, Kou03, LL06, LE09, LSCM03, LWH03, LYA13, MFB11, MHK06, NFM08, NLM05, OMBH06, PT08, PDS^{+}07, PQML11, PYS03, Pop07, Rem04, RSPD12, RR06, RS03, SMT04, SKM06, SH08, SRHC13, TCMS04, TR09, UFF06, WS08, YO11, YS08, ZMCA05, ZKC03, ZDF10, Ziv10, ZCCD06, BCDH10, HG11]. human-delineated [Ano06m, GKK05]. Humans [Ano94f]. Hybrid [Bai88, CC96, Mor81, Peu83, WZ97, DWW^{+}12, Fon90, KSR^{+}12, KL11, VMP03]. HybridTree [AGCA06]. Hyperacuity [Kro86]. Hyperbolae [Ros98b]. Hyperbolas [MB79]. Hyperbolic [CS01, XWYY10, ZK05]. Hyperbolization [Fre77]. hypercomplex [AS09]. Hypercube [DRCF95, LHKC97, LP90a]. hypergraphs [BB13]. Hypermaps [BD94a]. Hypermatrix [AS93a, AS93b]. Hyperplanes [AAS97]. hyperquadric [CC96]. Hyperquadrics [Han88]. Hypervolume [PASS01]. Hypotheses [MS97b, YK95]. Hypothesis [HDM86, Kan94c, LVW97, MML87, PPK93]. Hypothesizing [Kan91b, RSB93]. Hysteresis [Ehr78].

IAPR [EHG^{+}10]. Iberian [CCR^{+}05]. ICA [DBBB03, Hu08]. ICA-based [Hu08]. ICDAR [Ano96g]. Iconic [CBD^{+}03, GL86, TCAC90]. ICP [FDMA97, PLH04, YB07]. ICP-based [YB07]. Ideal [Fiu89]. Idecs [HC77]. Idempotent [BDL92a]. identical [HBL^{+}11]. Identification [Bai88, CA86a, CTE95, FT84, GLR^{+}99, KH96, LCD97, TN08, YK97, ABEN09, ABC^{+}03, BRA^{+}10, BCM13, CTM^{+}13, CL08, FM91, ILRB04, KS91b, LY05, LSCM03, LN10, ML13, PGGM04, RCTV12, TDK10, WPK09, XYZH11, BDL92b, HH05]. identifier [WF05]. Identifying [AK78, Maa94, RL93a, Sau93, TN05, GS06]. Identikit [BM96]. Identity [YK86]. If [KL77]. IFS [BBC00]. IFTrace [MSF^{+}12]. Ignorance [JBS^{+}91]. II [BDL92b, CU10b, Dou92b, ELA91, Fiu91b, JRV83, MD95, RH91]. Illuminant [DC98, Dre94]. Illuminants [GM87, APB10]. Illumination [BFF97, BWL04, Chi97, FW97, GG09, KC94, KC95, Lai00, LZ97a, Max91, MCF10, OD99, OD01, SJ89, AC09a, AC09b, ARARCE11, CCYC12, DD11b, DL10, Hu11, Jea11, LCT09, LY06, MTVM04, OK04, YWZ11]. illumination-based [ARARCE11]. illumination-encoded [Jea11]. illumination-invariant [AC09a]. Illumination-robust [MCF10]. Image [Ahu83, AAS85, AR77, AGHN94, AK11, ABW97, APV99, Ano94g, Ano94f, Ano95g, Ano01r, Ano06m, Ant82, AS93a, ADRY94, AEH79, ACW96, BK01, BCZ93, BB87, BS99, Bie85, BP94, BFY00, BRW85, BW93, Bou79, BHF08, BACL97, BS88, BGS83, Bur81b, Bur83, CL83, CGL98, CMW^{+}97, CP80, CYH94, CM97, Cho88, CT93, Coh85, CF92, Col77, CH09, CC00, CL97b, Cre08, CWC94, CW00, D’H86, DB88, DT96a, DF02, Dav93, De 88, DCCL99, DPB00, DH00, DW87, DG01, Di 86, DR93, Dre94, DSH04, EK98, EA95, Enk88, FUS^{+}98, FF79, FRL^{+}98, Fog91, FST94, FSB85, Fre77, Fri80, Fu80, FL96, FM98, FC86, Gau92, GMG92, GFS04, GLD93, GGMV08, GMW12, GR85, Gos89, GHS95, GGR01, HM84, HE82, HC77, HM78, Har80a, HW81, HS85]. Image [HS87, HR99, Haw78, Haw82, HL76, HH96, Hey82, Hob97, Hor74, HLF^{+}97, HNR84, HMA10, HJS89, HAGR91, Hum77, HDM86, IP98, JWG04, JM92, JC93, KB98, iK87a, Kan91b, Kan94c, KSS97, KS95a, KWK84, KBZ96, KNJ84, Kis96a, KIF85, KD96, KVdG^{+}97, Kov89, Kri84, KJRA96, KH83b, KH83a, KH86, Lai00, Lan91, LCS84, LBD92, Lee81a, Lee83b, LB97, LI00, LBS80, Leu92, LL98, LPF78, LN85, LCH95, LL95, LMM95, LN98, LJ91, Liu97, LLE^{+}09, LB87, Lyn81, MBKB02, MS96b, MAP99, Mas85, MKK02, MS97b, McC82, McC80, MBDB88, MK01, MPC94, MS85, MM92, Mey88, MB94, MN95, Mil80a, MBMC11, ML78, MYLP98, Mis84, MY87b, MFV80, Mok92, MPPG98, Muk92, MR92, Nag83, ND92, NDN^{+}97, NVWV97, NLW13, OH81, OS87, OD97, OTL96, OYTY98, OBH04, PZ09]. Image [PS83, PP95, PBQ99, PM97a, PMV00, RW88, RMR85, RBA94, RP88, Ree92, Ree84b, RWWH00, Ric84, RC03, RWD90, RM98, Roh94, Ros88, Ros89, Ros90, Ros91, Ros92, Ros93a, Ros94, Ros95, Ros96a, Ros97, Ros98a, Ros99a, Ros00a, Ros01, Ros10a, STEK96, SUO00, SU01b, SK79, SP97a, SC97a, SB93, ST96, SC99b, SB85, Sch86, SLST99, SWH84, SF95, SHG^{+}88, Sha94, Shi99, Shi83, SP97c, Shu97, SBK^{+}99, SPK^{+}02, SA92, SJ89, SL99, SLCP85, SMB95, Ste01, SK83b, SW86, SS79, Sze93, TVLS08, TST^{+}83, TS00a, Tan79, Tay00, Thi92, TS86, TSK94, TZ00, Tou80, THT^{+}98, UZC97, US96, Ull81, VLR84, VKP98, WN99, WLD99, WWC82, WWW89b, Wal87, WVL83, WD96, WCZ02, WLH85, WF78, Wec78, WC79, WN86, WBB85, WALL00, Wil79, WB82, Yac83]. Image [Yan93b, YL94, YB95, YFZ98, ZW97, ZL01, ZFG08, ZCL99, dM92, vv92a, AM06, AQ09, AM93, Ang07, AC09a, AO04, ASFP03, ATC^{+}13, BBF^{+}11, BCL^{+}90, BTCH05, BK07, BP05, BF07, BCDH10, BT05, BvdHL^{+}13, BB04, BD94b, Big90, BSMK13, BPB13, Bor91, Bur81a, CG09, CFYU12, CH06a, Cel90, CT10, Cha91, CLC91, CYNO11, CUAT13, CLZZ13, CFM^{+}13, CU10a, CU10b, CU11, CG04, CKS^{+}05, DR03, DBF04, Dam08, DR04, Dem05, DSNN08, DAM12, DF91, DCS05, EL91, EW91, ELA91, FPC^{+}08, FY06, Fit88, FAB12, FYH11, GRGB^{+}13, GFL^{+}11, GSS12, GH08, GSST03, GS08, GCPF08, GDR04, HDS08, HMC10, HJ12, HC13a, Hei04, Her90, HH91, HS89, HC13b, HWW06, HGS08, HAKK91, IP91, JMPG11, KS02, KS89, KK13, KM94, KA08, KN03, KHH^{+}12, KB91b]. image [KH90, KMT11, LV03, LT05, LC11, LM91, LH95, LSC08, LEB07, LPZ08, LL12, LC88b, LFL08, LLC11, LS12, LPV07, MWF07, MVP06, MUS06, MSR07, MSG10, MMV06, Mar93, MMK04, Mas09, Mat89, MHW89, MJBR88, Mee89, MGPP11, MB05, MTAA11, MGPJ11, NHK08, NFU02, OC90, OTO06, OK04, PJW11, PSE^{+}11, PMR92, PS03, Pen03, Pen89, PA10b, PFGG09, PG13, PBG04, PMF90, Pun03, QAB^{+}11, RDM^{+}11, Reb89, Rem04, RFS03, Sah05, SCD11, SG11, Sez90, SB13, SKH08, SKU^{+}09, SCvW11, Sub90, TLT91b, TL05, TT91, TFB80, TLEF06, The83, TMB12, VMP03, WS89, WL89, WLZW04, WZ04, WLW06, WO10, WSSS13, WKP13, WB90, WW90, WW91, WWJ13a, XYW^{+}08, YZT^{+}13, YGH11, YB89, YCL07, ZC89, ZZZ06, ZTH^{+}11, ZYXZ13, ZLS^{+}13, ZUS06, ZU09, dMFU10, MSF^{+}12, Ros00b]. Image-Based [FL96, CG04, FPC^{+}08, WLZW04, WLW06]. Image-Flow [SA92]. image-guided [ASFP03]. Image-modeling [LCH95]. Image-Pair [DH00]. image-plane [PMF90]. Image-Space [LI00]. ImageCLEF [THL13]. Imaged [CB98]. Imagery [Agg83, BM99, BBD^{+}94, Chi81, CJC01, CP81, DRDKE13, EM96, FTW81, FY85, GT84, Hsu79, LV03, LH84, May99, MNSK98, MCPB00, NK00, PCJC98, PD83, SHJB^{+}83, SM93, SK85, TM86, BPHB91, DZL07, DS07, HVD^{+}89, HOH^{+}07, HO76, JTEA91, PSR08, SSN03, TSR89, YCH07, ZZZP09]. Images [AG00, Ano95g, AA93, AOR94, Bat84, BGT94, BJ86, Bid86, Big97, Boo97, Bor86, BM80, BM97, Bur80, Cag93, Cai88, CW94, CC97, CA97, CM95, Chi97, CCA92, CJC^{+}98, CH78, CN87b, CPD93, CGP85, DH92, Dav97, DUC97, DC86, Doe98, EC88, FH84a, FKL^{+}98, FS95, FMR01, FSB85, FM99, GPK99, GCB92, GM79, GM94, GSU00, GW01, Gro82, GBB98, GN98, GJP96, GB93, HdVL99, Har83b, HRS02, Hei99, Her80, HHI95, Hor77, HW79, Hor84, HN88, HNRR90, Iiz87, JGR85, JV97, JB99, JEK98, KZ93, KW99, Kas80, KCA81, KSG84, KCD00, KDRC98, KLK88, KS96, Kim97, KSI98, Kle80, KD86, KMA^{+}00, KK83, KU95, KdVL99, Lee81b, LR90, LK91, Lee91, LPR93, LF96, LH92, LSBG92, LAS94, Liu93, MW00, Maa94, MA84, MA85, Mai76, MS97a, MGMS01, Mar82]. Images [MY95, Mas02, MMP85, MCPB99, MWL99, MWLA99, MB85, MF77, ME98a, MAM97, MR96, Mor76, Muk97, NA85, NMP97, NC93, NL96, Nis98, OD99, OD02, OP96, Ols93, OY92, Pan78b, Pav80, PR92a, PH82, Pot87, Pud98, RJ94, RC97, Ram76, Ram84, RSB93, RBA94, Ree84a, RY98, RFL02, RY95, RMFB02, SA91, SAA93, SA96, Sam82a, Sel86, Sob78, SF97, Spi98, SLN95, SH84, SA85, SB02, SM99, TSP97, TK97, TD83, TL79, Udu82, WWHL88, Wam85, WB97, WH01, Wat87, Wei92, WN87, Whi93, Wil84, WD92, Yan93a, YTTT83, YK97, ZK81, ZSN96, ZT98, ZC93, dCCP12, vYB89, vv92a, ÁB13, BI10, BP95, BSMG05, Beu91, BDL92b, BDHM09, BGR89, BB11, CCTCR09, CCR^{+}05, CFG06, CTM^{+}13, CSS^{+}13a, CL90, CCH91, CLHW94, CP91, CGG91, Cum91, DCFM07]. images [FM91, FMGA^{+}12, FL09, GK03, GL82, GE08, GCEC07, GU89, Gou91, HVD^{+}89, HQN05, HSJS10, HH82, JEF^{+}12, JRH03, KL07, KM89a, KN04, KS91c, KS12, Kou03, KNO^{+}09, KSG^{+}13, LD90, LP91, LJHH07, LPS^{+}11, LB05, Li92, LT90a, LDD09, LcTT91, LTT91, LP90b, LS09, LMDB11, LBCA10, LP10, MN06, MR89, MSS90, Mar90, MJ11, McD81b, MAL10, Mig12, MDR91, MB95, MGPF08, Mos91, NKP11, NKPT13, O’G94, OJRT08, PE09, PBM^{+}11, PL10, Pey09, PS12, PCR^{+}04, QKH^{+}12, RSS07, RKH05, Sam89, Sau91, Sch06, SCS91, SW94, SS11b, SdB03, TAK09, TLT91a, TA13, TCH07, TS11, TP05, Tri90, UA90, ÜB05, WS03, WPK09, WLI08, WB11, YHR^{+}05, YWMS08, YZ06, YT13, ZMCA05, ZSCP08, ZRL^{+}11, ZZAA92, ZG91, BS89, SRTBS91, Vos88]. Imaging [AHRW87, Bra85, HH98, SGK00, Ver81, GHA10, GHMT09, GPC^{+}10, HGSM11, KZD^{+}11, KLL^{+}11, KLBP11, SGA12]. immersive [TCMS04]. impact [TM04]. Imperfect [DY98, RN93]. Imperfecta [TZ82]. Implementation [Bre03, BS87, BKA84, CWSI87, FSS94, GV84, GLR^{+}99, Gri83a, LHHC98, MNHO00, SRL82, TST^{+}83, YPVv81, AFH81, BTNS90, CRT90, LM91, MSI10, MFB11, MZC^{+}05, MAY^{+}10, NN04, SBB10, SS90b, SM10, WL89, dLAH07]. implementing [KL10]. Implicit [CG94, GA00, HSIW98, LDPD97, LSB^{+}00, RAH97, SAG84, SK01, Ül01, VTG95, ZOMK00, AGCA06, BGA05, BMM^{+}07, HUF05, MCQ05, TRS06, VCT09, WSKH13, ZQ11]. Implicitization [SAG85, SS95b, SZKD99]. Imposing [FB97]. Improve [ACB98, ZW97, FBF08, KB91b, dSdSF^{+}12]. Improved [Ano92b, CM12, Cap84, GPC^{+}10, JJM95, JB89, Mil99, MB05, OEK08, Sch76, UG92, HH07, LM12, Pec91, SW94, SZ07, SYPK13]. Improvement [Dan81b]. improves [BHMB10, FL92]. Improving [GBF12, IP91, RPG12, XJK12, YAK^{+}08, ZK81]. Improvisation [Hod95]. impulse [SB91]. impulsive [MGPF08]. inaccurate [GA91]. incenter [HMESI13]. including [WR08]. Incompatibility [Åst97, Col97, PRW97a, CR88]. Incomplete [Nai87, RN93, KBN12, MYC09]. incompressible [ACG^{+}09]. inconsistent [LPC08]. Incorporating [GW07, LHH97, dSdSF^{+}12, CSY08]. increment [NFM08]. Incremental [DHP08, GA00, GB08, iK84, Pen94, WR87, XG08a, Buz03, Dam08, FFFP07, KS02]. Independent [BKMSR98, DT96a, FD99, NFM08, Sin87, SJ89, WWW89b, AFSW03, CD11, EKY08, LT05, WWW89a]. independently [OCVV04]. Index [Ano93a, Ano93b, Ano93c, Ano94c, Ano94d, Ano94e, Ano95e, Ano95f, Ano95d, Ano96e, Ano96f, Ano96d, Ano97c, Ano97d, Ano97e, Ano97f, Ano97b, Ano98b, Ano98c, Ano98a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00a, Ano01d, Ano01e, Ano01f, Ano01g, Ano01c, Ano02c, Ano02d, Ano02e, Ano02f, Ano02k, Ano03r, Ano03t, Ano03u, Ano03v, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano06p, Ano06q, Ano06r, Ano06s, Ano07l, Ano09a, GM90, WCZ02, Ano03s, LZWP03, PBG04]. Index-based [GM90]. Indexed [Sch80b, Sch81, Sch82]. Indexing [BGSdVL98, CS98, CS00, DvLV08, Doe98, GFS04, GL86, MAP99, MLP97, Nis99, SO01, SBxx, YC98, BL04, Fly92, JN09, QT10, TCAC90]. indicators [CH06a]. individual [XFSC13]. Indoor [GW01, CGU11, DWB11, KPPK09]. indoor-sports [KPPK09]. Induction [PC99, VBS^{+}04]. Industrial [SOJ^{+}95, Wec81, ZZZ06]. Industry [Gar76]. Inertia [GPP88]. Infer [CCA92]. Inference [JvdBS99, LF79, SB95, Ham05, WKP13]. Inferring [KMB97, OGH04, KRK11]. infinite [Hub12, SB91]. Inflating [CM95]. Influence [HFKN97, BGPD09, GZP05]. Information [AHD94, AEM98, AD84, Art79, BEGB13, Boo97, CY83a, Chi81, CM97, Fis94, GS92, HB98a, Hob00, JM79, JB91, Kro86, LL98, MHN84, Mil80b, Nai87, iOKS80, PMV00, Pot77, Pra83a, RW88, Ree80, Sha94, SLCP85, SB02, Tan79, iTF78, Udu81, CSY08, DUSL94, GH08, HSSH89, Hei04, KK07, KT07, LL12, PW91, SKU^{+}09, WSSS13, ZYT10]. Information-Based [PMV00]. Information-Directed [RW88]. Information-theoretic [BEGB13, WSSS13]. Informational [Bri84]. informations [PS03]. informative [DL10]. Infrared [UA77, BBC^{+}07, DZL07, HASS10, KHA^{+}05, SSN03]. inhomogeneity [MUS06]. Inhomogeneous [GSP02, YHN11]. Initial [HSSB98, JB89]. Initialization [CYES00, NFSK97, SKSR08]. Injectivity [CL00b]. ink [HO76]. inpainting [BR12, CHSV08]. Input [PSM80]. Inscribed [BM98]. insect [GODC07]. insensitive [BWL04, GJ10, NB10, PV06]. insertion [YJC^{+}09]. Inspection [COW98, Chi88, EU85, HF80, MG95a, MEDT96, ME98b, NJ95, Oka84, Per81, SOJ^{+}95, TG95a, TG95b, LA11, SDR91]. inspired [BC10, BCDH10, EK12, HL13, MFG10]. Instabilities [ASZ99b]. instance [FBF08, YGC13]. Instantaneous [Pra81, PV06]. Instantiating [WRH97]. Integer [KD81]. Integrability [FW97, KS03]. Integrals [FS85, iIK85]. Integrated [BL09, Gro84, HF80, LD98, SA95, Tan81a, VZP^{+}09, ASFP03, HAKK91, KSS08, PBG04, TMB12, TG95a]. Integrating [AT89, BZ99, DCTO97, MNE00, SSdVL06, TCZ^{+}12, NT10, Nis96a, eGZW07]. Integration [BGR89, CT93, DL97, FTW81, HDM86, KMN11, MFJ95, Mas02, WL88, CUAT13, CJL06, DGG08, EDB12, dOSJVBS12, RFS03, SSL^{+}12, VSP06]. Intelligent [MB98, SO07, MFG10, RGA10, Tho10, VD10, Jon08]. Intelligible [SLCP85]. Intensified [Wam85]. Intensities [Leu92]. Intensity [CH85, CM94a, Chi97, CW00, De 93, FDMA97, GMA83, GJP96, JB89, KC94, LN98, LP77, Nag83, O’G88, Per81, Piz81, Ree92, TTIM96, WLH85, ZW93, ZU09, ZBV93, AS08b, CD13, Hil83, JC06, SKU^{+}09, SKSR08, WW91]. Intensity-Based [De 93, FDMA97]. Intensity-Dependent [Ree92]. Intensity-Guided [JB89]. intent [PSYZ13]. inter [GB08, JSRS08]. inter-camera [JSRS08]. Interacting [BGT94, PDS^{+}07, JBC08, KPPK09, PA06]. Interaction [MKS^{+}08, ZXK02, EK12, FR11, HSH07, JS07, JFS11, KPKH07, PYS03, SA04]. Interactions [Cav87, Ken86, PT08, ZNG^{+}13]. Interactive [AAV96, BB95, Bou79, BM96, DC04, GK95, Gud82, HM78, KSS08, LB06, MBKB02, MB98, Peu79, PSM80, PZV13, Udu82, UA77, CD95, CG04, DWB11, GVK06, KS02, LF04, Mar93, MCQ05, MO11, MM05, NFU02, SBS04, THL03, WWH07, WWLV11, dMFU10]. interactive-rate [CD95]. Interactively [PC99]. Interdisciplinary [MST00]. interest [CHMG12, GG09, ILRB04, KL10]. interest-based [ILRB04]. Interface [KH83a, LF04, NLM05]. interfaces [MĆK09]. Interfacing [Ull81]. interferometric [WB11]. Interframe [AM01, SJ93a]. Interior [AB88, ZH79, VSR12]. Interline [Yan93a]. Interlocking [Cho79]. Intermediate [IKS86, JC81, LT81, VCBC88]. Intermediate-Level [IKS86, JC81]. International [Ano94j, Ano96g, GP06, LV03, KB01, Wyv03]. interpolant [CR88, Klu78]. interpolated [WEY06, ZS11]. Interpolating [BG80, Bar84, KS04a, BG79]. Interpolation [AM01, ABMT87, BS96, BG91, BH83a, GY99, GKR02, Gri83a, GL98, HC96, KS95a, LCC89, Mid79, OY92, PMV00, SKS97, TL79, BGLSS04, CCH91, GC80, HN82, Kim04, LLXW13, OBS05, VSR12, XZWB06, YYF89]. Interpolation-Based [OY92]. Interpolators [Sch93]. interpolatory [BFRA12, HB91]. Interpretability [Hor77]. Interpretation [DUC97, DTG96, HB98a, KL93, LT81, MS00, Mun95, OMLL98, PC89, Pel79, SB00, Ste01, TN07, ARARCE11, BC10, Fly92, Har83a, KK07, LWH03, MHW89, PMF90, SM06, UN91, VZP^{+}09, XP11]. interpretations [OTO06]. Interpreting [Dou81, San77]. Interquartile [SWH84]. Interrogation [Elb01, HS05]. Intersect [San78]. Intersecting [OCON82, MD95]. Intersection [HHS^{+}01, KKO98, LP79, SAG85, Tan89, YY84a, LPR89, MG95b, YY84b]. Intersections [Lev79, PS94, Sar83, SDC04, JBK04, MD95, WJG02]. Interval [SSP01a]. intra [ASFP03]. intra-surgical [ASFP03]. intraoperative [LPR^{+}03]. Intrinsic [DAM12, Kim97, CYW04a, CYW04b, JXCZ13, LC11, QHXC12]. Introduction [Ano95h, BCM97, CFS98, KR89, LLE^{+}09, Ser86, GSST03, DCCL99, MT97]. intrusive [YC05]. Intuitive [Ley87a]. Invariance [Chu02, iK87a, SC00b, DS90]. Invariant [BJ86, BHNR93, DG01, EK88, GDIIHK11, KR98, KORC10, LS01, MST85, Mer81, MS85, MPPG98, PEFM98, SSS13, VKP98, WH78, ZC89, AC09a, AKC11, ASCF13, BT05, FB12, Fly92, HMF10, LSCM03, MPVF11, OMBH06, OBH04, OH04, Pun03, SCE04, SHC^{+}12, TVC09, WCZ^{+}07, XYZH11, ZZL13]. Invariant-Based [KR98, VKP98]. Invariants [Che96, KPH02, NG98b, NS91, PR92a, QV98, RW97, Sel81, SLL01, TC86, BPHB91, BG09, GBB98, HN95, MTVM04, PC05, ZCF13]. Inverse [AL11, Hsu79, TGB00, UPBS08, WVL81, LcTT91, OH06]. Inversion [SAG85]. invertible [ITF06]. Inverting [KC94]. Investigation [RWV95, Ull83, LL12]. Investigations [Enk88, Gau92]. Involving [KW00]. Io [McD81b]. IR [CFB05, MNSK98]. Iris [BKK11, Far11, GRGB^{+}13, BHBF10, BHF08, HBF09, HBL^{+}11, LDGS^{+}13, NFSD13, PS12, CJL06]. irises [HBL^{+}11]. irregular [GDIIHK11, KA12, VCT09, VRKL13].

Irregularly [GSP01, PPT06, TN05]. Iso [LM00, Las92]. Iso-surface [Las92]. Iso-Surfaces [LM00]. isogeometric [TJ12]. isointensity [TG95c]. Isolated [BBC00, NS98, Sup02]. Isolated-Object [BBC00]. Isolating [MGPF08]. Isoluminance [Sav87]. Isometric [Pen92]. isosurface [Kim13]. isosurfaces [GK04, VS08]. Isothetic [YLWY92, BR90]. Isotopic [FK99, CCS05]. Isotropic [Mid79, AdVDI05, BS92, HN82]. Isovist [DB79]. Isovists [DB79]. Issue [Agg83, Ano01h, Ano01q, Ano01r, BSW01, CFS98, CNK01, DRDKE13, FHP01, GHPW12, HB05, KB98, KS00, KCOTW06, RFL02, SCOG09, SRK02a, Tau02b, WP00, Wyv03, Ano05p, Ano06o, BPS10, CA10, CKB10, FPDK12, FYH11, GP06, GHMT09, HMC10, HTEB11, HGSM11, HCS03, HM13, JWDF05, Jon08, KSM^{+}06, KPKH07, KLBP11, LBK10, LLE^{+}09, MPF07, MYK03, NLW13, PS05a, STV09, SST06, SMHH04, THL13, Tho10]. Issues [Ros86b, SK86, Tho86]. iterated [Wil89]. iteration [HKM12, NKP11]. iteration-by-iteration [NKP11]. Iterative [CH99, CUSZ07, Col77, DM82, Ebe76, GSK02, Mai81, ODD96, RMR85, Sam82b, Van77, AL11, HQN05, LBNS09, MKA73, Ram72, TMB12]. IVIS [TG95a].

J [Ano93d, Ano94f, Ano95g, CV13, CM94b]. Jacobians [TG13]. jet [HO76]. Jigsaw [RB82]. Joe [GW93a]. Johansson [SGDP01]. Joining [NHK08]. Joint [KGFP10, MS97a, MMA06, QV98, SM06, ZBLS13, Gon09, HUF05, JLD13, YO11, ZZ07]. Jonathon [Ano94h]. Jordan [AHL96, Her92]. JPEG [Liu97]. JPEG2000 [BRSSAL11, TVLS08]. Jump [HH97]. Jump-Out [HH97]. Junction [ÁB13, LL97b]. Junctions [Dem96, LM99a, BB04]. Justification [KU92].

Kalman [Ano06m, GKK05, LK91, SK83b, YNCO11]. Kalman-particle [YNCO11]. Kandidats [HM78]. Karhunen [CC97, SB85]. Karhunen-Loeve [CC97, SB85]. Keeping [Gui99]. Kernel [PPK93, ZRL^{+}11, BB13, GGMV08, GCPF08, KK95, SB89, WZL^{+}03, WHM^{+}09, ZGLG12, ZCK09, DT10]. kernel-based [GCPF08, WZL^{+}03, ZCK09]. Kernel-edit [DT10]. kernel-predictability [GGMV08]. kernels [Hub12, JBR08, YLL12]. Key [AS83, ADDK99, PR03]. keyframe [TM07a]. Keyword [Sch80b, Sch81, Sch82]. Keyword-Indexed [Sch80b, Sch81, Sch82]. Kidney [SP81]. Kinematic [ZDF10]. kinematical [FLB06]. Kinematics [AL11, TGB00, OH06, UPBS08]. Kirchhoff [RH06]. Klinker [Ano94f]. knee [LPS^{+}11]. Knot [BKW96]. Knots [GW93a, Sup79, GW93b, Reb89]. Know [Hoc87]. Knowing [iK85]. Knowledge [AR77, CT93, CL97b, DTG96, EM96, FTW81, OD99, MR90b, Mat89, MHW89, XP11]. Knowledge-Based [CT93, CL97b, DTG96, Mat89]. Knowledge-Guided [EM96]. Korean [SHKP98]. Kropatsch [Ano92a]. Kruse [Dan81b]. Kulpa [Hor79].

L [Ano95g, CH11]. Labeled [FMRV94, San90, WDN^{+}12]. Labeling [BCZ93, Fau81, NHR81, THN92, YD94, YB95, ZH79, CPC08, CCL04, EyGS11, JLL13, MR89, Nic95, RCVA11, SMD^{+}08, SHS03]. Labelled [MRF96]. Labelling [ERW93, GLR^{+}99, LHB87, AHDM10, HQN05, SRS11, UA90]. labels [SYPK13]. laboratory [TN08]. Lack [Har86]. laden [CGAY13]. LAMP [ZH04]. Landmark [TW98, DDLP10, GSS12, RFS03, TLWT12, WR08]. Landmarks [HRS02, KSS92, HS06, SSM06]. Landsat [CN87b, HW79, Yam80]. Lane [FUS^{+}98, Gui99, Lee02, KS02, LY05]. Lane-Departure [Lee02, LY05]. Language [BKMSR98, Gud82, RW76, RW79, Sug78, WWW89a, OTO06, WCZ^{+}07, VM01]. Languages [Col81]. Laplace [vYB89]. Laplacian [Ber84, DvLV08, Ney93, SB89, ZHM11]. Laplacian-of-a-Gaussian [Ney93]. Laplacian-of-Gaussian [SB89]. Large [Abe84, Ber89, CGR13, Chi97, FPDK12, IZKB12, Mar07, PSK^{+}02, Peu83, Sch78, SA02, Tau02b, CPS10, DSS94, HBH10, KSR^{+}12, MPST08, YWZ11, YC05, ZTH^{+}11]. Large-scale [FPDK12, IZKB12, CPS10, YWZ11, ZTH^{+}11]. Laser [CZZS07, FK09, SRML09, ZG06, FRNS05]. Laser-based [CZZS07, FRNS05]. late [LDC^{+}13]. latent [SHC^{+}12, ZG10]. Lateral [BH83b]. Latin [HK93]. Lattice [Car96, Dav93, LL92, SP97c]. Lattices [BNG02, Ang07]. Laurent [Ano95g]. Lawrence [Ano93e]. Layer [GY01]. Layer-Based [GY01]. Layered [OGH04, GH03, ZH04]. layering [CLZZ13]. Layers [Wag76, CKS^{+}05]. Layout [Hab85, Hob00, ES06]. Layouts [Ahu86]. lazy [LK03]. LBPE [LY05]. Leading [Lin02]. leaf [NHK08]. Leafcodes [van86]. Learned [Alo94, KP00, NMP97, TMQM13]. learners [CWO^{+}11]. Learning [BBC00, BM79, COW98, CWH^{+}13, CKLP09, DC00b, FFFP07, GJH01, GK95, KN99, KK90, KLL84, LYSS12, PSR08, PSYZ13, PBQ99, PEF92, RAHT11, SCvW11, SC98, TMN06, USKB10, WZL^{+}03, XYZH11, XYW11, BSMK13, CSJ13, CC11, CMH13, CFM^{+}13, DD11b, EKY08, EL07, FKS10, FLHK08, GCPF08, HOH^{+}07, ML13, OGH04, RL13a, TA11, WRKP05, WS08, WKP13, XST04, YGC13, ZRKZ^{+}11, dSdSF^{+}12, Ano94h]. Learning-based [TMN06, ML13]. Learnt [OH06, CGH08]. Least [FM99, GSV05, Jos94, KSŽ96, PW86, Ber89, MP09b, ZZ10]. Least-Squares [FM99, KSŽ96, GSV05]. leaves [CTM^{+}13]. Lee [BP95, JS87]. left [BKR^{+}89, HOPA91, SG82, WSKH13, WWJ13b]. left-handed [HOPA91]. Legal [KABP98]. Legendre [KP97, SLB^{+}00]. LeMéHauté [Ano95g]. Length [BKW96, Cou81b, DS87, GJH01, Kan88, KLK88, Kis96b, LL97b, VS82, YJ84, BS92, Che08, EL91, Kle13, SGH07, SCCP05]. Lengths [EPRR79, Ell81]. Lens [Gos89, KKH96, PM97b, TTG94]. lenses [BHBF10]. Leonard [Ano93d]. Level [AGW85, DPB00, DG01, DR93, Fel85, Gud82, Ham77, IKS86, JC81, KSW85, KSKB95, KB95b, KU95, LT81, LLSV00, Mar80, ME98b, PA00, RL93a, SLK86, SK83b, SW83b, THCG84, WWW89a, Wec78, WP88, ZOMK00, ZH79, Abu89, BC10, BCDH10, BB03, CLHW94, CU11, DGC12, Dem05, DCS05, FPC^{+}08, HSSH89, HFC96, KK13, KYM13, KS04b, LFL08, MMV06, PSE^{+}11, PD05, Pec91, RCG^{+}09, SM06, SS90b, VCBC88, WZ04, ZYT10, SRTBS91]. level-of-detail [RCG^{+}09]. Level-Set [LLSV00, FPC^{+}08]. Levels [Iiz87, PA97, dBD98, FKS10, KB91b, SSdVL06]. levelsets [TRG^{+}13]. Leveraging [MSI10]. Libraries [DCCL99]. Library [KMGC84, PBN^{+}09]. LIDAR [SO07, ZN13]. Ligature [ASZ99b]. Light [CVP10, HF93, HC94, LZ97a, Man86, Max86, OD97, OD01, VY94, BHSD^{+}13, CAF09, CF07, CFB05, CMD06, Dre96, HASS10, LF08, LB04, SW13, TMNM09, WS03, WNH05, YHS95]. light-field [CAF09, CMD06]. Light-Source [HF93]. Light-weight [CVP10]. Lighting [Bic98, GJ10, LCT09, ZJ05]. Lightness [Bla85, Hor74, LH93]. Lightpen [BN85, HBKN87]. Like [ST80, DAM12, XHJF12]. Likelihood [CT97, CF92, CHRM96, FY85, HD97, HNR84, RB92, HH07, HPR90]. Limb [UZC97, ZB05]. Limb/Terminator [UZC97]. Limbs [LRD99, TGB00]. Limited [Fog93, JGR85, SMD^{+}08, Wam85, Whi93, CD10]. Limited-View [JGR85]. limits [HUF05]. Lindeberg [Ano94i]. Line [AHD94, AHD98, AK85, BL92, BF87, Bid91, CA97, CH99, DLHT99, Ede87, ES81a, FS80, GMW83, Goo92, Gro82, GBB98, GO87, HK93, Ham77, ITN84, JV97, Jar77, JB99, iIK85, KL77, Kim82, KB00, KOY86, KP00, LC85a, LC79, LH88a, LH88b, LD98, MML87, MAN84b, OW83, PKP97, PLL00, Pat79, Peu79, PM97b, Ris89, Rob96b, SP97b, SBT85, Spe92, SM97, TDMT85, TD83, iTTF82, Tsa96, WV97, Wil81, Woj84, WL88, YY84a, ZK81, ZF94, ZZLZ13, dFP92, vvv88, Alb74, Boo79c, CAF09, CDT11, FS03, LF82, Mae90, MSS90, NDO09, PIK90, RL13a, Sha06, SDPO81, UN91, Wal89, WHHB12, YGH11, YY84b, ZXY^{+}12, ZS11]. Line-Based [PM97b, Spe92, ZZLZ13]. Line-Drawing [FS80, SP97b]. Line-Integration [WL88]. Line-Shaped [Woj84]. line-space [CAF09]. Linear [AS83, AM01, BS96, Bau85, BEPW00, BRW88, CT97, FMV93, FTW81, För87, Gar82, HD97, Hor77, HS79, Jac01, KH86, LPR89, LH88b, NB80, NN04, OCON82, Sch92, SS90a, SS91, SC92, Shn81b, Spe92, Stu76, Sur86, SHS03, Van77, Wal88, WZWT99, ZHAH88, Zhu89, And03, AK91, AC09b, Bar05, BS04b, Buz03, CCL04, CSS13b, Fon90, Fra81, IH91, ITNP12, KL07, KORC10, LY05, LX88, PDA03, PL08, QAB^{+}11, Wil78]. Linear-Time [WZWT99, SHS03, CCL04]. linearization [KH90]. Linearize [Piz81]. Lines [AM78a, Cag93, EPRR79, Gaa77, GL97, HE81, JvdBS99, KHB01, MGK00, MM88, MAM97, PR79, RW76, RW79, SLL01, TG96, VS82, Wil81, WR87, BA06, BS05, Buz03, MFA89, RL13b, Sch06, Ste13, Yui89]. lingual [WHN08]. Linguistic [ALK^{+}09, Shl83]. linguistics [JN09]. Link [Ced79a, Mor76, Sur86]. linked [AKC11]. Linking [KVdG^{+}97]. Lipschitz [MR96]. Liquids [FM96]. LISP [DM78]. List [Ano02j, Ano03w, Ano04w, Ano05v, Ano06t, Ano07m]. Literature [Ros00a]. Live [FUS^{+}98, ZB05]. LMMSE [dLAH07]. loadings [Reb89]. lobe [YSL11]. Lobula [MAY^{+}10]. Local [AS88, BW93, Bur83, Chi97, Col81, D’H86, GY99, GBB98, HGA86, HSD85, HS87, HB88, JJ94, KS95a, Kaw82, Kaw83, KP00, LCSL07, Lee81a, LBS80, LS09, Mil99, MB11, PA00, Pat79, PP95, Pen89, RB89, Ree82, SMR98, Sch93, Sel81, Sha79b, SKVS13, TG11, TS00b, Ull79, WN87, ZCL99, BCM13, BL89, BG09, CH06a, CHC11, CK09, DC04, ESS10, FL92, GCFMT12, HBG13, HSJS10, JBR08, LPS^{+}11, MPVF11, PV06, PG13, PTE12, Sah05, SHS03, TCZ^{+}12, TS11, UKH88, VR95, WPS03, WY11, WHHB12, XYW11, YZT^{+}13, YGC13, ZZL13, RK11]. localisation [AW09]. Localization [CYES00, HR99, LSB^{+}00, RAH97, Sto87, BDS12, JLD13, KA12, KMBH09, MN06, RAC^{+}13, ŠRDC09, SIT07, WR08]. Localized [Bat84, MK76, SB00, CLC91, XFSC13]. localizing [MAL10]. Locally [FLHK08, LR90, Yan93b, KL07, LLC11, PK05, dCCP12]. Locate [HdVL99]. Locating [GLD93, Kou03, RKK^{+}00, Sel86, SZ07, TDMT85, Mar89, NH89]. Location [AW98, BKW96, JC94, Shi99, Sug88, WSV91, YH81, LM12, PBG04, SZ03, SM13b, WCF10]. Loci [GK77, SWS11]. locomotion [LE09]. Locus [Rub80]. Loeve [CC97, SB85]. Log [MGMS01, Mas09, Sch06, TP05, UM90]. Log-Polar [MGMS01, Mas09, Sch06, TP05]. logarithm [Hu11]. Logarithmic [WC79]. logging [MR05]. Logic [GV84, HJS89, MCPB00, ALK^{+}09, XP11]. logo [PA10b]. Logotype [Spi98]. loin [CCR^{+}05]. Long [MB94, Sch80a, GBF12, PA10a, YAK^{+}08]. long-term [PA10a]. Look [Tso94]. looking [RP08]. Lookup [Bas81, GV84, Reb89]. Looming [RJ00]. Loop [SBK^{+}99, SS11a, WWLV11]. lossy [YWMS08]. Loveparade [KB12]. Low [ASVO12, DPB00, FTW81, LN10, Mar80, OS87, SLCP85, SK83b, WWW89a, Wec78, ZH79, BCDH10, CSS^{+}13a, DGC12, Dem05, KMBH09, RAC^{+}13, Sau91, WZ04, ZZ10, ZYT10]. Low- [OS87, ZYT10]. Low-dimensional [ASVO12]. low-grade [RAC^{+}13]. Low-Level [DPB00, Mar80, SK83b, WWW89a, Wec78, ZH79, DGC12, Dem05, WZ04]. low-rank [ZZ10]. Low-Resolution [FTW81, LN10]. Lowe [ACB98]. Lower [CT95, Kan98, TC95, Zha97b]. Luca [Ano01s]. Lucchese [Ano01s]. Lumen [SGHM00]. luminance [dLAH07]. lumped [NÇ10]. Lungs [LSB^{+}00]. Lynn [Ano94f].

M [CM94b, CGL94]. M. [CM94b]. Ma [Loh10]. Machine [Ano96c, Bec85, Bri86, BD02, CG87, DP88, EU85, Lee02, Pag92, Ros87b, Td93, WWW89b, BGTG04, Boy04, Dor89, YHS95]. Machine-Independent [WWW89b]. Machines [Hoc87, IKS86, MS94, CMBP09, LP90a, SB13]. macula [QKH^{+}12]. Made [SM93, JWL12]. Magnetic [RSB93, RMFB02, CCR^{+}05, KZD^{+}11]. magnification [YAK^{+}08]. Magnified [FH84a]. magnitude [KM89b]. magnitudes [LMDB11]. Mahadevan [Ano94h]. Mahalanobis [Hsu79]. Maintenance [FH84a]. malaria [TDK10]. Mallat [AM00]. Mammogram [CL97a]. mammograms [CSY08, SRP10]. Mammography [Ano94j, BRSSAL11, RC03]. Man [Bec85, Bri86, SM93, JWL12]. Man-Made [SM93, JWL12]. Manage [SB95]. Management [Tan81a, GM90]. maneuvering [MC09a]. Manifold [GHQ06, Gus07, KADS02, LY13, Pey09, ACS03, AC09b, DMMP03, EL07, GK04, SM13a]. Manifold-based [Gus07]. manifolds [AAASC11, HQ12a, HQ12b, LHYK05, WS08]. Manipulation [CS89, GR87a, JC93, Peu83, CH88, GWCO11, SMT04, SGA12]. Manipulations [BM80]. manual [KSG^{+}13]. Many [Lau97, DOSD11]. many-to-many [DOSD11]. Map [LK97, OMLL98, SB79, Td92, YTTT83, BI11, BB03, BR12, JC06, KS02, KGC05, KORC10, LSC08, SCS91, BR93, CMBV04, DBZ07]. Mapping [BIP00, CGL98, Dan80, DW87, HC96, IS02, KD86, MST85, Sch76, Sch78, SWYP00, Wah83, WD99, AFSW03, CKM11, OMW^{+}07, ŠRDC09]. mapping-independent [AFSW03]. Mappings [Piz81]. Maps [DTG96, GMA83, GSV00, HB98c, HF93, Jok98, KC94, KSKB95, KK94, OMLL98, Pie79, Td93, Yac83, Cou13, DR03, DSdlH^{+}11, DDLP10, FWWT13, GWT09, JBWK11, KS04a, LYSS12, Mas09, PY08b, PMC13, PCR^{+}04, SSL^{+}12, TÉSK11, TC11, WDN^{+}12]. Marching [MS94, TG96, HMA10, OK07]. margin [CGR13, GHZ^{+}13, KSR^{+}12, LLC11]. Markerless [KV06, SHK11, JBWK11]. Markers [GBR79]. Markov [BP05, BD94b, BCM06, GJH01, HE82, HS80, HPvB^{+}10, Kan80b, KBZ96, KH98, KABP98, MCPB00, MSM81, NN13, PJW11, SP97a, SGH07, WKP13, WD92, WB11, YK95, ZSN96, ZZAA92]. Markovian [KWK84, MCPB99, PCR^{+}04, RMFB02]. Marr [For88]. Marr-Hildreth [For88]. Mars [OMW^{+}07, SB13]. Martians [GB96]. Mask [WW88, Nad90]. Masking [OWW85, RCG^{+}09]. Masks [KK79, Rob77, WBB85, Fiu91a]. Mason [WR05]. Mass [FSB85, CSY08, Dem05, NÇ10]. mass-spring [NÇ10]. Massively [CG87, IKS86, MR89]. Match [GBB98, PA83, Shi99]. matched [Tri90]. matches [DLS^{+}09]. Matching [AM01, AG00, BAM87, BR95, Bid86, BDL^{+}06, CN87a, COW98, CL83, CTF^{+}98, CGP85, DC00a, ES81a, GGR01, GL86, HB98b, HS88, HH82, IAP^{+}11, Jok98, KS91a, KPMR91, KP96, KC99, Lai00, Mas02, MAN84b, MM88, MN85, MA83, NG98a, NMP97, PLL00, PC99, PM97a, Pri86b, RB82, RBA94, RH95, Ros93b, SHKP98, SM94, SA95, Tan81b, THT^{+}98, WCH98, WH78, YS06, YJ84, AKC11, BK89, BZS08, BL09, Bre03, CM12, CK11, CJ93, CC07, CK09, CWLJ13, CR03, DIOV06, DOSD11, DSH04, FWH13, Far11, Goh08, GS95, GDR04, HBG13, HQW^{+}12, HZW^{+}10, IHTA90, JKM07, KD10, KMP05, KZ05, KMBH09, LD90, LLC13, LZLP10, LS09, MR90a, MR90b, MAL10, Mil89, Mos91, OBH04, OH04, PLLL03, PFGG09, PMW05, PDTE06, RB89, SAS12, SZ03, SKH08, SBM^{+}06, SY11]. matching [TZY08, UBEP09, WPS03, XHW09, YS09, YK08, ZP11, tHV09, PE09, STLH08]. matching-recognizing [LLC13]. matchings [UKH88]. Materials [Sko86]. Mathematical [Ali77, Ano95g, BDL92a, BB13, Hua80, Lev79, Moo77, Ser86, BDL92b, Fiu89, Fiu91b, Gho88, HR90, RH91]. Mathematics [Åst97, Col97, PEFM98, PRW97a, PRW97b]. Matrices [KP96, ZT80, Gol05, GK90, LPVM13]. Matrix [BGK98, CP79, CZZF97, NR88b, SB98b, SW83b, TI01, TZM98, WC92, ZL01, LLL13, MSI10, Pec91, ZZ10]. Matter [AM94]. matting [HKS06, LYL10]. Max [MS96b, CGR13]. max-flow [ZSCP08]. max-margin [CGR13]. Maxima [AS88, DM81]. Maximizing [WCZ02]. Maximum [BGS83, CT97, CF92, CHRM96, FY85, GHHX04, HD97, Hor77, LL98, Min79, MR92, ND97, RB92, CKK^{+}12, HPR90, LLC11]. Maximum-Likelihood [HD97, RB92]. Maxshift [TVLS08]. May [Bog88]. MCCD [TMT10]. MDS [Mig12]. MDS-based [Mig12]. Mean [KON87, LLR10, LD95, MS96b, MHMO09, ZLS^{+}13, Dou92a, Dou92b, HW06, MSR07, ZYS09]. Mean-absolute-error [LD95]. mean-square [Dou92a, Dou92b]. Means [Boo79a, Mee94, Sha94, BBC^{+}07, CFG06, HS06, JLD12, KH90, Mar90, MJ11]. Measure [ALK99, APV99, De 93, Oka81, Sha94, WWHL88, KN11, LMRMJ08, MGW10, PDK96, RM06, Ros08, TH04, WDN^{+}12, YK08]. Measurement [EPRR79, Kul77, Nev76, OD02, PJ88, PA82, SGK00, TI01, Wah83, FJJ91, HSSH89, JJT91, XFSC13, ZZZ06]. Measurements [BKW96, JJ94, YC78b, BS92, BHMB10, EL91, ELA91, YC78a]. Measures [CY83a, Neg96, RPTB01, SB98a, SK84, VLR84, WV78, YYL96, BAP08, DGZ12, KY06, MM06, Pec91, RKG03, Got08]. Measuring [Car01, CK11, KT08, Ros99c, RŽ05, WHN08]. Mechanical [CLD96, LCD97, DLP13, HW94]. mechanism [GS08]. Mechanisms [YYL96, XO93]. media [HM13]. Medial [BM95, CCMW97, CS01, Sam85, SB98c, SPW96, WBR86, WBR88, CLK09, CK11, CL05, LKC94, SWS11, MDFS11a]. Median [BM86, Dan81a, DN91, Ree82, SWH84, YH81, CDLD77, FKV^{+}11, Wil89]. Median-based [DN91]. Medical [Boo97, BM97, CS89, DUC97, GCB92, GR87a, MAM97, NLW13, SPK^{+}02, TK97, CCH91, CUAT13, HVD^{+}89, KLBP11, KSG^{+}13, MJ11, Mos91, WPK09, YZT^{+}13, YJKK91]. Medicine [CP81]. Mellin [DG01]. Membranes [Pen99]. Memories [MM81, MM80]. Memory [LSZ83, HAKK91]. Mensuration [För87]. Ment [Min79]. Merge [CMVM86, CP79, CLC91, DR03, Lap88, LK03]. Merging [BL00, BS00b, SCvW11]. meridional [Tri90]. Mesh [BVL02, JH98, KG01, LHKC97, ROH88, Tau02a, TTIM96, TGSH98, WSZL13, ACS03, ACH^{+}13, BPG05, FML12, FWWT13, GYH13, HAKK91, dOSJVBS12, Kan80b, LLXW13, LVM04, MR89, MBH^{+}12, MWTN04, MPVF11, SY10, TH12, Vás11, WY11, ZHM11, ZZC^{+}13, ZZLZ13, ZZZY13]. Mesh-Connected [ROH88]. Meshes [Fuj97, IS02, KADS02, LBSP02, MKY01, PSK^{+}02, SRK02b, TC87, Tan95, Tau02b, WH00, AGCA06, BGTG04, CL95, DCL^{+}08, IA03, LBM04, MSR07, PSF07, QHXC12, RCG^{+}09, She03, SBA13, UCB13]. meshing [BO05, OBS06a]. Meshless [MMS^{+}07, SOG09]. meshSIFT [SKVS13]. meta [TFL^{+}09]. meta-data [TFL^{+}09]. Metals [Oka84]. Metamers [DF92]. metamorphosis [CYW04a, CYW04b]. Method [Bid86, Bid91, BRW85, BO91, BGS83, CH87, Ced79b, CMVM86, Cre99, HY98, HW94, KK88b, iK84, KSW85, KB91a, KB95b, KB00, KD82, LBSP02, Mar82, MY95, MY87b, OPR78, OD02, PD83, Pie79, PM97a, PB96, SO01, SY98, SLB^{+}00, SRT01, TB99, WD84, WJW94, WL88, Yam78, Yam79, ZOMK00, ACG^{+}09, AM93, AK91, BYN^{+}04, CYE91, DMW10, Eva06, Eva11, FL09, HDS08, HMA10, KIK89, KK13, Liu10, MCT10, MKA73, NH89, PW06, Pec91, RR06, RL13a, Sal90, SAS12, SSL^{+}12, SM90, SDR91, SCCP05, SDPO81, TM07b, Tri90, Wu02, YB89, YCL07, ZS11, ZCF13]. Methodologies [Jol94]. Methodology [HSSB98, AC09a, DL10, LMRMJ08, LFMP13]. Methods [Bar84, Car01, CR97, CH80, EOS84, FKW98, For72, Gle01, HdVL99, HNR84, II86, JM79, Ku84, KH94, NF06, NA79, OS87, Pat79, Ree80, RFC97, Sar83, SB85, WWL92, vv92a, BPHB91, Bre03, CCTCR09, CMH13, CU11, DFS08, DSY10, GL82, HNB04, KZD^{+}11, Lan91, LTT91, MSR07, MW91, MK05, OEK08, PD05, PBSG12, RN12, SCD11, TT91, WRB11, WB90, XYZH11, ZFG08, ZCK09, RC13]. Metric [KK11, MT84, Por00, SW83a, WPR85, CGU11, FLHK08, FK09, LFL08, MTG07, SMD^{+}08, SCvW11, ZGLG12, ZZZ06]. metric-based [MTG07]. Metrically [KP00]. Metrication [PR79, VS82]. Metrics [Ste01]. Mexican [HQ12b, HQ12a]. micro [TDWH07]. Microbathymetric [SWYP00]. Microcalcification [CL97a]. microcomputers [CYE91]. Micrograph [Sel86]. microscopy [ZMCA05]. Microstructure [WH01]. microstructures [PFV^{+}11]. Mid [PCJC98, KYM13, ZYT10]. Mid- [PCJC98, ZYT10]. mid-level [KYM13]. MIMD [ERW93]. Min [MS96b, ZSCP08]. min-cut [ZSCP08]. min-cut/max-flow [ZSCP08]. Min/Max [MS96b]. Minicomputers [LPF78]. minima [PV06]. Minimal [DBB83, FSS84, GK03, ND97, NSEA13, O’R85, ACS03, BN90, KBJ^{+}10, RL13b, SDC04]. Minimal-delay [NSEA13]. Minimax [KD82]. Minimization [HPB94, Hob97, O’R82, SE11]. Minimizes [KD81]. Minimizing [HJK02, Far02]. Minimum [Fog84, GPP88, LL97b, MRF96, SS84b, Sur86, CXY^{+}09, Kle13, MEYD11, SCMS13]. minimum-cost [MEYD11]. Minimum-Energy [MRF96]. minimum-length [Kle13]. mining [PHY^{+}11]. Minkowski [Gho88, Gho90, LKE98, MP03, PS07, VM06, YPVv81]. Minutiae [UBEP09]. Minutiae-based [UBEP09]. MIRFLICKR [THL13]. MIRFLICKR/ImageCLEF [THL13]. mirror [PA13]. Mirroring [Wec79, TLGS05]. Misregistrations [BBD^{+}94]. Missing [Jac01, MC09b, ZZ10]. mit-es [DM82]. MITES [DM82]. Mixed [SHKP98, WWC82, PV13]. Mixing [AGCA06]. Mixture [MK01, PP95, EKY08, FL09, JWG04, AQ09]. mixtures [LWGP08]. MLESAC [TZ00]. MLS [WHHB12]. Mobile [Ano93d, Ano94g, KK92, KK93, GLOC10, HSH07, MLH13, ST10, ZKRH04]. modal [ABI^{+}04, BCF06, CA10, LYKL12, NT10, RKG03]. modalities [LHJ^{+}09, WHN08]. Mode [Peu79, DAM12, WSC^{+}12]. Model [BCA98, BLd95, BR95, Bri98, BS00b, BS87, CKB96, Car96, CM95, CG04, CS82, DM82, Dou81, GSS00, GPK99, GM94, GY88, GBB98, Gro84, GL97, Gui99, HW81, HKD95, HY98, Jur99, KWK84, KBZ96, KC94, KC95, KABP98, KK92, KK93, KMA^{+}00, LZ97a, LK97, LH93, LHHC98, MS97a, Man86, MWLA99, Mok92, Muk97, Mur87, NFJ93, Nis96b, NH92, PSWH84, RH95, Roh94, SK02, SMK02, SIK92, SI96, Ser80, SSP01b, SLK86, SHD86, SLL01, SH08, SM97, TW98, VV02, VBH97, WW94, WC99, WLI08, YXYW00, YC98, YB01, YH78, ZCCD06, Zuc76b, And03, AC09b, AZN11, AO03, BvdHL^{+}13, BD94b, BCM06, BPB13, BS04b, BH12, Bur81a, CTM^{+}13, CUAT13, CP09, CC03, CC96, DBF04, Dam08, DUSL94, DD11a, DN82, DQ04, EyGS11, FMGA^{+}12, FFY^{+}04, FAB12, GBHS06]. model [Gho88, GHHX04, GH03, GPDR13, HL13, HH07, HG11, HKK08, KK07, KHH^{+}12, KH90, KNO^{+}09, LT05, LM89, LA11, LYG07, LWGP08, LYCG08, LBCA10, LN10, LPR^{+}03, ML13, MAY^{+}10, Mig12, OC90, PE09, PL07, RH06, RLC^{+}11, SKH08, SKU^{+}09, SM13a, SFWG08, TCH07, VAWW10, WB12, WMBY12, WWJ13b, XHW09, YZY11, AQ09, HH05]. Model-Based [GM94, HKD95, HY98, KK92, KK93, KMA^{+}00, MS97a, Mur87, NFJ93, Roh94, SK02, SLK86, SLL01, YC98, YB01, CG04, SH08, WLI08, AZN11, CTM^{+}13, FAB12, GBHS06, GHHX04, KK07, LBCA10]. Model-Driven [CKB96, SM97, DM82]. Modelbase [SB98b]. ModelCamera [PBSM06]. Modeler [BD94a]. modelers [ACS03]. Modeling [Acu92, ACF00, Ano01h, AK96, BY01, CNK01, CMRS98, CP91, CJC^{+}98, EK98, EPB05, Fis94, FPDK12, FS80, Fu80, FJP06, GA13, GP06, Gol13, HF01, HFR06, HF93, IK89, JSRS08, KH98, KB01, LK01, LSB^{+}00, LB98, Man84a, Mas02, MKK02, Mea82, MCPB00, Mil80a, NLW13, PASS01, PF01, RWV95, Rot82, SC00a, SY98, SL96, TTIM96, Tou80, TDT12, TF84b, TGSH98, VFV93, Wyv03, YB99, ZW93, ZTH^{+}11, ZK01, ZNG^{+}13, AAASC11, AGCA06, ACWK06, BGA05, BCDH10, BB11, CLCO13, CD13, CSG^{+}03, DLP13, ES04, FG89, FF09, GHMT09, LCH95, MCQ05, MMP09, SCD11, TFB80, TÉSK11, TF84a, THL03, TA11, WY07, WKP13, YT13, YYF89, YLL12, GHPW12]. Modelled [HFKN97, MB85]. Modelling [GLD93, LVM04, HGSM11, KMN11, LRLB11, PFV^{+}11, PBN^{+}09, PZV13, SKBS13]. Models [AHZ96, BM95, BL98a, BD02, CYW04a, CSDC96, Coh85, Dav97, DB79, DF01, DUC97, EFF98, Fel85, FST94, För87, Fri80, FB97, GS01, GJH01, GSP02, GMT00, GB93, HB98a, HD97, HHI95, HQ12b, Hua80, IP98, Kas80, KCA81, KVdG^{+}97, LVW97, Lev79, LT81, LSBG92, LK00, LT97, MBK81, McC80, MB94, MN95, MFV80, NFSK97, NDC86, Nis97, Nis99, Pha01, Pie88, PA98, Pot87, Rew84, SSP01a, SGS01, Sch80a, SF95, SSJ86, SP97b, SRS11, Spe94, SB00, TML00, TS01, TGSH98, VTG95, WRH97, YKA01, ÁB13, ARARCE11, BGA05, BMM^{+}07, BLH91, BF10, CGH08, CFCP11, CHSV08, CVB09, CSS13b, CMD06, Coh91, CTCG95, CNC03, DCH12, DB03, DSY10, ESS10, Eva06, FFFP07, GCFMT12, HE82, HS80, JEF^{+}12, JBC08, JWL12, Kan80b, KS91b, KMBG09, LSD^{+}07]. models [LKC94, Mar93, MJ11, MCB13, MMA06, MKS^{+}08, NN13, NÇ10, OJRT08, Pec07, Pey09, PHK92, QAB^{+}11, Ros10b, SIK92, SI03, SKM06, SGH07, SRHC13, TMT10, UK12a, UFF06, VM06, XG08b, YK95, ZZC^{+}13, ZZAA92, ZGLP12, DGG08, KS00, KB01, TRG^{+}13]. Modification [Ham77, HW79, KNJ84, Ku84, MBDB88, Peu79, RY95, Dre96]. Modifications [MC95]. Modified [McC82, KHS94, MAY^{+}10, ZGLP12]. Modules [PEF92, DUSL94]. molecular [RG12]. Moment [CT88, DPB00, MTVM04, PR92b, TC86, Tsa85]. Moment-Based [PR92b]. Moment-Preserving [CT88, Tsa85]. Moments [CH87, GY99, KŽ99, LS92, Muk92, SC99b, SLB^{+}00, WH78, YAT97, Dem05, Sal90, San90]. Momentum [ALP06]. Momentum-based [ALP06]. Mondrian [Bla85]. monitoring [ESS10, HMEB07]. Monitors [MV86]. Monocular [CN95, CWC94, FMV93, KSS92, SM93, SGDP01, WN99, WLD99, ÁB13, CC03, DN82, RSPD12, UFF06, dP10]. Monoid [SZKD99]. Morphing [GY01, BGA05, XS04]. Morphological [Ang07, CNDS13, Gad91, GHS95, HT91, Hei99, JC98, LAS94, SH09, Sko86, ZH86, vv92a, Dou92a, Dou92b, LD95, Mor90, SW05, SD90, WR05]. Morphology [Ano95g, BDL92a, HKD95, PLS97, PS95, Ser86, Ste86, BDL92b, BB13, GE08, HR90, RH91]. Morphometric [Boo97, Sah05]. Morse [AC07, CD11]. Mosaic [HMD93, AWK04, SP06]. mosaic-based [AWK04]. Mosaicing [LDD09, CPS10]. Mosaics [GSV00]. Most [Ano12q, Ano13r, Ano07j, Ano07k, Ano08p]. Motion [ACLS98, AC99, Agg83, AS09, AEM98, Ano93d, AT83, BK83, BDVK10, BEPW00, BH83b, CV92, Cav87, CSC96, CP86, Che91, CPK99, CWC94, D’H86, DT96a, Dan97, DWS83, DH00, DC98, DC00a, FH84b, FD99, Gle01, GM94, GB97, HZ86, HA93, HH97, IF99, Jac01, JMA79, KS91a, KG94, iIK85, iK85, iK86, iK87b, KN03, KJRA96, KH86, KC99, Law83, Lee76, LK91, LS01, LNY83, Lin02, LJ89, LH88a, LH88b, LHHC98, MNE00, MS97a, MBH^{+}12, MB94, MG01, MS96c, NK00, NP92, Nev76, Oli00, Oli01, Pen92, Pen94, Pen99, PCP02, PSM80, Pot77, Pra81, RS91a, SA91, SA96, Sav87, SP97d, SGDP01, SF97, Spe92, Spe94, SM94, SK83b, SW86, SBZ97, TO99, TP92, TS01, TTIM96, VF96, WLD99, WF02, WD96, WA87, WY91, XL98]. Motion [YH83, YK97, ZF94, ZHAH88, Zhu89, ALP06, AS08a, ACG^{+}09, BS05, BF07, BC10, BT05, CG09, CMBV04, CFCP11, CMBP09, CRH05, CH06b, DGC12, FM91, FLB06, Fit88, FSF07, GZP05, GBHS06, GODC07, GW07, GWT09, HSH07, HMF10, HH91, HRC09, HC13c, KBN12, KM89a, KHK10, KH90, KL10, LCSL07, LMRMJ08, LB06, LH90, LL06, Lhu08, LcTT91, LTT91, LX88, LZWP03, LWH03, LYA13, MPF07, MU11, MHK06, MP09b, NFM08, NT10, Neg12, Pan03, PD05, PW06, PY08b, PV06, Pop07, RLS06, RN12, RSPD12, SKM06, TLT91a, TMQM13, TCMS04, TP05, TR09, TLMT^{+}05, UK12a, UFF06, VSP06, WRB06, WS06, WW91, XYW11, YWZ11, YS06, YNCO11, YC05, YSD03, YR06, ZBLS13, ZB05, ZS09, ZT09, LY13]. Motion-Based [NK00, WF02, MBH^{+}12, KL10]. motion-blurred [CG09, LcTT91, LTT91, TLT91a]. Motion-Egomotion [DH00]. Motion-Model-Based [LHHC98]. Motions [BA96, HJK02, Bar05, KV06, KL13, RRR11]. Motivated [BL98a]. Motor [BH86]. mouse [TTH07]. Movement [BL01, Gav99, HF01, HFR06, ITNP12, PQML11, WS08, MAY^{+}10]. Movements [KS95b, Roh94, SFWG08]. movies [SZ03]. Moving [AN84, FT79, GBR79, HJ83, HC96, JN93, Pra81, RRS83, SMK02, WD96, WZ97, WN86, WN87, YMA82, ACH^{+}13, BP09, CD95, CYC10, CCYC12, DN82, JKM07, KG90, LD90, OCVV04, QC04]. MPEG [ADDK99, FSF07]. MPEG-4 [FSF07]. MPM [CMBV04]. MPU [BMM^{+}07]. MR [BvdHL^{+}13, CFYU12, DCS05, HRS02, LPS^{+}11, LSB^{+}00, ZU09]. MR-image [CFYU12]. MRF [GJP96, KL11, SKH08]. MRFs [AKC11, KTP08]. MRI [GPP88, GPDR13, RAH97, WSKH13, WWJ13b, ZRL^{+}11]. MSS [HW79]. Muller [RP88]. Multi [AMMV99, BDS12, BF10, CJ93, CPT07, CPS10, Di 86, Dre94, HFC96, ITNP12, KK13, LS08, MFB11, OS87, PLL12, Pat13, Pen03, PMC13, RJ94, She96, SCL13, TMT10, WJ07, WP93b, ABI^{+}04, Ano06m, AKC11, BKK11, BSMK13, BCF06, CA10, CPP^{+}11, CD10, CWO^{+}11, DR04, DD11b, DCS05, FBF08, GKK05, GCEC07, HC13c, IJDAB13, KD10, KW12, KL10, LBNS09, MB11, NN13, NT10, RM03, RCTV12, RKG03, SSL^{+}12, Sta05, TRS06, UM05, VRKL13, YWZ11, YGC13, YCKA10, ZRL^{+}11, ZH04, ZNG^{+}13]. Multi-agent [KK13]. Multi-camera [MFB11, CA10, HC13c, KD10, RCTV12, YCKA10]. multi-channel [IJDAB13, NN13]. Multi-class [Pen03]. multi-colored [DR04]. Multi-core [TMT10, KL10]. Multi-Edge [She96]. multi-feature [CWO^{+}11]. Multi-illuminant [Dre94]. Multi-Image [Di 86]. multi-instance [FBF08, YGC13]. multi-Kalman [Ano06m, GKK05]. Multi-level [HFC96]. Multi-matching [CJ93]. multi-modal [ABI^{+}04, BCF06, CA10, NT10, RKG03]. Multi-object [SCL13, ZNG^{+}13]. multi-perspective [CPT07, ZH04]. multi-phase [DCS05, IJDAB13]. Multi-Resolution [OS87, AKC11]. Multi-resolutive [Pat13]. Multi-scale [AMMV99, BDS12, LS08, PLL12, WP93b, BKK11, LBNS09, SSL^{+}12, Sta05, TRS06, VRKL13]. multi-scale/irregular [VRKL13]. multi-sensored [CD10]. Multi-spectral [CPT07, GCEC07, ZRL^{+}11]. multi-subspace [DD11b]. Multi-target [PMC13, KW12, UM05, YCKA10]. multi-task [BSMK13]. multi-user [YWZ11]. Multi-view [BF10, CJ93, CPS10, ITNP12, RJ94, WJ07, CPP^{+}11, MB11, RM03]. Multicamera [Mur95, TA11]. Multichannel [FSB85, RDM^{+}11, ZGK95]. Multicolored [MS00]. Multicriterion [WL89]. Multidimensional [HD97, Han93, Her92, KP96, KTNO97, Udu94, WPR85, MJ11]. Multidirectional [Lam84]. Multifactor [PQML11]. Multifingered [SKOS95]. Multiframe [TO99]. Multigrid [Enk88]. multilayered [KK07]. Multilevel [HS88, OMLL98, PG94, Ter83, HDS08, KMT11, OBS05]. Multilocal [LLSV00]. Multimedia [STLH08]. Multimodal [JS07, LDC^{+}13, MKK02, PY08a, YKA01, KT07, OH05]. Multimodality [GCB92, Mos91]. Multinary [BAM87]. Multiobject [LC88a]. Multiocular [LRD99]. Multipart [BLP95]. Multiperson [IB01]. Multiphase [WSKH13, NHSC09]. Multiple [BA96, CFM02, CM95, CA86a, CCS01, CJC^{+}98, CM99b, EFF98, FW97, FMR01, GA91, GK95, HS87, HH12, JN93, Jok98, LT90b, LV96, MFJ95, MY95, Mas02, MS97b, MKY01, Nis95, OD99, OD02, PA10a, SU01a, SU01b, SCS99, Spi98, SB94, SA95, TA88, WD96, WH01, WB01, YSD03, BSMG05, BL09, BPB13, CKM11, CHH09, CYP^{+}10, CS10, CH11, CUSZ07, CZZS07, DUSL94, Gol05, JRH03, JBC08, KV06, KN03, KK95, KN04, KHK10, KPPK09, LF08, LLR10, LHJ^{+}09, MD95, MMV06, MMA06, Mas09, OGH04, PA06, PT08, PD11, ROJX09, SSdVL06, SD90, SYPK13, SH05, SH08, TRG^{+}13, UK12a, WRKP05, WS03, WDB12, WS90, WHN08, XST04, YSL11, ZC89, ZQ11, dSdSF^{+}12]. Multiple-Attribute [GK95]. multiple-lobe [YSL11]. Multiple-Order [HS87]. Multiple-seed [GA91]. multiple-view [CH11]. Multiprocessor [ERW93, BTNS90, HAKK91]. Multiregion [MMV06]. Multiresolution [BK89, BBF^{+}11, BMZB02, CKB96, CW94, CL97a, FKW98, GM94, GR87b, Har85, HMD93, JTEA91, Kri84, LS01, LH92, MPC94, NDC86, OP96, PA97, Rei96, SL96, SMB95, TW98, TM86, VBH97, YW99, CGL92, UCB13]. Multiscale [BM98, BRdBS99, DT97, FWH13, GJP96, HPB94, Hu11, KVdG^{+}97, Mok97, NDN^{+}97, NVWV97, PM89, PB99, Ros93b, BA89, BNG03, BNG05, DAM12, NBDB04, SH09]. Multisensor [LMM95, MN95]. Multisensory [MA85]. Multisource [FTW81]. Multispectral [AM06, Coh85, DWX83, Cum91, PCR^{+}04, ÜB05]. Multitemperature [KBZ96]. multitemplate [BPW91]. Multitemporal [DWX83]. Multithreshold [WH84]. multithresholding [O’G94]. Multivariate [Kas80, TLEF06, AQ09]. Multiview [DF01, BY12, LYA13, UFF06]. Must [Hoc87]. Mutual [KT07, PMV00, PC05, ZKRH04]. My [Uhr86]. myopia [JBS^{+}91].

N [Ano94g, And03, ZSCP08]. naive [Buz03]. naivete [JBS^{+}91]. Narrow [AS08a, Mil09, MBMC11]. Natural [AK96, AOR94, DF92, GM85, HWW06, LK01, LW85, MMN83, Rub80, SHJB^{+}83, CTM^{+}13, LBNS09, Mig12, MG95b, Sta05, YWMS08, YYF89]. Navier [CD95]. Navigation [BH83c, GSV00, KK92, KK93, KR99, RJ00, Sug88, ILRB04, ŠRDC09, TDWH07, Ano93d]. Navigational [RR95]. Near [KADS02, CHC11, HASS10, JN09, KK90, TMNM09, ZTH^{+}11]. near-duplicate [CHC11, JN09]. near-duplicated [ZTH^{+}11]. Near-Optimal [KADS02, KK90]. Nearest [CGU11, KHH^{+}12]. Nearest-neighbor [CGU11]. Necessary [VK92]. Necklaces [GSP02]. Needs [WN87]. Negative [Gho91, LLL13]. neglect [HH05]. Neighbor [Sam82a, Sam89, BMR91, CGU11, KHH^{+}12]. Neighborhood [JR86, MMS97, MKK02, MC95, Sob78, YK80, GHZ^{+}13, Hu08, MLF^{+}12, NSEA13, PMR92, SW04, TBN95]. neighborhood-sequence [NSEA13]. Neighborhood-Uniform [JR86]. Neighborhoods [CM99b, DW87, Rag92, Yac83]. Neighboring [SW83b]. Neighbors [KCA81, Sch92]. Neighbourhoods [SB02]. Nested [Ley87a, TS00b]. Nesting [AGW85]. Net [WRH97]. Nets [AMMV99, GLM78b, MAM97, GLM78a, TLEF06]. Network [Ano94g, CGL98, AVBK10, GFW13, SG82, Ziv10]. Networks [OM84, SB95, SC00a, SC97b, SC98, BSM10, BPS10, BBB96, DDLP10, KPKPW90, Li92, LWH03, MCT10, See89, SJ12, SST06, TN07, ÜB05]. Neural [Ano94g, CG87, CGL98, GM87, SC97b, SC98, Wat87, WRH97, BBB96, GFW13, KPKPW90, TLEF06]. Neuron [Sel86]. neutrosophic [SG11]. News [FM84, TL78, WHN08]. newspaper [WS89]. Ninth [SRK02a]. Nlc [JR86]. No [Pri86a, Tso94]. Node [AK78, JRV82, JRV83]. Node-Rewriting [JRV82, JRV83]. Nodes [SC97b, PL08]. Noise [CY83b, Chi97, Her80, HNRR90, Imm96, KWK84, KD86, KC01, Lee81a, Liu93, Mas85, Ols93, SRML09, TO99, Wil98, WWL92, ZH79, Imm91, LJ90, MGPF08, RK11, YL08, ZG91]. Noise-Smoothing [CY83b]. Noisy [BP84, GM94, GW01, KSG84, LR02, MM92, MF77, PSK^{+}02, RRS83, Sha78, SH84, Tas77, TSK94, WD92, ZK81, vYB89, FM91, HK93, KGC05, KHS94, LC88a, LBCA10, Sha75, Tri90, VRKL13, ZZAA92]. Non [BY12, CMD06, Con88, HBA93, LBCA10, PRR03, Pen92, PDA03, BHBF10, BPS10, BDS12, BS04a, CR03, DMMP03, FB05, Fon90, GK03, GRB13, GW07, HSJS10, HC13c, JSRS08, KORC10, LJHH07, LLL13, Loh10, MMK04, MPVF11, NLM05, PA13, RKG03, Sha06, SKH08, SAC09, SB05, TMQM13, TLCH05, WR08, YC05, ZZZ06]. Non-accidentalness [HBA93]. non-central [PA13]. Non-Concentric [Con88]. non-contact [NLM05]. non-conventional [BPS10]. non-cosmetic [BHBF10]. non-cubic [SB05]. Non-Gaussian [LBCA10]. non-intrusive [YC05]. Non-linear [PDA03, Fon90, KORC10]. non-local [HSJS10, MPVF11]. non-manifold [DMMP03]. non-metric [ZZZ06]. non-motion [GW07]. non-negative [LLL13]. non-overlapped [LJHH07]. non-overlapping [HC13c, JSRS08]. Non-parametric [CMD06, BDS12, MMK04]. non-radial [WR08]. Non-rigid [BY12, PRR03, Pen92, CR03, GRB13, RKG03, SKH08, TMQM13]. non-simple [GK03]. non-SVP [FB05]. non-topology [Loh10]. non-uniform [SAC09, TLCH05]. non-vanishing [BS04a]. non-voting [Sha06]. Nonanalytic [SCS99]. noncentral [GA09]. Nonconvex [Bd96, WN86]. Noncoplanar [CRC97]. Nonfuzziness [WCZ02]. Nonlinear [CRC97, CBM01, Ein83, EL07, FS95, GMG92, KS96, NVWV97, Pan78a, Sch76, Sch78, TGSH98, YBDC93, vYB89, CCH91, DAM12, LV11, PW06, SCvW11]. Nonorthogonal [RBA94]. Nonparametric [LB97, ZOMK00, BCMCB09, YHN11]. Nonperiodic [AF81]. Nonplanar [LTS93]. Nonrigid [ACLS98, Ano01r, CP86, FDMA97, FT98, GSST03, KG94, LPR^{+}03, Pen94, Pen99, TGSH98, CBD^{+}03]. NonSelf [KM00]. NonSelf-Occluding [KM00]. Nontriangular [KG01]. Nonuniform [Chi97, WP93b]. Nonuniformly [MM92]. norm [DOSD11, ZW03]. Normal [CD92, PSK^{+}02, CRT90, HC13c, TH12, YA12, ZQ11]. Normalization [RY98, CM12, Hu11, LDGS^{+}13]. Normalized [CA84, LGJ82, WWHL88, WM93, GH08, WW80]. Normals [WL88]. Norman [Ano94f]. nose [NB10]. Note [Ano01l, Ano01m, Ano01n, Ano02i, Ano03q, Ano13s, BS89, BM95, DC88, Di 86, EU85, FS85, HZ86, Hen84, Man84a, MBK81, O’G88, Rab92, Rub82, SBT85, TC89, Vos88, WWW89a, WWW89b, WX91, Bor91, Hor79, Kul79a, Nad90, Wal89, Ano06n]. Notes [LR12]. Notion [Gho91, Sab76]. Novel [APV99, CCP97, KHS94, KR99, LBSP02, SK01, CKLP09, CU10b, DK13, KBN12, PCC13, YC05, ZSCP08, ZCF13]. novelty [WHN08]. NP [HW94]. Nuclear [CP81]. nuclei [BEH^{+}81]. nucleotides [RG12]. Number [Ano92b, Ano92a, Ano01s, GW93a, KK93, Oli01, SC97b, WSV91, APB10]. numbers [HY11, LP91]. Numeral [HY98]. Numerical [DFS08, FH84b, Mai76, HKM12, KBJ^{+}10]. NURB [Ano95h]. NURBS [HKM12, KP12].

Object [ACF00, AKL93, AW09, AA93, AW98, BBC00, BB03, BJ86, BZ99, BSF02, CF01, CGL98, CK87, CJ93, CS98, CEC^{+}80, CS00, DUC97, DCTO97, DC00b, GLD93, GV78, GBL08, GK95, GR87b, Haa82, HR99, Hod95, HP96, ILRB04, KMB97, KA94, KP96, KP00, Lau97, LI00, LT90b, LD98, LLC12, LWH03, MDFS11b, MFJ95, Mas02, MKK02, May99, MNSK98, NG98b, Nag78, NDC86, NA79, OG98, PS05b, PF87, PR92b, QV98, RW97, SU01a, San78, SF95, SN99, SY98, SGB01, SHD86, SLL01, SB94, Sta95, SBxx, Sto87, SKBS13, THN92, TTIM96, THCG84, TCC90, US96, Wag76, WWB84, XAB07, YMA82, YK87, YK86, YT99, YC98, You86, ZZZP09, ZYS09, ACAAC^{+}08, AT13, AHDM10, BSM10, BL04, BPB13, BH12, CHH09, CS04, CWO^{+}11, CL08, CYC10, CCYC12, DPR92, DHP08, ES04, FFM05]. object [FFFP07, FLCdA06, FR11, Fit88, Fly92, GB10, GM90, HYJ11, IK89, IH91, JEF^{+}12, JBR08, Kar89, KRK11, KBD^{+}12, KS04b, KH13, LV03, LMRMJ08, LH90, LD90, LL12, LC09, LAL^{+}10, MHSP10, MM90, NDO09, PE09, PSE^{+}11, PSR08, Pen89, PL10, PZV13, STV09, SZ07, SCL13, ST10, SA90, SIT07, SFWG08, TG11, TAK09, TID07, TC11, UA90, WDB12, XYZH11, XST04, YZY11, YNCO11, YJA96, ZYT10, ZNG^{+}13, ZCK09]. Object- [LI00]. object-action [KRK11]. Object-Background [KA94]. Object-based [LWH03, LMRMJ08]. object-centered [SCL13]. Object-level [BB03, PSE^{+}11]. Object-Process [LD98]. Objective [SJST07, SYPK13]. Objects [AN84, BU93, BAM87, BL92, BLP95, BH99, Bid92, BN84, Bie87, BC85, BH95, CM95, CA86a, CA86b, FHMB84, FT79, GR87a, GR85, GESB95, Gro82, GB93, HCHD01, IE99, JT80, KK88b, KII98, KMGC84, Kul83, LF96, LM99b, LCC89, LK00, MA85, MS97b, MS00, Mer81, MKW94, MB85, NL96, ND97, Nur86, PF87, Pot87, RRS83, Rew84, RGC87, SK02, SU01b, SMK02, SB93, SCS99, Tas77, Tay00, TF84b, TGSH98, VKP98, WD96, Wec81, WA87, WN86, WN87, Woj84, ZF94, AVBK10, And03, Ano06m, BL08, BS04b, BBB11, BB11, BP09, CKLP09, CUSZ07, DR04, DGC12, DMMP03, DBB13, GKK05, Gar82, GB08, GRB13, HRC09, HH82, JKM07, KMP05, KS12, Kri92, LA11, LYKL12, Mar89, MHMO09, MSF^{+}12, NÇ10, OCVV04, PA10a, PLLL03, PS07, PHK92, RNDA13]. objects [Ros80a, Sal90, SIK92, SOG09, TLGS05, TF84a, UKH88, VK91, VZP^{+}09, WRKP05, XOF05, XL88, YHN11]. Oblique [HGv87, LSC08]. observable [HPvB^{+}10, ZT09]. observation [Dou92a, Dou92b]. Observer [Pra81]. Obstacle [LB98, CSS13b, MTAA11, VD90]. obstacles [CD95, KG90, LHM06]. Obtain [Che98, SSL^{+}12]. Obtaining [CJ82, KM03, RJ94]. Occluded [HFKN97, Rut82, WH96, OBH04, OH04, PLLL03]. Occluders [ASZ99a]. Occluding [KM00, Sau99, ZM96, Beu91, SD92]. Occlusion [BHNR93, CLZZ13, CTE95, CN95, FK00, HKA13, Lai00, Ull83, BTCH05, CH11, HH12, LST13, MSSS09, YG07]. occlusions [PA10a]. Occurrence [CP79, ZT80, GK90, KP96, LPVM13, Pec91, PA10b, THCG84]. Ocean [SWYP00]. OCR [CB98, LZ97b]. Oct [JT80]. Oct-Trees [JT80]. Octants [AGW86]. OCTOR [JR09]. Octree [AN84, Mea82, Pot87, SA90, Sze93, ZZZL13]. Octree-based [ZZZL13]. Octrees [CA86a, CA86b, Sch92, WA87, Yau84, CH88, LPR89, Sam89, VS08]. octtrees [Gar82]. Off [AHD94, AHD98, DLHT99, MR96, BK07, HER81, KK11]. Off-Line [AHD94, AHD98, DLHT99]. off-the-plane [KK11]. offensive [AO04]. office [OGH04]. offline [KSR^{+}12]. offs [LHH^{+}98]. Offsets [PLR96]. Old [PB96]. Omnidirectional [BI10, OYTY98, SS09, BPS10, CYP^{+}10, PBSG12, SST06]. on-board [GSPL10]. On-Line [ITN84, dFP92, NDO09, RL13a]. One [CWSI87, ES81a, Eva11, FL87, Lyn81, RRS83, Tso94, Ull81, Ull83, Fra89, GSV05, GBW89, SCS91, WSV05, Eva06]. One-Dimensional [ES81a, Lyn81, RRS83, Ull81]. One-Pass [CWSI87, Fra89]. One-to-One [FL87]. Online [BSM10, NHY10, WWLV11, TMQM13, USKB10, YCKA10]. Only [CN87a]. onto [WD99]. Opaque [Sau99, WR93, WR96]. open [DSdlH^{+}11, NRJ11]. OpenCV [SM10]. Opening [HKD95]. Openings [BJ96, RH91]. operates [Fra81]. Operating [KH83a]. Operations [Bau85, CD93, KDK78, Kaw82, Kaw83, Kle80, MK76, NK00, Ree82, SS90a, WW88, YK87, Elb05, Gho90, SHS03, TBN95]. Operator [For88, Gra78, Gro82, Sab76, SC92, Shi86, vYB89, UM90]. Operators [BH83a, CTH84, GV84, GHS95, HRS02, Hei99, Man84a, PD79, PC89, Ull81, Yam79, Yam80, YPVv81, ACS03, Ang07, Gho88, GR05, KM89b, MW91, VBS^{+}04]. opti [NT10]. opti-acoustic [NT10]. Optic [PA83, ZHAH88, CSS13b, Mar07, QKH^{+}12]. Optical [BK83, CV92, CH87, CM94a, Enk88, FSA01, FSV07, HZ86, HJS89, JW87, Jea11, JM09a, iK86, iK87b, LSVD85, LHH^{+}98, Mai81, MNCG01, Mit88, Muk97, NDBT95, Pra81, Pra83a, RDM^{+}11, SP97d, Spe97, SB00, SS84a, TS00a, TA88, XS98, ZI87, BC91, BL09, DRAB08, GYTL09, GPY^{+}07, HMF10, IM06, KN11, LB10, MN06, MZC^{+}05, MCF10, RPG12, SM06, TLCH05, TDWH07]. Optical-flow [JM09a, DRAB08]. Optical-Model-Based [Muk97]. Optically [Nag78]. optics [FB05, kWwZ13]. Optimal [ADDK99, ACDB12, BAM87, BP84, BR95, Dou92a, Dou92b, Jos99, Kas94, KADS02, LH99, MEYD11, O’G88, PV06, RSv89, RBA94, SB91, Sbe00, SS87, SC92, SK84, THT^{+}98, WW88, YHS95, ZM94, Bha91, CL91, DBF04, KK90, KMP05, LS91, SC93, SS11b, WLMG08, Dou92a, Dou92b]. Optimal-flow [MEYD11]. optimally [HKK08]. optimisation [LM12]. Optimised [Vás11]. Optimising [XG08b]. Optimization [AEH79, BVL02, BS92, FB97, GKR02, HL76, IW97, Jon97, KKK99, LPS^{+}11, MK05, Pel84, TGSH98, Ull79, AS09, BRA^{+}10, BPB11, CMH13, HG11, HZLM11, KIK89, KL11, KLBP11, Li92, LD95, OEK08, PB11, PZ08, PZ09, PW06, TL05, YSL11]. optimized [SM10]. Optimizing [BW98, PKP97, Wil98, KTP08]. Optimum [KA94, CFYU12, WY11, dSdSF^{+}12]. optimum-path [CFYU12, dSdSF^{+}12]. options [TVLS08]. ORASSYLL [KP00]. Order [CV92, HS87, Kri84, Nag83, PA98, RM02, VF96, WW95, DS90, DD11a, KA08, PL08, Sub90, ZZP12]. Ordered [Pud98, Rag92, Zha97a, Ang07, See89]. Ordering [MMS99]. orders [CBK03].

Ordinary [FM99, Sug93]. Organ [NSK^{+}97, SP81, BvdHL^{+}13]. Organization [ACF00, AD93, ASZ99b, BSF02, MHN84, SB98a, SMK02, Sau99, Smi85, HGS08]. Organizations [SSN78, Tam83]. Organized [KP00]. Organizing [CG87, KS04a, TLEF06]. Organs [HL79]. Orientation [AD86, BH95, Sav87, SL85, TS86, TCCK90, YK86, Zuc85, ZI87, CF07, Dre96, JWL12, RFS03, WZ04]. orientation-from-color [Dre96]. Orientations [SK83a, ZJ05]. Oriented [Her93, KW87, OS95, SJ93b, SHD86, You86, FYH11, GZJ05, HL13, PCC13, RS91b]. orienting [Sal90]. original [AK91]. Orthogonal [CL00a, FB97, HAKK91, Kub84, Ros98b, WW88, BKR^{+}89, KA12, nLPR91, LFMP13, MPV13, VPAM12, YGH11]. Orthogonality [KL93]. orthogonally [DBB13]. Orthographic [Che91, iK86, TA88, LCT09]. Orthographically [SS84a]. Orthography [SKK83]. oscillations [Boy04]. Osteogenesis [TZ82]. Other [For72, GMG92, Ree82, Gho90]. out-of-focus [FM91]. Outdoor [BD02, CPC08]. Outlier [DF02, LE09]. Outlines [Hob97, NA79, Got08, LYG07]. Output [Lee76, SB89]. Oval [Ros98b]. overhead [PE09]. Overlap [GCB92, MSW96]. Overlapped [Eng83, LJHH07]. Overlapping [LAS94, NS98, BEH^{+}81, EKY08, Gol05, HC13c, JSRS08]. overview [Pop07, RWD90].

P [Ano95g, Hor79, Kul79a]. P.-J [Ano95g]. Pacific [SRK02a, Ano04q, BSW01, HCS03, KS00]. Package [TST^{+}83]. Packaging [Oka88]. Packet [TS00a]. Page [Ant98, GT84, KSI98, PZ92]. pages [Ano01s, GW93a, KK93, Oli01]. paintbrush [ZG06]. Painterly [PY08b, XTLP04]. painting [LF04]. paintings [CHL05]. Pair [DF02, DH00, SA91, SA96, YC78b, Tri90, YC78a]. Pair-Wise [DF02]. Pairs [RFC97, TS86]. pairwise [Gol05, RM03]. Palette [SS95a]. palm [ABEN09]. Pan [CC00, SP06, DDLP10]. Pan-tilt-zoom [SP06]. panorama [Che08, DWB11, ZH04]. panoramas [BDL^{+}06]. Panoramic [FB05, HH98, KW99, MAL10, ZKRH04]. Paper [Ano07j, Ano07k, Ano08p, Ano12q, BKMSR98, CGL94, Nag86, Rab92, Ano13r, Hor79, Kul79a]. Papers [Ano01h, Ano01q, Ano01r, GHPW12]. Parabolas [Boo79a]. parabolic [Ste13]. paracatadioptric [BA06]. Paradigm [MML87, ZN08]. Paradigms [FUS^{+}98]. Parallel [AS93a, AM78a, AW98, BP84, BRW88, BCG95, CG87, CD93, Che98, CWSI87, CCS95, CDLD77, DB88, DRCF95, DW87, ER96, Eng83, FK83, GH90, GH92, HFC96, IKS86, IW97, JN93, JRV82, JRV83, KPS76, KSS97, Kas94, KBZ96, Kaw82, Kaw83, Kle80, LHKC97, LH99, LHB87, MS96a, MW00, MNHO00, Mon84, OR81, PK99, Pog85, Ree80, Ree84b, RF02, ROH88, RW76, RW79, SKS11, SM97, Tan95, THT^{+}98, TF81, WBR88, YK87, CL91, GS90, LM91, Li92, MR89, MHSP10, SS78, SBA13]. Parallelepiped [KK79]. Parallelepipeds [Med84]. Parallelism [Fel85]. parallelograms [KK09]. Parameter [CD92, För87, GS99, SP97a, SC00a, SCS99, ZZAA92, HD09, LM89, SRTBS91, Sah05, SS11b, UTB^{+}11, XWYY10]. Parameter-Controlled [GS99]. parameterization [ALP06, CHZ^{+}13]. parameterizations [NESP10, TJ12]. Parameterized [WSSD96, YB99, DB03]. Parameterizing [ANM98]. Parameters [CT97, CSC96, CL00a, KC95, KC01, Muk92, QY02, SMR98, Udu81, WSV91, YK97, BF07, CYW04a, CYW04b, FM91, GA09, KM94, KM89a, KY06, LMC09, NÇ10, PA13, RAC^{+}13, TA11]. Parametric [AEM98, BCA98, BA96, DM01, Far86, GBHS06, Gui99, KM84, LVW97, LL95, Pag99, PS83, QAB^{+}11, SAG84, Ül01, WF02, AFSW03, BDS12, Bot78, CMD06, KA08, KGC05, KH90, KNO^{+}09, MD95, MMK04, MP09b, ZQ11]. Parametrically [LC79]. Parametrization [BGK95, Cha83, KSŽ96]. Paraperspective [Chu02]. Parasite [TDK10]. park [CPC08]. Parsing [BC88b, LP77, DGG08, MDFS11a, PSYZ13]. Part [BL92, EU85, KS04b, CWO^{+}11, LAL^{+}10, PS05b]. Part-level [KS04b]. part-sense [CWO^{+}11]. Partial [BHNR93, Lai00, Pla96, KS03, KMP05, LPR^{+}03, MB95, SKVS13, XOF05]. partial-surface [XOF05]. Partially [HFKN97, GB13, HPvB^{+}10, Oli91, OBH04, OH04, PLLL03]. Particle [DD11a, LST13, BW11, BL09, BKMV07, DUSL94, HBB^{+}12, Mar93, MEYD11, MHSP10, SBB10, YNCO11, RRR11]. Particular [Lin02]. Partition [CCTCR09, ABD11, BW11, MWF07, OBS06b, TRS06]. Partition-distance [CCTCR09]. partitioned [WDB12]. Partitioning [KM00, SB98b, DBB13, MMV06, MMK04, TL05]. Partitions [FSS84]. Parts [LF96, RJ94, RDR95, DHP08, LLC12, PA06, PYS03, ZZZ06]. Pascal [LS92]. PASHA [CBD^{+}03]. Pass [Cha81, CWSI87, CCS95, OS87, DV82, Fra89]. passers [MLH13]. passers-by [MLH13]. Passive [BH83c, TS86, BS04b]. Patch [Dub76, VV02, GFL^{+}11, GC80]. Patches [BM97, CMPP99, Gos89, SZ96a, Whi93, Far82, PZV13, TCCK90, XYW11]. Path [DJG01, SU01a, YYL96, CFYU12, DGZ12, KS02, MZB^{+}10, dSdSF^{+}12]. Paths [BV99, Sur86]. Pattern [Arc81, Bai88, Big97, CN87a, CG87, CCP97, D’H86, GL86, HB98c, HS83, Jar77, KC99, LSMS85, McC80, MT00, Sab76, SA79, iTTF82, Woj87, BRP04, Fla89, JR09, Kaw78, LC88a, MGPP11, YR06, ZC89]. Patterns [AD86, AHZ96, Bd96, KPS76, KW87, Kod76, Kub84, MS85, Mil80b, ME98a, Nis97, SSN78, SJ93b, SRL82, SK88, TDMT85, Wah83, WR93, AT89, BMR91, BHSD^{+}13, GWT09, IO09, MB05, MB11, TLGS05, UKH88, YLM11]. PCA [DBBB03]. PCB [MEDT96]. PDE [DQ05, MPST08]. PDE-based [DQ05]. PDQ [SBA13]. peak [Imm91, Sez90]. peaks [FS03]. Pedestrian [BBC^{+}07, DZL07, GSPL10, KRJ^{+}08]. pedestrians [ST07]. peer [MGPF08]. pentahedron [LYCG08]. Pentland [Dre96]. Penumbral [GM79]. People [HCHD01, HF01, Hoc87, MJD^{+}00, PF01, CHP^{+}11, CZZS07, GLOC10, HFR06, HH12, DFP^{+}13, PMC13, SH05, TMB12]. Perceived [Hab85]. Perceptible [Nag78]. Perception [Ano94g, BC92, Bec85, Bec86, Bri84, Bri86, Ger85, GM87, Ku84, MJS97, RL93a, SL85, Smi85, SGDP01, Wam85, Boy04, FY06, OH05, SB96a]. Perceptual [AD93, ASZ99b, Bar85, BSI87, CH96, CCP97, JDP97, SB95, SMK02, Sau99, SN99, SPK^{+}02, WH96, AT89, GZP05, LBNS09]. Perceptually [IW97, SM99]. Percolation [PA98]. Perfecting [CLD96]. Performance [BS00a, BG09, Car01, CR97, DB94, Har94a, Har94b, HNR88, HNRR90, KTP08, LPH01, MM06, PDK96, PEF92, SGB01, Shi94, Tam83, TCB^{+}08, TS01, TSK94, VD10, WWW89b, WVL81, WH94, WWL92, YJ84, Ano05p, BHBF10, BGPD09, BLH91, DRAB08, FBF08, HBF09, HC13b, KB91b, LCP90, TM07a, UPBS08, Wu93]. performance-based [TM07a]. Perimeter [Kul77, Wec81]. Perimeters [Kul83]. Period [GLR^{+}99]. Periodic [LAS94, RSPD12]. periodicity [SDC04]. peripheral [HO76]. permutation [TAK09]. Perpendicular [ALSR11]. Person [HF01, ALK^{+}09, KT07, PY08a, RCJ^{+}13, VZP^{+}09, HFR06]. Personal [RCJ^{+}13, MFS^{+}07]. Personalized [CD10, CSJ13]. Persons [WN99, HPvB^{+}10, MW13, PA06]. Perspective [BR95, Che96, CM94b, GLD93, Gui99, Har80b, HMD93, HCLL89, iK87b, KWK94, LK94, MA84, MSH86, PR92a, SKK83, SA92, SB90, TB94a, CPT07, DWW^{+}12, Gol13, HN95, LH90, Pen89, SA90, UN91, YHR^{+}05, ZH04]. Perturbation [KMI79, OK07]. perturbations [MJBR88]. Perturbing [ZW03]. PG [SRK02a]. PG2004 [KCOTW06]. phantoms [AO03]. Phase [Bat84, Ekl79, FJJ91, JJ94, OH81, PS97, TM94, AS09, DCS05, IJDAB13, MKS^{+}08, WB11]. Phase-Based [TM94, FJJ91]. Phillips [Ano94f]. Photo [HMD93, WSCO^{+}12]. Photo-Mosaic [HMD93]. Photogrammetric [WD99]. Photographs [AMA79, NMI79, UA77, Che08, CHL05, Lap88]. photography [NFA04]. Photometric [APB10, KC95, KB91a, KP97, NG98b, OD01, Td92, GCFMT12, HASS10, HJ12, JC06, JMPG11, YA12]. Photometry [CJ82]. Photomotion [ZTS96]. Photos [FM84, TL78, IZKB12, PHY^{+}11]. Physical [Ano94f, Bri84, DF01, Hod95, Man86, RWV95, ZW93, DQ05]. Physically [CBC^{+}07, VFV93]. Physician [SBK^{+}99]. Physician-in-the-Loop [SBK^{+}99]. Physics [Ano93e, BY01, Bra97, MS97b, WR08, DLP13]. Physics-Based [Ano93e, BY01, Bra97, MS97b, WR08, DLP13]. physiology [PDS^{+}07]. pictogram [BRA^{+}10]. Pictorial [Gro84, JC93, KR98, SK88, Tan76, Tou80, GM90]. Picture [AK78, Alg83, Ali77, Bic98, Ble84, BW76, CY83a, CMVM86, CT88, Col81, Gar76, GLM78a, Gra78, Gud82, GLM78b, HP78, KDK78, KA94, KSW85, KY86, KD76, KD85, Lee76, LT81, Mai81, MS78, MK76, NR88b, OPR78, Per76, Ros76, Ros77, Ros79, Ros80b, Ros81, Ros82, Ros83, Ros84, Ros85, Ros86c, Ros87a, Sch76, Shi83, Sug78, SH77, TP75, YK80, LX88, Mar89, Moo77, Sha79a].

Pictures [Ano92b, BKA84, Cha79, FK83, Ham77, HE81, HS79, ITN84, JP78, KR85b, Kul77, Lat93, Lat97, LSBG92, MML87, NA85, Oka81, Pan78a, Pet85, RYN98, Sha78, Tan81a, iTF78, TF81, UG92, Wal88, van86, Abu89, DSS94, JJN76, Sha75]. Piecewise [BR93, BS96, BA96, CCF01, SRK02b, AK91, Bar07, BL08, Klu78, Kri92, PZV13, Wil78]. Piecewise-Linear [BS96]. Piecewise-Smooth [BA96]. Pile [YK86]. piles [TN08]. Pipelined [OTL96]. pitted [PK05]. PIV [ACG^{+}09]. Pixel [Che98, Tam84, Yam80, ACDB12, GBF12, GGO10, JLL13, LFL08, VMP03, XJK12]. pixel-labeling [JLL13]. pixel-level [LFL08]. Pixels [Leu92, MGPF08]. Pizlo [HM97, May97, Ver97]. Placement [MG95a, CYP^{+}10]. plan [ES06]. plan-specific [ES06]. Planar [ATN83, Ahu86, AD86, BCL96, BH99, Bra94, BHNR93, GBB98, iIK85, iK85, Maz87, MS96c, NG98a, Nai87, NR88a, PR92a, Rob85, ST96, SAG85, SLY89, SY11, SKB98, SW86, TS86, WV97, WD99, Wil81, ACAAC^{+}08, Bar07, Boi88, CW02, HY11, HT89, PZV13, TSR89, Wil78]. Plane [CCMW97, Con88, EKH01, LB98, Mil80a, MFV80, OW86, CKS^{+}05, For89, HS91, HN95, KK11, LHM06, Neg12, OK04, PMF90, Ram72, WJG02]. Plane-Sweep [OW86]. Planes [KSŽ96, MBK81, ZF94, Buz03, KPKPW90, KK11, Tri90]. Planetary [UZC97]. Plang [SK88]. Planned [IB01]. Planning [SKOS95, TTA94, TG95b, YT99, Gho90, ZKRH04]. plant [PBN^{+}09]. PlantGL [PBN^{+}09]. plants [FJP06]. platform [MZB^{+}10]. platforms [VAWW10]. Plausibility [CPC99]. playback [SBS04]. players [FLB06]. Plots [Fra79]. plotter [HO76]. plotting [Bot78]. POCS [AM06]. Point [AD86, BO91, CPC99, GSP02, GSK02, HRS02, HCLL89, Kal82, KG94, KWK94, Kle85, KSS92, LMR84, LK00, Mid79, OD97, PJ88, RKG03, Ros99b, Spe92, SBZ97, Tay00, TD83, TML00, TS01, WB01, ABD11, AGCA06, ATC^{+}13, BHSD^{+}13, CGAY13, CLK09, CDT11, CS04, CXY^{+}09, CK09, CR03, For89, GG09, HN82, HY11, HT89, JXCZ13, KPKPW90, Kim04, LH90, LZLP10, MLF^{+}12, MS09, NBPF11, PB11, PLL12, RAC^{+}13, SAS12, Wu02, YK08, ZN13]. Point-Based [LK00]. Point-Enhanced [GSP02]. Point-Set [TD83, SAS12]. Pointer [DRCF95, MM80]. Pointer-Based [DRCF95]. Points [DT96a, FT98, JK02, LJ87, MA84, MA83, OG98, OW84, PM97a, Shi99, SHD86, SLL01, Vel95, WZ97, WW93, ZL01, ZH79, AK91, CHMG12, Eva11, HS91, Kui08, LM95, LB10, Loh10, MPST08, ODD96, SS78, TY05, Tri90, UTB^{+}11]. Poisson [XZWB06]. Polar [MGMS01, RB82, Ül01, KORC10, Mas09, Sch06, TP05]. Polar/Spherical [Ül01]. polarimetric [ZZZP09]. polarisation [AH08]. Polarization [LL97a]. policies [OH05]. Polya [BG91]. Polygon [CD93, DBB83, IS02, JS87, KADS02, LR02, Lee83a, LL86, O’R82, Sur86, Van84, WL85, YLWY92, Gho90, GS90, GBW89]. Polygonal [AHD94, Ahu83, BS96, ET94, Goo92, HB98b, HPR90, II86, JH98, KD82, MKW94, MH98, Tau02b, WD84, WW93, DGZ12, HS91, KPMR91, LBM04, LVM04, Ram72]. Polygonization [GA00]. Polygons [BW98, BKW96, BM98, Cou81b, Fra83, Gho91, KK81, MB79, MSW96, OCON82, WD99, WW88, BR90, Hub12, Kle13, RSFdM04]. Polyhedra [Fra83, JB92, Kal82, SP97b, SB90, Sug78, KM03]. Polyhedral [FHMB84, Han88, KCD00, TC87, Kar89, VM06]. Polyhedron [Ish84, LMR84, Bha91, HT89, KIK89]. Polynomial [Bar84, CMPP99, DSdlH^{+}11, DCL^{+}08, KS95a, LKE98, Far82, KM94, Mar90, XWYY10]. Polynomial/Rational [LKE98]. Polynomials [KP97, BG91, KA12]. Pomerleau [Ano94g]. Pooling [ATC^{+}13, KYM13]. population [Ham05]. population-based [Ham05]. Pores [CEP84]. pork [CCR^{+}05]. porosimetry [RCVA11]. Portable [HT98]. Portraits [PS97]. Pose [AKC11, ACB98, AW98, BK01, CS10, CH99, CS00, FMV93, HDF12, Jos99, Jur99, KSS92, KH94, NB10, RY98, Sto87, ÁB13, AC09b, CDT11, CYNO11, DGC12, DLF06, EBN^{+}07, HF11, HH12, KZ05, KMN11, LST13, LY06, LSTF12, ODD96, PD11, PHK92, PDTE06, RP08, SO07, SHC^{+}12, SRHC13, TAK09, ZDF10, Ziv10, dP10]. pose-based [PD11]. pose-contour [PDTE06]. Pose-Estimation [ACB98]. Pose-insensitive [NB10]. Pose-invariant [AKC11]. pose-wise [AC09b]. poses [Kri92]. Position [KK88b, Kro86, Lam84, Nag86, YK86, IHTA90, PA13]. Positional [CPK99]. positioning [UKH88, YHS95]. Positions [CD92, Haa82, Uhr86]. Positive [Bat84, BB13]. post [CRH05]. post-production [CRH05]. Posteriori [Cou81b]. postproduction [MR05]. Postures [LC85b]. Potential [GESB95, GBW89, KS91c]. Potentials [RM02]. Power [QV98]. Powerful [BGS83]. pp [Ano92b, Ano92a]. Practical [Ano95h, BS88, SLN95, dLAH07]. practice [PBSG12, Ano93e]. practices [TCB^{+}08]. pre [KSS08]. pre-integrated [KSS08]. Preattentive [Tre85]. Precise [GCEC07, AS08b, HSSH89, dOSJVBS12]. Precision [Kro86, KSS08]. Precomputing [LL06]. preconditioners [KMT11]. predict [CCR^{+}05]. predictability [GGMV08]. Prediction [KK92, KK93, MML87, PA83, RWV95, TS01, MKA73, PSYZ13, QAB^{+}11, TDT12]. Prediction/Verification [MML87]. Predictive [SYF99]. Preface [Ano01o, DHK12, HM13]. Prefiltering [YBDC93]. Preimages [AK85].

Preprocessing [Ble84, LSVD85, Pag97, RY98, BYN^{+}04]. prescription [BHBF10]. presence [CXFS06, LF08, PA10a, YS06]. present [MSS90]. presentation [TD04]. Preservation [ASS97, KDK78, Kaw82, Kaw83, LH92, Ma94, Ma96, ML00, Loh10, SS06]. preserved [ZZC^{+}13]. Preserves [Ehr78, Lio91]. Preserving [CT88, CY83b, GL95, HFF93, NM79, Pra83b, RM02, RYN98, SP97d, SBZ97, Tsa85, VB98, AK91, BDHM09, CK09, Eva11, GYH13, Hu08, LLL13, MGPJ11, TKPR09, WHHB12, ZSCP08]. Presmoothing [HC13a]. Primal [eGZW07, Sel86, KTP08]. primal-dual [KTP08]. primitive [AFSW03, UKH88]. Primitives [AGHN94, BH86, FMV93, RL93b, Sel86, Vee97, Wal87]. Principal [Ehr78]. Principles [Ano93e, BB92, Ku84]. Printed [Alg83, ME98b, ME98a, Por00, RD77, SDR91]. Printer [Ham77]. prior [PW91, TMQM13, WSKH13, YZT^{+}13]. prioritized [LB06]. priority [BRSSAL11]. priors [CC11, LCH95]. Prism [PF87]. Probabilistic [CH96, Cre99, GGR01, HD09, HBA93, KD10, KHH^{+}12, KF86, LT97, MGK00, PBQ99, Tsa96, WC99, ZKC03, CDT11, FSV07, GRGB^{+}13, Har83a, HNB04, HS89, HW07, KMN11, Lea92, LHYK05, MR90b, PJW11, PLLL03, SM12, SYK96, TID07, ZG10, TC11]. Probabilities [GV78, Sbe00, YY84a, LPS^{+}11, MR90a, YY84b]. Probability [CH78, Mil80a, VS82, LH95, TC11, XP11]. probes [BFR13]. Problem [AS83, ABMT87, AM78b, Bra85, BSB87, CP80, Fri80, GO87, HN91, HCLL89, HW94, JN93, JT86, Jur99, KB95b, KB00, LKE00, Mer88, OWW85, AL11, Dre96, IKST05, Kar89, NESP10, Oli91]. Problems [Bat84, För87, GMG92, GMW83, HPB94, HHS^{+}01, HW94, Kas94, Sug88, BTCH05, Gho90, JLL13, KL11, KMT11, KBJ^{+}10, OEK08, VW80]. Procedural [PFV^{+}11, QY02]. Procedure [CMVM86, LP79, OG98, Wec78, JM09a, Ram72, SS78, SB89]. Procedures [CT93, PW86]. Process [IF99, Liu97, LD98, MRF96, ABD11, HPvB^{+}10, UK12a, RRR11]. Processes [CA97, Pie79, SB95, SA79, Ter83, Ull79, Mat89, Mil89, NFM08]. Processing [AHD94, Ano94g, AS93a, BRW88, BW76, BACL97, Bur81b, CKB96, CL83, CPK99, CT93, CDLD77, Dav93, De 88, DRCF95, DPB00, Eng83, Gar76, GMG92, Gra78, Gud82, HM84, HM78, Haw78, Haw82, Hey82, HP78, KDK78, KNJ84, KDRC98, KY86, KM84, KJRA96, KH83a, LCS84, Law83, LH99, MS96b, McC82, MS85, MK76, ND92, OS87, Per76, Ree79, Ree84b, RM98, Ros76, Ros77, Ros79, Ros80b, Ros81, Ros82, Ros83, Ros84, Ros85, Ros86c, Ros87a, Sob78, SS79, TST^{+}83, TP75, Tau02b, Thi92, Tre85, UZC97, WWW89b, Wal87, WF78, WC79, WBB85, Wil79, YK80, dM92, AM93, AC09a, BCDH10, Big90, Dem05, EW91, FFY^{+}04, Gar82, Ham05, Jea11, JM09b, KMT11, LC11, LEB07, LPV07, Mar93, Mat89, MPVF11, NL90, Reb89, Ros10a, Ros10b]. processing [Sah05, Sha79a, SDPO81, Ano95g]. Processor [GR87a, HJS89, Mar80, OTL96, THT^{+}98, MR89]. Processors [DW87, OR81, Ree80, ROH88]. Procrustes [CLCO13]. produce [KGC05]. Produced [Her80]. producing [HO76]. Product [Kri84]. production [CD10, CRH05]. products [PG13]. Professor [CV13]. Profile [Jok98]. Profiles [Gro82, HH77, LP77, Per81, CD13, Ste13]. program [IK89]. programmable [Gon09]. Programming [BEPW00, FC86, Hor77, LHB87, OTL96, WWW89a, CM92, ES81b, HQW^{+}12, LZLP10, MSI10]. Programs [SHG^{+}88, DM78, IH91]. Progress [CFS98, IF95]. Progression [GW93a, GW93b]. Progressive [AM01, FM84, JEK98, MGK00, RG10]. Projected [PA82, SS84a]. projecting [BHSD^{+}13]. Projection [Chu02, Gui99, Gul79, iK85, iK86, iK87b, LK94, Min79, OD97, SB90, TA88, YZZ^{+}10, YH83, ZT98, Bar06, DMW10, Gol05, Lan91, LZLP10, UN91, WY11]. Projections [Ano01s, Col77, DB76, HGv87, KWK94, KVW94, Kub84, Luc01, MKW94, Tas77, Wag76, BKR^{+}89, BA06, Gol13, HN95, Sal90, San90, SM90, SG82, TP05]. Projective [Ano95h, ACB98, CDH99, FAB97, FMV93, GHMQ97, Kan91a, LV96, NS91, RFC97, ZL01, BPHB91, OBH04, OH04, SY11, TH06]. projector [CR88]. Proof [HW83]. Propagation [CM99b, Egg98, KK83, LL92, Rag92, BCMCB09, CS07, FF09, HER81, PL08]. Properties [ASS97, BC85, Bur83, CV92, Dor84, GL95, KDK78, Kan91b, Kis96a, Kul79b, OD02, Pag92, Pan78b, Pet00, Ros99c, Rut81, SSP01a, Sha78, SPW96, Shn81a, WN87, WBR86, WBR88, ABD11, BY08, CW02, CKS^{+}05, Eva06, GFW13, JFS11, MVP06, RFLSA11, SDC04, TJ12]. Property [IS02, Kaw82, OD99, Ron86, SB98b, vvv88]. Proportional [AGW85]. prostate [SM13a]. prostatic [TRG^{+}13]. Prototype [Sch76]. prototypes [RAHT11]. Provably [BO05]. provide [RGA10]. Proximity [MHN84, JN09]. Pruning [SB98c]. Pseudo [AS88, YK95]. pseudo- [YK95]. Pseudo-Euclidean [AS88]. Pseudocolor [MM81]. Pseudohexagonal [HY94]. pseudomanifolds [VPAM12]. PSF [Cha91]. Psychological [CPC99]. Psychophysical [Sav87]. PTZ [WZ08]. Publisher [Ano02i, Ano03q, Ano06n, Ano13s]. Publishing [Gar76]. Pulmonary [WW97]. Pulse [SSJ86, TSK94, GFW13]. Pupil [HBF09, KA12, YWZ11]. Purpose [TB94b, Wal87]. purposes [CNC03].

Purposive [Alo92, CM94b, SG94]. Puzzle [RB82]. Pyramid [BMR91, Bur80, BKA84, Ede87, JM92, OP96, Tan76, WZWT99, BCL^{+}90, CWLJ13, CR90, BK03]. Pyramidal [HF93, SAA93, TCAC90]. Pyramids [Ant82, BT88, SZ96a, Shn81b, Tan81b, BBB96, GDIIHK11, MJBR88, Mee89]. Pythagorean [Far02]. Pythagorean-hodograph [Far02]. Python [PBN^{+}09]. Python-based [PBN^{+}09].

QMF [KC92]. Quad [HS79, LLXW13]. quadrangulations [HA03]. Quadrants [AGW86]. Quadratic [AEH79, BM97, FMV93, HL76, TC87, YKC^{+}86, BPB11, LZLP10, MG95b, OEK08, SRTBS91, YZZ^{+}10]. Quadric [CCF01, JBK04, Lev79, NFJ93, Sar83, WJG02]. Quadtree [CA84, Kas94, LMKG85, Maz87, Sam85, SC97b, SS87, Wal88, LS91]. Quadtrees [Abe84, Bau85, BRW88, CA86a, DRCF95, Pet85, Sam82a, Sam84, Sam85, Sch92, SS90a, Shn81a, Shn81b, Yau84, Dye82, LGJ82, MSS90, Sam80, SS91]. Quadtrees/Octrees [CA86a]. Qualitative [Alo92, Ano93d, BB91, Got08, NP92, TP92, Wei90, FMGA^{+}12]. qualities [ZB05]. Quality [Alg83, DT96b, WY11, HAGR91, KLL^{+}11, ZZC^{+}13]. quality-sensitive [KLL^{+}11]. quantification [LSCM03]. Quantitative [CY83b, Iiz87, SB98a, LFL08]. Quantization [SYF99, San78, SP97c, Shu97, CS07, DF91, JO11, JWG04]. Quantized [GV78, Kul83, Ros80a]. Quantizing [CH78]. Quantum [Wam85, GS12]. Quasi [HFF93, HW94, Iiz87, IE99, Kaw83, Por00]. Quasi-circular [HFF93]. Quasi-Gray [Iiz87]. Quasi-mechanical [HW94]. Quasi-Metric [Por00]. Quasi-Objects [IE99]. Quasi-Preservation [Kaw83]. Quaternion [Muk92, SF07]. quaternionic [DCFM07]. Quaternions [HB98b, WSV91, Gol11]. Queries [SBA13].

Querying [SL99].

R [Ano95g, AM94]. R. [CGL94]. R.G.B. [GB96]. racquet [LHJ^{+}09]. Radar [FY85, Lee81b, YTTT83, Zit88]. Radial [Ada93, Ano01s, ADRY94, GPP88, KD81, Luc01, PM97b, BSM10, FSF07, KBJ^{+}10, OBS06b, RB89, TM04, WR08]. radiance [RH06, RKH05]. Radii [Con88]. Radiographic [CH78]. Radiographs [De 83a, WF78, FLCdA06]. Radiological [PV97, OTO06]. Radiometric [AF81, KGFP10]. Radiopaque [GBR79]. Radiosity [Sbe00]. Radius [CT95, Lan87, TC95, Ber89, TC89]. Radon [TWS06, ZS11]. Raised [KSG84]. Random [DC86, HD97, IF99, Kas80, KCA81, KBZ96, KH83b, MCPB00, MFV80, MRF96, NR88a, NC93, NI82, PV13, SP97a, Sbe00, SA79, Ser80, Smi85, Tas77, WKP13, WD92, YY84a, ZSN96, Bar07, BD94b, CJL06, Gou91, HS80, WB11, YY84b]. randomization [RG10]. Randomized [CC01, XO93]. Range [AK96, AA93, BLP95, BJ86, BR12, BS00b, CFM02, CM95, DF02, EFF98, GMA83, GW01, GJP96, GB93, HBH10, HHI95, JB99, KC94, LF96, Lyn81, MY95, Mas02, Mur95, NL96, O’G88, OD02, PM89, PS00, RJ94, RF02, RFL02, SA91, SAA93, SA96, ST96, SK85, SF97, SJB02, SB00, YD94, YL94, ZW93, ASFP03, BGR89, FK09, GBF12, HF11, HH91, HSJS10, KM89a, Li92, LT90a, LSKK10, LS12, LS09, MSR07, Mas09, MB05, RSS07, Rub82, SY10, SKU^{+}09, SKSR08, TG11, TSR89, TS11, VD90, YAK^{+}08, YW07, ZG06]. range-sensing [ASFP03]. Rank [HSD85, Hey82, KY86, TR09, ZZ10]. ranked [WDB12]. Rao [CT95, Kan98, TC95]. Rapid [Sze93, Wat87, AC09a, YCH07]. rarity [WSZL13]. Raster [BN85, Cap84, Ced79a, Fra79, LL92, Pav79, Peu79, Bad77, DV82, GL82]. Raster-Mode [Peu79]. Rasterization [vvv88, BBB11]. Rasterizing [RSFdM04]. Rasters [Sam84]. rate [CD95, Sau91, TVC09]. Rates [SLCP85]. Ratio [BB83, CH85, Kan94a, NS91, ACDB12, YC05]. Rational [CG94, EKH01, KP11, LKE98, PLR96, SAG85, SS95b, WW95, BFRA12, CW02, HB91, WX91]. Rational-Ruled [SS95b]. rationale [Pec07]. Rationalizing [Bro78]. Ratios [LF98, ASCF13]. Ray [Bid91, GBR79, GM79, HC94, Mar82, Pag97, Pag99, Rot82, Wag76, You86, AS08b, KS91c]. raycasting [Kim13]. Rays [Fra83, KHB01]. RBFs [OBS05]. re [BCM13]. re-identification [BCM13]. reactive [TM07b]. Reading [KABP98]. Real [BIP00, BEPW00, CGH08, GR87a, Gon09, HT98, HMD93, LB98, LHHC98, MWTN04, MBDB88, MS85, MTAA11, OYTY98, PGGM04, SJ89, TGB00, UM05, Wam85, WWB84, ZZZY13, ZXK02, AM04, BCMCB09, BDS12, BGTG04, DLS^{+}09, FFM05, HZW^{+}10, DFP^{+}13, MZB^{+}10, MFS^{+}07, Nic95, PDA03, RSS07, RL13a, SM12, SS90b, SGH07, SIT07, TLT91a, WWLV11, YWZ11, ZJ05, Ziv10]. Real-Time [BIP00, BEPW00, GR87a, HT98, HMD93, LB98, LHHC98, MBDB88, OYTY98, TGB00, Wam85, ZXK02, CGH08, Gon09, MWTN04, MTAA11, UM05, ZZZY13, AM04, BCMCB09, BDS12, BGTG04, HZW^{+}10, MZB^{+}10, MFS^{+}07, PDA03, RL13a, SM12, SGH07, SIT07, WWLV11, YWZ11, ZJ05, Ziv10]. real-world [TLT91a]. Realistic [FM96, GL97, Thi92, YB01]. reality [CKM11, WS03]. Realization [Ree79, SS79]. realtime [ZZZL13]. Reasoning [GESB95, KN99, KK92, KK93, DFP^{+}13, MM90]. Received [Ano97g, Ano98d]. Receptive [BK03, CS82, LL12]. reckoning [Gre04]. Recoding [FK83]. Recognition [AHD98, Ano96g, Ano01q, AA93, AD86, BAM87, BH99, BM79, BJ86, BF87, Big97, BP94, BB95, BZ99, BSF02, BHNR93, CN87a, CF01, CG87, CGL98, CH87, CJ93, CTF^{+}98, CS98, CCS01, CEC^{+}80, CS00, CW00, DL97, DCTO97, DV98, DC00b, DT97, EK97, GESB95, GK95, HH77, HR99, Hod95, Jar77, JRH03, KH96, KABP98, KP96, KP00, KMGC84, LSVD85, LB00, LT90b, LTS93, Mai76, MFJ95, MLP97, MST85, MKK02, Mer81, MNSK98, MYLP98, MT00, Mur87, NSK^{+}97, NG98b, NMP97, NP92, NDC86, PLLL03, Pav86, PR92a, Pla96, PR92b, QV98, RD77, Ris89, RDR95, RW97, Roh94, Rub80, Rut82, Sab76, SN99, Shi99, SGB01, SHD86, SLL01, SB94, Sta95, SBxx, Sto87, SH77, Tou80, UKH88, VKP98, VV92b, Woj84, Woj87, YB99, YC98, YF80, YFZ98, ZXK02]. recognition [AAASC11, AT13, ALSR11, AC09a, AC09b, AKC11, ASCF13, BHBF10, BRA^{+}10, BKK11, BL04, BWL04, BRP04, BEGB13, BCF06, BH12, Buz03, CGU11, CMBP09, CGR13, CCFC13, CS04, CFB05, CKLP09, CSG^{+}03, CNC03, DT10, DPR92, Dor89, EKY08, EK12, FWH13, FG89, FBF08, FFY^{+}04, Far11, FLCdA06, FR11, Fla89, Fly92, FAB12, GJ10, GBL08, GM90, GZJ05, HHWP03, HOH^{+}07, HMF10, HNB04, Hu08, Hu11, IK89, IH91, ITNP12, JLD12, JLD13, JM09b, KRK11, KHA^{+}05, KDV12, KS04b, LRW08, LCSL07, LHYK05, LY06, LLC13, LT90a, LL12, LC88a, LL08, LYSS12, LLC12, LDC^{+}13, MS09, MYK03, MU11, MTVM04, MB11, NFM08, NN13, NFSD13, Nis96a, NDO09, PC05, PQML11, PPT06, PS05b, PTE12, PS12, RAHT11, RM03, RR06, RS03, RCJ^{+}13, SM12, SIK92, STV09, SHC^{+}12]. recognition [SSM06, SKVS13, SKM06, SSN03, SSS13, TG11, TFL^{+}09, VK91, WW80, WRKP05, WY07, WCZ^{+}07, WS08, WRB06, WRB11, YS09, YAK^{+}08, YK95, ZC89, ZKC03, BGPD09, TFL^{+}09]. Recognizable [ITN84]. Recognizing [DBBB03, IB01, Por00, VM01, CU10b, LLC13, PD11]. Reconfigurable [Fuj97, THT^{+}98, CL95]. Reconstruct [Lau97]. Reconstructing [BGT94, Cav87, Gol05, Ish84, KIK89, KS03, KVW94, KL13, LCC89, MKW94, Min79, OCVV04, RSPD12, SCS91, TRS06]. Reconstruction [ABE98, AK96, AEH79, BKR^{+}89, BM99, BCL96, BCZ93, BKW96, BL01, BW93, BSB87, CFM02, CPC99, CCH91, CD92, CCS01, CM94b, Col77, DG01, DC00a, DB76, EW87, Ede94, FW97, FRL^{+}98, FKW98, FK99, GM79, GW01, Gri84, GP85, Gui98, Gui99, Gul79, GJP96, HGv87, Hen98, HL76, JK02, JB89, KWK94, Kd88, KSd88, KSS00, Kub84, LDPD97, LSHT02, Mer88, MM92, Nai87, O’R94, OG98, OD97, PCJC98, PS83, PS00, RFC97, Sam85, SGHM00, Tan95, TB94b, Tas77, Tay00, Ter83, TM94, VB98, VBH97, Wag76, ZW97, ZM96, ZOMK00, ZK01, BI10, BR12, Boi88, BMM^{+}07, CGAY13, CLK09, CPP^{+}11, CC11, CC03, CCD11, DWB11, DQ04, FPC^{+}08, FB05, GRGB^{+}13, GSV05, GPC^{+}10, IZKB12, JRH03, dOSJVBS12, KK11, KNO^{+}09, LHM06, LM91, LB08, LY13]. reconstruction [MPST08, MWTN04, OC90, OBS06b, PW91, PCR^{+}04, Rem04, SY10, SCL13, SHK11, SMD^{+}08, SG82, SH08, SS11b, TH06, Tan11, UK12b, VBN11, WSCO^{+}12, WL89, WHHB12, XL88, YHR^{+}05, YW07, ZZZL13, ZN13, Ziv10]. Reconstructions [CDH99, THO94, HASS10]. Reconstructive [Ver81]. Recorder [Lee76]. Recording [Lee76]. Recover [FL96, GR05]. Recovering [ACAAC^{+}08, CG09, JJ94, LR02, Mur95, SB93, SP97b, WD96, WC99, WALL00]. Recovery [AD86, CP99, CJC01, DC98, HPB94, KG94, PCV94, RC97, SF97, SA02, SS84a, TI01, WL88, YFZ98, BF07, CYNO11, KLL^{+}11, KZ05, SKBS13]. Rectangle [AS83, HW94, VW80]. Rectangles [Ell81]. Rectangular [FSS84, Fog84, LL92, Med84, SS76, KZ05]. Rectification [FSB85, SSP01b, CCD11]. Rectilinear [EOW84, FMRV94]. rectilinearity [RŽ05, Ros08]. recursion [HQN05]. Recursive [Ano94g, BL94, CSC96, CWC94, DC98, FS84, KK79, Kle13, Lee86, LP79, MK79, MF77, OPR78, TMQM13, FKV^{+}11, JR09, LC88b, MDR91, NHSC09]. Recursive-Batch [CWC94]. Red [PS94]. Red/Blue [PS94]. Reduce [KD86, MR90a, MR90b]. Reduced [Che98, EC88]. Reducing [RMD08]. Reduction [BL98a, Ein83, KAES99, KS95a, PA00, Pel79, Rob85, WWL92, CP09, LLL13, RRR11, Sez90, ZW03]. Redundancy [CM99a, WHN08]. Reeb [HA03]. Reed [RP88]. Reference [UK12b, LLR10]. referencing [AWK04]. Refined [Lee81a]. Refinement [Sam82b, SSF94, BI11, CD11, LK03]. Refining [THO94, UM90]. Reflectance [LK97, OD99, OD01, PK05, SI96, SP97b, Td92, Td93, WN87, LMC09, YA12]. reflections [LF08, NNT11, SW13]. Reflectivity [Bas81]. Refraction [RC06]. Refresh [MM81, MM80]. Region [AEM98, BB88, BL00, CWH^{+}13, CC97, Cha81, Fre76, GCB92, Har80a, HL84, IP98, KS91a, Kan80a, KLL^{+}11, Lee86, Lem79a, Lem79b, MB94, Mil79b, Mil79a, MK79, Mon84, NMI79, OPR78, iOKS80, PM97a, PSWH84, PBG04, RW88, SYF99, Sam80, SL99, TL88, YK80, Zuc76a, AT89, BA89, CKK^{+}12, EyGS11, IJDAB13, MSS90, MMV06, MJ11, Mil09, MBMC11, PFGG09, SI03, SO07, SCvW11, KL10]. Region-Based [AEM98, CC97, MB94, PM97a, SYF99, KLL^{+}11, PBG04, Mil09, MBMC11, SI03]. region-labeling [EyGS11]. region-merging [SCvW11]. Regional [CD13, Sha79b]. Regions [AB88, AGW85, BRW88, BV99, GSP01, KŽ99, KM00, LM99a, Rob96b, SLY89, SM99, ABD11, BE11, CKM11, GS95, nLPR91, Lio91, PD05, SH09, TSR89, TN05]. Registering [BLP95, TS11]. Registration [Ano01r, BKLO87, CFM02, DF02, Dav97, EFF98, FDMA97, FAB97, GCB92, HLF^{+}97, Jok98, KD86, KPH02, LHS01, MY95, Mas02, Mer88, MA83, OD02, PMV00, PLH04, RC03, RF02, RFL02, SK02, SKSR08, TB99, TH86, VV02, VLR84, WB01, AS08b, ASFP03, BI10, BBF^{+}11, BT05, BvdHL^{+}13, Bur81a, CBD^{+}03, Che08, CHZ^{+}13, CFM^{+}13, CR03, GGMV08, GSST03, HVD^{+}89, HY11, IP91, JBWK11, KT07, LV11, Liu10, LS12, LPR^{+}03, MMA06, Mas09, MDdMG09, NESP10, NBDB04, PB11, PRR03, RKG03, RFS03, SCD11, Tan11, TA13, TMB12, TZY08, WR08, XOF05]. Regression [Rew84, CFM^{+}13, LY05]. Regular [BM98, KD76, LAS94, SRK02b, YK87, Gus07]. Regularity [Kis96a, TJ12]. Regularization [RM02, LEB07, SM13a]. Regularized [ZGK95, BvdHL^{+}13, YLA09]. Regularly [MSN82]. Reillumination [Wor05]. Reillumination-driven [Wor05]. Reincarnation [PB96]. Related [BN84, Bie87, GK98, GK77, HHS^{+}01, RBA94, Ros00a, Gho90]. Relation [PA98, Nac82]. Relational [COW98, CS00, Haa82, HH82, PLLL03]. Relations [AD93, BAM87, KK88a, FAB12]. Relationship [HNR88, MAN84b, Lea92]. Relationships [KW00, JSRS08]. Relative [Chu02, Fre76, SU01b, CUSZ07, KPE90]. Relaxation [DWX83, DW87, Fau81, Hen84, KF86, LL95, LHB87, NHR81, Pri86b, Rut82, SM94, Ull79, Yam79, Yam80, Har83a, HS89, LPZ08, OEK08]. relaxed [WS06]. Relevance [MBKB02, PBQ99, MW13, Pen03]. Relevant [JDP97]. Reliability [För87]. Reliable [CDT11, CMPP99, LRW08]. Reliably [De 93]. relief [BGR89, ZZLZ13, ZZZY13]. relighting [WLZW04]. remeshing [AdVDI05, Gus07, LM12, ZGLP12]. Removal [Mar82, NI82]. Removing [LH84, CYC10, LB05]. Rendering [BIP00, Bid92, CRH05, EK98, RCG^{+}09, BSMG05, BS04b, Fiu89, Fiu91b, GY05, KSS08, RL13b, WLW06, kWwZ13, XTLP04]. reparameterization [YZZ^{+}10]. Repeated [CCS01, GS06, PGGM04]. Replacing [TDMT85]. Replicated [ND92]. Replies [Har94a]. Reply [Åst97, Bec86, CM94b, Col97, HM97, May97, Ver97]. Reporting [PS94]. represent [GW90]. Representation [Ahu83, AAS85, Alb74, AGHN94, AK85, Ano94g, AT83, Bie87, BB95, BRdBS99, BV99, BH83b, CF01, CWH^{+}13, CM99a, CS89, CA84, CA86b, CM94b, DFP85, DFP89, DT97, Ede94, Fis94, For72, Fra83, GK98, HGB98, KCD00, KD76, KD96, Kri84, Kro86, LMKG85, nLPR91, LB04, LT90a, MY87a, Mok97, Mon84, PJ88, PA97, PF87, PR79, PR92b, RJ94, Rei96, Ros93b, SMR98, SAG84, SW83a, SSP01b, TB94a, VS82, Wal88, Wei88, WA87, YLWY92, ZT98, ZXK02, AQ09, AWK04, ATC^{+}13, Bar06, BSMK13, CPP^{+}11, CG04, DBF04, Dam08, DSS94, FPC^{+}08, FG89, GK04, HPR90, HNB04, KM03, LL13, LD95, PD11, RK11, Rub82, Sam80, STV09, SS06, SBM^{+}06, SSS13, SY11, SWS11, UCB13, VBS^{+}04, WRB11, ZT09, ZH04, BS05, LV03]. Representations [AN84, FPDK12, GK98, GJP96, HTEB11, KP00, LV96, NVWV97, Ros86a, Shn81b, Ül01, BKK11, CAF09, HS06, MPV13, NBPF11, OGH04, ZC89]. representative [GDIIHK11]. Represented [BRW88, HS79, Pet85, Pie79, RGC87, Sam82a, van86, BB11, Sam89]. Representing [FS95, JT80, MB79, NL96, Ram84, TAK09, MSS90, YS08]. Reproduction [Hor84, MV86, LMC09]. repulsion [RM03]. Requantization [WP88]. requirements [ES06].

Resampling [HY94, Tau02a]. Research [AR77, Bie85]. Reservoir [CEP84]. residential [ÜB05, ZN13]. resistant [RK11]. Resolution [Alg83, Ano92b, CMRS98, CTH84, CJC01, FCG01, FTW81, Han93, JH98, Lee91, MCPB99, OS87, PE09, PCJC98, RY95, Stu76, TTA94, UG92, Ver81, WZWT99, AM06, AKC11, CSS^{+}13a, CD10, FSV07, HSJS10, IP91, KZD^{+}11, LT05, LN10, NFSD13, SP06, ZH04]. Resolution-First [Han93]. resolutive [Pat13]. resolved [JC06]. Resonance [RSB93, RMFB02, CCR^{+}05, KZD^{+}11]. resource [MFG10]. resource-constrained [MFG10]. respect [BFR13, LGJ82]. respiration [ZCCD06]. Response [DB94, KM89b, SB91]. Rest [RM02]. Restoration [ACW96, BW93, Bou79, Bra85, BS88, BGS83, Cai88, Fri80, GMG92, KWK84, KSG84, KH86, LX88, Mai81, McD81b, MR92, OH81, RMR85, SK79, SS95a, SMB95, TLT91a, ZSN96, BD94b, CL90, HMA10, HAGR91, Lan91, LC88b, MWF07, SW94, SM90, TLT91b, TT91, ZZAA92, ZGK95]. restorations [LcTT91]. restoring [LTT91]. restricted [LWLS12]. Results [Bas81, BNG02, FH84b, HN91]. retargeting [ZDF10]. Retina [ST80]. Retina-Like [ST80]. retinal [NBDB04]. Retrieval [AS83, APV99, BS99, Car01, CFG06, CY83a, Doe98, GFS04, JEK98, KB98, KR98, MBKB02, MKK02, MK01, PBQ99, SLST99, SBK^{+}99, SPK^{+}02, Sup02, ÁB13, ABI^{+}04, CHC11, CWLJ13, DSY10, FLHK08, GSS12, GH08, GCPF08, HMC10, Hei04, HC13b, HGS08, ILRB04, JWG04, JN09, KHH^{+}12, LK03, LZWP03, LC09, MSG10, NHK08, Pen03, PA10b, PFGG09, PR03, PBG04, Pun03, SLS03, TLEF06, ZTH^{+}11, ZYXZ13]. Retrieving [LF08]. Retrospective [KW12]. Reverse [EFF98, SOJ^{+}95, SS11a]. Review [AC99, Ano93d, Ano93e, Ano94g, Ano94f, Ano94h, Ano94i, Ano95g, Ano95h, Ano97g, BL98a, MA85, Pav78, RD93, SA91, EBN^{+}07, KHA^{+}05, RN12, Ano98d]. Reviewer [Ano94a, Ano95a, Ano96a, Ano97h, Ano98f, Ano00f, Ano02j, Ano03w, Ano04w, Ano05v, Ano06t, Ano07m, Ano09s, Ano10q, Ano11q, Ano12r, Ano12s, Ano13t, Ano13u, Ano99g, Ano01p]. Reviewing [Jon97]. Revolution [YL94, JBK04]. reward [KS12]. Rewriting [JRV82, JRV83]. RFID [GLOC10]. RGB [GH03]. RGBN [PBM^{+}11]. RGC [AO03]. RGC-sm [AO03]. RHT [XO93]. Rib [De 83a, WF78]. Ribbon [MWL99]. Ribbon-Based [MWL99]. Ribbons [MWLA99, Pon90]. Ricci [CHZ^{+}13]. richness [EK12]. Ridges [Har83b]. Riemann [Lil97]. Riemannian [AAASC11, ZGLG12, ZRKZ^{+}11]. rig [HC13c, KD10]. rigging [CBC^{+}07]. right [AO03, HOPA91]. right-handed [HOPA91]. Rigid [BK83, HZ86, HH91, LHH97, LH88a, SS84a, YH83, BY12, CR03, GRB13, KSS92, LH90, LST13, NKPT13, NESP10, PRR03, PV06, Pen92, RKG03, SKH08, TMQM13]. rigidly [ACH^{+}13]. Ringed [HHS^{+}01, ZQ11]. Rings [Dor79, Kul79b]. RKLT [SYF99]. Road [BW11, Gui98, Gui99, Gui00, LJ87, ZC93, BRA^{+}10, FFY^{+}04, SJ12, VCBC88]. road-sign [BRA^{+}10]. Roads [FTW81]. roadway [MZB^{+}10]. Robot [Ano93d, Ano94g, Ano94h, CN87a, KK92, KK93, SLK86, Sug88, SIT07, GLOC10, LHM06, MFS^{+}07, MLH13, ST10, VK91]. Robot-vision [SIT07]. Robotic [BL98a, TTG94]. robots [ZKRH04]. Robust [AM04, Ano01s, BD94b, BA96, BGK98, CSY08, CMVM86, CTE95, CK09, CCYC12, CV94, DB03, DG01, Dre94, FM91, FR11, JBR08, KGC05, KK07, KB00, KH94, Lai00, LB00, Lin02, Luc01, MY95, MGK00, MK01, MFS^{+}07, MST00, Min94, MN94, NDBT95, PS03, PYS03, RNDA13, SMK02, SHC^{+}12, SK01, TB99, TZM98, TZ00, VSP06, VLR84, WLD99, WWWM12, Wil98, WWW95, XFSC13, YWZ11, YGH11, ZHM11, ZYXZ13, ZJ05, ACS03, BSM10, BI10, Cou13, FS03, GG09, GCFMT12, HBH10, HBH11, HDF12, KBJ^{+}10, LRW08, MPVF11, MCF10, PB11, SSL^{+}12, SS90b, WB12, ZXY^{+}12, BETV08]. Robustly [BFY00, TS11]. Robustness [MJBR88, MN06, RPG12]. ROC [BKD01, SJST07]. rock [TN08]. Rocks [CEP84, TN08]. ROI [BRSSAL11, TVLS08]. Role [AM94, CM94b, Hen98, TB94a, Ham05]. Ron [GW93a]. room [GPC^{+}10]. Root [McC82]. Rosenfeld [HM97, May97, Ver97]. rotating [TAK09, TM04]. Rotation [CK00, DS90, DH92, EA95, FH84b, iK87a, Mer81, MS85, Pun03, WEY06, BDVK10, Far02, ZZL13]. Rotation-Invariant [Mer81, Pun03]. rotation-minimizing [Far02]. rotational [Sub90]. Rotationally [BH83a, SK02, MPVF11]. Rotations [TQ97, Gol13, OK04, RFLSA11]. Rough [Woj87, SB13]. round [HER81]. round-off [HER81]. route [MSSS09]. Routines [KH83b]. Rubber [LS94]. Rule [BPW91, DY98, KW00, LL99, LN85, MSH86, She86, DK13]. Rule-Based [DY98, KW00, LN85, MSH86, BPW91]. Ruled [PLR96, SS95b, MP03]. Rules [BS00b, KCA81, SYK96, Wil89]. Run [Cou81b, Ell81, KLK88, LMKG85]. Run-Encoding [LMKG85]. Run-Length [KLK88]. Runlength [KU95]. Running [FS85]. runways [HCN90].

S [CHC11, JC81, JC81]. S-Cube [CHC11]. S.O.M. [BLT05]. saddle [Kui08]. Salience [RW95]. Saliency [BSF02, WSZL13]. Salient [CM99a, Sau93, SM99, CVP10]. Same [KL77, Kar89]. same-object [Kar89]. Sample [CM99a, Ber89, HBB^{+}12]. Sampled [Bur80, SWS11, PPT06]. Samples [HD97]. Sampling [HGv87, IF99, Tan95, BW11, Bar07, BO05, CCD11, Fiu91a, HT91, HMA10, KL11, OC90, WDB12]. SAR [BP95, HMEB07, Zit88]. SAR-Theory [HMEB07]. Satellite [AF81, LH84, MAM97, QAB^{+}11, SO07, ÜB05]. Satisfaction [BZ99]. Satisfy [vvv88, ES06]. Saturation [LH93]. SCA [CPOO09]. scaffold [CLK09]. scaffolds [CK11]. Scalable [KSC97, CFCP11, GB08, MĆK09, ŠRDC09]. Scale [Ano94i, BC88b, FT98, JC98, Kim97, KD86, Mer81, MS85, Nis98, PS95, RS88, SUO00, SA02, Taj83, TSK94, WP93a, XHJF12, AMMV99, BKK11, BDS12, BDL^{+}06, CGR13, CHC11, CPS10, Dou92b, DSH04, EL91, ELA91, FPDK12, GE08, GPY^{+}07, IZKB12, KL07, Kui08, LS08, LBNS09, MUS06, PLL12, Sah05, SSL^{+}12, Sta05, TRS06, WP93b, XSD12, YWZ11, YK95, ZTH^{+}11, ZUS06]. Scale-And [Mer81]. Scale-Based [SUO00, ZUS06]. Scale-Space [Ano94i, BC88b, TSK94, XHJF12, BDL^{+}06]. scale-spaces [GE08]. scale/irregular [VRKL13]. Scales [BL98b, MKY01, PBN^{+}09, WS90]. Scaling [Cou81b, O’R94, KMP05]. Scan [Cap84, Ced79a, CD93, ES81a, JB99, LC79, LL92, YYL96, NESP10]. Scanned [Ell81]. scanner [FK09, MDR91, ZG06]. scanners [SRML09]. Scanning [Han93, Ull81, BLT05, LCT09, SO07, WWLV11, YGH11]. Scans [SP81, TSK94, CPS10, NB10, SW04, SKSR08]. Scattered [Boo79b, GKR02, OG98, Sam82b, Vel95, Kim04, KS04a, MK05, OBS05, OBS06a]. Scatterplots [BC88a]. Scene [BB91, Bic98, CFM02, Che00, CTH84, CBB95, DC00b, GW01, HFKN97, Har80b, KW00, MNE00, MJS97, MA78, MMP09, Pot77, RKK^{+}00, Rub80, ST80, SK85, SB00, Ste01, TY05, TFB80, WH78, XL98, ZT98, Bar07, BC10, BCM06, CGU11, CSS^{+}13a, CLZZ13, CG04, DCH12, DN82, HL13, KK07, Lhu08, LS08, RM91, STV09, YT13, ZH04, XP11]. Scene-Based [Che00]. Scene-consistent [TY05]. Scenes [AT83, BM99, BFF97, CS89, CCS01, Dou81, FRL^{+}98, FT79, Hab85, HGB98, JJ83, RN93, RA77, SHJB^{+}83, SA02, Bar05, BP09, DWB11, GY05, HCN90, MMP09, Rub82, SCL13, TN07, WRKP05, YR06]. Schematic [Jar77]. Schematics [Ble84]. Scheme [KCM85, SYF99, WVL81, YW99, ŽA98, GW90, HMESI13, LLXW13, LDC^{+}13, LBNS09, NHK08, NBDB04, San77, WNH05, ZZ07]. Schemes [FS80, KOY86, PR79, San78, SKS97, BFRA12]. Schumaker [Ano95g]. Science [Åst97, Col97, PRW97a, PRW97b]. Scientific [Ano95h, Her72]. Scintigraphic [PH82]. Scissors [MB98]. Scope [Lee86]. Screen [RGC87]. Screening [Mey86]. Script [Ali77, Moo77]. Sculpting [FCG01]. Seal [FT84]. Seaming [HH98]. Search [AM01, BAM87, Gra78, Rub80, YT99, YLA09, KHH^{+}12, MU11, RSS07, ST10, SM13b, XST04, LEA^{+}10]. Searching [HP96, KAES99, MRF96, DR04]. Second [Ano94j, Ano95b, Nag83, PA98, RM02]. Second-Order [Nag83, PA98]. secret [CJL06]. Secrets [HBG13]. Section [CV13, SS95b]. Sectional [CP99, KSd88]. Sections [BCL96, Boo79b, CEP84, EU85, Far86, LCC89, Sam82b, Sor81, Yau84, Boi88, MG95b, NRJ11, Tan11]. security [CJL06]. See [Hoc87]. seed [GA91]. seedling [KM03]. Seeds [SU01a, ZC93, CUSZ07]. Seeing [RG10]. Segment [CCA92, GMW83, MN85, MNHO00, PS94, FS03, LK03, SC93, DGG08]. Segment-Based [MN85]. Segmentation [AEM98, Ant98, Ant82, ACH^{+}13, BEH^{+}81, BM98, BL00, Ble84, BRW85, BS00b, Ced79a, CMVM86, CP79, CP80, CL97a, CM97, CF92, CTH84, CEP84, Dan81b, DM82, DH00, DC86, DW87, DV98, DCS05, EM96, FUS^{+}98, GB93, HS85, HL84, HGR^{+}13, HY98, JMA79, JJ83, JW94, Jon99, Kan80a, KH98, KSI98, Koh81, KVdG^{+}97, KSC97, LN85, LM99b, LL97b, LJ91, MNE00, MY95, MS97b, MS00, MCPB99, ME98a, MB98, Nad84, NC93, NVWV97, OPR78, iOKS80, PZ92, Pel79, PP95, PSWH84, Pot77, Pra83b, PB99, RW88, RD93, RWWH00, RA77, RMFB02, Rut81, SAA93, SUO00, SU01b, SP97a, SMK02, SJ93a, Shl83, SA95, SC98, TK97, Udu82, US96, WWC82, WV97, WF02, Wec78, WHL84, WD92, WWJ13b, YHN11, YYL98, AS09, ABEN09, AHDM10, ASFP03, BCL^{+}90, Bar07]. segmentation [BP05, BvdHL^{+}13, BPB13, BP09, BF10, CMBV04, CFYU12, Cel90, CT10, CLC91, CUAT13, CMB^{+}12, CU10a, CU10b, CU11, Cre08, DR03, DBZ07, DUSL94, FAB12, GFL^{+}11, GBHS06, GCEC07, GB13, GBL08, GDR04, GPDR13, GW07, HDS08, HE82, HC13a, HBH10, HS89, IJDAB13, JLD13, JMPG11, KS02, KBN12, KS89, KK13, KCC89, KGU10, Lap88, LV11, LCP90, LPS^{+}11, ML13, MVP06, MBH^{+}12, MMK04, MO11, MGPP11, Mig12, Mil09, MBMC11, MB05, MSF^{+}12, Mor90, NRJ11, NHSC09, NN04, NFU02, PJW11, PCR^{+}04, QAB^{+}11, SCE04, SM06, SG11, Sha05, SF07, SMD^{+}08, SCvW11, TA13, TSR89, The83, TN08, TRG^{+}13, TC11, VMP03, WO10, WSSS13, WRB11, WW90, WS06, WSKH13, WWJ13a, XST04, XAB07, XYW11, YZT^{+}13, YWMS08, YGC13, YB89, YJA96, ZBLS13, ZSCP08, ZFG08, ZRL^{+}11]. segmentation [ZLS^{+}13, ZUS06, ZU09, dMFU10]. Segmentation-based [HGR^{+}13]. segmentations [CCTCR09, KSG^{+}13, LH95]. Segmented [LL98, Pla96, EHG^{+}10]. Segmenting [RN93]. Segments [AK85, Cre99, FMRV94, GBB98, Kim82, MAN84b, OW83, Pha86, TD83, WV97, WP93a, YY84a, ZF94, ŽA98, Alb74, dFCS93, For89, GS90, Mae90, YY84b]. Segregation [BSI87, JKM07]. Seidel [CRC97]. Selectable [DT96b]. selected [GHPW12, HKK08]. Selection [BL98b, BS00b, DM81, KIF85, LSPV04, Mey88, MH79, Oli94, PG94, Sch76, SM97, Wal87, WH84, Wes78, Wil98, Zuc85, ZI87, BPBS13, BEGB13, CYNO11, GBHS06, GFW13, HG11, JR09, KY06, LK03, PZX13, SO07, SB13, TG11, Wha91, ZRL^{+}11]. Selective [CHMG12, HH05, LF83, OH05, WRKP05, DL05, GZJ05, LDC^{+}13, MTG07]. Self [CXFS06, CG87, DWW^{+}12, DC01, LWLS12, FK09, GB13, KS04a, QC04, RSL10, TLEF06, TM04, ZDF10]. self-avoiding [GB13]. Self-Calibration [DC01, CXFS06, DWW^{+}12, LWLS12, FK09, QC04, RSL10, TM04]. Self-Organizing [CG87, KS04a, TLEF06]. Semantic [ABC^{+}03, GMW12, GLM78b, KSd88, ABI^{+}04, DCH12, GYTL09, GLM78a, ILRB04, IJDAB13, JN09, LYSS12, LSTARMB11, PSE^{+}11, SM12, VZP^{+}09, XST04, ZG10, ZTH^{+}11]. semantic-based [SM12]. Semantics [Kan80a, Kod76, Fiu89, Fiu91b, FYH11, HS05]. Semi [TLWT12, WHM^{+}09, DWB11, Gus07, KS12, NN13, Pen89, Tri90]. semi-analytic [Tri90]. semi-interactive [DWB11]. semi-local [Pen89]. semi-regular [Gus07]. Semi-supervised [TLWT12, WHM^{+}09]. semi-transparent [KS12]. semiautomatic [SDPO81]. seminar [GHPW12]. sense [CWO^{+}11]. Sensing [Ano93d, ASFP03, GZJ05, LSKK10, OH05, SB96a]. Sensitive [RW76, RW79, KLL^{+}11]. Sensitivity [KH94, LFMP13, LP10]. Sensor [CH85, MG95a, TTA94, TTG94, TG95b, WW94, YT99, AZSVK05, CA10, LSKK10, SIK92, TDWH07, TMB12, YHS95]. sensored [CD10]. sensorial [CCR^{+}05]. sensors [IK89, IKST05]. Sensory [BH86, OGH04]. Sensory-Motor [BH86]. Separable [Gad91, WBB85]. Separating [BW98]. Separation [Sav87, AS09, ZZZP09]. Sequence [CA97, LK91, LZ97b, NDN^{+}97, Pot87, WALL00, XS98, FR11, GS06, JM09b, LD90, NSEA13, PGGM04, Rem04, WZL^{+}03, ZZZ06]. Sequences [ALK99, CWC94, CW00, Enk88, FRL^{+}98, Fog91, GMW12, GHS95, HNR84, IP98, KSS97, Law83, MB94, Nag83, Pie79, PM97a, PF01, RWWH00, Roh94, SJ93a, SF95, SPS81, SJ89, SLCP85, SBZ97, Sze93, TPR^{+}00, WN99, WLD99, WN86, Yac83, ZW97, ACH^{+}13, BF07, CXFS06, CSG^{+}03, DCS05, DN82, DHP08, HJ12, HH91, KH90, LSC08, LS08, LWH03, MC09b, NT10, Neg12, PS03, RG12, RM03, SH05, TY05, TVC09]. Sequential [AK77, BSF02, Cap84, FAB12, GM85, HW06, JRV82, JRV83, PLS97, RW76, RW79, WD84, ES81b, LC88b, MD82, SYK96, SAC09, SHS03, WS08, UPBS08]. Serial [LCC89, Sor81, TV99, Tan11]. Series [MRW^{+}97, LEA^{+}10]. service [MFS^{+}07]. Set [ACF00, BG80, Bau85, Bic98, BP94, BDL92a, GAD01, HD97, LLSV00, ND97, Nur86, OW83, SS90a, Tan89, TS00b, TD83, ZOMK00, ACS03, BG79, BDL92b, CDT11, CBT^{+}04, CU11, DM12, FPC^{+}08, HS91, KK13, MMV06, PB11, PD05, PW91, SAS12, SG11, SRS11, TT91]. set-theoretic [TT91]. Sets [DL97, GW01, KSKB95, KB95b, LER95, NG98a, OW84, Peu83, Ron86, SR00, Sch78, Ser80, Shi99, WB97, WB01, AGCA06, BFR13, CGAY13, Cre08, DCS05, GK03, HY11, KPKPW90, SM06, Sha11, WSCO^{+}12, dCCP12]. Setting [KTP08]. Seven [SOD10]. Seventh [Ano96c]. Several [FT79, KPKPW90, LCP90]. SFM [CX11]. Shaded [RGC87]. Shading [BHMB10, GR85, Gri84, GM87, HB86, HLKF95, KB91a, KP97, KB95a, KB95b, LK94, LK97, MR96, OD97, SKB96, Bru88, DFS08, FL92, KN03, KK90, Oli91, Pen89, Sze91, Wor05]. Shading-Off [MR96]. Shadow [JW94, CYC10, SCE04, WCF10, YZ06]. Shadows [Max84, Max91, SK83a, Thi92, CF07, JF10, LP90b]. Shafer [Ano93e]. Shape [AK77, ASZ99b, BCL96, BH99, Boi88, BCG95, Boo97, BHNR93, COW98, Car01, Cav87, CPC99, CCP97, CP86, CTF^{+}98, CFA98, CJ82, CCD11, DT10, Dan78b, DM01, DC98, DY98, DT97, EK97, FW97, FK99, GP06, GM87, HB88, HF01, Hob00, HB86, HFF93, HLKF95, JC98, JC90, JEK98, JMPG11, KB91a, KP97, KB95a, KB95b, KR98, KMGC84, LPC08, LL99, LF83, LK94, LK97, LW85, LSBG92, LYG07, LYCG08, LK00, MWL86, MST85, Mas02, Mil89, MA83, Mok97, MPPG98, MSH86, Mur87, NSK^{+}97, NMI79, NR88a, Nis96a, Nis99, OD97, Oli91, OBH04, OH04, PJ88, Pav78, PEFM98, PV97, Ros86a, Rut79, Rut81, SKB96, SP97b, SKB98, Td92, TI01, TSP97, TFL^{+}09, TZY08, Udu81, VY94, VTG95, Wah83, Web83, Wec79, Wei92, Wei88, Wyv03, YF80, YJ84, YFZ98, ZOMK00]. shape [AAASC11, BSMG05, BF07, BvdHL^{+}13, BY12, BGK95, Bru88, BF10, CH06a, CK11, CYW04a, CYW04b, CC11, CUAT13, CL08, CLCO13, CGL92, Coe12, CTCG95, DZL07, DFS08, EL07, FPC^{+}08, FG89, Gho88, Goh08, GPDR13, HFR06, HS05, HG11, HC13c, IH91, KŽ12, KK90, KNO^{+}09, LE09, LPS^{+}11, LPZ08, Liu10, MDFS11a, MC09b, MWTN04, MD82, NHK08, Pen89, PBG04, PW91, PS12, RK11, RAHT11, Rem04, SBM^{+}06, SM13a, SY11, SH08, SWS11, SKBS13, Sze91, TG11, TWS06, TMQM13, TÉSK11, TH04, TC11, WB12, WSKH13, Wor05, WWJ13b, XZWB06, XWYY10, YB07, YZT^{+}13, YLA09, ZZC^{+}13, Zit88, NLW13]. Shape-based [JMPG11, FG89]. shape-constrained [WWJ13b]. Shape-from-recognition [TFL^{+}09]. Shape-From-Shading [GM87, DFS08]. shape-interrogation [HS05]. Shape-Preserving [HFF93]. Shape-Specific [MA83]. shape-texture [HG11]. Shaped [DFP85, GSP01, Gro82, Woj84, TA13, VK91]. shaped-based [TA13]. Shapes [ANM98, BRdBS99, Cho79, CGP85, Gho91, Han88, KH98, KS96, NWP97, PR92a, Pla96, RH85, Rut82, SC99a, SSN78, ST96, Sup02, Thi92, Woj87, YL94, AC07, CKK^{+}12, GR05, GW90, HOPA91, HW06, IAP^{+}11, KP12, LL13, LT90a, LBNS09, MPV13, Sha05, YYF89]. Shared [ASZ99a]. sharing [HAKK91]. Sharp [BMZB02, WHHB12]. Sharpening [CT88, MY87b, Mor90]. Sharpness [FH84a]. Shear [CK00]. Shears [TQ97, WEY06]. Sheeting [LS94]. sheetmetal [ZZZ06]. SHGC [SB93]. Shielding [KK88a]. Shift [SK79, ZYS09, ZLS^{+}13, LLR10]. Shifted [Ano92b, UG92]. ship [IHTA90]. Shock [SKS97]. Shock-Capturing [SKS97]. shockwaves [SGTL09]. shorelines [BKP10]. short [PS03]. Shortcuts [Wei92]. Shortest [DJG01]. Shot [Che00, YW99, SOD10]. Should [Hoc87]. Shrinking [KPS76, GH90]. Sides [DBB83]. SIFT [DIOV06, LS09, XHJF12, ZYS09]. SIFT-like [XHJF12]. Sigma [Lee83b]. Sign [CW00, OD99, VM01, BRA^{+}10, FFY^{+}04, WCZ^{+}07, YS09]. Signal [Chi97, HNR88, HNRR90, JN93, Kan80a, KWK84, EW91, Jea11]. Signal-Dependent [KWK84]. Signals [BR93, BL94, RB92, Pey09]. Signature [DLHT99, MKK02]. Signatures [BHNR93, Hob00, SC00b, PG13]. Signed [Mas02, Gre04]. Silhouette [AAASC11, BL01, ES04, LHS01, LPC08, LYG07, TCMS04]. Silhouette-Based [LHS01, AAASC11, TCMS04]. Silhouettes [CA86a, HCHD01, Lau97, Pot87, DT09, LYCG08, MBH^{+}12, SY10, SA90, YW07]. SIMD [IKS86, LP90a, MHSP10, Ree82, ROH88, TV99]. SIMD-based [MHSP10]. Similar [Iiz87, MHMO09]. similarities [PG13]. Similarity [BJ97, Car01, CY83a, Hen98, KAES99, STLH08, SK84, TP05, VLR84, YK08, BB13, BAP08, CK11, DL05, FLHK08, GCPF08, Got08, HBL^{+}11, MGW10, NHK08, RKG03, TH04]. similarity-based [NHK08]. Simple [ASS97, ASZ99a, BSB87, CD93, HM84, JB92, KK88b, KIF85, Lee83a, LR90, Ree84a, Sur86, BPG05, Eva11, GK03, GH03, HE82, KM89b, KA12, Loh10, TC89]. Simplicity [LM96]. Simplification [BVL02, KKK99, KSS00, Zhu89, CD11, DGZ12, VS08, WR05]. Simplified [ZHAH88, BC10]. Simplifies [Dan97, ZU09]. Simplifying [AM97, SdB03]. Simply [Her98]. Simulated [BCG95, Wat87]. Simulating [HH05]. Simulation [Gou84, LB97, LWGP08, Thi92, VCVQ^{+}98, Wag76, ZYP09, CD95, IHTA90, KMBG09, PDA03, SGTL09]. simulation-and-matching [IHTA90]. simulations [HMEB07, WBOL07]. Simultaneous [DC98, EFF98, Jok98, JC06, Jur99, KPKPW90, LM99b, PA06, Td92, THN92, TRG^{+}13, VM01, WB01, CHH09]. Single [BK01, BL92, Cha81, CC11, CCS95, Dre94, GLD93, Gui98, HR99, LA11, LC85b, LN98, Sug88, Tay00, YL94, CG09, CH06a, DMW10, HJ12, HQW^{+}12, KSR^{+}12, KK95, KTP08, KS12, KM03, LVM04, MDdMG09, Pen89, RKH05, XYW^{+}08, ZZ07]. single-data [KK95]. single-direction [HQW^{+}12]. Single-Pass [Cha81, CCS95]. Single-View [YL94, HJ12, KM03]. singular [SCCP05, TJ12]. Sinusoidal [GLR^{+}99]. SIT [Wam85]. Site [CJC^{+}98]. sites [AO04]. six [Sha11]. Size [Sch92, CFG06, MGW10]. Sizes [Shi99]. sizing [TN08]. Skeletal [MPV13, Sug78, AGCA06, TH04]. Skeleton [ABE98, BM95, BA92, Bra94, CYW04a, CYW04b, JXCZ13, SS06, SPS81, Sin87, TF84b, Boo79c, HXS09, JXC^{+}13, LKC94, San90, TF84a, WWWM12]. Skeleton-based [JXCZ13]. Skeleton-driven [CYW04a, CYW04b]. Skeletonization [HP84, KSKB95, Pud98, TTF04]. Skeletons [AM97, Che98, NSK^{+}97, NGC92, TSP97, BGLSS04, Cou13, Goh08, Sha05, SdB03]. Sketch [FML12, PBM^{+}11, Sel86, ZDL^{+}11, eGZW07, HC13b]. Sketch-based [FML12, PBM^{+}11]. Sketching [Haa82]. Skew [Spi98, Yan93a]. Skewed [Fri86, VMUO95, Pon90]. skin [CGW^{+}07, SJST07, XYW^{+}08]. Skinning [HB91, She03]. skull [YLL12]. Sky [Max91]. SLAM [KD10, SE11]. Slice [RGC87, MDdMG09, YG07]. Slice-Represented [RGC87]. slice-wise [YG07]. Slices [BS96, O’R94]. Sliding [de 83b, Cho88]. Slope [BF87]. Sloped [Coh85]. slow [yKL11]. slow-in [yKL11]. slow-out [yKL11]. sm [AO03]. Small [Chi97, FH84b, FT98, MS94, CDT11, KPE90]. Small-Rotation [FH84b]. Smart [BKMV07, CVP10, GPC^{+}10, MCT10, MHSP10, WMBY12, Ziv10]. smart-room [GPC^{+}10]. Smear [HL84]. Smears [WHL84]. SMI [Wyv03, PS05a, SCOG09]. Smooth [BU93, BA96, NWP97, SS11a, BL08, GR05, Kri92, Pen89, SA81, UK12a]. Smoothed [KNJ84]. smoother [LV11]. Smoothing [BB87, CBM01, CY83b, Ehr78, JC98, Lee81b, Lee83b, LL95, LBSP02, Mas85, MPC94, NM79, Pan78a, Pra83b, WVL81, BI11, GYH13, GS08, HS89, NKP11, WY11]. Smoothly [Han88]. Smoothness [CM94a, SM94]. Snake [Pet99, WWJ13b]. Snakes [RAH97, Sap97, SZ07]. snooker [DRK03]. soccer [ABC^{+}03, DLS^{+}09, FLB06, MSSS09, ROJX09, VMP03]. Soft [PS95, YLM11]. Softassign [SAS12]. Software [KH83a, TST^{+}83, BLT05]. solar [CF07, JF10]. Solid [EPB05, NH92, SSP01a, San78, SS06, ZDL^{+}11, ZLH13, ZGLP12]. Solids [HP84, PCR86, RAH97, Rot82, YL94, Kar89, San77]. Solution [CP99, HCLL89, HA93, Jur99, BLT05, CRT90, DK13, Dre96, Gho90]. Solutions [Bat84, KMI79, LKE00, OD01, SW86, KT08, KBJ^{+}10, LPR^{+}03]. solver [AL11]. solvers [KMT11]. Solving [FH84b, HW94, KB95b]. Some [CY83b, Dor84, EOS84, FR80, GK98, GMW83, HPB94, HHS^{+}01, HN91, KDK78, Kol83, O’L88, PD79, Ros86b, Sha79b, Sug88, Tho86, WWL92, Yam80, Big90, TT91]. Sonar [Ano93d, MCPB99, MCPB00, TS00a, TPR^{+}00, Neg12]. Sonka [Loh10]. Sort [LK03]. Sort-Merge [LK03]. Sorting [WSCO^{+}12]. Sound [Lee76]. Source [BLd95, CJ82, HF93, Man86, Min79, OD97, OD01, SHG^{+}88, VY94, CF07, Dre96, RAC^{+}13, TMNM09, YHS95]. Sources [GM79, LZ97a, LF08, WS03]. Space [Åst97, Ano94i, Bar85, BC88b, BL98a, BB83, Col97, DB79, FT98, Gau92, GR85, HR99, Hob97, HGB98, JC98, KKO98, Kim97, Kle85, Kod76, KC87, LI00, LL97a, LJ89, Man84a, Mit88, Mok97, Nur86, OW86, Pet99, PRW97a, PRW97b, PA82, RC97, RS88, SC00a, SO01, SCS99, SKK83, SS79, TSK94, VV92b, WW88, Yan93b, ZL01, AQ09, BTCH05, BT05, BDL^{+}06, CAF09, CHC11, Dye82, FS03, GPY^{+}07, HKK08, JSRS08, KH13, Kui08, LH95, LL08, LN10, OC90, SHC^{+}12, TH06, Thü03, VMP03, WMBY12, XHJF12]. Space-Economical [OW86]. Space-Efficient [Hob97]. Space-Encoded [PA82]. Space-Optimal [WW88]. Space-Variant [BL98a, RC97]. spaceborne [HMEB07]. spaced [DN91, Klu78, TN05]. Spaces [Han93, Her93, Her98, CS07, EL03, Eva06, Eva11, GE08, QT10, dSdSF^{+}12, dLAH07]. SPAMM [RAH97]. Spanning [SS84b]. Sparse [CWH^{+}13, KP00, MM92, MS10, OBS06b, PCV94, WR93, WR96, BR12, CC11, CS07, FB12, LHM06, LY13, Pat13]. sparsely [PPT06]. Spatial [AHZ96, BSI87, BC88b, BL98b, BH83b, CGL98, CA97, Chi81, Dav97, Dav79, DCFM07, GS92, Hab85, Kal82, KW00, LW85, Lyn81, MHN84, PA00, Peu83, Pha01, SO01, WF02, WWB84, ZD01, CSY08, CCTCR09, CHC11, FMGA^{+}12, Far02, FAB12, Fra89, Gho90, Hei04, HT89, HGS08, KM89b, KY06, MPF07, PSE^{+}11, TP05, TCC90, WSSS13, WDB12, YSD03, ZTH^{+}11]. spatial-domain [TP05]. Spatial-Feature [WF02]. spatial-scale [CHC11]. Spatially [Lai00, SB96a]. Spatio [LJ89, NDO09, PD83, Pet99, Yac83, CHMG12, CWLJ13, DLF06, LCSL07, RL13a, SA04, WXRA07, XYW11]. Spatio-Temporal [LJ89, PD83, Yac83, NDO09, CHMG12, CWLJ13, DLF06, LCSL07, RL13a, WXRA07, XYW11]. Spatio-Velocity [Pet99, SA04]. Spatiotemporal [AD93, DIMT12, JW87, TI01, BZS08, JYTK11]. Spatiotemporal-Frequency [JW87]. Special [Agg83, Ano01h, Ano01q, Ano01r, Ano05p, Ano06o, BPS10, BSW01, CFS98, CA10, CKB10, CV13, CNK01, DRDKE13, FHP01, FPDK12, FYH11, GP06, GHMT09, GHPW12, Gro82, GO87, HB05, HMC10, HTEB11, HGSM11, HCS03, JWDF05, Jon08, KB98, KS00, KCOTW06, KSM^{+}06, KPKH07, KLBP11, LBK10, MPF07, MYK03, NLW13, RFL02, STV09, SCOG09, SST06, SRK02a, Tau02b, THL13, Tho10, WP00, Wyv03, HM13, LLE^{+}09, PS05a, SMHH04]. species [CTM^{+}13]. Specific [Bas81, BF87, DC00b, MA83, ES06]. Specific-Line [BF87]. Specification [BD94a, LD98, RW76, RW79, SK88]. Specified [Ram76, GS95]. specimen [MSG10]. Specimens [Mey86, KORC10]. Speckle [Lee81b, BP95]. Spectra [SB98b, DvLV08]. Spectral [BL04, Chi81, BEGB13, CHP^{+}11, CPT07, DCFM07, GCEC07, OEK08, PTE12, TCC90, YSD03, ZRL^{+}11, ZZZP09]. Spectral-Spatial [Chi81]. Spectrometry [SGK00]. Spectrum [HKD95, Mai76, HD07]. Specular [CTE95, CJ82, CKS^{+}05, LF08, SIK92, ZMCA05]. Specularities [LKK00, LB05, OJRT08]. Specularity [Dre94, HB88, LL97a]. speech [PY08a]. Speechreading [LT97]. Speed [AHRW87, DT96b, EA95, LI00, SSF94, THT^{+}98, Cha74]. Speed-Up [THT^{+}98]. Speeded [BETV08]. Speeded-Up [BETV08]. Speeding [For88]. speeds [NFU02]. Spetsakis [Zha97b]. Spetsakis-Aloimonos [Zha97b]. Sphere [KKO98, Lil97, BBHF10, Fra81, SW13, TMNM09]. Spheres [Ada93, LP10, ZQ11]. Spherical [BH95, KHK10, Ül01, YH83, BI10, Beu91, CHZ^{+}13, CPS10, RDM^{+}11, WLZW04]. Spider [TST^{+}83]. Spiral [BE11, WC79, Big90]. Splat [CGAY13]. Splat-based [CGAY13]. Splatting [LI00]. Spline [ABMT87, BG80, Bar84, BLH91, CL00b, FSSL86, GW93a, HJK02, KM00, PW86, Rab92, RFS03, TC87, YKC^{+}86, BG79, CXY^{+}09, CL91, GW93b, GCB90, HB91, SDPO81, XWYY10, YZZ^{+}10, Rab92]. Spline-based [RFS03]. Splines [CLR80, FSS94, GSS00, HFF93, MY87a, Olk95, Pha89, SB84, WLH85, DCL^{+}08, GHQ06, HN82, KP11, KP12, ZK05, ZQ11]. Split [CMVM86, CP79, CLC91, DR03, Lap88]. Split-And-Merge [CP79, CMVM86, CLC91, DR03, Lap88]. Splitting [Lee86, Lem79a, OPR78, SOG09, HZLM11]. SPM [KSM^{+}06]. sport [MP09a]. sports [KPPK09, LHJ^{+}09, LWH03]. Spot [BT88]. SpotIt [BM96]. Spots [NS98]. Spotting [ZXK02]. Spread [Ree92]. spring [NÇ10]. Square [Dor84, KON87, WS91, Dou92a, Dou92b, ZZ10]. Square- [WS91]. Squares [FM99, Jos94, KSŽ96, PW86, YY84a, Ber89, GSV05, HK93, MP09b, NA90, YY84b]. squares-based [MP09b]. Sridhar [Ano94h]. Stabbing [GS90]. Stability [BC88b, FT98, QV98, Ric84, LM89]. Stabilization [CC00]. Stabilizing [FF09]. Stable [YL08, Kri92]. Stack [YBDC93]. Stage [CSDC96, SP97d, WLMG08]. Stained [HGA86]. Staircase [SSJ86]. stance [NF06]. standard [And03, Buz03, KMBH09]. standardization [ZU09]. State [SS79, AO03, BGTG04, JM09b, KTP08, LN10, Ros10a, SCD11]. state-of-the-art [JM09b, SCD11]. State-Space [SS79, LN10]. states [FR11]. Static [RA77, WY07, Bar05, CSG^{+}03, GY05, Rem04]. stationary [RSPD12]. Stationing [EOW84]. Statistic [KIF85]. Statistical [Bai88, CH80, Fri80, HT88, HNR88, Kan94b, Kan94c, KSG^{+}13, LK00, MKW94, ML78, Pan78b, PD79, SM13a, Spe94, WC92, de 83b, BvdHL^{+}13, BF10, CLC91, GPDR13, HKK08, JTEA91, KGC05, KY06, Nis96a, TT91, WS06]. Statistics [FSA01, Lee81a, PA98, BMR91, KS89, TLEF06]. steerable [AS08a]. Steered [FUS^{+}98]. steganographic [YCL07]. Step [SC92, WR87, BYN^{+}04, KM89b, TCC90]. step-wise [TCC90]. Stepping [Boo79a]. Stepping-Stone [Boo79a]. Steps [Ano92a, KT89]. Stereo [AM01, BM99, BH83b, CN95, CL86, CHRM96, DC00a, EW87, GY88, HQW^{+}12, HH98, JT86, JB91, KC95, KS95b, KB91a, KP97, LK91, LL97a, LSHT02, LHB87, MS97a, MN85, Mur95, Nev76, OD01, PW06, Td92, WZ08, YC78b, AK10, AK11, APB10, BCMCB09, BBC^{+}07, CPP^{+}11, CC07, DBZ07, ES04, FB05, GBF12, HASS10, HBG13, HZW^{+}10, HKA13, HH82, JMPG11, KPE90, KN03, KGFP10, KT07, LS08, MSI10, MCT10, NT10, SE11, Tri90, Wei90, YC78a, YA12, YK08, ZN08, ZKRH04]. stereo-based [MCT10, SE11]. Stereo-Motion [DC00a]. stereocomparison [MKA73]. Stereokinetic [CPD93]. stereophotogrammetry [BNL90]. Stereoscopic [Jon97]. stereotactic [MDdMG09]. stereovision [PCC13]. Steven [Ano93e]. Stick [WLW06]. still [PL10]. Stimulating [Ano94f]. Stimulus [Smi85]. stippling [SLKL11]. Stirling [Ros98b]. Stitching [LHS01]. Stochastic [ADDK99, CEC^{+}80, Fu80, LRLB11, LF79, Mee89, Mil80a, MFV80, PB11, SGHM00, SSS82, TF84a, TF84b, VB98, WZWT99, Yan93b, AO03, KK13, KL11]. Stokes [CD95]. stomach [KS91c]. stomachs [KS91c]. Stone [Boo79a]. stopping [SYK96]. Storage [FSB85, KMGC84, Peu83]. Straight [AM78a, Bid91, Cag93, Gaa77, GL97, Kim82, LH88a, LH88b, MM88, Pha86, RW76, RW79, SBT85, VS82, WV97, WP93a, Wil81, Alb74, BGLSS04, LF82, Mae90, MFA89, Sch06, Sha06, SC93, ZS11]. Straight-Line [WV97, Wil81, Alb74]. Straightness [Ano92a, Kis96b, MMS97, KT89]. Strategies [AA93, CJ93, CEP84, Goh08, LVW97, SLK86, CUAT13, KTP08, KYM13, YLA09]. Strategy [BM99, LN85, MST85, YB95, Bar07, GCPF08, MFB11]. Streak [Mar82]. Streams [DH00, OYTY98, GGO10, KLV06]. street [DN82, ÜB05]. Strength [SU01a]. Striate [CS82]. strict [ZK08]. String [CTF^{+}98, JC81, LAS94, ZNG^{+}13]. Strings [HY98]. Strip [BW98, LY90]. striped [TLGS05]. Stripes [Mon84]. Striping [AF81]. Strips [RCG^{+}09]. strokes [Pha91]. Strong [Ron86]. Structural [Bai88, CH80, HBA93, KLL84, MLP97, MSN82, MMN83, MAN84b, Nis95, Nis96b, Nis97, Nis99, Oka88, SA85, iTF78, Wec79, WCH98, BEGB13, MJBR88, Nis96a, TFB80, ZG10]. Structural/Statistical [Bai88]. Structure [AS83, Abe84, BS05, BKA84, CJC01, CWC94, DT96b, Dav79, GS92, HA93, Jac01, JJ94, KMB97, iK86, iK87b, KWK94, LW85, LLL13, LPH01, MA85, MS97a, Maz87, MS96c, Oli00, Oli01, Peu83, Pog85, ST80, SLKL11, Spe92, SB87, SBZ97, SS84a, TP75, TO99, TP92, WD96, XS98, YH83, ZT80, AT89, Big90, DSS94, FWWT13, eGZW07, KD10, KN03, KGK10, Kui08, Lhu08, LCZ09, MSI10, NKPT13, RLS06, TMQM13, TN07, WW90, XYZH11, YZT^{+}13, YT13, YG07, ZBLS13, LY13]. Structure-from-Motion [Jac01, Oli00, Oli01, TP92, BS05, RLS06, LY13]. Structured [DFP89, BHSD^{+}13, BB03, HK93, HW06, LCT09, WNH05]. structured-light [BHSD^{+}13]. Structures [AK78, Bur80, CCA92, FTW81, GMW83, JDP97, KMA^{+}00, Ley87a, LHH97, MS78, MMP85, Nis98, RSB93, SAA93, SM93, BBF^{+}11, BCL^{+}90, FPC^{+}08, FAB12, KZ05, KSG^{+}13, LVM04, RC13, Sha79a, VW80, WWWM12, YJA96]. Structuring [ZH86, SW05, SD90]. studies [FRDC06]. studio [RP08]. Study [CGP85, DF02, GMT00, HSSB98, LCZ^{+}01, Lin02, NESP10, Sup79, SH77, THO94, Wag76, BFRA12, DBZ07, FML12, GCFMT12, HS06, HF11, JM09b, LCP90, PSE^{+}11, PM82, SCD11, SYPK13, TT91, VD10]. studying [CU11]. style [CRH05]. stylization [LYKL12]. Stylus [MWL99, MWLA99]. Stylus-Generated [MWL99, MWLA99]. sub [GBF12, NRJ11, XJK12]. sub-pixel [GBF12, XJK12]. sub-sections [NRJ11]. Subband [KC92, Liu97]. Subdivision [BMZB02, CLR80, HWJ96, JK02, DQ04, HMESI13, LLXW13, MK05, MCQ05, MMS^{+}07, SS11a, UCB13, WX91]. subdivision-based [DQ04, MCQ05]. subfields [GH90]. Subgroup [HB98b]. subisomorphism [DSdlH^{+}11]. Subiteration [Ma96, PK99]. subject [LY06]. subjects [SSS13]. Submarine [CC00]. Submersible [NK00]. submersion [ZRKZ^{+}11]. suboptimal [GYH13]. Subpattern [ME98b]. Subpixel [Ano92b, BKLO87, BW93, CL00a, SKS97, TH86, UG92]. Subsea [TPR^{+}00]. Subset [JR09, LS94, vvv88, MVP06, YO11]. subsets [BRP04, DSNN08]. Subspace [DSY10, DD11b, FLHK08, MMP09]. Subspaces [FB97]. Substrate [HT98]. subtraction [BT05, DS07, ZCF13]. Successive [Liu97]. Successively [ZZ10]. Sufficient [Egg98, HA93]. suitable [CYE91, HZW^{+}10]. Sum [LKE98, GO94, MP03, PS07, VM06]. summaries [AWK04]. Summarization [CB98, ALK^{+}09, LHJ^{+}09, SLS03]. summarize [CH09]. Summarizing [PHY^{+}11]. Sun [Max91]. sunken [ZZLZ13]. super [AM06, FSV07, JC06, NFSD13]. super-resolution [AM06, FSV07, NFSD13]. super-resolved [JC06]. supercoupling [AKC11]. Superpipelined [DRAB08]. Superquadratic [GB93]. Superquadrics [ZK01, KS04b]. superresolution [BR12]. superresolution-inpainting [BR12]. supervised [SCvW11, TLWT12, WHM^{+}09]. supervision [FKS10]. Supplementary [Ano13v]. Support [GK98, KF86, SMB95, Zuc85, CMBP09, HGR^{+}13, HBG13, Hub12, SB13]. supported [OBS05]. supporting [OTO06]. suppression [BP95]. SURF [BETV08]. Surface [Ano95g, AK96, BG80, BM95, BJ86, BF87, BSF02, BH83a, Bri98, BM97, CLK09, CYH94, CLL^{+}99, CA86b, CMPP99, COK95, CP99, Dub76, EW87, FW97, Far86, FKW98, GSS00, Goo92, Gri83a, Gri83b, Gri84, GP85, GM87, GL98, HB98a, HL78, HW83, HSIW98, JC81, JJ94, JK02, JB89, iIK85, iK85, Kd88, KSd88, KZD^{+}11, KP97, KKK99, KU92, KPH02, LSB^{+}00, LM95, LM99b, LCC89, LTS93, LKE00, MK02, Mil99, Nis98, OG98, OD99, OD01, Oka84, PA82, PS00, QL96, SA96, SSF94, SZ96a, SL85, SK83a, SL96, SF97, Ter83, TS86, TH12, Udu82, VB98, VBH97, WH01, WH00, WL88, XL88, YA12, ZM96, AdVDI05, AFH81, BI11, BG79, BGR89, BMM^{+}07, CGAY13, CHSV08, CXY^{+}09, CHZ^{+}13, CL91, DQ05, DQ04, Far82, FR80, GYH13, GBHS06, GU89, HUF05, HB91]. surface [IO09, JXC^{+}13, KK90, KH90, Las92, LF04, LKC94, LY13, LT90a, MPST08, MMA06, MPVF11, MCQ05, MB05, MB95, OBS06b, PMW05, PZV13, RKH05, SY10, SKVS13, SOG09, TL05, TN05, TN08, TCCK90, UK12b, VSR12, WJG02, WPS03, WY11, WHHB12, WF05, XOF05, YW07, YYF89, YHS89, ELA91]. Surface-Based [HSIW98, OG98, KZD^{+}11]. Surface/Axis [BM95, LKC94]. Surfaces [Ano95h, BU93, BMZB02, BPYA85, CLL^{+}99, CCF01, CG94, CCA92, CJ82, DFP85, EK97, Elb01, FAB97, For72, Fra95, FL96, GA00, HHS^{+}01, Her92, Her93, HWJ96, Kim97, KSS00, KR85a, KM84, KM00, LM00, LC79, Lan84, Lev79, LTS93, LKK00, ML00, NFSK97, NFJ93, Pet00, PCV94, PLR96, Rab92, Ree84a, Rei96, Sar83, Sau99, SAG84, SS95b, SZKD99, SK01, SW86, TC87, WLH85, WH96, Zha99, AGCA06, AFSW03, BG91, BO05, BGK95, CP91, Eva06, FH12, GCB90, HKM12, HA03, Hub12, JBK04, KS03, LC11, LYA13, MK05, Mil09, MBMC11, MG95b, MP03, PJW11, PS07, PK05, RB89, RKW91, TG95c, TRS06, WSC^{+}12, XWYY10, YZZ^{+}10, ZQ11]. Surfaces-From [Ano95h]. surfel [CPP^{+}11]. surgery [ASFP03, PDA03]. surgical [ASFP03]. Surround [LCT09, EK12]. Surveillance [Fog93, CPC08, CHH09, GMW12, GWT09, MFB11, MW13, RCTV12, TMB12, VD10, WMBY12, YCKA10, Jon08]. Survey [Boe82, CF01, CL97b, Doe98, Gav99, Haw78, HL01, May99, Mil80a, MG01, MEDT96, NJ95, PR92b, Ros76, Ros86c, Ros89, SSWC88, VDO85, WVL83, WP84, Wes78, BCF06, BHF08, CCFC13, CH88, Dav75, Dav76, DFS08, GB10, HS06, IK88, JS07, JJN76, KR89, MHK06, Ros92, Sha79a, TA13, WKP13, WRB11, ZFG08]. suspicious [WMBY12]. Svd [SK79, DIOV06, ZZP12]. SVD-matching [DIOV06]. SVMs [BRA^{+}10]. SVP [FB05]. swarms [GA13]. Sweep [AKL93, OW86]. sweepers [ACWK06, AWC06]. Swept [AWC06]. Swimming [TML00]. Swirling [ACWK06]. Swirling-sweepers [ACWK06]. Switching [O’R82]. Sylvester [CS10]. Symbolic [Ano95h, KDRC98, KP00, MM90, SB87]. Symbols [BM79]. Symmetric [BH83a, Ehr78, GLD93, SK02, LA11, Nac82, RM06]. symmetrical [YJA96]. Symmetries [AK78, Big97, BKA84, JB92, ST96, Pon90]. Symmetry [BCM13, Fri86, HBA93, Ley87b, Per81, Rob96b, TS00b, VMUO95, YHR^{+}05, ZW97, ZBV93, JXCZ13, Mar89, YL90]. Symmetry-based [YHR^{+}05]. Symmetry-Curvature [Ley87b, YL90]. Symmetry-driven [BCM13]. Symposium [CPOO09, Jam09]. Synchronization [Boy04, TR09]. Synchronous [BTNS90]. synergies [PT08]. Synergistic [CUAT13, dMFU10]. synonyms [GSS12]. Syntactic [BM79, FF79, FK83, Fu80, KSd88, Lu78, Dor89, Fla89, IJDAB13]. syntactic/geometric [Dor89]. Syntactic/Semantic [KSd88]. Syntactical [SRL82]. Synthesis [Boo97, GM85, LF79, MN95, MSM81, Nis97, PS97, SLN95, YH78, AO03, CP91, CH06b, CCD11, GODC07, HKS06, OH06, SHK11, Sta05, SH05, SRML09, UBEP09, WS03, WZL^{+}03, ZK08, ZLH13]. Synthesizing [KJRA96, LPR^{+}03]. Synthetic [FY85, Lee81b, Mey88, TTIM96, BGR89, DM12, Zit88]. System [BKMSR98, BS99, BM96, CN95, CMRS98, CJC^{+}98, Eng83, Gud82, HC77, HM78, Her72, HF80, HH98, Koh81, KH83a, LCS84, Lee02, MFJ95, Mis84, ME98b, Oka88, PE92, Ree79, SBK^{+}99, TM86, Tan81a, THT^{+}98, Wam85, YC78b, YYL96, AGCA06, ABI^{+}04, AZSVK05, CJL06, CGW^{+}07, DLS^{+}09, DR04, DUSL94, DM78, ESS10, FG89, FFY^{+}04, FY06, FLCdA06, GSPL10, GODC07, HSKH07, HWW06, ILRB04, KPE90, KGFP10, Lhu08, MSG10, NL90, NKB11, PFGG09, RGA10, ÜB05, VK91, VD10, VZP^{+}09, YC78a, BCDH10, FRNS05, TG95a]. Systematic [Nag78, WP84, LS12]. Systems [BBC00, CL97b, EA95, FS95, GS92, HDM86, KS95b, Ken86, Kov86, LH99, Ley85, MHN84, NR88b, PA82, Ros86b, SC00a, She86, SS79, Tho86, Uhr86, Bar06, BHSD^{+}13, BRP04, CYP^{+}10, DLP13, GA09, HD07, Hil83, HZW^{+}10, JBS^{+}91, LM89, LFMP13, Mat89, OH05, PA13, SBB10, Tho10, TA11, WMBY12, YCA^{+}10, Nag86]. Systolic [CRT90, Nic95, SLY89].

T [DCL^{+}08]. T-meshes [DCL^{+}08]. Table [Bas81, GV84, GK95, SSS82, CXFS06, Reb89]. tables [Fiu91a, Fly92, HH82]. Tablet [BM79]. Tackling [HW94]. Tagging [CWH^{+}13]. Take [Lau97]. Taking [FL96]. tampering [KLL^{+}11]. Tangent [Zuc85]. Tangential [LKK00]. Tape [SHG^{+}88]. Target [IKST05, MYC09, JBC08, KW12, PMC13, UM05, VSP06, YCKA10]. Targets [CSR83, JN93, KPPK09, MC09a]. Tarr [CM94b]. Task [BH86, CL83, DC00b, GZJ05, SGB01, BRA^{+}10, BSMK13, ES06, HL13, RGA10]. task-driven [RGA10]. Task-Specific [DC00b, ES06]. Tasks [IKS86, KR99, Wec81]. Taylor [BKK11]. TBS [PT08]. TC [EHG^{+}10]. TC-12 [EHG^{+}10]. Teacher [EKY08]. Teacher-directed [EKY08]. team [PKK^{+}09]. Technical [OMLL98]. Technique [Ano01s, AAV96, BL01, CYH94, FTW81, HFC96, HH97, KSd88, LBD92, LCC89, Liu97, Luc01, McC82, Mok92, OD97, PLL00, RGC87, SH84, BL89, Cel90, CCH91, CCL04, DM12, KA12, MWF07, RC03, TM07a, YW07]. Techniques [Ano98e, Bas81, BY98, BS00b, CN87a, CF01, CY83b, CLR80, Dav76, DWX83, Fau81, Gar76, GKR02, GL86, HM84, HS85, Jar77, JJT91, KM84, MAP99, McD81a, MNSK98, Nad84, NHR81, PR92b, RD93, RA77, SSWC88, Sam82a, Shl83, SB79, SK01, TGB00, Wes78, YF80, AS09, BTCH05, Bre03, Dav75, FK09, HSSH89, HBG13, HS89, JJN76, JM09b, LCP90, MGPF08, MM05, OTO06, PSE^{+}11, PR03, SM13b, TLT91b, TA13].

Telepresence [OYTY98]. Television [MV86, SK83b]. Tell [KL77]. Telugu [RD77]. Template [CYES00, MSS90, RBA94, Tan81b, THT^{+}98, MR90a, MR90b, UBEP09, AW09]. Templates [DJG01, Gad91, LSB^{+}00, SL99, DLF06, GRGB^{+}13]. Temporal [AHD94, AHRW87, CA97, JT86, KSC97, LJ89, NP92, PD83, SA04, UFF06, Yac83, ACH^{+}13, CHMG12, CWLJ13, CSG^{+}03, DLF06, HDF12, LCSL07, NDO09, RL13a, WXRA07, XYW11]. Temporally [LYL10]. tennis [DGG08, YJC^{+}09]. Tension [Bar84, Dub76]. Tensor [AG00, Dod98, Kri84, LLC11, Sah05, XSD12, GYTL09, LR12, LBNS09, MGPJ11, Nor09, PLL12, PG13, RPG12, TH12]. Tensor-based [LLC11]. term [PA10a]. Terminal [KD85, SA81]. termination [WX91]. Terminator [UZC97]. Terms [Kis96b, Alb74]. Terrain [AK96, KK94, LPZ08, OMW^{+}07, The83, YYF89]. Terrains [dBD98]. Tessellation [AAS85, ITN84]. Tessellations [MFV80, RSFdM04]. Test [HNR84, LM96]. tested [FFFP07]. Testing [Kan91b, Kan94c, KK81, PPK93, RH06, Wu93]. Tests [ML78, de 83b, CLC91]. Tetra [JFS11]. Tetra-trees [JFS11]. tetrahedra [OK07]. Tetrahedral [ITF06, SMR98, BGTG04, BPG05, ZGLP12]. Text [AHD94, BKMSR98, DV98, Hob00, LAS94, WWC82, YT13, MTG07, YK95]. Text/Background [LAS94]. Text/Image [WWC82]. texton [ZZL13]. textons [XHJF12]. Textual [SLST99, LDC^{+}13]. Textural [AM00, CH80]. Texture [Acu92, BIP00, BL76, BSI87, Cav87, CP79, CH78, CTH84, DM82, EF78, EK88, Ekl79, GSP01, GPK99, GR92, GK90, HGA86, HSD85, Hsu79, HC96, JC94, KCC89, KC92, LSD^{+}07, LHS01, LJ90, Lu78, LF79, MSM81, NP92, PPT06, PB99, RW88, RL93a, RD93, RPTB01, SW83a, Shu97, SA02, SW83b, SM99, SC98, TZ82, VDO85, WR93, WR96, WH01, WD99, WC92, YH78, Zuc76b, ZT80, ASVO12, CCD11, DL10, GFL^{+}11, GB13, eGZW07, HS80, HOH^{+}07, HG11, HBL^{+}11, JC90, KORC10, LF08, LPVM13, MGPP11, Mig12, Pec91, Pun03, QAB^{+}11, RS91b, SG11, SF07, VBS^{+}04, WS89, XHJF12, ZK08, ZZL13]. texture-based [MGPP11]. Textured [Bid92, CF92, CJ82, DC86, JJ83, NC93, WD92, YK87, BGA05, CP91, HS89]. Textures [Dav79, DM80, DM81, GM85, LB97, LW85, MSN82, MMN83, PA97, PA98, KS91b, Sta05, WLW06, ZLH13]. Texturing [QY02, BI10, ZDL^{+}11]. Their [AK85, AGW85, Bie87, BC85, CS89, JT80, Kub84, LTS93, NSK^{+}97, Pot87, PLR96, SC00b, Wag76, Zuc85, vdWvO96, CTCG95, CKS^{+}05, FLB06, GCFMT12, LcTT91, SSM06, TG13, YYF89]. Theorem [TL88, YAT97, BFR13, HT89]. theorems [YL90]. Theoretic [HP78, RSB93, SA92, BEGB13, Kaw78, TT91, The83, WSSS13]. Theoretical [CL86, KOY86, KON87, THO94]. Theory [Ano94i, AS93b, Bie85, EPB05, Gri83a, Hab85, Har86, HT88, HNR88, HNRR90, HKA13, JN93, KR85b, McC80, Mok97, SUO00, SU01b, SWG02, Sch93, US96, VY94, AC07, AFH81, BC91, Bot78, DB03, KLBP11, NRJ11, XP11, HMEB07, KGK10, MUS06]. There [Tso94, Ver97, AQ09]. Thereof [Har86]. Thermal [SHJB^{+}83, DS07, HOH^{+}07, SSN03, TMB12, YCH07]. thermal-visible [TMB12]. Thermophysical [MNSK98]. thickness [Coe12, HSSH89, RNDA13]. Thin [AMMV99, CEP84, MAM97, TDK10]. Thinning [Arc81, BM95, Che98, CWSI87, CCS95, FK83, GS99, GH92, Ma94, Ma96, MS96a, MW00, MPJN87, MWL99, NKP11, PK99, Pav80, Pud98, TF81, LKC94, O’G90]. Thinnings [BJ96]. Third [Cav87, Ros87b]. Thoracic [LSB^{+}00, ML13]. thoroughly [PK05]. Threat [KR99]. Three [Art79, AT83, AD86, BSMG05, Bor96, CR89, CA86b, Col77, Dou81, ET94, FH84b, HP84, HGv87, HL79, JT80, Jos99, KK79, KK88b, Kri84, LD90, LJ91, Liu77, Lum83, MA85, MS09, MNHO00, MCPB99, Mit88, MH98, Mul92, OD01, PS97, PCR86, PW86, Ree84a, SGS01, SF95, TK97, THN92, TQ97, WR93, WR96, WD96, WN86, YC78b, YAT97, ZM94, ZM96, AFH81, BS92, Gar82, GU89, HQN05, LB08, Nac82, PJW11, SB05, TBN95, UA90, UKH88, WW80]. Three- [Kri84]. Three-Class [MCPB99]. Three-Dimensional [Art79, AT83, AD86, CA86b, Col77, Dou81, FH84b, HP84, HGv87, HL79, JT80, KK79, KK88b, Liu77, Lum83, MA85, MNHO00, Mit88, PS97, PCR86, Ree84a, SF95, TK97, THN92, TQ97, WR93, WR96, WD96, WN86, YC78b, YAT97, ZM94, ZM96, BSMG05, CR89, MS09, AFH81, Gar82, GU89, HQN05, LB08, Nac82, PJW11, SB05, TBN95, UA90, UKH88, WW80]. Three-frame [LD90]. Three-Light-Source [OD01]. Threshold [KIF85, MH79, Oli94, PG94, PS95, Wes78, Wil98, VR95, Wha91]. Thresholded [Pan78b]. Thresholding [BLd95, CDLD77, Gla93, HFC96, HS88, KSW85, Koh81, LL98, Pun81, Ros02, SSWC88, Sha94, Tsa85, WCZ02, WHL84, Abu89, GFL^{+}11, HDS08, LCP90]. Through-the-Lens [KKH96]. Tighter [Zha97b]. Tiles [ZK08]. Tilings [Mil99]. Tilt [CC00, DDLP10, SP06]. Time [Agg83, AGW85, BIP00, BEPW00, CBM01, CGP85, Eng83, FF79, Fog91, GR87a, HT98, HMD93, Lee91, LB98, LSKK10, LHHC98, MBDB88, MS85, Nag78, OYTY98, PD83, SKOS95, Sch92, SJ84, SY98, Sur86, TM86, TGB00, VV92b, Wam85, WW88, WZWT99, WL85, WBR86, ZXK02, AM04, BT05, BCMCB09, BDS12, BGTG04, BHMB10, CGH08, CCL04, DLS^{+}09, DDWZ12, FFM05, Fra81, Gon09, HZW^{+}10, JSRS08, DFP^{+}13, MZB^{+}10, MWTN04, MFS^{+}07, MTAA11, Nic95, PDA03, PGGM04, RAC^{+}13, RL13a, SM12, SS91, SS90b, SGH07, Sub90, SIT07, SHS03, UM05, WWLV11, YWZ11, ZZZY13, ZJ05, Ziv10, LBK10]. Time- [WW88]. Time-Dependent [SY98]. Time-of-Flight [LSKK10, BHMB10, LBK10]. time-to-collision [Sub90]. Time-Varying [Agg83, CBM01, CGP85, FF79, Fog91, Lee91, PD83, SKOS95, SJ84, TM86]. timing [TM07a]. tissue [CFYU12, DCS05, SRP10]. Tockner [Ano92a]. TOF [NB10, GPC^{+}10]. TOF-scans [NB10]. Tokens [SB87]. Tomograms [HL79, JGR85]. Tomographic [BKW96, Col77, VBN11]. Tomography [Art79, Her80, KMI79, SP81, SHS79, Ver81, BPBS13, IM06, XL88]. Tonal [Alg83]. Tone [GT84, JJN76]. Tony [Ano94i]. tool [DAM12, MR05]. Tools [PSM80]. top-down [KMN11]. Topmost [YK86]. Topographic [KK94, WLH85, BDL92b, TCC90, WY07]. Topography [WP93a]. Topological [ACF00, ASS97, AC07, Cou13, DBF04, Dam08, Eva06, GL95, HW83, HA03, LM00, Lat93, MST85, PA97, SMR98, SSP01a, SPW96, SA85, TTF04, VS08, ABD11, DR03, GFW13]. Topologically [CMPP99, EK88]. Topologies [EL03]. Topology [BCL96, BP94, Bre01, DM01, EF78, FWL88, Kaw82, Kaw83, Kov89, Ma94, ML00, NS96, OS95, RYN98, SR00, ZSCP08, DQ04, Eva11, Her90, KR89, LV03, Loh10, SC96, SS06, UCB13]. Topology-Based [FWL88]. Topology-Oriented [OS95]. Topology-Preserving [RYN98, Eva11]. Toppsy [Eng83]. Torsion [Mok97]. Torsion-Based [Mok97]. Torus [KKO98]. Torus/Sphere [KKO98]. Total [Kis96b]. totally [Ang07]. tourist [PHY^{+}11]. tower [XP11]. Trace [Lem79b]. traced [NRJ11]. Tracing [Arc81, Bid91, GR81, HC94, iK85, Pag97, Pag99, You86, CCL04, Lio91, MW13, WPK09]. Track [MW13, AVBK10, PT08]. Tracker [KSS97, TS01, AM04, SGH07]. trackers [TMN06]. Tracking [BF87, BL98b, DF01, Dem96, DJG01, FLB06, HFKN97, HW83, IP98, KS95b, KH98, KB95b, KU92, KH13, LRD99, MJ11, MJD^{+}00, MB94, ML78, MB85, PV13, Pet99, PF01, QL96, RAH97, RRS83, ROJX09, TTIM96, TPR^{+}00, VV92b, WN99, WS06, Ano06m, BSM10, BW11, BBH^{+}12, BCMCB09, BL09, BY12, BKMV07, CGH08, CKM11, CYP^{+}10, CPT07, CZZS07, DZL07, DBZ07, DD11a, DG11, DPT07, EDB12, GKK05, GLOC10, GB08, GRB13, GU89, GCFMT12, HD09, HYJ11, HP05, HH07, HGR^{+}13, HUF05, HW07, HDF12, HH12, IKST05, JSRS08, JBR08, JWDF05, JBC08, KBN12, KV06, KSR^{+}12, KGFP10, KW12, KPPK09, KT07, DFP^{+}13, LHYK05, LST13, LLR10, LSTF12, LA05, LN10, MYC09, MC09a, MZB^{+}10, MEYD11, MHSP10, MHMO09, MLH13, MD82, MM05, NHY10, NKB11, NLM05, OMBH06, PA10a, PD05]. tracking [PA06, PMC13, PYS03, RMD08, RRR11, RCTV12, SA04, SOG09, TID07, TMB12, TM07b, TP05, TTH07, UM05, UFF06, VSP06, WDB12, YWZ11, YNCO11, YJC^{+}09, ZN08, ZT09, ZYS09, ZJ05, ZCK09]. Tracks [Lee76]. Trade [LHH^{+}98]. Trade-offs [LHH^{+}98]. trademark [CFG06]. trademarks [PA10b]. Traffic [HMEB07, CSJ13, SJ12]. training [CHH09, CTCG95, FFFP07]. Trajectories [Bid91, CM94a, AAASC11, BN90, CHP^{+}11, KBN12, KL13, OCVV04, WCF10]. Trajectory [CK84, LB08, DN82, PKK^{+}09, YGC13]. trajectory-based [PKK^{+}09]. Transfer [ACW96, PKD07, TTF04, TFL^{+}09]. Transferral [SHG^{+}88]. Transfiguration [TTIM96]. Transfinite [VSR12]. Transform [AM00, AS88, BL94, BM00, BM02, CC97, CR97, CCMW97, CS01, Con88, CSR83, DGH98, DG01, EC88, FMRV94, GR92, HBA93, HNR88, HNRR90, KB00, KTNO97, Lea93, LHKC97, LH99, Lem79b, LLL86, LS92, LMM95, Liu97, MGK00, MNHO00, Mis84, MY87b, Mul92, NGC92, Oli94, OP96, PPK93, PKP97, RRS83, RP88, Ris89, ROH88, SWG02, SJ01, SK98, TV99, TS00a, TD83, Wat87, WBR86, WBR88, AKC11, BTNS90, ÇÖD08, CT10, CT12, CS04, CL95, dFCS93, Gre04, Hu11, IK88, IAP^{+}11, KB91c, Lea92, LY05, NSEA13, PIK90, SA04, SYK96, TWS06, XO93, ZS11, MSF^{+}12, PCC13, Sha06]. Transform-Coding [EC88]. Transformation [CM99b, Dav97, ER96, GLR^{+}99, Hum77, HS79, LB98, MMN83, iTTF82, WLH85, CGR13, DDWZ12, Fra89, KLV06, Mor90, OBH04, OH04, RK11, SC96, SG11, SW04, SY11]. Transformations [Ano01s, BS89, Big97, Bor84, Bor86, CK00, DC88, Dav93, Egg98, FL87, Har80b, Kis96a, KC87, LSBG92, Luc01, MS78, Nis96b, Pet85, Rag92, SC99b, Sha75, Sko86, Vos88, van86, BS92, Bor91, BDHM09, DL05, Eva11, Fit88, NKPT13, NESP10]. Transforming [ZL01, CLK09]. Transforms [BDL92a, Bor96, Bur81b, LBS80, OS87, Ols99, PJ88, Pag92, Pag97, RW95, Sam85, Sha78, SPW96, SK84, SB02, vv92a, AM93, CGL92, HQ12a, HQ12b, KM94, Lea92, Mar90, MS10, Nis96a, SS91, SB05]. Transition [YW99]. transitions [UK12a]. translate [Bha91]. Translated [MSW96]. Translating [DT96b, Nur86, OW83]. Translation [Wal88, WC99, BDVK10]. Translational [HJ12, Law83]. translations [LGJ82]. Translucency [Bri84]. Transmission [FM84, Tan79]. Transparency [Bec85, Bec86, Bri84, Bri86, Max84]. Transparent [YK87, KS12]. Transportable [KH83a, TST^{+}83]. Transposition [FSB85]. Trapezoidal [Wil84]. TRASMIL [YGC13]. travelogues [PHY^{+}11]. traversal [Vás11]. treatment [LcTT91]. TRECVid [SOD10]. Tree [Abe84, AK78, Bur80, Fu80, IKS86, LF79, PF87, WW97, BBF^{+}11, ÇÖD08, CT10, CTM^{+}13, GW90, Hu11, HQW^{+}12, JLD13, LZWP03, LVM04, RC13, TN07]. tree-based [JLD13]. tree-structure [TN07]. Trees [Ahu86, HdVL99, HS79, JT80, Jon99, LHKC97, Max91, Mil79b, Mun95, SS84b, Tam84, DSS94, GK04, JFS11, MU11, QT10]. Trends [De 88, SK86]. Tri [XS04]. Tri-view [XS04]. Triangle [LS92, AGCA06, LLXW13, QHXC12]. triangle/quad [LLXW13]. triangles [OK07, Žun03]. triangular [GC80, MSR07, Shi81]. Triangulated [KPH02, HA03]. Triangulation [DFP89, HS97, SP92, SL96, Tan95, dFP92, AFSW03, BS05, CH11, GYH13, Nor09]. Triangulations [WP84, WCH98]. Tribute [Kak97]. trigonometric [ZK05]. trihedral [San77]. Trilinear [Zha97b]. Trimmed [CMPP99, KM00]. Triplet [QV98, BP05]. truly [CU10b]. truncation [AM93]. trunk [TCH07]. truth [Cre08, SYPK13]. Tubes [VV92b]. Tubular [KMA^{+}00]. Tumor [RAC^{+}13, ZRL^{+}11]. Tuning [SHS79]. tunnel [RCTV12]. Turbulence [VV92b]. turn [CXFS06]. turn-table [CXFS06]. Turning [VFV93]. Tutor [FKS10]. Tutor-based [FKS10]. TV [ACDB12, DN82, LX88, YC78a, YC78b]. Twist [BBK78, Sel81]. Twisted [Buc88]. Two [AH08, BW98, Bat84, Bid86, CSDC96, CDH99, DM12, Egg98, ET94, FS84, FSSL86, FL87, GY99, HD97, HHS^{+}01, ITN84, Jos99, KWK94, KL77, Kri84, Kub84, LMKG85, LB97, LT81, LP79, Liu77, Liu97, MBK81, MZ96, Mer81, MH98, Muk92, Mul92, NDC86, Pie88, RYN98, SSN78, Shn81b, SP97d, Sko86, SK84, SA95, SS79, Ull81, Wei92, WB90, WLMG08, Zhu89, Abu89, ACAAC^{+}08, BKR^{+}89, BI10, Bha91, BYN^{+}04, DBF04, DV82, Fra89, GHZ^{+}13, Got08, JM09b, KNO^{+}09, LH90, LWGP08, MKS^{+}08, MG95b, Ros08, Sha11, SW04, SG82, SCCP05, TCH07, WZ08, WCF10, YGH11]. Two- [Kri84, Liu77]. two-component [Ros08]. Two-Dimensional [Bat84, FS84, FSSL86, FL87, GY99, HD97, ITN84, LMKG85, LB97, LT81, MZ96, Mer81, NDC86, Pie88, SSN78, Sko86, SS79, Ull81, AH08, Abu89, Bha91, DBF04, GHZ^{+}13, Got08]. Two-Directionally [Kub84]. two-fluid [LWGP08]. two-orthogonal [YGH11]. two-pass [DV82, Fra89]. two-phase [MKS^{+}08]. Two-Stage [CSDC96, SP97d, WLMG08]. two-step [BYN^{+}04]. Two-Valued [RYN98]. Two-View [Zhu89]. Types [KDK78, RWV95]. typical [MB95]. Typography [Her72].

Ultrasonic [KMI79]. Unaligned [SS90a]. Unbiased [Jos94, Ste13]. Uncalibrated [BK01, Tay00, VF96]. Uncertain [KN99, Pel79, PS05b]. Uncertainties [KK92, KK93, LM89, WR08]. Uncertainty [Cag93, CZZF97, GSS00, Shi99, CP04, CC03, DD11a, KG90, KT08, KN11, SS11b, TM07b]. unconstrained [DCH12, NKB11, PA10b]. Understand [MBMC11]. Understanding [AR77, AK11, Ano93d, Ano06m, BB91, Bie85, Bra97, CGL98, CTM^{+}13, CBB95, CL97b, DC00b, FWL88, Gol11, GMW12, HQ82, HF01, HDM86, KB98, LNY83, LSMS85, OBH04, PZ09, PT08, PEF92, ZT98, BHF08, HFR06, HAKK91, WKP13, Ano94f, LLE^{+}09]. Underwater [CFM02, GSV00, MCPB00, MT00, NK00, SWYP00, MN06]. Unexpected [AHZ96]. Unification [KKK99]. Unified [CWH^{+}13, Max91, MN95, RJ00, SA92, JLD13, KS89, LBM04, LH03, YZY11, ZK05]. Uniform [CL00b, JR86, KH86, Taj83, LX88, SAC09, TLCH05, XWYY10, ZK05]. Unifying [Pha89, SLST99, Bar06]. Unilateral [Gou91]. Unique [HA93, RAC^{+}13]. Uniqueness [Bat84, CM99a, FH84b, Lee91, OD01]. Unit [HB98b, MBDB88]. unity [OBS06b, TRS06]. Univariate [Dub77, Kas80]. Unix [LCS84].

Unix-Based [LCS84]. Unknown [FW97, OD99, Vel95, DQ04, GS06, Sal90, SSS13]. unlabeled [CHH09, Wu93]. Unmanned [NK00]. Unoccluded [PR92b]. unordered [MAL10]. Unorganized [JK02, ZOMK00, MK02, WSCO^{+}12]. unprepared [LA05]. Unregistered [HHI95]. Unresolved [CSR83]. Unscented [DG11]. unseen [RG10]. Unstructured [BCA98, CPS10, PLL12]. Unsupervised [BP05, BCM06, CHH09, CT10, CF92, HS89, JWL12, MGPP11, NC93, NHSC09, PP95, PB99, RM03, TVC09, TA11, WD92, YWMS08, GCEC07, ZFG08]. untextured [ÁB13]. Unwrapping [OH81]. up/top [KMN11]. Update [CMW^{+}97]. Updating [MS96c]. Upon [GY88]. upright [JWL12]. upsampling [XJK12]. Ur [Ano92b]. Urban [BM99, CTH84, FRL^{+}98, FMR01, HB98a, CM12, LS12, SJ12, WSCO^{+}12, YG07, ZN13]. Usage [NSK^{+}97]. Use [BBC00, Bid92, CN95, DR93, EFF98, Ekl79, GPK99, JT80, LP90b, Oka88, RWV95, SC97b, SGB01, TL88, Wah83, CU11, HSSH89, HS80, Loh10, SDPO81, Ano95h]. Used [SB85]. Useful [Coh85, GHMQ97]. User [CYES00, FUS^{+}98, IZKB12, KDV12, PJW11, PHY^{+}11, YWZ11]. user-assisted [PJW11]. user-contributed [IZKB12]. user-generated [PHY^{+}11]. User-Steered [FUS^{+}98]. Using [AHRW87, Acu92, AAS85, AEM98, APV99, Ant98, AD84, AA93, AMMV99, BKP10, BVL02, Bas81, BLd95, BCDH10, BH99, BB88, Ble84, BM86, BKD01, Bri98, BH95, COW98, Cai88, CW94, CC97, CM94a, CP79, CD93, CM95, CL97a, CS98, Che98, CLL^{+}99, CL00a, CK00, CA86a, CCA92, CJ82, CTH84, CSR83, CM99b, DT96a, DT96b, Dav97, DWX83, DUC97, DC86, DW87, DJG01, EW87, EU85, FBF08, FD99, FSSL86, FKL^{+}98, FTW81, Fog84, FSB85, FK99, GCB92, GV84, GR92, Gos89, GKR02, Gro82, GBB98, GL86, GLM78b, GJP96, GB93, GSK02, HB98a, HD97, Har80b, HCHD01, HR99, HB98b, HS88, Hob00, HN95, Hor77, HF93, HLF^{+}97, HBA93, Hum79, JW87, JC94, JM79, Jar77, Jon99, JB91, Jur99, KSW85, KS95a, KWK84, KC95, KSG84, KLK88]. Using [KB91a, KP97, KSI98, KHB01, KD76, KK92, KK93, KTNO97, KSS92, KMGC84, KC92, Lam84, LVW97, Lee76, Lee81a, LF82, LI00, LB00, LT81, LL95, LMM95, LL97a, LSHT02, LL97b, LJ87, LH88a, LH88b, Liu97, LZ97b, LF98, MBKB02, MA85, Mai81, MR90a, MR90b, Mar80, Mar89, MGK00, MAN84b, MS97b, Mea82, MK01, MY87a, MB94, Mil79a, Mis84, MB85, MF77, MB95, MS94, Mur87, Mur95, NG98b, NMP97, NP92, NDC86, NL96, Nis95, OPR78, Oli94, OJRT08, PKP97, PA00, PC99, Per81, PS97, PA97, PR92a, PSWH84, Pot77, Pra83b, Pri86b, RB82, Rag92, RJ94, RSB93, Rew84, RM98, Ros93b, Ros98b, Rub80, Rut81, Rut82, SAA93, STEK96, SK79, SYF99, SB95, SC00a, SP92, SWH84, SZKD99, Sel86, SO01, SGHM00, SB98b, SK83a, SY98, SP97b, Shi86, Shl83]. Using [Shn81a, SPK^{+}02, SHD86, SHKP98, SL99, SLL01, SF97, Spe97, SYPK13, SM94, Sug88, SS84b, SB02, SM97, SC98, TM94, TD83, TML00, TTIM96, TS86, Tsa96, Ül01, VB98, VK92, WSV91, WLH85, Wec79, WM93, WW97, WZWT99, WD92, YK87, Yam78, Yam80, Yan93a, YAT97, YKA01, YH83, YC98, YH78, YF80, ZW97, ZOMK00, de 83b, vv92a, Abu89, AM06, AS09, AW09, AC07, ABEN09, ALK^{+}09, AC09a, AC09b, ASCF13, ARARCE11, BW11, BG91, BS05, BRA^{+}10, BZS08, BP05, Beu91, BL09, BD94b, BWL04, BMM^{+}07, BF10, CGH08, CHP^{+}11, CFCP11, CMBP09, CH06a, CKM11, ÇÖD08, CT10, CT12, CGR13, CCL04, CPP^{+}11, CYW04a, CYW04b, CD95, Cho88, CFM^{+}13, CC03, Cre08, CKS^{+}05, DK13, DZL07, DT09, DBZ07, DM12, DGC12, DS07, DIOV06, DLF06]. using [DCS05, Dre96, DQ05, EKY08, ESS10, Eva06, FPC^{+}08, FB05, Fly92, FKS10, Fu80, FK09, GHZ^{+}13, GS06, GBHS06, Gho88, Gho90, GL82, GLM78a, Goh08, GH90, GA09, GDIIHK11, GK04, GFW13, GPC^{+}10, HK93, HKM12, HASS10, HE82, HA03, HY11, HPvB^{+}10, HMF10, HS89, Hu11, HQW^{+}12, HC13c, HH82, HKK08, IAP^{+}11, JKM07, JC90, JWG04, JBC08, JYTK11, JBWK11, JC06, KL07, KIK89, KPKPW90, KS03, KS89, KMBG09, KM94, Kim04, KLL^{+}11, KS04a, KM03, KS04b, KMN11, KNO^{+}09, LRW08, Lan91, DFP^{+}13, LHYK05, LY06, Lhu08, LCZ09, LB10, LC88a, LYG07, LHJ^{+}09, Liu10, LLC12, LDC^{+}13, LVM04, LPVM13, LAL^{+}10, LT97, LYA13, MGW10, ML13, MSI10, MDFS11b, MBH^{+}12, MZC^{+}05, Mil89, MSF^{+}12, MM06, MCF10, MDR91, Ney93, NNT11, O’G94, ODD96, OCVV04, PY08a, PZX13, PRR03, PC05, PLLL03, PW06]. using [PY08b, PA10b, PG13, PKD07, PHK92, PL08, PBG04, PW91, RB89, RG12, RRR11, ROJX09, RL13a, Ros10a, SY10, SCE04, SAS12, SIK92, Sau91, SJST07, SHC^{+}12, SW04, SZ07, SKU^{+}09, ST10, SAC09, SGH07, SKS11, SRHC13, SM13b, TMT10, TLGS05, TS11, TN07, TRG^{+}13, TR09, TKL^{+}09, VS08, VD90, WW80, WZ08, WS89, WJ07, WRB06, Wha91, WMBY12, WS90, WSKH13, WR08, WWJ13b, XYZH11, XAB07, YZZ^{+}10, YGH11, YK95, YG07, YC05, YLL12, ZC89, ZK08, ZHM11, ZXY^{+}12, ZZC^{+}13, ZZAA92, ZT09, ZYT10, ZS11, ZYS09, ZQ11, ZGK95, ZNG^{+}13, dLAH07, dMFU10]. Utility [DTG96]. Utilization [Chi81, O’G88, Sha94]. Utilizing [LHB87, KK11].

V [Ano94g, JC81]. V-S-S [JC81]. validating [ZB05]. Validation [SUO00, BY08, ZGK95]. Valleys [Har83b]. Valued [FS84, RYN98]. Values [MH79, iTTF82]. Vanishing [BO91, LJ87, MA84, BS04a]. variability [Dem05]. Variable [BCL^{+}90, GJH01, KB00, MGW10, SGH07, ZJ05]. Variable-Length [GJH01, SGH07]. variables [BW11, CLCO13]. Variance [Imm96, WH00]. Variant [BL98a, RC97, SK79]. variants [HF11, RH06]. Variation [Con88, TD83, GHZ^{+}13]. Variational [FKW98, HB86, ZOMK00, CHSV08, HW06, LJHH07, MCF10, RPG12, TRS06, dP10]. Variations [Gau92, Nag78, Nag83, NHR81, PAA^{+}87, SK88, TTIM96, LY06, SKVS13, TLCH05]. Various [RWV95, YWZ11, ZJ05]. Varying [Agg83, BFF97, Bic98, CBM01, CGP85, DW87, FF79, Fog91, Lai00, Lee91, PD83, SKOS95, SJ84, TM86, DL10, OK04, SB96a]. Vascular [WW97, WWWM12]. Vector [APV99, Ced79b, Che98, Fog91, Fra79, Mul92, PSK^{+}02, PS97, SYF99, SJ01, SM90, TA88, VS82, WW97, WSSD96, CMBP09, DF91, JWG04, LSPV04, MWF07, Pos77, See89, SB13, ZLS^{+}13]. Vector-Based [APV99]. Vector-City [SJ01]. Vectorgraph [LC85a]. Vectorial [ZUS06]. Vectorization [GL82, JV97, NL90, VRKL13]. Vectorized [CLD96, DL97, LCD97]. Vectorizer [Pav86]. Vectors [GK77, Nag83, Sel81]. Vegetation [Fog93]. Vegetation-Limited [Fog93]. Vehicle [KS95b, BKP10, GVK06, RCTV12]. Vehicles [HFKN97, NK00, SHJB^{+}83, SWYP00, JBC08, MFG10, TDWH07]. Velcro [NFSK97]. Velocimetry [FS95]. Velocity [CPD93, DB88, FT79, Fog91, Pet99, SM94, TA88, Yac83, LCSL07, SA04]. velocity-adapted [LCSL07]. Velocity-Based [CPD93]. venation [NHK08]. ventricle [BKR^{+}89, SG82, WSKH13, WWJ13b]. Vergence [CTE95, MGMS01, TM94, SB96a]. Verification [DLHT99, Gaa77, LVW97, MML87, ABEN09, CJL06, DM12, RSS07, SKSR08]. versatile [MZB^{+}10]. versus [HHWP03, KŽ12]. vertebra [ML13]. Vertex [Hob97, JC81, KLV06, LM12, San77]. Vertex-String-Surface [JC81]. Vertex-transformation [KLV06]. Vertices [BG80, BG79, nLPR91]. Very [Peu83, Sam82b, SGHM00]. Vessel [SGHM00, TKL^{+}09]. Via [FC86, TL79, YJ84, AAASC11, ANM98, BI11, BM95, CFYU12, CYH94, EK12, GYH13, GWT09, Hob97, KA08, KSKB95, KORC10, LKC94, LYSS12, Lyn81, MS09, MMS99, SC99a, SC97a, SMD^{+}08, Sto87, Tou80, TGSH98, WY11, YMA82, YWMS08, ZRKZ^{+}11]. Video [ALK99, AWK04, ADDK99, Bas81, CSJ13, CH06b, DCCL99, GSV00, HB05, HR99, HNB04, HH97, KSC97, LC09, Mai81, MSF^{+}12, NK00, OYTY98, PF01, SJ93a, SLS03, SOD10, TR09, TPR^{+}00, XL98, YYL98, YW99, ABI^{+}04, ALK^{+}09, Ano06m, AHDM10, AC09b, BZS08, BY12, CHH09, CCFC13, CPT07, CWLJ13, CC03, CSG^{+}03, CRH05, DK13, DCH12, DGG08, DRK03, DHP08, ESS10, FRDC06, FYH11, GKK05, GYTL09, GS06, GMW12, GWCO11, HMC10, HPvB^{+}10, JN09, JYTK11, KB12, KGU10, LYL10, LYKL12, LK03, LHJ^{+}09, LLE^{+}09, LLC11, LWH03, MWTN04, MSSS09, MĆK09, PSYZ13, PR03, PGGM04, RR06, SM12, SBS04, SYPK13, SH05, SMHH04, TD04, TY05, TCMS04, TMB12, TVC09, USKB10, VD10, WXRA07, WHM^{+}09, XG08b, YJC^{+}09, ZKC03]. Video-based [CSJ13, HNB04, ESS10]. video-hermeneutics [GMW12]. video-surveillance [GMW12]. videos [ABC^{+}03, CCTCR09, CD10, GBL08, KT07, LYA13, MW13, NDO09, RL13a, RCJ^{+}13, TD04]. Videoshop [WXRA07]. View [ASCF13, BL92, EK98, EF78, Gui98, HMF10, JGR85, KHB01, LC85b, Ley87a, MKW94, OD02, OYTY98, TTA94, YL94, Zhu89, ATC^{+}13, BF10, CPP^{+}11, CJ93, CC11, CH11, CCD11, CPS10, EKY08, HJ12, HKS06, HDF12, ITNP12, KM03, LHM06, MB11, RJ94, RM03, SMD^{+}08, SH05, TAK09, TVC09, WJ07, XS04, ZKRH04]. view-based [HDF12, TAK09]. View-Dependent [OYTY98]. view-independent [EKY08]. View-invariant [HMF10]. Viewed [Rab92, GCB90]. Viewer [SL85]. Viewer-Centered [SL85]. Viewing [CFA98, Chu02]. Viewpoint [DCTO97, OMBH06, SD92, WCZ^{+}07, CM12, DL10, LA11, MTVM04, WRB06]. Viewpoints [RWV95]. Views [BGSdVL98, BLP95, CFM02, Che91, EFF98, Lee86, LT90b, LV96, MFJ95, RFC97, SA95, ACAAC^{+}08, CKLP09, DPR92, Gol05, GSV05, JSRS08, KV06, LH90, PT08, RSPD12, SA90, SH08, SCCP05]. Virtual [EK98, LSZ83, Mur95, SH05, XTLP04, CCD11, DC04, HSKH07, RP08, RCVA11, YJC^{+}09, ZKRH04, FPDK12]. virtual-endoscopic [HSKH07]. Vis [AK11, MBMC11, PZ09]. Visibility [CD93, Fuj97, JS87, KM00, Lee83a, LL86, WR96]. Visible [GL98, HC94, JB89, MLF^{+}12, RWV95, CFB05, DS07, HD07, HASS10, PS12, SSN03, TN08, TMB12]. Visible-Surface [JB89]. Vision [AM94, Alo92, Ano93e, Ano96c, Ano98e, Ano06m, AA93, AAV96, BB92, BL98a, BY98, Bro94, BD02, BS87, BS88, CFS98, DB94, EBN^{+}07, Fel85, FMV93, FHP01, För87, GLOC10, Ger85, Gri83b, GY88, HT98, HB05, HTEB11, HSH07, Har86, Har94b, HF01, IKS86, IF95, JC81, Jai94, JBC08, JH98, Jol94, Kan94a, Ken86, KK92, KK93, Kov86, KR99, LVW97, Lee02, LRD99, LSHT02, LLE^{+}09, Mar80, Mee94, MST00, MG01, MPPG98, Mul88, MT00, Nag86, OBH04, Pag92, Per81, PEFM98, Pog85, Pop07, Ros86b, Ros87b, Ros88, Ros92, Ros93a, Ros94, Ros95, Ros96a, Ros97, Ros98a, Ros99a, Ros00a, Ros00b, Ros01, SG94, SB95, SC00a, SK86, She86, Shi94, SLK86, Td93, TTA94, TTG94, TB94a, TM94, Tho86, TS92, Tre85, Tso94, Uhr86, Ver97]. Vision [WWW89a, Wec81, WH94, YYL96, AK10, AK11, Ano05p, BPS10, BDVK10, BC10, BBC^{+}07, CKB10, CM92, DBZ07, Ham05, HD07, Hil83, HBH11, JBS^{+}91, KPKH07, KLBP11, KMT11, LBK10, Li92, MP09a, MR05, MFS^{+}07, MFG10, MHK06, PZ08, PZ09, PL07, Ros89, Ros90, Ros91, SGS^{+}10, Sah05, SBB10, SKS11, SST06, SIT07, SFWG08, TCB^{+}08, Tho10, UM05, VAWW10, VK91, VZP^{+}09, WWH07, WZ08, WKP13, WRB11, YHS95, ZKRH04, Ano94i, Ano95b, CM94b, HFR06, STLH08]. Vision-Based [HF01, KR99, MG01, EBN^{+}07, HSH07, Pop07, MR05, MHK06, WRB11, HFR06]. Vision-Guided [KK92, KK93]. Vision-Principles [Ano93e]. Visual [Åst97, Ano98e, BC88a, BY98, Bra97, Chi88, Col97, CS82, Fis94, Gav99, GSS12, GSV00, Gri83a, Gri84, GP85, GAD01, HPB94, HOH^{+}07, HF80, JN09, KRK11, Kol83, Kro86, Ku84, KR99, LHM06, LHYK05, Neg12, NJ95, Oka88, OMW^{+}07, PRW97a, PRW97b, PEF92, RJ00, RA77, SVS97, SLST99, SGTL09, ST10, Spe94, Sup02, TW98, TY01, Ter83, Wal87, WS08, YR06, AK91, ATC^{+}13, BBH^{+}12, BBHF10, BL08, BF05, CGR13, CYNO11, DLS^{+}09, DDLP10, DD11a, FMGA^{+}12, Far82, FFFP07, FAB12, FKS10, FLHK08, GCPF08, GBL08, GBW89, HD09, HYJ11, HH05, HM13, HWW06, HAGR91, ILRB04, JOvW^{+}05, KD10, KHA^{+}05, KYM13, LDC^{+}13, LSTARMB11, LN10, MPF07, NT10, NHY10, PY08a, PL10, SFWG08, THL13, TLMT^{+}05, TTH07, WRKP05, WZ04, vGSV^{+}10, BCDH10, Jon08]. visual-context-aware [PL10]. visual-object-based [SFWG08]. Visualization [CC00, Dod98, FST94, KK88a, LKE00, SK01, ACDB12, CBT^{+}04, CG04, MWTN04, Ros10b, YG07]. Visualizing [FS95, TN05]. Vlsi [HS83, Ahu86, LM91, RS88, RM91]. vocabulary [LSTARMB11]. Volume [Ano92b, Ano92a, Ano93a, Ano93b, Ano93c, Ano94c, Ano94d, Ano94e, Ano95e, Ano95f, Ano95d, Ano96e, Ano96f, Ano96d, Ano97c, Ano97d, Ano97e, Ano97f, Ano97b, Ano98b, Ano98c, Ano98a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00a, Ano01d, Ano01e, Ano01f, Ano01g, Ano01c, Ano01h, Ano01s, Ano02c, Ano02d, Ano02e, Ano02f, Ano02k, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano07l, Ano07m, Ano09a, BM97, CK00, CNK01, CA86b, FCG01, GS99, GW93a, KKK99, KK93, LBSP02, ACWK06, BYN^{+}04, BF05, ELA91, FWWT13, IA03, KSS08, LPR89]. volume [LB08, LPR^{+}03, RC06, SdB03, TTF04, Tan11, WBOL07, YJKK91, Oli01]. Volume/Surface [CA86b]. Volumes [FDMA97, LSB^{+}00, BZS08, DPR92, WRB06, WEY06]. Volumetric [AAV96, DN82, GS01, GSU00, GB93, NWP97, SBS04, TG95a, TK97, YL94, PFV^{+}11, THL03, VS08, YW07]. Voronoi [AAS85, AdVDI05, BBB96, BA92, KSI98, KZW12, NSK^{+}97, OS95, Sug93]. Vorticity [VV92b]. Voting [IF99, LZ97b, PSK^{+}02, LBNS09, MGPJ11, PLL12, RPG12, RC13, Sha06, SKBS13]. Voxel [GR87a, ALK^{+}09]. Voxelization [COK95, SK01]. Voxels [SR00, SB05]. Voyager [McD81b]. VRML [FPDK12]. Vs [Kan80a, KTP08, LHH^{+}98, TS00a].

W [Ano93d]. W. [Ano92a]. Walk [Sbe00]. walks [GB13]. Walsh [AM93, Mis84]. Walsh-Hadamard [Mis84]. Wang [ZK08]. Warping [ADRY94, CL98, FM98, HH98, LS94, YFZ98, LJHH07, PBM^{+}11]. water [PCR^{+}04]. Watershed [BL00]. Watershed-Based [BL00]. Wave [ACF00, KMI79, Sch80a, kWwZ13]. Waveform [Pra83b]. Wavefront [BB88]. Wavelength [Mey88]. wavelengths [PS12]. Wavelet [AM00, BL94, CL97a, DLHT99, HLKF95, LMM95, MAP99, Oli94, OP96, Rei96, SP97c, TS00a, ÇÖD08, CT10, CT12, HQ12a, HQ12b, Hu11, LBCA10, SG11]. Wavelet-Based [DLHT99, HLKF95]. Wavelets [Ano95g, Far11, WLZW04]. Way [Tso94]. weak [LH90]. wearable [NKB11, SE11]. Web [SLST99]. Webber [Ano94f]. Webs [GM87]. weight [CVP10, CKK^{+}12, HBG13, dMFU10]. Weighted [Har85, KD85, WVL81, ASCF13, HQW^{+}12, JBR08, LDC^{+}13, LSPV04, TL05, dCCP12]. Weiss [HM97, May97, Ver97]. weld [FLCdA06]. Well [LER95, Lat97, WB97, Oli91]. Well-Composed [LER95, Lat97, WB97]. well-constrained [Oli91]. What-and-Where [CGL98]. Where [CGL98, VZP^{+}09]. Which [Lea93]. while [TZM98]. White [Mor76]. whole [KL13]. whole-body [KL13]. Whyte [Ano93d]. Wide [Bid91, CKM11, SLST99, Rub82]. Wide-area [CKM11]. Width [AD84, Sin87]. Width-Independent [Sin87]. Wiener [LcTT91]. Wigner [JC90]. wild [HP05]. Winding [LP91]. Window [JLL13, KU95, Ree82, Bha91, GS08, YHN11]. Window-Based [KU95]. Windowing [TLT91b, LcTT91]. Wire [FUS^{+}98]. Wireframes [Las92]. Wireless [Ziv10, LWLS12]. Wise [DF02, SJ93a, AC09b, TCC90, YG07]. Within [ML00, BK03, Kou03]. without [CB98, CYES00, Ede94, FHMB84, iK85, OD99, PLH04, Rob96a, RSFdM04]. Wolff [Ano93e]. Word [KH96, KABP98, JN09]. workflows [KDV12]. Workshop [Ano94j, KB01, Ros87b, Uhr86]. workspace [RGA10]. World [Bla85, LSHT02, SLST99, SJ89, KPPK09, TLT91a]. wrinkle [CGW^{+}07]. Wrinkles [YB01]. wrist [KL13].

X [AS08b, GBR79, GM79, HT98, HC94, KS91c, KHB01, Mar82, Wag76]. X-ray [KS91c, GBR79, GM79, HC94, Mar82, Wag76, AS08b].

Year [Sch81, Sch82]. Years [AT13, Sch80b, SOD10]. YIQ [LL08].

ZDF [DBZ07]. Zenon [Hor79]. Zero [Mai76, MZ96, Yui89, ZM94, DN91, SB91, WW91]. zero-crossing [DN91, WW91]. Zero-Crossing-Based [MZ96, ZM94]. Zeta [CMRS98]. Zonal [CHB86]. Zone [BC88a]. zones [KG90, TRG^{+}13]. Zoom [MPPG98, PEFM98, TTG94, CXFS06, DDLP10, SP06, SSdVL06, TM07b]. Zoom-Invariant [MPPG98, PEFM98]. Zooming [LDPD97, ZZ07].

References

Arman:1993:CBV

[AA93] Farshid Arman and J. K.

Aggarwal. CAD-based vision: Object recognition in cluttered range images using recognition strategies. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(1):33–48, July 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1030/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1030/production/

pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1031/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1031/production/
pdf.

Abdelkader:2011:SBG

[AAASC11] Mohamed F. Abdelkader, Wael Abd-Almageed, Anuj Srivastava, and Rama Chellappa. Silhouette-based gesture and action recognition via modeling trajectories on Riemannian shape manifolds. Computer Vision and Image Understanding: CVIU, 115 (3):439–455, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ahuja:1985:IRU

[AAS85] Narendra Ahuja, Byong An, and Bruce Schachter. Image representation using Voronoi tessellation. Computer Vision, Graphics, and Image Processing, 29(3):286– 295, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Andres:1997:DAH

[AAS97] Eric Andres, Raj Acharya, and Claudio Sibata. Discrete analytical hyperplanes. Graphical Models and Image Processing: GMIP, 59 (5):302–309, September 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0427/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0427/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0427/production/ref.

Astruc:1996:CVN

[AAV96] Dominique Astruc, Jean-Régis Angilella, and Alain Vincent. The cone of vision: a new technique for interactive volumetric display. Graphical Models and Image Processing: GMIP, 58(4):387–393, July 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0031/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0031/production/
pdf.

Ali:1988:NAE

[AB88] S. M. Ali and R. E. Burge.

A new algorithm for extracting the interior of bounded regions based on chain coding. Computer Vision, Graphics, and Image Processing, 43(2):256–264, August 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Alvarez:2013:JAP

[ÁB13] Hugo Álvarez and Diego Borro. Junction assisted 3D pose retrieval of untextured 3D models in monocular images. Computer Vision and Image Understanding: CVIU, 117(10):1204–1214, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001671.

Assfalg:2003:SAS

[ABC^{+}03] Jürgen Assfalg, Marco Bertini, Carlo Colombo, Alberto Del Bimbo, and Walter Nunziati. Semantic annotation of soccer videos: automatic highlights identification. Computer Vision and Image Understanding: CVIU, 92(2–3):285–305, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alata:2011:GDP

[ABD11] O. Alata, S. Burg, and A. Dupas. Grouping/degrouping point process, a point process driven by geometrical and topological properties of a partition in regions. Computer Vision and Image Understanding: CVIU, 115(9): 1324–1339, September 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001251.

Abel:1984:BTS

[Abe84] D. J. Abel. A B^{+}-tree structure for large quadtrees. Computer Vision, Graphics, and Image Processing, 27(1):19– 31, July 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Amenta:1998:CSC

[ABE98] N. Amenta, M. Bern, and D. Eppstein. The crust and the β-skeleton: Combinatorial curve reconstruction. Graphical Models and Image Processing: GMIP, 60(2):125–??, ???? 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic).

Amayeh:2009:HBV

[ABEN09] Gholamreza Amayeh, George Bebis, Ali Erol, and Mircea Nicolescu. Hand-based verification and identification using palm–finger segmentation and fusion. Computer Vision and Image Understanding: CVIU, 113(4):477–501, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Amir:2004:MMS

[ABI^{+}04] Arnon Amir, Sankar Basu, Giridharan Iyengar, Ching-Yung Lin, Milind Naphade, John R. Smith, Savitha Srinivasan, and Belle Tseng. A multi-modal system for the retrieval of semantic video events. Computer Vision and Image Understanding: CVIU, 96(2):216–236, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alia:1987:ASN

[ABMT87] G. Alia, F. Barsi, E. Martinelli, and N. Tani. Angular spline: a new approach to the interpolation problem in computer graphics. Computer Vision, Graphics, and Image Processing, 39(1):56– 72, July 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Abutaleb:1989:ATG

[Abu89] Ahmed S. Abutaleb. Automatic thresholding of graylevel pictures using twodimensional entropy. Computer Vision, Graphics, and Image Processing, 47(1):22– 32, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Amini:1997:BIA

[ABW97] Amir A. Amini, Fred L. Bookstein, and David C. Wilson. Biomedical image analysis. Computer Vision and Image Understanding: CVIU, 66(2):95–96, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0616/production;   http:
                                        //www.idealibrary.com/links/

artid/cviu.1997.0616/production/
pdf.

Aggarwal:1999:HMA

[AC99] J. K. Aggarwal and Q. Cai.

Human motion analysis: a review. Computer Vision and Image Understanding: CVIU, 73(3):428–440, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0744/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0744/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0744/production/ref.

Allili:2007:TAS

[AC07] M. Allili and D. Corriveau.

Topological analysis of shapes using Morse theory. Computer Vision and Image Understanding: CVIU, 105(3): 188–199, March 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arandjelovic:2009:MRI

[AC09a] Ognjen Arandjelović and Roberto Cipolla. A methodology for rapid illuminationinvariant face recognition using image processing filters. Computer Vision and Image Understanding: CVIU, 113 (2):159–171, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arandjelovic:2009:PWL

[AC09b] Ognjen Arandjelović and Roberto Cipolla. A posewise linear illumination manifold model for face recognition using video. Computer Vision and Image Understanding: CVIU, 113(1): 113–125, January 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alberich-Carraminana:2008:RED

[ACAAC^{+}08] Maria Alberich-Carramiñana, Guillem Alenyà, Juan Andrade-Cetto, Elisa Martínez, and Carme Torras. Recovering epipolar direction from two affine views of a planar object. Computer Vision and Image Understanding: CVIU, 112 (2):195–209, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Araujo:1998:FPF

[ACB98] Helder Araújo, Rodrigo L. Carceroni, and Christopher M. Brown. A fully projective formulation to improve the accuracy of Lowe’s pose-estimation algorithm. Computer Vision and Image Understanding: CVIU, 70(2):227–238, May 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0632/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0632/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0632/production/ref.

Azari:2012:OPA

[ACDB12] Hossein Azari, Irene Cheng, Kostas Daniilidis, and Anup Basu. Optimal pixel aspect ratio for enhanced 3D TV visualization. Computer Vision and Image Understanding: CVIU, 116(1):38–53, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001871.

Adan:2000:MWS

[ACF00] Antonio Adán, Carlos Cerrada, and Vicente Feliu. Modeling wave set: Definition and application of a new topological organization for 3D object modeling. Computer Vision and Image Understanding: CVIU, 79(2):281–307, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0855;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0855/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0855/ref.

Alvarez:2009:NEB

[ACG^{+}09] L. Alvarez, C. A. Castaño, M. García, K. Krissian, L. Mazorra, A. Salgado, and J. Sánchez. A new energybased method for 3D motion estimation of incompressible PIV flows. Computer Vision and Image Understanding: CVIU, 113(7):802–810, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arcila:2013:STM

[ACH^{+}13] Romain Arcila, Cédric Cagniart, Franck Hétroy, Edmond Boyer, and Florent Dupont. Segmentation of temporal mesh sequences into rigidly moving components. Graphical Models, 75(1):10–22, January 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1524070312000720.

Aggarwal:1998:NMA

[ACLS98] J. K. Aggarwal, Q. Cai, W. Liao, and B. Sabata. Nonrigid motion analysis: Articulated and elastic motion. Computer Vision and Image Understanding: CVIU, 70(2):142–156, May 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0620/production;   http:

//www.idealibrary.com/links/
artid/cviu.1997.0620/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0620/production/ref.

Akleman:2003:MCS

[ACS03] Ergun Akleman, Jianer Chen, and Vinod Srinivasan. A minimal and complete set of operators for the development of robust manifold mesh modelers. Graphical Models, 65(5):286– 304, September 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Acuna:1992:TMU

[Acu92] Carmen O. Acuna. Texture modeling using Gibbs distributions. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(3): 210–222, May 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Azencott:1996:TFE

[ACW96] Robert Azencott, Bernard Chalmond, and Jia-Ping Wang. Transfer function estimation, film fusion and image restoration. Graphical Models and Image Processing: GMIP, 58(1):65–74, January 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.

                                        0004/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0004/production/
                                        pdf.

Angelidis:2006:SSC

[ACWK06] Alexis Angelidis, Marie-Paule Cani, Geoff Wyvill, and Scott King. Swirlingsweepers: Constant-volume modeling. Graphical Models, 68(4):324–332, July 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070306000038.

Arcelli:1984:AFD

[AD84] Carlo Arcelli and Gabriella Sanniti Di Baja. An approach to figure decomposition using width information. Computer Vision, Graphics, and Image Processing, 26(1):61– 72, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Augusteijn:1986:RRT

[AD86] Marijke F. Augusteijn and Charles R. Dyer. Recognition and recovery of the three-dimensional orientation of planar point patterns. Computer Vision, Graphics, and Image Processing, 36(1):76–99, October 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Allman:1993:CSR

[AD93] Mark Allman and Charles R.

Dyer. Computing spatiotemporal relations for dynamic perceptual organization. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3):338–351, November 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1046/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1046/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1048/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1048/production/
pdf.

Adams:1993:RDD

[Ada93]  R. Adams.  Radial de-

composition of disks and spheres. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(5):325–332, September 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1024/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1024/production/
pdf.

Avrithis:1999:SFO

[ADDK99] Yannis S. Avrithis, Anastasios D. Doulamis, Nikolaos D. Doulamis, and Stefanos D. Kollias. A stochastic framework for optimal key frame extraction from MPEG video databases. Computer Vision and Image Understanding: CVIU, 75(1– 2):3–24, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0761/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0761/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0761/production/ref.

Arad:1994:IWR

[ADRY94] Nur Arad, Nira Dyn, Daniel Reisfeld, and Yehezkel Yeshurun. Image warping by radial basis functions: Application to facial expressions. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(2):161–172, March 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1015/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1015/production/
pdf.

Alliez:2005:CVD

[AdVDI05] Pierre Alliez, Éric Colin de Verdière, Olivier Devillers, and Martin Isenburg. Centroidal Voronoi diagrams for isotropic surface remeshing. Graphical Models, 67 (3):204–231, May 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Artzy:1979:QOI

[AEH79] Ehud Artzy, Tommy Elfving, and Gabor T. Herman. Quadratic optimization for image reconstruction — 2. Computer Graphics and Image Processing, 11(3):242– 261, November 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Altunbasak:1998:RBP

[AEM98] Yucel Altunbasak, P. Erhan Eren, and A. Murat Tekalp. Region-based parametric motion segmentation using color information. Graphical Models and Image Processing: GMIP, 60(1):013–023, January 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0453/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0453/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0453/production/ref.

Algazi:1981:REN

[AF81] V. Ralph Algazi and Gary E. Ford. Radiometric equalization of nonperiodic striping in satellite data. Computer Graphics and Image Processing, 16(3):287–295, July 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Artzy:1981:TDI

[AFH81] Ehud Artzy, Gideon Frieder, and Gabor T. Herman. The theory, design, implementation and evaluation of a threedimensional surface detection algorithm. Computer Graphics and Image Processing, 15 (1):1–24, January 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Attene:2003:MIP

[AFSW03] Marco Attene, Bianca Falcidieno, Michela Spagnuolo, and Geoff Wyvill. A mappingindependent primitive for the triangulation of parametric surfaces. Graphical Models, 65 (5):260–273, September 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Alexander:2000:EMD

[AG00] Daniel C. Alexander and James C. Gee. Elastic matching of diffusion tensor images. Computer Vision and Image Understanding:

CVIU, 77(2):233–250, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0817;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0817/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0817/ref.

Allegre:2006:HMS

[AGCA06] Rémi Allègre, Eric Galin, Raphaëlle Chaine, and Samir Akkouche. The HybridTree: Mixing skeletal implicit surfaces, triangle meshes, and point sets in a free-form modeling system. Graphical Models, 68(1):42–64, January 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000561.

Aggerwal:1983:GES

[Agg83] J. K. Aggerwal. Guest editorial: Special issue on motion and time-varying imagery. Computer Vision, Graphics, and Image Processing, 21(1):1–2, January 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Alter-Gartenberg:1994:CIR

[AGHN94] Rachel Alter-Gartenberg, Friedrich O.

Huck, and Ramkumar Narayanswamy. Compact image representation by edge primitives. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):1–7, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1001/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1001/production/
                                        pdf.

Atkinson:1985:CRH

[AGW85] H. H. Atkinson, I. Gargantini, and T. R. S. Walsh. Counting regions, holes, and their nesting level in time proportional to the border. Computer Vision, Graphics, and Image Processing, 29 (2):196–215, February 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Atkinson:1986:FQO

[AGW86] H. H. Atkinson, I. Gargantini, and T. R. S. Walsh. Filling by quadrants or octants. Computer Vision, Graphics, and Image Processing, 33 (2):138–155, February 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Atkinson:2008:TDB

[AH08] Gary A. Atkinson and Edwin R. Hancock. Twodimensional BRDF estimation from polarisation. Computer Vision and Image Understanding: CVIU, 111(2): 126–141, August 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Abuhaiba:1994:PLH

[AHD94] I. S. I. Abuhaiba, M. J. J.

Holt, and S. Datta. Processing of off-line handwritten text: Polygonal approximation and enforcement of temporal information. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(4):324–335, July 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1029/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1029/production/
pdf.

Abuhaiba:1998:RLC

[AHD98] I. S. I. Abuhaiba, M. J. J.

Holt, and S. Datta. Recognition of off-line cursive handwriting. Computer Vision and Image Understanding: CVIU, 71(1):19–38, July 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0629/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1997.0629/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0629/production/ref.

Appiah:2010:AHV

[AHDM10] Kofi Appiah, Andrew Hunter, Patrick Dickinson, and Hongying Meng. Accelerated hardware video object segmentation: From foreground detection to connected components labelling. Computer Vision and Image Understanding: CVIU, 114(11):1282– 1291, November 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aharoni:1996:JG

[AHL96] Ron Aharoni, Gabor T. Herman, and Martin Loebl. Jordan graphs. Graphical Models and Image Processing: GMIP, 58(4):345–359, July 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0028/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0028/production/
                                        pdf.

Acharya:1987:HSI

[AHRW87] R. S. Acharya, P. B. Heffernan, R. A. Robb, and H. Wechsler. High-speed 3D imaging of the beating heart using temporal estimation. Computer Vision, Graphics, and Image Processing, 39 (3):279–290, September 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ahuja:1983:APD

[Ahu83] Narendra Ahuja. On approaches to polygonal decomposition for hierarchical image representation. Computer Vision, Graphics, and Image Processing, 21(3):200– 214, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ahuja:1986:EPE

[Ahu86] Narendra Ahuja. Efficient planar embedding of trees for VLSI layouts. Computer Vision, Graphics, and Image Processing, 34(2):189– 203, May 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Aykroyd:1996:USP

[AHZ96] R. G. Aykroyd, J. G. B.

Haigh, and S. Zimeras. Unexpected spatial patterns in exponential family auto models. Graphical Models and Image Processing: GMIP, 58(5):452–463, September 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0037/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0037/production/
                                        pdf.

Agrawala:1977:SAE

[AK77] Ashok K. Agrawala and Ashok V. Kulkarni. Sequential approach to the extraction of shape features. Computer Graphics and Image Processing, 6(6):538– 557, December 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Alexandridis:1978:PDT

[AK78] Nikitas Alexandridis and Allen Klinger. Picture decomposition, tree data-structures, and identifying directional symmetries as node combinations. Computer Graphics and Image Processing, 8 (1):43–77, August 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Anderson:1985:RDL

[AK85] Timothy A. Anderson and Chul E. Kim. Representation of digital line segments and their preimages. Computer Vision, Graphics, and Image Processing, 30(3):279–288, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Aoyama:1991:PLA

[AK91] H. Aoyama and M. Kawagoe.

A piecewise linear approximation method preserving visual feature points of original figures. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):435– 446, September 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Arakawa:1996:FMN

[AK96] Kenichi Arakawa and Eric Krotkov. Fractal modeling of natural terrain: Analysis and surface reconstruction with range data. Graphical Models and Image Processing: GMIP, 58(5):413–436, September 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0035/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0035/production/
pdf.

Ambrosch:2010:AHB

[AK10] Kristian Ambrosch and Wilfried Kubinger. Accurate hardware-based stereo vision. Computer Vision and Image Understanding: CVIU, 114(11):1303–1316, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ambrosch:2011:CAH

[AK11] Kristian Ambrosch and Wilfried Kubinger. Corrigendum to “Accurate hardwarebased stereo vision” [Comput. Vis. Image Understanding 114 (2010) 1303–1316]. Computer Vision and Image Understanding: CVIU, 115 (2):287, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arashloo:2011:PIF

[AKC11] Shervin Rahimzadeh Arashloo, Josef Kittler, and William J. Christmas. Pose-invariant face recognition by matching on multi-resolution MRFs linked by supercoupling transform. Computer Vision and Image Understanding: CVIU, 115(7):1073–1083, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000749.

Ahn:1993:AGS

[AKL93] Jae-Woo W. Ahn, Myung-Soo S. Kim, and Soon-Bum B. Lim. Approximate general sweep boundary of a 2D curved object. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(2):98–128, March 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1008/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1008/production/
pdf.

Amir:1999:GFD

[AL99] Arnon Amir and M. Lindenbaum. Ground from figure discrimination. Computer Vision and Image Understanding: CVIU, 76(1):7–18, October 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0786/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0786/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0786/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0797/production;
 http://www.idealibrary.
com/links/artid/cviu.1999.
0797/production/pdf;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0797/production/
ref.

Aristidou:2011:FFI

[AL11] Andreas Aristidou and Joan Lasenby. FABRIK: a fast, iterative solver for the inverse kinematics problem. Graphical Models, 73(5):243– 260, September 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070311000178.

Albano:1974:RDC

[Alb74] A. Albano. Representation of digitized contours in terms of conic arc and straight-line segments. Computer Graphics and Image Processing, 3 (1):23–33, March 1974. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Algie:1983:RTC

[Alg83] Stephen H. Algie. Resolution and tonal continuity in bilevel printed picture quality. Computer Vision, Graphics, and Image Processing, 24 (3):329–346, December 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ali:1977:MPG

[Ali77] Moonis Ali. Mathematical picture grammar applied to script generation. Computer Graphics and Image Processing, 6(1):93–102, February 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Adjeroh:1999:DMV

[ALK99] Donald A. Adjeroh, M. C. Lee, and Irwin King. A distance measure for video sequences. Computer Vision and Image Understanding: CVIU, 75(1–2): 25–45, July/August 1999.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0764/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0764/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0764/production/ref.

Anderson:2009:LSV

[ALK^{+}09] Derek Anderson, Robert H.

Luke, James M. Keller, Marjorie Skubic, Marilyn Rantz, and Myra Aud. Linguistic summarization of video for fall detection using voxel person and fuzzy logic. Computer Vision and Image Understanding: CVIU, 113(1): 80–89, January 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aloimonos:1992:PQA

[Alo92] Yiannis Aloimonos. Purposive, qualitative, active vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):1–??, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Aloimonos:1994:WL

[Alo94]  Y. Aloimonos.  What I

have learned. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):74–85, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1032/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1032/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1037/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1037/production/
pdf.

Abe:2006:MBP

[ALP06] Yeuhi Abe, C. Karen Liu, and Zoran Popović. Momentumbased parameterization of dynamic character motion. Graphical Models, 68(2):194– 211, March 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000299.

Andres:2011:GPB

[ALSR11] Eric Andres, Gaëlle Largeteau-Skapin, and Marc Rodríguez. Generalized perpendicular bisector and exhaustive discrete circle recognition. Graphical Models, 73(6):354–364, November 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000233.

Arcelli:1978:PGS

[AM78a] Carlo Arcelli and Antonio Massarotti. On the parallel generation of straight digital lines. Computer Graphics and Image Processing, 7(1): 67–83, February 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ashkar:1978:CEP

[AM78b] G. P. Ashkar and J. W. Modestino. Contour extraction problem with biomedical applications. Computer Graphics and Image Processing, 17 (3):331–335, June 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Anguh:1993:TMC

[AM93] M. M. Anguh and R. R. Martin. A truncation method for computing Walsh transforms with applications to image processing. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(6):482–493, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1036/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1036/production/
pdf.

Aggarwal:1994:RRR

[AM94] J. K. Aggarwal and W. N. Martin. The role of R and R in vision: Is it a matter of definition? Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):100–102, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1038/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1038/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1043/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1043/production/
pdf.

Attali:1997:CSC

[AM97] Dominique Attali and Annick Montanvert. Computing and simplifying 2D and 3D continuous skeletons. Computer Vision and Image Understanding: CVIU, 67(3):261–273, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0536/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0536/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0536/production/ref.

Angel:2000:AMW

[AM00] Paul Angel and Colin Morris. Analyzing the Mallat wavelet transform to delineate contour and textural features. Computer Vision and Image Understanding: CVIU, 80(3):267–288, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0877;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0877/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0877/ref.

Aguado:2001:PLS

[AM01] Alberto S. Aguado and Eugenia Montiel. Progressive linear search for stereo matching and its application to interframe interpolation. Computer Vision and Image Understanding: CVIU, 81(1):46–71, January 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0886;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0886/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0886/ref.

Anderson:2004:RRT

[AM04] Keith Anderson and Peter W.

McOwan. Robust real-time face tracker for cluttered environments. Computer Vision and Image Understanding: CVIU, 95(2):184–200, August 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aguena:2006:MID

[AM06] Marcia L. S. Aguena and Nelson D. A. Mascarenhas. Multispectral image data fusion using POCS and superresolution. Computer Vision and Image Understanding: CVIU, 102(2):178–187, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ali:1979:CBC

[AMA79] M. Ali, W. N. Martin, and J. K. Aggarwal. Colorbased computer analysis of aerial photographs. Computer Graphics and Image Processing, 9(3):282–293, March 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Amir:1990:AFC

[Ami90] Israel Amir. Algorithm for finding the center of circular fiducials. Computer Vision, Graphics, and Image Processing, 49(3):398– 406, March 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Armande:1999:TNE

[AMMV99] N. Armande, P. Montesinos, O. Monga, and Guy Vaysseix. Thin nets extraction using a multi-scale approach. Computer Vision and Image Understanding: CVIU, 73(2):248–257, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0658/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0658/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0658/production/ref.

Ahuja:1984:ORM

[AN84] Narendra Ahuja and Charles Nash. Octree representations of moving objects. Computer Vision, Graphics, and Image Processing, 26(2):207– 216, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Andres:2003:DLO

[And03] Eric Andres. Discrete linear objects in dimension n: the standard model. Graphical Models, 65(1–3):92–111, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Angulo:2007:MCO

[Ang07] Jesús Angulo. Morphological colour operators in totally ordered lattices based on distances: Application to image filtering, enhancement and analysis. Computer Vision and Image Understanding: CVIU, 107(1– 2):56–73, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aguado:1998:PAS

[ANM98] Alberto S. Aguado, Mark S. Nixon, and M. Eugenia Montiel. Parameterizing arbitrary shapes via Fourier descriptors for evidence-gathering extraction. Computer Vision and Image Understanding: CVIU, 69(2):202–221, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0558/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0558/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0558/production/ref.

Anonymous:1992:VNW

[Ano92a]  Anonymous.  Volume 45,

number 1 (1989): W. G. Kropatsch and H. Tockner, “Detecting the Straightness of Digital Curves in O(N) Steps,” pp. 1–21. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(2): 269–??, September 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). See [KT89].

Anonymous:1992:VNH

[Ano92b] Anonymous. Volume 54, number 2 (1992): Hanoch Ur and Daniel Gross, “Improved Resolution from Subpixel Shifted Pictures,” pp. 181–186. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4):365–??, July 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). See [UG92].

Anonymous:1993:AIVa

[Ano93a]  Anonymous.  Author in-

dex for volume 55. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(6):544, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1042/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1042/production/
pdf.

Anonymous:1993:AIVb

[Ano93b] Anonymous. Author index for volume 57. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):401, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1027/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1027/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1028/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1028/production/
pdf.

Anonymous:1993:AIVc

[Ano93c] Anonymous. Author index for volume 58. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3):399, November 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1050/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1050/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1052/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1052/production/
pdf.

Anonymous:1993:RDS

[Ano93d] Anonymous. Review of Directed Sonar Sensing for Mobile Robot Navigation, by J. J. Leonard and H. F. Durrant-Whyte and Review of Qualitative Motion Understanding, by W. Burger and B. Bhanu. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(1):136, July 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1993.
1035/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1035/production/
pdf.

Anonymous:1993:RPB

[Ano93e]  Anonymous.  Review of

Physics-Based Vision-Principles and Practice, by Steven A. Shafer, Glenn E. Healey, and Lawrence B. Wolff. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):400, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1993.
1027/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1027/production/
pdf.

Anonymous:1994:RA

[Ano94a]  Anonymous.  1993–1994

reviewer acknowledgment. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(6):510– 511, November 1994. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.

                                        com/links/artid/cgip.1994.
                                        1046/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1046/production/
                                        pdf.

Anonymous:1994:A

[Ano94b]  Anonymous.  Announce-

ment. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):136, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1010/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1010/production/
                                        pdf.

Anonymous:1994:AIVa

[Ano94c]  Anonymous.  Author in-

dex for volume 56. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(6):513, November 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1047/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1047/production/
                                        pdf.

Anonymous:1994:AIVb

[Ano94d] Anonymous. Author index for volume 59. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):406, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1027/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1027/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1032/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1032/production/
pdf.

Anonymous:1994:AIVc

[Ano94e] Anonymous. Author index for volume 60. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3):399, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1066/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1066/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1072/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1072/production/
pdf.

Anonymous:1994:RPA

[Ano94f] Anonymous. Review of A Physical Approach to Color Image Understanding, by Gudrun J. Klinker and Review of Stimulating Humans: Computer Graphics, Animation, and Control, by Norman I. Badler, Cary B. Phillips, and Bonnie Lynn Webber. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):135, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1009/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1009/production/
                                        pdf.

Anonymous:1994:RIR

[Ano94g] Anonymous. Review of Image Representation and Processing — A Recursive Approach, by V. V. Alexandrov and N. D. Gorsky and Review of Neural Network Perception for Mobile Robot Guidance, by Dean A. Pomerleau. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):262, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1020/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1020/production/
                                        pdf.

Anonymous:1994:RRL

[Ano94h] Anonymous. Review of Robot Learning, by Jonathon H.

Connell and Sridhar Mahadevan. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):405, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1994.
1031/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1031/production/
pdf.

Anonymous:1994:RSS

[Ano94i]  Anonymous.  Review of

Scale-Space Theory in Computer Vision, by Tony Lindeberg. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):266, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1994.
1062/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1062/production/
pdf.

Anonymous:1994:SIW

[Ano94j] Anonymous. Second international workshop on digital mammography. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):263, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1021/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1021/production/
                                        pdf.

Anonymous:1995:RA

[Ano95a]  Anonymous.  1994-1995

reviewer acknowledgment. Graphical Models and Image Processing: GMIP, 57 (6):539, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1046/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1046/production/
                                        pdf.

Anonymous:1995:ASA

[Ano95b] Anonymous. ACCV ’95 Second Asian Conference on Computer Vision. Computer Vision and Image Understanding: CVIU, 61(1): 151, January 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/artid/
                                        cviu.1995.1011/production;
                                         http://www.idealibrary.
                                        com/links/artid/cviu.1995.
                                        1011/production/pdf.

Anonymous:1995:A

[Ano95c]  Anonymous.  Announce-

ment. Graphical Models and Image Processing:

GMIP, 57(1):79, January 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1009/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1009/production/
pdf.

Anonymous:1995:AIVc

[Ano95d] Anonymous. Author index for volume 57. Graphical Models and Image Processing: GMIP, 57(6):540, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1047/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1047/production/
pdf.

Anonymous:1995:AIVa

[Ano95e] Anonymous. Author index for volume 61. Computer Vision and Image Understanding: CVIU, 61(3):475, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1036/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1036/production/
pdf.

Anonymous:1995:AIVb

[Ano95f] Anonymous. Author index for volume 62. Computer Vision and Image Understanding: CVIU, 62(3):392, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1063/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1063/production/
pdf.

Anonymous:1995:RMM

[Ano95g] Anonymous. Review of Mathematical Morphology in Image Processing, by E. R. Dougherty and review of Wavelets, Images and Surface Fitting, by P.-J. Laurent, A. LeMéHauté, and L. L. Schumaker. Computer Vision and Image Understanding: CVIU, 61(2):292, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1022/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1022/production/
pdf.

Anonymous:1995:RNC

[Ano95h] Anonymous. Review of NURB Curves and Surfaces-From Projective Geometry to Practical Use, by G. Farin and review of An Introduction to Scientific, Symbolic, and Graphical Computation, by E. Fiume. Computer Vision and Image Understanding: CVIU, 62(1):144, July 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1045/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1045/production/
pdf.

Anonymous:1996:RA

[Ano96a]  Anonymous.  1995–1996

reviewer acknowledgment. Graphical Models and Image Processing: GMIP, 58 (6):593, November 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0048/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0048/production/
pdf.

Anonymous:1996:A

[Ano96b] Anonymous. ANNOUNCE-MENT. Graphical Models and Image Processing: GMIP, 58(1):99, January 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0008/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0008/production/
pdf.

Anonymous:1996:ABS

[Ano96c] Anonymous. ANNOUNCE-MENT: BMVC96 Seventh British Machine Vision Conference. Computer Vision and Image Understanding: CVIU, 63(2):397, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0030/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0030/production/
                                        pdf.

Anonymous:1996:AIVc

[Ano96d] Anonymous. Author index for volume 58. Graphical Models and Image Processing: GMIP, 58(6):594, November 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0049/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0049/production/
                                        pdf.

Anonymous:1996:AIVa

[Ano96e] Anonymous. Author index for volume 63. Computer Vision and Image Understanding: CVIU, 63(3):613, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0042/production;   http:

//www.idealibrary.com/links/
artid/cviu.1996.0042/production/
pdf.

Anonymous:1996:AIVb

[Ano96f] Anonymous. Author index for volume 64. Computer Vision and Image Understanding: CVIU, 64(3):443, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0071/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0071/production/
pdf.

Anonymous:1996:IFI

[Ano96g] Anonymous. ICDAR ’97:

Fourth international conference on document analysis and recognition. Computer Vision and Image Understanding: CVIU, 64(2): 303–304, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0061/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0061/production/
pdf.

Anonymous:1997:A

[Ano97a]  Anonymous.  Announce-

ment. Computer Vision and Image Understanding: CVIU, 67(3):324–??, September 1997. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0642/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0642/production/
                                        pdf.

Anonymous:1997:AIVe

[Ano97b] Anonymous. Author index for volume 59. Graphical Models and Image Processing: GMIP, 59(6):496–??, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0458/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0458/production/
                                        pdf.

Anonymous:1997:AIVa

[Ano97c] Anonymous. Author index for volume 65. Computer Vision and Image Understanding: CVIU, 65(3):455, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0622/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0622/production/
                                        pdf.

Anonymous:1997:AIVb

[Ano97d] Anonymous. Author index for volume 66. Computer Vision and Image Understanding: CVIU, 66(3):347–??, June 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0636/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0636/production/
pdf.

Anonymous:1997:AIVc

[Ano97e] Anonymous. Author index for volume 67. Computer Vision and Image Understanding: CVIU, 67(3):325–??, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0654/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0654/production/
pdf.

Anonymous:1997:AIVd

[Ano97f] Anonymous. Author index for volume 68. Computer Vision and Image Understanding: CVIU, 68(3):363–??, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0671/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0671/production/
pdf.

Anonymous:1997:BRR

[Ano97g] Anonymous. Books received for review. Computer Vision and Image Understanding: CVIU, 65(1):109, January 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0588/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0588/production/
pdf.

Anonymous:1997:RA

[Ano97h] Anonymous. Reviewer acknowledgment. Graphical Models and Image Processing: GMIP, 59(6):495–??, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0457/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0457/production/
pdf.

Anonymous:1998:AIVc

[Ano98a] Anonymous. Author index for volume 60. Graphical Models and Image Processing: GMIP, 60(6):489, November 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0486/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0486/production/
pdf.

Anonymous:1998:AIVa

[Ano98b] Anonymous. Author index for volume 71. Computer Vision and Image Understanding: CVIU, 71(3):449, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0725/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0725/production/
pdf.

Anonymous:1998:AIVb

[Ano98c] Anonymous. Author index for volume 72. Computer Vision and Image Understanding: CVIU, 72(3):414, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0740/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0740/production/
pdf.

Anonymous:1998:BRR

[Ano98d] Anonymous. BOOKS RE-CEIVED FOR REVIEW. Computer Vision and Image Understanding: CVIU, 70(1):120, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1998.0677/production;
http://www.idealibrary.

com/links/artid/cviu.1998.
0677/production/pdf.

Anonymous:1998:CVV

[Ano98e]  Anonymous.  Computer

vision for visual computing: Techniques and applications. Computer Vision and Image Understanding: CVIU, 71(2):153, August 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0714/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0714/production/
pdf.

Anonymous:1998:RA

[Ano98f] Anonymous. Reviewer acknowledgment. Graphical Models and Image Processing: GMIP, 60(6):488, November 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0485/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0485/production/
pdf.

Anonymous:1999:AIVe

[Ano99a] Anonymous. Author index for volume 61. Graphical Models and Image Processing: GMIP, 61(6):375, November 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0508/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0508/production/
pdf.

Anonymous:1999:AIVa

[Ano99b] Anonymous. Author index for volume 73. Computer Vision and Image Understanding: CVIU, 73(3):455, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0749/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0749/production/
pdf.

Anonymous:1999:AIVb

[Ano99c] Anonymous. Author index for volume 74. Computer Vision and Image Understanding: CVIU, 74(3):236, June 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0774/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0774/production/
pdf.

Anonymous:1999:AIVc

[Ano99d] Anonymous. Author index for volume 75. Computer Vision and Image Understanding: CVIU, 75(3):319, September 1999. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0796/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0796/production/
                                        pdf.

Anonymous:1999:AIVd

[Ano99e] Anonymous. Author index for volume 76. Computer Vision and Image Understanding: CVIU, 76(3):298, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0820/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0820/production/
                                        pdf.

Anonymous:1999:E

[Ano99f]  Anonymous.  ERRATUM.

Graphical Models and Image Processing: GMIP, 61 (6):369–371, November 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0509/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1999.0509/production/
                                        pdf.

Anonymous:1999:RA

[Ano99g] Anonymous. REVIEWER ACKNOWLEDGMENT. Graphical Models and Image Processing: GMIP, 61(6):373, November 1999. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0510/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0510/production/
pdf.

Anonymous:2000:AIVe

[Ano00a] Anonymous. Author index for volume 62. Graphical Models, 62(6):447, November 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0535;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0535/
pdf.

Anonymous:2000:AIVa

[Ano00b] Anonymous. Author index for volume 77. Computer Vision and Image Understanding: CVIU, 77(3):388, March 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0839;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0839/
pdf.

Anonymous:2000:AIVb

[Ano00c] Anonymous. Author index for volume 78. Computer Vision and Image Understanding: CVIU, 78(3):374, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0863;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0863/
pdf.

Anonymous:2000:AIVc

[Ano00d] Anonymous. Author index for volume 79. Computer Vision and Image Understanding: CVIU, 79(3):442, September 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0873;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0873/
pdf.

Anonymous:2000:AIVd

[Ano00e] Anonymous. Author index for volume 80. Computer Vision and Image Understanding: CVIU, 80(3):384, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0898;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0898/
pdf.

Anonymous:2000:RA

[Ano00f] Anonymous. Reviewer acknowledgment. Graphical Models, 62(6):445, November 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0534;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0534/
pdf.

Anonymous:2001:Aa

[Ano01a] Anonymous. Announcement.

Computer Vision and Image Understanding: CVIU, 83 (2):??, August 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:Ab

[Ano01b] Anonymous. Announcement.

Computer Vision and Image Understanding: CVIU, 84(1): ??, October 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:AIVe

[Ano01c] Anonymous. Author index for volume 63. Graphical Models, 63(6):480, November 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2001:AIVa

[Ano01d] Anonymous. Author index for volume 81. Computer Vision and Image Understanding: CVIU, 81(3):446, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0911;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0911/
pdf.

Anonymous:2001:AIVb

[Ano01e] Anonymous. Author index for volume 82. Computer Vision and Image Understanding: CVIU, 82(3):255, June 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0928;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0928/
pdf.

Anonymous:2001:AIVc

[Ano01f] Anonymous. Author index for volume 83. Computer Vision and Image Understanding: CVIU, 83(3):296, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0942;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0942/
pdf.

Anonymous:2001:AIVd

[Ano01g] Anonymous. Author index for volume 84. Computer Vision and Image Understanding: CVIU, 84(3):409, December 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:CPS

[Ano01h] Anonymous. Call for papers: Special issue on volume modeling. Graphical Models, 63(1):61–62, January 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0538;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0538/
pdf.

Anonymous:2001:EAb

[Ano01i] Anonymous. Editorial announcement. Graphical Models, 63(2):63–64, March 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0545;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0545/
pdf.

Anonymous:2001:EAa

[Ano01j] Anonymous. Editorial announcement. Graphical Models, 63(2):??, March 2001.

CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2001:GE

[Ano01k] Anonymous. Guest editorial. Computer Vision and Image Understanding: CVIU, 84(1):??, October 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:Na

[Ano01l] Anonymous. Note. Computer Vision and Image Understanding: CVIU, 82(3):??, June 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:Nb

[Ano01m] Anonymous. Note. Computer Vision and Image Understanding: CVIU, 83(1):??, July 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:Nc

[Ano01n] Anonymous. Note. Computer Vision and Image Understanding: CVIU, 83(3): ??, September 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2001:P

[Ano01o] Anonymous. Preface. Graphical Models, 63(2):??, March 2001. CODEN GRMOFM.

ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2001:RA

[Ano01p] Anonymous. REVIEWER ACKNOWLEDGMENT. Graphical Models, 63(6):479, November 2001. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2001:SIF

[Ano01q] Anonymous. Special issue on face recognition: Call for papers. Computer Vision and Image Understanding: CVIU, 84(1):198–199, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0947;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0947/
pdf.

Anonymous:2001:SIN

[Ano01r]  Anonymous.  Special is-

sue on nonrigid image registration: Call for papers. Computer Vision and Image Understanding: CVIU, 83(2):192–193, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0929;   http://www.
idealibrary.com/links/doi/

                                        10.1006/cviu.2001.0929/
                                        pdf.

Anonymous:2001:VNA

[Ano01s]  Anonymous.  Volume 81,

number 1 (2001), in the article “A Frequency Domain Technique Based on Energy Radial Projections for Robust Estimation of Global 2D Affine Transformations,” by Luca Lucchese, pages 72–116. Computer Vision and Image Understanding: CVIU, 82(1):82–83, April 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0913;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0913/
                                        pdf.

Anonymous:2002:Aa

[Ano02a] Anonymous. Announcement. Graphical Models, 64(5):333, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:Ab

[Ano02b] Anonymous. Announcement. Graphical Models, 64(6):396, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:AIVa

[Ano02c] Anonymous. Author index for volume 85. Computer Vision and Image Understanding: CVIU, 85(3):232, March 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2002:AIVb

[Ano02d] Anonymous. Author index for volume 86. Computer Vision and Image Understanding: CVIU, 86(3):191, June 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2002:AIVc

[Ano02e] Anonymous. Author index for volume 87. Computer Vision and Image Understanding: CVIU, 87(1–3):131, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2002:AIVd

[Ano02f] Anonymous. Author index for volume 88. Computer Vision and Image Understanding: CVIU, 88(3):189, December 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2002:EBa

[Ano02g] Anonymous. Editorial Board.

Graphical Models, 64(5):C2, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:EBb

[Ano02h] Anonymous. Editorial Board. Graphical Models, 64(6):C2, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:PN

[Ano02i] Anonymous. Publisher’s note. Graphical Models, 64(5):iii– iv, September 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:RAL

[Ano02j] Anonymous. Reviewer acknowledgement list. Graphical Models, 64(6):397, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2002:VAI

[Ano02k] Anonymous. Volume author index. Graphical Models, 64 (6):398, November 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:Aa

[Ano03a] Anonymous. Announcement. Computer Vision and Image Understanding: CVIU, 89(1): 108, January 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:Ab

[Ano03b] Anonymous. Announcement.

Computer Vision and Image Understanding: CVIU, 89(2– 3):318, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:Ac

[Ano03c] Anonymous. Announcement.

Computer Vision and Image Understanding: CVIU, 90 (1):128, April 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBa

[Ano03d] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 89(1): C02, January 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBb

[Ano03e] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 89 (2–3):C02, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBc

[Ano03f] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 90 (1):C02, April 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBd

[Ano03g] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 90 (2):C02, May 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBe

[Ano03h] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 90 (3):C02, June 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBf

[Ano03i] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 91 (1–2):C02, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBg

[Ano03j] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 91(3): C02, September 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBh

[Ano03k] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 92(1):

C02, October 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBi

[Ano03l] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 92(2– 3):CO2, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:EBj

[Ano03m] Anonymous. Editorial Board.

Graphical Models, 65(1–3): C2, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:EBk

[Ano03n] Anonymous. Editorial Board.

Graphical Models, 65(4): C2, July 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:EBl

[Ano03o] Anonymous. Editorial Board.

Graphical Models, 65(5):C2, September 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:EBm

[Ano03p] Anonymous. Editorial Board.

Graphical Models, 65(6):CO2, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:PN

[Ano03q] Anonymous. Publisher’s note. Computer Vision and Image Understanding: CVIU, 89(1): iii–iv, January 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAIa

[Ano03r] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 89(2–3):319, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAIb

[Ano03s] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 90(3):313, June 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAIc

[Ano03t] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 91(3):368, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAId

[Ano03u]  Anonymous.  Volume au-

thor index. Computer Vision and Image Understanding: CVIU, 92(2–3):306, November/December 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2003:VAIe

[Ano03v] Anonymous. Volume author index. Graphical Models, 65 (6):405, November 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2003:VRA

[Ano03w]  Anonymous.  Volume re-

viewer acknowledgement list. Graphical Models, 65(6):406, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBg

[Ano04a] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 95(3): CO2, September 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBh

[Ano04b] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 96(1): CO2, October 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBi

[Ano04c] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 96(2): CO2, November 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBj

[Ano04d] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 96(3): CO2, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBa

[Ano04e] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 93(1): CO2, January 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBb

[Ano04f] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 93(2): CO2, February 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBc

[Ano04g] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 93(3): CO2, March 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBd

[Ano04h] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 94 (1–3):CO2, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBe

[Ano04i] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 95 (1):CO2, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBf

[Ano04j] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 95(2): CO2, August 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:EBk

[Ano04k] Anonymous. Editorial Board.

Graphical Models, 66(1):CO2, January 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBl

[Ano04l] Anonymous. Editorial Board.

Graphical Models, 66(2): CO2, March 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBm

[Ano04m] Anonymous. Editorial Board. Graphical Models, 66(3): CO2, May 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBn

[Ano04n] Anonymous. Editorial Board. Graphical Models, 66(4): CO2, July 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBo

[Ano04o] Anonymous. Editorial Board. Graphical Models, 66(5):CO2, September 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:EBp

[Ano04p] Anonymous. Editorial Board. Graphical Models, 66(6):CO2, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:PG

[Ano04q] Anonymous. Pacific Graphics 2003. Graphical Models, 66 (6):331–332, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:VAIa

[Ano04r] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 93(3):347, March 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:VAIb

[Ano04s] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 94(1–3):311, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:VAIc

[Ano04t] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 95(3):354, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:VAId

[Ano04u]  Anonymous.  Volume au-

thor index. Computer Vision and Image Understanding: CVIU, 96(3):472, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2004:VAIe

[Ano04v] Anonymous. Volume author index. Graphical Models, 66 (6):440, November 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2004:VRA

[Ano04w]  Anonymous.  Volume re-

viewer acknowledgement list. Graphical Models, 66(6):439, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBa

[Ano05a] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 97(1): CO2, January 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBb

[Ano05b] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 97(2): CO2, February 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBc

[Ano05c] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 97(3): CO2, March 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBd

[Ano05d] Anonymous. Ed. board. Computer Vision and Image Understanding: CVIU, 98(1): CO2, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBe

[Ano05e] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 98 (2):CO2, May 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBf

[Ano05f] Anonymous. Editorial board.

Computer Vision and Image Understanding: CVIU, 98 (3):CO2, June 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBg

[Ano05g] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 99 (1):CO2, July 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBh

[Ano05h] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 99(2): CO2, August 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBi

[Ano05i] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 100 (1–2):??, October/November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:EBj

[Ano05j] Anonymous. Editorial Board. Graphical Models, 67(1):CO2, January 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBk

[Ano05k] Anonymous. Editorial Board. Graphical Models, 67(2): CO2, March 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBl

[Ano05l] Anonymous. Editorial board. Graphical Models, 67(3): CO2, May 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBm

[Ano05m] Anonymous. Editorial board. Graphical Models, 67(4): CO2, July 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Anonymous:2005:EBn

[Ano05n] Anonymous. Editorial Board. Graphical Models, 67(5): C2, September 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000469.

Anonymous:2005:EBo

[Ano05o] Anonymous. Editorial Board.

Graphical Models, 67(6): C2, November 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000603.

Anonymous:2005:SIA

[Ano05p] Anonymous. Special issue:

Attention and performance in computer vision. Computer Vision and Image Understanding: CVIU, 100(1–2): 1–2, October/November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAIa

[Ano05q] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 97(3):384, March 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAIb

[Ano05r] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 98(3):513, June 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAIc

[Ano05s] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 99(3):527, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAId

[Ano05t] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 100(3):458, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2005:VAIe

[Ano05u]  Anonymous.  Volume au-

thor index. Graphical Models, 67(6):622, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000640.

Anonymous:2005:VRA

[Ano05v]  Anonymous.  Volume re-

viewer acknowledgement list. Graphical Models, 67(6): 621, November 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000639.

Anonymous:2006:CEB

[Ano06a]  Anonymous.  C2 — ed-

itorial board. Graphical Models, 68(5–6):??, September/November 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000701.

Anonymous:2006:EBf

[Ano06b]  Anonymous.  Ed. Board.

Computer Vision and Image Understanding: CVIU, 103 (2):??, August 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBg

[Ano06c]  Anonymous.  Ed. Board.

Computer Vision and Image Understanding: CVIU, 104 (1):??, October 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBa

[Ano06d] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 101 (1):??, January 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBb

[Ano06e] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 101 (2):??, February 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBc

[Ano06f] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 102 (1):??, April 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBd

[Ano06g] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 102 (2):??, May 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBe

[Ano06h] Anonymous. Editorial board. Computer Vision and Image Understanding: CVIU, 103 (1):??, July 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:EBh

[Ano06i] Anonymous. Editorial Board. Graphical Models, 68(1): C2, January 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000858.

Anonymous:2006:EBi

[Ano06j] Anonymous. Editorial Board. Graphical Models, 68(2): C2, March 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000063.

Anonymous:2006:EBj

[Ano06k] Anonymous. Editorial Board.

Graphical Models, 68(3): C2, May 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000257.

Anonymous:2006:EBk

[Ano06l] Anonymous. Editorial board.

Graphical Models, 68(4): ??, July 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1524070306000452.

Anonymous:2006:EMK

[Ano06m]  Anonymous.  Erratum to

“A multi-Kalman filtering approach for video tracking of human-delineated objects in cluttered environments” [Comput. Vision Image Understanding 99 (2005) 1–57]. Computer Vision and Image Understanding: CVIU, 102 (3):259, June 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [GKK05].

Anonymous:2006:PN

[Ano06n]  Anonymous.  Publisher’s

Note. Computer Vision and Image Understanding: CVIU, 101(1):??, January 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:SI

[Ano06o]  Anonymous.  Special is-

sue. Graphical Models, 68 (2):65, March 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000354.

Anonymous:2006:VAIa

[Ano06p] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 101(3):204, March 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:VAIb

[Ano06q] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 102(3):260–316, June 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:VAIc

[Ano06r] Anonymous. Volume author index. Computer Vision and Image Understanding: CVIU, 103(3):229, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:VAId

[Ano06s]  Anonymous.  Volume au-

thor index. Computer Vision and Image Understanding: CVIU, 104(2–3):258, November/December 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2006:VRA

[Ano06t]  Anonymous.  Volume re-

viewer acknowledgement list. Graphical Models, 68(5–6): 496, September/November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000786.

Anonymous:2007:CEBd

[Ano07a] Anonymous. C2 — editorial board. Graphical Models, 69(1):??, January 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000889.

Anonymous:2007:CEBe

[Ano07b] Anonymous. C2 — editorial board. Graphical Models, 69(2):??, March 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070307000057.

Anonymous:2007:CEBf

[Ano07c] Anonymous. C2 — editorial board. Graphical Models, 69(3–4):??, May/July 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070307000185.

Anonymous:2007:CEBg

[Ano07d]  Anonymous.  C2 — ed-

itorial board. Graphical Models, 69(5–6):??, September/November 2007. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S152407030700029X.

Anonymous:2007:CEBa

[Ano07e] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 105(1):??, January 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:CEBb

[Ano07f] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 107(1–2):??, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:CEBc

[Ano07g] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 108(1–2):??, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:EBa

[Ano07h] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 105 (2):??, February 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:EBb

[Ano07i] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 106 (1):??, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:MCPa

[Ano07j] Anonymous. Most Cited Paper Award. Computer Vision and Image Understanding: CVIU, 106(1):1–2, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2007:MCPb

[Ano07k] Anonymous. Most Cited Paper Award. Graphical Models, 69(3–4):159, May/July 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070307000136.

Anonymous:2007:VAI

[Ano07l]  Anonymous.  Volume au-

thor index. Graphical Models, 69(5–6):276, September/November 2007. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070307000331.

Anonymous:2007:VRA

[Ano07m]  Anonymous.  Volume re-

viewer acknowledgement list. Graphical Models, 69(5–6): 275, September/November 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S152407030700032X.

Anonymous:2008:CEBk

[Ano08a] Anonymous. C2 — editorial board. Graphical Models, 70(1–2):??, January/March 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070308000076.

Anonymous:2008:CEBl

[Ano08b] Anonymous. C2 — editorial board. Graphical Models, 70(3):??, May 2008. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070308000106.

Anonymous:2008:CEBm

[Ano08c] Anonymous. C2 — editorial board. Graphical Models, 70(4):??, July 2008. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000118.

Anonymous:2008:CEBn

[Ano08d] Anonymous. C2 — editorial board. Graphical Models, 70(5):??, September 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407030800012X.

Anonymous:2008:CEBo

[Ano08e] Anonymous. C2 — editorial board. Graphical Models, 70(6):??, November 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000131.

Anonymous:2008:CEBa

[Ano08f] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 109(1):??, January 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBb

[Ano08g] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 109(2):??, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBc

[Ano08h] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 110(1):??, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBd

[Ano08i] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 110(2):??, May 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBe

[Ano08j] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 111(1):??, July 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBf

[Ano08k] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 111(2):??, August 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBg

[Ano08l] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 111(3):??, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBh

[Ano08m] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 112(1):??, October 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBi

[Ano08n] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 112(2):??, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:CEBj

[Ano08o] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 112(3):??, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2008:MCP

[Ano08p]  Anonymous.  Most Cited

Paper Award. Computer Vision and Image Understanding: CVIU, 111(3):247– 248, September 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:AIV

[Ano09a] Anonymous. Author index for volume 71. Graphical Models, 71(6):242, November 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000459.

Anonymous:2009:CEBl

[Ano09b] Anonymous. C2 — editorial board. Graphical Models, 71(4):??, July 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000228.

Anonymous:2009:CEBa

[Ano09c] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(1):??, January 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBb

[Ano09d] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(2):??, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBc

[Ano09e] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(3):??, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBd

[Ano09f] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(4):??, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBe

[Ano09g] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(5):??, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBf

[Ano09h] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(6):??, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBg

[Ano09i] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 113(7):??, July 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBh

[Ano09j] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(8):??, August 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBi

[Ano09k] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(9):??, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBj

[Ano09l] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(10):??, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:CEBk

[Ano09m] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 113(11):??, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2009:EBa

[Ano09n] Anonymous. Editorial Board. Graphical Models, 71(1):

C2, January 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000083.

Anonymous:2009:EBb

[Ano09o] Anonymous. Editorial Board.

Graphical Models, 71(2): C2, March 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000113.

Anonymous:2009:EBc

[Ano09p] Anonymous. Editorial Board.

Graphical Models, 71(3): C2, May 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000162.

Anonymous:2009:EBd

[Ano09q] Anonymous. Editorial Board.

Graphical Models, 71(5): ??, September 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000241.

Anonymous:2009:EBe

[Ano09r] Anonymous. Editorial Board.

Graphical Models, 71(6): ??, November 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070309000629.

Anonymous:2009:RA

[Ano09s] Anonymous. Reviewer acknowledgment. Graphical Models, 71(6):240–241, November 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070309000307.

Anonymous:2010:CEBl

[Ano10a] Anonymous. C2 — editorial board. Graphical Models, 72(3):??, May 2010. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000093.

Anonymous:2010:CEBm

[Ano10b] Anonymous. C2 — editorial board. Graphical Models, 72(4):i–ii, July 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000111.

Anonymous:2010:CEBa

[Ano10c] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(1):??, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBb

[Ano10d] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(2):??, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBc

[Ano10e] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(3):??, March 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBd

[Ano10f] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(4):??, April 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBe

[Ano10g] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(5):??, May 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBf

[Ano10h] Anonymous. COV2: Ed.

Board. Computer Vision and Image Understanding: CVIU, 114(6):??, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBg

[Ano10i] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(7):??, July 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBh

[Ano10j] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(8):??, August 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBi

[Ano10k] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(9):??, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBj

[Ano10l] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(10):??, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:CEBk

[Ano10m] Anonymous. COV2: Ed. Board. Computer Vision and Image Understanding: CVIU, 114(11):??, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2010:EBa

[Ano10n] Anonymous. Editorial Board.

Graphical Models, 72(1): ??, January 2010. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070310000044.

Anonymous:2010:EBb

[Ano10o] Anonymous. Editorial Board.

Graphical Models, 72(2): ??, March 2010. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031000007X.

Anonymous:2010:EBc

[Ano10p] Anonymous. Editorial board.

Graphical Models, 72(6):i– ii, November 2010. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031000038X.

Anonymous:2010:RA

[Ano10q] Anonymous. Reviewer acknowledgment. Graphical Models, 72(6):74–75, November 2010. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1524070310000378.

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 115 (1):??, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 115 (2):??, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 115 (3):??, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 115 (4):??, April 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBe

[Ano11e] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 115 (5):??, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Anonymous:2011:EBf

[Ano11f] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 115(7):??, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001111.

Anonymous:2011:EBg

[Ano11g] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 115 (8):??, August 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001299.

Anonymous:2011:EBh

[Ano11h] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 115(9):??, September 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001469.

Anonymous:2011:EBi

[Ano11i] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 115(10):??, October 2011.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421100172X.

Anonymous:2011:EBj

[Ano11j] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 115(11):??, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001901.

Anonymous:2011:EBk

[Ano11k] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 115(12):??, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002189.

Anonymous:2011:EBl

[Ano11l] Anonymous. Editorial Board. Graphical Models, 73(1):i– ii, January 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000445.

Anonymous:2011:EBm

[Ano11m] Anonymous. Editorial Board. Graphical Models, 73(2):i– ii, March 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000038.

Anonymous:2011:EBn

[Ano11n] Anonymous. Editorial Board.

Graphical Models, 73(3): i–ii, May 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000130.

Anonymous:2011:EBo

[Ano11o] Anonymous. Editorial Board.

Graphical Models, 73(5):i– ii, September 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000269.

Anonymous:2011:EBp

[Ano11p] Anonymous. Editorial Board.

Graphical Models, 73(6):i– ii, November 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000300.

Anonymous:2011:RA

[Ano11q] Anonymous. Reviewer acknowledgment. Graphical Models, 73(6):376, November 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000440.

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 116(1):??, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002426.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial board. Computer Vision and Image Understanding: CVIU, 116(2):??, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002657.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 116 (3):??, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000057.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 116(4):??, April 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000240.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 116(5):??, May 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000409.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial board.

Computer Vision and Image Understanding: CVIU, 116(6):??, June 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000550.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial board.

Computer Vision and Image Understanding: CVIU, 116(7):??, July 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000689.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial board.

Computer Vision and Image Understanding: CVIU, 116 (8):??, August 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000811.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 116(9):??, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000926.

Anonymous:2012:EBm

[Ano12j] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 116(10):??, October 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001087.

Anonymous:2012:EBq

[Ano12k] Anonymous. Editorial board. Computer Vision and Image Understanding: CVIU, 116(11):??, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421200121X.

Anonymous:2012:EBr

[Ano12l] Anonymous. Editorial board. Computer Vision and Image Understanding: CVIU, 116(12):??, December 2012.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001373.

Anonymous:2012:EBj

[Ano12m] Anonymous. Editorial Board.

Graphical Models, 74(2):i– ii, March 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000070.

Anonymous:2012:EBk

[Ano12n] Anonymous. Editorial Board.

Graphical Models, 74(4): i–ii, July 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000392.

Anonymous:2012:EBl

[Ano12o] Anonymous. Editorial Board.

Graphical Models, 74(5):i– ii, September 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000409.

Anonymous:2012:EBp

[Ano12p] Anonymous. Editorial Board.

Graphical Models, 74(6):i– ii, November 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000434.

Anonymous:2012:MCP

[Ano12q] Anonymous. Most cited paper award. Computer Vision and Image Understanding: CVIU, 116(5):661–662, May 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000343.

Anonymous:2012:RAa

[Ano12r] Anonymous. Reviewer acknowledgment. Computer Vision and Image Understanding: CVIU, 116(1):??, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002542.

Anonymous:2012:RAb

[Ano12s] Anonymous. Reviewer acknowledgment. Graphical Models, 74(6):374, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000641.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial board. Computer Vision and Image Understanding: CVIU, 117(1):??, January 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001567.

Anonymous:2013:EBh

[Ano13b] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 117(2):C2, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001993.

Anonymous:2013:EBb

[Ano13c] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 117(2):C2, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001993.

Anonymous:2013:EBi

[Ano13d] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 117 (3):??, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000064.

Anonymous:2013:EBc

[Ano13e] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 117 (3):C2, March 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000064.

Anonymous:2013:EBj

[Ano13f] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 117 (4):C2, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000180.

Anonymous:2013:EBd

[Ano13g] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 117 (4):C2, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000180.

Anonymous:2013:EBe

[Ano13h] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 117(5):C2, May 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000398.

Anonymous:2013:EBl

[Ano13i] Anonymous. Editorial Board. Computer Vision and Image Understanding: CVIU, 117(6):C2, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421300057X.

Anonymous:2013:EBn

[Ano13j] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 117(7):C2, July 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000817.

Anonymous:2013:EBo

[Ano13k] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 117(8):C2, August 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000982.

Anonymous:2013:EBp

[Ano13l] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 117(9):C2, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001112.

Anonymous:2013:EBq

[Ano13m] Anonymous. Editorial Board.

Computer Vision and Image Understanding: CVIU, 117(10):C2, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213001434.

Anonymous:2013:EBr

[Ano13n] Anonymous. Editorial board. Computer Vision and Image Understanding: CVIU, 117(11):C2, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421300163X.

Anonymous:2013:EBf

[Ano13o] Anonymous. Editorial Board. Graphical Models, 75(1):i– ii, January 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000040.

Anonymous:2013:EBg

[Ano13p] Anonymous. Editorial Board. Graphical Models, 75(2):i– ii, March 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000106.

Anonymous:2013:EBm

[Ano13q] Anonymous. Editorial Board. Graphical Models, 75(3): i–ii, May 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031300009X.

Anonymous:2013:MCP

[Ano13r] Anonymous. Most cited paper award. Computer Vision and Image Understanding: CVIU, 117(5):571, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000441.

Anonymous:2013:PNa

[Ano13s]  Anonymous.  Publisher’s

note. Graphical Models, 75 (1):1, January 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070313000052.

Anonymous:2013:RAa

[Ano13t] Anonymous. Reviewer acknowledgment. Computer Vision and Image Understanding: CVIU, 117(1):??, January 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001543.

Anonymous:2013:RAb

[Ano13u] Anonymous. Reviewer acknowledgment. Graphical Models, 75(6):371, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070313000568.

Anonymous:2013:SCI

[Ano13v] Anonymous. Supplementary content. Graphical Models, 75(2):??, March 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031200077X#
MMCvFirst.

Antonisse:1982:ISP

[Ant82]  H. J. Antonisse.  Image

segmentation in pyramids. Computer Graphics and Image Processing, 19(4):367– 383, August 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Antonacopoulos:1998:PSU

[Ant98] Apostolos Antonacopoulos. Page segmentation using the description of the background. Computer Vision and Image Understanding: CVIU, 70(3):350–369, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0691/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0691/production/
pdf;   http://www.idealibrary.

com/links/artid/cviu.1998.
0691/production/ref.

Azencot:2003:DSS

[AO03] Jacques Azencot and Maciej Orkisz. Deterministic and stochastic state model of right generalized cylinder (RGC-sm): application in computer phantoms synthesis. Graphical Models, 65(6):323– 350, November 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Arentz:2004:COS

[AO04] Will Archer Arentz and Bjørn Olstad. Classifying offensive sites based on image content. Computer Vision and Image Understanding: CVIU, 94(1– 3):295–310, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aw:1994:CDF

[AOR94] Y. K. Aw, R. Owens, and J. Ross. A catalog of 1-D features in natural images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(2):173–181, March 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1016/production;   http:
//www.idealibrary.com/links/

artid/cgip.1994.1016/production/
pdf.

Argyriou:2010:PSA

[APB10] Vasileios Argyriou, Maria Petrou, and Svetlana Barsky. Photometric stereo with an arbitrary number of illuminants. Computer Vision and Image Understanding: CVIU, 114(8):887–900, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Androutsos:1999:NVB

[APV99] D. Androutsos, K. N. Plataniotis, and A. N. Venetsanopoulos. A novel vectorbased approach to color image retrieval using a vector angular-based distance measure. Computer Vision and Image Understanding: CVIU, 75(1–2):46–58, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0767/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0767/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0767/production/ref.

Alata:2009:TBC

[AQ09] Olivier Alata and Ludovic Quintard. Is there a best color space for color image characterization or representation based on Multivariate Gaussian Mixture Model? Computer Vision and Image Understanding: CVIU, 113(8): 867–877, August 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Akin:1977:KAI

[AR77] Omer Akin and Raj Reddy.

Knowledge acquisition for image understanding research. Computer Graphics and Image Processing, 6(4):307– 334, August 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ayala-Raggi:2011:AFI

[ARARCE11] Salvador E. Ayala-Raggi, Leopoldo Altamirano-Robles, and Janeth Cruz-Enriquez. Automatic face interpretation using fast 3 D illuminationbased AAM models. Computer Vision and Image Understanding: CVIU, 115(2): 194–210, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arcelli:1981:PTC

[Arc81] Carlo Arcelli. Pattern thinning by contour tracing. Computer Graphics and Image Processing, 17(2):130– 144, October 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Artzy:1979:DTD

[Art79] E. Artzy. Display of threedimensional information in computed tomography. Computer Graphics and Image Processing, 9(2):196– 198, February 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Abel:1983:DSA

[AS83] D. J. Abel and J. L. Smith. Data structure and algorithm based on a linear key for a rectangle retrieval problem. Computer Vision, Graphics, and Image Processing, 24(1):1–13, October 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Arcelli:1988:FLM

[AS88] Carlo Arcelli and Gabriella Sanniti Di Baja. Finding local maxima in a pseudo-Euclidean distance transform. Computer Vision, Graphics, and Image Processing, 43 (3):361–367, September 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Antzoulatos:1993:HAA

[AS93a] Daniel G. Antzoulatos and Alexander A. Sawchuk. Hypermatrix algebra: Applications in parallel image processing. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):42–62, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1003/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1003/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1003/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1003/production/
pdf.

Antzoulatos:1993:HAT

[AS93b] Daniel G. Antzoulatos and Alexander A. Sawchuk. Hypermatrix algebra: Theory. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):24–41, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1002/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1002/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1002/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1002/production/
pdf.

Alexiadis:2008:NDS

[AS08a] Dimitrios S. Alexiadis and George D. Sergiadis. Narrow directional steerable filters in motion estimation. Computer Vision and Image Understanding: CVIU, 110(2): 192–211, May 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Aouadi:2008:APR

[AS08b] Souha Aouadi and Laurent Sarry. Accurate and precise 2D–3D registration based on X-ray intensity. Computer Vision and Image Understanding: CVIU, 110(1):134–151, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Alexiadis:2009:MES

[AS09] Dimitrios S. Alexiadis and George D. Sergiadis. Motion estimation, segmentation and separation, using hypercomplex phase correlation, clustering techniques and graphbased optimization. Computer Vision and Image Understanding: CVIU, 113(2): 212–234, February 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ashraf:2013:VIA

[ASCF13] Nazim Ashraf, Yuping Shen, Xiaochun Cao, and Hassan Foroosh. View invariant action recognition using weighted fundamental ratios. Computer Vision and Image Understanding: CVIU, 117(6):587–602, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000131.

Audette:2003:IRS

[ASFP03] Michel A. Audette, Kaleem Siddiqi, Frank P. Ferrie, and Terry M. Peters. An integrated range-sensing, segmentation and registration framework for the characterization of intra-surgical brain deformations in imageguided surgery. Computer Vision and Image Understanding: CVIU, 89(2–3):226–251, February/March 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Agrawal:1997:PTP

[ASS97] R. C. Agrawal, S. C. Sahasrabudhe, and R. K. Shevgaonkar. Preservation of topological properties of a simple closed curve under digitalization. Computer Vision and Image Understanding: CVIU, 67(2):99–111, August 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0514/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0514/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/cviu.1996.
                                        0514/production/ref.

Aastrom:1997:GVS

[Åst97] Kalle Åström. The geometry of visual space: About the incompatibility between science and mathematics — reply. Computer Vision and Image Understanding: CVIU, 65(3):436–438, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0494/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0494/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0494/production/ref.

Alvarez:2012:LDC

[ASVO12] Susana Alvarez, Anna Salvatella, Maria Vanrell, and Xavier Otazu. Low-dimensional and comprehensive color texture description. Computer Vision and Image Understanding: CVIU, 116 (1):54–67, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421100186X.

August:1999:CFG

[ASZ99a] Jonas August, Kaleem Siddiqi, and Steven W. Zucker. Contour fragment grouping and shared, simple occluders. Computer Vision and Image Understanding: CVIU, 76(2):146–162, November 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0795/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0795/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0795/production/ref.

August:1999:LIP

[ASZ99b] Jonas August, Kaleem Siddiqi, and Steven W. Zucker. Ligature instabilities in the perceptual organization of shape. Computer Vision and Image Understanding: CVIU, 76(3):231–243, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0802/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0802/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0802/production/ref.

Asada:1983:RTD

[AT83] Minoru Asada and Saburo Tsuji. Representation of three-dimensional motion in dynamic scenes. Computer Vision, Graphics, and Image Processing, 21(1):118– 144, January 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ahuja:1989:EEP

[AT89] Narendra Ahuja and Mihran Tuceryan. Extraction of early perceptual structure in dot patterns. integrating region, boundary, and component Gestalt. Computer Vision, Graphics, and Image Processing, 48(3):304– 356, December 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Andreopoulos:2013:YOR

[AT13] Alexander Andreopoulos and John K. Tsotsos. 50 years of object recognition: Directions forward. Computer Vision and Image Understanding: CVIU, 117(8):827–891, August 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421300091X.

Avila:2013:PIR

[ATC^{+}13] Sandra Avila, Nicolas Thome, Matthieu Cord, Eduardo Valle, and Arnaldo de A. Araújo. Pooling in image representation: the visual codeword point of view. Computer Vision and Image Understanding: CVIU, 117(5):453–465, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001737.

Agui:1983:APF

[ATN83] Takeshi Agui, Masahiro Takeda, and Masayuki Nakajima. Animating planar folds by computer. Computer Vision, Graphics, and Image Processing, 24(2):244– 254, November 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Alahi:2010:CDD

[AVBK10] Alexandre Alahi, Pierre Vandergheynst, Michel Bierlaire, and Murat Kunt. Cascade of descriptors to detect and track objects across any network of cameras. Computer Vision and Image Understanding: CVIU, 114(6): 624–640, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Austin:1998:OLP

[AW98] W. J. Austin and A. M.

Wallace. Object location by parallel pose clustering. Computer Vision and Image Understanding: CVIU, 72(3):304–327, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0672/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1997.0672/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0672/production/ref.

Allan:2009:OLU

[AW09] Moray Allan and Christopher K. I. Williams. Object localisation using the Generative Template of Features. Computer Vision and Image Understanding: CVIU, 113(7): 824–838, July 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Angelidis:2006:SSD

[AWC06] Alexis Angelidis, Geoff Wyvill, and Marie-Paule Cani. Sweepers: Swept deformation defined by gesture. Graphical Models, 68(1):2–14, January 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S152407030500055X.

Aner-Wolf:2004:VSC

[AWK04] Aya Aner-Wolf and John R. Kender. Video summaries and cross-referencing through mosaic-based representation. Computer Vision and Image Understanding: CVIU, 95(2): 201–237, August 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Arbab-Zavar:2011:GMB

[AZN11] Banafshe Arbab-Zavar and Mark S. Nixon. On guided model-based analysis for ear biometrics. Computer Vision and Image Understanding: CVIU, 115(4):487–502, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Amir:2005:ESE

[AZSVK05] Arnon Amir, Lior Zimet, Alberto Sangiovanni-Vincentelli, and Sean Kao. An embedded system for an eye-detection sensor. Computer Vision and Image Understanding: CVIU, 98(1):104–123, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Blostein:1989:MRD

[BA89] Dorothea Blostein and Narendra Ahuja. A multiscale region detector. Computer Vision, Graphics, and Image Processing, 45(1):22–41, January 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Brandt:1992:CSC

[BA92] Jonathan W. Brandt and V. Ralph Algazi. Continuous skeleton computation by Voronoi diagram. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(3):329–338, May 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Black:1996:REM

[BA96] Michael J. Black and P. Anandan. The robust estimation of multiple motions: Parametric and piecewise-smooth flow fields. Computer Vision and Image Understanding: CVIU, 63(1):75–104, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0006/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0006/production/
pdf.

Barreto:2006:FCP

[BA06] João P. Barreto and Helder Araujo. Fitting conics to paracatadioptric projections of lines. Computer Vision and Image Understanding: CVIU, 101(3):151–165, March 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bretto:1997:CIP

[BACL97] A. Bretto, J. Azema, H. Cherifi, and B. Laget. Combinatorics and image processing. Graphical Models and Image Processing: GMIP, 59(5):265–277, September 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0437/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0437/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0437/production/ref.

Badler:1977:DGR

[Bad77] N. I. Badler. Disk generators for a raster display device. Computer Graphics and Image Processing, 6(??): 589–593, ???? 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Baird:1988:FIH

[Bai88] Henry S. Baird. Feature identification for hybrid structural/statistical pattern classification. Computer Vision, Graphics, and Image Processing, 42(3):318–333, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ben-Arie:1987:ORO

[BAM87] Jezekiel Ben-Arie and A. Zvi Meiri. 3D objects recognition by optimal matching search of multinary relations graphs. Computer Vision, Graphics, and Image Processing, 37(3):345–361, March 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Brooks:2008:ASM

[BAP08] Rupert Brooks, Tal Arbel, and Doina Precup. Anytime similarity measures for faster alignment. Computer Vision and Image Understanding: CVIU, 110(3):378–389, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barsky:1984:EPM

[Bar84] B. A. Barsky. Exponential and polynomial methods for applying tension to an interpolating spline curve. Computer Vision, Graphics, and Image Processing, 27(1):1–18, July 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Barnard:1985:CBP

[Bar85] Stephen T. Barnard. Choosing a basis for perceptual space. Computer Vision, Graphics, and Image Processing, 29(1):87–99, January 1985. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bartoli:2005:GDS

[Bar05] Adrien Bartoli. The geometry of dynamic scenes—on coplanar and convergent linear motions embedded in 3D static scenes. Computer Vision and Image Understanding: CVIU, 98(2):223–238, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barreto:2006:UGR

[Bar06] João P. Barreto. A unifying geometric representation for central projection systems. Computer Vision and Image Understanding: CVIU, 103 (3):208–217, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bartoli:2007:RSS

[Bar07] Adrien Bartoli. A random sampling strategy for piecewise planar scene segmentation. Computer Vision and Image Understanding: CVIU, 105(1):42–59, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bass:1981:UVL

[Bas81] Daniel H. Bass. Using the video lookup table for reflectivity calculations: Specific techniques and graphic results. Computer Graphics and Image Processing, 17(3): 249–261, November 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Bates:1984:UST

[Bat84] R. H. T. Bates. Uniqueness of solutions to two-dimensional Fourier phase problems for localized and positive images.

Computer Vision, Graphics, and Image Processing, 25 (2):205–217, February 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bauer:1985:SOL

[Bau85] Michael A. Bauer. Set operations on linear quadtrees. Computer Vision, Graphics, and Image Processing, 29 (2):248–258, February 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Brill:1983:CFE

[BB83] Michael H. Brill and Eamon B. Barrett. Closedform extension of the anharmonic ratio to N-space. Computer Vision, Graphics, and Image Processing, 23(1): 92–98, July 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Basu:1987:AHE

[BB87] Anup Basu and Christopher M. Brown. Algorithms and hardware for efficient image smoothing. Computer Vision, Graphics, and Image Processing, 40(2):131– 146, November 1987. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bhanu:1988:ADF

[BB88] Bir Bhanu and Wilhelm Burger. Approximation of displacement fields using wavefront region growing. Computer Vision, Graphics, and Image Processing, 41(3):306– 322, March 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bhanu:1991:QAD

[BB91] Bir Bhanu and Wilhelm Burger. A qualitative approach to dynamic scene understanding. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(2):184–205, September 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Ballard:1992:PAV

[BB92] Dana H. Ballard and Christopher M. Brown. Principles of animate vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):3–21, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Bogoni:1995:IRR

[BB95] Luca Bogoni and Ruzena Bajcsy. Interactive recognition and representation of functionality. Computer Vision and Image Understanding: CVIU, 62(2):

194–214, September 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1050/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1050/production/
pdf.

Bergevin:2003:OLS

[BB03] R. Bergevin and A. Bubel. Object-level structured contour map extraction. Computer Vision and Image Understanding: CVIU, 91(3): 302–334, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bergevin:2004:DCJ

[BB04] R. Bergevin and A. Bubel. Detection and characterization of junctions in a 2D image. Computer Vision and Image Understanding: CVIU, 93(3):288–309, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Brimkov:2011:CMO

[BB11] Valentin E. Brimkov and Reneta P. Barneva. Computational modeling of objects represented in images. Graphical Models, 73(6):311– 312, November 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000208.

Bloch:2013:MMH

[BB13] Isabelle Bloch and Alain Bretto. Mathematical morphology on hypergraphs, application to similarity and positive kernel. Computer Vision and Image Understanding: CVIU, 117(4):342–354, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001828.

Bertin:1996:VPC

[BBB96] Etienne Bertin, Horst Bischof, and Pascal Bertolino. Voronoi pyramids controlled by Hopfield neural networks. Computer Vision and Image Understanding: CVIU, 63(3): 462–475, May 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1996.0035/production;
http://www.idealibrary.
com/links/artid/cviu.1996.
0035/production/pdf.

Brimkov:2011:CDB

[BBB11] Valentin E. Brimkov, Reneta P.

Barneva, and Boris Brimkov. Connected distance-based rasterization of objects in arbitrary dimension. Graphical Models, 73(6):323–334, November 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000191.

Baldoni:2000:UIC

[BBC00] Matteo Baldoni, Cristina Baroglio, and Davide Cavagnino. Use of IFS codes for learning 2D isolatedobject classification systems. Computer Vision and Image Understanding: CVIU, 77(3):371–387, March 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0823;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0823/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0823/ref.

Bertozzi:2007:PDM

[BBC^{+}07] M. Bertozzi, A. Broggi, C. Caraffi, M. Del Rose, M. Felisa, and G. Vezzoni. Pedestrian detection by means of far-infrared stereo vision. Computer Vision and Image Understanding: CVIU, 106(2–3):194–204, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Berman:1994:EBB

[BBD^{+}94] M. Berman, L. M. Bischof, S. J. Davies, A. A. Green, and M. Craig. Estimating band-to-band misregistrations in aliased imagery. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(6):479–493, November 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1043/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1043/production/
pdf.

Bardera:2011:MIR

[BBF^{+}11] Anton Bardera, Imma Boada, Miquel Feixas, Jaume Rigau, and Mateu Sbert. Multiresolution image registration based on tree data structures. Graphical Models, 73 (4):111–126, July 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031100004X.

Bellotto:2012:CVT

[BBH^{+}12] Nicola Bellotto, Ben Benfold, Hanno Harland, Hans-Hellmut Nagel, Nicola Pirlo, Ian Reid, Eric Sommerlade, and Chuan Zhao. Cognitive visual tracking and camera control. Computer Vision and Image Understanding: CVIU, 116(3):457–471, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421100213X.

Bogdanova:2010:DVA

[BBHF10] Iva Bogdanova, Alexandre Bur, Heinz Hügli, and Pierre-André Farine. Dynamic visual attention on the sphere. Computer Vision and Image Understanding: CVIU, 114(1): 100–110, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barnhill:1978:NTC

[BBK78] R. E. Barnhill, J. H. Brown, and I. M. Klucewicz. A new twist in computer aided geometric design. Computer Graphics and Image Processing, 8(1):78–91, August 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Biswas:1985:GDC

[BC85] S. N. Biswas and B. B. Chaudhuri. On the generation of discrete circular objects and their properties. Computer Vision, Graphics, and Image Processing, 32 (2):158–170, November 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bezdek:1988:CZS

[BC88a] James C. Bezdek and Er-Woon Chiou. Core zone scatterplots: a new approach to feature extraction for visual displays. Computer Vision, Graphics, and Image Processing, 41(2):186– 209, February 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bischof:1988:PSS

[BC88b] Walter F. Bischof and Terry Caelli. Parsing scale-space and spatial stability analysis. Computer Vision, Graphics, and Image Processing, 42(2):192–205, May 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bergholm:1991:TOF

[BC91] F. Bergholm and S. Carlsson. A theory of optical flow. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (2):171–188, March 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Bajcsy:1992:AEP

[BC92] Ruzena Bajcsy and Mario Campos. Active and exploratory perception. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):31–40, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Benoit:2010:FBI

[BC10] A. Benoit and A. Caplier. Fusing bio-inspired vision data for simplified high level scene interpretation: Application to face motion analysis. Computer Vision and Image Understanding: CVIU, 114(7): 774–789, July 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bardinet:1998:PDM

[BCA98] Eric Bardinet, Laurent D. Cohen, and Nicholas Ayache. A parametric deformable model to fit unstructured 3D data. Computer Vision and Image Understanding: CVIU, 71(1):39–54, July 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0595/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0595/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0595/production/ref.

Benoit:2010:UHV

[BCDH10] A. Benoit, A. Caplier, B. Durette, and J. Herault. Using Human Visual System modeling for bio-inspired low level image processing. Computer Vision and Image Understanding: CVIU, 114(7):758–773, July 2010. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Bowyer:2006:SAC

[BCF06] Kevin W. Bowyer, Kyong Chang, and Patrick Flynn. A survey of approaches and challenges in 3D and multimodal 3D + 2D face recognition. Computer Vision and Image Understanding: CVIU, 101(1):1–15, January 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bongiovanni:1995:PSA

[BCG95] G. Bongiovanni, P. Crescenzi, and C. Guerra. Parallel simulated annealing for shape detection. Computer Vision and Image Understanding: CVIU, 61(1):60–69, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1005/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1005/production/
pdf.

Baronti:1990:VPS

[BCL^{+}90] S. Baronti, A. Casini, F. Lotti, L. Favaro, and V. Roberto. Variable pyramid structures for image segmentation. Computer Vision, Graphics, and Image Processing, 49(3):346– 356, March 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bajaj:1996:ATS

[BCL96] Chandrajit L. Bajaj, Edward J. Coyle, and Kwun-Nan Lin. Arbitrary topology shape reconstruction from planar cross sections. Graphical Models and Image Processing: GMIP, 58 (6):524–543, November 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0044/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0044/production/
pdf.

Badler:1997:EI

[BCM97] Norman Badler, Ingrid Carlbom, and Dimitris Metaxas. Editorial introduction. Graphical Models and Image Processing: GMIP, 59(1):iii, January 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0415/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0415/production/
pdf.

Bicego:2006:USA

[BCM06] Manuele Bicego, Marco Cristani, and Vittorio Murino. Unsupervised scene analysis: a hidden Markov model approach.

Computer Vision and Image Understanding: CVIU, 102 (1):22–41, April 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bazzani:2013:SDA

[BCM13] Loris Bazzani, Marco Cristani, and Vittorio Murino. Symmetrydriven accumulation of local features for human characterization and re-identification. Computer Vision and Image Understanding: CVIU, 117 (2):130–144, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001464.

Bernier:2009:FNB

[BCMCB09] Olivier Bernier, Pascal Cheung-Mon-Chan, and Arnaud Bouguet. Fast nonparametric belief propagation for realtime stereo articulated body tracking. Computer Vision and Image Understanding: CVIU, 113(1):29–47, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barth:1993:IEL

[BCZ93] Erhardt Barth, Terry Caelli, and Christoph Zetzsche. Image encoding, labeling, and reconstruction from differential geometry. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(6):428–446, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1033/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1033/production/
pdf.

Bertrand:1994:ASM

[BD94a] Yves Bertrand and Jean-Francois F. Dufourd. Algebraic specification of a 3Dmodeler based on hypermaps. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):29–60, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1005/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1005/production/
pdf.

Bhatt:1994:RIR

[BD94b] M. R. Bhatt and U. B. Desai. Robust image restoration algorithm using Markov random field model. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):61–74, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1006/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1006/production/
pdf.

Borgefors:1996:ANP

[Bd96] Gunilla Borgefors and Gabriella Sanniti di Baja. Analyzing nonconvex 2D and 3D patterns. Computer Vision and Image Understanding: CVIU, 63(1):145–157, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0010/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0010/production/
pdf.

Buluswar:2002:CMO

[BD02] Shashi D. Buluswar and Bruce A. Draper. Color models for outdoor machine vision. Computer Vision and Image Understanding: CVIU, 85(2):71–99, February 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bose:2009:CPT

[BDHM09] Prosenjit Bose, Vida Dujmović, Ferran Hurtado, and Pat Morin. Connectivitypreserving transformations of binary images. Computer Vision and Image Understanding: CVIU, 113(10):1027– 1038, October 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bleau:1992:NSFa

[BDL92a] Andre Bleau, Jacques De Guise, and A.-Robert LeBlanc. A new set of fast algorithms for mathematical morphology. I. idempotent geodesic transforms. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (2):178–209, September 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Bleau:1992:NSFb

[BDL92b] Andre Bleau, Jacques De Guise, and A.-Robert LeBlanc. A new set of fast algorithms for mathematical morphology. II. Identification of topographic features on grayscale images. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (2):210–229, September 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Briggs:2006:MSS

[BDL^{+}06] Amy J. Briggs, Carrick Detweiler, Yunpeng Li, Peter C. Mullen, and Daniel Scharstein. Matching scalespace features in 1D panoramas. Computer Vision and Image Understanding: CVIU, 103(3):184–195, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bertini:2012:MSR

[BDS12] Marco Bertini, Alberto Del Bimbo, and Lorenzo Seidenari. Multi-scale and real-time non-parametric approach for anomaly detection and localization. Computer Vision and Image Understanding: CVIU, 116 (3):320–329, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002104.

Bazin:2010:MED

[BDVK10] J. C. Bazin, C. Demonceaux, P. Vasseur, and I. S. Kweon. Motion estimation by decoupling rotation and translation in catadioptric vision. Computer Vision and Image Understanding: CVIU, 114(2): 254–273, February 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barton:2011:SFA

[BE11] Michael Barton and Gershon Elber. Spiral fat arcs — bounding regions with cubic convergence. Graphical Models, 73(2):50–57, March 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000214.

Beck:1985:PTM

[Bec85] Jacob Beck. Perception of transparency in man and machine. Computer Vision, Graphics, and Image Processing, 31(2):127–138, August 1985. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Beck:1986:RPT

[Bec86] Jacob Beck. Reply: The perception of transparency: a reply to Brill. Computer Vision, Graphics, and Image Processing, 35(2):272– 273, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bonev:2013:ITS

[BEGB13] Boyan Bonev, Francisco Escolano, Daniela Giorgi, and Silvia Biasotti. Informationtheoretic selection of highdimensional spectral features for structural recognition. Computer Vision and Image Understanding: CVIU, 117(3):214–228, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001919.

Bengtsson:1981:SCC

[BEH^{+}81] E. Bengtsson, O. Eriksson, J. Holmquist, T. Jarkrans, B. Nordin, and B. Stenkvist. Segmentation of cervical cells: Detection of overlapping cell nuclei. Computer Graphics and Image Processing, 16(4): 382–394, August 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ben-Ezra:2000:RTM

[BEPW00] Moshe Ben-Ezra, Shmuel Peleg, and Michael Werman. Real-time motion analysis with linear programming. Computer Vision and Image Understanding: CVIU, 78(1):32–52, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0826;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0826/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0826/ref.

Berzins:1984:ALE

[Ber84] Valdis Berzins. Accuracy of Laplacian edge detectors. Computer Vision, Graphics, and Image Processing, 27 (2):195–210, August 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Berman:1989:LSB

[Ber89] Mark Berman. Large sample bias in least squares estimators of a circular arc center and its radius. Computer Vision, Graphics, and Image Processing, 45(1):126– 128, January 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bay:2008:SRF

[BETV08] Herbert Bay, Andreas Ess, Tinne Tuytelaars, and Luc Van Gool. Speeded-up Robust Features (SURF). Computer Vision and Image Understanding: CVIU, 110(3): 346–359, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Beusmans:1991:COC

[Beu91] J. M. H. Beusmans. Computing occluding contours using spherical images. Computer Vision, Graphics, and Image Processing. Image Understanding, 53(1):97–111, January 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Bevacqua:1987:SSL

[BF87] Giuseppe Bevacqua and Ruhal Floris. A surface specificline tracking and slope recognition algorithm. Computer Vision, Graphics, and Image Processing, 40(2):219– 227, November 1987. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Breckon:2005:AVC

[BF05] Toby P. Breckon and Robert B.

Fisher. Amodal volume completion: 3D visual completion. Computer Vision and Image Understanding: CVIU, 99 (3):499–526, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Benoit:2007:TDR

[BF07] Stephen Benoit and Frank P.

Ferrie. Towards direct recovery of shape and motion parameters from image sequences. Computer Vision and Image Understanding: CVIU, 105(2):145– 165, February 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Butakoff:2010:MVF

[BF10] Constantine Butakoff and Alejandro F. Frangi. Multiview face segmentation using fusion of statistical shape and appearance models. Computer Vision and Image Understanding: CVIU, 114(3): 311–321, March 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barnard:1997:CCS

[BFF97] Kobus Barnard, Graham Finlayson, and Brian Funt. Color constancy for scenes with varying illumination. Computer Vision and Image Understanding: CVIU, 65 (2):311–321, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0567/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0567/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0567/production/ref.

Battaglino:2013:DTH

[BFR13] D. Battaglino, A. Frosini, and S. Rinaldi. A decomposition theorem for homogeneous sets with respect to diamond probes. Computer Vision and Image Understanding: CVIU, 117(4):319–325, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001804.

Boschiroli:2012:RBI

[BFRA12] Maria Boschiroli, Christoph Fünfzig, Lucia Romani, and Gudrun Albrecht. G^{1} rational blend interpolatory schemes: a comparative study. Graphical Models, 74(1):29–49, January 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031100066X.

Black:2000:REC

[BFY00] Michael J. Black, David J.

Fleet, and Yaser Yacoob. Robustly estimating changes in image appearance. Computer Vision and Image Understanding: CVIU, 78(1):8–31, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0825;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0825/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0825/ref.

Barsky:1979:DSB

[BG79] B. Barsky and D. Greenberg. Determining a set of B-spline control vertices to generate an interpolating surface. Computer Graphics and Image Processing, 11(??): 203–209, ???? 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Barsky:1980:DSB

[BG80] Brian A. Barsky and Donald P. Greenberg. Determining a set of B-spline control vertices to generate an interpolating surface. Computer Graphics and Image Processing, 14(3):203– 226, November 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Barry:1991:IAC

[BG91] P. J. Barry and R. N. Goldman. Interpolation and approximation of curves and surfaces using Polya polynomials. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2):137–148, March 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Burghouts:2009:PEL

[BG09] Gertjan J. Burghouts and Jan-Mark Geusebroek. Performance evaluation of local colour invariants. Computer Vision and Image Understanding: CVIU, 113(1): 48–62, January 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barbier:2005:FMA

[BGA05] Aurélien Barbier, Eric Galin, and Samir Akkouche. A framework for modeling, animating, and morphing textured implicit models. Graphical Models, 67(3):166–188, May 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Brechbuhler:1995:PCS

[BGK95] Ch. Brechbühler, G. Gerig, and O. Kübler. Parametrization of closed surfaces for 3-D shape description. Computer Vision and Image Understanding: CVIU, 61 (2):154–170, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1013/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1013/production/
pdf.

Bober:1998:ARE

[BGK98] M. Bober, N. Georgis, and J. Kittler. On accurate and robust estimation of fundamental matrix. Computer Vision and Image Understanding: CVIU, 72(1):39–53, October 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0670/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0670/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0670/production/ref.

Barequet:2004:CIS

[BGLSS04] Gill Barequet, Michael T.

Goodrich, Aya Levi-Steiner, and Dvir Steiner. Contour interpolation by straight skeletons. Graphical Models, 66 (4):245–260, July 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Beveridge:2009:FIA

[BGPD09] J. Ross Beveridge, Geof H. Givens, P. Jonathon Phillips, and Bruce A. Draper. Factors that influence algorithm performance in the Face Recognition Grand Challenge. Computer Vision and Image Understanding: CVIU, 113(6): 750–762, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Boulanger:1989:ISS

[BGR89] Pierre Boulanger, André Gagalowicz, and Marc Rioux. Integration of synthetic surface relief in range images. Computer Vision, Graphics, and Image Processing, 47 (3):361–372, September 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Burch:1983:IRP

[BGS83] S. F. Burch, S. F. Gull, and J. Skilling. Image restoration by a powerful maximum entropy method. Computer Vision, Graphics, and Image Processing, 23(2):113– 128, August 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bebis:1998:IBA

[BGSdVL98] George Bebis, Michael Georgiopoulos, Mubarak Shah, and Niels da Vitoria Lobo. Indexing based on algebraic functions of views. Computer Vision and Image Understanding: CVIU, 72(3): 360–378, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0679/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0679/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0679/production/ref.

Bedini:1994:DAR

[BGT94] L. Bedini, I. Gerace, and A. Tonazzini. A deterministic algorithm for reconstructing images with interacting discontinuities. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(2):109–123, March 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1011/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1011/production/
pdf.

Bielser:2004:SMR

[BGTG04] D. Bielser, P. Glardon, M. Teschner, and M. Gross. A state machine for real-time cutting of tetrahedral meshes. Graphical Models, 66(6):398– 417, November 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Brady:1983:RSO

[BH83a] Michael Brady and Berthold K. P. Horn. Rotationally symmetric operators for surface interpolation. Computer Vision, Graphics, and Image Processing, 22(1):70– 94, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bridwell:1983:DSR

[BH83b] Nelson J. Bridwell and Thomas S. Huang. Discrete spatial representation for lateral motion stereo. Computer Vision, Graphics, and Image Processing, 21(1):33–57, January 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bruss:1983:PN

[BH83c] Anna R. Bruss and Berthold K. P. Horn. Passive navigation. Computer Vision, Graphics, and Image Processing, 21(1):3–20, January 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ballard:1986:TFP

[BH86] Dana H. Ballard and Leo Hartman. Task frames: Primitives for sensory-motor coordination. Computer Vision, Graphics, and Image Processing, 36(2/3):274–297, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Burel:1995:DOO

[BH95] Gilles Burel and Hugues Henoco. Determination of the orientation of 3D objects using spherical harmonics. Graphical Models and Image Processing: GMIP, 57(5):400–408, September 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1034/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1034/production/
pdf.

Berthilsson:1999:RPO

[BH99] Rikard Berthilsson and Anders Heyden. Recognition of planar objects using the density of affine shape. Computer Vision and Image Understanding: CVIU, 76(2):135–145, November 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.

                                        0781/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0781/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0781/production/ref.

Bui:2012:ECB

[BH12] T. T. Quyen Bui and Keum-Shik Hong. Evaluating a color-based active basis model for object recognition. Computer Vision and Image Understanding: CVIU, 116(11): 1111–1120, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421200104X.

Bhattacharya:1991:OAT

[Bha91] B. K. Bhattacharya. An optimal algorithm to translate a convex polyhedron through a two-dimensional convex window. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(3): 269–270, May 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Baker:2010:DIR

[BHBF10] Sarah E. Baker, Amanda Hentz, Kevin W. Bowyer, and Patrick J. Flynn. Degradation of iris recognition performance due to non-cosmetic prescription contact lenses. Computer Vision and Image Understanding: CVIU, 114(9):1030–1044, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bowyer:2008:IUI

[BHF08] Kevin W. Bowyer, Karen Hollingsworth, and Patrick J. Flynn. Image understanding for iris biometrics: a survey. Computer Vision and Image Understanding: CVIU, 110 (2):281–307, May 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bohme:2010:SCI

[BHMB10] Martin Böhme, Martin Haker, Thomas Martinetz, and Erhardt Barth. Shading constraint improves accuracy of time-of-flight measurements. Computer Vision and Image Understanding: CVIU, 114(12):1329–1335, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bruckstein:1993:ISP

[BHNR93] Alfred M. Bruckstein, Robert J.

Holt, Arun N. Netravali, and Thomas J. Richardson. Invariant signatures for planar shape recognition under partial occlusion. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(1):49–65, July 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1031/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1031/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1032/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1032/production/
pdf.

Ben-Hamadou:2013:FCS

[BHSD^{+}13] Achraf Ben-Hamadou, Charles Soussen, Christian Daul, Walter Blondel, and Didier Wolf. Flexible calibration of structured-light systems projecting point patterns. Computer Vision and Image Understanding: CVIU, 117(10): 1468–1481, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001069.

Banno:2010:OTB

[BI10] Atsuhiko Banno and Katsushi Ikeuchi. Omnidirectional texturing based on robust 3D registration through Euclidean reconstruction from two spherical images. Computer Vision and Image Understanding: CVIU, 114(4): 491–499, April 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Banno:2011:DMR

[BI11] Atsuhiko Banno and Katsushi Ikeuchi. Disparity map refinement and 3D surface smoothing via directed anisotropic diffusion. Computer Vision and Image Understanding: CVIU, 115(5):611–619, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bichsel:1998:ASP

[Bic98]  Martin Bichsel.  Analyz-

ing a scene’s picture set under varying lighting. Computer Vision and Image Understanding: CVIU, 71(3): 271–280, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0627/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0627/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0627/production/ref.

Bidasaria:1986:MAE

[Bid86] H. B. Bidasaria. A method for almost exact histogram matching for two digitized images. Computer Vision, Graphics, and Image Processing, 34(1):93–98, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bidasaria:1991:MRT

[Bid91] H. B. Bidasaria. A method for ray tracing a wide class of generalized cylinders with straight line trajectories. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2):101–107, March 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Bidasaria:1992:DRT

[Bid92] H. B. Bidasaria. Defining and rendering of textured objects through the use of exponential functions. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(2): 97–102, March 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Biederman:1985:HIU

[Bie85] Irving Biederman. Human image understanding: Recent research and a theory. Computer Vision, Graphics, and Image Processing, 32(1):29–73, October 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bieri:1987:CEC

[Bie87] H. Bieri. Computing the Euler characteristic and related additive functionals of digital objects from their bintree representation. Computer Vision, Graphics, and Image Processing, 40(1):115– 126, October 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bigun:1990:SFS

[Big90] Josef Bigun. A structure feature for some image processing applications based on spiral functions. Computer Vision, Graphics, and Image Processing, 51(2):166– 194, August 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bigun:1997:PRI

[Big97] Josef Bigün. Pattern recognition in images by symmetries and coordinate transformations. Computer Vision and Image Understanding: CVIU, 68(3):290–307, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0556/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0556/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0556/production/ref.

Bajaj:2000:CBT

[BIP00] Chandrajit Bajaj, Insung Ihm, and Sanghun Park. Compression-based 3D texture mapping for real-time rendering. Graphical Models, 62(6):391–410, November 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0532;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0532/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0532/ref.

Besl:1986:ISC

[BJ86] Paul J. Besl and Ramesh C. Jain. Invariant surface characteristics for 3D object recognition in range images. Computer Vision, Graphics, and Image Processing, 33(1):33– 80, January 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Breen:1996:AOT

[BJ96] Edmond J. Breen and Ronald Jones. Attribute openings, thinnings, and granulometries. Computer Vision and Image Understanding: CVIU, 64(3):377–389, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0066/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0066/production/
pdf.

Basri:1997:CS

[BJ97] Ronen Basri and David Jacobs. Constancy and similarity. Computer Vision and Image Understanding: CVIU, 65(3):447–449, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0497/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0497/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0497/production/ref.

Ballard:1983:RBM

[BK83] Dana H. Ballard and O. A.

Kimball. Rigid body motion from depth and optical flow. Computer Vision, Graphics, and Image Processing, 22(1):95–115, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See note [HZ86].

Bajcsy:1989:MEM

[BK89] R. Bajcsy and S. Kovacic.

Multiresolution elastic matching. Computer Vision, Graphics, and Image Processing, 46 (1):1–21, April 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Barron:2001:EAP

[BK01] Carlos Barrón and Ioannis A. Kakadiaris. Estimating anthropometry and pose from a single uncalibrated image. Computer Vision and Image Understanding: CVIU, 81(3):269–284, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0888;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0888/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0888/ref.

Brun:2003:RFW

[BK03] Luc Brun and Walter Kropatsch. Receptive fields within the Combinatorial Pyramid framework. Graphical Models, 65 (1–3):23–42, May 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Bassiou:2007:CIH

[BK07] Nikoletta Bassiou and Constantine Kotropoulos. Color image histogram equalization by absolute discounting back-off. Computer Vision and Image Understanding: CVIU, 107(1–2):108– 122, July/August 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Burton:1984:IEP

[BKA84] F. Warren Burton, John G. Kollias, and Nikitas A. Alexandridis. An implementation of the exponential pyramid data structure with application to determination of symmetries in pictures. Computer Vision, Graphics, and Image Processing, 25 (2):218–225, February 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bowyer:2001:EDE

[BKD01] Kevin Bowyer, Christine Kranenburg, and Sean Dougherty. Edge detector evaluation using empirical ROC curves. Computer Vision and Image Understanding: CVIU, 84(1):77–103, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0931;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0931/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0931/ref.

Bastys:2011:IRF

[BKK11] Algirdas Bastys, Justas Kranauskas, and Volker Krüger. Iris recognition by fusing different representations of multiscale Taylor expansion. Computer Vision and Image Understanding: CVIU, 115(6): 804–816, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Berenstein:1987:GAS

[BKLO87] Carlos A. Berenstein, Laveen N. Kanal, David Lavine, and Eric C. Olson. A geometric approach to subpixel registration accuracy. Computer Vision, Graphics, and Image Processing, 40 (3):334–360, December 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bayer:1998:CPD

[BKMSR98] T. Bayer, U. Kressel, H. Mogg-Schneider, and I. Renz. Categorizing paper documents. A generic system for domain and language independent text categorization. Computer Vision and Image Understanding: CVIU, 70(3): 299–??, ???? 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bray:2007:SPF

[BKMV07] M. Bray, E. Koller-Meier, and L. Van Gool. Smart particle filtering for highdimensional tracking. Computer Vision and Image Understanding: CVIU, 106(1): 116–129, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Baker:2010:USA

[BKP10] Patrick Baker and Behrooz Kamgar-Parsi. Using shorelines for autonomous air vehicle guidance. Computer Vision and Image Understanding: CVIU, 114(6):723–729, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bai:1989:RLV

[BKR^{+}89] Z. D. Bai, P. R. Krishnaiah, C. R. Rao, P. S. Reddy, Y. N. Sun, and L. C. Zhao. Reconstruction of the left ventricle from two orthogonal projections. Computer Vision, Graphics, and Image Processing, 47 (2):165–188, August 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Belcastro:1996:TRP

[BKW96] Lori Belcastro, William C.

Karl, and Alan S. Willsky. Tomographic reconstruction of polygons from knot location and chord length measurements. Graphical Models and Image Processing: GMIP, 58(3):233–245, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0020/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0020/production/
pdf.

Bajcsy:1976:TGD

[BL76] Ruzena Bajcsy and Lawrence Lieberman. Texture gradient as a depth cue. Computer Graphics and Image Processing, 5(1):52–67, March 1976. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). BAJESY76a.

Beghdadi:1989:CET

[BL89] Azeddine Beghdadi and Alain Le Négrate. Contrast enhancement technique based on local detection of edges. Computer Vision, Graphics, and Image Processing, 46(2):162–174, May 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bergevin:1992:PDO

[BL92] Robert Bergevin and Martin D. Levine. Part decomposition of objects from single view line drawings. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(1):73–83, January 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Barrat:1994:RWT

[BL94] M. Barrat and O. Lepetit. Recursive wavelet transform for 2D signals. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):106–108, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1010/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1010/production/
pdf.

Bolduc:1998:RBM

[BL98a] Marc Bolduc and Martin D. Levine. A review of biologically motivated space-variant data reduction models for robotic vision. Computer Vision and Image Understanding: CVIU, 69(2):170–184, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0560/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0560/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0560/production/ref.

Bretzner:1998:FTA

[BL98b] Lars Bretzner and Tony Lindeberg. Feature tracking with automatic selection of spatial scales. Computer Vision and Image Understanding: CVIU, 71(3):385–392, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0650/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0650/production/

                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0650/production/ref.

Bleau:2000:WBS

[BL00] André Bleau and L. Joshua Leon. Watershed-based segmentation and region merging. Computer Vision and Image Understanding: CVIU, 77(3):317–370, March 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0822;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.1999.0822/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0822/ref.

Bottino:2001:SBT

[BL01] Andrea Bottino and Aldo Laurentini. A silhouette based technique for the reconstruction of human movement. Computer Vision and Image Understanding: CVIU, 83(1):79–95, July 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0918;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0918/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0918/ref.

Berwick:2004:SGC

[BL04] Daniel Berwick and Sang Wook Lee. Spectral gradients for color-based object recognition and indexing. Computer Vision and Image Understanding: CVIU, 94(1–3): 28–43, April/June 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bottino:2008:VHP

[BL08] Andrea Bottino and Aldo Laurentini. The visual hull of piecewise smooth objects. Computer Vision and Image Understanding: CVIU, 110 (1):7–18, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bhandarkar:2009:IDT

[BL09] Suchendra M. Bhandarkar and Xingzhi Luo. Integrated detection and tracking of multiple faces using particle filtering and optical flowbased elastic matching. Computer Vision and Image Understanding: CVIU, 113(6): 708–725, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Blake:1985:BCL

[Bla85] Andrew Blake. Boundary conditions for lightness computation in Mondrian world. Computer Vision, Graphics, and Image Processing, 32 (3):314–327, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Beghdadi:1995:ETU

[BLd95] Azeddine Beghdadi, Alain Le Négrate, and Patrick Viaris de Lesegno. Entropic thresholding using a block source model. Graphical Models and Image Processing: GMIP, 57(3):197–205, May 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1019/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1019/production/
pdf.

Bley:1984:SPE

[Ble84] Heinrich Bley. Segmentation and preprocessing of electrical schematics using picture graphs. Computer Vision, Graphics, and Image Processing, 28(3):271– 288, December 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Brzakovic:1991:SMB

[BLH91] D. Brzakovic, A. Liakopoulos, and L. Hong. Spline models for boundary detection description: formulation and performance evaluation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):392–401, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Bergevin:1995:RRV

[BLP95] Robert Bergevin, Denis Laurendeau, and Denis Poussart. Registering range views of multipart objects. Computer Vision and Image Understanding: CVIU, 61(1):1–16, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1001/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1001/production/
pdf.

Baumberg:2005:MCS

[BLT05] Adam Baumberg, Alex Lyons, and Richard Taylor. 3D S.O.M. —- a commercial software solution to 3D scanning. Graphical Models, 67 (6):476–495, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407030500010X.

Berthod:1979:LSR

[BM79] M. Berthod and J. P. Maroy.

Learning in syntactic recognition of symbols drawn on a graphic tablet. Computer Graphics and Image Processing, 9(2):166–182, February 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Braccini:1980:FGM

[BM80] Carlo Braccini and Giuseppe Marino. Fast geometrical manipulations of digital images. Computer Graphics and Image Processing, 13(2): 127–141, June 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Bovik:1986:EDU

[BM86] Alan C. Bovik and David C. Munson, Jr. Edge detection using median comparisons. Computer Vision, Graphics, and Image Processing, 33(3):377–389, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bertrand:1995:NBS

[BM95] Gilles Bertrand and Grégoire Malandain. A note on “Building Skeleton Models via 3-D Medial Surface/ Axis Thinning Algorithms”. Graphical Models and Image Processing: GMIP, 57 (6):537–538, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1045/production;   http:

//www.idealibrary.com/links/
artid/gmip.1995.1045/production/

pdf. See [LKC94].

Brunelli:1996:SII

[BM96] R. Brunelli and O. Mich.

SpotIt! an interactive identikit system. Graphical Models and Image Processing: GMIP, 58(5):399– 404, September 1996. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0033/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0033/production/
pdf.

Bricault:1997:VMI

[BM97] Ivan Bricault and Olivier Monga. From volume medical images to quadratic surface patches. Computer Vision and Image Understanding: CVIU, 67(1):24–38, July 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0501/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0501/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0501/production/ref.

Bergevin:1998:MCS

[BM98] Robert Bergevin and Marielle Mokhtari. Multiscale contour segmentation and approximation: An algorithm based on the geometry of regular inscribed polygons. Computer Vision and Image Understanding: CVIU, 71(1):55–73, July 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0634/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0634/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0634/production/ref.

Baillard:1999:RUS

[BM99] C. Baillard and H. Maître. 3D reconstruction of urban scenes from aerial stereo imagery: a focusing strategy. Computer Vision and Image Understanding: CVIU, 76 (3):244–258, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0793/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0793/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0793/production/ref.

Boxer:2000:ECE

[BM00] Laurence Boxer and Russ Miller. Efficient computation of the Euclidean distance transform. Computer Vision and Image Understanding: CVIU, 80(3):379–383, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0880;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0880/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0880/ref. See corrigen-

dum [BM02].

Boxer:2002:CEC

[BM02] Laurence Boxer and Russ Miller. Corrigendum to “Efficient Computation of the Euclidean Distance Transform”. Computer Vision and Image Understanding: CVIU, 86(2): 137–140, May 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [BM00].

Braude:2007:CBS

[BMM^{+}07] Ilya Braude, Jeffrey Marker, Ken Museth, Jonathan Nissanov, and David Breen. Contour-based surface reconstruction using MPU implicit models. Graphical Models, 69(2):139–157, March 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000683.

Banerjee:1991:PCN

[BMR91] S. Banerjee, D. Mount, and A. Rosenfeld. Pyramid computation of neighbor distance statistics in dot patterns. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):373–381, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Biermann:2002:SFM

[BMZB02] Henning Biermann, Ioana M. Martin, Denis Zorin, and Fausto Bernardini. Sharp features on multiresolution subdivision surfaces. Graphical Models, 64(2):61–77, March 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Bieri:1984:AEC

[BN84] H. Bieri and W. Nef. Algorithms for the Euler characteristic and related additive functionals of digital objects. Computer Vision, Graphics, and Image Processing, 28 (2):166–175, November 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Budrikis:1985:NLR

[BN85] Zigmantas L. Budrikis and Arun N. Netravali. New lightpen for raster displays. Computer Vision, Graphics, and Image Processing, 32 (3):359–383, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bruckstein:1990:MET

[BN90] Alfred M. Bruckstein and Arun N. Netravali. On minimal energy trajectories. Computer Vision, Graphics, and Image Processing, 49(3):283–296, March 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Braga-Neto:2002:CCL

[BNG02] Ulisses Braga-Neto and John Goutsias. Connectivity on complete lattices: New results. Computer Vision and Image Understanding: CVIU, 85(1):22–53, January 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Braga-Neto:2003:MAC

[BNG03] Ulisses Braga-Neto and John Goutsias. A multiscale approach to connectivity. Computer Vision and Image Understanding: CVIU, 89(1):70– 107, January 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Braga-Neto:2005:CMC

[BNG05] Ulisses Braga-Neto and John Goutsias. Constructing multiscale connectivities. Computer Vision and Image Understanding: CVIU, 99(1):

126–150, July 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Brookshire:1990:AS

[BNL90] G. Brookshire, M. Nadler, and Choon Lee. Automated stereophotogrammetry. Computer Vision, Graphics, and Image Processing, 52 (2):276–296, November 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Brillault-OMahoney:1991:NMV

[BO91] B. Brillault-O’Mahoney. New method for vanishing point detection. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (2):289–300, September 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Boissonnat:2005:PGS

[BO05] Jean-Daniel Boissonnat and Steve Oudot. Provably good sampling and meshing of surfaces. Graphical Models, 67 (5):405–451, September 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000056.

Boehm:1982:CS

[Boe82] Wolfgang Boehm. On cubics: a survey. Computer Graphics and Image Processing, 19 (3):201–226, July 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Bogomolny:1988:DGM

[Bog88] A. Bogomolny. Digital geometry may not be discrete. Computer Vision, Graphics, and Image Processing, 43 (2):205–220, August 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Boissonnat:1988:SRP

[Boi88] Jean-Daniel Boissonnat. Shape reconstruction from planar cross sections. Computer Vision, Graphics, and Image Processing, 44(1):1–29, October 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bookstein:1979:CGG

[Boo79a] F. L. Bookstein. Closing gaps, and gaps with a steppingstone, by means of parabolas. Computer Graphics and Image Processing, 10(4):372– 374, August 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Bookstein:1979:FCS

[Boo79b] Fred L. Bookstein. Fitting conic sections to scattered data. Computer Graphics and Image Processing, 9(1):

56–71, January 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Bookstein:1979:LS

[Boo79c] Fred L. Bookstein. The line skeleton. Computer Graphics and Image Processing, 11(2): 123–137, October 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Bookstein:1997:SIM

[Boo97] Fred L. Bookstein. Shape and the information in medical images: a decade of the morphometric synthesis. Computer Vision and Image Understanding: CVIU, 66(2):97–118, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0607/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0607/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0607/production/ref.

Borgefors:1984:DTA

[Bor84]  Gunilla Borgefors.  Dis-

tance transformations in arbitrary dimensions. Computer Vision, Graphics, and Image Processing, 27(3):321– 345, September 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See also [DC88].

Borgefors:1986:DTD

[Bor86] Gunilla Borgefors. Distance transformations in digital images. Computer Vision, Graphics, and Image Processing, 34(3):344–371, June 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See also [Vos88, BS89].

Borgefors:1991:ACN

[Bor91] G. Borgefors. Another comment on ‘A note on distance transformations in digital images’. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (2):301–306, September 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Borgefors:1996:DDT

[Bor96] Gunilla Borgefors. On digital distance transforms in three dimensions. Computer Vision and Image Understanding: CVIU, 64(3):368–376, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0065/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0065/production/
pdf.

Botting:1978:TPC

[Bot78]  R. J. Botting.  A the-

ory of parametric curve plotting. Computer Graphics and Image Processing, 7(1):139– 145, February 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Boulter:1979:IDI

[Bou79] J. F. Boulter. Interactive digital image restoration and enhancement. Computer Graphics and Image Processing, 11 (4):301–312, December 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Boyd:2004:SOM

[Boy04] Jeffrey E. Boyd. Synchronization of oscillations for machine perception of gaits. Computer Vision and Image Understanding: CVIU, 96(1): 35–59, October 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bertolazzi:1984:PAO

[BP84] P. Bertolazzi and M. Pirozzi. A parallel algorithm for the optimal detection of a noisy curve. Computer Vision, Graphics, and Image Processing, 27(3):380–386, September 1984. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bischel:1994:HFR

[BP94] M. Bischel and A. P. Pentland. Human face recognition and the face image set’s topology. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):254–261, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1017/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1017/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1019/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1019/production/
pdf.

Baraldi:1995:AFL

[BP95] A. Baraldi and F. Parmiggiani.

Alternative form of the Lee filter for speckle suppression in SAR images. Graphical Models and Image Processing: GMIP, 57(1):75–78, January 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1008/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1008/production/
pdf.

Benboudjema:2005:UIS

[BP05] Dalila Benboudjema and Wojciech Pieczynski. Unsupervised image segmentation using triplet Markov fields. Computer Vision and Image Understanding: CVIU, 99 (3):476–498, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bugeau:2009:DSM

[BP09] Aurélie Bugeau and Patrick Pérez. Detection and segmentation of moving objects in complex scenes. Computer Vision and Image Understanding: CVIU, 113(4):459–476, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bulo:2011:GBQ

[BPB11] Samuel Rota Bulò, Marcello Pelillo, and Immanuel M. Bomze. Graph-based quadratic optimization: a fast evolutionary approach. Computer Vision and Image Understanding: CVIU, 115(7): 984–995, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000713.

Bogovic:2013:MOG

[BPB13] John A. Bogovic, Jerry L. Prince, and Pierre-Louis Bazin. A multiple object geometric deformable model for image segmentation. Computer Vision and Image Understanding: CVIU, 117 (2):145–157, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001440.

Batenburg:2013:DAS

[BPBS13] K. Joost Batenburg, Willem Jan Palenstijn, Péter Balázs, and Jan Sijbers. Dynamic angle selection in binary tomography. Computer Vision and Image Understanding: CVIU, 117(4):306–318, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001774.

Bloch:2005:NCS

[BPG05] Isabelle Bloch, Jérémie Pescatore, and Line Garnero. A new characterization of simple elements in a tetrahedral mesh. Graphical Models, 67 (4):260–284, July 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Barrett:1991:GMD

[BPHB91] E. B. Barrett, P. M. Payton, N. N. Haag, and M. H. Brill. General methods for determining projective invariants in imagery. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (1):46–65, January 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Barreto:2010:SIO

[BPS10] Joao P. Barreto, Tomas Pajdla, and Akihiro Sugimoto. Special issue on omnidirectional vision, camera networks and non-conventional cameras. Computer Vision and Image Understanding: CVIU, 114(2):167, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Brzakovic:1991:RBM

[BPW91] D. Brzakovic, R. Patton, and R. L. Wang. Rule-based multitemplate edge detector. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(3):258–268, May 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Brady:1985:DS

[BPYA85] Michael Brady, Jean Ponce, Alan Yuille, and Haruo Asada. Describing surfaces. Computer Vision, Graphics, and Image Processing, 32(1):1–28, October 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bhattacharya:1990:CCI

[BR90] Prabir Bhattacharya and Azriel Rosenfeld. Contour codes of isothetic polygons.

Computer Vision, Graphics, and Image Processing, 50(3):353–363, June 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Banerjee:1993:MEP

[BR93] Saibal Banerjee and Azriel Rosenfeld. MAP estimation of piecewise constant digital signals. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):63–80, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1004/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1004/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1004/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1004/production/
pdf.

Beveridge:1995:OGM

[BR95] J. Ross Beveridge and Edward M. Riseman. Optimal geometric model matching under full 3D perspective. Computer Vision and Image Understanding: CVIU, 61(3):351–364, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.

                                        1028/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1995.1028/production/
                                        pdf.

Bhavsar:2012:RMS

[BR12] Arnav V. Bhavsar and Ambasamudram N. Rajagopalan. Range map superresolutioninpainting, and reconstruction from sparse data. Computer Vision and Image Understanding: CVIU, 116(4): 572–591, April 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002724.

Bracewell:1985:IPR

[Bra85] R. N. Bracewell. An imaging problem: Restoration of blurred digital characters. Computer Vision, Graphics, and Image Processing, 29(3):329–335, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Brandt:1994:CCC

[Bra94] Jonathan W. Brandt. Convergence and continuity criteria for discrete approximations of the continuous planar skeleton. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):116–124, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1007/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1007/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1007/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1007/production/
pdf.

Brand:1997:PBV

[Bra97] Matthew Brand. Physicsbased visual understanding. Computer Vision and Image Understanding: CVIU, 65 (2):192–205, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0572/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0572/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0572/production/ref.

Bascon:2010:OPI

[BRA^{+}10] S. Maldonado Bascón, J. Acevedo Rodríguez, S. Lafuente Arroyo, A. Fernndez Caballero, and F. López-Ferreras. An optimization on pictogram identification for the road-sign recognition task using SVMs. Computer Vision and Image Understanding: CVIU, 114 (3):373–383, March 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Borgefors:1999:MRS

[BRdBS99] Gunilla Borgefors, Giuliana Ramella, Gabriella Sanniti di Baja, and Stina Svensson. On the multiscale representation of 2D and 3D shapes. Graphical Models and Image Processing: GMIP, 61(1):44–62, January 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0489/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1999.0489/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0489/production/ref.

Bretto:2001:CGD

[Bre01]  Alain Bretto.  Compara-

bility graphs and digital topology. Computer Vision and Image Understanding: CVIU, 82(1):33–41, April 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0901;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0901/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0901/ref.

Breuel:2003:ITG

[Bre03] Thomas M. Breuel. Implementation techniques for geometric branch-and-bound matching methods. Computer Vision and Image Understanding: CVIU, 90(3): 258–294, June 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Brill:1984:PIC

[Bri84] Michael H. Brill. Physical and informational constraints on the perception of transparency and translucency. Computer Vision, Graphics, and Image Processing, 28 (3):356–362, December 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Brill:1986:PTM

[Bri86] Michael H. Brill. Perception of transparency in man and machine: a comment on Beck. Computer Vision, Graphics, and Image Processing, 35 (2):270–271, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bribiesca:1998:DEM

[Bri98] Ernesto Bribiesca. Digital elevation model data analysis using the contact surface area. Graphical Models and Image Processing: GMIP, 60(2):166–172, March 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0463/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0463/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0463/production/ref.

Brooks:1978:RED

[Bro78] M. J. Brooks. Rationalizing edge detectors. Computer Graphics and Image Processing, 8(2):277–285, October 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Brown:1994:TGV

[Bro94] Christopher M. Brown. Toward general vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):89–91, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1034/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1034/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1039/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1039/production/
pdf.

Bolle:2004:EAP

[BRP04] Ruud M. Bolle, Nalini K.

Ratha, and Sharath Pankanti. Error analysis of pattern recognition systems—the subsets bootstrap. Computer Vision and Image Understanding: CVIU, 93(1):1–33, January 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bartrina-Rapesta:2011:JRC

[BRSSAL11] Joan Bartrina-Rapesta, Joan Serra-Sagristà, and Francesc Aulí-Llinàs. JPEG2000 ROI coding through component priority for digital mammography. Computer Vision and Image Understanding: CVIU, 115(1):59–68, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bruckstein:1988:SS

[Bru88] Alfred M. Bruckstein. On shape from shading. Computer Vision, Graphics, and Image Processing, 44(2):139– 154, November 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Boukharouba:1985:ASM

[BRW85] S. Boukharouba, J. M. Rebordao, and P. L. Wendel. An amplitude segmentation method based on the distribution function of an image. Computer Vision, Graphics, and Image Processing, 29(1):47–59, January 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bhaskar:1988:PPR

[BRW88] S. K. Bhaskar, Azriel Rosenfeld, and Angela Y. Wu. Parallel processing of regions represented by linear quadtrees. Computer Vision, Graphics, and Image Processing, 42(3):371–380, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bumbaca:1987:DIC

[BS87] Federico Bumbaca and Kenneth C. Smith. Design and implementation of a colour vision model for computer vision applications. Computer Vision, Graphics, and Image Processing, 39(2):226– 245, August 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bumbaca:1988:PAI

[BS88] Federico Bumbaca and Kenneth C. Smith. Practical approach to image restoration for computer vision. Computer Vision, Graphics, and Image Processing, 42(2):220– 233, May 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Beckers:1989:CND

[BS89] A. L. D. Beckers and A. W. M.

Smeulders. A comment on “A Note on Distance Transformations in Digital Images”. Computer Vision, Graphics, and Image Processing, 47(1):89–91, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See [Bor86, Vos88].

Beckers:1992:OLM

[BS92] A. L. D. Beckers and A. W. M.

Smeulders. Optimization of length measurements for isotropic distance transformations in three dimension. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(3): 296–306, May 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Barequet:1996:PLI

[BS96] Gill Barequet and Micha Sharir. Piecewise-linear interpolation between polygonal slices. Computer Vision and Image Understanding: CVIU, 63(2):251–272, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0018/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0018/production/
pdf.

Berman:1999:FID

[BS99] Andrew P. Berman and Linda G. Shapiro. A flexible image database system for content-based retrieval. Computer Vision and Image Understanding: CVIU, 75 (1–2):175–195, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0772/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0772/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0772/production/ref.

Brejl:2000:DED

[BS00a] Marek Brejl and Milan Sonka. Directional 3D edge detection in anisotropic data: Detector design and performance assessment. Computer Vision and Image Understanding: CVIU, 77(2):84–110, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0811;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.1999.0811/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0811/ref.

Bubna:2000:MST

[BS00b] Kishore Bubna and Charles V. Stewart. Model selection techniques and merging rules for range data segmentation algorithms. Computer Vision and Image Understanding: CVIU, 80(2):215–245, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0871;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0871/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0871/ref.

Biswas:2004:ADF

[BS04a] Arpan Biswas and Vadim Shapiro. Approximate distance fields with non-vanishing gradients. Graphical Models, 66(3):133–159, May 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Braquelaire:2004:CMR

[BS04b] Achille Braquelaire and Robert Strandh. A color model for rendering linear passive graphic 2D objects. Graphical Models, 66(2):71–88, March 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Bartoli:2005:SMU

[BS05] Adrien Bartoli and Peter Sturm. Structure-from-motion using lines: Representation, triangulation, and bundle adjustment. Computer Vision and Image Understanding: CVIU, 100(3): 416–441, December 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Brevdo:1987:SAP

[BSB87] L. Brevdo, S. Sideman, and R. Beyar. Simple approach to the problem of 3-D reconstruction. Computer Vision, Graphics, and Image Processing, 37(3):420– 427, March 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Boyer:2002:SSS

[BSF02] Kim L. Boyer, Ravi Srikantiah, and Patrick J. Flynn. Saliency sequential surface organization for free-form object recognition. Computer Vision and Image Understanding: CVIU, 88(3):152–188, December 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Beck:1987:SFC

[BSI87] Jacob Beck, Anne Sutter, and Richard Ivry. Spatial frequency channels and perceptual grouping in texture segregation. Computer Vision, Graphics, and Image Processing, 37(2):299– 325, February 1987. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Babu:2010:OAR

[BSM10] R. Venkatesh Babu, S. Suresh, and Anamitra Makur. Online adaptive radial basis function networks for robust object tracking. Computer Vision and Image Understanding: CVIU, 114(3):297–310, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bartesaghi:2005:TDS

[BSMG05] Alberto Bartesaghi, Guillermo Sapiro, Tom Malzbender, and Dan Gelb. Three-dimensional shape rendering from multiple images. Graphical Models, 67 (4):332–346, July 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Binder:2013:ERM

[BSMK13] Alexander Binder, Wojciech Samek, Klaus-Robert Müller, and Motoaki Kawanabe. Enhanced representation and multi-task learning for image annotation. Computer Vision and Image Understanding: CVIU, 117(5):466–478, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001713.

Barsky:2001:SIP

[BSW01] Brian Barsky, Yoshihisa Shinagawa, and Wenping Wang. Special issue on Pacific Graphics 2000. Graphical Models, 63(4):211, July 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Blanford:1988:BSD

[BT88] Ronald P. Blanford and Steven L. Tanimoto. Brightspot detection in pyramids. Computer Vision, Graphics, and Image Processing, 43 (2):133–149, August 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Bentoutou:2005:DST

[BT05] Y. Bentoutou and N. Taleb. A 3-D space–time motion detection for an invariant image registration approach in digital subtraction angiography. Computer Vision and Image Understanding: CVIU, 97(1): 30–50, January 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Barsky:2005:EAD

[BTCH05] Brian A. Barsky, Michael J. Tobias, Derrick P. Chu, and Daniel R. Horn. Elimination of artifacts due to occlusion and discretization problems in image space blurring techniques. Graphical Models, 67(6):584–599, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000093.

Ben-Tzvi:1990:SMI

[BTNS90] D. Ben-Tzvi, A. Naqvi, and M. Sandler. Synchronous multiprocessor implementation of the Hough transform. Computer Vision, Graphics, and Image Processing, 52 (3):437–446, December 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Basri:1993:AOS

[BU93] Ronen Basri and Shimon Ullman. The alignment of objects with smooth surfaces. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):331–345, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1022/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1022/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1022/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1022/production/
pdf.

Buchanan:1988:TCC

[Buc88]  Thomas Buchanan.  The

twisted cubic and camera calibration. Computer Vision, Graphics, and Image Processing, 42(1):130– 132, April 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Burt:1980:TPS

[Bur80] Peter J. Burt. Tree and pyramid structures for coding hexagonally sampled binary images. Computer Graphics and Image Processing, 14(3): 271–280, November 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Burr:1981:DMI

[Bur81a] D. J. Burr. A dynamic model for image registration. Computer Graphics and Image Processing, 15(2):102– 112, February 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Burt:1981:FFT

[Bur81b] Peter J. Burt. Fast filter transforms for image processing. Computer Graphics and Image Processing, 16 (1):20–51, May 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Burt:1983:FAE

[Bur83]  Peter J. Burt.  Fast al-

gorithms for estimating local image properties. Computer Vision, Graphics, and Image Processing, 21(3):368– 382, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Buzer:2003:LIA

[Buz03] Lilian Buzer. A linear incremental algorithm for naive and standard digital lines and planes recognition. Graphical Models, 65(1–3):61–76, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Braquelaire:1999:EPN

[BV99] Jean-Pierre Braquelaire and Anne Vialard. Euclidean paths: a new representation of boundary of discrete regions. Graphical Models and Image Processing: GMIP, 61(1):16–43, January 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0488/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0488/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0488/production/ref.

Berendsen:2013:FFI

[BvdHL^{+}13] Floris F. Berendsen, Uulke A.

van der Heide, Thomas R. Langerak, Alexis N. T. J. Kotte, and Josien P. W. Pluim. Free-form image registration regularized by a statistical shape model: application to organ segmentation in cervical MR. Computer Vision and Image Understanding: CVIU, 117(9): 1119–1127, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000660.

Balmelli:2002:MOU

[BVL02] Laurent Balmelli, Martin Vetterli, and Thomas M. Liebling. Mesh optimization using global error with application to geometry simplification. Graphical Models, 64(3– 4):230–257, May 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Bonczek:1976:PPA

[BW76] Robert H. Bonczek and Andrew B. Whinston. Picture processing and automatic data base design. Computer Graphics and Image Processing, 5(4):484– 495, December 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Boult:1993:LIR

[BW93] Terrance E. Boult and George Wolberg. Local image reconstruction and subpixel restoration algorithms. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(1):63–77, January 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1005/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1005/production/
pdf.

Barequet:1998:OSS

[BW98] Gill Barequet and Barbara Wolfers. Optimizing a strip separating two polygons. Graphical Models and Image Processing: GMIP, 60(3):214–221, May 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0470/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0470/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0470/production/ref.

Bai:2011:RTU

[BW11] Li Bai and Yan Wang. Road tracking using particle filters with partition sampling and auxiliary variables. Computer Vision and Image Understanding: CVIU, 115(10): 1463–1471, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001421.

Bischof:2004:IIR

[BWL04] Horst Bischof, Horst Wildenauer, and Aleš Leonardis. Illumination insensitive recognition using eigenspaces. Computer Vision and Image Understanding: CVIU, 95 (1):86–104, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bolle:1998:CVV

[BY98] R. Bolle and B.-L. Yeo. Computer vision for visual computing: Techniques and applications. Computer Vision and Image Understanding: CVIU, 71(2):153–??, ???? 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bashforth:2001:PBE

[BY01] Byron Bashforth and Yee-Hong Yang. Physics-based explosion modeling. Graphical Models, 63(1):21–44, January 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2000.0536;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/gmod.2000.0536/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2000.0536/ref.

Bastanlar:2008:CVB

[BY08] Yalin Bastanlar and Yasemin Yardimci. Corner validation based on extracted corner properties. Computer Vision and Image Understanding: CVIU, 112(3):243– 261, December 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Bilir:2012:NRS

[BY12] S. C. Bilir and Y. Yemez.

Non-rigid 3D shape tracking from multiview video. Computer Vision and Image Understanding: CVIU, 116(11): 1121–1134, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001026.

Bing:2004:TSM

[BYN^{+}04] Cheng Bing, Wang Ying, Zheng Nanning, Bian Zhengzhong, and Zhang Yongping. A two-step method for preprocessing volume data. Computer Vision and Image Understanding: CVIU, 95(2): 150–164, August 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Boshra:1999:CSA

[BZ99] Michael Boshra and Hong Zhang. A constraint-satisfaction approach for 3D object recognition by integrating 2D and 3D data. Computer Vision and Image Understanding: CVIU, 73(2):200–214, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0730/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0730/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0730/production/ref.

Basharat:2008:CBV

[BZS08] Arslan Basharat, Yun Zhai, and Mubarak Shah. Content based video matching using spatiotemporal volumes. Computer Vision and Image Understanding: CVIU, 110 (3):360–377, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chien:1984:NQR

[CA84] C. H. Chien and J. K. Aggarwal. A normalized quadtree representation. Computer Vision, Graphics, and Image Processing, 26(3):331– 346, June 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chien:1986:IOM

[CA86a] C. H. Chien and J. K. Aggarwal. Identification of 3D objects from multiple silhouettes using quadtrees/octrees.

Computer Vision, Graphics, and Image Processing, 36(2/3):256–273, November/ December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chien:1986:VSO

[CA86b] C. H. Chien and J. K. Aggarwal. Volume/surface octrees for the representation of three-dimensional objects. Computer Vision, Graphics, and Image Processing, 36 (1):100–113, October 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chang:1997:LCC

[CA97] Yuh-Lin Chang and J. K.

Aggarwal. Line correspondences from cooperating spatial and temporal grouping processes for a sequence of images. Computer Vision and Image Understanding: CVIU, 67(2):186–201, August 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0527/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0527/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0527/production/ref.

Cavallaro:2010:SIM

[CA10] Andrea Cavallaro and Hamid Aghajan. Special issue on multi-camera and multimodal sensor fusion. Computer Vision and Image Understanding: CVIU, 114(6): 609–610, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Camahort:2009:LSA

[CAF09] Emilio Camahort, Francisco Abad, and Don Fussell. A line-space analysis of lightfield representations. Graphical Models, 71(5):169–183, September 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070309000101.

Caglioti:1993:USL

[Cag93] Vincenzo Caglioti. On the uncertainty of straight lines in digital images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(4):255–270, July 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1993.
                                        1018/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1993.1018/production/
                                        pdf.

Cai:1988:RBI

[Cai88] Zuguang Cai. Restoration of binary images using contour direction chain codes description. Computer Vision, Graphics, and Image Processing, 41(1):101–106, January 1988. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Capson:1984:IAS

[Cap84] David W. Capson. An improved algorithm for the sequential extraction of boundaries from a raster scan. Computer Vision, Graphics, and Image Processing, 28 (1):109–125, October 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Carstensen:1996:ALM

[Car96] Jens Michael Carstensen. An active lattice model in a Bayesian framework. Computer Vision and Image Understanding: CVIU, 63 (2):380–387, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0027/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0027/production/
pdf.

Carlin:2001:MPS

[Car01]  Mats Carlin.  Measuring

the performance of shape similarity retrieval methods.

Computer Vision and Image Understanding: CVIU, 84(1):44–61, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0935;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0935/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0935/ref.

Cavanagh:1987:RTD

[Cav87] Patrick Cavanagh. Reconstructing the third dimension: Interactions between color, texture, motion, binocular disparity, and shape. Computer Vision, Graphics, and Image Processing, 37 (2):171–195, February 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chen:1998:SID

[CB98] Francine R. Chen and Dan S. Bloomberg. Summarization of imaged documents without OCR. Computer Vision and Image Understanding: CVIU, 70(3):307–320, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0688/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0688/production/
pdf;   http://www.idealibrary.

com/links/artid/cviu.1998.
0688/production/ref.

Cooper:1995:CSU

[CBB95] Paul R. Cooper, Lawrence A.

Birnbaum, and Matthew E. Brand. Causal scene understanding. Computer Vision and Image Understanding: CVIU, 62(2): 215–231, September 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1051/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1051/production/
pdf.

Capell:2007:PBR

[CBC^{+}07] Steve Capell, Matthew Burkhart, Brian Curless, Tom Duchamp, and Zoran Popović. Physically based rigging for deformable characters. Graphical Models, 69(1):71–87, January 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S1524070306000634.

Cachier:2003:IFB

[CBD^{+}03] Pascal Cachier, Eric Bardinet, Didier Dormont, Xavier Pennec, and Nicholas Ayache. Iconic feature based nonrigid registration: the PASHA algorithm. Computer Vision and Image Understanding: CVIU, 89(2– 3):272–298, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Couprie:2003:DO

[CBK03] Michel Couprie, Gilles Bertrand, and Yukiko Kenmochi. Discretization in 2D and 3D orders. Graphical Models, 65 (1–3):77–91, May 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Chen:2001:SED

[CBM01] Y. Chen, C. A. Z. Barcelos, and B. A. Mair. Smoothing and edge detection by time-varying coupled nonlinear diffusion equations. Computer Vision and Image Understanding: CVIU, 82(2):85–100, May 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0903;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0903/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0903/ref.

Chaudhuri:2004:ESE

[CBT^{+}04] A. Ray Chaudhuri, A. Basu, K. Tan, S. Bhandari, and B. B. Chaudhuri. An efficient set estimator in high dimensions: consistency and applications to fast data visualization. Computer Vision and Image Understanding: CVIU, 93(3):260–287, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cohen:1996:HHM

[CC96] Isaac Cohen and Laurent D.

Cohen. Hybrid hyperquadric model for 2-D and 3-D data fitting. Computer Vision and Image Understanding: CVIU, 63(3):527–541, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0039/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0039/production/
pdf.

Carevic:1997:RBC

[CC97] Dragana Carevic and Terry Caelli. Region-based coding of color images using Karhunen-Loeve transform. Graphical Models and Image Processing: GMIP, 59(1):27–38, January 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0402/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0402/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1996.
0402/production/ref.

Cretual:2000:DSP

[CC00] Armel Crétual and François Chaumette. Dynamic stabilization of a pan and tilt camera for submarine image visualization. Computer Vision and Image Understanding: CVIU, 79(1):47–65, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0849;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0849/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0849/ref.

Chen:2001:ERA

[CC01] Teh-Chuan Chen and Kuo-Liang Chung. An efficient randomized algorithm for detecting circles. Computer Vision and Image Understanding: CVIU, 83(2):172–191, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0923;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0923/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0923/ref.

Chowdhury:2003:FRM

[CC03] Amit K. Roy Chowdhury and Rama Chellappa. Face reconstruction from monocular video using uncertainty analysis and a generic model. Computer Vision and Image Understanding: CVIU, 91(1–2): 188–213, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheng:2007:BSM

[CC07] Li Cheng and Terry Caelli.

Bayesian stereo matching. Computer Vision and Image Understanding: CVIU, 106 (1):85–96, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2011:SSV

[CC11] Yu Chen and Roberto Cipolla.

Single and sparse view 3D reconstruction by learning shape priors. Computer Vision and Image Understanding: CVIU, 115(5):586–602, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cohen:1992:UDS

[CCA92] Isaac Cohen, Laurent D.

Cohen, and Nicholas Ayache. Using deformable surfaces to segment 3-D images and infer differential structures. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (2):242–263, September 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Colombo:2011:SRT

[CCD11] Carlo Colombo, Dario Comanducci, and Alberto Del Bimbo. Shape reconstruction and texture sampling by active rectification and virtual view synthesis. Computer Vision and Image Understanding: CVIU, 115(2): 161–176, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2001:BQS

[CCF01] Changsong Chen, Falai Chen, and Yuyu Feng. Blending quadric surfaces with piecewise algebraic surfaces. Graphical Models, 63(4):212– 227, July 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Chaquet:2013:SVD

[CCFC13] Jose M. Chaquet, Enrique J. Carmona, and Antonio Fernández-Caballero. A survey of video datasets for human action and activity recognition. Computer Vision and Image Understanding: CVIU, 117(6):633–659, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000295.

Chang:1991:RMI

[CCH91] Long-Wen Chang, Hown-Wen Chen, and Ju-Rone Ho. Reconstruction of 3D medical images: a nonlinear interpolation technique for reconstruction of 3D medical images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):382–391, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Chang:2004:LTC

[CCL04] Fu Chang, Chun-Jen Chen, and Chi-Jen Lu. A lineartime component-labeling algorithm using contour tracing technique. Computer Vision and Image Understanding: CVIU, 93(2):206–220, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Choi:1997:NAM

[CCMW97] Hyeong In Choi, Sung Woo Choi, Hwan Pyo Moon, and Nam-Sook Wee. New algorithm for medial axis transform of plane domain. Graphical Models and Image Processing: GMIP, 59 (6):463–483, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0444/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0444/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/gmip.1997.
                                        0444/production/ref.

Chaudhuri:1997:NAC

[CCP97] A. Ray Chaudhuri, B. B. Chaudhuri, and S. K. Parui. A novel approach to computation of the shape of a dot pattern and extraction of its perceptual border. Computer Vision and Image Understanding: CVIU, 68(3):257–275, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0550/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0550/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0550/production/ref.

Cernadas:2005:AMR

[CCR^{+}05] E. Cernadas, P. Carrión, P. G. Rodriguez, E. Muriel, and T. Antequera. Analyzing magnetic resonance images of Iberian pork loin to predict its sensorial characteristics. Computer Vision and Image Understanding: CVIU, 98(2): 344–360, May 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Choy:1995:NSP

[CCS95] Steven S. O. Choy, Clifford Sze-Tsan Choy, and Wan-Chi Siu. New singlepass algorithm for parallel thinning. Computer Vision and Image Understanding: CVIU, 62(1):69–77, July 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1042/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1042/production/
pdf.

Choudhury:2001:RBR

[CCS01] Ragini Choudhury, Santanu Chaudhury, and J. B. Srivastava. Reconstruction based recognition of scenes with multiple repeated components. Computer Vision and Image Understanding: CVIU, 84(3):325–360, December 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chazal:2005:CIA

[CCS05] Frédéric Chazal and David Cohen-Steiner. A condition for isotopic approximation. Graphical Models, 67 (5):390–404, September 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000044.

Cardoso:2009:PDM

[CCTCR09] Jaime S. Cardoso, Pedro Carvalho, Luís F. Teixeira, and Luís Corte-Real. Partitiondistance methods for assessing spatial segmentations of images and videos. Computer Vision and Image Understanding: CVIU, 113(7): 811–823, July 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Choi:2012:RMO

[CCYC12] JinMin Choi, Hyung Jin Chang, Yung Jun Yoo, and Jin Young Choi. Robust moving object detection against fast illumination change. Computer Vision and Image Understanding: CVIU, 116 (2):179–193, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002347.

Chattopadhyay:1992:PER

[CD92] S. Chattopadhyay and P. P. Das. Parameter estimation and reconstruction of digital conics in normal positions. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(5):385– 395, September 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Chen:1993:PAV

[CD93] Ling Tony Chen and Larry S. Davis. A parallel algorithm for the visibility of a simple polygon using scan operations. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(3):192–202, May 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1014/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1014/production/
pdf.

Chen:1995:TIR

[CD95] Jim X. Chen and Niels Da Vitoria Lobo. Toward interactive-rate simulation of fluids with moving obstacles using Navier– Stokes equations. Graphical Models and Image Processing: GMIP, 57(2):107–116, March 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1012/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1012/production/
pdf.

Chen:2010:PPB

[CD10] Fan Chen and Christophe De Vleeschouwer. Personalized production of basketball videos from multi-sensored data under limited display resolution. Computer Vision and Image Understanding: CVIU, 114(6):667–680, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Comic:2011:DIS

[CD11] Lidija Comić and Leila De Floriani. Dimensionindependent simplification and refinement of Morse complexes. Graphical Models, 73 (5):261–285, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070311000154.

Chung:2013:RAM

[CD13] François Chung and Hervé Delingette. Regional appearance modeling based on the clustering of intensity profiles. Computer Vision and Image Understanding: CVIU, 117(6):705–717, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000271.

Csurka:1999:FCB

[CDH99] Gabriella Csurka, David Demirdjian, and Radu Horaud. Finding the collineation between two projective reconstructions. Computer Vision and Image Understanding: CVIU, 75(3): 260–268, September 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0782/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0782/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0782/production/ref.

Cordella:1977:TCP

[CDLD77] L. Cordella, M. J. B. Duff, S. Levialdi, and P. E. Danielsson. Thresholding: a challenge for parallel processing getting the median faster. Computer Graphics and Image Processing, 6(3):207–220, 71–78, June 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chaperon:2011:RCP

[CDT11] Thomas Chaperon, Jacques Droulez, and Guillaume Thibault. Reliable camera pose and calibration from a small set of point and line correspondences: a probabilistic approach. Computer Vision and Image Understanding: CVIU, 115(5):576–585, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cooper:1980:SBE

[CEC^{+}80] D. B. Cooper, H. Elliott, F. Cohen, L. Reiss, and P. Symoser. Stochastic boundary estimation and object recognition. Computer Graphics and Image Processing, 12(4):326–356, April 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Cederberg:1979:CLC

[Ced79a] Roger L. T. Cederberg. Chain-link coding and segmentation for raster scan devices. Computer Graphics and Image Processing, 10(3): 224–234, July 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Cederberg:1979:NMV

[Ced79b] Roger L. T. Cederberg. A new method for vector generation. Computer Graphics and Image Processing, 9(2): 183–195, February 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Celenk:1990:CCT

[Cel90] M. Celenk. A color clustering technique for image segmentation. Computer Vision, Graphics, and Image Processing, 52(2):145– 170, November 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Crabtree:1984:ESS

[CEP84] Sterling J. Crabtree, Jr., Robert Ehrlich, and Christopher Prince. Evaluation of strategies for segmentation of blue-dyed pores in thin sections of reservoir rocks. Computer Vision, Graphics, and Image Processing, 28(1):1–18, October 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Cohen:1992:MLU

[CF92] Fernand S. Cohen and Zhigang Fan. Maximum likelihood unsupervised textured image segmentation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(3):239–251, May 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Campbell:2001:SFF

[CF01] Richard J. Campbell and Patrick J. Flynn. A survey of free-form object representation and recognition techniques. Computer Vision and Image Understanding: CVIU, 81(2):166–210, February 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0889;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0889/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0889/ref.

Cao:2007:CCL

[CF07] Xiaochun Cao and Hassan Foroosh. Camera calibration and light source orientation from solar shadows. Computer Vision and Image Understanding: CVIU, 105(1): 60–72, January 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheong:1998:EEV

[CFA98] LoongFah Cheong, Cornelia Fermüller, and Yiannis Aloimonos. Effects of errors in the viewing geometry on shape estimation. Computer Vision and Image Understanding: CVIU, 71(3): 356–372, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0649/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0649/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0649/production/ref.

Chen:2005:IVL

[CFB05] Xin Chen, Patrick J. Flynn, and Kevin W. Bowyer. IR and visible light face recognition. Computer Vision and Image Understanding: CVIU, 99 (3):332–358, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Canton-Ferrer:2011:HMC

[CFCP11] Cristian Canton-Ferrer, Josep R.

Casas, and Montse Pardàs. Human motion capture using scalable body models. Computer Vision and Image Understanding: CVIU, 115(10):1363–1374, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421100138X.

Cerri:2006:RTI

[CFG06] A. Cerri, M. Ferri, and D. Giorgi. Retrieval of trademark images by means of size functions. Graphical Models, 68(5–6):451–471, September/November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000592.

Castellani:2002:RMA

[CFM02] Umberto Castellani, Andrea Fusiello, and Vittorio Murino. Registration of multiple acoustic range views for underwater scene reconstruction. Computer Vision and Image Understanding: CVIU, 87(1–3):78–89, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chou:2013:SIR

[CFM^{+}13] Chen-Rui Chou, Brandon Frederick, Gig Mageras, Sha Chang, and Stephen Pizer. 2D/3D image registration using regression learning. Computer Vision and Image Understanding: CVIU, 117(9): 1095–1106, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000647.

Camps:1998:RPC

[CFS98] Octavia Camps, Patrick J. Flynn, and George C. Stockman. Recent progress in CAD-based computer vision: An introduction to the special issue. Computer Vision and Image Understanding: CVIU, 69(3):251–252, March 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0662/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0662/production/
pdf.

Cappabianco:2012:BTM

[CFYU12] Fábio A. M. Cappabianco, Alexandre X. Falcão, Clarissa L. Yasuda, and Jayaram K. Udupa. Brain tissue MRimage segmentation via optimumpath forest clustering. Computer Vision and Image Understanding: CVIU, 116(10): 1047–1059, October 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000999.

Carpenter:1987:MPA

[CG87] Gail A. Carpenter and Stephen Grossberg. A massively parallel architecture for a self-organizing neural pattern recognition machine. Computer Vision, Graphics, and Image Processing, 37(1):54– 115, January 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). CARPENTER87.

Chionh:1994:ECI

[CG94] Eng-Wee W. Chionh and Ronald N. Goldman. On the existence and the coefficients of the implicit equation of rational surfaces. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):19–24, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1003/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1003/production/
pdf.

Cortelazzo:2004:MBI

[CG04] Guido M. Cortelazzo and Concettina Guerra. Modelbased and image-based 3D scene representation for interactive visualization. Computer Vision and Image Understanding: CVIU, 96(3): 269–273, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Caglioti:2009:RBM

[CG09] Vincenzo Caglioti and Alessandro Giusti. Recovering ball motion from a single motionblurred image. Computer Vision and Image Understanding: CVIU, 113(5):590–597, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Campos:2013:SBS

[CGAY13] Ricard Campos, Rafael Garcia, Pierre Alliez, and Mariette Yvinec. Splat-based surface reconstruction from defect-laden point sets. Graphical Models, 75(6):346–361, November 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000258.

Cumani:1991:EBD

[CGG91] A. Cumani, P. Grattoni, and A. Guiducci. An edge-based description of color images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):313–323, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Caillette:2008:RTD

[CGH08] Fabrice Caillette, Aphrodite Galata, and Toby Howard. Real-time 3-D human body tracking using learnt models of behaviour. Computer Vision and Image Understanding: CVIU, 109(2): 112–125, February 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cinque:1992:CSD

[CGL92] L. Cinque, C. Guerra, and S. Levialdi. Computing shape description transforms on a multiresolution architecture. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(3): 287–295, May 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Cinque:1994:PRM

[CGL94] L. Cinque, C. Guerra, and S. Levialdi. On the paper by R. M. Haralick. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2): 250–252, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1051/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1051/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/cviu.1994.
                                        1056/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1056/production/
                                        pdf.

Carpenter:1998:WWF

[CGL98] G. A. Carpenter, S. Grossberg, and G. W. Lesher. The what-and-where filter. A spatial mapping neural network for object recognition and image understanding. Computer Vision and Image Understanding: CVIU, 69(1): 1–??, ???? 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Costabile:1985:MSC

[CGP85] M. F. Costabile, C. Guerra, and G. G. Pieroni. Matching shapes: a case study in time-varying images. Computer Vision, Graphics, and Image Processing, 29(3):296– 310, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chakraborty:2013:LSC

[CGR13] Bhaskar Chakraborty, Jordi Gonzàlez, and F. Xavier Roca. Large scale continuous visual event recognition using max-margin Hough transformation framework. Computer Vision and Image Understanding: CVIU, 117(10): 1356–1368, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001920.

Cakir:2011:NNB

[CGU11] Fatih Cakir, Ugur Güdükbay, and Özgür Ulusoy. Nearestneighbor based metric functions for indoor scene recognition. Computer Vision and Image Understanding: CVIU, 115(11):1483–1492, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001780.

Cutler:2007:ADW

[CGW^{+}07] Lawrence D. Cutler, Reid Gershbein, Xiaohuan Corina Wang, Cassidy Curtis, Erwan Maigret, Luca Prasso, and Peter Farson. An art-directed wrinkle system for CG character clothing and skin. Graphical Models, 69(5–6):219–230, September/November 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000622.

Conners:1978:EPQ

[CH78] Richard W. Conners and Charles A. Harlow. Equal probability quantizing and texture analysis of radiographic images. Computer Graphics and Image Processing, 8(3):447–463, December 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Conners:1980:TST

[CH80] Richard W. Conners and Charles A. Harlow. Toward a structural textural analyzer based on statistical methods. Computer Graphics and Image Processing, 12 (3):224–256, March 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Carrihill:1985:EIR

[CH85] Brian Carrihill and Robert Hummel. Experiments with the intensity ratio depth sensor. Computer Vision, Graphics, and Image Processing, 32 (3):337–358, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Cash:1987:OCR

[CH87] Glenn L. Cash and Mehdi Hatamian. Optical character recognition by the method of moments. Computer Vision, Graphics, and Image Processing, 39(3):291–310, September 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chen:1988:SCM

[CH88] Homer H. Chen and Thomas S. Huang. A survey of construction and manipulation of octrees. Computer Vision, Graphics, and Image Processing, 43(3):409–431, September 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Castano:1996:PAP

[CH96] Rebecca L. Castaño and Seth Hutchinson. A probabilistic approach to perceptual grouping. Computer Vision and Image Understanding: CVIU, 64(3):399–419, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0068/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0068/production/
pdf.

Christy:1999:IPC

[CH99] Stéphane Christy and Radu Horaud. Iterative pose computation from line correspondences. Computer Vision and Image Understanding: CVIU, 73(1):137–144, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0717/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0717/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0717/production/ref.

Castelan:2006:AHD

[CH06a] Mario Castelán and Edwin R. Hancock. Acquiring height data from a single image of a face using local shape indicators. Computer Vision and Image Understanding: CVIU, 103(1):64–79, July 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Collomosse:2006:VMA

[CH06b] J. P. Collomosse and P. M. Hall. Video motion analysis for the synthesis of dynamic cues and futurist art. Graphical Models, 68(5–6):402–414, September/November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070306000403.

Corso:2009:IDF

[CH09] J. J. Corso and G. D. Hager. Image description with features that summarize. Computer Vision and Image Understanding: CVIU, 113(4): 446–458, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chesi:2011:FMV

[CH11] G. Chesi and Y. S. Hung. Fast multiple-view L _{2} triangulation with occlusion handling. Computer Vision and Image Understanding: CVIU, 115 (2):211–223, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chaikin:1974:AHS

[Cha74] G. Chaikin. An algorithm for high speed curve generation. Computer Graphics and Image Processing, 3(??): 346–349, ???? 1974. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chassery:1979:CCD

[Cha79] Jean Marc Chassery. Connectivity and consecutivity in digital pictures. Computer Graphics and Image Processing, 9(3):294–300, March 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chakravarty:1981:SPC

[Cha81] Indranil Chakravarty. Singlepass, chain generating algorithm for region boundaries. Computer Graphics and Image Processing, 15(2):182– 193, February 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chassery:1983:DCD

[Cha83] Jean-Marc Chassery. Discrete convexity: Definition, parametrization, and compatibility with continuous convexity. Computer Vision, Graphics, and Image Processing, 21(3):326–344, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chalmond:1991:PEI

[Cha91] B. Chalmond. PSF estimation for image deblurring. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):364–372, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Cheng:1986:AAZ

[CHB86] L. M. Cheng, A. S. Ho, and R. E. Burge. An adaptive asteroid zonal filter for data compression. Computer Vision, Graphics, and Image Processing, 34(3):292– 301, June 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Cheng:2011:ELD

[CHC11] Xiangang Cheng, Yiqun Hu, and Liang-Tien Chia. Exploiting local dependencies with spatial-scale space (S-Cube) for near-duplicate retrieval. Computer Vision and Image Understanding: CVIU, 115 (6):750–758, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:1991:DMD

[Che91] Homer H. Chen. Determining motion and depth from binocular orthographic views. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(1):47–55, July 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Cheng:1996:AAI

[Che96]  Yu Cheng.  Analysis of

affine invariants as approximate perspective invariants. Computer Vision and Image Understanding: CVIU, 63(2):197–207, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0014/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0014/production/
pdf.

Chen:1998:HDP

[Che98]  Yung-Sheng Chen.  Hid-

den deletable pixel detection using vector analysis in parallel thinning to obtain bias-reduced skeletons. Computer Vision and Image Understanding: CVIU, 71(3):294–311, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0647/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0647/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0647/production/ref.

Cheong:2000:SBS

[Che00] Loong-Fah Cheong. Scenebased shot change detection and comparative evaluation. Computer Vision and Image Understanding: CVIU, 79(2):224–235, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0858;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0858/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0858/ref.

Chen:2008:FLR

[Che08]  Hui Chen.  Focal length

and registration correction for building panorama from photographs. Computer Vision and Image Understanding: CVIU, 112(2):225– 230, November 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Celik:2009:UST

[CHH09] Hasan Celik, Alan Hanjalic, and Emile A. Hendriks. Unsupervised and simultaneous training of multiple object detectors from unlabeled surveillance video. Computer Vision and Image Understanding: CVIU, 113(10): 1076–1094, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chittineni:1981:USS

[Chi81] C. B. Chittineni. Utilization of spectral-spatial information in the classification of imagery data. Computer Graphics and Image Processing, 16(4): 305–340, August 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). CHITTINENI81.

Chin:1988:AVI

[Chi88]  Roland T. Chin.  Auto-

mated visual inspection: 1981 to 1987. Computer Vision, Graphics, and Image Processing, 41(3):346– 381, March 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chitti:1997:DSL

[Chi97]  Yasmina Chitti.  Detec-

tion of small local intensity changes in CCD images with nonuniform illumination and large signal dependent noise. Graphical Models and Image Processing: GMIP, 59(3):139–148, May 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.

                                        com/links/artid/gmip.1997.
                                        0426/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0426/production/
                                        pdf.

Cutzu:2005:DPP

[CHL05] Florin Cutzu, Riad Hammoud, and Alex Leykin. Distinguishing paintings from photographs. Computer Vision and Image Understanding: CVIU, 100(3):249– 273, December 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chakraborty:2012:SST

[CHMG12] Bhaskar Chakraborty, Michael B. Holte, Thomas B. Moeslund, and Jordi Gonzàlez. Selective spatio-temporal interest points. Computer Vision and Image Understanding: CVIU, 116(3):396–410, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002128.

Chow:1979:AGI

[Cho79] William W. Chow. Automatic generation of interlocking shapes. Computer Graphics and Image Processing, 9 (4):333–353, April 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chochia:1988:IEU

[Cho88]  Pavel A. Chochia.  Im-

age enhancement using sliding histograms. Computer Vision, Graphics, and Image Processing, 44(2):211– 229, November 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Calderara:2011:DAP

[CHP^{+}11] Simone Calderara, Uri Heinemann, Andrea Prati, Rita Cucchiara, and Naftali Tishby. Detecting anomalies in people’s trajectories using spectral graph analysis. Computer Vision and Image Understanding: CVIU, 115(8): 1099–1111, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000919.

Cox:1996:MLS

[CHRM96] Ingemar J. Cox, Sunita L.

Hingorani, Satish B. Rao, and Bruce M. Maggs. A maximum likelihood stereo algorithm. Computer Vision and Image Understanding: CVIU, 63(3):542–567, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0040/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1996.0040/production/
                                        pdf.

Caselles:2008:GVM

[CHSV08] V. Caselles, G. Haro, G. Sapiro, and J. Verdera. On geometric variational models for inpainting surface holes. Computer Vision and Image Understanding: CVIU, 111(3):351– 373, September 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chung:1977:CGA

[Chu77] W. L. Chung. On circle generation algorithms. Computer Graphics and Image Processing, 6(??):196–198, ???? 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chung:2002:RVD

[Chu02] Ronald Chung. Relative viewing distance: a correspondence invariance under paraperspective projection. Computer Vision and Image Understanding: CVIU, 86(1): 1–31, April 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2013:RFB

[CHZ^{+}13] X. Chen, H. He, G. Zou, X. Zhang, X. Gu, and J. Hua. Ricci flow-based spherical parameterization and surface registration. Computer Vision and Image Understanding: CVIU, 117(9): 1107–1118, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000659.

Coleman:1982:ODS

[CJ82] E. North Coleman, Jr. and Ramesh Jain. Obtaining 3dimensional shape of textured and specular surfaces using four-source photometry. Computer Graphics and Image Processing, 18(4): 309–328, April 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chen:1993:SMV

[CJ93] Sei-Wang W. Chen and Anil K. Jain. Strategies of multi-view and multimatching for 3D object recognition. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):121–130, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1008/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1008/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1008/production;   http:
//www.idealibrary.com/links/

artid/cviu.1993.1008/production/
pdf.

Collins:1998:ASA

[CJC^{+}98] Robert T. Collins, Christopher O. Jaynes, Yong-Qing Cheng, Xiaoguang Wang, Frank Stolle, Edward M. Riseman, and Allen R. Hanson. The ascender system: Automated site modeling from multiple aerial images. Computer Vision and Image Understanding: CVIU, 72(2):143–162, November 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0729/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0729/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0729/production/ref.

Cord:2001:ABS

[CJC01] Matthieu Cord, Michel Jordan, and Jean-Pierre Cocquerez. Accurate building structure recovery from high resolution aerial imagery. Computer Vision and Image Understanding: CVIU, 82(2):138–173, May 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0905;   http://www.
idealibrary.com/links/doi/

10.1006/cviu.2001.0905/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0905/ref.

Chin:2006:HSI

[CJL06] Chong Siew Chin, Andrew Teoh Beng Jin, and David Ngo Chek Ling. High security Iris verification system based on random secret integration. Computer Vision and Image Understanding: CVIU, 102 (2):169–177, May 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chang:1984:TDE

[CK84] Edward S. H. Chang and Ludwik Kurz. Trajectory detection and experimental designs. Computer Vision, Graphics, and Image Processing, 27(3):346–368, September 1984. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chang:1987:ODE

[CK87] Edward S. H. Chang and Ludwik Kurz. Object detection and experimental designs. Computer Vision, Graphics, and Image Processing, 40 (2):147–168, November 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chen:2000:VRU

[CK00] Baoquan Chen and Arie Kaufman. 3D volume rotation using shear transformations. Graphical Models, 62(4):308–322, July 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0525;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0525/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0525/ref.

Choi:2009:RFP

[CK09] Ouk Choi and In So Kweon. Robust feature point matching by preserving local geometric consistency. Computer Vision and Image Understanding: CVIU, 113(6): 726–742, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chang:2011:MSS

[CK11] Ming-Ching Chang and Benjamin B. Kimia. Measuring 3D shape similarity by graphbased matching of the medial scaffolds. Computer Vision and Image Understanding: CVIU, 115(5):707–720, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Califano:1996:DMD

[CKB96] Andrea Califano, Rick Kjeldsen, and Ruud M. Bolle.

Data- and model-driven multiresolution processing. Computer Vision and Image Understanding: CVIU, 63 (1):27–49, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0003/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0003/production/
pdf.

Chai:2010:SIE

[CKB10] Sek Chai, Branislav Kisačanin, and Nikolaos Bellas. Special issue on embedded vision. Computer Vision and Image Understanding: CVIU, 114(11):1115, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chun:2012:ACM

[CKK^{+}12] Jinhee Chun, Natsuda Kaothanthong, Ryosei Kasai, Matias Korman, Martin Nöllenburg, and Takeshi Tokuyama. Algorithms for computing the maximum weight region decomposable into elementary shapes. Computer Vision and Image Understanding: CVIU, 116(7):803–814, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000501.

Chiu:2009:LGN

[CKLP09] Han-Pang Chiu, Leslie Pack Kaelbling, and Tomás Lozano-Pérez. Learning to generate novel views of objects for class recognition. Computer Vision and Image Understanding: CVIU, 113(12): 1183–1197, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Castle:2011:WAA

[CKM11] Robert O. Castle, Georg Klein, and David W. Murray. Wide-area augmented reality using camera tracking and mapping in multiple regions. Computer Vision and Image Understanding: CVIU, 115(6):854–867, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Criminisi:2005:ELA

[CKS^{+}05] Antonio Criminisi, Sing Bing Kang, Rahul Swaminathan, Richard Szeliski, and P. Anandan. Extracting layers and analyzing their specular properties using epipolar-plane-image analysis. Computer Vision and Image Understanding: CVIU, 97(1):51–85, January 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cantoni:1983:MTI

[CL83] V. Cantoni and S. Levialdi. Matching the task to an image processing architecture. Computer Vision, Graphics, and Image Processing, 22(2):301–309, May 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Clark:1986:TBD

[CL86] James J. Clark and Peter D. Lawrence. A theoretical basis for diffrequency stereo. Computer Vision, Graphics, and Image Processing, 35(1):1–19, July 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chang:1990:FAR

[CL90] Long-Wen Chang and Kuen-Long Leu. A fast algorithm for the restoration of images based on chain codes description and its applications. Computer Vision, Graphics, and Image Processing, 50(3):296–307, June 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chung:1991:COP

[CL91] Kuo-Liang Chung and Ferng-Ching Lin. A cost-optimal parallel algorithm for B-spline surface fitting. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6): 601–605, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Chung:1995:HTR

[CL95] Kuo-Liang Chung and Horn-Yi Lin. Hough transform on reconfigurable meshes. Computer Vision and Image Understanding: CVIU, 61(2):278–284, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1020/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1020/production/
pdf.

Chen:1997:DMS

[CL97a] C. H. Chen and G. G. Lee. On digital mammogram segmentation and microcalcification detection using multiresolution wavelet analysis. Graphical Models and Image Processing: GMIP, 59(5):349–364, September 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0443/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0443/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0443/production/ref.

Crevier:1997:KBI

[CL97b] Daniel Crevier and Richard Lepage. Knowledge-based image understanding systems: a survey. Computer Vision and Image Understanding: CVIU, 67(2):161–185, August 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0520/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0520/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0520/production/ref.

Chan:1998:CBW

[CL98] Kwai Hung Chan and Rynson W. H. Lau. Contour-based warping. Graphical Models and Image Processing: GMIP, 60(5):331–348, September 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0476/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0476/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0476/production/ref.

Chen:2000:SEC

[CL00a] Fei-Long Chen and Shiaur-Wehn Lin. Subpixel estimation of circle parameters using orthogonal circular detector. Computer Vision and Image Understanding: CVIU, 78(2):206–221, May 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0836;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0836/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0836/ref.

Choi:2000:ICU

[CL00b] Yongchoel Choi and Seungyong Lee. Injectivity conditions of 2D and 3D uniform cubic B-spline functions. Graphical Models, 62(6):411–427, November 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2000.0531;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/gmod.2000.0531/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2000.0531/ref.

Chazal:2005:MA

[CL05] Frédéric Chazal and André Lieutier. The “λ-medial axis”. Graphical Models, 67 (4):304–331, July 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Chi:2008:GSC

[CL08] Yanling Chi and Maylor K. H. Leung. A general shape context framework for object identification. Computer Vision and Image Understanding: CVIU, 112(3): 324–336, December 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:1991:SMI

[CLC91] Shiuh-Yung Chen, Wei-Chung Lin, and Chin-Tu Chen. Splitand-merge image segmentation based on localized feature analysis and statistical tests. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):457– 475, September 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Cho:2013:EGG

[CLCO13] Jungchan Cho, Minsik Lee, Chong-Ho Choi, and Songhwai Oh. EM–GPA: Generalized Procrustes analysis with hidden variables for 3D shape modeling. Computer Vision and Image Understanding: CVIU, 117(11): 1549–1559, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001392.

Chen:1996:PVM

[CLD96] Yuan Chen, Noshir A. Langrana, and Atish K. Das. Perfecting vectorized mechanical drawings. Computer Vision and Image Understanding: CVIU, 63(2):273–286, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0019/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0019/production/
pdf.

Chow:1994:NDA

[CLHW94] L. R. Chow, H. C. Liu, S. Y. Hsu, and D. W. Wu. A new dynamic approach for finding the contour of bi-level images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(6):507–509, November 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1045/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1045/production/
pdf.

Chang:2009:SRP

[CLK09] Ming-Ching Chang, Frederic Fol Leymarie, and Benjamin B. Kimia. Surface reconstruction from point clouds by transforming the medial scaffold. Computer Vision and Image Understanding: CVIU, 113(11):1130– 1146, November 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:1999:SAU

[CLL^{+}99] H.-Y. Chen, I.-K. Lee, S. Leopoldseder, H. Pottmann, T. Randrup, and J. Wallner. On surface approximation using developable surfaces. Graphical Models and Image Processing: GMIP, 61(2):110–124, March 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0487/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0487/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0487/production/ref.

Cohen:1980:DBS

[CLR80] Elaine Cohen, Tom Lyche, and Richard F. Riesenfeld. Discrete B-splines and subdivision techniques in computer-aided geometric design and computer graphics. Computer Graphics and Image Processing, 14(2):87– 111, October 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chen:2013:OCI

[CLZZ13] Xiaowu Chen, Qing Li, Dongyue Zhao, and Qinping Zhao. Occlusion cues for image scene layering. Computer Vision and Image Understanding: CVIU, 117(1):42–55, January 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001300.

Chen:1992:AVP

[CM92] Chien-Huei Chen and P. G. Mulgaonkar. Automatic vision programming. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(2):170– 183, March 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Chaudhury:1994:OFE

[CM94a] Krishnendu Chaudhury and Rajiv Mehrotra. Optical flow estimation using smoothness of intensity trajectories. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):230–244, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1049/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1049/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1054/production;   http:

//www.idealibrary.com/links/
artid/cviu.1994.1054/production/
pdf.

Christensen:1994:PRR

[CM94b] Henrik I. Christensen and Claus B. Madsen. Purposive reconstruction: a reply to “A Computational and Evolutionary Perspective on the Role of Representation in Vision” by M. J. Tarr and M. J. Black. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):103–108, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1039/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1039/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1044/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1044/production/

pdf. See [TB94a].

Chen:1995:DCO

[CM95] Yang Chen and Gérard Medioni. Description of complex objects from multiple range images using an inflating balloon model. Computer Vision and Image Understanding: CVIU, 61(3):325–334, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1995.
                                        1026/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1995.1026/production/
                                        pdf.

Cho:1997:ISC

[CM97] Kyujin Cho and Peter Meer. Image segmentation from consensus information. Computer Vision and Image Understanding: CVIU, 68 (1):72–89, October 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0546/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0546/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0546/production/ref.

Casadei:1999:BUA

[CM99a] Stefano Casadei and Sanjoy Mitter. Beyond the uniqueness assumption: Ambiguity representation and redundancy elimination in the computation of a covering sample of salient contour cycles. Computer Vision and Image Understanding: CVIU, 76(1):19–35, October 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0790/production;   http:

//www.idealibrary.com/links/
artid/cviu.1999.0790/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0790/production/ref.

Cuisenaire:1999:FED

[CM99b] O. Cuisenaire and B. Macq.

Fast Euclidean distance transformation by propagation using multiple neighborhoods. Computer Vision and Image Understanding: CVIU, 76 (2):163–172, November 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0783/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0783/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0783/production/ref.

Cao:2012:IFE

[CM12] Yanpeng Cao and John Mc-Donald. Improved feature extraction and matching in urban environments based on 3D viewpoint normalization. Computer Vision and Image Understanding: CVIU, 116(1):86–101, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001974.

Chica:2012:EGS

[CMB^{+}12] Antoni Chica, Eva Monclús, Pere Brunet, Isabel Navazo, and Àlvar Vinacua. Exampleguided segmentation. Graphical Models, 74(6):302–310, November 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000094.

Cao:2009:HMR

[CMBP09] Dongwei Cao, Osama T. Masoud, Daniel Boley, and Nikolaos Papanikolopoulos. Human motion recognition using support vector machines. Computer Vision and Image Understanding: CVIU, 113(10):1064–1075, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Calderon:2004:MMA

[CMBV04] Felix Calderon, Jose L. Marroquin, Salvador Botello, and Baba C. Vemuri. The MPM– MAP algorithm for motion segmentation. Computer Vision and Image Understanding: CVIU, 95(2):165– 183, August 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Christoudias:2006:NPL

[CMD06] C. Mario Christoudias, Louis-Philippe Morency, and Trevor Darrell. Non-parametric and light-field deformable models. Computer Vision and Image Understanding: CVIU, 104 (1):16–35, October 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheriet:2013:LFO

[CMH13] Mohamed Cheriet, Reza Farrahi Moghaddam, and Rachid Hedjam. A learning framework for the optimization and automation of document binarization methods. Computer Vision and Image Understanding: CVIU, 117(3):269–280, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001877.

Cho:1999:TRA

[CMPP99] Wonjoon Cho, Takashi Maekawa, Nicholas M. Patrikalakis, and Jaime Peraire. Topologically reliable approximation of trimmed polynomial surface patches. Graphical Models and Image Processing: GMIP, 61(2):84–109, March 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0483/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0483/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0483/production/ref.

Cignoni:1998:ZRM

[CMRS98] P. Cignoni, C. Montani, C. Rocchini, and R. Scopigno. Zeta: a resolution modeling system. Graphical Models and Image Processing: GMIP, 60(5):305–329, September 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0477/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1998.0477/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0477/production/ref.

Cheevasuvit:1986:RMP

[CMVM86] Fusak Cheevasuvit, Henri Maitre, and Daniel Vidal-Madjar. A robust method for picture segmentation based on a split-and-merge procedure. Computer Vision, Graphics, and Image Processing, 34(3):268–281, June 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chandrasekaran:1997:EUA

[CMW^{+}97] S. Chandrasekaran, B. S. Manjunath, Y. F. Wang, J. Winkeler, and H. Zhang. An eigenspace update algorithm for image analysis. Graphical Models and Image Processing: GMIP, 59(5):321–332, September 1997. CODEN GMIPF4.

ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0425/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0425/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0425/production/ref.

Caelli:1987:FEO

[CN87a] Terry Caelli and Shyamala Nagendran. Fast edge-only matching techniques for robot pattern recognition. Computer Vision, Graphics, and Image Processing, 39(2):131– 143, August 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Conradsen:1987:DDF

[CN87b] Knut Conradsen and Gert Nilsson. Data dependent filters for edge enhancement of Landsat images. Computer Vision, Graphics, and Image Processing, 38(2):101– 121, May 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chung:1995:UMG

[CN95] Ronald Chung and Ramakant Nevatia. Use of monocular groupings and occlusion analysis in a hierarchical stereo system. Computer Vision and Image Understanding: CVIU, 62(3):245–268, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1995.
                                        1053/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1995.1053/production/
                                        pdf.

Cunado:2003:AED

[CNC03] David Cunado, Mark S. Nixon, and John N. Carter. Automatic extraction and description of human gait models for recognition purposes. Computer Vision and Image Understanding: CVIU, 90 (1):1–41, April 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cousty:2013:MFG

[CNDS13] Jean Cousty, Laurent Najman, Fabio Dias, and Jean Serra. Morphological filtering on graphs. Computer Vision and Image Understanding: CVIU, 117(4):370–385, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421200183X.

Chen:2001:SIV

[CNK01] Min Chen, Gregory M. Nielson, and Arie E. Kaufman. Special issue on volume modeling. Graphical Models, 63 (6):385–386, November 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Celik:2008:FFE

[ÇÖD08] Turgay Çelik, Hüseyin Özkaramanlı, and Hasan Demirel. Facial feature extraction using complex dual-tree wavelet transform. Computer Vision and Image Understanding: CVIU, 111(2):229–246, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Coeurjolly:2012:FAA

[Coe12] David Coeurjolly. Fast and accurate approximation of digital shape thickness distribution in arbitrary dimension. Computer Vision and Image Understanding: CVIU, 116(12):1159–1167, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001178.

Cohen:1985:GSF

[Coh85] Edgar A. Cohen, Jr. Generalized sloped facet models useful in multispectral image analysis. Computer Vision, Graphics, and Image Processing, 32(2):171– 190, November 1985. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Cohen:1991:ACM

[Coh91] L. D. Cohen. On active contour models and balloons. Computer Vision, Graphics, and Image Processing. Image Understanding, 53(2): 211–218, March 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Cohen-Or:1995:FSV

[COK95] Daniel Cohen-Or and Arie Kaufman. Fundamentals of surface voxelization. Graphical Models and Image Processing: GMIP, 57(6):453– 461, November 1995. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1039/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1039/production/
pdf.

Colsher:1977:ITD

[Col77]  James G. Colsher.  Iter-

ative three-dimensional image reconstruction from tomographic projections. Computer Graphics and Image Processing, 6(6):513– 537, December 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Collard:1981:LPL

[Col81] R. F. A. Collard. Local picture languages. Computer Graphics and Image Processing, 17(4):332– 344, December 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Collins:1997:GVS

[Col97] Robert T. Collins. The geometry of visual space: About the incompatibility between science and mathematics — reply. Computer Vision and Image Understanding: CVIU, 65(3):434–435, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0493/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0493/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0493/production/ref.

Conker:1988:DPV

[Con88] Robert S. Conker. A dual plane variation of the Hough transform for detecting nonconcentric circles of different radii. Computer Vision, Graphics, and Image Processing, 43(2):115–132, August 1988. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Coueignoux:1981:CGC

[Cou81a]  Ph. Coueignoux.  Char-

acter generation by computer. Computer Graphics and Image Processing, 16(3): 240–269, July 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Coueignoux:1981:PSR

[Cou81b] Ph. Coueignoux. A posteriori scaling of run length encoded polygons. Computer Graphics and Image Processing, 17(1): 84–89, September 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Couprie:2013:TMR

[Cou13] Michel Couprie. Topological maps and robust hierarchical Euclidean skeletons in cubical complexes. Computer Vision and Image Understanding: CVIU, 117(4):355–369, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001762.

Caelli:1998:SMI

[COW98] Terry Caelli, Erol Osman, and Geoff West. 3D shape matching and inspection using geometric features and relational learning. Computer Vision and Image Understanding: CVIU, 72(3):340–350, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.

0659/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0659/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0659/production/ref.

Chen:1979:STU

[CP79] Patrick C. Chen and Theodosios Pavlidis. Segmentation by texture using a co-occurrence matrix and a split-and-merge algorithm. Computer Graphics and Image Processing, 10(2):172– 182, June 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). CHEN79.

Chen:1980:ISE

[CP80] P. C. Chen and T. Pavlidis.

Image segmentation as an estimation problem. Computer Graphics and Image Processing, 12(2):153– 172, February 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). CHEN80.

Costabile:1981:BDA

[CP81] M. F. Costabile and G. G.

Pieroni. Boundary detection algorithms in nuclear medicine imagery. Computer Graphics and Image Processing, 17(4):362– 374, December 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chen:1986:SMN

[CP86] Su-Shing S. Chen and Michael Penna. Shape and motion of nonrigid bodies. Computer Vision, Graphics, and Image Processing, 36(2/3):175–207, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Cohen:1991:MSI

[CP91] F. S. Cohen and A. S. Patel. Modeling and synthesis of images of 3D textured surfaces. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6):501– 510, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Cong:1999:ASS

[CP99] G. Cong and B. Parvin. An algebraic solution to surface recovery from cross-sectional contours. Graphical Models and Image Processing: GMIP, 61(4):222–243, July 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0499/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0499/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0499/production/ref.

Cheong:2004:DDU

[CP04] Loong-Fah Cheong and Chin-Hwee Peh. Depth distortion under calibration uncertainty. Computer Vision and Image Understanding: CVIU, 93(3): 221–244, March 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cho:2009:FAD

[CP09] Minkook Cho and Hyeyoung Park. A feature analysis for dimension reduction based on a data generation model with class factors and environment factors. Computer Vision and Image Understanding: CVIU, 113(9):1005–1016, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chan:1999:BSR

[CPC99] Moses W. Chan, Zygmunt Pizlo, and David M. Chelberg. Binocular shape reconstruction: Psychological plausibility of the 8-point algorithm. Computer Vision and Image Understanding: CVIU, 74(2):121–137, May 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0748/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0748/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/cviu.1999.
                                        0748/production/ref.

Calderara:2008:HHE

[CPC08] Simone Calderara, Andrea Prati, and Rita Cucchiara. HECOL: Homography and epipolar-based consistent labeling for outdoor park surveillance. Computer Vision and Image Understanding: CVIU, 111(1):21–42, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cornilleau-Peres:1993:VBC

[CPD93] Valérie Cornilleau-Pérès and Jacques Droulez. Velocitybased correspondence in stereokinetic images. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(2):137–146, September 1993. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1993.
                                        1034/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1993.1034/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1993.
                                        1036/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1993.1036/production/
                                        pdf.

Choi:1999:PMC

[CPK99] Kwang-Jin Choi, Sang-Hyun Park, and Hyeong-Seok Ko.

Processing motion capture data to achieve positional accuracy. Graphical Models and Image Processing: GMIP, 61(5):260–273, September 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0505/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0505/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0505/production/ref.

Cani:2009:SS

[CPOO09] Marie-Paule Cani, Fred Pighin, James F. O’Brien, and Carol O’Sullivan. SCA 2006 Symposium. Graphical Models, 71(6):197, November 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000277.

Chang:2011:GFM

[CPP^{+}11] Ju Yong Chang, Haesol Park, In Kyu Park, Kyoung Mu Lee, and Sang Uk Lee. GPUfriendly multi-view stereo reconstruction using surfel representation and graph cuts. Computer Vision and Image Understanding: CVIU, 115 (5):620–634, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chum:2005:GEH

[CPS05] Ondřej Chum, Tomáš Pajdla, and Peter Sturm. The geometric error for homographies. Computer Vision and Image Understanding: CVIU, 97(1):86–102, January 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Craciun:2010:MVS

[CPS10] Daniela Craciun, Nicolas Paparoditis, and Francis Schmitt. Multi-view scans alignment for 3D spherical mosaicing in large-scale unstructured environments. Computer Vision and Image Understanding: CVIU, 114(11):1248–1263, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cheng:2007:MSM

[CPT07] Shinko Y. Cheng, Sangho Park, and Mohan M. Trivedi. Multi-spectral and multiperspective video arrays for driver body tracking and activity analysis. Computer Vision and Image Understanding: CVIU, 106(2–3):245– 257, May/June 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cohen:1988:IPB

[CR88] E. Cohen and R. F. Riesenfeld. An incompatibility projector based on an interpolant of Gregory. Computer Graphics and Image Processing, 8 (2):294–298, March 1988. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Cappelletti:1989:TDB

[CR89] John Danilo Cappelletti and Azriel Rosenfeld. Threedimensional boundary following. Computer Vision, Graphics, and Image Processing, 48(1):80–92, October 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Connelly:1990:PAF

[CR90] Steven Connelly and Azriel Rosenfeld. A pyramid algorithm for fast curve extraction. Computer Vision, Graphics, and Image Processing, 49(3):332–345, March 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chinen:1997:PAF

[CR97] Troy T. Chinen and Todd R.

Reed. A performance analysis of fast Gabor transform methods. Graphical Models and Image Processing: GMIP, 59(3):117–127, May 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0421/production;   http:
//www.idealibrary.com/links/

                                        artid/gmip.1997.0421/production/
                                        pdf.

Chui:2003:NPM

[CR03] Haili Chui and Anand Rangarajan. A new point matching algorithm for non-rigid registration. Computer Vision and Image Understanding: CVIU, 89(2–3):114–141, February/March 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chatterjee:1997:NGS

[CRC97] Chanchal Chatterjee, Vwani P. Roychowdhury, and Edwin K. P. Chong. A nonlinear Gauss–Seidel algorithm for noncoplanar and coplanar camera calibration with convergence analysis. Computer Vision and Image Understanding: CVIU, 67(1):58–80, July 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0516/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0516/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0516/production/ref.

Crevier:1999:PME

[Cre99] Daniel Crevier. A probabilistic method for extracting chains of collinear segments. Computer Vision and Image Understanding:

CVIU, 76(1):36–53, October 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0785/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0785/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0785/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0787/production;
 http://www.idealibrary.
com/links/artid/cviu.1999.
0787/production/pdf;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0787/production/
ref.

Crevier:2008:ISA

[Cre08] Daniel Crevier. Image segmentation algorithm development using ground truth image data sets. Computer Vision and Image Understanding: CVIU, 112(2): 143–159, November 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Collomosse:2005:RCS

[CRH05] J. P. Collomosse, D. Rowntree, and P. M. Hall. Rendering cartoon-style motion cues in post-production video. Graphical Models, 67(6):549– 564, November 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000111.

Comon:1990:SIA

[CRT90] P. Comon, Y. Robert, and D. Trystram. Systolic implementation of the adaptive solution to normal equations. Computer Vision, Graphics, and Image Processing, 52 (3):402–408, December 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Crettez:1982:MCR

[CS82] J. P. Crettez and J. C. Simon. Model for cell receptive fields in the visual striate cortex. Computer Graphics and Image Processing, 20(4): 299–318, December 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Chen:1989:RDM

[CS89] Lih-Shyang S. Chen and Marc R. Sontag. Representation, display, and manipulation of 3D digital scenes and their medical applications. Computer Vision, Graphics, and Image Processing, 48 (2):190–216, November 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chen:1998:FFO

[CS98] Jin-Long Chen and George C. Stockman. 3D free-form object recognition using indexing by contour features. Computer Vision and Image Understanding: CVIU, 71(3):334–355, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0648/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0648/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0648/production/ref.

Costa:2000:ORP

[CS00] Mauro S. Costa and Linda G.

Shapiro. 3D object recognition and pose with relational indexing. Computer Vision and Image Understanding: CVIU, 79(3): 364–407, September 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0865;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0865/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0865/ref.

Choi:2001:HHD

[CS01] Sung Woo Choi and Hans-Peter Seidel. Hyperbolic Hausdorff distance for medial axis transform. Graphical Models, 63(5):369–384, September 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Chau:2004:ADP

[CS04] Chun-Pong Chau and Wan-Chi Siu. Adaptive dual-point Hough transform for object recognition. Computer Vision and Image Understanding: CVIU, 96(1):1–16, October 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Coughlan:2007:DQB

[CS07] James Coughlan and Huiying Shen. Dynamic quantization for belief propagation in sparse spaces. Computer Vision and Image Understanding: CVIU, 106(1): 47–58, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2010:PEM

[CS10] Chong Chen and Dan Schonfeld. Pose estimation from multiple cameras based on Sylvester’s equation. Computer Vision and Image Understanding: CVIU, 114(6): 652–666, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chaudhuri:1996:REM

[CSC96] S. Chaudhuri, S. Sharma, and S. Chatterjee. Recursive estimation of motion parameters. Computer Vision and Image Understanding: CVIU, 64(3):434–442, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0070/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0070/production/
pdf.

Chen:1996:TSD

[CSDC96] S. W. Chen, G. Stockman, C. Y. Dai, and C. P. Chuang. Two-stage dynamic deformation for construction of 3D models. Graphical Models and Image Processing: GMIP, 58(5):484–493, September 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0040/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0040/production/
pdf.

Cohen:2003:FER

[CSG^{+}03] Ira Cohen, Nicu Sebe, Ashutosh Garg, Lawrence S. Chen, and Thomas S. Huang. Facial expression recognition from video sequences: temporal and static modeling. Computer Vision and Image Understanding: CVIU, 91(1–2): 160–187, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chao:2013:VBP

[CSJ13] Qianwen Chao, Jingjing Shen, and Xiaogang Jin. Videobased personalized traffic learning. Graphical Models, 75(6):305–317, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000234.

Cowart:1983:DUT

[CSR83] Alan E. Cowart, Wesley E. Snyder, and W. Howard Ruedger. Detection of unresolved targets using the Hough transform. Computer Vision, Graphics, and Image Processing, 21(2):222– 238, February 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chamveha:2013:HDE

[CSS^{+}13a] Isarun Chamveha, Yusuke Sugano, Daisuke Sugimura, Teera Siriteerakul, Takahiro Okabe, Yoichi Sato, and Akihiro Sugimoto. Head direction estimation from low resolution images with scene adaptation. Computer Vision and Image Understanding: CVIU, 117(10):1502–1511, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001094.

Chessa:2013:ALM

[CSS13b] Manuela Chessa, Fabio Solari, and Silvio P. Sabatini. Adjustable linear models for optic flow based obstacle avoidance. Computer Vision and Image Understanding: CVIU, 117(6):603–619, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000283.

Cao:2008:RIC

[CSY08] Aize Cao, Qing Song, and Xulei Yang. Robust information clustering incorporating spatial information for breast mass detection in digitized mammograms. Computer Vision and Image Understanding: CVIU, 109(1): 86–96, January 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:1988:MPS

[CT88] Ling-Hwei Chen and Wen-Hsiang Tsai. Momentpreserving sharpening — a new approach to digital picture deblurring. Computer Vision, Graphics, and Image Processing, 41(1):1–13, January 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Clement:1993:KBA

[CT93] Véronique Clément and Monique Thonnat. A knowledgebased approach to integration of image processing procedures. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):166–184, March 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1011/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1011/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1011/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1011/production/
pdf.

Chan:1995:CRL

[CT95] Y. T. Chan and Samuel M. Thomas. Cramer-Rao lower bounds for estimation of a circular arc center and its radius. Graphical Models and Image Processing: GMIP, 57 (6):527–532, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic).

Chan:1997:AML

[CT97] Y. T. Chan and S. M. Thomas. An approximate maximum likelihood linear estimator of circle parameters. Graphical Models and Image Processing:

GMIP, 59(3):173–178, May 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0424/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0424/production/
pdf.

Celik:2010:UCI

[CT10] Turgay Celik and Tardi Tjahjadi. Unsupervised colour image segmentation using dualtree complex wavelet transform. Computer Vision and Image Understanding: CVIU, 114(7):813–826, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Celik:2012:ACC

[CT12] Turgay Celik and Tardi Tjahjadi. Adaptive colour constancy algorithm using discrete wavelet transform. Computer Vision and Image Understanding: CVIU, 116(4):561–571, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002712.

Cootes:1995:ASM

[CTCG95] T. F. Cootes, C. J. Taylor, D. H. Cooper, and J. Graham. Active shape models — their training and application. Computer Vision and Image Understanding: CVIU, 61(1):38–59, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1004/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1004/production/
pdf.

Ching:1995:RVC

[CTE95] Wee-Soon Ching, Peng-Seng Toh, and Meng-Hwa Er. Robust vergence with concurrent identification of occlusion and specular highlights. Computer Vision and Image Understanding: CVIU, 62(3):298–308, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1056/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1056/production/
pdf.

Chen:1998:EAS

[CTF^{+}98] S. W. Chen, S. T. Tung, C. Y. Fang, Shen Cherng, and Anil K. Jain. Extended attributed string matching for shape recognition. Computer Vision and Image Understanding: CVIU, 70(1):36–50, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.

com/links/artid/cviu.1998.
0599/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0599/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0599/production/ref.

Conners:1984:SHR

[CTH84] Richard W. Conners, Mohan M. Trivedi, and Charles A. Harlow. Segmentation of a high-resolution urban scene using texture operators. Computer Vision, Graphics, and Image Processing, 25(3):273– 310, March 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Cerutti:2013:ULN

[CTM^{+}13] Guillaume Cerutti, Laure Tougne, Julien Mille, Antoine Vacavant, and Didier Coquin. Understanding leaves in natural images

— a model-based approach for tree species identification. Computer Vision and Image Understanding: CVIU, 117(10):1482–1501, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001331.

Ciesielski:2010:AFFa

[CU10a] Krzysztof Chris Ciesielski and Jayaram K. Udupa. Affinity functions in fuzzy connectedness based image segmentation I: Equivalence of affinities. Computer Vision and Image Understanding: CVIU, 114(1):146–154, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ciesielski:2010:AFFb

[CU10b] Krzysztof Chris Ciesielski and Jayaram K. Udupa. Affinity functions in fuzzy connectedness based image segmentation II: Defining and recognizing truly novel affinities. Computer Vision and Image Understanding: CVIU, 114(1): 155–166, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ciesielski:2011:FCD

[CU11] Krzysztof Chris Ciesielski and Jayaram K. Udupa. A framework for comparing different image segmentation methods and its use in studying equivalences between level set and fuzzy connectedness frameworks. Computer Vision and Image Understanding: CVIU, 115(6):721–734, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2013:GAS

[CUAT13] Xinjian Chen, Jayaram K. Udupa, Abass Alavi, and Drew A. Torigian. GC-ASM: Synergistic integration of graph-cut and active shape model strategies for medical image segmentation. Computer Vision and Image Understanding: CVIU, 117(5):513–524, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001956.

Cumani:1991:EDM

[Cum91] A. Cumani. Edge detection in multispectral images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1):40–51, January 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Ciesielski:2007:IRF

[CUSZ07] Krzysztof Chris Ciesielski, Jayaram K. Udupa, Punam K. Saha, and Ying Zhuge. Iterative relative fuzzy connectedness for multiple objects with multiple seeds. Computer Vision and Image Understanding: CVIU, 107(3):160– 182, September 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Campani:1992:MAF

[CV92] Marco Campani and Alessandro Verri. Motion analysis from first-order properties of optical flow. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):90–107, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Cole:1994:DRC

[CV94] R. Cole and U. Vishkin. On the detection of robust curves. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(3):189–204, May 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1018/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1018/production/
                                        pdf.

Chellappa:2013:SSC

[CV13] Rama Chellappa and Baba Vemuri. Special section in celebration of Professor J. K. Aggarwal. Computer Vision and Image Understanding: CVIU, 117(10):1203, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213001410.

Chaouch:2009:AM

[CVB09] Mohamed Chaouch and Anne Verroust-Blondet. Alignment of 3D models. Graphical Models, 71(2):63–76, March 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000046.

Casares:2010:LWS

[CVP10] Mauricio Casares, Senem Velipasalar, and Alvaro Pinto. Light-weight salient foreground detection for embedded smart cameras. Computer Vision and Image Understanding: CVIU, 114(11): 1223–1237, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Calway:1994:CEI

[CW94] A. D. Calway and R. Wilson. Curve extraction in images using a multiresolution framework. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):359–366, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1025/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1025/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1029/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1029/production/
pdf.

Cui:2000:ABH

[CW00] Yuntao Cui and Juyang Weng. Appearance-based hand sign recognition from intensity image sequences. Computer Vision and Image Understanding: CVIU, 78(2):157–176, May 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0837;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0837/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0837/ref.

Chen:2002:BPR

[CW02] Falai Chen and Wenping Wang. The μ-basis of a planar rational curve–properties and computation. Graphical Models, 64(6):368–381, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Cui:1994:RBE

[CWC94] Ning Cui, John Juyang Weng, and Paul Cohen. Recursive-batch estimation of motion and structure from monocular image sequences. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):154–170, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1010/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1010/production/

pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1012/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1012/production/
pdf.

Cao:2013:UDL

[CWH^{+}13] Xiaochun Cao, Xingxing Wei, Yahong Han, Yi Yang, Nicu Sebe, and Alexander Hauptmann. Unified dictionary learning and region tagging with hierarchical sparse representation. Computer Vision and Image Understanding: CVIU, 117(8):934–946, August 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000623.

Choi:2013:STP

[CWLJ13] Jaesik Choi, Ziyu Wang, Sang-Chul Lee, and Won J. Jeon. A spatio-temporal pyramid matching for video retrieval. Computer Vision and Image Understanding: CVIU, 117(6):660–669, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000325.

Chen:2011:BPS

[CWO^{+}11] Shi Chen, Jinqiao Wang, Yi Ouyang, Bo Wang, Changsheng Xu, and Hanqing Lu. Boosting part-sense multifeature learners toward effective object detection. Computer Vision and Image Understanding: CVIU, 115(3): 364–374, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chin:1987:OPT

[CWSI87] Roland T. Chin, Hong-Khoon Wan, D. L. Stover, and R. D. Iverson. A one-pass thinning algorithm and its parallel implementation. Computer Vision, Graphics, and Image Processing, 40(1):30– 40, October 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Cheong:2011:BSA

[CX11] Loong-Fah Cheong and Xu Xiang. Behaviour of SFM algorithms with erroneous calibration. Computer Vision and Image Understanding: CVIU, 115(1):16–30, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Cao:2006:SCT

[CXFS06] Xiaochun Cao, Jiangjian Xiao, Hassan Foroosh, and Mubarak Shah. Self-calibration from turn-table sequences in presence of zoom and focus. Computer Vision and Image Understanding: CVIU, 102 (3):227–237, June 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2009:CMD

[CXY^{+}09] Xiao-Diao Chen, Gang Xu, Jun-Hai Yong, Guozhao Wang, and Jean-Claude Paul. Computing the minimum distance between a point and a clamped B-spline surface. Graphical Models, 71 (3):107–112, May 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000071.

Chang:1983:PIM

[CY83a] Shi-Kuo Chang and Chung-Chun Yang. Picture information measures for similarity retrieval. Computer Vision, Graphics, and Image Processing, 23(3):366– 375, September 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Chin:1983:QES

[CY83b] Roland T. Chin and Chia-Lung Yeh. Quantitative evaluation of some edgepreserving noise-smoothing techniques. Computer Vision, Graphics, and Image Processing, 23(1):67–91, July 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Choi:2010:ASE

[CYC10] JinMin Choi, Yung Jun Yoo, and Jin Young Choi. Adaptive shadow estimator for removing shadow of moving object. Computer Vision and Image Understanding: CVIU, 114(9):1017–1029, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Crabtree:1991:FAE

[CYE91] J. Crabtree, Jr., Li-Ping Yuan, and R. Ehrlich. A fast and accurate erosion-dilation method suitable for microcomputers. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53 (3):283–290, May 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Coughlan:2000:EDT

[CYES00] James Coughlan, Alan Yuille, Camper English, and Dan Snow. Efficient deformable template detection and localization without user initialization. Computer Vision and Image Understanding: CVIU, 78(3):303–319, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0842;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0842/

pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0842/ref.

Chen:1994:CCI

[CYH94] Yung-Sheng S. Chen, Hung-Tien T. Yen, and Wen-Hsing H. Hsu. Compression of color image via the technique of surface fitting. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(3):272–279, May 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1024/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1024/production/
pdf.

Chen:2011:HPR

[CYNO11] Cheng Chen, Yi Yang, Feiping Nie, and Jean-Marc Odobez. 3D human pose recovery from image by efficient visual feature selection. Computer Vision and Image Understanding: CVIU, 115(3): 290–299, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Chen:2010:CHP

[CYP^{+}10] Chung-Hao Chen, Yi Yao, David Page, Besma Abidi, Andreas Koschan, and Mongi Abidi. Camera handoff and placement for automated tracking systems with multiple omnidirectional cameras. Computer Vision and Image Understanding: CVIU, 114 (2):179–197, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Che:2004:CSD

[CYW04a] WuJun Che, XunNian Yang, and GuoZhao Wang. Corrigendum to “Skeleton-driven 2D distance field metamorphosis using intrinsic shape parameters” [Graphical Models 66 (2004) 102–126]. Graphical Models, 66(4):261, July 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Che:2004:SDD

[CYW04b] WuJun Che, XunNian Yang, and GuoZhao Wang. Skeletondriven 2D distance field metamorphosis using intrinsic shape parameters. Graphical Models, 66(2):102–126, March 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Csurka:1997:CUF

[CZZF97] Gabriella Csurka, Cyril Zeller, Zhengyou Zhang, and Olivier D. Faugeras. Characterizing the uncertainty of the fundamental matrix. Computer Vision and Image Understanding: CVIU, 68(1):18–36, October 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0531/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0531/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0531/production/ref.

Cui:2007:LBD

[CZZS07] Jinshi Cui, Hongbin Zha, Huijing Zhao, and Ryosuke Shibasaki. Laser-based detection and tracking of multiple people in crowds. Computer Vision and Image Understanding: CVIU, 106(2– 3):300–312, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Damiand:2008:TMI

[Dam08] Guillaume Damiand. Topological model for 3D image representation: Definition and incremental extraction algorithm. Computer Vision and Image Understanding: CVIU, 109(3):260–289, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Diop:2012:INM

[DAM12] El Hadji S. Diop, Radjesvarane Alexandre, and Lionel Moisan. Intrinsic nonlinear multiscale image decomposition: a 2D empirical mode decomposition-like tool. Computer Vision and Image Understanding: CVIU, 116 (1):102–119, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001986.

Danielsson:1978:CCG

[Dan78a] P.-E. Danielsson. Comments on circle generator for display devices. Computer Graphics and Image Processing, 7(??): 300–301, ???? 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Danielsson:1978:NSF

[Dan78b] Per Erik Danielsson. New shape factor. Computer Graphics and Image Processing, 7(2):292–299, April 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Danielsson:1980:EDM

[Dan80] Per Erik Danielsson. Euclidean distance mapping. Computer Graphics and Image Processing, 14(3):227– 248, November 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Danielsson:1981:GMF

[Dan81a] Per Erik Danielsson. Getting the median faster. Computer Graphics and Image Processing, 17(1):71–78, September 1981. CODEN CGIPBG.

ISSN 0146-664X (print), 1557-9697 (electronic).

Danielsson:1981:IKS

[Dan81b] Per Erik Danielsson. Improvement of Kruse’s segmentation algorithm. Computer Graphics and Image Processing, 17 (4):394–396, December 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Daniilidis:1997:FSM

[Dan97] Konstantinos Daniilidis. Fixation simplifies 3D motion estimation. Computer Vision and Image Understanding: CVIU, 68(2):158–169, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0535/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0535/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0535/production/ref.

Davies:1975:SED

[Dav75] L. S. Davies. A survey of edge detection techniques. Computer Graphics and Image Processing, 4(??):248– 270, ???? 1975. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). DAVIES75.

Davis:1976:SED

[Dav76] L. Davis. A survey of edge detection techniques. Computer Graphics and Image Processing, ??(??):??, ???? 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Davis:1979:CSS

[Dav79] Larry S. Davis. Computing the spatial structure of cellular textures. Computer Graphics and Image Processing, 11(2):111–122, October 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Davidson:1993:CLT

[Dav93] Jennifer L. Davidson. Classification of lattice transformations in image processing. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):283–306, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1020/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1020/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1020/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1020/production/
pdf.

Davatzikos:1997:STR

[Dav97] Christos Davatzikos. Spatial transformation and registration of brain images using elastically deformable models. Computer Vision and Image Understanding: CVIU, 66(2):207–222, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0605/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0605/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0605/production/ref.

Dreike:1976:CRF

[DB76] Philip Dreike and Douglas P.

Boyd. Convolution reconstruction of fan beam projections. Computer Graphics and Image Processing, 5(4): 459–469, December 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Davis:1979:CMS

[DB79] Larry S. Davis and Michael L.

Benedikt. Computational models of space: Isovists and isovist fields. Computer Graphics and Image Processing, 11(1):49–72, September 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

DHaeyer:1988:PCI

[DB88] Johan P. F. D’Haeyer and Ignace Bruyland. Parallel computation of image curve velocity fields. Computer Vision, Graphics, and Image Processing, 43(2):239– 255, August 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Draper:1994:RPC

[DB94] Bruce A. Draper and J. Ross Beveridge. Response to “Performance Characterization in Computer Vision”. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2): 262–263, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1055/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1055/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1060/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1060/production/

pdf. See [Har94a, Har94b, Shi94].

DelaTorre:2003:RPC

[DB03] Fernando De la Torre and Michael J. Black. Robust parameterized component analysis: theory and applications to 2D facial appearance models. Computer Vision and Image Understanding: CVIU, 91(1–2):53–71, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dori:1983:CCP

[DBB83] Dov Dori and Moshe Ben-Bassat. Circumscribing a convex polygon by a polygon of fewer sides with minimal area addition. Computer Vision, Graphics, and Image Processing, 24(2):131– 159, November 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Dutt:2013:APO

[DBB13] Mousumi Dutt, Arindam Biswas, and Partha Bhowmick. Approximate partitioning of 2D objects into orthogonally convex components. Computer Vision and Image Understanding: CVIU, 117(4):326–341, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001853.

Draper:2003:RFP

[DBBB03] Bruce A. Draper, Kyungim Baek, Marian Stewart Bartlett, and J. Ross Beveridge. Recognizing faces with PCA and ICA. Computer Vision and Image Understanding: CVIU, 91(1–2):115–137, July/August 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

deBerg:1998:LDT

[dBD98] Mark de Berg and Katrin T. G. Dobrindt. On levels of detail in terrains. Graphical Models and Image Processing: GMIP, 60(1):001–012, January 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0460/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0460/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0460/production/ref.

Damiand:2004:TMT

[DBF04] Guillaume Damiand, Yves Bertrand, and Christophe Fiorio. Topological model for two-dimensional image representation: definition and optimal extraction algorithm. Computer Vision and Image Understanding: CVIU, 93(2): 111–154, February 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dankers:2007:MZS

[DBZ07] Andrew Dankers, Nick Barnes, and Alex Zelinsky. MAP ZDF segmentation and tracking using active stereo vision: Hand tracking case study. Computer Vision and Image Understanding: CVIU, 108(1– 2):74–86, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Derin:1986:STI

[DC86] Haluk Derin and William S.

Cole. Segmentation of textured images using Gibbs random fields. Computer Vision, Graphics, and Image Processing, 35(1):72–98, July 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Das:1988:NDT

[DC88] P. P. Das and B. N. Chatterji. A note on “Distance Transformations in Arbitrary Dimensions”. Computer Vision, Graphics, and Image Processing, 43(3):368– 385, September 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See [Bor84].

Deshpande:1998:REI

[DC98] S. G. Deshpande and S. Chaudhuri. Recursive estimation of illuminant motion from flow field and simultaneous recovery of shape. Computer Vision and Image Understanding: CVIU, 72(1):10–20, October 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.

com/links/artid/cviu.1997.
0657/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0657/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0657/production/ref.

Dornaika:2000:CSM

[DC00a] F. Dornaika and R. Chung. Cooperative stereo-motion: Matching and reconstruction. Computer Vision and Image Understanding: CVIU, 79(3):408–427, September 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0867;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0867/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0867/ref.

Drummond:2000:LTS

[DC00b] Tom Drummond and Terry Caelli. Learning taskspecific object recognition and scene understanding. Computer Vision and Image Understanding: CVIU, 80(3): 315–348, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0882;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0882/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/cviu.
2000.0882/ref.

Dornaika:2001:AAC

[DC01] F. Dornaika and R. Chung.

An algebraic approach to camera self-calibration. Computer Vision and Image Understanding: CVIU, 83(3): 195–215, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0925;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0925/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0925/ref.

Dewaele:2004:IGL

[DC04] Guillaume Dewaele and Marie-Paule Cani. Interactive global and local deformations for virtual clay. Graphical Models, 66(6):352–369, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

DelBimbo:1999:GEI

[DCCL99] Alberto Del Bimbo, Vittorio Castelli, Shih-Fu Chang, and Chung-Sheng Li. GUEST ED-ITORS’ INTRODUCTION: Content-based access of image and video libraries. Computer Vision and Image Understanding: CVIU, 75(1– 2):1–2, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0775/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0775/production/
                                        pdf.

deCampos:2012:ISL

[dCCP12] Teófilo de Campos, Gabriela Csurka, and Florent Perronnin. Images as sets of locally weighted features. Computer Vision and Image Understanding: CVIU, 116(1):68–85, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001822.

Denis:2007:SSQ

[DCFM07] Patrice Denis, Philippe Carre, and Christine Fernandez-Maloigne. Spatial and spectral quaternionic approaches for colour images. Computer Vision and Image Understanding: CVIU, 107(1– 2):74–87, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dee:2012:BSS

[DCH12] Hannah M. Dee, Anthony G. Cohn, and David C. Hogg. Building semantic scene models from unconstrained video. Computer Vision and Image Understanding: CVIU, 116(3):446–456, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002025.

Deng:2008:PSH

[DCL^{+}08] Jiansong Deng, Falai Chen, Xin Li, Changqi Hu, Weihua Tong, Zhouwang Yang, and Yuyu Feng. Polynomial splines over hierarchical T-meshes. Graphical Models, 70(4):76–86, July 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000039.

Drapaca:2005:STB

[DCS05] Corina S. Drapaca, Valerie Cardenas, and Colin Studholme. Segmentation of tissue boundary evolution from brain MR image sequences using multiphase level sets. Computer Vision and Image Understanding: CVIU, 100(3):312– 329, December 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dickinson:1997:AOR

[DCTO97] Sven J. Dickinson, Henrik I.

Christensen, John K. Tsotsos, and Göran Olofsson. Active object recognition integrating attention and viewpoint control. Computer Vision and Image Understanding: CVIU, 67(3):239–260, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0532/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0532/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0532/production/ref.

DelBimbo:2011:PFB

[DD11a] Alberto Del Bimbo and Fabrizio Dini. Particle filter-based visual tracking with a first order dynamic model and uncertainty adaptation. Computer Vision and Image Understanding: CVIU, 115(6): 771–786, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dong:2011:ALM

[DD11b] Y. Dong and G. N. DeSouza. Adaptive learning of multi-subspace for foreground detection under illumination changes. Computer Vision and Image Understanding: CVIU, 115(1): 31–49, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DelBimbo:2010:EDV

[DDLP10] A. Del Bimbo, F. Dini, G. Lisanti, and F. Pernici.

Exploiting distinctive visual landmark maps in pan–tilt– zoom camera networks. Computer Vision and Image Understanding: CVIU, 114(6): 611–623, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Deng:2012:CTG

[DDWZ12] Yue Deng, Qionghai Dai, Ruiping Wang, and Zengke Zhang. Commute time guided transformation for feature extraction. Computer Vision and Image Understanding: CVIU, 116(4):473–483, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002578.

DeSouza:1983:ARD

[De 83a] Peter De Souza. Automatic rib detection in chest radiographs. Computer Vision, Graphics, and Image Processing, 23(2):129–161, August 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

deSouza:1983:EDU

[de 83b] Peter de Souza. Edge detection using sliding statistical tests. Computer Vision, Graphics, and Image Processing, 23(1):1–14, July 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

DeBiase:1988:TAI

[De 88] G. A. De Biase. Trends in astronomical image processing. Computer Vision, Graphics, and Image Processing, 43 (3):347–360, September 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

DeVleeschauwer:1993:IBC

[De 93] D. De Vleeschauwer. An intensity-based, coarse-to-fine approach to reliably measure binocular disparity. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):204–218, March 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1013/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1013/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1013/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1013/production/
pdf.

Demi:1996:CTE

[Dem96]  Marcello Demi.  Contour

tracking by enhancing corners and junctions. Computer Vision and Image Understanding: CVIU, 63 (1):118–134, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0008/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0008/production/
pdf.

Demi:2005:GLC

[Dem05] M. Demi. On the gray-level central and absolute central moments and the mass center of the gray-level variability in low-level image processing. Computer Vision and Image Understanding: CVIU, 97(2): 180–208, February 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dixit:1991:HAV

[DF91] S. S. Dixit and Y. Feng. Hierarchical address vector quantization for image coding. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1):63–70, January 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Drew:1992:NM

[DF92] Mark S. Drew and Brian V.

Funt. Natural metamers. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(2): 139–151, September 1992.

CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Delamarre:2001:AMM

[DF01] Quentin Delamarre and Olivier Faugeras. 3D articulated models and multiview tracking with physical forces. Computer Vision and Image Understanding: CVIU, 81(3):328–357, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0892;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0892/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0892/ref.

Dalley:2002:PWR

[DF02] Gerald Dalley and Patrick Flynn. Pair-wise range image registration: a study in outlier classification. Computer Vision and Image Understanding: CVIU, 87(1–3):104–115, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Costa:1993:EDD

[dFCS93] Luciano da F. Costa and Mark B. Sandler. Effective detection of digital bar segments with Hough transform. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(3):180–191, May 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1013/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1013/production/
pdf.

DeFloriani:1985:DBR

[DFP85] L. De Floriani, B. Falcidieno, and C. Pienovi. Delaunaybased representation of surfaces defined over arbitrarily shaped domains. Computer Vision, Graphics, and Image Processing, 32(1):127– 140, October 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

DeFloriani:1989:SGR

[DFP89] L. De Floriani, B. Falcidieno, and C. Pienovi. Structured graph representation of a hierarchical triangulation. Computer Vision, Graphics, and Image Processing, 45 (2):215–226, February 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

deFloriani:1992:LAC

[dFP92] L. de Floriani and E. Puppo.

An on-line algorithm for constrained Delaunay triangulation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4): 290–300, July 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Lascio:2013:RTA

[DFP^{+}13] Rosario Di Lascio, Pasquale Foggia, Gennaro Percannella, Alessia Saggese, and Mario Vento. A real time algorithm for people tracking using contextual reasoning. Computer Vision and Image Understanding: CVIU, 117(8):892–908, August 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000908.

Durou:2008:NMS

[DFS08] Jean-Denis Durou, Maurizio Falcone, and Manuela Sagona. Numerical methods for shapefrom-shading: a new survey with benchmarks. Computer Vision and Image Understanding: CVIU, 109(1): 22–43, January 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Derrode:2001:REF

[DG01] Stéphane Derrode and Faouzi Ghorbel. Robust and efficient Fourier-Mellin transform approximations for graylevel image reconstruction and complete invariant description. Computer Vision and Image Understanding: CVIU, 83(1):57–78, July 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0922;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0922/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0922/ref.

Dorini:2011:UFT

[DG11] Leyza Baldo Dorini and Siome Klein Goldenstein. Unscented feature tracking. Computer Vision and Image Understanding: CVIU, 115 (1):8–15, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Daubney:2012:EPA

[DGC12] Ben Daubney, David Gibson, and Neill Campbell. Estimating pose of articulated objects using low-level motion. Computer Vision and Image Understanding: CVIU, 116(3):330–346, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002049.

Delakis:2008:AIS

[DGG08] Manolis Delakis, Guillaume Gravier, and Patrick Gros. Audiovisual integration with Segment Models for tennis video parsing. Computer Vision and Image Understanding: CVIU, 111(2):142– 154, August 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Daul:1998:HTN

[DGH98] Christian Daul, Pierre Graebling, and Ernest Hirsch. From the Hough transform to a new approach for the detection and approximation of elliptical arcs. Computer Vision and Image Understanding: CVIU, 72(3):215–236, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0696/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0696/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0696/production/ref.

Daneshpajouh:2012:CPP

[DGZ12] Shervin Daneshpajouh, Mohammad Ghodsi, and Alireza Zarei. Computing polygonal path simplification under area measures. Graphical Models, 74(5):283–289, September 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000264.

DHaeyer:1986:DMI

[D’H86] Johan D’Haeyer. Determining motion of image curves from local pattern changes. Computer Vision, Graphics, and Image Processing, 34(2):166–188, May 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Danielsson:1992:HAR

[DH92] Per-Erik Danielsson and Magnus Hammerin. Highaccuracy rotation of images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4):340–344, July 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Demirdjian:2000:MED

[DH00] David Demirdjian and Radu Horaud. Motion-egomotion discrimination and motion segmentation from imagepair streams. Computer Vision and Image Understanding: CVIU, 78(1):53–68, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0827;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0827/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0827/ref.

Deng:2012:P

[DHK12] Jiansong Deng, Kai Hormann, and Misha Kazhdan. Preface. Graphical Models, 74(4):75, July 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000355.

Drouin:2008:IDO

[DHP08] Stéphane Drouin, Patrick Hébert, and Marc Parizeau. Incremental discovery of object parts in video sequences. Computer Vision and Image Understanding: CVIU, 110 (1):60–74, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DiZenzo:1986:NNG

[Di 86] Silvano Di Zenzo. Note: a note on the gradient of a multi-image. Computer Vision, Graphics, and Image Processing, 33(1):116– 125, January 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Drosou:2012:SAH

[DIMT12] Anastasios Drosou, Dimosthenis Ioannidis, Konstantinos Moustakas, and Dimitrios Tzovaras. Spatiotemporal analysis of human activities for biometric authentication. Computer Vision and Image Understanding: CVIU, 116(3):411–421, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002098.

Delponte:2006:SMU

[DIOV06] Elisabetta Delponte, Francesco Isgrò, Francesca Odone, and Alessandro Verri. SVDmatching using SIFT features. Graphical Models, 68(5–6):415–431, September/November 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000579.

Dubuisson-Jolly:2001:TDT

[DJG01] Marie-Pierre Dubuisson-Jolly and Alok Gupta. Tracking deformable templates using a shortest path algorithm. Computer Vision and Image Understanding: CVIU, 81(1):26–45, January 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0883;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0883/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0883/ref.

Danielsson:1979:DCA

[DK79] Per Erik Danielsson and Bjørn Kruse. Distance checking algorithms. Computer Graphics and Image Processing, 11(4): 349–376, December 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Dadashi:2013:AFN

[DK13] Roghayeh Dadashi and Hamidreza Rashidy Kanan. AVCD-FRA: a novel solution to automatic video cut detection using fuzzy-rule-based approach. Computer Vision and Image Understanding: CVIU, 117(7):807–817, July 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000544.

Das:1997:RID

[DL97] Atish K. Das and Noshir A.

Langrana. Recognition and integration of dimension sets in vectorized engineering drawings. Computer Vision and Image Understanding: CVIU, 68(1):90–108, October 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0537/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0537/production/
                                        pdf;   http://www.idealibrary.

com/links/artid/cviu.1997.
0537/production/ref.

Draper:2005:ESA

[DL05] Bruce A. Draper and Albert Lionelle. Evaluation of selective attention under similarity transformations. Computer Vision and Image Understanding: CVIU, 100(1–2): 152–171, October/November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Drbohlav:2010:TCI

[DL10] Ondřej Drbohlav and Aleš Leonardis. Towards correct and informative evaluation methodology for texture classification under varying viewpoint and illumination. Computer Vision and Image Understanding: CVIU, 114(4):439–449, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

deLavarene:2007:PIL

[dLAH07] Brice Chaix de Lavarène, David Alleysson, and Jeanny Hérault. Practical implementation of LMMSE demosaicing using luminance and chrominance spaces. Computer Vision and Image Understanding: CVIU, 107(1–2): 3–13, July/August 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dimitrijevic:2006:HBP

[DLF06] M. Dimitrijevic, V. Lepetit, and P. Fua. Human body pose detection using Bayesian spatio-temporal templates. Computer Vision and Image Understanding: CVIU, 104(2–3):127–139, November/December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Deng:1999:WBL

[DLHT99] Peter Shaohua Deng, Hong-Yuan Mark Liao, Chin Wen Ho, and Hsiao-Rong Tyan. Wavelet-based off-line handwritten signature verification. Computer Vision and Image Understanding: CVIU, 76(3):173–190, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0799/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0799/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0799/production/ref.

Duan:2013:GEP

[DLP13] Ye Duan, Dong Li, and P. Frank Pai. Geometrically exact physics-based modeling and computer animation of highly flexible 1D mechanical systems. Graphical Models, 75(2):56–68, March 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070313000039.

DOrazio:2009:VSR

[DLS^{+}09] T. D’Orazio, M. Leo, P. Spagnolo, M. Nitti, N. Mosca, and A. Distante. A visual system for real time detection of goal events during soccer matches. Computer Vision and Image Understanding: CVIU, 113 (5):622–632, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dionne:1978:ASA

[DM78] Mark S. Dionne and Alan K.

Mackworth. ANTICS: a system for animating LISP programs. Computer Graphics and Image Processing, 7(1): 105–119, February 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Davis:1980:EDT

[DM80] Larry S. Davis and Amar Mitiche. Edge detection in textures. Computer Graphics and Image Processing, 12 (1):25–39, January 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Davis:1981:EDT

[DM81] Larry S. Davis and Amar Mitiche. Edge detection in textures — maxima selection. Computer Graphics and Image Processing, 16(2): 158–165, June 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Davis:1982:MMD

[DM82] Larry S. Davis and Amar Mitiche. MITES (mites): a model-driven, iterative texture segmentation algorithm. Computer Graphics and Image Processing, 19 (2):95–110, June 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

deSaintPierre:1992:NEC

[dM92] T. de Saint Pierre and M. Milgram. New and efficient cellular algorithms for image processing. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (3):261–274, May 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Delingette:2001:STC

[DM01] H. Delingette and J. Montagnat. Shape and topology constraints on parametric active contours. Computer Vision and Image Understanding: CVIU, 83(2):140–171, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.

com/links/doi/10.1006/cviu.
2001.0920;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0920/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0920/ref.

Datta:2012:TDS

[DM12] Madhura Datta and C. A.

Murthy. Two dimensional synthetic face generation and verification using set estimation technique. Computer Vision and Image Understanding: CVIU, 116(9): 1022–1031, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000793.

deMiranda:2010:SAW

[dMFU10] P. A. V. de Miranda, A. X.

Falcão, and J. K. Udupa. Synergistic arc-weight estimation for interactive image segmentation using graphs. Computer Vision and Image Understanding: CVIU, 114(1): 85–99, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DeFloriani:2003:DNM

[DMMP03] Leila De Floriani, Mostefa M.

Mesmoudi, Franco Morando, and Enrico Puppo. Decomposing non-manifold objects in arbitrary dimensions. Graphical Models, 65(1–3):

2–22, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Dunne:2010:EGC

[DMW10] Aubrey K. Dunne, John Mallon, and Paul F. Whelan. Efficient generic calibration method for general cameras with single centre of projection. Computer Vision and Image Understanding: CVIU, 114(2):220– 233, February 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dreschler:1982:VMD

[DN82] L. Dreschler and H. H. Nagel. Volumetric model and 3-D trajectory of a moving car derived from monocular TV frame sequences of a street scene. Computer Graphics and Image Processing, 20(3): 199–228, November 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Defee:1991:MBZ

[DN91] I. Defee and Y. Neuvo. Median-based zero-crossing edge detectors for closely spaced edges. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2): 196–203, March 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Dodd:1998:NAV

[Dod98] R. K. Dodd. A new approach to the visualization of tensor fields. Graphical Models and Image Processing: GMIP, 60(4):286–303, July 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0473/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0473/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0473/production/ref.

Doermann:1998:IRD

[Doe98] David Doermann. The indexing and retrieval of document images: a survey. Computer Vision and Image Understanding: CVIU, 70(3):287–298, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0687/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0687/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0687/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0692/production;
 http://www.idealibrary.
com/links/artid/cviu.1998.

                                        0692/production/pdf;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0692/production/
                                        ref.

Doros:1979:AGD

[Dor79] Marek Doros. Algorithms for generation of discrete circles, rings, and disks. Computer Graphics and Image Processing, 10(4):366– 371, August 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Doros:1984:SPG

[Dor84]  Marek Doros.  On some

properties of the generation of discrete circular arcs on a square grid. Computer Vision, Graphics, and Image Processing, 28(3):377– 383, December 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Dori:1989:SGA

[Dor89]  Dov Dori.  A syntactic/

geometric approach to recognition of dimensions in engineering machine drawings. Computer Vision, Graphics, and Image Processing, 47 (3):271–291, September 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Demirci:2011:EMM

[DOSD11] M. Fatih Demirci, Yusuf Osmanlioglu, Ali Shokoufandeh, and Sven Dickinson. Efficient many-to-many feature matching under the l_{1} norm. Computer Vision and Image Understanding: CVIU, 115(7):976–983, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000804.

Junior:2012:RCH

[dOSJVBS12] Jurandir de Oliveira Santos Junior, Alexandre Vrubel, Olga R. P. Bellon, and Luciano Silva. 3D reconstruction of cultural heritages: Challenges and advances on precise mesh integration. Computer Vision and Image Understanding: CVIU, 116(12): 1195–1207, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001166.

Douglass:1981:ITD

[Dou81]  Robert J. Douglass.  In-

terpreting three-dimensional scenes: a model building approach. Computer Graphics and Image Processing, 17(2): 91–113, October 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Dougherty:1992:OMSa

[Dou92a] Edward R. Dougherty. Optimal mean-square N-observation digital morphological filters. I. Optimal binary filters. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(1):36– 54, January 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Dougherty:1992:OMSb

[Dou92b] Edward R. Dougherty. Optimal mean-square N-observation digital morphological filters. II. Optimal gray-scale filters. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(1):55– 72, January 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Dori:1988:GDM

[DP88] Dov Dori and Amir Pnueli. The grammar of dimensions in machine drawings. Computer Vision, Graphics, and Image Processing, 42(1):1– 18, April 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

deLaGorce:2010:VAM

[dP10] Martin de La Gorce and Nikos Paragios. A variational approach to monocular hand-pose estimation. Computer Vision and Image Understanding: CVIU, 114(3): 363–372, March 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Demi:2000:FAC

[DPB00] M. Demi, M. Paterni, and A. Benassi. The first absolute central moment in low-level image processing. Computer Vision and Image Understanding: CVIU, 80(1):57–87, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0861;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0861/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0861/ref.

Dickinson:1992:VVA

[DPR92] S. J. Dickinson, A. P. Pentland, and A. Rosenfeld. From volumes to views: an approach to 3-D object recognition. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (2):130–154, March 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Dowdall:2007:CT

[DPT07] Jonathan Dowdall, Ioannis T.

Pavlidis, and Panagiotis Tsiamyrtzis. Coalitional tracking. Computer Vision and Image Understanding: CVIU, 106(2–3):205–219, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Duan:2004:SBD

[DQ04] Ye Duan and Hong Qin. A subdivision-based deformable model for surface reconstruction of unknown topology. Graphical Models, 66(4):181– 202, July 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Du:2005:DPB

[DQ05] Haixia Du and Hong Qin. Dynamic PDE-based surface design using geometric and physical constraints. Graphical Models, 67(1):43– 71, January 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Donahue:1993:ULC

[DR93] M. J. Donahue and S. I. Rokhlin. On the use of level curves in image analysis. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):185–203, March 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1993.
                                        1012/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1993.1012/production/
                                        pdf;   http://www.idealibrary.

com/links/artid/cviu.1993.
1012/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1012/production/
pdf.

Damiand:2003:SMA

[DR03] Guillaume Damiand and Patrick Resch. Split-and-merge algorithms defined on topological maps for 3D image segmentation. Graphical Models, 65(1–3):149–167, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Das:2004:FSS

[DR04] M. Das and E. M. Riseman.

FOCUS: a system for searching for multi-colored objects in a diverse image database. Computer Vision and Image Understanding: CVIU, 94(1– 3):168–192, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Diaz:2008:SHP

[DRAB08] Javier Díaz, Eduardo Ros, Rodrigo Agís, and Jose Luis Bernier. Superpipelined highperformance optical-flow computation architecture. Computer Vision and Image Understanding: CVIU, 112(3): 262–273, December 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dehne:1995:HAP

[DRCF95] Frank Dehne, Andrew Rau-Chaplin, and Afonso G. Ferreira. Hypercube algorithms for parallel processing of pointer-based quadtrees. Computer Vision and Image Understanding: CVIU, 62(1):1–10, July 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1995.1037/production;
 http://www.idealibrary.
com/links/artid/cviu.1995.
1037/production/pdf.

Debled-Rennesson:2013:SID

[DRDKE13] Isabelle Debled-Rennesson, Eric Domenjoud, Bertrand Kerautret, and Philippe Even. Special issue on discrete geometry for computer imagery. Computer Vision and Image Understanding: CVIU, 117(4):305, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000167.

Drew:1994:RSD

[Dre94]  Mark S. Drew.  Robust

specularity detection from a single multi-illuminant color image. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):320–327, May 1994. CODEN CIUNEJ.

ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1022/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1022/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1026/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1026/production/
pdf.

Drew:1996:DSO

[Dre96] Mark S. Drew. Direct solution of orientation-from-color problem using a modification of Pentland’s light source direction estimator. Computer Vision and Image Understanding: CVIU, 64(2):286–299, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0059/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0059/production/
pdf.

Denman:2003:CBA

[DRK03] H. Denman, N. Rea, and A. Kokaram. Contentbased analysis for video from snooker broadcasts. Computer Vision and Image Understanding: CVIU, 92(2–3):176– 195, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dorst:1987:LED

[DS87] Leo Dorst and Arnold W. M. Smeulders. Length estimators for digitized contours. Computer Vision, Graphics, and Image Processing, 40 (3):311–333, December 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Danielsson:1990:RIG

[DS90] Per-Erik E. Danielsson and Olle Seger. Rotation invariance in gradient and higher order derivative detectors. Computer Vision, Graphics, and Image Processing, 49 (2):198–221, February 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Davis:2007:BSU

[DS07] James W. Davis and Vinay Sharma. Background-subtraction using contour-based fusion of thermal and visible imagery. Computer Vision and Image Understanding: CVIU, 106(2–3):162–182, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Damiand:2011:PAS

[DSdlH^{+}11] Guillaume Damiand, Christine Solnon, Colin de la Higuera, Jean-Christophe Janodet, and Émilie Samuel.

Polynomial algorithms for subisomorphism of nD open combinatorial maps. Computer Vision and Image Understanding: CVIU, 115(7): 996–1010, July 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000816.

daSilva:2012:IMD

[dSdSF^{+}12] André Tavares da Silva, Jefersson Alex dos Santos, Alexandre Xavier Falcão, Ricardo da S. Torres, and Léo Pini Magalhães. Incorporating multiple distance spaces in optimum-path forest classification to improve feedback-based learning. Computer Vision and Image Understanding: CVIU, 116(4):510–523, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421100261X.

Dufournaud:2004:IMS

[DSH04] Yves Dufournaud, Cordelia Schmid, and Radu Horaud. Image matching with scale adjustment. Computer Vision and Image Understanding: CVIU, 93(2):175–194, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Denton:2008:CSI

[DSNN08] Trip Denton, Ali Shokoufandeh, John Novatnack, and Ko Nishino. Canonical subsets of image features. Computer Vision and Image Understanding: CVIU, 112(1): 55–66, October 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DeJonge:1994:TES

[DSS94] Wiebren De Jonge, Peter Scheuermann, and Ardie Schijf. S^{+}-trees: an efficient structure for the representation of large pictures. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):265–280, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1018/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1994.1018/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1022/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1022/production/
                                        pdf.

Dutagaci:2010:SMR

[DSY10] Helin Dutağacı, Bülent Sankur, and Yücel Yemez. Subspace methods for retrieval of general 3D models. Computer Vision and Image Understanding: CVIU, 114(8):

865–886, August 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DaVitoriaLobo:1996:CED

[DT96a] Niels Da Vitoria Lobo and John K. Tsotsos. Computing egomotion and detecting independent motion from image motion using collinear points. Computer Vision and Image Understanding: CVIU, 64(1):21–52, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0044/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0044/production/
pdf.

Dalmia:1996:HSE

[DT96b] Arun K. Dalmia and Mohan Trivedi. High-speed extraction of 3D structure of selectable quality using a translating camera. Computer Vision and Image Understanding: CVIU, 64(1):97–110, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0047/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0047/production/
pdf.

Dudek:1997:SRR

[DT97] Gregory Dudek and John K. Tsotsos. Shape representation and recognition from multiscale curvature. Computer Vision and Image Understanding: CVIU, 68(2): 170–189, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0533/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0533/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0533/production/ref.

Daliri:2009:CSU

[DT09] Mohammad Reza Daliri and Vincent Torre. Classification of silhouettes using contour fragments. Computer Vision and Image Understanding: CVIU, 113(9):1017– 1025, September 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Daliri:2010:SRB

[DT10] Mohammad Reza Daliri and Vincent Torre. Shape recognition based on Kernel-edit distance. Computer Vision and Image Understanding: CVIU, 114(10):1097–1103, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

DenHartog:1996:KBI

[DTG96] J. E. Den Hartog, T. K.

Ten Kate, and J. J. Gerbrands. Knowledge-based interpretation of utility maps. Computer Vision and Image Understanding: CVIU, 63(1):105–117, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0007/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0007/production/
pdf.

Dube:1976:TBS

[Dub76] R. Peter Dube. Tension in a bicubic surface patch. Computer Graphics and Image Processing, 5(4):496– 502, December 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Dube:1977:UBF

[Dub77] R. Peter Dube. Univariate blending functions and alternatives. Computer Graphics and Image Processing, 6(4): 394–408, August 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Delibasis:1997:DFD

[DUC97] K. Delibasis, P. E. Undrill, and G. G. Cameron. Designing Fourier descriptorbased geometric models for object interpretation in medical images using genetic algorithms. Computer Vision and Image Understanding: CVIU, 66(3):286–300, June 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0505/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0505/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0505/production/ref.

Dayanand:1994:PSM

[DUSL94] S. Dayanand, W. R. Uttal, T. Shepherd, and C. Lunskis, Jr. A particle system model for combining edge information from multiple segmentation modules. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(3):219–230, May 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1020/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1020/production/
pdf.

Distante:1982:TPF

[DV82] A. Distante and N. Veneziani. A two-pass filling algorithm for raster graphics. Computer Graphics and Image Processing, 20(3):288– 295, November 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Dori:1998:SRD

[DV98] Dov Dori and Yelena Velkovitch.

Segmentation and recognition of dimensioning text from engineering drawings. Computer Vision and Image Understanding: CVIU, 69 (2):196–201, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0585/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0585/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0585/production/ref.

Demirci:2008:ITL

[DvLV08] M. Fatih Demirci, Reinier H.

van Leuken, and Remco C. Veltkamp. Indexing through Laplacian spectra. Computer Vision and Image Understanding: CVIU, 110(3): 312–325, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Derin:1987:PIS

[DW87] Haluk Derin and Chee-Sun S.

Won. A parallel image segmentation algorithm using relaxation with varying neighborhoods and its mapping to array processors. Computer Vision, Graphics, and Image Processing, 40(1):54– 78, October 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Dang:2011:SIP

[DWB11] Trung Kien Dang, Marcel Worring, and The Duy Bui. A semi-interactive panorama based 3D reconstruction framework for indoor scenes. Computer Vision and Image Understanding: CVIU, 115(11): 1516–1524, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001664.

Davis:1983:CBM

[DWS83] Larry S. Davis, Zhongquan Wu, and Hanfang Sun. Contour-based motion estimation. Computer Vision, Graphics, and Image Processing, 23(3):313–326, September 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Deng:2012:SCH

[DWW^{+}12] Xiaoming Deng, Fuchao Wu, Yihong Wu, Fuqing Duan, Liang Chang, and Hongan Wang. Self-calibration of hybrid central catadioptric and perspective cameras. Computer Vision and Image Understanding: CVIU, 116(6):715–729, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000355.

Davis:1983:EMM

[DWX83] Larry S. Davis, Cheng-Ye Wang, and Hu-Chen Xie. Experiment in multispectral, multitemporal crop classification using relaxation techniques. Computer Vision, Graphics, and Image Processing, 23(2):227–235, August 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ding:1998:CSI

[DY98] Yuan Ding and Tzay Y.

Young. Complete shape from imperfect contour: a rule-based approach. Computer Vision and Image Understanding: CVIU, 70(2): 197–211, May 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1997.0635/production;
 http://www.idealibrary.
com/links/artid/cviu.1997.
0635/production/pdf;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0635/production/
ref.

Dyer:1982:SEQ

[Dye82]  Charles R. Dyer.  The

space efficiency of quadtrees. Computer Graphics and Image Processing, 19(4):335– 348, August 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Dai:2007:PDT

[DZL07] Congxia Dai, Yunfei Zheng, and Xin Li. Pedestrian detection and tracking in infrared imagery using shape and appearance. Computer Vision and Image Understanding: CVIU, 106(2–3):288– 299, May/June 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Eggers:1995:HSI

[EA95] Daniel D. Eggers and Eugene Ackerman. High speed image rotation in embedded systems. Computer Vision and Image Understanding: CVIU, 61(2):270–277, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1995.
                                        1019/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1995.1019/production/
                                        pdf.

Eberlein:1976:IGE

[Ebe76] Robert B. Eberlein. Iterative gradient edge detection algorithm. Computer Graphics and Image Processing, 5 (2):245–253, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Erol:2007:VBH

[EBN^{+}07] Ali Erol, George Bebis, Mircea Nicolescu, Richard D. Boyle, and Xander Twombly. Vision-based hand pose estimation: a review. Computer Vision and Image Understanding: CVIU, 108(1– 2):52–73, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ersoy:1988:TCI

[EC88] O. K. Ersoy and C. H.

Chen. Transform-coding of images with reduced complexity. Computer Vision, Graphics, and Image Processing, 42(1):19–31, April 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Erdem:2012:FBT

[EDB12] Erkut Erdem, Séverine Dubuisson, and Isabelle Bloch. Fragments based tracking with adaptive cue integration. Computer Vision and Image Understanding: CVIU, 116(7):827–841, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000525.

Edelman:1987:LCA

[Ede87] Shimon Edelman. Line connectivity algorithms for an asynchronous pyramid computer. Computer Vision, Graphics, and Image Processing, 40(2):169–187, November 1987. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Edelman:1994:RR

[Ede94] Shimon Edelman. Representation without reconstruction. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):92–94, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1035/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1035/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1040/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1040/production/
pdf.

Ehrich:1978:VTT

[EF78] R. W. Ehrich and J. P. Foith. View of texture topology and texture description. Computer Graphics and Image Processing, 8(2):174– 202, October 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Eggert:1998:SRM

[EFF98] David W. Eggert, Andrew W.

Fitzgibbon, and Robert B. Fisher. Simultaneous registration of multiple range views for use in reverse engineering of CAD models. Computer Vision and Image Understanding: CVIU, 69(3):253–272, March 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0667/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0667/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0667/production/ref.

Eggers:1998:TFE

[Egg98] Hinnik Eggers. Two fast Euclidean distance transformations in Z^{2} based on sufficient propagation. Computer Vision and Image Understanding: CVIU, 69 (1):106–116, January 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0596/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0596/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/cviu.1997.
                                        0596/production/ref.

Guo:2007:PSI

[eGZW07] Cheng en Guo, Song-Chun Zhu, and Ying Nian Wu. Primal sketch: Integrating structure and texture. Computer Vision and Image Understanding: CVIU, 106(1): 5–19, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Escalante:2010:SAI

[EHG^{+}10] Hugo Jair Escalante, Carlos A. Hernández, Jesus A. Gonzalez, A. López-López, Manuel Montes, Eduardo F. Morales, L. Enrique Sucar, Luis Villaseñor, and Michael Grubinger. The segmented and annotated IAPR TC-12 benchmark. Computer Vision and Image Understanding: CVIU, 114(4):419–428, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ehrich:1978:SHS

[Ehr78] R. W. Ehrich. Symmetric hysteresis smoothing algorithm that preserves principal features. Computer Graphics and Image Processing, 8(1): 121–126, August 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Einbu:1983:NRD

[Ein83] John M. Einbu. Nonlinear reduction of data. Computer Vision, Graphics, and Image Processing, 22(2):260– 267, May 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Eichmann:1988:TIT

[EK88] G. Eichmann and T. Kasparis. Topologically invariant texture descriptors. Computer Vision, Graphics, and Image Processing, 41(3):267–281, March 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Elber:1997:GSR

[EK97] Gershon Elber and Myung-Soo Kim. Geometric shape recognition of freeform curves and surfaces. Graphical Models and Image Processing: GMIP, 59(6):417– 433, November 1997. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0441/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0441/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0441/production/ref.

EbroulIzquierdo:1998:IAM

[EK98] M. Ebroul Izquierdo and Silko Kruse. Image analysis for 3D modeling, rendering, and virtual view generation. Computer Vision and Image Understanding: CVIU, 71(2):231–253, August 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0706/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0706/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0706/production/ref.

Escobar:2012:ARB

[EK12] María-José Escobar and Pierre Kornprobst. Action recognition via bio-inspired features: The richness of center-surround interaction. Computer Vision and Image Understanding: CVIU, 116(5):593–605, May 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000185.

Elber:2001:CHR

[EKH01] Gershon Elber, Myung-Soo Kim, and Hee-Seok Heo. The convex hull of rational plane curves. Graphical Models, 63(3):151–162, May 2001. CODEN GRMOFM.

ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0546;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0546/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0546/ref.

Eklundh:1979:UFP

[Ekl79] J. O. Eklundh. On the use of Fourier phase features for texture discrimination. Computer Graphics and Image Processing, 9(2):199– 201, February 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). EKLUNDH79.

Ebrahimpour:2008:TDL

[EKY08] Reza Ebrahimpour, Ehsanollah Kabir, and Mohammad Reza Yousefi. Teacherdirected learning in viewindependent face recognition with mixture of experts using overlapping eigenspaces. Computer Vision and Image Understanding: CVIU, 111(2):195–206, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Eberley:1991:GSI

[EL91] D. Eberley and J. Lancaster.

On gray scale image measurements. I. Arc length and area. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6):538– 549, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Eckhardt:2003:TDS

[EL03] Ulrich Eckhardt and Longin Jan Latecki. Topologies for the digital spaces and. Computer Vision and Image Understanding: CVIU, 90(3): 295–312, June 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Elgammal:2007:NML

[EL07] Ahmed Elgammal and Chan-Su Lee. Nonlinear manifold learning for dynamic shape and dynamic appearance. Computer Vision and Image Understanding: CVIU, 106(1):31–46, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Eberly:1991:GSI

[ELA91] D. Eberly, J. Lancaster, and A. Alyassin. On gray scale image measurements. II. Surface area and volume. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6): 550–562, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Elber:2001:CEI

[Elb01] Gershon Elber. Curve evaluation and interrogation on surfaces. Graphical Models, 63(3):197–210, May 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0541;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0541/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0541/ref.

Elber:2005:GFB

[Elb05]  Gershon Elber.  General-

ized filleting and blending operations toward functional and decorative applications. Graphical Models, 67(3):189– 203, May 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ellis:1981:DRL

[Ell81] J. R. Ellis, Jr. Distribution of run lengths over scanned rectangles. Computer Graphics and Image Processing, 15 (3):246–264, March 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ezquerra:1996:KGS

[EM96] Norberto Ezquerra and Rakesh Mullick. Knowledge-guided segmentation of 3D imagery.

Graphical Models and Image Processing: GMIP, 58 (6):510–523, November 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0043/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0043/production/
pdf.

Engbersen:1983:TTO

[Eng83] Antonius Paulus Johannes Engbersen. Toppsy: a time overlapped parallel processing system. Computer Vision, Graphics, and Image Processing, 24(1):97–106, October 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Enkelmann:1988:IMA

[Enk88]  Wilfried Enkelmann.  In-

vestigations of multigrid algorithms for the estimation of optical flow fields in image sequences. Computer Vision, Graphics, and Image Processing, 43(2):150– 177, August 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Edelsbrunner:1984:SMC

[EOS84] H. Edelsbrunner, M. H. Overmars, and R. Seidel. Some methods of computational geometry applied to computer graphics. Computer Vision, Graphics, and Image Processing, 28(1):92–108, October 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Edelsbrunner:1984:SGR

[EOW84] Herbert Edelsbrunner, Joseph O’Rourke, and Emmerich Welzl. Stationing guards in rectilinear art galleries. Computer Vision, Graphics, and Image Processing, 27 (2):167–176, August 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Elber:2005:SMT

[EPB05] Gershon Elber, Nick Patrikalakis, and Pere Brunet. Solid modeling theory and applications. Graphical Models, 67(5):371–372, September 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407030500007X.

Ellis:1979:MLD

[EPRR79] T. J. Ellis, D. Proffitt, D. Rosen, and W. Rutkowski. Measurement of the lengths of digitized curved lines. Computer Graphics and Image Processing, 10(4):333– 347, August 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Embrechts:1996:PED

[ER96] Hugo Embrechts and Dirk Roose. A parallel Euclidean distance transformation algorithm. Computer Vision and Image Understanding: CVIU, 63(1):15–26, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0002/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0002/production/
                                        pdf.

Embrechts:1993:CLM

[ERW93] Hugo Embrechts, Dirk Roose, and Patrick Wambacq. Component labelling on a MIMD multiprocessor. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):155–165, March 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/artid/
                                        ciun.1993.1010/production;
                                         http://www.idealibrary.
                                        com/links/artid/ciun.1993.
                                        1010/production/pdf;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1993.1010/production;
                                         http://www.idealibrary.
                                        com/links/artid/cviu.1993.
                                        1010/production/pdf.

Ehrich:1981:CBF

[ES81a] R. W. Ehrich and F. H. Schroeder. Contextual boundary formation by one-dimensional edge detection and scan line matching. Computer Graphics and Image Processing, 16 (2):116–149, June 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Elliott:1981:ADP

[ES81b] H. Elliott and L. Srinivasan. An application of dynamic programming to sequential boundary estimation. Computer Graphics and Image Processing, 17(4):291– 314, December 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Esteban:2004:SSF

[ES04] Carlos Hernández Esteban and Francis Schmitt. Silhouette and stereo fusion for 3D object modeling. Computer Vision and Image Understanding: CVIU, 96(3): 367–392, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Erdem:2006:ACL

[ES06] Ugur Murat Erdem and Stan Sclaroff. Automated camera layout to satisfy task-specific and floor plan-specific coverage requirements. Computer Vision and Image Understanding: CVIU, 103(3):156– 169, September 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ekenel:2010:VBD

[ESS10] Hazım Kemal Ekenel, Johannes Stallkamp, and Rainer Stiefelhagen. A video-based door monitoring system using local appearance-based face models. Computer Vision and Image Understanding: CVIU, 114(5):596–608, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Eu:1994:APC

[ET94] D. Eu and G. T. Toussaint. On approximating polygonal curves in two and three dimensions. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(3):231–246, May 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1021/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1021/production/
                                        pdf.

Etesami:1985:NAD

[EU85] Faryar Etesami and John J. Uicker, Jr. Note: Automatic dimensional inspection of machine part crosssections using Fourier analysis. Computer Vision, Graphics, and Image Processing, 29 (2):216–247, February 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Evako:2006:TPC

[Eva06] Alexander V. Evako. Topological properties of closed digital spaces: One method of constructing digital models of closed continuous surfaces by using covers. Computer Vision and Image Understanding: CVIU, 102(2):134–144, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Evako:2011:CSP

[Eva11] Alexander V. Evako. Characterizations of simple points, simple edges and simple cliques of digital spaces: One method of topologypreserving transformations of digital spaces by deleting simple points and edges. Graphical Models, 73(1):1– 9, January 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070310000159.

Eastman:1987:UDF

[EW87] Roger D. Eastman and Allen M. Waxman. Using disparity functionals for stereo correspondence and surface reconstruction. Computer Vision, Graphics, and Image Processing, 39(1):73– 101, July 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Eberly:1991:AGA

[EW91] D. Eberly and D. Wenzel. Adaption of group algebras to signal and image processing. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):340–348, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Escalante:2011:EBM

[EyGS11] Hugo Jair Escalante, Manuel Montes y Goméz, and Luis Enrique Sucar. An energy-based model for region-labeling. Computer Vision and Image Understanding: CVIU, 115 (6):787–803, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Feldmar:1997:PRF

[FAB97] Jacques Feldmar, Nicholas Ayache, and Fabienne Betting. 3D-2D projective registration of free-form curves and surfaces. Computer Vision and Image Understanding: CVIU, 65(3):403–424, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0499/production;   http:

//www.idealibrary.com/links/
artid/cviu.1996.0499/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0499/production/ref.

Fouquier:2012:SMB

[FAB12] Geoffroy Fouquier, Jamal Atif, and Isabelle Bloch. Sequential model-based segmentation and recognition of image structures driven by visual features and spatial relations. Computer Vision and Image Understanding: CVIU, 116(1):146–165, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001998.

Farin:1982:CVC

[Far82] G. Farin. A construction for visual C1 continuity of polynomial surface patches. Computer Graphics and Image Processing, 20(3):272– 282, November 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Farouki:1986:CPS

[Far86] R. T. Farouki. The characterization of parametric surface sections. Computer Vision, Graphics, and Image Processing, 33(2):209– 236, February 1986. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Farouki:2002:ERM

[Far02]  Rida T. Farouki.  Exact

rotation-minimizing frames for spatial Pythagoreanhodograph curves. Graphical Models, 64(6):382–395, November 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Farouk:2011:IRB

[Far11] R. M. Farouk. Iris recognition based on elastic graph matching and Gabor wavelets. Computer Vision and Image Understanding: CVIU, 115 (8):1239–1244, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001081.

Faugeras:1981:DDT

[Fau81] Olivier D. Faugeras. Decomposition and decentralization techniques in relaxation labeling. Computer Graphics and Image Processing, 16(4): 341–355, August 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Fua:1997:IHC

[FB97] P. Fua and C. Brechbühler. Imposing hard constraints on deformable models through optimization in orthogonal subspaces. Computer Vision and Image Understanding:

CVIU, 65(2):148–162, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0568/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0568/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0568/production/ref.

Fiala:2005:PSR

[FB05] Mark Fiala and Anup Basu.

Panoramic stereo reconstruction using non-SVP optics. Computer Vision and Image Understanding: CVIU, 98(3): 363–397, June 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ferraz:2012:SCB

[FB12] Luis Ferraz and Xavier Binefa. A sparse curvaturebased detector of affine invariant blobs. Computer Vision and Image Understanding: CVIU, 116(4):524–537, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002621.

Faltemier:2008:UMI

[FBF08] Timothy C. Faltemier, Kevin W.

Bowyer, and Patrick J. Flynn. Using multi-instance enrollment to improve performance of 3D face recognition. Computer Vision and Image Understanding: CVIU, 112(2): 114–125, November 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Furst:1986:EDI

[FC86] M. A. Furst and P. E. Caines. Edge detection with image enhancement via dynamic programming. Computer Vision, Graphics, and Image Processing, 33(3):263– 279, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ferley:2001:RAV

[FCG01] Eric Ferley, Marie-Paule Cani, and Jean-Dominique Gascuel. Resolution adaptive volume sculpting. Graphical Models, 63(6):459–478, November 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Fejes:1999:DIM

[FD99] Sándor Fejes and Larry S. Davis. Detection of independent motion using directional motion estimation. Computer Vision and Image Understanding: CVIU, 74(2):101–120, May 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0751/production;   http:

//www.idealibrary.com/links/
artid/cviu.1999.0751/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0751/production/ref.

Feldmar:1997:EIA

[FDMA97] J. Feldmar, J. Declerck, G. Malandain, and N. Ayache. Extension of the ICP algorithm to nonrigid intensity-based registration of 3D volumes. Computer Vision and Image Understanding: CVIU, 66(2):193–206, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0606/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0606/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0606/production/ref.

Feldman:1985:CMP

[Fel85] Jerome A. Feldman. Connectionist models and parallelism in high level vision. Computer Vision, Graphics, and Image Processing, 31 (2):178–200, August 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fan:1979:SAT

[FF79] T. I. Fan and K. S. Fu.

Syntactic approach to timevarying image analysis. Computer Graphics and Image Processing, 11(2):138– 149, October 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Farenzena:2009:SMG

[FF09] Michela Farenzena and Andrea Fusiello. Stabilizing 3D modeling with geometric constraints propagation. Computer Vision and Image Understanding: CVIU, 113(11): 1147–1157, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fei-Fei:2007:LGV

[FFFP07] Li Fei-Fei, Rob Fergus, and Pietro Perona. Learning generative visual models from few training examples: an incremental Bayesian approach tested on 101 object categories. Computer Vision and Image Understanding: CVIU, 106(1):59–70, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fasel:2005:GFR

[FFM05] Ian Fasel, Bret Fortenberry, and Javier Movellan. A generative framework for real time object detection and classification. Computer Vision and Image Understanding: CVIU, 98(1):182–210, April 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fang:2004:ARS

[FFY^{+}04] C. Y. Fang, C. S. Fuh, P. S.

Yen, S. Cherng, and S. W. Chen. An automatic road sign recognition system based on a computational model of human recognition processing. Computer Vision and Image Understanding: CVIU, 96(2): 237–268, November 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Falcidieno:1989:ARR

[FG89] Bianca Falcidieno and Franca Giannini. Automatic recognition and representation of shape-based features in a geometric modeling system. Computer Vision, Graphics, and Image Processing, 48 (1):93–123, October 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fahnestock:1984:MSM

[FH84a] J. D. Fahnestock and B. R.

Hunt. The maintenance of sharpness in magnified digital images. Computer Vision, Graphics, and Image Processing, 27(1):32–45, July 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fang:1984:STD

[FH84b] J.-Q. Fang and T. S. Huang.

Solving three-dimensional small-rotation motion equations: Uniqueness, algorithms, and numerical results. Computer Vision, Graphics, and Image Processing, 26(2):183–206, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Farin:2012:AGS

[FH12] Gerald Farin and Dianne Hansford. Agnostic G^{1} Gregory surfaces. Graphical Models, 74(6):346–350, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000367.

Faugeras:1984:PAD

[FHMB84] O. D. Faugeras, M. Hebert, P. Mussi, and J. D. Boissonnat. Polyhedral approximation of 3-D objects without holes. Computer Vision, Graphics, and Image Processing, 25(2):169– 183, February 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Flynn:2001:SIE

[FHP01] Patrick J. Flynn, Adam Hoover, and P. Jonathon Phillips. Special issue on empirical evaluation of computer vision algorithms. Computer Vision and Image Understanding: CVIU, 84(1):1–4, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0948;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0948/
pdf.

Field:1986:ADA

[Fie86] Dan Field. Algorithms for drawing anti-aliased circles and ellipses. Computer Vision, Graphics, and Image Processing, 33(1):1–15, January 1986. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fischler:1994:MRV

[Fis94] Martin A. Fischler. The modeling and representation of visual information. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):98–99, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1037/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1037/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1042/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1042/production/
pdf.

Fitzpatrick:1988:EGD

[Fit88] J. Michael Fitzpatrick. The existence of geometrical densityimage transformations corresponding to object motion. Computer Vision, Graphics, and Image Processing, 44 (2):155–174, November 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fiume:1989:MSR

[Fiu89] Eugene Fiume. A mathematical semantics of rendering. I. Ideal rendering. Computer Vision, Graphics, and Image Processing, 48 (3):281–303, December 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fiume:1991:CMC

[Fiu91a] E. Fiume. Coverage masks and convolution tables for the fast area sampling. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1): 25–30, January 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Fiume:1991:MSR

[Fiu91b] E. Fiume. A mathematical semantics of rendering. II. Approximation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1):

19–24, January 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Fleet:1991:PBD

[FJJ91] D. J. Fleet, A. D. Jepson, and M. R. M. Jenkin. Phase-based disparity measurement. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (2):198–210, March 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Fuhrer:2006:MHP

[FJP06] Martin Fuhrer, Henrik Wann Jensen, and Przemyslaw Prusinkiewicz. Modeling hairy plants. Graphical Models, 68(4):333–342, July 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407030600004X.

Favre:1983:PST

[FK83] A. Favre and Hj. Keller.

Parallel syntactic thinning by recoding of binary pictures. Computer Vision, Graphics, and Image Processing, 23(1):99–112, July 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fujimura:1999:SRC

[FK99] Kikuo Fujimura and Eddy Kuo. Shape reconstruction from contours using isotopic deformation. Graphical Models and Image Processing: GMIP, 61(3):127–147, May 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0494/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0494/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0494/production/ref;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0502/production;
 http://www.idealibrary.
com/links/artid/gmip.1999.
0502/production/pdf.

Fielding:2000:CCO

[FK00] Gabriel Fielding and Moshe Kam. Computing the cost of occlusion. Computer Vision and Image Understanding: CVIU, 79(2):324–329, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0853;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0853/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0853/ref.

Furukawa:2009:LRS

[FK09] Ryo Furukawa and Hiroshi Kawasaki. Laser range scanner based on self-calibration techniques using coplanarities and metric constraints. Computer Vision and Image Understanding: CVIU, 113(11): 1118–1129, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fischer:1998:EBA

[FKL^{+}98] André Fischer, Thomas H.

Kolbe, Felicitas Lang, Armin B. Cremers, Wolfgang Förstner, Lutz Plümer, and Volker Steinhage. Extracting buildings from aerial images using hierarchical aggregation in 2D and 3D. Computer Vision and Image Understanding: CVIU, 72(2):185–203, November 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0721/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0721/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0721/production/ref.

Fritz:2010:TBL

[FKS10] Mario Fritz, Geert-Jan M.

Kruijff, and Bernt Schiele. Tutor-based learning of visual categories using different levels of supervision. Computer Vision and Image Understanding: CVIU, 114(5): 564–573, May 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ferrer:2011:GFM

[FKV^{+}11] M. Ferrer, D. Karatzas, E. Valveny, I. Bardaji, and H. Bunke. A generic framework for median graph computation based on a recursive embedding approach. Computer Vision and Image Understanding: CVIU, 115(7):919–928, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000786.

Fieguth:1998:EMC

[FKW98] Paul W. Fieguth, William C. Karl, and Alan S. Willsky. Efficient multiresolution counterparts to variational methods for surface reconstruction. Computer Vision and Image Understanding: CVIU, 70(2):157–176, May 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0630/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0630/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0630/production/ref.

Fitzpatrick:1987:COO

[FL87] J. Michael Fitzpatrick and Michael R. Leuze. A class of one-to-one two-dimensional transformations. Computer Vision, Graphics, and Image Processing, 39(3):369– 382, September 1987. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ferrie:1992:CCI

[FL92] F. P. Ferrie and J. Lagarde. Curvature consistency improves local shading analysis. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(1): 95–105, January 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Fua:1996:TAI

[FL96] P. Fua and Y. G. Leclerc.

Taking advantage of imagebased and geometry-based constraints to recover 3-D surfaces. Computer Vision and Image Understanding: CVIU, 64(1):111–127, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0048/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0048/production/
pdf.

Fan:2009:FEM

[FL09] Shu-Kai S. Fan and Yen Lin.

A fast estimation method for the generalized Gaussian mixture distribution on complex images. Computer Vision and Image Understanding: CVIU, 113(7):839–853, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Flasinski:1989:CEG

[Fla89]  Mariusz Flasinski.  Char-

acteristics of edNLC-graph grammar for syntactic pattern recognition. Computer Vision, Graphics, and Image Processing, 47(1):1–21, July 1989. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Figueroa:2006:TSP

[FLB06] Pascual J. Figueroa, Neucimar J. Leite, and Ricardo M. L. Barros. Tracking soccer players aiming their kinematical motion analysis. Computer Vision and Image Understanding: CVIU, 101(2): 122–135, February 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Felisberto:2006:ODR

[FLCdA06] Marcelo Kleber Felisberto, Heitor Silvério Lopes, Tania Mezzadri Centeno, and Lúcia Valéria Ramos de Arruda. An object detection and recognition system for weld bead extraction from digital radiographs. Computer Vision and Image Understanding: CVIU, 102(3):238–249, June 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fu:2008:LAS

[FLHK08] Yun Fu, Zhu Li, Thomas S.

Huang, and Aggelos K. Katsaggelos. Locally adaptive subspace and similarity metric learning for visual data clustering and retrieval. Computer Vision and Image Understanding: CVIU, 110(3): 390–402, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Flynn:1992:ORU

[Fly92] P. J. Flynn. 3D object recognition using invariant feature indexing of interpretation tables. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (2):119–129, March 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Fanelli:1984:PTN

[FM84] A. M. Fanelli and B. Marangelli. Progressive transmission of news photos. Computer Vision, Graphics, and Image Processing, 27(2):239– 245, August 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fabian:1991:RIM

[FM91] R. Fabian and D. Malah. Robust identification of motion and out-of-focus blur parameters from blurred and noisy images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):403– 412, September 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Foster:1996:RAL

[FM96] Nick Foster and Dimitri Metaxas. Realistic animation of liquids. Graphical Models and Image Processing: GMIP, 58(5):471–483, September 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0039/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0039/production/
pdf.

Fujimura:1998:FFI

[FM98] Kikuo Fujimura and Mihail Makarov. Foldover-free image warping. Graphical Models and Image Processing: GMIP, 60(2):100–111, March 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0454/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0454/production/
pdf;   http://www.idealibrary.

com/links/artid/gmip.1998.
0454/production/ref.

Fua:1999:AHO

[FM99] P. Fua and C. Miccio. Animated heads from ordinary images: a least-squares approach. Computer Vision and Image Understanding: CVIU, 75(3):247–259, September 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0778/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0778/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0778/production/ref.

Falomir:2012:MQD

[FMGA^{+}12] Zoe Falomir, Lledó Museros, Luis Gonzalez-Abril, M. Teresa Escrig, and Juan A. Ortega. A model for the qualitative description of images based on visual and spatial features. Computer Vision and Image Understanding: CVIU, 116(6):698–714, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421200032X.

Fan:2012:SBM

[FML12] Lubin Fan, Min Meng, and Ligang Liu. Sketch-based mesh cutting: a comparative study. Graphical Models, 74(6):292–301, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000082.

Fradkin:2001:BDM

[FMR01] M. Fradkin, H. Maître, and M. Roux. Building detection from multiple aerial images in dense urban areas. Computer Vision and Image Understanding: CVIU, 82(3):181–207, June 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0917;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0917/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0917/ref.

Foresti:1994:GRS

[FMRV94] Gianluca Foresti, Vittorio Murino, Carlo S. Regazzoni, and Gianni Vernazza. Grouping of rectilinear segments by the labeled Hough transform. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):22–42, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1002/production;   http:
//www.idealibrary.com/links/

artid/ciun.1994.1002/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1002/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1002/production/
pdf.

Ferri:1993:PPE

[FMV93] M. Ferri, F. Mangili, and G. Viano. Projective pose estimation of linear and quadratic primitives in monocular computer vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(1):66–84, July 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1032/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1032/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1033/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1033/production/
pdf.

Fogg:1984:CRG

[Fog84] D. A. Fogg. Contour to rectangular grid conversion using minimum curvature. Computer Vision, Graphics, and Image Processing, 28(1):85–91, October 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fogel:1991:EVV

[Fog91] Sergei V. Fogel. The estimation of velocity vector fields from time-varying image sequences. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (3):253–287, May 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Fogg:1993:VLG

[Fog93] D. A. Fogg. Vegetationlimited ground-to-air surveillance. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55 (6):419–427, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1993.
                                        1032/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1993.1032/production/
                                        pdf.

Fong:1990:AAC

[Fon90] Amelia Fong. Algorithms and architectures for a class of non-linear hybrid filters. Computer Vision, Graphics, and Image Processing, 50(1):101–111, April 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Forrest:1972:COM

[For72] A. Robin Forrest. On Coons’ and other methods for the representation of curved surfaces. Computer Graphics and Image Processing, 1(4):341– 359, December 1972. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Forstner:1987:RAP

[För87]  W. Förstner.  Reliability

analysis of parameter estimation in linear models with applications to mensuration problems in computer vision. Computer Vision, Graphics, and Image Processing, 40 (3):273–310, December 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Forshaw:1988:SMH

[For88] M. R. B. Forshaw. Speeding up the Marr-Hildreth edge operator. Computer Vision, Graphics, and Image Processing, 41(2):172– 185, February 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Forchhammer:1989:DPG

[For89] Soren Forchhammer. Digital plane and grid point segments. Computer Vision, Graphics, and Image Processing, 47(3):373–384, September 1989. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fablet:2008:IBR

[FPC^{+}08] R. Fablet, S. Pujolle, A. Chessel, A. Benzinou, and F. Cao. 2D image-based reconstruction of shape deformation of biological structures using a level-set representation. Computer Vision and Image Understanding: CVIU, 111 (3):295–306, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Frahm:2012:SIV

[FPDK12] Jan-Michael Frahm, Marc Pollefeys, Frank Dellaert, and Jana Kosecka. Special issue on Virtual Representations and Modeling of Large-scale environments (VRML). Computer Vision and Image Understanding: CVIU, 116(1):1, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002402.

Feng:1980:SNS

[FR80] D. Feng and F. Richard. Some new surface forms for computer aided geometric design. Computer Graphics and Image Processing, 23(4): 324–331, ???? 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Filipovych:2011:RSA

[FR11] Roman Filipovych and Eraldo Ribeiro. Robust sequence alignment for actor–object interaction recognition: Discovering actor–object states. Computer Vision and Image Understanding: CVIU, 115 (2):177–193, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Franklin:1979:EAD

[Fra79] W. Randolph Franklin. Evaluation of algorithms to display vector plots on raster devices. Computer Graphics and Image Processing, 11(4):377– 397, December 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Franklin:1981:EHS

[Fra81] W. Randolph Franklin. An exact hidden sphere algorithm that operates in linear time. Computer Graphics and Image Processing, 15 (4):364–379, April 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Franklin:1983:RNR

[Fra83] W. Randolph Franklin. Rays — new representation for polygons and polyhedra. Computer Vision, Graphics, and Image Processing, 22(3):327–338, June 1983.

CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fraser:1989:CHS

[Fra89] Donald Fraser. Comparison at high spatial frequencies of two-pass and one-pass geometric transformation algorithms. Computer Vision, Graphics, and Image Processing, 46(3):267–283, June 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Francon:1995:DCS

[Fra95] Jean Francon. Discrete combinatorial surfaces. Graphical Models and Image Processing: GMIP, 57(1):20–26, January 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1003/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1003/production/
pdf.

Favreau:2006:AGV

[FRDC06] Laurent Favreau, Lionel Reveret, Christine Depraz, and Marie-Paule Cani. Animal gaits from video: Comparative studies. Graphical Models, 68(2):212–234, March 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000330.

Freuder:1976:ARA

[Fre76] Eugene C. Freuder. Affinity:

a relative approach to region finding. Computer Graphics and Image Processing, 5(2): 254–264, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Frei:1977:IEH

[Fre77] Werner Frei. Image enhancement by histogram hyperbolization. Computer Graphics and Image Processing, 6 (3):286–294, June 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Frieden:1980:SMI

[Fri80] B. Roy Frieden. Statistical models for the image restoration problem. Computer Graphics and Image Processing, 12(1):40–59, January 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Friedberg:1986:FAS

[Fri86] Stuart A. Friedberg. Finding axes of skewed symmetry. Computer Vision, Graphics, and Image Processing, 34(2):138–155, May 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Faugeras:1998:DRU

[FRL^{+}98] Olivier Faugeras, Luc Robert, Stéphane Laveau, Gabriella Csurka, Cyril Zeller, Cyrille Gauclin, and Imad Zoghlami. 3-D reconstruction of urban scenes from image sequences. Computer Vision and Image Understanding: CVIU, 69(3):292–309, March 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0665/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0665/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0665/production/ref.

Frintrop:2005:BLB

[FRNS05] Simone Frintrop, Erich Rome, Andreas Nüchter, and Hartmut Surmann. A Bimodal Laser-Based Attention System. Computer Vision and Image Understanding: CVIU, 100(1–2):124–151, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Freeman:1980:CAL

[FS80] H. Freeman and J. A. Saghri. Comparative analysis of linedrawing modeling schemes. Computer Graphics and Image Processing, 12(3):203– 223, March 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ferrari:1984:FRA

[FS84] Leonard A. Ferrari and Jack Sklansky. A fast recursive algorithm for binaryvalued two-dimensional filters. Computer Vision, Graphics, and Image Processing, 26(3):292–302, June 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ferrari:1985:NDI

[FS85] Leonard A. Ferrari and Jack Sklansky. A note on Duhamel integrals and running average filters. Computer Vision, Graphics, and Image Processing, 29(3):358– 360, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ford:1995:RVF

[FS95] Ralph M. Ford and Robin N.

Strickland. Representing and visualizing fluid flow images and velocimetry data by nonlinear dynamical systems. Graphical Models and Image Processing: GMIP, 57 (6):462–482, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1040/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1040/production/
pdf.

Furukawa:2003:ARL

[FS03] Yasutaka Furukawa and Yoshihisa Shinagawa. Accurate and robust line segment extraction by analyzing distribution around peaks in Hough space. Computer Vision and Image Understanding: CVIU, 92(1): 1–25, October 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fermuller:2001:SOF

[FSA01] Cornelia Fermüller, David Shulman, and Yiannis Aloimonos. The statistics of optical flow. Computer Vision and Image Understanding: CVIU, 82(1):1–32, April 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0900;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0900/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0900/ref.

Fraser:1985:RMI

[FSB85] Donald Fraser, Robert A. Schowengerdt, and Ian Briggs. Rectification of multichannel images in mass storage using image transposition. Computer Vision, Graphics, and Image Processing, 29(1):23–36, January 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fratarcangeli:2007:FMC

[FSF07] Marco Fratarcangeli, Marco Schaerf, and Robert Forchheimer. Facial motion cloning with radial basis functions in MPEG-4 FBA. Graphical Models, 69(2):106–118, March 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000658.

Ferrari:1984:MRP

[FSS84] L. Ferrari, P. V. Sankar, and J. Sklansky. Minimal rectangular partitions of digitized blobs. Computer Vision, Graphics, and Image Processing, 28(1):58– 71, October 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ferrari:1994:EAI

[FSS94] L. A. Ferrari, M. J. Silbermann, and P. V. Sankar. Efficient algorithms for the implementation of general B-splines. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):102–105, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.

                                        1009/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1009/production/
                                        pdf.

Ferrari:1986:ETD

[FSSL86] Leonard A. Ferrari, P. V. Sankar, Jack Sklansky, and Sidney Leeman. Efficient two-dimensional filters using B-spline functions. Computer Vision, Graphics, and Image Processing, 35(2):152– 169, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ford:1994:IMD

[FST94] Ralph M. Ford, Robin N. Strickland, and Bruce A. Thomas. Image models for 2-D flow visualization and compression. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):75–93, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1007/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1007/production/
                                        pdf.

Fransens:2007:OFB

[FSV07] Rik Fransens, Christoph Strecha, and Luc Van Gool. Optical flow based superresolution: a probabilistic approach. Computer Vision and Image Understanding: CVIU, 106(1):106–115, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fennema:1979:VDS

[FT79] Claude L. Fennema and William B. Thompson. Velocity determination in scenes containing several moving objects. Computer Graphics and Image Processing, 9(4): 301–315, April 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Fan:1984:ACS

[FT84] Ting-Jun J. Fan and Wen-Hsiang H. Tsai. Automatic Chinese seal identification. Computer Vision, Graphics, and Image Processing, 25(3):311–330, March 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Fidrich:1998:SCP

[FT98] Márta Fidrich and Jean-Philippe Thirion. Stability of corner points in scale space: The effects of small nonrigid deformations. Computer Vision and Image Understanding: CVIU, 72(1):72–83, October 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.

                                        0661/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0661/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0661/production/ref.

Fischler:1981:DRL

[FTW81] M. A. Fischler, J. M. Tenenbaum, and H. C. Wolf. Detection of roads and linear structures in low-resolution aerial imagery using a multisource knowledge integration technique. Computer Graphics and Image Processing, 15(3): 201–223, March 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). FISCHLER81.

Fu:1980:SIM

[Fu80] K. S. Fu. Syntactic image modeling using stochastic tree grammars. Computer Graphics and Image Processing, 12 (2):136–152, February 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Fujimura:1997:VCR

[Fuj97]  Kikuo Fujimura.  Visibil-

ity computation on reconfigurable meshes. Graphical Models and Image Processing: GMIP, 59(6):395– 406, November 1997. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.

0440/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0440/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0440/production/ref.

Falcao:1998:USI

[FUS^{+}98] Alexandre X. Falcão, Jayaram K. Udupa, Supun Samarasekera, Shoba Sharma, Bruce Elliot Hirsch, and Roberto de A. Lotufo. Usersteered image segmentation paradigms: Live wire and live lane. Graphical Models and Image Processing: GMIP, 60(4):233–260, July 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0475/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0475/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0475/production/ref.

Fan:1997:SCS

[FW97] Joel Fan and Lawrence B.

Wolff. Surface curvature and shape reconstruction from unknown multiple illumination and integrability. Computer Vision and Image Understanding: CVIU, 65(2):347–359, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.

                                        com/links/artid/cviu.1996.
                                        0581/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0581/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0581/production/ref.

Fadaifard:2013:MFE

[FWH13] Hadi Fadaifard, George Wolberg, and Robert Haralick. Multiscale 3D feature extraction and matching with an application to 3D face recognition. Graphical Models, 75(4):157–176, July 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000076.

Fahn:1988:TBC

[FWL88] Chin-Shyurng Fahn, Jhing-Fa Wang, and Jau-Yien Lee. A topology-based component extractor for understanding electronic circuit diagrams. Computer Vision, Graphics, and Image Processing, 44 (2):119–138, November 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Feng:2013:CCM

[FWWT13] Xin Feng, Yuanzhen Wang, Yanlin Weng, and Yiying Tong. Compact combinatorial maps: a volume mesh data structure. Graphical Models, 75(3):149–156, May 2013. CODEN GRMOFM.

ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000690.

Frost:1985:MLC

[FY85] V. S. Frost and L. S.

Yurovsky. Maximum likelihood classification of synthetic aperture radar imagery. Computer Vision, Graphics, and Image Processing, 32 (3):291–313, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Feldman:2006:HIP

[FY06] Thomas Feldman and Laurent Younes. Homeostatic image perception: an artificial system. Computer Vision and Image Understanding: CVIU, 102(1):70–80, April 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Fu:2011:SIF

[FYH11] Yun (Raymond) Fu, Shuicheng Yan, and Thomas S. Huang. Special issue on featureoriented image and video computing for extracting contexts and semantics. Computer Vision and Image Understanding: CVIU, 115(3):289, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gargantini:1991:MSC

[GA91] I. Gargantini and H. H. Atkinson. Multiple-seed 3D connectivity filling for inaccurate borders. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6): 563–573, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Galin:2000:IPI

[GA00] Eric Galin and Samir Akkouche. Incremental polygonization of implicit surfaces. Graphical Models, 62(1):19–39, January 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/artid/gmod.1999.
0514/production;   http:
//www.idealibrary.com/links/
artid/gmod.1999.0514/production/
pdf;   http://www.idealibrary.
com/links/artid/gmod.1999.
0514/production/ref.

Goncalves:2009:EPN

[GA09] Nuno Gonçalves and Helder Araújo. Estimating parameters of noncentral catadioptric systems using bundle adjustment. Computer Vision and Image Understanding: CVIU, 113(1):11–28, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ghanem:2013:MDS

[GA13] Bernard Ghanem and Narendra Ahuja. Modeling dynamic swarms. Computer Vision and Image Understanding: CVIU, 117(1):1–11, January 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001282.

Gaafar:1977:CVB

[Gaa77]  Magdy Gaafar.  Convex-

ity verification, block-chords, and digital straight lines. Computer Graphics and Image Processing, 6(4):361– 370, August 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Gader:1991:SDA

[Gad91] Paul D. Gader. Separable decompositions and approximations of greyscale morphological templates. Computer Vision, Graphics, and Image Processing. Image Understanding, 53(3):288–296, May 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Gutierrez:2001:SDV

[GAD01] J. Gutiérrez, G. Ayala, and M. E. Díaz. Set descriptors for visual evaluation of human corneal endothelia. Computer Vision and Image Understanding:

CVIU, 84(2):249–263, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0936;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0936/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0936/ref.

Gard:1976:DPP

[Gar76]  Robert L. Gard.  Digi-

tal picture processing techniques for the publishing industry. Computer Graphics and Image Processing, 5(2): 151–171, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Gargantini:1982:LOF

[Gar82]  Irene Gargantini.  Linear

octtrees for fast processing of three-dimensional objects. Computer Graphics and Image Processing, 20(4):365– 374, December 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Gauch:1992:IIC

[Gau92] John M. Gauch. Investigations of image contrast space defined by variations on histogram equalization. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4):269–280, July 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Gavrila:1999:VAH

[Gav99] D. M. Gavrila. The visual analysis of human movement: a survey. Computer Vision and Image Understanding: CVIU, 73(1):82–98, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0716/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0716/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0716/production/ref.

Gupta:1993:VSR

[GB93] Alok Gupta and Ruzena Bajcsy. Volumetric segmentation of range images of 3D objects using superquadratic models. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3):302–326, November 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1044/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1044/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1046/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/cviu.1993.1046/production/
                                        pdf.

Golland:1996:WRG

[GB96] P. Golland and A. M. Bruckstein. Why R.G.B.? or how to design color displays for Martians. Graphical Models and Image Processing: GMIP, 58(5):405–412, September 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0034/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0034/production/
                                        pdf.

Golland:1997:MC

[GB97] P. Golland and A. M. Bruckstein. Motion from color. Computer Vision and Image Understanding: CVIU, 68 (3):346–362, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0553/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0553/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0553/production/ref.

Gilbert:2008:IST

[GB08] Andrew Gilbert and Richard Bowden. Incremental, scalable tracking of objects inter camera. Computer Vision and Image Understanding: CVIU, 111(1):43–58, July 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Galleguillos:2010:CBO

[GB10] Carolina Galleguillos and Serge Belongie. Context based object categorization: a critical survey. Computer Vision and Image Understanding: CVIU, 114(6):712–722, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goncalves:2013:DTS

[GB13] Wesley Nunes Gonçalves and Odemir Martinez Bruno. Dynamic texture segmentation based on deterministic partially self-avoiding walks. Computer Vision and Image Understanding: CVIU, 117(9):1163–1174, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000921.

Gros:1998:ULP

[GBB98] Patrick Gros, Olivier Bournez, and Edmond Boyer. Using local planar geometric invariants to match and model images of line segments. Computer Vision and Image Understanding: CVIU, 69(2):135–155, February 1998. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0565/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0565/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0565/production/ref.

Gehrig:2012:ISP

[GBF12] Stefan K. Gehrig, Hernán Badino, and Uwe Franke. Improving sub-pixel accuracy for long range stereo. Computer Vision and Image Understanding: CVIU, 116(1):16–24, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001792.

Gheissari:2006:PMB

[GBHS06] Niloofar Gheissari, Alireza Bab-Hadiashar, and David Suter. Parametric modelbased motion segmentation using surface selection criterion. Computer Vision and Image Understanding: CVIU, 102(2):214–226, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gouet-Brunet:2008:ORS

[GBL08] Valérie Gouet-Brunet and Bruno Lameyre. Object recognition and segmentation in videos by connecting heterogeneous visual features. Computer Vision and Image Understanding: CVIU, 111 (1):86–109, July 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gerbrands:1979:CAM

[GBR79] Jan J. Gerbrands, Fred Booman, and Johan H. C. Reiber. Computer analysis of moving radiopaque markers from X-ray cinefilms. Computer Graphics and Image Processing, 11(1):35– 48, September 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Gualtieri:1989:VPO

[GBW89] J. A. Gualtieri, S. Baugher, and M. Werman. The visual potential: one convex polygon. Computer Vision, Graphics, and Image Processing, 46(1):96–130, April 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gregory:1980:CTI

[GC80] J. A. Gregory and P. Charrot.

A C1 triangular interpolation patch for computer-aided geometric design. Computer Graphics and Image Processing, 13(1):80–87, May 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Goshtasby:1990:BSC

[GCB90] A. Goshtasby, Fuhua Cheng, and B. A. Barsky. B-spline curves and surfaces viewed as digital filters. Computer Vision, Graphics, and Image Processing, 52(2):264– 275, November 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See note [Rab92].

Gerlot-Chiron:1992:RMM

[GCB92] Pascale Gerlot-Chiron and Yves Bizais. Registration of multimodality medical images using a region overlap criterion. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(5):396– 406, September 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Gomez:2007:PAU

[GCEC07] David Delgado Gomez, Line Harder Clemmensen, Bjarne K. Ersbøll, and Jens Michael Carstensen. Precise acquisition and unsupervised segmentation of multi-spectral images. Computer Vision and Image Understanding: CVIU, 106(2–3):183–193, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gouiffes:2012:SLP

[GCFMT12] M. Gouiffès, C. Collewet, C. Fernandez-Maloigne, and A. Trémeau. A study on local photometric models and their application to robust tracking. Computer Vision and Image Understanding: CVIU, 116(8):896–907, August 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421200063X.

Gosselin:2008:CVD

[GCPF08] Philippe Henri Gosselin, Matthieu Cord, and Sylvie Philipp-Foliguet. Combining visual dictionary, kernel-based similarity and learning strategy for image category retrieval. Computer Vision and Image Understanding: CVIU, 110(3):403–417, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gonzalez-Diaz:2011:IRC

[GDIIHK11] Rocio Gonzalez-Diaz, Adrian Ion, Mabel Iglesias-Ham, and Walter G. Kropatsch. Invariant representative cocycles of cohomology generators using irregular graph pyramids. Computer Vision and Image Understanding: CVIU, 115(7):1011–1022, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000774.

Greenspan:2004:CDS

[GDR04] Hayit Greenspan, Guy Dvir, and Yossi Rubner. Contextdependent segmentation and matching in image databases. Computer Vision and Image Understanding: CVIU, 93(1): 86–109, January 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gimenez:2008:EAM

[GE08] David Gimenez and Adrian N. Evans. An evaluation of area morphology scale-spaces for colour images. Computer Vision and Image Understanding: CVIU, 110(1): 32–42, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gershon:1985:APC

[Ger85] Ron Gershon. Aspects of perception and computation in color vision. Computer Vision, Graphics, and Image Processing, 32(2):244– 277, November 1985. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Green:1995:GRA

[GESB95] Kevin Green, David Eggert, Louise Stark, and Kevin Bowyer. Generic recognition of articulated objects through reasoning about potential function. Computer Vision and Image Understanding: CVIU, 62(2): 177–193, September 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1049/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1049/production/
pdf.

Gao:2011:AIS

[GFL^{+}11] Xinbo Gao, Rong Fu, Xuelong Li, Dacheng Tao, Beichen Zhang, and Huigen Yang. Aurora image segmentation by combining patch and texture thresholding. Computer Vision and Image Understanding: CVIU, 115(3): 390–402, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gevers:2004:CII

[GFS04] Theo Gevers, Graham Finlayson, and Raimondo Schettini. Color for image indexing and retrieval. Computer Vision and Image Understanding: CVIU, 94(1–3): 1–2, April/June 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gu:2013:ASU

[GFW13] Xiaodong Gu, Yu Fang, and Yuanyuan Wang. Attention selection using global topological properties based on pulse coupled neural network. Computer Vision and Image Understanding: CVIU, 117(10):1400–1411, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213001045.

Gevrekci:2009:IRI

[GG09] Murat Gevrekci and Bahadir K. Gunturk. Illumination robust interest point detection. Computer Vision and Image Understanding: CVIU, 113(4):565–571, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gomez-Garcia:2008:IRB

[GGMV08] Héctor Fernando Gómez-García, José L. Marroquín, and Johan Van Horebeek. Image registration based on kernel-predictability. Computer Vision and Image Understanding: CVIU, 112(2): 160–172, November 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Grossmann:2010:DCC

[GGO10] Etienne Grossmann, José António Gaspar, and Francesco Orabona. Discrete camera calibration from pixel streams. Computer Vision and Image Understanding: CVIU, 114(2): 198–209, February 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Greenspan:2001:CPF

[GGR01] Hayit Greenspan, Jacob Goldberger, and Lenny Ridel. A continuous probabilistic framework for image matching. Computer Vision and Image Understanding: CVIU, 84 (3):384–406, December 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gokmen:1990:PSA

[GH90] M. Gokmen and R. W. Hall.

Parallel shrinking algorithms using 2-subfields approaches. Computer Vision, Graphics, and Image Processing, 52 (2):191–209, November 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Guo:1992:FFP

[GH92] Zicheng Guo and Richard W.

Hall. Fast fully parallel thinning algorithms. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(3):317–328, May 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Granier:2003:SLR

[GH03] Xavier Granier and Wolfgang Heidrich. A simple layered RGB BRDF model. Graphical Models, 65(4):171–184, July 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gondra:2008:CBI

[GH08] Iker Gondra and Douglas R. Heisterkamp. Content-based image retrieval with the normalized information distance. Computer Vision and Image Understanding: CVIU, 111(2):219–228, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gao:2010:HIC

[GHA10] Chunyu Gao, Hong Hua, and Narendra Ahuja. A hemispherical imaging camera. Computer Vision and Image Understanding: CVIU, 114 (2):168–178, February 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gong:2004:MEM

[GHHX04] Yihong Gong, Mei Han, Wei Hua, and Wei Xu. Maximum entropy model-based baseball highlight detection and classification. Computer Vision and Image Understanding: CVIU, 96(2): 181–199, November 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gros:1997:HUP

[GHMQ97] P. Gros, R. Hartley, R. Mohr, and L. Quan. How useful is projective geometry? Computer Vision and Image Understanding: CVIU, 65(3):442–446, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0496/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0496/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0496/production/ref.

Godin:2009:SIN

[GHMT09] Guy Godin, Patrick Hébert, Takeshi Masuda, and Gabriel Taubin. Special issue on new advances in 3D imaging and modeling. Computer Vision and Image Understanding: CVIU, 113(11): 1105–1106, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ghosh:1988:MMS

[Gho88] Pijush K. Ghosh. A mathematical model for shape description using Minkowski operators. Computer Vision, Graphics, and Image Processing, 44(3):239–269, December 1988. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ghosh:1990:SPC

[Gho90] Pijush K. Ghosh. A solution of polygon containment, spatial planning, and other related problems using Minkowski operations. Computer Vision, Graphics, and Image Processing, 49(1):1–35, January 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ghosh:1991:APT

[Gho91] Pijush K. Ghosh. An algebra of polygons through the notion of negative shapes. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(1):119–144, July 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Grandine:2012:SIS

[GHPW12] Thomas Grandine, Stefanie Hahmann, Jörg Peters, and Wenping Wang. Special issue of selected papers from the 8th Dagstuhl seminar on Geometric Modeling. Graphical Models, 74(6):291, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000422.

Gu:2006:MS

[GHQ06] Xianfeng Gu, Ying He, and Hong Qin. Manifold splines. Graphical Models, 68(3):237–254, May 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407030600021X.

Goutsias:1995:MOI

[GHS95] John Goutsias, Henk J. A. M.

Heijmans, and K. Sivakumar. Morphological operators for image sequences. Computer Vision and Image Understanding: CVIU, 62 (3):326–346, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1058/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1058/production/
pdf.

Gao:2013:FEU

[GHZ^{+}13] Quanxue Gao, Xiujuan Hao, Qijun Zhao, Weiguo Shen, and Jingjie Ma. Feature extraction using twodimensional neighborhood margin and variation embedding. Computer Vision and Image Understanding: CVIU, 117(5):525–531, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000027.

Gopalan:2010:CCL

[GJ10] Raghuraman Gopalan and David Jacobs. Comparing and combining lighting insensitive approaches for face recognition. Computer Vision and Image Understanding: CVIU, 114(1):135–145, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Galata:2001:LVL

[GJH01] Aphrodite Galata, Neil Johnson, and David Hogg. Learning variable-length Markov models of behavior. Computer Vision and Image Understanding: CVIU, 81 (3):398–413, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0894;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0894/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0894/ref.

Gunsel:1996:RBD

[GJP96] Bilge Günsel, Anil K. Jain, and Erdal Panayirci. Reconstruction and boundary detection of range and intensity images using multiscale MRF representations. Computer Vision and Image Understanding: CVIU, 63(2):353–366, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0025/production;   http:
//www.idealibrary.com/links/

artid/cviu.1996.0025/production/
pdf.

Giardina:1977:ACA

[GK77] Charles R. Giardina and Frank P. Kuhl. Accuracy of curve approximation by harmonically related vectors with elliptical loci. Computer Graphics and Image Processing, 6(3):277–285, June 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Gotlieb:1990:TDB

[GK90] Calvin C. Gotlieb and Herbert E. Kreyszig. Texture descriptors based on cooccurrence matrices. Computer Vision, Graphics, and Image Processing, 51(1):70– 86, July 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Grewe:1995:ILM

[GK95] Lynne Grewe and Avinash C.

Kak. Interactive learning of a multiple-attribute hash table classifier for fast object recognition. Computer Vision and Image Understanding: CVIU, 61(3):387–416, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1030/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1995.1030/production/
                                        pdf.

Ghosh:1998:SFR

[GK98] Pijush K. Ghosh and K. Vinod Kumar. Support function representation of convex bodies, its application in geometric computing, and some related representations. Computer Vision and Image Understanding: CVIU, 72(3): 379–403, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0674/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0674/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0674/production/ref.

Gau:2003:MNS

[GK03] C. J. Gau and T. Yung Kong. Minimal non-simple sets in 4D binary images. Graphical Models, 65(1–3):112–130, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gress:2004:ERE

[GK04] Alexander Greß and Reinhard Klein. Efficient representation and extraction of 2manifold isosurfaces using kdtrees. Graphical Models, 66 (6):370–397, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gao:2005:MKF

[GKK05] Jean Gao, Akio Kosaka, and Avinash C. Kak. A multi-Kalman filtering approach for video tracking of humandelineated objects in cluttered environments. Computer Vision and Image Understanding: CVIU, 99(1):1–57, July 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See erratum [Ano06m].

Greiner:2002:SDI

[GKR02] Günther Greiner, Andreas Kolb, and Angela Riepl. Scattered data interpolation using data dependant optimization techniques. Graphical Models, 64(1):1–18, January 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gibson:1982:VRI

[GL82] L. Gibson and D. Lucas.

Vectorization of raster images using hierarchical methods. Computer Graphics and Image Processing, 20(1):82– 89, September 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Grosky:1986:IIU

[GL86] William I. Grosky and Yi Lu.

Iconic indexing using generalized pattern matching techniques. Computer Vision, Graphics, and Image Processing, 35(3):383–403, September 1986. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gross:1995:DPT

[GL95] Ari Gross and Longin Latecki. Digitizations preserving topological and differential geometric properties. Computer Vision and Image Understanding: CVIU, 62 (3):370–381, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1061/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1061/production/
pdf.

Gross:1997:RDM

[GL97] Ari Gross and Longin Jan Latecki. A realistic digitization model of straight lines. Computer Vision and Image Understanding: CVIU, 67(2):131–142, August 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0530/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0530/production/
pdf;   http://www.idealibrary.

com/links/artid/cviu.1997.
0530/production/ref.

Guo:1998:DVS

[GL98] Baining Guo and Joseph Liu.

Direct visible surface interpolation. Computer Vision and Image Understanding: CVIU, 72(3):328–339, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0668/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0668/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0668/production/ref.

Glasbey:1993:AHB

[Gla93] C. A. Glasbey. An analysis of histogram-based thresholding algorithms. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(6):532–537, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1040/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1040/production/
pdf.

Glachet:1993:LMF

[GLD93] R. Glachet, J. T. Lapreste, and M. Dhome. Locating and modelling a flat symmetric object from a single perspective image. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):219–226, March 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1014/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1014/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1014/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1014/production/
pdf.

Gleicher:2001:CCB

[Gle01] Michael Gleicher. Comparing constraint-based motion editing methods. Graphical Models, 63(2):107–134, March 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0549;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0549/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0549/ref.

Giustini:1978:PGU

[GLM78a] M. D. Giustini, M. D. Levine, and A. S. Malowany. Picture generation using semantic nets. Computer Graphics and Image Processing, 7 (??):1–29, ???? 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Guistini:1978:PGU

[GLM78b] R. D. Guistini, M. D. Levine, and A. S. Malowany. Picture generation using semantic nets. Computer Graphics and Image Processing, 7 (1):1–29, February 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Germa:2010:VRD

[GLOC10] T. Germa, F. Lerasle, N. Ouadah, and V. Cadenat. Vision and RFID data fusion for tracking people in crowds by a mobile robot. Computer Vision and Image Understanding: CVIU, 114(6):641–651, June 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Glossop:1999:IHT

[GLR^{+}99] K. Glossop, P. J. G. Lisboa, P. C. Russell, A. Siddans, and G. R. Jones. An implementation of the Hough transformation for the identification and labelling of fixed period sinusoidal curves. Computer Vision and Image Understanding: CVIU, 74(1):96–100, April 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.

                                        0747/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0747/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0747/production/ref.

Gillman:1979:RXR

[GM79] Geoff Gillman and Iain Macleod. Reconstruction of X-ray sources from penumbral images. Computer Graphics and Image Processing, 11(3): 227–241, November 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Gagalowicz:1985:SSN

[GM85] Andre Gagalowicz and Song De Ma. Sequential synthesis of natural textures. Computer Vision, Graphics, and Image Processing, 30(3):289– 315, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Grossberg:1987:NDS

[GM87] Stephen Grossberg and Ennio Mingolla. Neural dynamics of surface perception: Boundary webs, illuminants, and shape-from-shading. Computer Vision, Graphics, and Image Processing, 37(1):116– 165, January 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Grosky:1990:IBO

[GM90] W. I. Grosky and R. Mehrotra. Index-based object recognition in pictorial data management. Computer Vision, Graphics, and Image Processing, 52(3):416–436, December 1990. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Goh:1994:MBM

[GM94] Wooi Boon Goh and Graham R. Martin. Model-based multiresolution motion estimation in noisy images. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):307–319, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1021/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1021/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1025/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1025/production/
pdf.

Gil:1983:ECI

[GMA83] Baldemar Gil, Amar Mitiche, and J. K. Aggarwal. Experiments in combining intensity and range edge maps. Computer Vision, Graphics, and Image Processing, 21(3):395–411, March 1983.

CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Geman:1992:NFF

[GMG92] Stuart Geman, Donald E. McClure, and Donand Geman. A nonlinear filter for film restoration and other problems in image processing. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4):281–289, July 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Guimond:2000:ABM

[GMT00] Alexandre Guimond, Jean Meunier, and Jean-Philippe Thirion. Average brain models: a convergence study. Computer Vision and Image Understanding: CVIU, 77 (2):192–210, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0815;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.1999.0815/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0815/ref.

Gonnet:1983:DDS

[GMW83] Gaston H. Gonnet, J. Ian Munro, and Derick Wood. Direct dynamic structures for some line segment problems.

Computer Vision, Graphics, and Image Processing, 23 (2):178–186, August 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gonzalez:2012:SUH

[GMW12] Jordi Gonzàlez, Thomas B.

Moeslund, and Liang Wang. Semantic understanding of human behaviors in image sequences: From videosurveillance to video-hermeneutics. Computer Vision and Image Understanding: CVIU, 116(3):305–306, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000100.

Gruen:1998:ABE

[GN98] A. Gruen and R. Nevatia. Automatic building extraction from aerial images. Computer Vision and Image Understanding: CVIU, 72(2):99–100, November 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0731/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0731/production/
pdf.

Guting:1987:NAS

[GO87] Ralf Hartmut Güting and Thomas Ottmann. New algorithms for special cases of the hidden line elimination problem. Computer Vision, Graphics, and Image Processing, 40(2):188– 204, November 1987. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Goshtasby:1994:CFS

[GO94] A. Goshtasby and W. D. O’Neill. Curve fitting by a sum of Gaussians. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(4):281–288, July 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1025/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1025/production/
pdf.

Gibson:2007:SCS

[GODC07] D. P. Gibson, D. J. Oziem, C. J. Dalton, and N. W. Campbell. A system for the capture and synthesis of insect motion. Graphical Models, 69(5–6):231–245, September/November 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000671.

Goh:2008:SSM

[Goh08]  Wooi-Boon Goh.  Strate-

gies for shape matching using skeletons. Computer Vision and Image Understanding: CVIU, 110(3):326–345, June 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goldberger:2005:RCP

[Gol05] Jacob Goldberger. Reconstructing camera projection matrices from multiple pairwise overlapping views. Computer Vision and Image Understanding: CVIU, 97(3): 283–296, March 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goldman:2011:UQ

[Gol11] Ron Goldman. Understanding quaternions. Graphical Models, 73(2):21–49, March 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070310000172.

Goldman:2013:MPP

[Gol13]  Ron Goldman.  Modeling

perspective projections in 3dimensions by rotations in 4dimensions. Graphical Models, 75(2):41–55, March 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000707.

Gong:2009:RTJ

[Gon09] Minglun Gong. Real-time joint disparity and disparity flow estimation on programmable graphics hardware. Computer Vision and Image Understanding: CVIU, 113(1):90–100, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goodrich:1992:PAH

[Goo92] Michael T. Goodrich. A polygonal approach to hiddenline and hidden-surface elimination. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1):1– 12, January 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Goshtasby:1989:CID

[Gos89] Ardeshir Goshtasby. Correction of image deformation from lens distortion using Bézier patches. Computer Vision, Graphics, and Image Processing, 47(3):385– 394, September 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gottfried:2008:QSM

[Got08] Björn Gottfried. Qualitative similarity measures—The case of two-dimensional outlines. Computer Vision and Image Understanding: CVIU, 110(1):117–133, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gourlay:1984:GSC

[Gou84] A. R. Gourlay. Grayscale simulation on a color display. Computer Vision, Graphics, and Image Processing, 27(1):92–96, July 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Goutsias:1991:UAG

[Gou91] J. Goutsias. Unilateral approximation of Gibbs random field images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(3): 240–257, May 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Grimson:1985:DDV

[GP85] W. Eric L. Grimson and Theo Pavlidis. Discontinuity detection for visual surface reconstruction. Computer Vision, Graphics, and Image Processing, 30(3):316–330, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Giannini:2006:SIS

[GP06] Franca Giannini and Alexander Pasko. Special issue: Shape modeling international 2004. Graphical Models, 68 (1):1, January 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000585.

Gumundsson:2010:IRS

[GPC^{+}10] Sigurjón Árni Guðmundsson, Montse Pardàs, Josep R. Casas, Jóhannes R. Sveinsson, Henrik Aanæs, and Rasmus Larsen. Improved 3D reconstruction in smart-room environments using ToF imaging. Computer Vision and Image Understanding: CVIU, 114(12):1376–1384, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Grosgeorge:2013:GCS

[GPDR13] D. Grosgeorge, C. Petitjean, J.-N. Dacher, and S. Ruan. Graph cut segmentation with a statistical shape model in cardiac MRI. Computer Vision and Image Understanding: CVIU, 117(9): 1027–1035, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000738.

Garcia:1999:UBM

[GPK99] P. García, M. Petrou, and S. Kamata. The use of Boolean model for texture analysis of grey images. Computer Vision and Image Understanding: CVIU, 74(3):227–235, June 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0760/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0760/production/
pdf.

Grattoni:1988:CDA

[GPP88] Paolo Grattoni, Fabrizio Pollastri, and Amedeo Premoli. A contour detection algorithm based on the minimum radial inertia (MRI) criterion. Computer Vision, Graphics, and Image Processing, 43(1):22–36, July 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gong:2007:GOF

[GPY^{+}07] Haifeng Gong, Chunhong Pan, Qing Yang, Hanqing Lu, and Songde Ma. Generalized optical flow in the scale space. Computer Vision and Image Understanding: CVIU, 105 (1):86–92, January 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Grant:1981:EAB

[GR81] G. Grant and A. F. Reid. Efficient algorithm for boundary tracing and feature extraction. Computer Graphics and Image Processing, 17(3): 225–237, November 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Gordon:1985:ISS

[GR85] Dan Gordon and R. Anthony Reynolds. Image space shading of 3-dimensional objects. Computer Vision, Graphics, and Image Processing, 29(3):361–376, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Goldwasser:1987:RTD

[GR87a] S. M. Goldwasser and R. A. Reynolds. Real-time display and manipulation of 3-D medical objects: the voxel processor architecture. Computer Vision, Graphics, and Image Processing, 39(1):1– 27, July 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gross:1987:MOD

[GR87b] Ari David Gross and Azriel Rosenfeld. Multiresolution object detection and delineation. Computer Vision, Graphics, and Image Processing, 39(1):102–115, July 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Goresnic:1992:TCU

[GR92] C. Goresnic and S. R. Rotman. Texture classification using the cortex transform. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4):329–339, July 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Gil:2005:EAO

[GR05] Debora Gil and Petia Radeva.

Extending anisotropic operators to recover smooth shapes. Computer Vision and Image Understanding: CVIU, 99(1): 110–125, July 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Granlund:1978:SGP

[Gra78] Goesta H. Granlund. In search of a general picture processing operator. Computer Graphics and Image Processing, 8(2):155– 173, October 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Godec:2013:HBT

[GRB13] M. Godec, P. M. Roth, and H. Bischof. Hough-based tracking of non-rigid objects. Computer Vision and Image Understanding: CVIU, 117(10):1245–1256, October 2013. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001890.

Grevera:2004:DRS

[Gre04] George J. Grevera. The “dead reckoning” signed distance transform. Computer Vision and Image Understanding: CVIU, 95(3):317– 333, September 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Galbally:2013:IIR

[GRGB^{+}13] Javier Galbally, Arun Ross, Marta Gomez-Barrero, Julian Fierrez, and Javier Ortega-Garcia. Iris image reconstruction from binary templates: an efficient probabilistic approach based on genetic algorithms. Computer Vision and Image Understanding: CVIU, 117(10):1512–1525, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001070.

Grimson:1983:ICT

[Gri83a] W. E. L. Grimson. An implementation of a computational theory of visual surface interpolation. Computer Vision, Graphics, and Image Processing, 22(1):39–69, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Grimson:1983:SCC

[Gri83b] W. E. L. Grimson. Surface consistency constraints in vision. Computer Vision, Graphics, and Image Processing, 24(1):28–51, October 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Grimson:1984:BSV

[Gri84] W. E. L. Grimson. Binocular shading and visual surface reconstruction. Computer Vision, Graphics, and Image Processing, 28(1):19– 43, October 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Groch:1982:ELS

[Gro82]  Wolf Dieter Groch.  Ex-

traction of line shaped objects from aerial images using a special operator to analyze the profiles of functions. Computer Graphics and Image Processing, 18(4): 347–358, April 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Grosky:1984:TDM

[Gro84]  William I. Grosky.  To-

ward a data model for integrated pictorial databases. Computer Vision, Graphics, and Image Processing, 25(3):371–382, March 1984.

CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Goodrich:1990:SPS

[GS90] Michael T. Goodrich and Jack Scott Snoeyink. Stabbing parallel segments with a convex polygon. Computer Vision, Graphics, and Image Processing, 49(2):152– 170, February 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Goodchild:1992:HSD

[GS92] Michael F. Goodchild and Yang Shiren. A hierarchical spatial data structure for global geographical information systems. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1): 31–44, January 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Gong:1995:DRM

[GS95] Yihong Gong and Masao Sakauchi. Detection of regions matching specified chromatic features. Computer Vision and Image Understanding: CVIU, 61(2):263–269, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1018/production;   http:

//www.idealibrary.com/links/
artid/cviu.1995.1018/production/
pdf.

Gagvani:1999:PCV

[GS99] Nikhil Gagvani and Deborah Silver. Parameter-controlled volume thinning. Graphical Models and Image Processing: GMIP, 61(3):149–164, May 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0495/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0495/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0495/production/ref.

Gagvani:2001:AVM

[GS01] Nikhil Gagvani and Deborah Silver. Animating volumetric models. Graphical Models, 63 (6):443–458, November 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gauch:2006:FIU

[GS06] John M. Gauch and Abhishek Shivadas. Finding and identifying unknown commercials using repeated video sequence detection. Computer Vision and Image Understanding: CVIU, 103(1): 80–88, July 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goshtasby:2008:AWM

[GS08] Ardeshir Goshtasby and Martin Satter. An adaptive window mechanism for image smoothing. Computer Vision and Image Understanding: CVIU, 111(2):155–169, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goldman:2012:FAQ

[GS12] Ron Goldman and Plamen Simeonov. Formulas and algorithms for quantum differentiation of quantum Bernstein bases and quantum Bézier curves based on quantum blossoming. Graphical Models, 74(6):326–334, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000240.

Gupta:2002:CDG

[GSK02] Sumit Gupta, Kuntal Sengupta, and Ashraf A. Kassim. Compression of dynamic 3D geometry data using iterative closest point algorithm. Computer Vision and Image Understanding: CVIU, 87(1– 3):116–130, July 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Garcia-Sevilla:2001:AIS

[GSP01] Pedro García-Sevilla and Maria Petrou. Analysis of irregularly shaped texture regions. Computer Vision and Image Understanding: CVIU, 84(1):62–76, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0941;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0941/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0941/ref.

Ghebreab:2002:NIP

[GSP02] S. Ghebreab, A. W. M. Smeulders, and P. R. Pfluger. Necklaces: Inhomogeneous and point-enhanced deformable models. Computer Vision and Image Understanding: CVIU, 86(2):96–117, May 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Geronimo:2010:BBP

[GSPL10] David Gerónimo, Angel D.

Sappa, Daniel Ponsa, and Antonio M. López. 2D–3Dbased on-board pedestrian detection system. Computer Vision and Image Understanding: CVIU, 114(5):583–595, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gallo:2000:FBS

[GSS00] Giovanni Gallo, Michela Spagnuolo, and Salvatore Spinello. Fuzzy B-splines: a surface model encapsulating uncertainty. Graphical Models, 62(1):40–55, January 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/artid/gmod.1999.
0512/production;   http:
//www.idealibrary.com/links/
artid/gmod.1999.0512/production/
pdf;   http://www.idealibrary.
com/links/artid/gmod.1999.
0512/production/ref.

Gavves:2012:VSL

[GSS12] Efstratios Gavves, Cees G. M. Snoek, and Arnold W. M. Smeulders. Visual synonyms for landmark image retrieval. Computer Vision and Image Understanding: CVIU, 116 (2):238–249, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002153.

Goshtasby:2003:NIR

[GSST03] Ardeshir Goshtasby, Lawrence Staib, Colin Studholme, and Demetri Terzopoulos. Nonrigid image registration: guest editors’ introduction. Computer Vision and Image Understanding: CVIU, 89(2– 3):109–113, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goshtasby:2000:AVI

[GSU00] A. Ardeshir Goshtasby, Milan Sonka, and Jayaram Udupa. Analysis of volumetric images. Computer Vision and Image Understanding: CVIU, 77(2):79–83, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0819;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0819/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0819/ref.

Gracias:2000:UVM

[GSV00] Nuno Gracias and José Santos-Victor. Underwater video mosaics as visual navigation maps. Computer Vision and Image Understanding: CVIU, 79(1):66–91, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0848;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0848/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0848/ref.

Grossmann:2005:LSR

[GSV05] Etienne Grossmann and José Santos-Victor. Least-squares 3D reconstruction from one or more views and geometric clues. Computer Vision and Image Understanding: CVIU, 99(2):151–174, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Goldwasser:1984:PCC

[GT84] Samuel M. Goldwasser and Donald E. Troxel. Page composition of continuous tone imagery. Computer Vision, Graphics, and Image Processing, 26(1):30–44, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gordon:1989:FST

[GU89] Dan Gordon and Jayaram K. Udupa. Fast surface tracking in three-dimensional binary images. Computer Vision, Graphics, and Image Processing, 45(2):196– 214, February 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gudmundsson:1982:IHL

[Gud82] Bjorn Gudmundsson. Interactive high-level language system for picture processing. Computer Graphics and Image Processing, 18(4): 392–340, April 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Guiducci:1998:RRS

[Gui98] Antonio Guiducci. 3D road reconstruction from a single view. Computer Vision and Image Understanding: CVIU, 70(2):212–226, May 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0633/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0633/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0633/production/ref.

Guiducci:1999:PMP

[Gui99]  Antonio Guiducci.  Para-

metric model of the perspective projection of a road with applications to lane keeping and 3D road reconstruction. Computer Vision and Image Understanding: CVIU, 73(3):414–427, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0737/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0737/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0737/production/ref.

Guiducci:2000:CCR

[Gui00] Antonio Guiducci. Camera calibration for road applications. Computer Vision and Image Understanding: CVIU, 79(2):250–266, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0857;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0857/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0857/ref.

Gullberg:1979:RFB

[Gul79]  Grant T. Gullberg.  Re-

construction of fan-beam data by filtering the backprojection. Computer Graphics and Image Processing, 10 (1):30–47, May 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Guskov:2007:MBA

[Gus07]  Igor Guskov.  Manifold-

based approach to semiregular remeshing. Graphical Models, 69(1):1–18, January 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070306000385.

Goren:1978:FCP

[GV78] F. C. A. Goren and P. W. Verbeek. Freeman-code probabilities of object boundary quantized contours. Computer Graphics and Image Processing, 17(3):391–402, June 1978.

CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Gerritsen:1984:ICL

[GV84] Frans A. Gerritsen and Piet W. Verbeek. Implementation of cellular-logic operators using 3 × 3 convolution and table lookup hardware. Computer Vision, Graphics, and Image Processing, 27(1):115–123, July 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Go:2006:ABI

[GVK06] Jared Go, Thuc D. Vu, and James J. Kuffner. Autonomous behaviors for interactive vehicle animations. Graphical Models, 68(2):90– 112, March 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000342.

Gunther:1990:ATA

[GW90] Oliver Günther and Eugene Wong. The arc tree: an approximation scheme to represent arbitrary curved shapes. Computer Vision, Graphics, and Image Processing, 51 (3):313–337, September 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Goldman:1993:VNA

[GW93a] R. Goldman and J. Warren. Volume 55, number 1 (1993) in the article ‘An Extension of Chaiken’s Algorithm to B-Spline Curves with Knots in Geometric Progression,’ by Ron Goldman and Joe Warren, pages 58–

62. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(4):324, July 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1023/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1023/production/
pdf.

Goldman:1993:ECA

[GW93b] Ron Goldman and Joe Warren. An extension of Chaiken’s algorithm to Bspline curves with knots in geometric progression. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(1):58–62, January 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1004/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1004/production/
pdf.

Gregor:2001:ISR

[GW01] J. Gregor and R. T. Whitaker.

Indoor scene reconstruction from sets of noisy range images. Graphical Models, 63 (5):304–332, September 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gruber:2007:INM

[GW07] Amit Gruber and Yair Weiss.

Incorporating non-motion cues into 3D motion segmentation. Computer Vision and Image Understanding: CVIU, 108(3):261–271, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Guttmann:2011:CAV

[GWCO11] Moshe Guttmann, Lior Wolf, and Danny Cohen-Or. Content aware video manipulation. Computer Vision and Image Understanding: CVIU, 115(12):1662–1678, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001597.

Gryn:2009:DMP

[GWT09] Jacob M. Gryn, Richard P.

Wildes, and John K. Tsotsos. Detecting motion patterns via direction maps with application to surveillance. Computer Vision and Image Understanding: CVIU, 113(2):

291–307, February 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Griswold:1988:NSV

[GY88] N. C. Griswold and C. P. Yeh. A new stereo vision model based upon the binocular fusion concept. Computer Vision, Graphics, and Image Processing, 41(2):153– 171, February 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Gao:1999:TDD

[GY99] Qinghuai Gao and Fang-Fang Yin. Two-dimensional direction-based interpolation with local centered moments. Graphical Models and Image Processing: GMIP, 61 (6):323–339, November 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0504/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0504/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0504/production/ref.

Gong:2001:LBM

[GY01] Minglun Gong and Yee-Hong Yang. Layer-based morphing. Graphical Models, 63(1):45–59, January 2001. CODEN GRMOFM.

ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0537;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0537/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0537/ref.

Gong:2005:CFR

[GY05] Minglun Gong and Yee-Hong Yang. Camera field rendering for static and dynamic scenes. Graphical Models, 67 (2):29, March 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Gao:2013:FPS

[GYH13] Zhanheng Gao, Zeyun Yu, and Michael Holst. Featurepreserving surface mesh smoothing via suboptimal Delaunay triangulation. Graphical Models, 75(1):23–38, January 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000756.

Gao:2009:DOF

[GYTL09] Xinbo Gao, Yimin Yang, Dacheng Tao, and Xuelong Li. Discriminative optical flow tensor for video semantic analysis. Computer Vision and Image Understanding: CVIU, 113(3):372–383, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gu:2005:TOF

[GZJ05] Haisong Gu, Yongmian Zhang, and Qiang Ji. Task oriented facial behavior recognition with selective sensing. Computer Vision and Image Understanding: CVIU, 100 (3):385–415, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Gao:2005:IPG

[GZP05] Qigang Gao, Yun Zhang, and Alan Parslow. The influence of perceptual grouping on motion detection. Computer Vision and Image Understanding: CVIU, 100(3): 442–457, December 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:1993:SCD

[HA93] Xiaoping P. Hu and Narendra Ahuja. Sufficient conditions for double or unique solution of motion and structure. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(2):161–176, September 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1993.
                                        1036/production;   http:
                                        //www.idealibrary.com/links/

artid/ciun.1993.1036/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1038/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1038/production/
pdf.

Hetroy:2003:TQC

[HA03] Franck Hétroy and Dominique Attali. Topological quadrangulations of closed triangulated surfaces using the Reeb graph. Graphical Models, 65(1–3):131–148, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Haar:1982:SEO

[Haa82] Robert L. Haar. Sketching: Estimating object positions from relational descriptions. Computer Graphics and Image Processing, 19(3): 227–247, July 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Haber:1985:TTP

[Hab85] Ralph Norman Haber. Toward a theory of the perceived spatial layout of scenes. Computer Vision, Graphics, and Image Processing, 31 (3):282–321, September 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Huck:1991:IGD

[HAGR91] F. O. Huck, R. Alter-Gartenberg, and Z.-U. Rahman. Image gathering and digital restoration for fidelity and visual quality. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1): 71–84, January 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Hwang:1991:OMS

[HAKK91] Kai Hwang, H. M. Alnuweiri, V. K. Prasanna Kumar, and Dongseung Kim. Orthogonal multiprocessor sharing memory with an enhanced mesh for integrated image understanding. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (1):31–45, January 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Hamill:1977:LPM

[Ham77] Philip Hamill. Line printer modification for better gray level pictures. Computer Graphics and Image Processing, 6(5):485–491, October 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hamker:2005:EAP

[Ham05]  Fred H. Hamker.  The

emergence of attention by population-based inference and its role in distributed processing and cognitive control of vision. Computer Vision and Image Understanding: CVIU, 100(1–2):64–106, October/November 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hanson:1988:HSD

[Han88] Andrew J. Hanson. Hyperquadrics: Smoothly deformable shapes with convex polyhedral bounds. Computer Vision, Graphics, and Image Processing, 44(2):191– 210, November 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hannaford:1993:RFS

[Han93] B. Hannaford. Resolutionfirst scanning of multidimensional spaces. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(5):359–369, September 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1027/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1027/production/
pdf.

Haralick:1980:ERA

[Har80a] Robert M. Haralick. Edge and region analysis for digital image data. Computer Graphics and Image Processing, 12 (1):60–73, January 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). HARALICK80.

Haralick:1980:UPT

[Har80b] Robert M. Haralick. Using perspective transformations in scene analysis. Computer Graphics and Image Processing, 13(3):191–221, July 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Haralick:1983:IPR

[Har83a] Robert M. Haralick. An interpretation for probabilistic relaxation. Computer Vision, Graphics, and Image Processing, 22(3):388–395, June 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). HARALICK83a.

Haralick:1983:RVD

[Har83b] Robert M. Haralick. Ridges and valleys on digital images. Computer Vision, Graphics, and Image Processing, 22(1):28–38, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hartley:1985:GWM

[Har85] Ralph Hartley. A Gaussianweighted multiresolution edge detector. Computer Vision, Graphics, and Image Processing, 30(1):70–83, April 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Haralick:1986:CVT

[Har86] Robert M. Haralick. Computer vision theory: the lack thereof. Computer Vision, Graphics, and Image Processing, 36(2/3):372–386, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Haralick:1994:CPC

[Har94a] Robert M. Haralick. Comments on performance characterization replies. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):264–265, September 1994. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1056/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1056/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1061/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1061/production/

pdf. See [Har94b, DB94, Shi94].

Haralick:1994:PCC

[Har94b] Robert M. Haralick. Performance characterization in computer vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):245–249, September 1994. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1050/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1050/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1055/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1055/production/

pdf. See also [DB94, Har94a, Shi94].

Hansen:2010:FRP

[HASS10] Mark F. Hansen, Gary A. Atkinson, Lyndon N. Smith, and Melvyn L. Smith. 3D face reconstructions from photometric stereo using near infrared and visible light. Computer Vision and Image Understanding: CVIU, 114(8): 942–951, August 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hawkes:1978:EIP

[Haw78] P. W. Hawkes. Electron image processing: a survey. Computer Graphics and Image Processing, 8(3):406– 446, December 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hawkes:1982:EIP

[Haw82] P. W. Hawkes. Electron image processing: 1978-1980. Computer Graphics and Image Processing, 18(1):58– 96, January 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Horn:1986:VAS

[HB86] Berthold K. P. Horn and Michael J. Brooks. The variational approach to shape from shading. Computer Vision, Graphics, and Image Processing, 33(2):174– 208, February 1986. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Healey:1988:LSS

[HB88] Glenn Healey and Thomas O.

Binford. Local shape from specularity. Computer Vision, Graphics, and Image Processing, 42(1):62–86, April 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hohmeyer:1991:SRB

[HB91] M. E. Hohmeyer and B. A.

Barsky. Skinning rational Bspline curves to construct an interpolatory surface. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6):511–521, November 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Haala:1998:IUS

[HB98a] Norbert Haala and Claus Brenner. Interpretation of urban surface models using 2D building information. Computer Vision and Image Understanding: CVIU, 72(2):204–214, November 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0720/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0720/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0720/production/ref.

Heisterkamp:1998:MPA

[HB98b] Douglas R. Heisterkamp and Prabir Bhattacharya. Matching 2D polygonal arcs by using a subgroup of the unit quaternions. Computer Vision and Image Understanding: CVIU, 69(2):246–249, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0566/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0566/production/
pdf;   http://www.idealibrary.

com/links/artid/cviu.1997.
0566/production/ref.

Ho:1998:PCC

[HB98c] Tin Kam Ho and Henry S.

Baird. Pattern classification with compact distribution maps. Computer Vision and Image Understanding: CVIU, 70(1):101–110, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0624/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0624/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0624/production/ref.

Hall:2005:SIV

[HB05] Peter Hall and Brian Barsky.

Special issue: Vision, video and graphics 2003. Graphical Models, 67(6):475, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000305.

Huddleston:1993:GES

[HBA93] James N. Huddleston and Jezekiel Ben-Arie. Grouping edgels into structural entities using circular symmetry, the distributed Hough transform, and probabilistic non-accidentalness. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):227– 242, March 1993. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1015/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1015/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1015/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1015/production/
pdf.

Hassan:2012:ASC

[HBB^{+}12] Waqas Hassan, Nagachetan Bangalore, Philip Birch, Rupert Young, and Chris Chatwin. An adaptive sample count particle filter. Computer Vision and Image Understanding: CVIU, 116(12): 1208–1222, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001191.

Hollingsworth:2009:PDD

[HBF09] Karen Hollingsworth, Kevin W. Bowyer, and Patrick J. Flynn. Pupil dilation degrades iris biometric performance. Computer Vision and Image Understanding: CVIU, 113(1): 150–157, January 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hosni:2013:SAS

[HBG13] Asmaa Hosni, Michael Bleyer, and Margrit Gelautz. Secrets of adaptive support weight techniques for local stereo matching. Computer Vision and Image Understanding: CVIU, 117(6):620–632, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000143.

Hesami:2010:RSL

[HBH10] Reyhaneh Hesami, Alireza BabHadiashar, and Reza HosseinNezhad. Range segmentation of large building exteriors: a hierarchical robust approach. Computer Vision and Image Understanding: CVIU, 114(4):475–490, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hoseinnezhad:2011:EHB

[HBH11] Reza Hoseinnezhad and Alireza Bab-Hadiashar. An Mestimator for high breakdown robust estimation in computer vision. Computer Vision and Image Understanding: CVIU, 115(8):1145–1156, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000956.

Hatamian:1987:AL

[HBKN87] M. Hatamian, Z. L. Budrikis, P. S. Kubik, and A. N. Netravali. Accurate lightpen. Computer Vision, Graphics, and Image Processing, 39 (2):246–257, August 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hollingsworth:2011:GII

[HBL^{+}11] Karen Hollingsworth, Kevin W. Bowyer, Stephen Lagree, Samuel P. Fenker, and Patrick J. Flynn. Genetically identical irises have texture similarity that is not detected by iris biometrics. Computer Vision and Image Understanding: CVIU, 115(11):1493–1502, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421100155X.

Haralick:1977:IDE

[HC77] Robert M. Haralick and Phil Currier. Image discrimination enhancement combination system (idecs). Computer Graphics and Image Processing, 6(4):371–381, August 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hsu:1994:VLX

[HC94] J. Hsu and D. M. Chelberg. Visible light and X-ray ray tracing of generalized cylinders. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(5):392–401, September 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1035/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1035/production/
pdf.

Huang:1996:DMA

[HC96] Chung-Lin Huang and Kou-Chang Chen. Directional moving averaging interpolation for texture mapping. Graphical Models and Image Processing: GMIP, 58 (4):301–313, July 1996. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0025/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0025/production/
pdf.

Heidary:2013:PEA

[HC13a] Kaveh Heidary and H. John Caulfield. Presmoothing effects in artificial color image segmentation. Computer Vision and Image Understanding: CVIU, 117 (3):195–201, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001932.

Hu:2013:PEG

[HC13b] Rui Hu and John Collomosse. A performance evaluation of gradient field HOG descriptor for sketch based image retrieval. Computer Vision and Image Understanding: CVIU, 117(7):790–806, July 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000349.

Hui:2013:DSM

[HC13c] Tak-Wai Hui and Ronald Chung. Determining shape and motion from non-overlapping multi-camera rig: a direct approach using normal flows. Computer Vision and Image Understanding: CVIU, 117(8):947–964, August 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000933.

Haritaoglu:2001:BDP

[HCHD01] Ismail Haritaoglu, Ross Cutler, David Harwood, and Larry S. Davis. Backpack: Detection of people carrying objects using silhouettes. Computer Vision and Image Understanding: CVIU, 81(3):385–397, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0893;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0893/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0893/ref.

Horaud:1989:ASP

[HCLL89] Radu Horaud, Bernard Conio, Olivier Leboulleux, and Bernard Lacolle. An analytic solution for the perspective 4-point problem. Computer Vision, Graphics, and Image Processing, 47(1):33– 44, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Huertas:1990:DRC

[HCN90] A. Huertas, W. Cole, and R. Nevatia. Detecting runways in complex airport scenes. Computer Vision, Graphics, and Image Processing, 51(2):107–145, August 1990. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hu:2003:SIP

[HCS03] Shi-Min Hu, Sabine Coquillart, and Heung-Yeung Shum. Special issue on Pacific Graphics 2002. Graphical Models, 65(4):169–170, July 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Handley:1997:MLE

[HD97] John C. Handley and Edward R. Dougherty. Maximumlikelihood estimation for the two-dimensional discrete Boolean random set and function models using multidimensional linear samples. Graphical Models and Image Processing: GMIP, 59(4):221–231, July 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0432/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0432/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0432/production/ref.

Hammoud:2007:AVA

[HD07] Riad Ibrahim Hammoud and James W. Davis. Advances in vision algorithms and systems beyond the visible spectrum. Computer Vision and Image Understanding: CVIU, 106(2–3):145–147, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Han:2009:PFB

[HD09] Bohyung Han and Larry S. Davis. Probabilistic fusionbased parameter estimation for visual tracking. Computer Vision and Image Understanding: CVIU, 113(4): 435–445, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Huang:2012:PRF

[HDF12] Chen Huang, Xiaoqing Ding, and Chi Fang. Pose robust face tracking by combining view-based AAMs and temporal filters. Computer Vision and Image Understanding: CVIU, 116(7):777–792, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000471.

Hwang:1986:HII

[HDM86] Vincent Shang-Shouq Hwang, Larry S. Davis, and Takashi Matsuyama. Hypothesis integration in image understanding systems. Computer Vision, Graphics, and Image Processing, 36(2/3):321–371, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hammouche:2008:MAT

[HDS08] Kamal Hammouche, Moussa Diaf, and Patrick Siarry. A multilevel automatic thresholding method based on a genetic algorithm for a fast image segmentation. Computer Vision and Image Understanding: CVIU, 109(2):

163–175, February 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Haering:1999:FCM

[HdVL99] Niels Haering and Niels da Vitoria Lobo. Features and classification methods to locate deciduous trees in images. Computer Vision and Image Understanding: CVIU, 75 (1–2):133–149, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0769/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0769/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0769/production/ref.

Herbison-Evans:1981:FAF

[HE81] D. Herbison-Evans. Fast algorithm for finding lines in pictures. Computer Graphics and Image Processing, 17(3):281– 289, November 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hanson:1982:ISU

[HE82] F. R. Hanson and H. Elliott. Image segmentation using simple Markov field models. Computer Graphics and Image Processing, 20(2):101– 132, October 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). HANSON82.

Heijmans:1999:CMO

[Hei99] Henk J. A. M. Heijmans.

Connected morphological operators for binary images. Computer Vision and Image Understanding: CVIU, 73(1):99–120, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0703/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0703/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0703/production/ref.

Heidemann:2004:CSC

[Hei04] Gunther Heidemann. Combining spatial and colour information for content based image retrieval. Computer Vision and Image Understanding: CVIU, 94(1–3):234– 270, April/June 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Henderson:1984:NDR

[Hen84] Thomas C. Henderson. A note on discrete relaxation. Computer Vision, Graphics, and Image Processing, 28 (3):384–388, December 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Henricsson:1998:RCA

[Hen98] Olof Henricsson. The role of color attributes and similarity grouping in 3D building reconstruction. Computer Vision and Image Understanding: CVIU, 72(2): 163–184, November 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0718/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0718/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0718/production/ref.

Hershey:1972:CSS

[Her72] Allen V. Hershey. A computer system for scientific typography. Computer Graphics and Image Processing, 1(3): 373–385, November 1972. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Herman:1980:NIP

[Her80]  Gabor T. Herman.  On

the noise in images produced by computed tomography. Computer Graphics and Image Processing, 12(3): 271–285, March 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Herbison-Evans:1981:CRP

[HER81] D. Herbison-Evans and D. S.

Richardson. Control of round-off propagation in articulating the human figure. Computer Graphics and Image Processing, 17(4):386– 393, December 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Herman:1990:TAI

[Her90] G. T. Herman. On topology as applied to image analysis. Computer Vision, Graphics, and Image Processing, 52 (3):409–415, December 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Herman:1992:DMJ

[Her92]  Gabor T. Herman.  Dis-

crete multidimensional Jordan surfaces. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(6): 507–515, November 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Herman:1993:OSD

[Her93] G. T. Herman. Oriented surfaces in digital spaces. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(5):381–396, September 1993. CODEN CGMPE5.

ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1029/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1029/production/
pdf.

Herman:1998:FSC

[Her98] Gabor T. Herman. Finitary 1-simply connected digital spaces. Graphical Models and Image Processing: GMIP, 60(1):046–056, January 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0456/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0456/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0456/production/ref.

Heygster:1982:RFD

[Hey82] Georg Heygster. Rank filters in digital image processing. Computer Graphics and Image Processing, 19(2): 148–164, June 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hsieh:1980:AVI

[HF80] Y. Y. Hsieh and K. S. Fu. Automatic visual inspection system for integrated circuit chips. Computer Graphics and Image Processing, 14(4):293– 343, December 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Houle:1993:LSM

[HF93] C. Houle and E. Fiume. Lightsource modeling using pyramidal light maps. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(5):346–358, September 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1026/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1026/production/
pdf.

Hilton:2001:MPT

[HF01] Adrian Hilton and Pascal Fua. Modeling people toward vision-based understanding of a person’s shape, appearance, and movement. Computer Vision and Image Understanding: CVIU, 81(3):227–230, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0907;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0907/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0907/ref.

Hillenbrand:2011:ESF

[HF11] Ulrich Hillenbrand and Alexander Fuchs. An experimental study of four variants of pose clustering from dense range data. Computer Vision and Image Understanding: CVIU, 115(10):1427–1448, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001445.

Hayat:1996:CFP

[HFC96] L. Hayat, M. Fleury, and A. F. Clark. Candidate functions for a parallel multi-level thresholding technique. Graphical Models and Image Processing: GMIP, 58(4):360–381, July 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0029/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0029/production/
                                        pdf.

Howell:1993:QCS

[HFF93] Gary W. Howell, Donald W. Fausett, and Laurene V. Fausett. Quasicircular splines: a shapepreserving approximation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(2):89–97, March 1993. CODEN CGMPE5.

ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1007/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1007/production/
pdf.

Haag:1997:IEM

[HFKN97] Michael Haag, Thomas Frank, Henner Kollnig, and Hans-Hellmut Nagel. Influence of an explicitly modelled 3D scene on the tracking of partially occluded vehicles. Computer Vision and Image Understanding: CVIU, 65(2):206–225, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0575/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0575/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0575/production/ref.

Hilton:2006:MPV

[HFR06] A. Hilton, P. Fua, and R. Ronfard. Modeling people: Vision-based understanding of a personś shape, appearance, movement, and behaviour. Computer Vision and Image Understanding: CVIU, 104(2–3):87–89, November/December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hofmann:2011:HMA

[HG11] Michael Hofmann and Dariu M. Gavrila. 3D Human model adaptation by frame selection and shape-texture optimization. Computer Vision and Image Understanding: CVIU, 115(11):1559–1570, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001846.

Harms:1986:CLC

[HGA86] H. Harms, U. Gunzer, and H. M. Aus. Combined local color and texture analysis of stained cells. Computer Vision, Graphics, and Image Processing, 33(3):364– 376, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hoover:1998:SER

[HGB98] Adam Hoover, Dmitry Goldgof, and Kevin W. Bowyer. The space envelope: a representation for 3D scenes. Computer Vision and Image Understanding: CVIU, 69(3):310–329, March 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0666/production;   http:

//www.idealibrary.com/links/
artid/cviu.1998.0666/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0666/production/ref.

Heber:2013:SBT

[HGR^{+}13] Markus Heber, Martin Godec, Matthias Rüther, Peter M. Roth, and Horst Bischof. Segmentation-based tracking by support fusion. Computer Vision and Image Understanding: CVIU, 117(6): 573–586, June 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000301.

Hurtut:2008:AIR

[HGS08] Thomas Hurtut, Yann Gousseau, and Francis Schmitt. Adaptive image retrieval based on the spatial organization of colors. Computer Vision and Image Understanding: CVIU, 112(2):101–113, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hilton:2011:SII

[HGSM11] Adrian Hilton, Guy Godin, Chang Shu, and Takeshi Masuda. Special issue on 3D imaging and modelling. Computer Vision and Image Understanding: CVIU, 115(5): 559–560, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Harauz:1987:OSP

[HGv87] George Harauz, Richard Gordon, and Marin van Heel. Oblique sampling of projections for direct threedimensional reconstruction. Computer Vision, Graphics, and Image Processing, 38(1): 81–89, April 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Harmon:1977:ARH

[HH77] Leon D. Harmon and Willard F. Hunt. Automatic recognition of human face profiles. Computer Graphics and Image Processing, 6(2): 135–156, April 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hwang:1982:MFO

[HH82] J. J. Hwang and E. L. Hall. Matching of featured objects using relational tables from stereo images. Computer Graphics and Image Processing, 20(2):22– 42, September 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Horn:1991:RBM

[HH91] Berthold K. P. Horn and J. G. Harris. Rigid body motion from range image sequences. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (1):1–13, January 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Hernandez:1996:CAE

[HH96] Gonzalo Hernández and Hans J.

Herrmann. Cellular automata for elementary image enhancement. Graphical Models and Image Processing: GMIP, 58(1):82–89, January 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0006/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0006/production/
pdf.

Huang:1997:AEJ

[HH97] Ho-Chao Huang and Yi-Ping Hung. Adaptive early jumpout technique for fast motion estimation in video coding. Graphical Models and Image Processing: GMIP, 59 (6):388–394, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0449/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0449/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/gmip.1997.
                                        0449/production/ref.

Huang:1998:PSI

[HH98] Ho-Chao Huang and Yi-Ping Hung. Panoramic stereo imaging system with automatic disparity warping and seaming. Graphical Models and Image Processing: GMIP, 60(3):196–208, May 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0467/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1998.0467/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0467/production/ref.

Heinke:2005:SAI

[HH05] Dietmar Heinke and Glyn W. Humphreys. Selective Attention for Identification Model: Simulating visual neglect. Computer Vision and Image Understanding: CVIU, 100(1–2):172–197, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hansen:2007:ILM

[HH07] Dan Witzner Hansen and Riad I. Hammoud. An improved likelihood model for eye tracking. Computer Vision and Image Understanding: CVIU, 106(2–3):220– 230, May/June 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Huo:2012:MPT

[HH12] Feifei Huo and Emile A.

Hendriks. Multiple people tracking and pose estimation with occlusion estimation. Computer Vision and Image Understanding: CVIU, 116(5):634–647, May 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000033.

Higuchi:1995:BDM

[HHI95] K. Higuchi, M. Hebert, and K. Ikeuchi. Building 3-D models from unregistered range images. Graphical Models and Image Processing: GMIP, 57(4):315–333, July 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1028/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1028/production/
pdf.

Heo:2001:ITR

[HHS^{+}01] Hee-Seok Heo, Sung Je Hong, Joon-Kyung Seong, Myung-Soo Kim, and Gershon Elber. The intersection of two ringed surfaces and some related problems. Graphical Models, 63(4):228–244, July 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Heisele:2003:FRC

[HHWP03] Bernd Heisele, Purdy Ho, Jane Wu, and Tomaso Poggio. Face recognition: componentbased versus global approaches. Computer Vision and Image Understanding: CVIU, 91(1–2):6–21, July/August 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hildreth:1983:DIC

[Hil83] Ellen C. Hildreth. The detection of intensity changes by computer and biological vision systems. Computer Vision, Graphics, and Image Processing, 22(1):1– 27, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Haynes:1983:DME

[HJ83] Susan M. Haynes and Ramesh Jain. Detection of moving edges. Computer Vision, Graphics, and Image Processing, 21(3):345– 367, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Harrison:2012:TPA

[HJ12] Adam P. Harrison and Dileepan Joseph. Translational photometric alignment of single-view image sequences. Computer Vision and Image Understanding: CVIU, 116(6):765–776, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000227.

Hyun:2002:MDA

[HJK02] Dae-Eun Hyun, Bert Jüttler, and Myung-Soo Kim. Minimizing the distortion of affine spline motions. Graphical Models, 64(2):128–144, March 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Huang:1989:BIA

[HJS89] K. S. Huang, B. K. Jenkins, and A. A. Sawchuk. Binary image algebra and optical cellular logic processor design. Computer Vision, Graphics, and Image Processing, 45(3):295–345, March 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Haberstroh:1993:LDN

[HK93] Richard Haberstroh and Ludwik Kurz. Line detection in noisy and structured backgrounds using Graeco-Latin squares. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(3):161–179, May 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1012/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1012/production/
pdf.

Huq:2013:OFS

[HKA13] Shafik Huq, Andreas Koschan, and Mongi Abidi. Occlusion filling in stereo: Theory and experiments. Computer Vision and Image Understanding: CVIU, 117(6):688–704, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000155.

Haralick:1995:MBM

[HKD95] Robert M. Haralick, Philip L. Katz, and Edward R. Dougherty. Model-based morphology: The opening spectrum. Graphical Models and Image Processing: GMIP, 57(1):1– 12, January 1995. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1001/production;   http:
//www.idealibrary.com/links/

artid/gmip.1995.1001/production/
pdf.

Hwang:2008:CDU

[HKK08] Youngbae Hwang, Jun-Sik Kim, and In-So Kweon. Change detection using a statistical model in an optimally selected color space. Computer Vision and Image Understanding: CVIU, 112(3): 231–242, December 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hanniel:2012:CHD

[HKM12] Iddo Hanniel, Adarsh Krishnamurthy, and Sara McMains. Computing the Hausdorff distance between NURBS surfaces using numerical iteration on the GPU. Graphical Models, 74(4):255–264, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000343.

Hasinoff:2006:BMV

[HKS06] Samuel W. Hasinoff, Sing Bing Kang, and Richard Szeliski. Boundary matting for view synthesis. Computer Vision and Image Understanding: CVIU, 103(1):22–32, July 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hummel:1987:DGB

[HKZ87] Robert A. Hummel, B. Kimia, and Steven W. Zucker. Deblurring Gaussian blur. Computer Vision, Graphics, and Image Processing, 38(1):66– 80, April 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Herman:1976:QOI

[HL76] Gabor T. Herman and Arnold Lent. Quadratic optimization for image reconstruction — 1. Computer Graphics and Image Processing, 5(3):319– 332, September 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Herman:1978:DBS

[HL78] G. T. Herman and H. K. Liu. Dynamic boundary surface detection. Computer Graphics and Image Processing, 7(1): 130–138, February 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Herman:1979:TDD

[HL79] Gabor T. Herman and Hsun Kao Liu. Three-dimensional display of human organs from computed tomograms. Computer Graphics and Image Processing, 9(1):1–21, January 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Haussmann:1984:REA

[HL84] G. Haussmann and C. E.

Liedtke. A region extraction approach to blood smear segmentation. Computer Vision, Graphics, and Image Processing, 25(2):133– 150, February 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hjelmaas:2001:FDS

[HL01] Erik Hjelmås and Boon Kee Low. Face detection: a survey. Computer Vision and Image Understanding: CVIU, 83(3):236–274, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0921;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0921/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0921/ref.

Han:2013:BIT

[HL13] Yina Han and Guizhong Liu. Biologically inspired task oriented gist model for scene classification. Computer Vision and Image Understanding: CVIU, 117 (1):76–95, January 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421200135X.

Hsieh:1997:IRU

[HLF^{+}97] Jun-Wei Hsieh, Hong-Yuan Mark Liao, Kuo-Chin Fan, Ming-Tat Ko, and Yi-Ping Hung. Image registration using a new edge-based approach. Computer Vision and Image Understanding: CVIU, 67(2):112–130, August 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0517/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0517/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0517/production/ref.

Hsieh:1995:WBS

[HLKF95] Jun-Wei Hsieh, Hong-Yuan Mark Liao, Ming-Tat Ko, and Kuo-Chin Fan. Wavelet-based shape from shading. Graphical Models and Image Processing: GMIP, 57(4):343–362, July 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1030/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1030/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1995.
1031/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1031/production/
pdf.

Haralick:1978:KII

[HM78] Robert M. Haralick and Gary Minden. Kandidats: an interactive image processing system. Computer Graphics and Image Processing, 8 (1):1–15, August 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hahn:1984:SET

[HM84] Saul Hahn and Eugenio E.

Mendoza. Simple enhancement techniques in digital image processing. Computer Vision, Graphics, and Image Processing, 26(2):233– 241, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hartley:1997:RPR

[HM97] Richard I. Hartley and Roger Mohr. Reply to pizlo, rosenfeld, and weiss. Computer Vision and Image Understanding: CVIU, 67(3): 320–323, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0644/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0644/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0644/production/ref.

Hu:2013:PSI

[HM13] Shi-Min Hu and Ralph R. Martin. Preface of special issue on computational visual media. Graphical Models, 75(3):103, May 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000088.

Huan:2010:IRB

[HMA10] Xiaoli Huan, Beddhu Murali, and Adel L. Ali. Image restoration based on the fast marching method and block based sampling. Computer Vision and Image Understanding: CVIU, 114(8): 847–856, August 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hanbury:2010:SII

[HMC10] Allan Hanbury, Henning Müller, and Paul Clough. Special issue on image and video retrieval evaluation. Computer Vision and Image Understanding: CVIU, 114(4): 409–410, April 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hooks:1993:RTP

[HMD93] J. T. Hooks, Jr., G. J. Martinsen, and V. Devarajan. On 3D real-time perspective generation from a multiresolution photo-mosaic data base. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(5):333–345, September 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1025/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1025/production/
pdf.

Hinz:2007:TMS

[HMEB07] Stefan Hinz, Franz Meyer, Michael Eineder, and Richard Bamler. Traffic monitoring with spaceborne SAR-Theory, simulations, and experiments. Computer Vision and Image Understanding: CVIU, 106(2–3):231–244, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hernandez-Mederos:2013:GIS

[HMESI13] Victoria Hernández-Mederos, Jorge Estrada-Sarlabous, and Ioannis Ivrissimtzis. Generalization of the incenter subdivision scheme. Graphical Models, 75(2):79–89, March 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000781.

Holte:2010:VIG

[HMF10] M. B. Holte, T. B. Moeslund, and P. Fihl. Viewinvariant gesture recognition using 3D optical flow and harmonic motion context. Computer Vision and Image Understanding: CVIU, 114(12): 1353–1361, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Harada:1982:IFP

[HN82] K. Harada and E. Nakamae. An isotropic four-point interpolation based on cubic splines. Computer Graphics and Image Processing, 20(3): 283–287, November 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Huertas:1988:DBA

[HN88] A. Huertas and R. Nevatia. Detecting buildings in aerial images. Computer Vision, Graphics, and Image Processing, 41(2):131– 152, February 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Holt:1991:CCP

[HN91] Robert J. Holt and Arun N. Netravali. Camera calibration problem: Some new results. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (3):368–383, November 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Holt:1995:UAI

[HN95] Robert J. Holt and Arun N.

Netravali. Using affine invariants on perspective projections of plane curves. Computer Vision and Image Understanding: CVIU, 61(1):112–121, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1008/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1008/production/
pdf.

Hongeng:2004:VBE

[HNB04] Somboon Hongeng, Ram Nevatia, and Francois Bremond. Video-based event recognition: activity representation and probabilistic recognition methods. Computer Vision and Image Understanding: CVIU, 96(2): 129–162, November 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hurley:2005:FFF

[HNC05] David J. Hurley, Mark S.

Nixon, and John N. Carter. Force field feature extraction for ear biometrics. Computer Vision and Image Understanding: CVIU, 98(3): 491–512, June 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hsu:1984:NLT

[HNR84] Y. Z. Hsu, H. H. Nagel, and G. Rekers. New likelihood test methods for change detection in image sequences. Computer Vision, Graphics, and Image Processing, 26(1):73–106, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hunt:1988:PHT

[HNR88] Douglas J. Hunt, Loren W. Nolte, and W. Howard Ruedger. Performance of the Hough transform and its relationship to statistical signal detection theory. Computer Vision, Graphics, and Image Processing, 43 (2):221–238, August 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hunt:1990:HTS

[HNRR90] D. J. Hunt, L. W. Nolte, A. R. Reibman, and W. H. Ruedger. Hough transform and signal detection theory performance for images with additive noise. Computer Vision, Graphics, and Image Processing, 52 (3):386–401, December 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hertz:1976:IJP

[HO76] C. H. Hertz and T. Orhaug.

The ink jet plotter: a computer peripheral for producing hard copy color imagery. Computer Graphics and Image Processing, 5(1): 1–12, March 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hobby:1997:SEO

[Hob97]  John D. Hobby.  Space-

efficient outlines from image data via vertex minimization and grid constraints. Graphical Models and Image Processing: GMIP, 59(2):73–88, March 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0419/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0419/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0419/production/ref.

Hobby:2000:USL

[Hob00] John D. Hobby. Using shape and layout information to find signatures, text, and graphics. Computer Vision and Image Understanding: CVIU, 80(1):88–110, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0868;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0868/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0868/ref.

Hochberg:1987:MSS

[Hoc87] Julian Hochberg. Machines should not see as people do, but must know how people see. Computer Vision, Graphics, and Image Processing, 37 (2):221–237, February 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hodges:1995:FPO

[Hod95]  Jack Hodges.  Functional

and physical object characteristics and object recognition in improvisation. Computer Vision and Image Understanding: CVIU, 62(2): 147–163, September 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1046/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1046/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1995.
1047/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1047/production/
pdf.

Hernandez:2007:VLT

[HOH^{+}07] Benjamín Hernández, Gustavo Olague, Riad Hammoud, Leonardo Trujillo, and Eva Romero. Visual learning of texture descriptors for facial expression recognition in thermal imagery. Computer Vision and Image Understanding: CVIU, 106(2–3):258– 269, May/June 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hel-Or:1991:CRH

[HOPA91] Y. Hel-Or, S. Peleg, and D. Avnir. Characterization of right-handed and left-handed shapes. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (3):297–302, May 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Horn:1974:DLI

[Hor74] Berthold K. P. Horn. Determining lightness from an image. Computer Graphics and Image Processing, 3(1): 277–299, December 1974. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Horn:1976:CGD

[Hor76] Berthold K. P. Horn. Circle generators for display devices. Computer Graphics and Image Processing, 5(2):

280–288, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). See also [Hor79, Kul79a].

Hord:1977:EDI

[Hor77] R. Michael Hord. Enhancing digital images for maximum interpretability using linear programming. Computer Graphics and Image Processing, 6(3):295–306, June 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Horn:1979:CNP

[Hor79] Berthold K. P. Horn. Comment on: “A note on the paper by B. K. P. Horn” by Zenon Kulpa. Computer Graphics and Image Processing, 9(1):104–??, January 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). See [Hor76, Kul79a].

Horn:1984:ERC

[Hor84] Berthold K. P. Horn. Exact reproduction of colored images. Computer Vision, Graphics, and Image Processing, 26(2):135–167, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Horowitz:1978:GTA

[HP78] Steven L. Horowitz and Theodosios Pavlidis. Graphtheoretic approach to picture processing. Computer Graphics and Image Processing, 7 (2):282–291, April 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hafford:1984:TDS

[HP84] Kimberly Jyl Hafford and Kendall Preston, Jr. Threedimensional skeletonization of elongated solids. Computer Vision, Graphics, and Image Processing, 27(1):78– 91, July 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hu:1996:HAE

[HP96] Jianying Hu and Theo Pavlidis.

A hierarchical approach to efficient curvilinear object searching. Computer Vision and Image Understanding: CVIU, 63(2):208–220, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0015/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0015/production/
pdf.

Hansen:2005:ETW

[HP05] Dan Witzner Hansen and Arthur E. C. Pece. Eye tracking in the wild. Computer Vision and Image Understanding: CVIU, 98(1):155–181, April 2005. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-

235X (electronic).

Heitz:1994:MMG

[HPB94] F. Heitz, P. Perez, and P. Bouthemy. Multiscale minimization of global energy functions in some visual recovery problems. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):125–134, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1008/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1994.1008/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1008/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1008/production/
                                        pdf.

Hemminger:1990:PRM

[HPR90] T. L. Hemminger and C. A. Pomalaza-Raez. Polygonal representation: a maximum likelihood approach. Computer Vision, Graphics, and Image Processing, 52 (2):239–247, November 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hoey:2010:AHA

[HPvB^{+}10] Jesse Hoey, Pascal Poupart, Axel von Bertoldi, Tammy Craig, Craig Boutilier, and Alex Mihailidis. Automated handwashing assistance for persons with dementia using video and a partially observable Markov decision process. Computer Vision and Image Understanding: CVIU, 114 (5):503–519, May 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Haralick:1982:UED

[HQ82] Robert M. Haralick and David Queeney. Understanding engineering drawings. Computer Graphics and Image Processing, 20(3):244– 258, November 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hou:2012:CDM

[HQ12a] Tingbo Hou and Hong Qin.

Continuous and discrete Mexican hat wavelet transforms on manifolds. Graphical Models, 74(4):221–232, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000306.

See corrigendum [HQ12b].

Hou:2012:CSD

[HQ12b] Tingbo Hou and Hong Qin.

Corrigendum to “Continuous and discrete Mexican hat wavelet transforms on manifolds” [Graphical Models 74 (2012) 221–232]. Graphical Models, 74(6):373, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000410.

See [HQ12a].

Hu:2005:FCC

[HQN05] Qingmao Hu, Guoyu Qian, and Wieslaw L. Nowinski. Fast connected-component labelling in three-dimensional binary images based on iterative recursion. Computer Vision and Image Understanding: CVIU, 99(3):414– 434, September 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:2012:SMU

[HQW^{+}12] Tingbo Hu, Baojun Qi, Tao Wu, Xin Xu, and Hangen He. Stereo matching using weighted dynamic programming on a single-direction four-connected tree. Computer Vision and Image Understanding: CVIU, 116 (8):908–921, August 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000641.

Heijmans:1990:ABM

[HR90] H. J. A. M. Heijmans and C. Ronse. The algebraic basis of mathematical morphology. I Dilations and erosions. Computer Vision, Graphics, and Image Processing, 50(3):245–295, June 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hausler:1999:FBO

[HR99] G. Häusler and D. Ritter.

Feature-based object recognition and localization in 3Dspace, using a single video image. Computer Vision and Image Understanding: CVIU, 73(1):64–81, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0704/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0704/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0704/production/ref.

Hu:2009:AMF

[HRC09] Yiqun Hu, Deepu Rajan, and Liang-Tien Chia. Attentionfrom-motion: a factorization approach for detecting attention objects in motion. Computer Vision and Image Understanding: CVIU, 113(3): 319–331, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hartkens:2002:EOD

[HRS02] Thomas Hartkens, Karl Rohr, and H. Siegfried Stiehl. Evaluation of 3D operators for the detection of anatomical point landmarks in MR and CT images. Computer Vision and Image Understanding: CVIU, 86(2):118–136, May 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hunter:1979:LTP

[HS79] G. M. Hunter and K. Steiglitz. Linear transformation of pictures represented by quad trees. Computer Graphics and Image Processing, 10(3): 289–296, July 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hassner:1980:UMR

[HS80] Martin Hassner and Jack Sklansky. The use of Markov random fields as models of texture. Computer Graphics and Image Processing, 12(4): 357–370, April 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). HASSNER80.

Hwang:1983:VAF

[HS83] Kai Hwang and Shun-Piao Su. Vlsi architectures for feature extraction and pattern classification. Computer Vision, Graphics, and Image Processing, 21(3):215– 228, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Haralick:1985:IST

[HS85] Robert M. Haralick and Linda G. Shapiro. Image segmentation techniques. Computer Vision, Graphics, and Image Processing, 29 (1):100–132, January 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). HARALICK85.

Hashimoto:1987:MOD

[HS87] M. Hashimoto and J. Sklansky. Multiple-order derivatives for detecting local image characteristics. Computer Vision, Graphics, and Image Processing, 39(1):28– 55, July 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hertz:1988:MTU

[HS88] Lois Hertz and Ronald W.

Schafer. Multilevel thresholding using edge matching. Computer Vision, Graphics, and Image Processing, 44 (3):279–295, December 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hsiao:1989:UTI

[HS89] John Y. Hsiao and Alexander A. Sawchuk. Unsupervised textured image segmentation using feature smoothing and probabilistic relaxation techniques. Computer Vision, Graphics, and Image Processing, 48(1):1–21, October 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hakimi:1991:FPF

[HS91] S. L. Hakimi and E. F. Schmeichel. Fitting polygonal functions to a set of points in the plane. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2): 132–136, March 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Hartley:1997:T

[HS97] Richard I. Hartley and Peter Sturm. Triangulation. Computer Vision and Image Understanding: CVIU, 68 (2):146–157, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0547/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0547/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0547/production/ref.

Hoffmann:2005:ASS

[HS05] C. M. Hoffmann and N. F. Stewart. Accuracy and semantics in shape-interrogation applications. Graphical Models, 67(5):373–389, September 2005. CODEN GRMOFM.

ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000032.

Heo:2006:FRM

[HS06] Giseon Heo and Christopher G. Small. Form representations and means for landmarks: a survey and comparative study. Computer Vision and Image Understanding: CVIU, 102(2):188–203, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Harwood:1985:TCL

[HSD85] David Harwood, Muralidhara Subbarao, and Larry S. Davis. Texture classification by local rank correlation. Computer Vision, Graphics, and Image Processing, 32 (3):404–411, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hannuksela:2007:VBM

[HSH07] Jari Hannuksela, Pekka Sangi, and Janne Heikkilä. Visionbased motion estimation for interaction with mobile devices. Computer Vision and Image Understanding: CVIU, 108(1–2):188–195, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hilton:1998:ISB

[HSIW98] A. Hilton, A. J. Stoddart, J. Illingworth, and T. Windeatt. Implicit surface-based geometric fusion. Computer Vision and Image Understanding: CVIU, 69(3):273–291, March 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0664/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0664/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0664/production/ref.

Huhle:2010:FRC

[HSJS10] Benjamin Huhle, Timo Schairer, Philipp Jenke, and Wolfgang Straßer. Fusion of range and color images for denoising and resolution enhancement with a non-local filter. Computer Vision and Image Understanding: CVIU, 114(12): 1336–1345, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Helferty:2007:CBS

[HSKH07] J. P. Helferty, A. J. Sherbondy, A. P. Kiraly, and W. E. Higgins. Computerbased system for the virtualendoscopic guidance of bronchoscopy. Computer Vision and Image Understanding: CVIU, 108(1–2):171– 187, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Heath:1998:CED

[HSSB98] M. Heath, S. Sarkar, T. Sanocki, and K. Bowyer. Comparison of edge detectors. A methodology and initial study. Computer Vision and Image Understanding: CVIU, 69(1): 38–??, ???? 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hachicha:1989:UGL

[HSSH89] A. Hachicha, S. Simon, J. Samson, and K. Hanna. The use of gray-level information and fitting techniques for precise measurement of corneal curvature and thickness. Computer Vision, Graphics, and Image Processing, 47(2):131– 164, August 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Hsu:1979:MCG

[Hsu79] Shin-Yi Hsu. Mahalanobis classifier with the generalized inverse approach for automated analysis of imagery texture data. Computer Graphics and Image Processing, 9(2):117–134, February 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Huang:1988:STE

[HT88] Jun S. Huang and Dong H. Tseng. Statistical theory of edge detection. Computer Vision, Graphics, and Image Processing, 43(3):337– 346, September 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Horn:1989:TDS

[HT89] William P. Horn and Dean L. Taylor. A theorem to determine the spatial containment of a point in a planar polyhedron. Computer Vision, Graphics, and Image Processing, 45(1):106– 116, January 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Heijmans:1991:MS

[HT91] H. J. A. M. Heijmans and A. Toet. Morphological sampling. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (3):384–400, November 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Hager:1998:XVP

[HT98] Gregory D. Hager and Kentaro Toyama. X vision: a portable substrate for real-time vision applications. Computer Vision and Image Understanding: CVIU, 69(1):023–037, January 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0561/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0561/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0561/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0586/production;
 http://www.idealibrary.
com/links/artid/cviu.1997.
0586/production/pdf;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0586/production/
ref;   http://www.idealibrary.
com/links/artid/cviu.1997.
0587/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0587/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0587/production/ref.

Hancock:2011:SIG

[HTEB11] Edwin R. Hancock, Andrea Torsello, Francisco Escolano, and Luc Brun. Special issue on graph-based representations in computer vision. Computer Vision and Image Understanding: CVIU, 115(7):903–904, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001093.

Hu:2008:IBN

[Hu08] Haifeng Hu. ICA-based neighborhood preserving analysis for face recognition. Computer Vision and Image Understanding: CVIU, 112(3): 286–295, December 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:2011:MIN

[Hu11] Haifeng Hu. Multiscale illumination normalization for face recognition using dualtree complex wavelet transform in logarithm domain. Computer Vision and Image Understanding: CVIU, 115(10):1384–1394, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421100141X.

Huang:1980:MMG

[Hua80] Thomas S. Huang. Mathematical models of graphics. Computer Graphics and Image Processing, 12(2):127– 135, February 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hubert:2012:CSB

[Hub12] Evelyne Hubert. Convolution surfaces based on polygons for infinite and compact support kernels. Graphical Models, 74(1):1–13, January 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000257.

Herda:2005:HIS

[HUF05] L. Herda, R. Urtasun, and P. Fua. Hierarchical implicit surface joint limits for human body tracking. Computer Vision and Image Understanding: CVIU, 99(2): 189–209, August 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hummel:1977:IEH

[Hum77]  Robert Hummel.  Image

enhancement by histogram transformation. Computer Graphics and Image Processing, 6(2):184–185, April 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Hummel:1979:FDU

[Hum79] Robert A. Hummel. Feature detection using basis functions. Computer Graphics and Image Processing, 9(1): 40–55, January 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Herbin:1989:ARD

[HVD^{+}89] M. Herbin, A. Venot, J. Y.

Devaux, E. Walter, J. F. Lebruchec, L. Dubertret, and J. C. Roucayrol. Automated registration of dissimilar images: Application to medical imagery. Computer Vision, Graphics, and Image Processing, 47(1):77–88, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Horn:1979:DLM

[HW79] Berthold K. P. Horn and Robert J. Woodham. Destriping Landsat MSS images by histogram modification. Computer Graphics and Image Processing, 10 (1):69–83, May 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Haralick:1981:FMI

[HW81] Robert M. Haralick and Layne Watson. A facet model for image data. Computer Graphics and Image Processing, 15(2): 113–129, February 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Herman:1983:TPS

[HW83] Gabor T. Herman and Dallas Webster. Topological proof of a surface tracking algorithm. Computer Vision, Graphics, and Image Processing, 23 (2):162–177, August 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Huang:1994:QMM

[HW94] Wenqi Q. Huang and Gang-Qiang Q. Wang. A quasimechanical method for solving the rectangle covering problem — an approach to tackling NP hard problems. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(3):267–271, May 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1023/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1023/production/
pdf.

Hua:2006:SMF

[HW06] Gang Hua and Ying Wu.

Sequential mean field variational analysis of structured deformable shapes. Computer Vision and Image Understanding: CVIU, 101(2): 87–99, February 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hua:2007:DPA

[HW07] Gang Hua and Ying Wu.

A decentralized probabilistic approach to articulated body tracking. Computer Vision and Image Understanding: CVIU, 108(3):272– 283, December 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Hu:1996:GSB

[HWJ96] Shi-Min Hu, Guo-Zhao Wang, and Tong-Guang Jin. Generalized subdivision of Bézier surfaces. Graphical Models and Image Processing: GMIP, 58(3):218–222, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0018/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0018/production/
                                        pdf.

Huang:2006:NCI

[HWW06] Kai-Qi Huang, Qiao Wang, and Zhen-Yang Wu. Natural color image enhancement and evaluation algorithm based on human visual system. Computer Vision and Image Understanding: CVIU, 103(1): 52–63, July 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

He:2009:HFB

[HXS09] Ying He, Xian Xiao, and Hock-Soon Seah. Harmonic 1-form based skeleton extraction from examples. Graphical Models, 71(2):49–62, March 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000349.

Her:1994:RPG

[HY94] Innchyn Her and Chi-Tseng T.

Yuan. Resampling on a pseudohexagonal grid. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(4):336–347, July 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1030/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1030/production/
pdf.

Hu:1998:MBS

[HY98] Jianming Hu and Hong Yan.

A model-based segmentation method for handwritten numeral strings. Computer Vision and Image Understanding: CVIU, 70(3):383–403, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0689/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0689/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0689/production/ref.

Ho:2011:ARP

[HY11] Jeffrey Ho and Ming-Hsuan Yang. On affine registration of planar point sets using complex numbers. Computer Vision and Image Understanding: CVIU, 115(1): 50–58, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Han:2011:CFE

[HYJ11] Zhenjun Han, Qixiang Ye, and Jianbin Jiao. Combined feature evaluation for adaptive visual object tracking. Computer Vision and Image Understanding: CVIU, 115 (1):69–80, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Haralick:1986:NRB

[HZ86] Robert M. Haralick and Xinhua Zhuang. A note on “Rigid Body Motion from Depth and Optical Flow”. Computer Vision, Graphics, and Image Processing, 34(3):372–387, June 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See [BK83].

Huang:2011:CSA

[HZLM11] Junzhou Huang, Shaoting Zhang, Hongsheng Li, and Dimitris Metaxas. Composite splitting algorithms for convex optimization. Computer Vision and Image Understanding: CVIU, 115(12): 1610–1622, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001585.

Humenberger:2010:FSM

[HZW^{+}10] Martin Humenberger, Christian Zinner, Michael Weber, Wilfried Kubinger, and Markus Vincze. A fast stereo matching algorithm suitable for embedded real-time systems. Computer Vision and Image Understanding: CVIU, 114(11):1180–1202, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Isenburg:2003:CHV

[IA03] Martin Isenburg and Pierre Alliez. Compressing hexahedral volume meshes. Graphical Models, 65(4):239–257, July 2003. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ion:2011:MAS

[IAP^{+}11] Adrian Ion, Nicole M. Artner, Gabriel Peyré, Walter G. Kropatsch, and Laurent D. Cohen. Matching 2D and 3D articulated shapes using the eccentricity transform. Computer Vision and Image Understanding: CVIU, 115(6): 817–834, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Intille:2001:RPM

[IB01] Stephen S. Intille and Aaron F. Bobick. Recognizing planned, multiperson action. Computer Vision and Image Understanding: CVIU, 81 (3):414–445, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0896;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0896/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0896/ref.

Imiya:1999:ECD

[IE99] Atsushi Imiya and Ulrich Eckhardt. The Euler characteristics of discrete objects and discrete quasi-objects. Computer Vision and Image Understanding: CVIU, 75(3):307–318, September 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0791/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0791/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0791/production/ref.

Ikeuchi:1995:RPC

[IF95] Katsushi Ikeuchi and Patrick J. Flynn. Recent progress in CAD-based vision. Computer Vision and Image Understanding: CVIU, 61(3): 293–294, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1995.1023/production;
 http://www.idealibrary.
com/links/artid/cviu.1995.
1023/production/pdf.

Imiya:1999:MAR

[IF99] Atsushi Imiya and Iris Fermin. Motion analysis by random sampling and voting process. Computer Vision and Image Understanding: CVIU, 73(3):309–328, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0734/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0734/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0734/production/ref.

Ikeuchi:1991:DLS

[IH91] K. Ikeuchi and Ki Sang Hong. Determining linear shape change: toward automatic generation of object recognition programs. Computer Vision, Graphics, and Image Processing. Image Understanding, 53(2): 154–170, March 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Ishimura:1990:AFS

[IHTA90] Nobumichi Ishimura, Takeshi Hashimoto, Shuichi Tsujimoto, and Suguru Arimoto. Automatic fixing of ship position by simulation-and-matching. Computer Vision, Graphics, and Image Processing, 51(1):38–53, July 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Imai:1986:CGM

[II86] Hiroshi Imai and Masao Iri. Computational-geometric methods for polygonal approximations of a curve. Computer Vision, Graphics, and Image Processing, 36(1):31– 41, October 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kanatani:1985:DMP

[iIK85] Ken ichi I. Kanatani. Detecting the motion of a planar surface by line and surface integrals. Computer Vision, Graphics, and Image Processing, 29(1):13–22, January 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Iizuka:1987:QES

[Iiz87]  Masayuki Iizuka.  Quan-

titative evaluation of similar images with quasi-gray levels. Computer Vision, Graphics, and Image Processing, 38(3):342–360, June 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Israel-Jost:2013:CSS

[IJDAB13] Vincent Israel-Jost, Jérôme Darbon, Elsa D. Angelini, and Isabelle Bloch. Conciliating syntactic and semantic constraints for multi-phase and multi-channel region segmentation. Computer Vision and Image Understanding: CVIU, 117(8):819–826, August 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000556.

Kanatani:1984:EIM

[iK84] Ken ichi Kanatani. Errors of the incremental method for curves. Computer Vision, Graphics, and Image Processing, 26(1):130– 133, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kanatani:1985:TPS

[iK85] Ken ichi Kanatani. Tracing planar surface motion from a projection without knowing the correspondence. Computer Vision, Graphics, and Image Processing, 29(1):1–12, January 1985.

CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kanatani:1986:SMO

[iK86] Ken ichi Kanatani. Structure and motion from optical flow under orthographic projection. Computer Vision, Graphics, and Image Processing, 35(2):181– 199, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kanatani:1987:CRI

[iK87a] Ken ichi Kanatani. Camera rotation invariance of image characteristics. Computer Vision, Graphics, and Image Processing, 39(3):328– 354, September 1987. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kanatani:1987:SMO

[iK87b] Ken ichi Kanatani. Structure and motion from optical flow under perspective projection. Computer Vision, Graphics, and Image Processing, 38(2):122–146, May 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Illingworth:1988:SHT

[IK88] J. Illingworth and J. Kittler. A survey of the Hough transform. Computer Vision, Graphics, and Image Processing, 44(1):87–116, October 1988. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ikeuchi:1989:MST

[IK89] Katsushi Ikeuchi and Takeo Kanade. Modeling sensors: toward automatic generation of object recognition program. Computer Vision, Graphics, and Image Processing, 48(1):50–79, October 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ibrahim:1986:AMP

[IKS86] Hussein A. H. Ibrahim, John R. Kender, and David Elliot Shaw. On the application of massively parallel SIMD tree machines to certain intermediate-level vision tasks. Computer Vision, Graphics, and Image Processing, 36(1):53–75, October 1986. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Isler:2005:TTD

[IKST05] Volkan Isler, Sanjeev Khanna, John Spletzer, and Camillo J. Taylor. Target tracking with distributed sensors: The focus of attention problem. Computer Vision and Image Understanding: CVIU, 100(1–2): 225–247, October/November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Idrissi:2004:OIB

[ILRB04] Khalid Idrissi, Guillaume Lavoué, Julien Ricard, and Atilla Baskurt. Object of interest-based visual navigation, retrieval, and semantic content identification system. Computer Vision and Image Understanding: CVIU, 94(1– 3):271–294, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ihrke:2006:AGO

[IM06] Ivo Ihrke and Marcus Magnor. Adaptive grid optical tomography. Graphical Models, 68(5–6):484–495, September/November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070306000610.

Imme:1991:NPE

[Imm91] M. Imme. A noise peak elimination filter. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2): 204–211, March 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Immerkaer:1996:FNV

[Imm96] John Immerkær. Fast noise variance estimation. Computer Vision and Image Understanding: CVIU, 64(2): 300–302, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0060/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0060/production/
pdf.

Iben:2009:GSC

[IO09] Hayley N. Iben and James F.

O’Brien. Generating surface crack patterns. Graphical Models, 71(6):198–208, November 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000058.

Ohta:1980:CIR

[iOKS80] Yu ichi Ohta, Takeo Kanade, and Toshiyuki Sakai. Color information for region segmentation. Computer Graphics and Image Processing, 13(3): 222–241, July 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Irani:1991:IRI

[IP91] M. Irani and S. Peleg. Improving resolution by image registration. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(3): 231–239, May 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Ivins:1998:CAR

[IP98] Jim Ivins and John Porrill. Constrained active region models for fast tracking in color image sequences. Computer Vision and Image Understanding: CVIU, 72(1):54–71, October 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0653/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0653/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0653/production/ref.

Isenburg:2002:CPM

[IS02] Martin Isenburg and Jack Snoeyink. Compressing the property mapping of polygon meshes. Graphical Models, 64 (2):114–127, March 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ishikawa:1984:RFP

[Ish84] Seiji Ishikawa. Reconstructing faces on a polyhedron from apparent gradients of edges.

Computer Vision, Graphics, and Image Processing, 28 (3):289–302, December 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Toriwaki:1978:ESI

[iTF78] Jun ichiro Toriwaki and Teruo Fukumura. Extraction of structural information from grey pictures. Computer Graphics and Image Processing, 7(1):30–51, February 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Irving:2006:THI

[ITF06] G. Irving, J. Teran, and R. Fedkiw. Tetrahedral and hexahedral invertible finite elements. Graphical Models, 68(2):66–89, March 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000317.

Inoue:1984:CPR

[ITN84] Katsushi Inoue, Itsuo Takanami, and Akira Nakamura. Connected pictures are not recognizable by deterministic twodimensional on-line tessellation acceptors. Computer Vision, Graphics, and Image Processing, 26(1):126– 129, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Iosifidis:2012:MVH

[ITNP12] Alexandros Iosifidis, Anastasios Tefas, Nikolaos Nikolaidis, and Ioannis Pitas. Multi-view human movement recognition based on fuzzy distances and linear discriminant analysis. Computer Vision and Image Understanding: CVIU, 116 (3):347–360, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002074.

Toriwaki:1982:GDT

[iTTF82] Jun ichiro Toriwaki, Masahiko Tanaka, and Teruo Fukumura. Generalized distance transformation of a line pattern with gray values and its applications. Computer Graphics and Image Processing, 20(4):319– 346, December 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ip:1997:DPP

[IW97] Horace H. S. Ip and W. H. Wong. Detecting perceptually parallel curves: Criteria and force-driven optimization. Computer Vision and Image Understanding: CVIU, 68(2):190–208, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0552/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0552/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0552/production/ref.

Irschara:2012:LSD

[IZKB12] Arnold Irschara, Christopher Zach, Manfred Klopschitz, and Horst Bischof. Largescale, dense city reconstruction from user-contributed photos. Computer Vision and Image Understanding: CVIU, 116(1):2–15, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001810.

Jacobs:2001:LFM

[Jac01]  David W. Jacobs.  Lin-

ear fitting with missing data for structure-from-motion. Computer Vision and Image Understanding: CVIU, 82(1):57–81, April 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0906;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0906/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0906/ref.

Jain:1994:EV

[Jai94]  Ramesh Jain.  Expansive

vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):86–88, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1033/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1994.1033/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1038/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1038/production/
                                        pdf.

James:2009:SCA

[Jam09] Doug James. Symposium on Computer Animation 2008. Graphical Models, 71 (4):125, July 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070309000186.

Jarvis:1977:LDE

[Jar77] John F. Jarvis. Line drawing editor: Schematic diagram editing using pattern recognition techniques. Computer Graphics and Image Processing, 6(5):452– 484, October 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Jou:1989:IIA

[JB89] Jinn-Yeu Y. Jou and Alan C.

Bovik. Improved initial approximation and intensityguided discontinuity detection in visible-surface reconstruction. Computer Vision, Graphics, and Image Processing, 47(3):292–326, September 1989. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Jordan:1991:UCI

[JB91] John R. Jordan, III and Alan C. Bovik. Using chromatic information in edgebased stereo correspondence. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(1):98–118, July 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Jiang:1992:SEA

[JB92] X. Y. Jiang and H. Bunke. A simple and efficient algorithm for determining the symmetries of polyhedra. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1): 91–95, January 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Jiang:1999:EDR

[JB99] Xiaoyi Jiang and Horst Bunke. Edge detection in range images based on scan line approximation. Computer Vision and Image Understanding: CVIU, 73 (2):183–199, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0715/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0715/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0715/production/ref.

Jia:2008:VBD

[JBC08] Zhen Jia, Arjuna Balasuriya, and Subhash Challa. Vision based data fusion for autonomous vehicles target tracking using interacting multiple dynamic models. Computer Vision and Image Understanding: CVIU, 109 (1):1–21, January 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jia:2004:QDC

[JBK04] Jinyuan Jia, George Baciu, and Ki-Wan Kwok. Quadric decomposition for computing the intersections of surfaces of revolution. Graphical Models, 66(5):303–330, September 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Jeyakar:2008:ROT

[JBR08] Jaideep Jeyakar, R. Venkatesh Babu, and K. R. Ramakrishnan. Robust object tracking with background-weighted local kernels. Computer Vision and Image Understanding: CVIU, 112(3):296– 309, December 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jain:1991:IMN

[JBS^{+}91] R. C. Jain, T. O. Binford, M. A. Snyder, Y. Aloimonos, A. Rosenfeld, T. S. Huang, K. W. Bowyer, and J. P. Jones. Ignorance, myopia, and naivete in computer vision systems. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (1):112–128, January 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Jiang:2011:GBM

[JBWK11] Xiaoyi Jiang, Klaus Broelemann, Steffen Wachenfeld, and Antonio Krüger. Graphbased markerless registration of city maps using geometric hashing. Computer Vision and Image Understanding: CVIU, 115(7):1032–1043, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000828.

Jacobus:1981:ILV

[JC81] C. Jacobus and R. T. Chien. Intermediate-level vision — building vertex-string-surface (V-S-S) graphs. Computer Graphics and Image Processing, 15(4):339–363, April 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Jau:1990:STU

[JC90] J. Y. Jau and R. T. Chin. Shape from texture using the Wigner distribution. Computer Vision, Graphics, and Image Processing, 52(2):248– 263, November 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Jungert:1993:IAP

[JC93] Erland Jungert and S. K. Chang. An image algebra for pictorial data manipulation. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(2):147–160, September 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1993.
                                        1035/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1993.1035/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1993.
                                        1037/production;   http:
                                        //www.idealibrary.com/links/

artid/cviu.1993.1037/production/
pdf.

Jain:1994:ABL

[JC94] Anil K. Jain and Yao Chen.

Address block location using color and texture analysis. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):179–190, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1046/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1046/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1051/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1051/production/
pdf.

Jang:1998:MSS

[JC98] Ben K. Jang and Roland T.

Chin. Morphological scale space for 2D shape smoothing. Computer Vision and Image Understanding: CVIU, 70(2):121–141, May 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0626/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0626/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/cviu.1997.
                                        0626/production/ref.

Joshi:2006:SES

[JC06] Manjunath V. Joshi and Subhasis Chaudhuri. Simultaneous estimation of superresolved depth map and intensity field using photometric cue. Computer Vision and Image Understanding: CVIU, 101(1):31–44, January 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jacot-Descombes:1997:APG

[JDP97] Alain Jacot-Descombes and Thierry Pun. Asynchronous perceptual grouping: From contours to relevant 2-D structures. Computer Vision and Image Understanding: CVIU, 66(1):1–24, April 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0509/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0509/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0509/production/ref.

Jean:2011:OSP

[Jea11] Yves D. Jean. Optical signal processing with illuminationencoded filters. Computer Vision and Image Understanding: CVIU, 115(5):561–575, May 2011. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Jamieson:2012:DHO

[JEF^{+}12] Michael Jamieson, Yulia Eskin, Afsaneh Fazly, Suzanne Stevenson, and Sven J. Dickinson. Discovering hierarchical object models from captioned images. Computer Vision and Image Understanding: CVIU, 116(7):842–853, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000495.

Jordan:1998:PCB

[JEK98] Corinne Le Buhan Jordan, Touradj Ebrahimi, and Murat Kunt. Progressive content-based shape compression for retrieval of binary images. Computer Vision and Image Understanding: CVIU, 71(2):198–212, August 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0707/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0707/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0707/production/ref.

Junejo:2010:GCE

[JF10] Imran N. Junejo and Hassan Foroosh. GPS coordinates estimation and camera calibration from solar shadows. Computer Vision and Image Understanding: CVIU, 114(9):991–1003, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jimenez:2011:TTP

[JFS11] Juan José Jiménez, Francisco Ramón Feito, and Rafael Jesús Segura. Tetratrees properties in graphic interaction. Graphical Models, 73(5):182–201, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070311000105.

Jaman:1985:DAI

[JGR85] Kirby A. Jaman, Richard Gordon, and Rangaraj M. Rangayyan. Display of 3D anisotropic images from limited-view computed tomograms. Computer Vision, Graphics, and Image Processing, 30(3):345–361, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Johnson:1998:CPM

[JH98] Andrew E. Johnson and Martial Hebert. Control of polygonal mesh resolution for 3-D computer vision. Graphical Models and Image Processing: GMIP, 60(4):261–285, July 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0474/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0474/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0474/production/ref.

Jayaramamurthy:1983:AST

[JJ83] S. N. Jayaramamurthy and Ramesh Jain. Approach to the segmentation of textured dynamic scenes. Computer Vision, Graphics, and Image Processing, 21(2):239– 261, February 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Jenkin:1994:RLS

[JJ94] Michael R. M. Jenkin and Allan D. Jepson. Recovering local surface structure through local phase difference measurements. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):72–93, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1005/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1005/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1005/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1994.1005/production/
                                        pdf.

Jie:1995:ICE

[JJM95] Zhou Jie, Peng Jiaxiong, and Ding Mingyue. Improved codebook edge detection. Graphical Models and Image Processing: GMIP, 57 (6):533–536, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1044/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1044/production/
                                        pdf.

Jarvis:1976:STD

[JJN76] J. F. Jarvis, C. N. Judice, and W. H. Ninke. A survey of techniques for the display of continuous tone pictures on bilevel displays. Computer Graphics and Image Processing, 5(1):13–40, March 1976. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Jenkin:1991:TDM

[JJT91] M. R. M. Jenkin, A. D. Jepson, and J. K. Tsotsos. Techniques for disparity measurement. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (1):14–30, January 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Jeong:2002:DRD

[JK02] Won-Ki Jeong and Chang-Hun Kim. Direct reconstruction of a displaced subdivision surface from unorganized points. Graphical Models, 64 (2):78–93, March 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Jain:2007:SMO

[JKM07] Vishal Jain, Benjamin B.

Kimia, and Joseph L. Mundy. Segregation of moving objects using elastic matching. Computer Vision and Image Understanding: CVIU, 108(3): 230–242, December 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jiang:2012:CCM

[JLD12] Zhuolin Jiang, Zhe Lin, and Larry S. Davis. Class consistent k-means: Application to face and action recognition. Computer Vision and Image Understanding: CVIU, 116(6):730–741, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000367.

Jiang:2013:UTB

[JLD13] Zhuolin Jiang, Zhe Lin, and Larry S. Davis. A unified tree-based framework for joint action localization, recognition and segmentation. Computer Vision and Image Understanding: CVIU, 117(10): 1345–1355, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001749.

Jung:2013:WAP

[JLL13] Ho Yub Jung, Kyoung Mu Lee, and Sang Uk Lee. Window annealing for pixellabeling problems. Computer Vision and Image Understanding: CVIU, 117 (3):289–303, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212002056.

Jaroslavski:1979:IDU

[JM79] L. P. Jaroslavski and N. S. Merzlyakov. Information display using the methods of digital holography. Computer Graphics and Image Processing, 10(1):1–29, May 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Jolion:1992:APF

[JM92] J. M. Jolion and A. Montanvert. The adaptive pyramid: a framework for 2D image analysis. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (3):339–349 (or 339–348??), May 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Jodoin:2009:OFB

[JM09a] Pierre-Marc Jodoin and Max Mignotte. Optical-flow based on an edge-avoidance procedure. Computer Vision and Image Understanding: CVIU, 113(4):511–531, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Just:2009:CST

[JM09b] Agnès Just and Sébastien Marcel. A comparative study of two state-of-the-art sequence processing techniques for hand gesture recognition. Computer Vision and Image Understanding: CVIU, 113 (4):532–543, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jain:1979:STD

[JMA79] Ramesh Jain, W. N. Martin, and J. K. Aggarwal. Segmentation through the detection of changes due to motion. Computer Graphics and Image Processing, 11(1):13– 34, September 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Julia:2011:SBI

[JMPG11] Carme Julià, Rodrigo Moreno, Domenec Puig, and Miguel Angel Garcia. Shape-based image segmentation through photometric stereo. Computer Vision and Image Understanding: CVIU, 115(1): 91–104, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jalihal:1993:SDT

[JN93] Devendra Jalihal and Loren W. Nolte. Signal detection theory approach to the multiple parallel moving targets problem. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(4):235–254, July 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1017/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1017/production/
pdf.

Jiang:2009:VWP

[JN09] Yu-Gang Jiang and Chong-Wah Ngo. Visual word proximity and linguistics for semantic video indexing and near-duplicate retrieval. Computer Vision and Image Understanding: CVIU, 113(3): 405–414, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jain:2011:GQ

[JO11] Brijnesh J. Jain and Klaus Obermayer. Graph quantization. Computer Vision and Image Understanding: CVIU, 115(7):946–961, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000920.

Jokinen:1998:ABM

[Jok98]  Olli Jokinen.  Area-based

matching for simultaneous registration of multiple 3D profile maps. Computer Vision and Image Understanding: CVIU, 71(3): 431–447, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0639/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0639/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0639/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0713/production;
http://www.idealibrary.
com/links/artid/cviu.1998.
0713/production/pdf.

Jolion:1994:CVM

[Jol94] Jean-Michel M. Jolion. Computer vision methodologies. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):53–71, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1004/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1004/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1004/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1004/production/
pdf.

Jones:1997:COH

[Jon97]  Graeme A. Jones.  Con-

straint, optimization, and hierarchy: Reviewing stereoscopic correspondence of complex features. Computer Vision and Image Understanding: CVIU, 65(1):57–78, January 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0482/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0482/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0482/production/ref.

Jones:1999:CFS

[Jon99] Ronald Jones. Connected filtering and segmentation using component trees. Computer Vision and Image Understanding: CVIU, 75(3): 215–228, September 1999.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0777/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0777/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0777/production/ref.

Jones:2008:SII

[Jon08]  Graeme A. Jones.  Spe-

cial issue on Intelligent Visual Surveillance. Computer Vision and Image Understanding: CVIU, 111(1):1, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Joseph:1994:ULS

[Jos94]  S. H. Joseph.  Unbiased

least squares fitting of circular arcs. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(5):424–432, September 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1039/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1039/production/
pdf.

Joseph:1999:OPE

[Jos99] S. H. Joseph. Optimal pose estimation in two and three dimensions. Computer Vision and Image Understanding: CVIU, 73(2):215–231, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0733/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0733/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0733/production/ref.

Jost:2005:ACC

[JOvW^{+}05] Timothée Jost, Nabil Ouerhani, Roman von Wartburg, René Müri, and Heinz Hügli. Assessing the contribution of color in visual attention. Computer Vision and Image Understanding: CVIU, 100(1–2):107–123, October/November 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jagoe:1978:GCD

[JP78] Roger Jagoe and Keith Paton. Generalized counting in digital pictures. Computer Graphics and Image Processing, 7(1):52–66, February 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Janssens:1986:NUN

[JR86] D. Janssens and G. Rozenberg. Neighborhood-uniform Nlc grammars. Computer Vision, Graphics, and Image Processing, 35(2):131– 151, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Jang:2009:OSS

[JR09] Justin Jang and Jarek Rossignac.

OCTOR: Subset selection in recursive pattern hierarchies. Graphical Models, 71(2):92–106, March 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000350.

Jaynes:2003:RRB

[JRH03] Christopher Jaynes, Edward Riseman, and Allen Hanson. Recognition and reconstruction of buildings from multiple aerial images. Computer Vision and Image Understanding: CVIU, 90(1):68–98, April 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Janssens:1982:SPN

[JRV82] D. Janssens, G. Rozenberg, and R. Verraedt. On sequential and parallel noderewriting graph grammars. Computer Graphics and Image Processing, 18(3):279– 304, March 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Janssens:1983:SPN

[JRV83] D. Janssens, G. Rozenberg, and R. Verraedt. On sequential and parallel noderewriting graph grammars, II. Computer Vision, Graphics, and Image Processing, 23 (3):295–312, September 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Joe:1987:CLV

[JS87] B. Joe and R. B. Simpson. Corrections to Lee’s visibility polygon algorithm. BIT, 27 (4):458–473, 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See [Lee83a].

Jaimes:2007:MHC

[JS07] Alejandro Jaimes and Nicu Sebe. Multimodal human– computer interaction: a survey. Computer Vision and Image Understanding: CVIU, 108(1–2):116–134, October/ November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Javed:2008:MIC

[JSRS08] Omar Javed, Khurram Shafique, Zeeshan Rasheed, and Mubarak Shah. Modeling inter-camera space–time and appearance relationships for tracking across non-overlapping views. Computer Vision and Image Understanding: CVIU, 109 (2):146–162, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jackins:1980:OTT

[JT80] Chris L. Jackins and Steven L.

Tanimoto. Oct-trees and their use in representing threedimensional objects. Computer Graphics and Image Processing, 14(3):249– 270, November 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Jenkin:1986:ATC

[JT86] Michael Jenkin and John K.

Tsotsos. Applying temporal constraints to the dynamic stereo problem. Computer Vision, Graphics, and Image Processing, 33(1):16– 32, January 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Jones:1991:MSA

[JTEA91] J. G. Jones, R. W. Thomas, P. G. Earwicker, and S. Addison. Multiresolution statistical analysis of computergenerated fractal imagery. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):349–363, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Jurie:1999:SSP

[Jur99]  Frederic Jurie.  Solution

of the simultaneous pose and correspondence problem using Gaussian error model. Computer Vision and Image Understanding: CVIU, 73(3):357–373, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0735/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0735/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0735/production/ref.

Janssen:1997:AVL

[JV97] Rik D. T. Janssen and Albert M. Vossepoel. Adaptive vectorization of line drawing images. Computer Vision and Image Understanding: CVIU, 65(1):38–56, January 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0484/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0484/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0484/production/ref.

Jonk:1999:GID

[JvdBS99] Arnold Jonk, Rein van den Boomgaard, and Arnold Smeulders. Grammatical inference of dashed lines. Computer Vision and Image Understanding: CVIU, 74(3):212–226, June 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0753/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0753/production/
pdf.

Jacobson:1987:DOF

[JW87] Lowell Jacobson and Harry Wechsler. Derivation of optical flow using a spatiotemporalfrequency approach. Computer Vision, Graphics, and Image Processing, 38(1):29– 65, April 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Jiang:1994:SSC

[JW94] Caixai X. Jiang and Matthew O.

Ward. Shadow segmentation and classification in a constrained environment. Computer Vision, Graphics, and Image Processing. Image Understanding, 59 (2):213–225, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1014/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1014/production/

                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1016/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1016/production/
                                        pdf.

Ji:2005:SIE

[JWDF05] Qiang Ji, Harry Wechsler, Andrew Duchowski, and Myron Flickner. Special issue: eye detection and tracking. Computer Vision and Image Understanding: CVIU, 98(1):1–3, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jeong:2004:IRU

[JWG04] Sangoh Jeong, Chee Sun Won, and Robert M. Gray. Image retrieval using color histograms generated by Gauss mixture vector quantization. Computer Vision and Image Understanding: CVIU, 94(1– 3):44–66, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Jin:2012:UUO

[JWL12] Yong Jin, Qingbiao Wu, and Ligang Liu. Unsupervised upright orientation of manmade models. Graphical Models, 74(4):99–108, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000148.

Jiang:2013:CSE

[JXC^{+}13] Wei Jiang, Kai Xu, Zhi-Quan Cheng, Ralph R. Martin, and Gang Dang. Curve skeleton extraction by coupled graph contraction and surface clustering. Graphical Models, 75(3):137–148, May 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000732.

Jiang:2013:SBI

[JXCZ13] Wei Jiang, Kai Xu, Zhi-Quan Cheng, and Hao Zhang. Skeleton-based intrinsic symmetry detection on point clouds. Graphical Models, 75(4):177–188, July 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070313000118.

Jiang:2011:AVE

[JYTK11] Fan Jiang, Junsong Yuan, Sotirios A. Tsaftaris, and Aggelos K. Katsaggelos. Anomalous video event detection using spatiotemporal context. Computer Vision and Image Understanding: CVIU, 115(3):323–333, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kammoun:1994:OED

[KA94] F. Kammoun and J. P. Astruc. Optimum edge detection for object-background picture. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):25–28, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1004/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1004/production/
pdf.

Keller:2008:GPI

[KA08] Y. Keller and A. Averbuch. Global parametric image alignment via highorder approximation. Computer Vision and Image Understanding: CVIU, 109(3): 244–259, March 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Krishnamoorthi:2012:SBE

[KA12] R. Krishnamoorthi and G. Annapoorani. A simple boundary extraction technique for irregular pupil localization with orthogonal polynomials. Computer Vision and Image Understanding: CVIU, 116 (2):262–273, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002013.

Knerr:1998:HMM

[KABP98] S. Knerr, E. Augustin, O. Baret, and D. Price. Hidden Markov model based word recognition and its application to legal amount reading on French checks. Computer Vision and Image Understanding: CVIU, 70(3):404–419, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0685/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0685/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0685/production/ref.

Khodakovsky:2002:NOC

[KADS02] Andrei Khodakovsky, Pierre Alliez, Mathieu Desbrun, and Peter Schröder. Near-optimal connectivity encoding of 2manifold polygon meshes. Graphical Models, 64(3–4): 147–168, May 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Kanth:1999:DRS

[KAES99] K. V. Ravi Kanth, Divyakant Agrawal, Amr El Abbadi, and Ambuj Singh. Dimensionality reduction for similarity searching in dynamic databases. Computer Vision and Image Understanding: CVIU, 75(1– 2):59–72, July/August 1999.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0762/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0762/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0762/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0763/production;
 http://www.idealibrary.
com/links/artid/cviu.1999.
0763/production/pdf;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0763/production/
ref.

Kak:1995:E

[Kak95] Avi Kak. Editorial. Computer Vision and Image Understanding: CVIU, 61(2): 153, March 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1995.1012/production;
 http://www.idealibrary.
com/links/artid/cviu.1995.
1012/production/pdf.

Kak:1997:THF

[Kak97] Avi Kak. Tribute: Herb Freeman. Computer Vision and Image Understanding: CVIU, 68(3):255–??, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.

0660/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0660/production/
pdf.

Kalay:1982:DSC

[Kal82] Yehuda E. Kalay. Determining the spatial containment of a point in general polyhedra. Computer Graphics and Image Processing, 19(4): 303–334, August 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kanade:1980:RSS

[Kan80a] Takeo Kanade. Region segmentation: Signal vs semantics. Computer Graphics and Image Processing, 13(4):279– 297, August 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kanal:1980:MMM

[Kan80b] Laveen N. Kanal. Markov mesh models. Computer Graphics and Image Processing, 12(4):371–375, April 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). KANAL80.

Kanatani:1988:CLA

[Kan88] Ken-Ichi Kanatani. Constraints on length and angle. Computer Vision, Graphics, and Image Processing, 41(1):28–42, January 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kanatani:1991:CPG

[Kan91a]  Kenichi Kanatani.  Com-

putational projective geometry. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(3): 333–348, November 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Kanatani:1991:HTG

[Kan91b] Kenichi Kanatani. Hypothesizing and testing geometric properties of image data. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(3): 349–357, November 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Kanatani:1994:CCR

[Kan94a] Kenichi Kanatani. Computational cross ratio for computer vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3):371–381, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1063/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1063/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.

1069/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1069/production/
pdf.

Kanatani:1994:SAG

[Kan94b]  Kenichi Kanatani.  Sta-

tistical analysis of geometric computation. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):286–306, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1020/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1020/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1024/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1024/production/
pdf.

Kanatani:1994:SFH

[Kan94c]  Kenichi Kanatani.  Sta-

tistical foundation for hypothesis testing of image data. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3):382–391, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1064/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1064/production/

                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1070/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1070/production/
                                        pdf.

Kanatani:1998:CRL

[Kan98] Kenichi Kanatani. Cramer-Rao lower bounds for curve fitting. Graphical Models and Image Processing: GMIP, 60(2):93–99, March 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0466/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1998.0466/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0466/production/ref.

Karasick:1989:SOP

[Kar89] M. Karasick. The sameobject problem for polyhedral solids. Computer Vision, Graphics, and Image Processing, 46(1):22–36, April 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kashyap:1980:UMR

[Kas80] R. L. Kashyap. Univariate and multivariate random field models for images. Computer Graphics and Image Processing, 12(3): 257–270, March 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kasif:1994:OPA

[Kas94] Simon Kasif. Optimal parallel algorithms for quadtree problems. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):281–285, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1019/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1019/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1023/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1023/production/
pdf.

Kawai:1978:GTF

[Kaw78] Satoru Kawai. A graph theoretic formulation of bit pattern algorithms for graphics. Computer Graphics and Image Processing, 7(1):84– 104, February 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kawai:1979:BCC

[Kaw79] Satoru Kawai. Boundary curve criterion. Computer Graphics and Image Processing, 11(3):281–289, November 1979. CODEN CGIPBG.

ISSN 0146-664X (print), 1557-9697 (electronic).

Kawai:1982:TPP

[Kaw82]  Satoru Kawai.  On the

topology preservation property of local parallel operations. Computer Graphics and Image Processing, 19(3): 265–280, July 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kawai:1983:TQP

[Kaw83]  Satoru Kawai.  Topology

quasi-preservation by local parallel operations. Computer Vision, Graphics, and Image Processing, 23(3):353– 365, September 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kim:1991:DSS

[KB91a] Byungil Kim and Peter Burger. Depth and shape from shading using the photometric stereo method. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(3):416–427, November 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Kiryati:1991:GLC

[KB91b] N. Kiryati and A. Bruckstein. Gray levels can improve the performance of binary image digitizers. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1): 31–39, January 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Kiryati:1991:AHT

[KB91c] N. Kiryati and A. M. Bruckstein. Antialiasing the Hough transform. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(3): 213–222, May 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Kimmel:1995:GSS

[KB95a] Ron Kimmel and Alfred M.

Bruckstein. Global shape from shading. Computer Vision and Image Understanding: CVIU, 62(3): 360–369, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1060/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1060/production/
pdf.

Kimmel:1995:TLS

[KB95b] Ron Kimmel and Alfred M.

Bruckstein. Tracking level sets by level sets: a method for solving the shape from shading problem. Computer Vision and Image Understanding: CVIU, 62(1):47–58, July 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1040/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1040/production/
pdf.

Kanai:1998:SID

[KB98] Junichi Kanai and Henry S. Baird. Special issue on document image understanding and retrieval. Computer Vision and Image Understanding: CVIU, 70(3):285–286, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0693/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0693/production/
pdf.

Kiryati:2000:HHT

[KB00] Nahum Kiryati and Alfred M. Bruckstein. Heteroscedastic Hough transform (HtHT): An efficient method for robust line fitting in the ‘errors in the variables’ problem. Computer Vision and Image Understanding: CVIU, 78(1):69–83, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.

1999.0828;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0828/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0828/ref.

Ko:2001:IWH

[KB01] Hyeong-Seok Ko and Norman I. Badler. The International Workshop on Human Modeling and Animation in Graphical Models. Graphical Models, 63(2):65, March 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0550;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0550/
pdf.

Krausz:2012:LAV

[KB12] Barbara Krausz and Christian Bauckhage. Loveparade 2010: Automatic video analysis of a crowd disaster. Computer Vision and Image Understanding: CVIU, 116(3):307–319, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002037.

Kontschieder:2012:EHG

[KBD^{+}12] Peter Kontschieder, Samuel Rota Bulò, Michael Donoser, Marcello Pelillo, and Horst Bischof. Evolutionary Hough Games for coherent object detection. Computer Vision and Image Understanding: CVIU, 116(11):1149–1158, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001142.

Kukelova:2010:FRN

[KBJ^{+}10] Zuzana Kukelova, Martin Byröd, Klas Josephson, Tomas Pajdla, and Kalle Åström. Fast and robust numerical solutions to minimal problems for cameras with radial distortion. Computer Vision and Image Understanding: CVIU, 114(2):234– 244, February 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Karavasilis:2012:NFM

[KBN12] Vasileios Karavasilis, Konstantinos Blekas, and Christophoros Nikou. A novel framework for motion segmentation and tracking by clustering incomplete trajectories. Computer Vision and Image Understanding: CVIU, 116(11): 1135–1148, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001063.

Kato:1996:HMR

[KBZ96] Zoltan Kato, Marc Berthod, and Josiane Zerubia. A hierarchical Markov random field model and multitemperature annealing for parallel image classification. Graphical Models and Image Processing: GMIP, 58(1):18–37, January 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0002/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0002/production/
pdf.

Krishnapuram:1987:HST

[KC87] Raghuram Krishnapuram and David Casasent. Hough space transformations for discrimination and distortion estimation. Computer Vision, Graphics, and Image Processing, 38(3):299–316, June 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kundu:1992:TCU

[KC92] Amlan Kundu and Jia-Lin Chen. Texture classification using QMF bank-based subband decomposition. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(5):369–384, September 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Kay:1994:IIM

[KC94] Greg Kay and Terry Caelli. Inverting an illumination model from range and intensity maps. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):183–201, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1012/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1994.1012/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1014/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1014/production/
                                        pdf.

Kay:1995:EPI

[KC95] Greg Kay and Terry Caelli. Estimating the parameters of an illumination model using photometric stereo. Graphical Models and Image Processing: GMIP, 57 (5):365–388, September 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1032/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1032/production/
                                        pdf.

Kyatkin:1999:PMC

[KC99] Alexander B. Kyatkin and Gregory S. Chirikjian. Pattern matching as a correlation on the discrete motion group. Computer Vision and Image Understanding: CVIU, 74(1):22–35, April 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0745/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0745/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0745/production/ref.

Kopparapu:2001:ENC

[KC01] Sunil Kopparapu and Peter Corke. The effect of noise on camera calibration parameters. Graphical Models, 63 (5):277–303, September 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Kashyap:1981:DRC

[KCA81] R. L. Kashyap, R. Chellappa, and N. Ahuja. Decision rules for choice of neighbors in random field models of images. Computer Graphics and Image Processing, 15 (4):301–318, April 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Keller:1989:TDS

[KCC89] James M. Keller, Susan Chen, and Richard M. Crownover.

Texture description and segmentation through fractal geometry. Computer Vision, Graphics, and Image Processing, 45(2):150–166, February 1989. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Khachan:2000:PRA

[KCD00] Mohammed Khachan, Patrick Chenin, and Hafsa Deddi. Polyhedral representation and adjacency graph in ndimensional digital images. Computer Vision and Image Understanding: CVIU, 79(3):428–441, September 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0859;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0859/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0859/ref.

Kundu:1985:GDC

[KCM85] M. K. Kundu, B. B. Chaudhuri, and D. Dutta Majumder. A generalised digital contour coding scheme. Computer Vision, Graphics, and Image Processing, 30(3):269–278, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ko:2006:SIP

[KCOTW06] Hyeong-Seok Ko, Daniel Cohen-Or, Demetri Terzopoulos, and Joe Warren. Special issue: PG2004. Graphical Models, 68(4):323, July 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000026.

Klinger:1976:EPR

[KD76] Allen Klinger and Charles R.

Dyer. Experiments on picture representation using regular decomposition. Computer Graphics and Image Processing, 5(1):68–105, March 1976. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kulpa:1981:FDI

[KD81] Zenon Kulpa and Marek Doros. Freeman digitization of integer circles minimizes the radial error. Computer Graphics and Image Processing, 17(2):181– 184, October 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kurozumi:1982:PAM

[KD82] Yoshisuke Kurozumi and Wayne A. Davis. Polygonal approximation by the minimax method. Computer Graphics and Image Processing, 19(3):248–264, July 1982.

CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Krithivasan:1985:TWG

[KD85] Kamala Krithivasan and Anindya Das. Terminal weighted grammars and picture description. Computer Vision, Graphics, and Image Processing, 30(1):13– 31, April 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Knoll:1986:AGS

[KD86] Thomas F. Knoll and Edward J. Delp. Adaptive gray scale mapping to reduce registration noise in difference images. Computer Vision, Graphics, and Image Processing, 33(2):129– 137, February 1986. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kehtarnavaz:1988:FSR

[Kd88] N. Kehtarnavaz and R. J. P. de Figueiredo. A framework for surface reconstruction from 3D contours. Computer Vision, Graphics, and Image Processing, 42(1):32– 47, April 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Koplowitz:1996:HRC

[KD96] Jack Koplowitz and Joseph DeLeone. Hierarchical representation of chain-encoded binary image contours. Computer Vision and Image Understanding: CVIU, 63 (2):344–352, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0024/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0024/production/
pdf.

Kaess:2010:PSM

[KD10] Michael Kaess and Frank Dellaert. Probabilistic structure matching for visual SLAM with a multi-camera rig. Computer Vision and Image Understanding: CVIU, 114(2): 286–296, February 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

KameswaraRao:1978:CCP

[KDK78] C. V. Kameswara Rao, P. E.

Danielsson, and B. Kruse. Checking connectivity preservation properties of some types of picture processing operations. Computer Graphics and Image Processing, 8(2): 299–309, October 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kia:1998:SCP

[KDRC98] Omid E. Kia, David S. Doermann, Azriel Rosenfeld, and Rama Chellapa. Symbolic compression and processing of document images. Computer Vision and Image Understanding: CVIU, 70(3):335–349, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0682/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0682/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0682/production/ref.

Kosmopoulos:2012:BFB

[KDV12] Dimitrios I. Kosmopoulos, Nikolaos D. Doulamis, and Athanasios S. Voulodimos. Bayesian filter based behavior recognition in workflows allowing for user feedback. Computer Vision and Image Understanding: CVIU, 116(3):422–434, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002050.

Kwon:1999:ACF

[KdVL99] Young H. Kwon and Niels da Vitoria Lobo. Age classification from facial images. Computer Vision and Image Understanding:

CVIU, 74(1):1–21, April 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0549/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0549/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0549/production/ref.

Kender:1986:VES

[Ken86] John R. Kender. Vision expert systems demand challenging expert interactions. Computer Vision, Graphics, and Image Processing, 34(1):102–104, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kittler:1986:CSF

[KF86] J. Kittler and J. Föglein.

On compatibility and support functions in probabilistic relaxation. Computer Vision, Graphics, and Image Processing, 34(3):257–267, June 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). KITTLER86a.

Kuhl:1982:EFF

[KG82] Frank P. Kuhl and Charles R.

Giardina. Elliptic Fourier features of a closed contour. Computer Graphics and Image Processing, 18(3): 236–258, March 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kehtarnavaz:1990:ECZ

[KG90] N. Kehtarnavaz and N. Griswold. Establishing collision zones for obstacles moving with uncertainty. Computer Vision, Graphics, and Image Processing, 49(1):95– 103, January 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kambhamettu:1994:CBA

[KG94] Chandra Kambhamettu and Dmitry B. Goldgof. Curvaturebased approach to point correspondence recovery in conformal nonrigid motion. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):26–43, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1029/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1994.1029/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1034/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1034/production/
                                        pdf.

Kronrod:2001:ECN

[KG01] Boris Kronrod and Craig Gotsman. Efficient coding of nontriangular mesh connectivity. Graphical Models, 63 (4):263–275, July 2001. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Kim:2005:RAD

[KGC05] Hae Yong Kim, Javier Giacomantone, and Zang Hee Cho. Robust anisotropic diffusion to produce enhanced statistical parametric map from noisy fMRI. Computer Vision and Image Understanding: CVIU, 99(3):435– 452, September 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2010:JRC

[KGFP10] Seon Joo Kim, David Gallup, Jan-Michael Frahm, and Marc Pollefeys. Joint radiometric calibration and feature tracking system with an application to stereo. Computer Vision and Image Understanding: CVIU, 114(5):574–582, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2010:ESC

[KGK10] Jun-Sik Kim, Pierre Gurdjos, and In So Kweon. Euclidean structure from confocal conics: Theory and application to camera calibration. Computer Vision and Image Understanding: CVIU, 114(7): 803–812, July 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kucuktunc:2010:FCH

[KGU10] Onur Küçüktunç, Uğur Güdükbay, and Özgür Ulusoy. Fuzzy color histogram-based video segmentation. Computer Vision and Image Understanding: CVIU, 114(1):125– 134, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Krusemark:1983:OSI

[KH83a] Scott Krusemark and R. M. Haralick. Operating system interface for transportable image processing software. Computer Vision, Graphics, and Image Processing, 23(1):42– 66, July 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Krusemark:1983:IRF

[KH83b] Scott Krusemark and Robert M. Haralick. Image random file access routines. Computer Vision, Graphics, and Image Processing, 23(3):239– 257, September 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ku:1986:NAR

[KH86] Fu-Nian Ku and Jian-Min Hu. A new approach to the restoration of an image blurred by a linear uniform motion. Computer Vision, Graphics, and Image Processing, 34(1):20–34, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Korsten:1990:EGM

[KH90] M. J. Korsten and Z. Houkes.

The estimation of geometry and motion of a surface from image sequences by means of linearization of a parametric model. Computer Vision, Graphics, and Image Processing, 50(1):1–28, April 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kumar:1994:RME

[KH94] Rakesh Kumar and Allen R.

Hanson. Robust methods for estimating pose and a sensitivity analysis. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3): 313–342, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1060/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1060/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1066/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1066/production/
pdf.

Khoubyari:1996:FFW

[KH96] Siamak Khoubyari and Jonathan J. Hull. Font and function word identification in document recognition. Computer Vision and Image Understanding: CVIU, 63(1):66–74, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0005/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0005/production/
                                        pdf.

Kervrann:1998:HMM

[KH98] Charles Kervrann and Fabrice Heitz. A hierarchical Markov modeling approach for the segmentation and tracking of deformable shapes. Graphical Models and Image Processing: GMIP, 60(3):173–195, May 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0469/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1998.0469/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0469/production/ref.

Kruger:2013:TOA

[KH13] Volker Krüger and Dennis Herzog. Tracking in object action space. Computer Vision and Image Understanding: CVIU, 117(7):764–789, July 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000313.

Kong:2005:RAV

[KHA^{+}05] Seong G. Kong, Jingu Heo, Besma R. Abidi, Joonki Paik, and Mongi A. Abidi. Recent advances in visual and infrared face recognition—a review. Computer Vision and Image Understanding: CVIU, 97(1):103–135, January 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kita:2001:CBD

[KHB01] Yasuyo Kita, Ralph Highnam, and Michael Brady. Correspondence between different view breast X rays using curved epipolar lines. Computer Vision and Image Understanding: CVIU, 83 (1):38–56, July 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0908;
http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0908/pdf;   http://
www.idealibrary.com/links/
doi/10.1006/cviu.2001.0908/
ref.

Kim:2012:PCM

[KHH^{+}12] Kunho Kim, Mohammad K.

Hasan, Jae-Pil Heo, Yu-Wing Tai, and Sung eui Yoon. Probabilistic cost model for nearest neighbor search in image retrieval. Computer Vision and Image Understanding: CVIU, 116(9):991–998, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000768.

Kim:2010:SAM

[KHK10] Jun-Sik Kim, Myung Hwangbo, and Takeo Kanade. Spherical approximation for multiple cameras in motion estimation: Its applicability and advantages. Computer Vision and Image Understanding: CVIU, 114(10):1068–1083, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Konstantopoulos:1994:NDN

[KHS94] C. Konstantopoulos, R. Hohlfeld, and G. Sandri. Novel deconvolution of noisy Gaussian filters with a modified Hermite expansion. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(6):433–441, November 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1040/production;   http:
//www.idealibrary.com/links/

artid/cgip.1994.1040/production/
pdf.

Kittler:1985:TSB

[KIF85] J. Kittler, J. Illingworth, and J. Föglein. Threshold selection based on a simple image statistic. Computer Vision, Graphics, and Image Processing, 30(2):125–147, May 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kenmochi:1998:BED

[KII98] Yukiko Kenmochi, Atsushi Imiya, and Akira Ichikawa. Boundary extraction of discrete objects. Computer Vision and Image Understanding: CVIU, 71(3): 281–293, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0652/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0652/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0652/production/ref.

Kamata:1989:REP

[KIK89] Seiichiro Kamata, Seiji Ishikawa, and Kiyoshi Kato. Reconstructing an edge on a polyhedron using an optimization method. Computer Vision, Graphics, and Image Processing, 47(1):92–104, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kim:1982:CSL

[Kim82] Chul E. Kim. On cellular straight line segments. Computer Graphics and Image Processing, 18(4):369– 381, April 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kimmel:1997:ISS

[Kim97] Ron Kimmel. Intrinsic scale space for images on surfaces: The geodesic curvature flow. Graphical Models and Image Processing: GMIP, 59(5):365–372, September 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0442/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0442/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0442/production/ref.

Kim:2004:EEU

[Kim04] Seung-Bum Kim. Eliminating extrapolation using point distribution criteria in scattered data interpolation. Computer Vision and Image Understanding: CVIU, 95(1): 30–53, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2013:GIR

[Kim13]  Minho Kim.  GPU iso-

surface raycasting of FCC datasets. Graphical Models, 75(2):90–101, March 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031200077X.

Kiselman:1996:RPD

[Kis96a] Christer O. Kiselman. Regularity properties of distance transformations in image analysis. Computer Vision and Image Understanding: CVIU, 64(3):390–398, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0067/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0067/production/
pdf.

Kishimoto:1996:CDC

[Kis96b]  Kazuo Kishimoto.  Char-

acterizing digital convexity and straightness in terms of “Length” and “Total Absolute Curvature”. Computer Vision and Image Understanding: CVIU, 63 (2):326–333, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0022/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0022/production/
                                        pdf.

Krueger:1996:SDF

[KJRA96] W. M. Krueger, S. D. Jost, K. Rossi, and U. Axen. On synthesizing discrete fractional Brownian motion with applications to image processing. Graphical Models and Image Processing: GMIP, 58(4):334–344, July 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0027/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0027/production/
                                        pdf.

Kadar:1979:CTD

[KK79] Ivan Kadar and Ludwik Kurz. Class of three-dimensional recursive parallelepiped masks. Computer Graphics and Image Processing, 11(3):262– 280, November 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Klette:1981:ATC

[KK81] Reinhard Klette and E. V. Krishnamurthy. Algorithms for testing convexity of digital polygons. Computer Graphics and Image Processing, 16 (2):177–184, June 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kulpa:1983:ACP

[KK83] Zenon Kulpa and Bjorn Kruse. Algorithms for circular propagation in discrete images. Computer Vision, Graphics, and Image Processing, 24(3):305–328, December 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kamada:1988:AGV

[KK88a] Tomihisa Kamada and Satoru Kawai. Advanced graphics for visualization of shielding relations. Computer Vision, Graphics, and Image Processing, 43(3):294– 312, September 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kamada:1988:SMC

[KK88b] Tomihisa Kamada and Satoru Kawai. A simple method for computing general position in displaying threedimensional objects. Computer Vision, Graphics, and Image Processing, 41(1):43– 56, January 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Knill:1990:LNO

[KK90] David C. Knill and Daniel Kersten. Learning a nearoptimal estimator for surface shape from shading. Computer Vision, Graphics, and Image Processing, 50(1):75– 100, April 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kosaka:1992:FVG

[KK92] Akio Kosaka and Avinash C. Kak. Fast vision-guided mobile robot navigation using model-based reasoning and prediction of uncertainties. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(3): 271–329, November 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). See erratum [KK93].

Kosaka:1993:EVN

[KK93] A. Kosaka and A. C. Kak. Erratum: Volume 56, number 3, (1992), in the article “Fast Vision-Guided Mobile Robot Navigation Using Model-Based Reasoning and Prediction of Uncertainties,” Akio Kosaka, Avinash C. Kak, pages 271–329. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):263, March 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1018/production;   http:
//www.idealibrary.com/links/

artid/ciun.1993.1018/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1018/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1018/production/

pdf. See [KK92].

Kweon:1994:ETT

[KK94] In So Kweon and Takeo Kanade. Extracting topographic terrain features from elevation maps. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):171– 182, March 1994. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1011/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1011/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1013/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1013/production/
pdf.

Kim:1995:EDC

[KK95] Jihong Kim and Yongmin Kim. Efficient 2-D convolution algorithm with the single-data multiple kernel approach. Graphical Models and Image Processing: GMIP, 57(2):175–182, March 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1017/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1017/production/
                                        pdf.

Kim:2007:RMB

[KK07] Sungho Kim and In So Kweon. Robust model-based scene interpretation by multilayered context information. Computer Vision and Image Understanding: CVIU, 105 (3):167–187, March 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2009:CCB

[KK09] Jun-Sik Kim and In So Kweon. Camera calibration based on arbitrary parallelograms. Computer Vision and Image Understanding: CVIU, 113(1):1–10, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2011:MRP

[KK11] Jun-Sik Kim and In So Kweon. Metric reconstruction of planes utilizing off-the-plane features. Computer Vision and Image Understanding: CVIU, 115(1):1–7, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kasaiezadeh:2013:MAS

[KK13] Alireza Kasaiezadeh and Amir Khajepour. Multi-agent stochastic level set method in image segmentation. Computer Vision and Image Understanding: CVIU, 117(9): 1147–1162, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000945.

Kyung:1996:NAT

[KKH96] Min-Ho Kyung, Myung-Soo Kim, and Sung Je Hong. A new approach to through-the-lens camera control. Graphical Models and Image Processing: GMIP, 58(3):262–285, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0022/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0022/production/
pdf.

Kim:1999:UDV

[KKK99] Dongryeol Kim, Jinsoo Kim, and Hyeong-Seok Ko. Unification of distance and volume optimization in surface simplification. Graphical Models and Image Processing: GMIP, 61 (6):363–367, November 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0506/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1999.0506/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0506/production/ref.

Kim:1998:TSI

[KKO98] Ku-Jin Kim, Myung-Soo Kim, and Kyungho Oh. Torus/sphere intersection based on a configuration space approach. Graphical Models and Image Processing: GMIP, 60(1):077–092, January 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0451/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0451/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0451/production/ref.

Keegan:1977:HCY

[KL77] James F. Keegan and A. M. Lesk. How can you tell if two line drawings are the same? Computer Graphics and Image Processing, 6(1): 90–92, February 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kanatani:1993:ICO

[KL93] Kenichi Kanatani and Wu Liu. 3D interpretation of conics and orthogonality. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3):286– 301, November 1993. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1043/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1043/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1045/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1045/production/
pdf.

Kadoury:2007:FDG

[KL07] Samuel Kadoury and Martin D. Levine. Face detection in gray scale images using locally linear embeddings. Computer Vision and Image Understanding: CVIU, 105(1): 1–20, January 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kumar:2010:IMB

[KL10] Avin Kumar and Baoxin Li.

On implementing motionbased Region of Interest detection on multi-core CELL. Computer Vision and Image Understanding: CVIU, 114(11):1139–1151, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2011:HAM

[KL11] Wonsik Kim and Kyoung Mu Lee. A hybrid approach for MRF optimization problems: Combination of stochastic sampling and deterministic algorithms. Computer Vision and Image Understanding: CVIU, 115(12): 1623–1637, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001640.

Kim:2013:RWB

[KL13] Hyejin Kim and Sung-Hee Lee. Reconstructing wholebody motions with wrist trajectories. Graphical Models, 75(6):328–345, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S152407031300026X.

Komodakis:2011:SIO

[KLBP11] Nikos Komodakis, Georg Langs, Horst Bischof, and Nikos Paragios. Special issue on optimization for vision, graphics and medical imaging: theory and applications. Computer Vision and Image Understanding: CVIU, 115(12):1597, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002165.

Klette:1980:POB

[Kle80]  Reinhard Klette.  Paral-

lel operations on binary images. Computer Graphics and Image Processing, 14(2):145– 158, October 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Klette:1985:DGP

[Kle85] Reinhard Klette. m-dimensional grid point space. Computer Vision, Graphics, and Image Processing, 30(1):1– 12, April 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Klette:2013:RCM

[Kle13] Gisela Klette. Recursive computation of minimum-length polygons. Computer Vision and Image Understanding: CVIU, 117(4):386–392, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001865.

Kim:1988:NCC

[KLK88] Seong-Dae D. Kim, Jeong-Hwan H. Lee, and Jae-Kyoon K. Kim. A new chain-coding algorithm for binary images using runlength codes. Computer Vision, Graphics, and Image Processing, 41(1):114– 128, January 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kodratoff:1984:LCS

[KLL84] Yves Kodratoff and Regine Lemerle-Loisel. Learning complex structural descriptions from examples. Computer Vision, Graphics, and Image Processing, 27(3):266– 290, September 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kim:2011:RBT

[KLL^{+}11] Kyung-Su Kim, Min-Jeong Lee, Ji-Won Lee, Tae-Woo Oh, and Hae-Yeoun Lee. Region-based tampering detection and recovery using homogeneity analysis in qualitysensitive imaging. Computer Vision and Image Understanding: CVIU, 115(9): 1308–1323, September 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001238.

Klucewicz:1978:PIA

[Klu78] I. M. Klucewicz. A piecewise C^{1} interpolant to arbitrarily spaced data. Computer Graphics and Image Processing, 8(??):92–112, August 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kim:2006:VTS

[KLV06] Youngmin Kim, Chang Ha Lee, and Amitabh Varshney. Vertex-transformation streams. Graphical Models, 68(4):371–383, July 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000373.

Koparkar:1984:CTP

[KM84] P. A. Koparkar and S. P.

Mudur. Computational techniques for processing parametric surfaces. Computer Vision, Graphics, and Image Processing, 28(3):303– 322, December 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kehtarnavaz:1989:FEM

[KM89a] N. Kehtarnavaz and S. Mohan. A framework for estimation of motion parameters from range images. Computer Vision, Graphics, and Image Processing, 45 (1):88–105, January 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kitchen:1989:ESD

[KM89b] L. J. Kitchen and J. A. Malin. The effect of spatial discretization on the magnitude and direction response of simple differential edge operators on a step edge. Computer Vision, Graphics, and Image Processing, 47 (2):243–258, August 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kayargadde:1994:EEP

[KM94] V. Kayargadde and J. B. Martens. Estimation of edge parameters and image blur using polynomial transforms. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(6):442–461, November 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1041/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1041/production/
pdf.

Krishnan:2000:PTS

[KM00] Shankar Krishnan and Dinesh Manocha. Partitioning trimmed spline surfaces into nonself-occluding regions for visibility computation. Graphical Models, 62 (4):283–307, July 2000. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0526;   http://www.
idealibrary.com/links/doi/

10.1006/gmod.2000.0526/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0526/ref.

Kowalczyk:2003:OCD

[KM03] M. Kowalczyk and W. S.

Mokrzycki. Obtaining complete 2 D view representation of polyhedra using concept of seedling single-view area. Computer Vision and Image Understanding: CVIU, 91 (3):280–301, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Krissian:2000:MBD

[KMA^{+}00] Karl Krissian, Grégoire Malandain, Nicholas Ayache, Régis Vaillant, and Yves Trousset. Model-based detection of tubular structures in 3D images. Computer Vision and Image Understanding: CVIU, 80(2):130–171, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0866;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0866/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0866/ref.

Kakadiaris:1997:IOS

[KMB97] Ioannis Kakadiaris, Dimitris Metaxas, and Ruzena Bajcsy. Inferring 2D object structure from the deformation of apparent contours. Computer Vision and Image Understanding: CVIU, 65(2):129–147, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0580/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0580/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0580/production/ref.

Kaufmann:2009:FSD

[KMBG09] Peter Kaufmann, Sebastian Martin, Mario Botsch, and Markus Gross. Flexible simulation of deformable models using discontinuous Galerkin FEM. Graphical Models, 71 (4):153–167, July 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000125.

Kroon:2009:ELL

[KMBH09] Bart Kroon, Sander Maas, Sabri Boughorbel, and Alan Hanjalic. Eye localization in low and standard definition content with application to face matching. Computer Vision and Image Understanding: CVIU, 113(8): 921–933, August 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kuhl:1984:GSR

[KMGC84] Frank P. Kuhl, O. Robert Mitchell, Marcus E. Glenn, and Didier J. Charpentier. Global shape recognition of 3-D objects using a differential library storage. Computer Vision, Graphics, and Image Processing, 27(1):97– 114, July 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kaveh:1979:UTB

[KMI79] M. Kaveh, R. K. Mueller, and R. D. Iverson. Ultrasonic tomography based on perturbation solutions of the wave equation. Computer Graphics and Image Processing, 9(2): 105–116, February 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kuo:2011:IBT

[KMN11] Paul Kuo, Dimitrios Makris, and Jean-Christophe Nebel. Integration of bottom-up/top-down approaches for 2D pose estimation using probabilistic Gaussian modelling. Computer Vision and Image Understanding: CVIU, 115(2): 242–255, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ko:2005:AOP

[KMP05] K. H. Ko, T. Maekawa, and N. M. Patrikalakis. Algorithms for optimal partial matching of free-form objects with scaling effects. Graphical Models, 67(2):29, March 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Koutis:2011:CPM

[KMT11] Ioannis Koutis, Gary L. Miller, and David Tolliver. Combinatorial preconditioners and multilevel solvers for problems in computer vision and image processing. Computer Vision and Image Understanding: CVIU, 115(12): 1638–1646, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001627.

Kim:1999:URL

[KN99] ZuWhan Kim and Ramakant Nevatia. Uncertain reasoning and learning for feature grouping. Computer Vision and Image Understanding: CVIU, 76(3):278–288, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0803/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0803/production/
pdf;   http://www.idealibrary.

com/links/artid/cviu.1999.
0803/production/ref.

Khamene:2003:MSM

[KN03] Ali Khamene and Shahriar Negahdaripour. Motion and structure from multiple cues; image motion, shading flow, and stereo disparity. Computer Vision and Image Understanding: CVIU, 90(1): 99–127, April 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:2004:ADC

[KN04] ZuWhan Kim and Ramakant Nevatia. Automatic description of complex buildings from multiple images. Computer Vision and Image Understanding: CVIU, 96(1): 60–95, October 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kybic:2011:BOF

[KN11] Jan Kybic and Claudia Nieuwenhuis. Bootstrap optical flow confidence and uncertainty measure. Computer Vision and Image Understanding: CVIU, 115(10): 1449–1462, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001536.

Kautsky:1984:SHM

[KNJ84] Jaroslav Kautsky, Nancy K. Nichols, and David L. B. Jupp. Smoothed histogram modification for image processing. Computer Vision, Graphics, and Image Processing, 26(3):271–291, June 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kurazume:2009:RFS

[KNO^{+}09] Ryo Kurazume, Kaori Nakamura, Toshiyuki Okada, Yoshinobu Sato, Nobuhiko Sugano, Tsuyoshi Koyama, Yumi Iwashita, and Tsutomu Hasegawa. 3D reconstruction of a femoral shape using a parametric model and two 2D fluoroscopic images. Computer Vision and Image Understanding: CVIU, 113(2): 202–211, February 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kodratoff:1976:GSP

[Kod76]  Yves Kodratoff.  Genera-

tion and semantics of patterns in a discrete space. Computer Graphics and Image Processing, 5(4):447– 458, December 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kohler:1981:SSB

[Koh81] Ralf Kohler. Segmentation system based on thresholding. Computer Graphics and Image Processing, 15(4): 319–338, April 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kolers:1983:SFV

[Kol83] Paul A. Kolers. Some features of visual form. Computer Vision, Graphics, and Image Processing, 23(1):15– 41, July 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kishimoto:1987:TAM

[KON87] K. Kishimoto, K. Onaga, and E. Nakamae. Theoretical assessments of mean square errors of antialiasing filters. Computer Vision, Graphics, and Image Processing, 37(3):428–437, March 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kumar:2010:ITC

[KORC10] S. Kumar, S. H. Ong, S. Ranganath, and F. T. Chew. Invariant texture classification for biomedical cell specimens via non-linear polar map filtering. Computer Vision and Image Understanding: CVIU, 114(1):44–53, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kouzani:2003:LHF

[Kou03] A. Z. Kouzani. Locating human faces within images. Computer Vision and Image Understanding: CVIU, 91 (3):247–279, September 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kovalevsky:1986:DEV

[Kov86]  V. A. Kovalevsky.  Dia-

log on “expert” vision systems: Comments. Computer Vision, Graphics, and Image Processing, 34(1):111– 114, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See [Ros86b, Nag86, Tho86, Uhr86].

Kovalevsky:1989:FTA

[Kov89] V. A. Kovalevsky. Finite topology as applied to image analysis. Computer Vision, Graphics, and Image Processing, 46(2):141–161, May 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kishimoto:1986:TEA

[KOY86] Kazuo Kishimoto, Kenji Onaga, and Kiyoshi Yamamoto. Theoretical error assessments of curved line digitization schemes on graphic displays. Computer Vision, Graphics, and Image Processing, 35 (2):170–180, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kovalev:1996:MCO

[KP96] Vassili Kovalev and Maria Petrou. Multidimensional cooccurrence matrices for object recognition and matching. Graphical Models and Image Processing: GMIP, 58(3):187–197, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0016/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0016/production/
pdf.

Kim:1997:SSP

[KP97] Bang-Hwan Kim and Rae-Hong Park. Shape from shading and photometric stereo using surface approximation by Legendre polynomials. Computer Vision and Image Understanding: CVIU, 66(3):255–270, June 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0515/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0515/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0515/production/ref.

Kruger:2000:OOR

[KP00] Norbert Krüger and Gabriele Peters. ORASSYLL: Object recognition with autonomously learned and sparse symbolic representations based on metrically organized local line detectors. Computer Vision and Image Understanding: CVIU, 77(1):48–77, January 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0794/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0794/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0794/production/ref.

Karciauskas:2011:RS

[KP11] Kestutis Karčiauskas and Jörg Peters. Rational G^{2} splines. Graphical Models, 73 (5):286–295, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S152407031100021X.

Karciauskas:2012:FFS

[KP12] Keştutis Karciauskas and Jörg Peters. Free-form splines combining NURBS and basic shapes. Graphical Models, 74(6):351–360, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000379.

Kamgar-Parsi:1990:CSS

[KPE90] Behrooz Kamgar-Parsi and Roger D. Eastman. Calibration of a stereo system with small relative angles. Computer Vision, Graphics, and Image Processing, 51(1):1–19, July 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Krsek:2002:DIB

[KPH02] Pavel Krsek, Tomá Pajdla, and Václav Hlavá. Differential invariants as the base of triangulated surface registration. Computer Vision and Image Understanding: CVIU, 87(1– 3):27–38, July 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kolsch:2007:SIV

[KPKH07] Mathias Kölsch, Vladimir Pavlović, Branislav Kisačanin, and Thomas S. Huang. Special issue on vision for human– computer interaction. Computer Vision and Image Understanding: CVIU, 108(1–2): 1–3, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kamgar-Parsi:1990:SFS

[KPKPW90] Br. Kamgar-Parsi, Bz. Kamgar-Parsi, and H. Wechsler. Simultaneous fitting of several planes to point sets using neural networks. Computer Vision, Graphics, and Image Processing, 52(3):341– 359, December 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kamgar-Parsi:1991:MGP

[KPMR91] B. Kamgar-Parsi, A. Margalit, and A. Rosenfeld. Matching general polygonal arcs. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (2):227–234, March 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Kristan:2009:CWT

[KPPK09] Matej Kristan, Janez Perš, Matej Perše, and Stanislav Kovačič. Closed-world tracking of multiple interacting targets for indoor-sports applications. Computer Vision and Image Understanding: CVIU, 113(5):598–611, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

KameswaraRao:1976:PSA

[KPS76] C. V. Kameswara Rao, B. Prasada, and K. R. Sarma. Parallel shrinking algorithm for binary patterns. Computer Graphics and Image Processing, 5(2):265–270, June 1976. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kong:1985:CAA

[KR85a] T. Y. Kong and A. W. Roscoe.

Continuous analogs of axiomatized digital surfaces. Computer Vision, Graphics, and Image Processing, 29(1):60–86, January 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kong:1985:TBD

[KR85b] T. Y. Kong and A. W.

Roscoe. A theory of binary digital pictures. Computer Vision, Graphics, and Image Processing, 32(2):221– 243, November 1985. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kong:1989:DTI

[KR89] T. Y. Kong and A. Rosenfeld. Digital topology. introduction and survey. Computer Vision, Graphics, and Image Processing, 48(3):357– 393, December 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kliot:1998:IBS

[KR98] Michael Kliot and Ehud Rivlin. Invariant-based shape retrieval in pictorial databases. Computer Vision and Image Understanding:

CVIU, 71(2):182–197, August 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0709/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0709/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0709/production/ref.

Kundur:1999:NAV

[KR99] Sridhar R. Kundur and Daniel Raviv. Novel active vision-based visual threat cue for autonomous navigation tasks. Computer Vision and Image Understanding: CVIU, 73(2):169–182, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0699/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0699/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0699/production/ref.

Krishnamurthy:1984:HOT

[Kri84] E. V. Krishnamurthy. Highorder tensor product approximation for two- and three- dimensional image blocks with application to multiresolution image representation. Computer Vision, Graphics, and Image Processing, 25(3):393–398, March 1984.

CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kriegman:1992:CSP

[Kri92] D. J. Kriegman. Computing stable poses of piecewise smooth objects. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(2):109– 118, March 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Kilambi:2008:EPC

[KRJ^{+}08] Prahlad Kilambi, Evan Ribnick, Ajay J. Joshi, Osama Masoud, and Nikolaos Papanikolopoulos. Estimating pedestrian counts in groups. Computer Vision and Image Understanding: CVIU, 110 (1):43–59, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kjellstrom:2011:VOA

[KRK11] Hedvig Kjellström, Javier Romero, and Danica Kragić. Visual object-action recognition: Inferring object affordances from human demonstration. Computer Vision and Image Understanding: CVIU, 115(1):81–90, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Krotkov:1986:VHR

[Kro86]  Eric P. Krotkov.  Visual

hyperacuity: Representation and computation of high precision position information. Computer Vision, Graphics, and Image Processing, 33 (1):99–115, January 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kartikeyan:1989:UAI

[KS89] B. Kartikeyan and A. Sarkar. A unified approach for image segmentation using exact statistics. Computer Vision, Graphics, and Image Processing, 48(2):217– 229, November 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kalivas:1991:RMM

[KS91a] Dimitrios S. Kalivas and Alexander A. Sawchuck. A region matching motion estimation algorithm. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(2):275–288, September 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Kartikeyan:1991:IAA

[KS91b] B. Kartikeyan and A. Sarkar. An identification approach for 2D autoregressive models in describing textures. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2): 121–131, March 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Kita:1991:EAS

[KS91c] Y. Kita and Y. Shirai. Extraction of accurate stomach contours from X-ray images of barium-filled stomachs and its application to detect potential abnormalities. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):447–456, September 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Karabassis:1995:AII

[KS95a] Evangelos Karabassis and Minas E. Spetsakis. An analysis of image interpolation, differentiation, and reduction using local polynomial fits. Graphical Models and Image Processing: GMIP, 57(3):183–196, May 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1018/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1018/production/
pdf.

Kehtarnavaz:1995:EAC

[KS95b] N. D. Kehtarnavaz and W. Sohn. Error analysis of camera movements in stereo vehicle tracking systems. Computer Vision and Image Understanding: CVIU, 62(3):347–359, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1059/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1059/production/
pdf.

Kimia:1996:GHE

[KS96] Benjamin B. Kimia and Kaleem Siddiqi. Geometric heat equation and nonlinear diffusion of shapes and images. Computer Vision and Image Understanding: CVIU, 64(3):305–322, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0062/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0062/production/
pdf.

Kim:2000:SIP

[KS00] Myung-Soo Kim and Hans-Peter Seidel. Special issue on Pacific Graphics ’99 in Graphical Models. Graphical Models, 62(6):389, November 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0533;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0533/
pdf.

Kang:2002:ELI

[KS02] Hyung Woo Kang and Sung Yong Shin. Enhanced lane: interactive image segmentation by incremental path map construction. Graphical Models, 64(5):282–303, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Karacal:2003:RDS

[KS03] Bilge Karaçal and Wesley Snyder. Reconstructing discontinuous surfaces from a given gradient field using partial integrability. Computer Vision and Image Understanding: CVIU, 92(1):78– 111, October 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Knopf:2004:ISD

[KS04a] George K. Knopf and Archana Sangole. Interpolating scattered data using 2D selforganizing feature maps. Graphical Models, 66(1):50– 69, January 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Krivic:2004:PLO

[KS04b] Jaka Krivic and Franc Solina. Part-level object recognition using superquadrics. Computer Vision and Image Understanding: CVIU, 95(1): 105–126, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kompella:2012:CRB

[KS12] Varun Raj Kompella and Peter Sturm. Collective-reward based approach for detection of semi-transparent objects in single images. Computer Vision and Image Understanding: CVIU, 116(4):484–499, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002591.

Kwok:1997:SAT

[KSC97] S. H. Kwok, W. C. Siu, and A. G. Constantinides. A scalable and adaptive temporal segmentation algorithm for video coding. Graphical Models and Image Processing: GMIP, 59(3):128–138, May 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0423/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0423/production/
pdf.

Kehtarnavaz:1988:SST

[KSd88] N. Kehtarnavaz, L. R. Simar, and R. J. P. de Figueiredo. A syntactic/semantic technique for surface reconstruction from cross-sectional contours. Computer Vision, Graphics, and Image Processing, 42(3):399–409, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kekre:1984:RNI

[KSG84] H. B. Kekre, S. C. Sahasrabudhe, and N. C. Goyal. Restoration of noisy images using a raised cosine function approximation. Computer Vision, Graphics, and Image Processing, 26(1):17– 29, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kurtek:2013:SAM

[KSG^{+}13] Sebastian Kurtek, Jingyong Su, Cindy Grimm, Michelle Vaughan, Ross Sowell, and Anuj Srivastava. Statistical analysis of manual segmentations of structures in medical images. Computer Vision and Image Understanding: CVIU, 117(9):1036–1050, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000714.

Kise:1998:SPI

[KSI98] Koichi Kise, Akinori Sato, and Motoi Iwata. Segmentation of page images using the area Voronoi diagram. Computer Vision and Image Understanding: CVIU, 70(3):370–382, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0684/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0684/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0684/production/ref.

Kimmel:1995:SDM

[KSKB95] Ron Kimmel, Doron Shaked, Nahum Kiryati, and Alfred M. Bruckstein. Skeletonization via distance maps and level sets. Computer Vision and Image Understanding: CVIU, 62(3):382–391, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1062/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1062/production/
pdf.

Kobbelt:2006:SIS

[KSM^{+}06] Leif Kobbelt, Vadim Shapiro, Botsch Mario, Cazals Frederic, Cohen-Or Danny, Hoppe Hugues, Hu Shimin, Jüttler Bert, Kim Myung-Soo, O’Brien James, Puppo Enrico, Velho Luiz, Wang Wenping, and Zeilfelder Frank. Special issue on SPM 05. Graphical Models, 68(3):235–236, May 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000233.

Khanloo:2012:LMF

[KSR^{+}12] Bahman Yari Saeed Khanloo, Ferdinand Stefanus, Mani Ranjbar, Ze-Nian Li, Nicolas Saunier, Tarek Sayed, and Greg Mori. A large margin framework for single camera offline tracking with hybrid cues. Computer Vision and Image Understanding: CVIU, 116(6):676–689, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000215.

Krishnan:1992:MPR

[KSS92] Radha Krishnan, H. J. Sommer, III, and Peter D. Spidaliere. Monocular pose of a rigid body using point landmarks. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (3):307–316, May 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Kang:1997:PFT

[KSS97] Sing Bing Kang, Richard Szeliski, and Heung-Yeung Shum. A parallel feature tracker for extended image sequences. Computer Vision and Image Understanding: CVIU, 67(3): 296–310, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0519/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0519/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0519/production/ref.

Klein:2000:RSS

[KSS00] Reinhard Klein, Andreas Schilling, and Wolfgang Straßer. Reconstruction and simplification of surfaces from contours. Graphical Models, 62(6):429–443, November 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2000.0530;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/gmod.2000.0530/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2000.0530/ref.

Kye:2008:ICP

[KSS08] Heewon Kye, Byeong-Seok Shin, and Yeong Gil Shin.

Interactive classification for pre-integrated volume rendering of high-precision volume data. Graphical Models, 70(6):125–132, November 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000088.

Kapur:1985:NMG

[KSW85] J. N. Kapur, P. K. Sahoo, and A. K. C. Wong. A new method for gray-level picture thresholding using the entropy of the histogram. Computer Vision, Graphics, and Image Processing, 29(3):273–285, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Klette:1996:PDP

[KSŽ96] Reinhard Klette, Ivan Stojmenovič, and Joviša Žunić. A parametrization of digital planes by least-squares fits and generalizations. Graphical Models and Image Processing: GMIP, 58(3):295–300, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0024/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0024/production/
pdf.

Kropatsch:1989:DSD

[KT89] W. G. Kropatsch and H. Tockner. Detecting the straightness of digital curves in O(N) steps. Computer Vision, Graphics, and Image Processing, 45(1):1–21, January 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See note [Ano92a].

Krotosky:2007:MIB

[KT07] Stephen J. Krotosky and Mohan M. Trivedi. Mutual information based registration of multimodal stereo videos for person tracking. Computer Vision and Image Understanding: CVIU, 106(2– 3):270–287, May/June 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kohli:2008:MUG

[KT08] Pushmeet Kohli and Philip H. S. Torr. Measuring uncertainty in graph cut solutions. Computer Vision and Image Understanding: CVIU, 112(1):30–38, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kozinska:1997:MAU

[KTNO97] Dorota Kozinska, Oleh J. Tretiak, Jonathan Nissanov, and Cengizhan Ozturk. Multidimensional alignment using the Euclidean distance transform. Graphical Models and Image Processing: GMIP, 59 (6):373–387, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0447/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0447/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0447/production/ref.

Komodakis:2008:PVC

[KTP08] Nikos Komodakis, Georgios Tziritas, and Nikos Paragios. Performance vs computational efficiency for optimizing single and dynamic MRFs: Setting the state of the art with primal-dual strategies. Computer Vision and Image Understanding: CVIU, 112 (1):14–29, October 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ku:1984:PMH

[Ku84] Fu-Nian N. Ku. The principles and methods of histogram modification adapted for visual perception. Computer Vision, Graphics, and Image Processing, 26(1):107– 117, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kong:1992:JFS

[KU92] T. Yung Kong and Jarayam K.

Udupa. A justification of a fast surface tracking algorithm. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(2):162– 170, March 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Kumaran:1995:DWB

[KU95] Muthu Kumaran and Scott E. Umbaugh. A dynamic window-based runlength coding algorithm applied to graylevel images. Graphical Models and Image Processing: GMIP, 57(4):267–282, July 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1025/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1025/production/
pdf.

Kuba:1984:RTD

[Kub84] Attila Kuba. The reconstruction of two-directionally connected binary patterns from their two orthogonal projections. Computer Vision, Graphics, and Image Processing, 27(3):249–265, September 1984. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kuijper:2008:EES

[Kui08] Arjan Kuijper. Exploring and exploiting the structure of saddle points in Gaussian scale space. Computer Vision and Image Understanding: CVIU, 112(3): 337–349, December 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kulpa:1977:APM

[Kul77] Zenon Kulpa. Area and perimeter measurement of blobs in discrete binary pictures. Computer Graphics and Image Processing, 6(5): 434–451, October 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kulpa:1979:NPB

[Kul79a] Z. Kulpa. A note on the paper by B. K. P. Horn: “Circle generator for display devices”. Computer Graphics and Image Processing, 9(1): 102–103, January 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). See [Hor76, Hor79].

Kulpa:1979:PDC

[Kul79b] Zenon Kulpa. On the properties of discrete circles, rings, and disks. Computer Graphics and Image Processing, 10(4): 348–365, August 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Kulpa:1983:MAA

[Kul83] Zenon Kulpa. More about areas and perimeters of quantized objects. Computer Vision, Graphics, and Image Processing, 22(2):268– 276, May 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kehl:2006:MTC

[KV06] Roland Kehl and Luc Van Gool. Markerless tracking of complex human motions from multiple views. Computer Vision and Image Understanding: CVIU, 104(2–3):190–209, November/December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Koster:1997:HLM

[KVdG^{+}97] André S. E. Koster, Koen L. Vincken, Cornelis N. de Graaf, Olaf C. Zander, and Max A. Viergever. Heuristic linking models in multiscale image segmentation. Computer Vision and Image Understanding: CVIU, 65 (3):382–402, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0490/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0490/production/
pdf;   http://www.idealibrary.

com/links/artid/cviu.1996.
0490/production/ref.

Karl:1994:REP

[KVW94] William C. Karl, George C.

Verghese, and Alan S. Willsky. Reconstructing ellipsoids from projections. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(2):124–139, March 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1012/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1012/production/
pdf.

Kass:1987:AOP

[KW87] Michael Kass and Andrew Witkin. Analyzing oriented patterns. Computer Vision, Graphics, and Image Processing, 37(3):362– 385, March 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kang:1999:CEC

[KW99] Sing Bing Kang and Richard Weiss. Characterization of errors in compositing panoramic images. Computer Vision and Image Understanding: CVIU, 73(2):269–280, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0727/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0727/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0727/production/ref.

Keller:2000:FRB

[KW00] James M. Keller and Xiaomei Wang. A fuzzy rulebased approach to scene description involving spatial relationships. Computer Vision and Image Understanding: CVIU, 80(1):21–41, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0872;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0872/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0872/ref.

Koppen:2012:BRM

[KW12] W. P. Koppen and M. Worring. Backtracking: Retrospective multi-target tracking. Computer Vision and Image Understanding: CVIU, 116 (9):967–980, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000653.

Kasturi:1984:IRS

[KWK84] Rangachar Kasturi, John F.

Walkup, and Thomas F. Krile. Image restoration in signal-dependent noise using a Markovian covariance model. Computer Vision, Graphics, and Image Processing, 28 (3):363–376, December 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kara:1994:SRP

[KWK94] Atsushi Kara, D. Mitchell Wilkes, and Kazuhiko Kawamura. 3D structure reconstruction from point correspondences between two perspective projections. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3):392– 397, November 1994. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1065/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1065/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1071/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1071/production/
pdf.

Wu:2013:CGO

[kWwZ13] Fu kun Wu and Chang wen Zheng. A comprehensive geometrical optics application for wave rendering. Graphical Models, 75 (6):318–327, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000246.

Kim:1986:RAP

[KY86] Vladimir Kim and Leonid Yaroslavskii. Rank algorithms for picture processing. Computer Vision, Graphics, and Image Processing, 35(2):234–258, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Koren:2006:ASE

[KY06] Raz Koren and Yitzhak Yitzhaky. Automatic selection of edge detector parameters based on spatial and statistical measures. Computer Vision and Image Understanding: CVIU, 102(2): 204–213, May 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Koniusz:2013:CML

[KYM13] Piotr Koniusz, Fei Yan, and Krystian Mikolajczyk. Comparison of mid-level feature coding approaches and pooling strategies in visual concept detection. Computer Vision and Image Understanding: CVIU, 117(5):479–492, May 2013. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001725.

Kamel:1993:EBC

[KZ93] Mohamed Kamel and Aiguo Zhao. Extraction of binary character/graphics images from grayscale document images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(3):203–217, May 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1015/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1015/production/
pdf.

Klette:1999:DAM

[KŽ99] Reinhard Klette and Joviša Žunić. Digital approximation of moments of convex regions. Graphical Models and Image Processing: GMIP, 61(5):274–298, September 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0501/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0501/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0501/production/ref.

Kosecka:2005:EMP

[KZ05] Jana Kosecká and Wei Zhang. Extraction, matching, and pose recovery based on dominant rectangular structures. Computer Vision and Image Understanding: CVIU, 100 (3):274–293, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Klette:2012:ASD

[KŽ12] Reinhard Klette and Joviša Žunić. ADR shape descriptor — distance between shape centroids versus shape diameter. Computer Vision and Image Understanding: CVIU, 116(6):690–697, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000331.

Khan:2011:SBA

[KZD^{+}11] Rez Khan, Qin Zhang, Shayan Darayan, Sankari Dhandapani, Sucharit Katyal, Clint Greene, Chandra Bajaj, and David Ress. Surface-based analysis methods for highresolution functional magnetic resonance imaging. Graphical Models, 73(6):313– 322, November 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000421.

Kwok:2012:CCB

[KZW12] Tsz-Ho Kwok, Yunbo Zhang, and Charlie C. L. Wang. Constructing common base domain by cues from Voronoi diagram. Graphical Models, 74(4):152–163, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000197.

Lourakis:2005:ECC

[LA05] Manolis I. A. Lourakis and Antonis A. Argyros. Efficient, causal camera tracking in unprepared environments. Computer Vision and Image Understanding: CVIU, 99(2): 259–290, August 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Law:2011:SVM

[LA11] Alvin J. Law and Daniel G.

Aliaga. Single viewpoint model completion of symmetric objects for digital inspection. Computer Vision and Image Understanding: CVIU, 115(5):603–610, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lai:2000:RIM

[Lai00]  Shang-Hong Lai.  Ro-

bust image matching under partial occlusion and spatially varying illumination change. Computer Vision and Image Understanding: CVIU, 78(1):84–98, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0829;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0829/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0829/ref.

Lu:2010:CBO

[LAL^{+}10] ChengEn Lu, Nagesh Adluru, Haibin Ling, Guangxi Zhu, and Longin Jan Latecki. Contour based object detection using part bundles. Computer Vision and Image Understanding: CVIU, 114(7): 827–834, July 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lam:1984:PDU

[Lam84] K. P. Lam. Position determination using generalized multidirectional gradient codes. Computer Vision, Graphics, and Image Processing, 28 (2):228–239, November 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Langridge:1984:DDF

[Lan84]  D. J. Langridge.  Detec-

tion of discontinuities in the first derivatives of surfaces.

Computer Vision, Graphics, and Image Processing, 27 (3):291–308, September 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Landau:1987:ECA

[Lan87] U. M. Landau. Estimation of a circular arc center and its radius. Computer Vision, Graphics, and Image Processing, 38(3):317–326, June 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Landraud:1991:IRE

[Lan91] A. M. Landraud. Image restoration and enhancement of characters, using convex projection methods. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1):85–92, January 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Laprade:1988:SMS

[Lap88] Robert H. Laprade. Splitand-merge segmentation of aerial photographs. Computer Vision, Graphics, and Image Processing, 44(1):77– 86, October 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Laszlo:1992:FGD

[Las92] Michael J. Laszlo. Fast generation and display of isosurface wireframes. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(6): 473–483, November 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Liang:1994:MAT

[LAS94] Su Liang, M. Ahmadi, and M. Shridhar. A morphological approach to text string extraction from regular periodic overlapping text/ background images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(5):402–413, September 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1036/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1036/production/
pdf.

Latecki:1993:TCC

[Lat93] Longin Latecki. Topological connectedness and 8connectedness in digital pictures. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):261–262, March 1993. CODEN CIUNEJ.

ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1017/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1017/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1017/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1017/production/
pdf.

Latecki:1997:WCP

[Lat97] Longin Jan Latecki. 3D wellcomposed pictures. Graphical Models and Image Processing: GMIP, 59(3):164–172, May 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0422/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0422/production/
pdf.

Laurentini:1997:HMS

[Lau97] Aldo Laurentini. How many 2D silhouettes does it take to reconstruct a 3D object? Computer Vision and Image Understanding: CVIU, 67(1):81–87, July 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0508/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1996.0508/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0508/production/ref.

Lawton:1983:PTM

[Law83] Daryl T. Lawton. Processing translational motion sequences. Computer Vision, Graphics, and Image Processing, 22(1):116–144, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lunscher:1987:FBI

[LB87] W. H. H. J. Lunscher and M. P. Beddoes. Fast binaryimage boundary extraction. Computer Vision, Graphics, and Image Processing, 38(3):229–257, June 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lee:1997:NES

[LB97] Thomas C. M. Lee and Mark Berman. Nonparametric estimation and simulation of twodimensional Gaussian image textures. Graphical Models and Image Processing: GMIP, 59(6):434–445, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0439/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0439/production/
                                        pdf;   http://www.idealibrary.

com/links/artid/gmip.1997.
0439/production/ref.

Li:1998:MGP

[LB98] Fuxing Li and Michael Brady. Modeling the ground plane transformation for real-time obstacle detection. Computer Vision and Image Understanding: CVIU, 71(1):137–152, July 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0645/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0645/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0645/production/ref.

Leonardis:2000:RRU

[LB00] Ale Leonardis and Horst Bischof. Robust recognition using eigenimages. Computer Vision and Image Understanding: CVIU, 78(1):99–118, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0830;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0830/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0830/ref.

Lelescu:2004:RCL

[LB04] Dan Lelescu and Frank Bossen. Representation and coding of light field data. Graphical Models, 66(4):203– 225, July 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Levine:2005:DRS

[LB05] Martin D. Levine and Jisnu Bhattacharyya. Detecting and removing specularities in facial images. Computer Vision and Image Understanding: CVIU, 100(3):330– 356, December 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

LeCallennec:2006:IMD

[LB06] Benoît Le Callennec and Ronan Boulic. Interactive motion deformation with prioritized constraints. Graphical Models, 68(2):175–193, March 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000263.

Lee:2008:TFT

[LB08] Sang-Chul Lee and Peter Bajcsy. Trajectory fusion for three-dimensional volume reconstruction. Computer Vision and Image Understanding: CVIU, 110(1):19–31, April 2008. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Lim:2010:EEU

[LB10] John Lim and Nick Barnes.

Estimation of the epipole using optical flow at antipodal points. Computer Vision and Image Understanding: CVIU, 114(2):245– 253, February 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Loza:2010:NGM

[LBCA10] Artur Loza, David Bull, Nishan Canagarajah, and Alin Achim. Non-Gaussian modelbased fusion of noisy images in the wavelet domain. Computer Vision and Image Understanding: CVIU, 114(1): 54–65, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

LeNegrate:1992:IET

[LBD92] Alain Le Négrate, Azeddine Beghdadi, and Henri Dupoisot. An image enhancement technique and its evaluation through bimodality analysis. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1): 13–22, January 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Larsen:2010:SIT

[LBK10] Rasmus Larsen, Erhardt Barth, and Andreas Kolb. Special issue on Time-of-Flight camera based computer vision. Computer Vision and Image Understanding: CVIU, 114(12):1317, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:2004:UAF

[LBM04] Guiqing Li, Hujun Bao, and Weiyin Ma. A unified approach for fairing arbitrary polygonal meshes. Graphical Models, 66(3):160–179, May 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Loss:2009:IMS

[LBNS09] Leandro Loss, George Bebis, Mircea Nicolescu, and Alexei Skurikhin. An iterative multi-scale tensor voting scheme for perceptual grouping of natural shapes in cluttered backgrounds. Computer Vision and Image Understanding: CVIU, 113(1): 126–149, January 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lester:1980:LTB

[LBS80] James M. Lester, John F. Brenner, and William D. Selles. Local transforms for biomedical image analysis. Computer Graphics and Image Processing, 13(1): 17–30, May 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Liu:2002:NVC

[LBSP02] Xinguo Liu, Hujun Bao, Heung-Yeung Shum, and Qunsheng Peng. A novel volume constrained smoothing method for meshes. Graphical Models, 64(3–4):169–182, May 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Lane:1979:GSL

[LC79] Jeff Lane and Loren Carpenter. A generalized scan line algorithm for the computer display of parametrically defined surfaces. Computer Graphics and Image Processing, 11(3): 290–297, November 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Landy:1985:VCE

[LC85a] Michael S. Landy and Yoav Cohen. Vectorgraph coding: Efficient coding of line drawings. Computer Vision, Graphics, and Image Processing, 30(3):331–344, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lee:1985:DHB

[LC85b] Hsi-Jian J. Lee and Zen Chen.

Determination of 3D human body postures from a single view. Computer Vision, Graphics, and Image Processing, 30(2):148–168, May 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liu:1988:MPR

[LC88a] Z.-Q. Liu and Terry M. Caelli. Multiobject pattern recognition and detection in noisy backgrounds using a hierarchical approach. Computer Vision, Graphics, and Image Processing, 44 (3):296–306, December 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liu:1988:SAR

[LC88b] Zhi-Qiang Liu and Terry Caelli. A sequential adaptive recursive filter for image restoration. Computer Vision, Graphics, and Image Processing, 44(3):332– 349, December 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liu:2009:VRB

[LC09] David Liu and Tsuhan Chen. Video retrieval based on object discovery. Computer Vision and Image Understanding: CVIU, 113(3):397– 404, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lai:2011:FII

[LC11] Rongjie Lai and Tony F.

Chan. A framework for intrinsic image processing on surfaces. Computer Vision and Image Understanding: CVIU, 115(12):1647–1661, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001603.

Lin:1989:NSI

[LCC89] Wei-Chung C. Lin, Shiuh-Yung Y. Chen, and Chin-Tu T. Chen. A new surface interpolation technique for reconstructing 3D objects from serial cross-sections. Computer Vision, Graphics, and Image Processing, 48 (1):124–143, October 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Langrana:1997:FIV

[LCD97] Noshir A. Langrana, Yuan Chen, and Atish K. Das. Feature identification from vectorized mechanical drawings. Computer Vision and Image Understanding: CVIU, 68(2):127–145, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0548/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0548/production/

                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0548/production/ref.

Levitan:1995:IMG

[LCH95] Emanuel Levitan, Michael Chan, and Gabor T. Herman. Image-modeling Gibbs priors. Graphical Models and Image Processing: GMIP, 57(2):117–130, March 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1013/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1013/production/
                                        pdf.

Lee:1990:CPS

[LCP90] Sang Uk Lee, Seok Yoon Chung, and Rae Hong Park. A comparative performance study of several global thresholding techniques for segmentation. Computer Vision, Graphics, and Image Processing, 52(2):171–190, November 1990. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Landy:1984:HUB

[LCS84] Michael S. Landy, Yoav Cohen, and George Sperling. Hips: a Unix-based image processing system. Computer Vision, Graphics, and Image Processing, 25(3):331– 347, March 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Laptev:2007:LVA

[LCSL07] Ivan Laptev, Barbara Caputo, Christian Schüldt, and Tony Lindeberg. Local velocityadapted motion events for spatio-temporal recognition. Computer Vision and Image Understanding: CVIU, 108 (3):207–229, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lanman:2009:SSL

[LCT09] Douglas Lanman, Daniel Crispell, and Gabriel Taubin. Surround structured lighting: 3-D scanning with orthographic illumination. Computer Vision and Image Understanding: CVIU, 113(11): 1107–1117, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lim:1991:EEI

[LcTT91] Hock Lim, Kah chye Tan, and B. T. G. Tan. Edge errors in inverse and Wiener filter restorations of motionblurred images and their windowing treatment. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2): 186–195, March 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Li:2001:EEF

[LCZ^{+}01] B. Li, R. Chellappa, Q. Zheng, S. Der, N. Nasrabadi, L. Chan, and L. Wang. Experimental evaluation of FLIR ATR approaches — A comparative study. Computer Vision and Image Understanding: CVIU, 84(1):5–24, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0938;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0938/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0938/ref.

Li:2009:FCS

[LCZ09] Shu-Xiao Li, Hong-Xing Chang, and Cheng-Fei Zhu. Fast curvilinear structure extraction and delineation using density estimation. Computer Vision and Image Understanding: CVIU, 113(6): 763–775, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:1990:TFC

[LD90] Hsi-Jian Lee and Hsi-Chou Deng. Three-frame corner matching and moving object extraction in a sequence of images. Computer Vision, Graphics, and Image Processing, 52(2):210– 238, November 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Loce:1995:MAE

[LD95] Robert P. Loce and Edward R. Dougherty. Meanabsolute-error representation and optimization of computational-morphological filters. Graphical Models and Image Processing: GMIP, 57(1):27–37, January 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1004/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1004/production/
pdf.

Liu:1998:GIL

[LD98] Wenyin Liu and Dov Dori.

A generic integrated line detection algorithm and its object-process specification. Computer Vision and Image Understanding: CVIU, 70(3):420–437, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0683/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0683/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/cviu.1998.
                                        0683/production/ref.

Liu:2013:MRV

[LDC^{+}13] Ningning Liu, Emmanuel Dellandréa, Liming Chen, Chao Zhu, Yu Zhang, Charles-Edmond Bichot, Stéphane Bres, and Bruno Tellez. Multimodal recognition of visual concepts using histograms of textual concepts and selective weighted late fusion scheme. Computer Vision and Image Understanding: CVIU, 117(5):493–512, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001695.

Liang:2009:MCC

[LDD09] Jian Liang, Daniel DeMenthon, and David Doermann. Mosaicing of camera-captured document images. Computer Vision and Image Understanding: CVIU, 113(4): 572–579, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lefevre:2013:EEF

[LDGS^{+}13] Thierry Lefevre, Bernadette Dorizzi, Sonia Garcia-Salicetti, Nadege Lemperiere, and Stephane Belardi. Effective elliptic fitting for iris normalization. Computer Vision and Image Understanding: CVIU, 117(6):732–745, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421300012X.

Lavest:1997:IRZ

[LDPD97] J. M. Lavest, C. Delherm, B. Peuchot, and N. Daucher. Implicit reconstruction by zooming. Computer Vision and Image Understanding: CVIU, 66(3):301–315, June 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0511/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0511/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0511/production/ref.

Lee:2009:DSO

[LE09] Chan-Su Lee and Ahmed Elgammal. Dynamic shape outlier detection for human locomotion. Computer Vision and Image Understanding: CVIU, 113(3):332–344, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Leavers:1992:DGH

[Lea92] V. F. Leavers. The dynamic generalized Hough transform: its relationship to the probabilistic Hough transforms and an application to the concurrent detection of circles and ellipses. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (3):381–398, November 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Leavers:1993:WHT

[Lea93] V. F. Leavers. Which Hough transform? Computer Vision, Graphics, and Image Processing. Image Understanding, 58(2):250–264, September 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1041/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1041/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1043/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1043/production/
pdf.

Lmaati:2010:DSE

[LEA^{+}10] Elmustapha Ait Lmaati, Ahmed El Oirrak, Driss Aboutajdine, Mohamed Daoudi, and Mohammed Najib Kaddioui. A 3-D Search engine based on Fourier series. Computer Vision and Image Understanding: CVIU, 114 (1):1–7, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lezoray:2007:GRC

[LEB07] Olivier Lezoray, Abderrahim Elmoataz, and Sébastien Bougleux. Graph regularization for color image processing. Computer Vision and Image Understanding: CVIU, 107(1–2):38–55, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:1976:RMP

[Lee76] Robert M. Lee. Recording motion picture sound tracks using a computer output film recorder. Computer Graphics and Image Processing, 5 (1):41–51, March 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Lee:1981:RFI

[Lee81a] Jong-Sen Lee. Refined filtering of image noise using local statistics. Computer Graphics and Image Processing, 15 (4):380–389, April 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Lee:1981:SAS

[Lee81b]  Jong Sen Lee.  Speckle

analysis and smoothing of synthetic aperture radar images. Computer Graphics and Image Processing, 17(1):24– 32, September 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Lee:1983:VSP

[Lee83a] D. T. Lee. Visibility of a simple polygon. Computer Vision, Graphics, and Image Processing, 22(2):207– 221, May 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See corrections [JS87].

Lee:1983:DIS

[Lee83b] Jong-Sen Lee. Digital image smoothing and the sigma filter. Computer Vision, Graphics, and Image Processing, 21(3):255–269, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lee:1986:RRS

[Lee86] Chin-Hwa Lee. Recursive region splitting at hierarchical scope views. Computer Vision, Graphics, and Image Processing, 33(2):237– 258, February 1986. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lee:1991:TVI

[Lee91]  Chia-Hoang Lee.  Time-

varying images: The effect of finite resolution on uniqueness. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (3):325–332, November 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Lee:2002:MVS

[Lee02] Joon Woong Lee. A machine vision system for lanedeparture detection. Computer Vision and Image Understanding: CVIU, 86(1): 52–78, April 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lemkin:1979:ARS

[Lem79a] Peter Lemkin. Approach to region splitting. Computer Graphics and Image Processing, 10(3):281–288, July 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Lemkin:1979:BTT

[Lem79b] Peter Lemkin. Boundary trace transform: an edge and region enhancement transform. Computer Graphics and Image Processing, 9(2):150– 165, February 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Latecki:1995:WCS

[LER95] Longin Latecki, Ulrich Eckhardt, and Azriel Rosenfeld. Well-composed sets. Computer Vision and Image Understanding: CVIU, 61(1):70–83, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.

                                        com/links/artid/cviu.1995.
                                        1006/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1995.1006/production/
                                        pdf.

Leu:1992:ICE

[Leu92] Jia-Guu Leu. Image contrast enhancement based on the intensities of edge pixels. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(6):497–506, November 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Levin:1979:MMD

[Lev79] Joshua Zev Levin. Mathematical models for determining the intersections of quadric surfaces. Computer Graphics and Image Processing, 11(1): 73–87, September 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Leyton:1985:GSA

[Ley85] Michael Leyton. Generative systems of analyzers. Computer Vision, Graphics, and Image Processing, 31 (2):201–241, August 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Leyton:1987:NSC

[Ley87a] Michael Leyton. Nested structures of control: an intuitive view. Computer Vision, Graphics, and Image Processing, 37(1):20–53, January 1987. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Leyton:1987:SCD

[Ley87b] Michael Leyton. Symmetrycurvature duality. Computer Vision, Graphics, and Image Processing, 38(3):327– 341, June 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lu:1979:STG

[LF79] S. Y. Lu and K. S. Fu.

Stochastic tree grammar inference for texture synthesis and discrimination. Computer Graphics and Image Processing, 9(3):234–245, March 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Lee:1982:UFD

[LF82] H. C. Lee and K. S. Fu.

Using the FFT to determine digital straight line chain codes. Computer Graphics and Image Processing, 18(4): 359–368, April 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Lee:1983:DSC

[LF83] H. C. Lee and K. S. Fu. 3-D shape from contour and selective confirmation. Computer Vision, Graphics, and Image Processing, 22(1):177– 193, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lejeune:1996:FPO

[LF96] André Lejeune and Frank P. Ferrie. Finding the parts of objects in range images. Computer Vision and Image Understanding: CVIU, 64(2):230–247, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0056/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0056/production/
pdf.

Luong:1998:DEU

[LF98] Q.-T. Luong and O. D. Faugeras. On the determination of epipoles using cross-ratios. Computer Vision and Image Understanding: CVIU, 71(1):1–18, July 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0621/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0621/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0621/production/ref.

Lawrence:2004:PII

[LF04] Jason Lawrence and Thomas Funkhouser. A painting interface for interactive surface deformations. Graphical Models, 66(6):418–438, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Lagger:2008:RML

[LF08] Pascal Lagger and Pascal Fua. Retrieving multiple light sources in the presence of specular reflections and texture. Computer Vision and Image Understanding: CVIU, 111(2):207–218, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2008:FBM

[LFL08] Zheng Liu, David S. Forsyth, and Robert Laganière. A feature-based metric for the quantitative evaluation of pixel-level image fusion. Computer Vision and Image Understanding: CVIU, 109(1): 56–68, January 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2013:SAB

[LFMP13] Yooyoung Lee, James J. Filliben, Ross J. Micheals, and P. Jonathon Phillips. Sensitivity analysis for biometric systems: a methodology based on orthogonal experiment designs. Computer Vision and Image Understanding: CVIU, 117(5):532–550, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000040.

Li:1982:NQR

[LGJ82] M. Li, W. I. Grosky, and R. Jain. Normalized quadtrees with respect to translations. Computer Graphics and Image Processing, 20(1):72– 81, September 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Liu:1984:NAR

[LH84] Z. K. Liu and B. R. Hunt. A new approach to removing cloud cover from satellite imagery. Computer Vision, Graphics, and Image Processing, 25(2):252– 256, February 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liu:1988:ERB

[LH88a] Yuncai Liu and Thomas S. Huang. Estimation of rigid body motion using straight line correspondences. Computer Vision, Graphics, and Image Processing, 43(1):37– 52, July 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liu:1988:LAM

[LH88b] Yuncai Liu and Thomas S.

Huang. A linear algorithm for motion estimation using straight line correspondences. Computer Vision, Graphics, and Image Processing, 44(1):35–57, October 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lee:1990:FPC

[LH90] Chia-Hoang Lee and T. Huang.

Finding point correspondences and determining motion of a rigid object from two weak perspective views. Computer Vision, Graphics, and Image Processing, 52 (3):309–327, December 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Li:1992:EPM

[LH92] Ze-Nian Li and Gongzhu Hu.

On edge preservation in multiresolution images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(6): 461–472, November 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Levkowitz:1993:GGL

[LH93] Haim Levkowitz and Gabor T. Herman. GLHS: a generalized lightness, hue, and saturation color model.

Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(4):271–285, July 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1019/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1019/production/
pdf.

Lavalle:1995:FCP

[LH95] Steven M. Lavalle and Seth A. Hutchinson. Framework for constructing probability distributions on the space of image segmentations. Computer Vision and Image Understanding: CVIU, 61(2):203–230, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1016/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1016/production/
pdf.

Lee:1999:OCC

[LH99] Yu-Hua Lee and Shi-Jinn Horng. Optimal computing the chessboard distance transform on parallel processing systems. Computer Vision and Image Understanding: CVIU, 73(3):374–390, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0741/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0741/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0741/production/ref.

Luo:2003:UFA

[LH03] Bin Luo and E. R. Hancock. A unified framework for alignment and correspondence. Computer Vision and Image Understanding: CVIU, 92(1):26–55, October 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lloyd:1987:PBS

[LHB87] S. A. Lloyd, E. R. Haddow, and J. F. Boyce. A parallel binocular stereo algorithm utilizing dynamic programming and relaxation labelling. Computer Vision, Graphics, and Image Processing, 39 (2):202–225, August 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Little:1997:DIR

[LHH97] J. A. Little, D. L. G. Hill, and D. J. Hawkes. Deformations incorporating rigid structures. Computer Vision and Image Understanding: CVIU, 66(2):223–232, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0608/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0608/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0608/production/ref.

Liu:1998:AVE

[LHH^{+}98] Hongche Liu, Tsai-Hong Hong, Martin Herman, Ted Camus, and Rama Chellappa. Accuracy vs efficiency trade-offs in optical flow algorithms. Computer Vision and Image Understanding: CVIU, 72(3):271–286, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0675/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0675/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0675/production/ref.

Liu:1998:MMB

[LHHC98] Hongche Liu, Tsai-Hong Hong, Martin Herman, and Rama Chellappa. Motionmodel-based boundary extraction and a real-time implementation. Computer Vision and Image Understanding: CVIU, 70(1):87–100, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0625/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0625/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0625/production/ref.

Liu:2009:FFS

[LHJ^{+}09] Chunxi Liu, Qingming Huang, Shuqiang Jiang, Liyuan Xing, Qixiang Ye, and Wen Gao. A framework for flexible summarization of racquet sports video using multiple modalities. Computer Vision and Image Understanding: CVIU, 113(3):415–424, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:1997:PCE

[LHKC97] Yu-Hua Lee, Shi-Jinn Horng, Tzong-Wann Kao, and Yuung-Jih Chen. Parallel computation of the Euclidean distance transform on the mesh of trees and the hypercube computer. Computer Vision and Image Understanding: CVIU, 68(1):109–119, October 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0539/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0539/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/cviu.1997.
                                        0539/production/ref.

Laganiere:2006:VRG

[LHM06] R. Laganière, H. Hajjdiab, and A. Mitiche. Visual reconstruction of ground plane obstacles in a sparse view robot environment. Graphical Models, 68(3):282–293, May 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070306000166.

Lensch:2001:SBA

[LHS01] Hendrik P. A. Lensch, Wolfgang Heidrich, and Hans-Peter Seidel. A silhouettebased algorithm for texture registration and stitching. Graphical Models, 63(4):245– 262, July 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Lhuillier:2008:ASS

[Lhu08]  Maxime Lhuillier.  Auto-

matic scene structure and camera motion using a catadioptric system. Computer Vision and Image Understanding: CVIU, 109(2):186– 203, February 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2005:VTR

[LHYK05] Kuang-Chih Lee, Jeffrey Ho, Ming-Hsuan Yang, and David Kriegman. Visual tracking and recognition using probabilistic appearance manifolds. Computer Vision and Image Understanding: CVIU, 99 (3):303–331, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:1992:TVR

[Li92] S. Z. Li. Toward 3D vision from range images: an optimization framework and parallel networks. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(3):231–260, May 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Lee:2000:ESS

[LI00] Rae Kyoung Lee and Insung Ihm. On enhancing the speed of splatting using both object- and imagespace coherence. Graphical Models, 62(4):263–282, July 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0524;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0524/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0524/ref.

Lillekjendlie:1997:CAF

[Lil97] Bjørn Lillekjendlie. Circular arcs fitted on a Riemann sphere. Computer Vision and Image Understanding: CVIU, 67(3):311–317, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0529/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0529/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0529/production/ref.

Lingrand:2002:ESP

[Lin02] Diane Lingrand. An exhaustive study of particular cases leading to robust and accurate motion estimation. Computer Vision and Image Understanding: CVIU, 85(3): 159–188, March 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liow:1991:CTA

[Lio91]  Yuh-Tay Liow.  A con-

tour tracing algorithm that preserves common boundaries between regions. Computer Vision, Graphics, and Image Processing. Image Understanding, 53(3):313–321, May 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Liu:1977:TTD

[Liu77] Hsun K. Liu. Two- and threedimensional boundary detection. Computer Graphics and Image Processing, 6(2): 123–134, April 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Liu:1993:CGN

[Liu93] Yong-Kui K. Liu. Comment on “Generation of Noise in Binary Images”. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(2):160, March 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1011/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1011/production/

pdf. See [ZG91].

Liu:1997:NTS

[Liu97] C. P. Liu. A new two successive process image compression technique using subband coding and JPEG discrete cosine transform coding. Graphical Models and Image Processing: GMIP, 59(3):179–191, May 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0430/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0430/production/
pdf.

Liu:2010:FFS

[Liu10] Yonghuai Liu. Free form shape registration using the barrier method. Computer Vision and Image Understanding: CVIU, 114(9):1004– 1016, September 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liou:1987:RFU

[LJ87] Shih-Ping P. Liou and Ramesh C. Jain. Road following using vanishing points. Computer Vision, Graphics, and Image Processing, 39(1):116– 130, July 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liou:1989:MDS

[LJ89] Shih-Ping P. Liou and Ramesh C. Jain. Motion detection in spatio-temporal space. Computer Vision, Graphics, and Image Processing, 45(2):227– 250, February 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liu:1990:TAD

[LJ90] Song-Sheng Liu and M. E. Jernigan. Texture analysis and discrimination in additive noise. Computer Vision, Graphics, and Image Processing, 49(1):52–67, January 1990. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liou:1991:ATD

[LJ91] Shih-Ping Liou and Ramesh C.

Jain. An approach to three dimensional image segmentation. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (3):237–252, May 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Lee:2007:VAB

[LJHH07] Chang-Ock Lee, Kiwan Jeon, Youngsoo Ha, and Jooyoung Hahn. A variational approach to blending based on warping for non-overlapped images. Computer Vision and Image Understanding: CVIU, 105 (2):112–120, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:1991:KFA

[LK91] Sukhan Lee and Youngchul Kay. A Kalman filter approach for accurate 3-D motion estimation from a sequence of stereo images. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(2): 244–258, September 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Lee:1994:SSP

[LK94] Kyoung Mu Lee and C.- C. Jay Kuo. Shape from shading with perspective projection. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):202–212, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1013/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1013/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1015/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1015/production/
pdf.

Lee:1997:SSG

[LK97] Kyoung Mu Lee and C.-C. Jay Kuo. Shape from shading with a generalized reflectance map model. Computer Vision and Image Understanding: CVIU, 67(2):143–160, August 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0522/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0522/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0522/production/ref.

Lorenz:2000:GPB

[LK00] Cristian Lorenz and Nils Krahnstöver. Generation of point-based 3D statistical shape models for anatomical objects. Computer Vision and Image Understanding: CVIU, 77(2):175–191, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0814;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0814/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0814/ref.

Lee:2001:NHM

[LK01] Doo-Won Lee and Hyeong-Seok Ko. Natural hairstyle modeling and animation. Graphical Models, 63(2):67– 85, March 2001. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0547;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0547/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0547/ref.

Liu:2003:FVS

[LK03] Yan Liu and John R. Kender.

Fast video segment retrieval by Sort-Merge feature selection, boundary refinement, and lazy evaluation. Computer Vision and Image Understanding: CVIU, 92(2–3):147–175, November/ December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:1994:BSM

[LKC94] Ta-Chih C. Lee, R. L. Kashyap, and Chong-Nam N. Chu. Building skeleton models via 3D medial surface/axis thinning algorithms. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(6):462–478, November 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1042/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1042/production/

pdf. See note [BM95].

Lee:1998:PRA

[LKE98] In-Kwon Lee, Myung-Soo Kim, and Gershon Elber. Polynomial/rational approximation of Minkowski sum boundary curves. Graphical Models and Image Processing: GMIP, 60(2):136–165, March 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0464/production;   http:

//www.idealibrary.com/links/
artid/gmip.1998.0464/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0464/production/ref.

Lurig:2000:HSD

[LKE00] Christoph Lürig, Leif Kobbelt, and Thomas Ertl. Hierarchical solutions for the deformable surface problem in visualization. Graphical Models, 62(1):2–18, January 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/artid/gmod.1999.
0515/production;   http:
//www.idealibrary.com/links/
artid/gmod.1999.0515/production/
pdf;   http://www.idealibrary.
com/links/artid/gmod.1999.
0515/production/ref.

Lu:2000:SST

[LKK00] Rong Lu, Jan J. Koenderink, and Astrid M. L. Kappers. Specularities on surfaces with tangential hairs or grooves. Computer Vision and Image Understanding: CVIU, 78(3):320–335, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0841;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0841/
pdf;   http://www.idealibrary.

                                        com/links/doi/10.1006/cviu.
                                        2000.0841/ref.

Lee:1986:CVP

[LL86] D. T. Lee and A. K. Lin. Computing the visibility polygon from an edge. Computer Vision, Graphics, and Image Processing, 34(1):1– 19, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Leymarie:1992:FRS

[LL92] F. Leymarie and Martin D. Levine. Fast raster scan distance propagation on the discrete rectangular lattice. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(1):84– 94, January 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Li:1995:ISU

[LL95] C. H. Li and C. K. Lee. Image smoothing using parametric relaxation. Graphical Models and Image Processing: GMIP, 57(2):161–174, March 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1016/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1016/production/
                                        pdf.

Lin:1997:DSU

[LL97a] Stephen Lin and Sang Wook Lee. Detection of specularity using stereo in color and polarization space. Computer Vision and Image Understanding: CVIU, 65 (2):336–346, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0577/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0577/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0577/production/ref.

Lindeberg:1997:SCE

[LL97b] Tony Lindeberg and Meng-Xiang Li. Segmentation and classification of edges using minimum description length approximation and complementary junction cues. Computer Vision and Image Understanding: CVIU, 67(1): 88–98, July 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1996.0510/production;
 http://www.idealibrary.
com/links/artid/cviu.1996.
0510/production/pdf;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0510/production/
ref.

Leung:1998:MSI

[LL98] C. K. Leung and E. K. Lam. Maximum segmented image information thresholding. Graphical Models and Image Processing: GMIP, 60(1):057–076, January 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0455/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0455/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0455/production/ref.

Latecki:1999:CRS

[LL99] Longin Jan Latecki and Rolf Lakämper. Convexity rule for shape decomposition based on discrete contour evolution. Computer Vision and Image Understanding: CVIU, 73(3):441–454, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0738/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0738/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0738/production/ref.

Leow:2004:AAA

[LL04] Wee Kheng Leow and Rui Li. The analysis and applications of adaptive-binning color histograms. Computer Vision and Image Understanding: CVIU, 94(1–3):67–91, April/ June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2006:PAB

[LL06] Jehee Lee and Kang Hoon Lee. Precomputing avatar behavior from human motion data. Graphical Models, 68(2):158–174, March 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000275.

Liu:2008:FCD

[LL08] Zhiming Liu and Chengjun Liu. Fusion of the complementary Discrete Cosine Features in the YIQ color space for face recognition. Computer Vision and Image Understanding: CVIU, 111 (3):249–262, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Linde:2012:CCC

[LL12] Oskar Linde and Tony Lindeberg. Composed complexcue histograms: An investigation of the information content in receptive field based image descriptors for object recognition. Computer Vision and Image Understanding: CVIU, 116(4):538–560, April 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002633.

Li:2013:GCC

[LL13] Guo Li and Ligang Liu. Geometry curves: a compact representation for 3D shapes. Graphical Models, 75 (5):265–278, September 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000179.

Liu:2011:TBL

[LLC11] Yang Liu, Yan Liu, and Keith C. C. Chan. Tensor-based locally maximum margin classifier for image and video classification. Computer Vision and Image Understanding: CVIU, 115(3):300–309, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2012:ORU

[LLC12] Ying-Ho Liu, Anthony J. T. Lee, and Fu Chang. Object recognition using discriminative parts. Computer Vision and Image Understanding: CVIU, 116(7):854–867, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000604.

Li:2013:HER

[LLC13] Yujian Li, Houjun Li, and Zhi Cai. Human eyebrow recognition in the matchingrecognizing framework. Computer Vision and Image Understanding: CVIU, 117 (2):170–181, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001452.

Liu:2009:ICV

[LLE^{+}09] Qingshan Liu, Xuelong Li, Ahmed Elgammal, Xian sheng Hua, Dong Xu, and Dacheng Tao. Introduction to Computer Vision and Image Understanding: the special issue on video analysis. Computer Vision and Image Understanding: CVIU, 113(3): 317–318, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Li:1986:FHT

[LLL86] Hungwen Li, Mark A. Lavin, and Ronald J. Le Master. Fast Hough transform: a hierarchical approach. Computer Vision, Graphics, and Image Processing, 36(2/3):139–161, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Li:2013:SPN

[LLL13] Zechao Li, Jing Liu, and Hanqing Lu. Structure preserving non-negative matrix factorization for dimensionality reduction. Computer Vision and Image Understanding: CVIU, 117(9):1175–1189, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000891.

Leichter:2010:MST

[LLR10] Ido Leichter, Michael Lindenbaum, and Ehud Rivlin. Mean Shift tracking with multiple reference color histograms. Computer Vision and Image Understanding: CVIU, 114 (3):400–408, March 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lopez:2000:MCB

[LLSV00] Antonio M. López, David Lloret, Joan Serrat, and Juan J. Villanueva. Multilocal creaseness based on the level-set extrinsic curvature. Computer Vision and Image Understanding: CVIU, 77(2):111–144, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0812;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0812/

pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0812/ref.

Lin:2013:NIS

[LLXW13] Shujin Lin, Xiaonan Luo, Songhua Xu, and Jianmin Wang. A new interpolation subdivision scheme for triangle/quad mesh. Graphical Models, 75(5):247–254, September 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031300012X.

Laughlin:1989:GSA

[LM89] Daniel L. Laughlin and Manfred Morari. Graphical stability analysis for control systems with model parameter uncertainties. Computer Vision, Graphics, and Image Processing, 47(1):59– 76, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lattard:1991:VIP

[LM91] D. Lattard and G. Mazare.

A VLSI implementation of parallel image reconstruction. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6):581– 591, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Liao:1995:SAC

[LM95] Chia-Wei W. Liao and Gerard Medioni. Surface approximation of a cloud of 3D points. Graphical Models and Image Processing: GMIP, 57(1):67–74, January 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1007/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1007/production/
                                        pdf.

Latecki:1996:AST

[LM96] Longin Latecki and C. Min Ma. An algorithm for a 3D simplicity test. Computer Vision and Image Understanding: CVIU, 63 (2):388–393, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0028/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0028/production/
                                        pdf.

Lee:1999:GLR

[LM99a] Mi-Suen Lee and Gérard Medioni. Grouping … into regions, curves, and junctions. Computer Vision and Image Understanding: CVIU, 76(1):54–69, October 1999. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0787/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0787/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0787/production/ref.

Liao:1999:SSA

[LM99b] Chia-Wei Liao and Gérard Medioni. Simultaneous surface approximation and segmentation of complex objects. Computer Vision and Image Understanding: CVIU, 73(1):43–63, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0694/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0694/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0694/production/ref.

Lachaud:2000:CAD

[LM00] Jacques-Olivier Lachaud and Annick Montanvert. Continuous analogs of digital boundaries: a topological approach to iso-surfaces. Graphical Models, 62(3):129–164, May 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.

                                        2000.0522;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/gmod.2000.0522/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2000.0522/ref.

Lai:2012:VLO

[LM12] Yu-Kun Lai and Ralph R. Martin. Vertex location optimisation for improved remeshing. Graphical Models, 74(4):233–243, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000318.

Li:2009:AER

[LMC09] Shiying Li, Yoshitsugu Manabe, and Kunihiro Chihara. Accurately estimating reflectance parameters for color and gloss reproduction. Computer Vision and Image Understanding: CVIU, 113(2): 308–316, February 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lopez-Molina:2011:GFE

[LMDB11] C. Lopez-Molina, B. De Baets, and H. Bustince. Generating fuzzy edge images from gradient magnitudes. Computer Vision and Image Understanding: CVIU, 115(11):1571–1580, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001688.

Lauzon:1985:TDR

[LMKG85] Jean Paul Lauzon, David M.

Mark, Lawrence Kikuchi, and J. Armando Guevara. Twodimensional run-encoding for quadtree representation. Computer Vision, Graphics, and Image Processing, 30(1): 56–69, April 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Li:1995:MIF

[LMM95] H. Li, B. S. Manjunath, and S. K. Mitra. Multisensor image fusion using the wavelet transform. Graphical Models and Image Processing: GMIP, 57(3):235–245, May 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1022/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1022/production/
pdf.

Lane:1984:EPP

[LMR84] Jeff Lane, Bob Magedson, and Mike Rarick. An efficient point in polyhedron algorithm. Computer Vision, Graphics, and Image Processing, 26(1):118– 125, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lazarevic-McManus:2008:OBC

[LMRMJ08] N. Lazarevic-McManus, J. R. Renno, D. Makris, and G. A. Jones. An object-based comparative methodology for motion detection based on the F -measure. Computer Vision and Image Understanding: CVIU, 111(1):74–85, July 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Levine:1985:RBI

[LN85] Martin D. Levine and Ahmed M. Nazif. Rule-based image segmentation: a dynamic control strategy approach. Computer Vision, Graphics, and Image Processing, 32 (1):104–126, October 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lin:1998:BDD

[LN98] Chungan Lin and Ramakant Nevatia. Building detection and description from a single intensity image. Computer Vision and Image Understanding: CVIU, 72(2): 101–121, November 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0724/production;   http:
//www.idealibrary.com/links/

artid/cviu.1998.0724/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0724/production/ref.

Lu:2010:LRC

[LN10] Xin Lu and Kiyoshi Nishiyama.

Low-resolution color-based visual tracking with state-space model identification. Computer Vision and Image Understanding: CVIU, 114(9): 1045–1054, September 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Levine:1983:UBC

[LNY83] Martin D. Levine, Peter B.

Noble, and Youssry M. Youssef. Understanding blood cell motion. Computer Vision, Graphics, and Image Processing, 21(1):58–84, January 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lohou:2010:DNT

[Loh10] Christophe Lohou. Detection of the non-topology preservation of Ma and Sonka’s algorithm, by the use of P -simple points. Computer Vision and Image Understanding: CVIU, 114(3):384–399, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lorre:1983:HDA

[Lor83] Jean J. Lorre. Histogram deconvolution: an aid to automated classifiers. Computer Vision, Graphics, and Image Processing, 23(3):334– 340, September 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lozano-Perez:1977:PIP

[LP77] Tomas Lozano-Perez. Parsing intensity profiles. Computer Graphics and Image Processing, 6(1):43–60, February 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Little:1979:RPF

[LP79] J. J. Little and T. K. Peucker. A recursive procedure for finding the intersection of two digital curves. Computer Graphics and Image Processing, 10 (2):159–171, June 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Lin:1990:EHH

[LP90a] Wei-Ming Lin and V. K. Prasanna Kumar. Efficient histogramming on hypercube SIMD machines. Computer Vision, Graphics, and Image Processing, 49(1):104– 120, January 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liow:1990:USE

[LP90b] Yuh-Tay Liow and Theo Pavlidis. Use of shadows for extracting buildings in aerial images. Computer Vision, Graphics, and Image Processing, 49(2):242– 277, February 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lee:1991:WEN

[LP91] Chung-Nim Lee and T. Poston. Winding and Euler numbers for 2D and 3D digital images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6): 522–537, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Lu:2010:SAC

[LP10] Yan Lu and Shahram Payandeh. On the sensitivity analysis of camera calibration from images of spheres. Computer Vision and Image Understanding: CVIU, 114(1): 8–20, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Landabaso:2008:SIS

[LPC08] Jose-Luis Landabaso, Montse Pardàs, and Josep Ramon Casas. Shape from inconsistent silhouette. Computer Vision and Image Understanding: CVIU, 112(2): 210–224, November 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Levialdi:1978:IAD

[LPF78] S. Levialdi, A. Pirri, and V. Franchina. Image acquisition device for minicomputers. Computer Graphics and Image Processing, 8(1): 113–120, August 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Liang:2001:PED

[LPH01] Jisheng Liang, Ihsin T. Phillips, and Robert M. Haralick. Performance evaluation of document structure extraction algorithms. Computer Vision and Image Understanding: CVIU, 84(1):144–159, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0933;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0933/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0933/ref.

Lavakusha:1989:LOV

[LPR89] Lavakusha, Arun K. Pujari, and P. G. Reddy. Linear octrees by volume intersection.

Computer Vision, Graphics, and Image Processing, 45(3):371–379, March 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lee:1993:HGD

[LPR93] Chung-Nim N. Lee, Timothy Poston, and Azriel Rosenfeld. Holes and genus of 2D and 3D digital images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(1):20–47, January 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1002/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1002/production/
pdf.

Lunn:2003:NBR

[LPR^{+}03] Karen E. Lunn, Keith D.

Paulsen, David W. Roberts, Francis E. Kennedy, Alex Hartov, and Leah A. Platenik. Nonrigid brain registration: synthesizing full volume deformation fields from model basis solutions constrained by partial volume intraoperative data. Computer Vision and Image Understanding: CVIU, 89(2–3):299–317, February/March 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2011:OLS

[LPS^{+}11] Soochahn Lee, Sang Hyun Park, Hackjoon Shim, Il Dong Yun, and Sang Uk Lee. Optimization of local shape and appearance probabilities for segmentation of knee cartilage in 3-D MR images. Computer Vision and Image Understanding: CVIU, 115(12): 1710–1720, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001639.

Lukac:2007:CIP

[LPV07] Rastislav Lukac, Konstantinos N. Plataniotis, and Anastasios N. Venetsanopoulos. Color image processing. Computer Vision and Image Understanding: CVIU, 107(1–2): 1–2, July/August 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Losson:2013:CTA

[LPVM13] O. Losson, A. Porebski, N. Vandenbroucke, and L. Macaire. Color texture analysis using CFA chromatic co-occurrence matrices. Computer Vision and Image Understanding: CVIU, 117(7):747–763, July 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000532.

Liao:2008:FBR

[LPZ08] Iman Yi Liao, Maria Petrou, and Rongchun Zhao. A fractal-based relaxation algorithm for shape from terrain image. Computer Vision and Image Understanding: CVIU, 109(3):227–243, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:1990:SCL

[LR90] Chung-Nim Lee and Azriel Rosenfeld. Simple connectivity is not locally computable for connected 3D images. Computer Vision, Graphics, and Image Processing, 51(1):87–95, July 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Latecki:2002:RPN

[LR02] Longin Jan Latecki and Azriel Rosenfeld. Recovering a polygon from noisy data. Computer Vision and Image Understanding: CVIU, 86(1): 32–51, April 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lehmann:2012:NCT

[LR12] Nicole Lehmann and Ulrich Reif. Notes on the curvature tensor. Graphical Models, 74(6):321–325, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000239.

Lerasle:1999:THL

[LRD99] F. Lerasle, G. Rives, and M. Dhome. Tracking of human limbs by multiocular vision. Computer Vision and Image Understanding: CVIU, 75(3):229–246, September 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0759/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0759/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0759/production/ref.

Lopez-Rubio:2011:SAB

[LRLB11] Ezequiel López-Rubio and Rafael Marcos Luque-Baena. Stochastic approximation for background modelling. Computer Vision and Image Understanding: CVIU, 115(6): 735–749, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lai:2008:RFR

[LRW08] Hung Lai, Venkatesh Ramanathan, and Harry Wechsler. Reliable face recognition using adaptive and robust correlation filters. Computer Vision and Image Understanding: CVIU, 111(3):

329–350, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lattanzi:1991:OBQ

[LS91] M. R. Lattanzi and C. A.

Shaffer. An optimal boundary to quadtree conversion algorithm. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (3):303–312, May 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Li:1992:PTT

[LS92] Bing-Cheng Li and Jun Shen.

Pascal triangle transform approach to the calculation of 3D moments. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4): 301–307, July 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Landau:1994:SWR

[LS94] P. Landau and E. Schwartz.

Subset warping: Rubber sheeting with cuts. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(3):247–266, May 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.

                                        1022/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1022/production/
                                        pdf.

Lee:2001:CIA

[LS01] Jehee Lee and Sung Yong Shin. A coordinate-invariant approach to multiresolution motion analysis. Graphical Models, 63(2):87–105, March 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2001.0548;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/gmod.2001.0548/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2001.0548/ref.

Li:2008:MSS

[LS08] Rui Li and Stan Sclaroff. Multi-scale 3D scene flow from binocular stereo sequences. Computer Vision and Image Understanding: CVIU, 110(1):75–90, April 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lo:2009:LFE

[LS09] Tsz-Wai Rachel Lo and J. Paul Siebert. Local feature extraction and matching on range images: 2.5D SIFT. Computer Vision and Image Understanding: CVIU, 113(12):1235–1250, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2012:SAI

[LS12] Lingyun Liu and Ioannis Stamos. A systematic approach for 2D-image to 3D-range registration in urban environments. Computer Vision and Image Understanding: CVIU, 116(1):25–37, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001809.

Lelieveldt:2000:AMF

[LSB^{+}00] Boudewijn P. F. Lelieveldt, Milan Sonka, Lizann Bolinger, Thomas D. Scholz, Hein Kayser, Rob van der Geest, and Johan H. C. Reiber. Anatomical modeling with fuzzy implicit surface templates: Application to automated localization of the heart and lungs in thoracic MR volumes. Computer Vision and Image Understanding: CVIU, 80(1):1–20, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0864;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0864/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0864/ref.

Li:1992:HMS

[LSBG92] Z. C. Li, C. Y. Suen, T. D. Bui, and Q. L. Gu. Harmonic models of shape transformations in digital images and pictures. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(3): 198–209, May 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

LeBesnerais:2008:DHM

[LSC08] G. Le Besnerais, M. Sanfourche, and F. Champagnat. Dense height map estimation from oblique aerial image sequences. Computer Vision and Image Understanding: CVIU, 109(2):204– 225, February 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2003:FAQ

[LSCM03] Yanxi Liu, Karen L. Schmidt, Jeffrey F. Cohn, and Sinjini Mitra. Facial asymmetry quantification for expression invariant human identification. Computer Vision and Image Understanding: CVIU, 91(1–2):138–159, July/August 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Larsen:2007:TEA

[LSD^{+}07] Rasmus Larsen, Mikkel B.

Stegmann, Sune Darkner, Søren Forchhammer, Timothy F. Cootes, and Bjarne Kjær Ersbøll. Texture enhanced appearance models. Computer Vision and Image Understanding: CVIU, 106(1): 20–30, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lin:2002:NAA

[LSHT02] Chung-Yi Lin, Sheng-Wen Shih, Yi-Ping Hung, and Gregory Y. Tang. A new approach to automatic reconstruction of a 3-D world using active stereo vision. Computer Vision and Image Understanding: CVIU, 85(2): 117–143, February 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lindner:2010:TFS

[LSKK10] Marvin Lindner, Ingo Schiller, Andreas Kolb, and Reinhard Koch. Time-of-flight sensor calibration for accurate range sensing. Computer Vision and Image Understanding: CVIU, 114(12):1318–1328, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lin:1985:EDU

[LSMS85] Xinggang Lin, Shigeyoshi Shimotsuji, Michihiko Minoh, and Toshiyuki Sakai. Efficient diagram understanding with characteristic pattern detection. Computer Vision, Graphics, and Image Processing, 30(1):84–106, April 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lukac:2004:SWV

[LSPV04] Rastislav Lukac, Bogdan Smolka, Konstantinos N. Plataniotis, and Anastasios N. Venetsanopoulos. Selection weighted vector directional filters. Computer Vision and Image Understanding: CVIU, 94(1–3):140–167, April/June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2013:PFO

[LST13] Jehoon Lee, Romeil Sandhu, and Allen Tannenbaum. Particle filters and occlusion handling for rigid 2D–3D pose tracking. Computer Vision and Image Understanding: CVIU, 117(8):922–933, August 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421300088X.

Lopez-Sastre:2011:TMD

[LSTARMB11] R. J. López-Sastre, T. Tuytelaars, F. J. Acevedo-Rodríguez, and S. Maldonado-Bascón. Towards a more discriminative and semantic visual vocabulary. Computer Vision and Image Understanding: CVIU, 115(3):415–425, March 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Livne:2012:HAP

[LSTF12] Micha Livne, Leonid Sigal, Nikolaus F. Troje, and David J. Fleet. Human attributes from 3D pose tracking. Computer Vision and Image Understanding: CVIU, 116(5):648–660, May 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000203.

Lashas:1985:OCR

[LSVD85] A. Lashas, R. Shurna, A. Verikas, and A. Dosinas. Optical character recognition based on analog preprocessing and automatic feature extraction. Computer Vision, Graphics, and Image Processing, 32 (2):191–207, November 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lumia:1983:NCC

[LSZ83] Ronald Lumia, Linda Shapiro, and Oscar Zuniga. New connected components algorithm for virtual memory computers. Computer Vision, Graphics, and Image Processing, 22(2):287–300, May 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Levine:1981:ILP

[LT81] Martin D. Levine and David Ting. Intermediate level picture interpretation using complete two-dimensional models. Computer Graphics and Image Processing, 16(3): 185–209, July 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Liang:1990:RRS

[LT90a] Ping Liang and John S. Todhunter. Representation and recognition of surface shapes in range images. A differential geometry approach. Computer Vision, Graphics, and Image Processing, 52 (1):78–109, October 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liu:1990:COR

[LT90b] Cheng-Hsiung Liu and Wen-Hsiang Tsai. 3D curved object recognition from multiple 2D camera views. Computer Vision, Graphics, and Image Processing, 50(2):177– 187, May 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Luettin:1997:SUP

[LT97] Juergen Luettin and Neil A. Thacker. Speechreading using probabilistic models. Computer Vision and Image Understanding: CVIU, 65 (2):163–178, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0570/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0570/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0570/production/ref.

Lachaud:2005:DMC

[LT05] J.-O. Lachaud and B. Taton.

Deformable model with a complexity independent from image resolution. Computer Vision and Image Understanding: CVIU, 99(3):453– 475, September 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lu:1993:CCN

[LTS93] H. Q. Lu, J. S. Todhunter, and T. W. Sze. Congruence conditions for nonplanar developable surfaces and their application to surface recognition. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3):265–285, November 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.

                                        1042/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1993.1042/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1993.
                                        1044/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1993.1044/production/
                                        pdf.

Lim:1991:NMR

[LTT91] Hock Lim, Kah-Chye Tan, and B. T. G. Tan. New methods for restoring motionblurred images derived from edge error considerations. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):479– 490, September 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Lu:1978:SAT

[Lu78] S. Y. Lu. Syntactic approach to texture analysis. Computer Graphics and Image Processing, 17(3):303–330, June 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Lucchese:2001:FDT

[Luc01]  Luca Lucchese.  A fre-

quency domain technique based on energy radial projections for robust estimation of global 2D affine transformations. Computer Vision and Image Understanding: CVIU, 81(1):72–116, January 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0885;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0885/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0885/ref.

Lumia:1983:NTD

[Lum83] Ronald Lumia. New threedimensional connected components algorithm. Computer Vision, Graphics, and Image Processing, 23(2):207– 217, August 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Luong:1996:CRG

[LV96] Q.-T. Luong and T. Viéville.

Canonical representations for the geometries of multiple projective views. Computer Vision and Image Understanding: CVIU, 64(2): 193–229, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0055/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0055/production/
pdf.

Lachaud:2003:ICD

[LV03] Jacques-Olivier Lachaud and Anne Vialard. 10th International Conference on Discrete Geometry for Computer Imagery: Discrete topology and geometry for image and object representation. Graphical Models, 65(1–3):1, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

LeGuyader:2011:CSR

[LV11] Carole Le Guyader and Luminita A. Vese. A combined segmentation and registration framework with a nonlinear elasticity smoother. Computer Vision and Image Understanding: CVIU, 115(12):1689–1709, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001573.

Lluch:2004:MTS

[LVM04] J. Lluch, R. Vivó, and C. Monserrat. Modelling tree structures using a single polygonal mesh. Graphical Models, 66(2):89–101, March 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Lam:1997:HVU

[LVW97] C. P. Lam, S. Venkatesh, and G. A. W. West. Hypothesis verification using parametric models and active vision strategies. Computer Vision and Image Understanding:

CVIU, 68(2):209–236, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0554/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0554/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0554/production/ref.

Leu:1985:DSS

[LW85] Jia-Guu G. Leu and William G.

Wee. Detecting the spatial structure of natural textures based on shape analysis. Computer Vision, Graphics, and Image Processing, 31(1):67–88, July 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Liu:2008:SAB

[LWGP08] Shiguang Liu, Zhangye Wang, Zheng Gong, and Qunsheng Peng. Simulation of atmospheric binary mixtures based on two-fluid model. Graphical Models, 70(6):117– 124, November 2008. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000064.

Luo:2003:OBA

[LWH03] Ying Luo, Tzong-Der Wu, and Jenq-Neng Hwang. Objectbased analysis and interpretation of human motion in sports video sequences by dynamic Bayesian networks. Computer Vision and Image Understanding: CVIU, 92(2–3):196–216, November/ December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2012:SCW

[LWLS12] Junbin Liu, Tim Wark, Ruan Lakemond, and Sridha Sridharan. Self-calibration of wireless cameras with restricted degrees of freedom. Computer Vision and Image Understanding: CVIU, 116(10):1033–1046, October 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000902.

Liu:1988:RBT

[LX88] Z. K. Liu and J. Y. Xiao. Restoration of blurred TV picture caused by uniform linear motion. Computer Vision, Graphics, and Image Processing, 44(1):30–34, October 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Leung:1990:DSA

[LY90] Maylor K. Leung and Yee-Hong Yang. Dynamic strip algorithm in curve fitting. Computer Vision, Graphics, and Image Processing, 51 (2):146–165, August 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Lee:2005:LDI

[LY05] Joon Woong Lee and Un Kun Yi. A lane-departure identification based on LBPE, Hough transform, and linear regression. Computer Vision and Image Understanding: CVIU, 99(3):359– 383, September 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Levine:2006:FRS

[LY06] Martin David Levine and Yingfeng Yu. Face recognition subject to variations in facial expression, illumination and pose using correlation filters. Computer Vision and Image Understanding: CVIU, 104 (1):1–15, October 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lhuillier:2013:MSR

[LY13] Maxime Lhuillier and Shuda Yu. Manifold surface reconstruction of an environment from sparse Structurefrom-Motion data. Computer Vision and Image Understanding: CVIU, 117(11): 1628–1644, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213001525.

Luo:2013:CEH

[LYA13] Weilan Luo, Toshihiko Yamasaki, and Kiyoharu Aizawa. Cooperative estimation of human motion and surfaces using multiview videos. Computer Vision and Image Understanding: CVIU, 117(11): 1560–1574, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213001367.

Liu:2008:SSB

[LYCG08] Xin Liu, Hongxun Yao, Xilin Chen, and Wen Gao. Shape from silhouettes based on a centripetal pentahedron model. Graphical Models, 70 (6):133–148, November 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S152407030800009X.

Liu:2007:SSO

[LYG07] Xin Liu, Hongxun Yaoand, and Wen Gao. Shape from silhouette outlines using an adaptive dandelion model. Computer Vision and Image Understanding: CVIU, 105 (2):121–130, February 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:2012:CCS

[LYKL12] Sun-Young Lee, Jong-Chul Yoon, Ji-Yong Kwon, and In-Kwon Lee. CartoonModes: Cartoon stylization of video objects through modal analysis. Graphical Models, 74 (2):51–60, March 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000033.

Lee:2010:TCV

[LYL10] Sun-Young Lee, Jong-Chul Yoon, and In-Kwon Lee. Temporally coherent video matting. Graphical Models, 72 (3):25–33, May 2010. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070310000081.

Lynch:1981:RIE

[Lyn81] David K. Lynch. Range image enhancement via onedimensional spatial filtering. Computer Graphics and Image Processing, 15(2):194– 200, February 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Liu:2012:LSF

[LYSS12] Jingen Liu, Yang Yang, Imran Saleemi, and Mubarak Shah. Learning semantic features for action recognition via diffusion maps. Computer Vision and Image Understanding: CVIU, 116(3):361–377, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002116.

Langer:1997:CLI

[LZ97a] Michael S. Langer and Steven W. Zucker. Casting light on illumination: a computational model and dimensional analysis of sources. Computer Vision and Image Understanding: CVIU, 65 (2):322–335, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0574/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0574/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0574/production/ref.

Lopresti:1997:UCS

[LZ97b] Daniel Lopresti and Jiangying Zhou. Using consensus sequence voting to correct OCR errors. Computer Vision and Image Understanding: CVIU, 67(1):39–47, July 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0502/production;   http:
//www.idealibrary.com/links/

artid/cviu.1996.0502/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0502/production/ref.

Lian:2010:QPB

[LZLP10] Wei Lian, Lei Zhang, Yan Liang, and Quan Pan. A quadratic programming based cluster correspondence projection algorithm for fast point matching. Computer Vision and Image Understanding: CVIU, 114(3):322– 333, March 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Liu:2003:MRM

[LZWP03] Feng Liu, Yueting Zhuang, Fei Wu, and Yunhe Pan. 3D motion retrieval with motion index tree. Computer Vision and Image Understanding: CVIU, 92(2–3):265–284, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Martin:1978:DSA

[MA78] W. N. Martin and J. K.

Aggarwal. Dynamic scene analysis. Computer Graphics and Image Processing, 17 (3):356–374, June 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Mitiche:1983:CRS

[MA83] Amar Mitiche and J. K. Aggarwal. Contour registration by shape-specific points for shape matching. Computer Vision, Graphics, and Image Processing, 22(3):396– 408, June 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Magee:1984:DVP

[MA84] M. J. Magee and J. K. Aggarwal. Determining vanishing points from perspective images. Computer Vision, Graphics, and Image Processing, 26(2):256–267, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Magee:1985:UMI

[MA85] M. J. Magee and J. K. Aggarwal. Using multisensory images to derive the structure of three-dimensional objects — a review. Computer Vision, Graphics, and Image Processing, 32(2):145– 157, November 1985. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ma:1994:TPT

[Ma94] Cherng Min Ma. On topology preservation in 3D thinning. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):328–339, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1023/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1023/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1027/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1027/production/
pdf.

Ma:1996:CPS

[Ma96]  C. Min Ma.  Connectiv-

ity preservation of 3D 6subiteration thinning algorithms. Graphical Models and Image Processing: GMIP, 58(4):382–386, July 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0030/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0030/production/
pdf.

Maa:1994:IEB

[Maa94] Chia-Yiu Y. Maa. Identifying the existence of bar codes in compressed images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(4):352–356, July 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1032/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1032/production/
                                        pdf.

Maes:1990:DSL

[Mae90] M. Maes. Digitization of straight line segments closeness and convexity. Computer Vision, Graphics, and Image Processing, 52(2):297– 305, November 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Maitre:1976:DRN

[Mai76] Henri Maitre. Defect recognition in numerical images by spectrum zero detection. Computer Graphics and Image Processing, 5(2): 238–244, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Maitre:1981:IPR

[Mai81] Henri Maitre. Iterative picture restoration using video optical feedback. Computer Graphics and Image Processing, 16(2):95–115, June 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Michel:2010:HML

[MAL10] Damien Michel, Antonis A. Argyros, and Manolis I. A.

Lourakis. Horizon matching for localizing unordered panoramic images. Computer Vision and Image Understanding: CVIU, 114(2): 274–285, February 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Monga:1997:TNC

[MAM97] Olivier Monga, Nasser Armande, and Philippe Montesinos. Thin nets and crest lines: Application to satellite data and medical images. Computer Vision and Image Understanding: CVIU, 67(3):285–295, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0507/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0507/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0507/production/ref.

Mantyla:1984:NMS

[Man84a] Martti Mantyla. A note on the modeling space of Euler operators. Computer Vision, Graphics, and Image Processing, 26(1):45–60, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Matsuyama:1984:SML

[MAN84b] Takashi Matsuyama, Hidekazu Arita, and Makoto Nagao. Structural matching of line drawings using the geometric relationship between line segments. Computer Vision, Graphics, and Image Processing, 27(2):177–194, August 1984. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Mansbach:1986:CCL

[Man86] Peter Mansbach. Calibration of a camera and light source by fitting to a physical model. Computer Vision, Graphics, and Image Processing, 35 (2):200–219, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Mandal:1999:FWH

[MAP99] M. K. Mandal, T. Aboulnasr, and S. Panchanathan. Fast wavelet histogram techniques for image indexing. Computer Vision and Image Understanding: CVIU, 75(1–2):99–110, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0766/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0766/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0766/production/ref.

Marks:1980:LLV

[Mar80] Philip Marks. Low-level vision using an array processor. Computer Graphics and Image Processing, 14(3):281– 292, November 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Margosian:1982:SRM

[Mar82] Paul M. Margosian. Streak removal method for CT based digital X-ray images. Computer Graphics and Image Processing, 19(1):76–80, May 1982. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Marola:1989:USD

[Mar89] Giovanni Marola. Using symmetry for detecting and locating objects in a picture. Computer Vision, Graphics, and Image Processing, 46(2):179–195, May 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Martens:1990:DDI

[Mar90] Jean-Bernard Martens. Deblurring digital images by means of polynomial transforms. Computer Vision, Graphics, and Image Processing, 50(2):157–176, May 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Marroquin:1993:DIP

[Mar93] J. L. Marroquin. Deterministic interactive particle models for image processing and computer graphics. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(5):408–417, September 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1031/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1031/production/
pdf.

Markussen:2007:LDD

[Mar07] Bo Markussen. Large deformation diffeomorphisms with application to optic flow. Computer Vision and Image Understanding: CVIU, 106 (1):97–105, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mastin:1985:AFD

[Mas85]  Gary A. Mastin.  Adap-

tive filters for digital image noise smoothing: an evaluation. Computer Vision, Graphics, and Image Processing, 31(1):103–121, July 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Masuda:2002:RIM

[Mas02] Takeshi Masuda. Registration and integration of multiple range images by matching signed distance fields for object shape modeling. Computer Vision and Image Understanding: CVIU, 87(1– 3):51–65, July 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Masuda:2009:LPH

[Mas09]  Takeshi Masuda.  Log-

polar height maps for multiple range image registration. Computer Vision and Image Understanding: CVIU, 113(11):1158–1169, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Matsuyama:1989:ESI

[Mat89] Takashi Matsuyama. Expert systems for image processing: knowledge-based composition of image analysis processes. Computer Vision, Graphics, and Image Processing, 48(1):22–49, October 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Max:1984:ATS

[Max84]  Nelson L. Max.  Atoms

with transparency and shadows. Computer Vision, Graphics, and Image Processing, 27(1):46–63, July 1984.

CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Max:1986:LDT

[Max86] Nelson L. Max. Light diffusion through clouds and haze. Computer Vision, Graphics, and Image Processing, 33(3):280–292, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Max:1991:USS

[Max91] Nelson L. Max. Unified sun and sky illumination for shadows under trees. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(3): 223–230, May 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Maybank:1997:RPR

[May97] S. J. Maybank. Reply to pizlo, rosenfeld, and weiss. Computer Vision and Image Understanding: CVIU, 67(3):318–319, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0643/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0643/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0643/production/ref.

Mayer:1999:AOE

[May99] Helmut Mayer. Automatic object extraction from aerial imagery — a survey focusing on buildings. Computer Vision and Image Understanding: CVIU, 74(2):138–149, May 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0750/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0750/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0750/production/ref.

Meng:2010:MML

[MAY^{+}10] Hongying Meng, Kofi Appiah, Shigang Yue, Andrew Hunter, Mervyn Hobden, Nigel Priestley, Peter Hobden, and Cy Pettit. A modified model for the Lobula Giant Movement Detector and its FPGA implementation. Computer Vision and Image Understanding: CVIU, 114(11): 1238–1247, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mazumder:1987:PDQ

[Maz87] Pinaki Mazumder. Planar decomposition for quadtree data structure. Computer Vision, Graphics, and Image Processing, 38(3):258–274, June 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Maxwell:1979:GPR

[MB79] P. C. Maxwell and P. W. Baker. The generation of polygons representing circles, ellipses and hyperbolas. Computer Graphics and Image Processing, 10 (1):84–93, May 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Mitiche:1985:TMO

[MB85] Amar Mitiche and Patrick Bouthemy. Tracking modelled objects using binocular images. Computer Vision, Graphics, and Image Processing, 32(3):384–396, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Meyer:1994:RBT

[MB94] François G. Meyer and Patrick Bouthemy. Regionbased tracking using affine motion models in long image sequences. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):119–140, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1042/production;   http:
                                        //www.idealibrary.com/links/

artid/ciun.1994.1042/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1047/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1047/production/
pdf.

Monga:1995:UPD

[MB95] Olivier Monga and Serge Benayoun. Using partial derivatives of 3D images to extract typical surface features. Computer Vision and Image Understanding: CVIU, 61(2):171–189, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1014/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1014/production/
pdf.

Mortensen:1998:ISI

[MB98] Eric N. Mortensen and William A. Barrett. Interactive segmentation with intelligent scissors. Graphical Models and Image Processing: GMIP, 60(5):349– 384, September 1998. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0480/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0480/production/
pdf;   http://www.idealibrary.

                                        com/links/artid/gmip.1998.
                                        0480/production/ref.

Min:2005:IRI

[MB05] Jaesik Min and Kevin W. Bowyer. Improved range image segmentation by analyzing surface fit patterns. Computer Vision and Image Understanding: CVIU, 97(2): 242–258, February 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Moore:2011:LBP

[MB11] S. Moore and R. Bowden. Local binary patterns for multiview facial expression recognition. Computer Vision and Image Understanding: CVIU, 115(4):541–558, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

McCollum:1988:HMU

[MBDB88] A. J. McCollum, C. C. Bowman, P. A. Daniels, and B. G. Batchelor. A histogram modification unit for real-time image enhancement. Computer Vision, Graphics, and Image Processing, 42(3):387–398, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Marras:2012:MBM

[MBH^{+}12] Stefano Marras, Michael M. Bronstein, Kai Hormann, Riccardo Scateni, and Roberto Scopigno. Motion-based mesh segmentation using augmented silhouettes. Graphical Models, 74(4):164–172, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000215.

Martins:1981:NCM

[MBK81] H. A. Martins, J. R. Birk, and R. B. Kelley. Note: Camera models based on data from two calibration planes. Computer Graphics and Image Processing, 17(2):173– 180, October 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

MacArthur:2002:ICB

[MBKB02] Sean D. MacArthur, Carla E.

Brodley, Avinash C. Kak, and Lynn S. Broderick. Interactive content-based image retrieval using relevance feedback. Computer Vision and Image Understanding: CVIU, 88(2):55–75, November 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mille:2011:CNB

[MBMC11] Julien Mille, Romuald Boné, Pascal Makris, and Hubert Cardot. Corrigendum to “Narrow band region-based active contours and surfaces for 2D and 3D segmentation” [Comput. Vis. Image Understand. 113 (2009) 946–965]. Computer Vision and Image Understanding: CVIU, 115 (2):286, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mukherjee:1995:ANE

[MC95] Debargha Mukherjee and B. N. Chatterji. Adaptive neighborhood extended contrast enhancement and its modifications. Graphical Models and Image Processing: GMIP, 57(3):254–265, May 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1024/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1024/production/
                                        pdf.

Maggio:2009:AAB

[MC09a] Emilio Maggio and Andrea Cavallaro. Accurate appearance-based Bayesian tracking for maneuvering targets. Computer Vision and Image Understanding: CVIU, 113(4):544–555, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Marques:2009:ESD

[MC09b] Manuel Marques and João Costeira. Estimating 3D shape from degenerate sequences with missing data.

Computer Vision and Image Understanding: CVIU, 113 (2):261–272, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Martins:2013:GFA

[MCB13] Pedro Martins, Rui Caseiro, and Jorge Batista. Generative face alignment through 2.5D active appearance models. Computer Vision and Image Understanding: CVIU, 117(3):250–268, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001944.

McClure:1980:IMP

[McC80] Donald E. McClure. Image models in pattern theory. Computer Graphics and Image Processing, 12(4): 309–325, April 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

McClellan:1982:MAR

[McC82] James H. McClellan. Modified alpha-root technique for image processing. Computer Graphics and Image Processing, 19(1):18–34, May 1982. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

McDonnell:1981:BFT

[McD81a]  M. J. McDonnell.  Box-

filtering techniques. Computer Graphics and Image Processing, 17(1):65– 70, September 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

McDonnell:1981:RVI

[McD81b] M. J. McDonnell. Restoration of Voyager 1 images of Io. Computer Graphics and Image Processing, 15(1): 79–86, January 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Molnar:2010:IRV

[MCF10] József Molnár, Dmitry Chetverikov, and Sándor Fazekas. Illuminationrobust variational optical flow using cross-correlation. Computer Vision and Image Understanding: CVIU, 114(10): 1104–1114, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mrak:2009:FAS

[MĆK09] Marta Mrak, Janko Ćalić, and Ahmet Kondoz. Fast analysis of scalable video for adaptive browsing interfaces. Computer Vision and Image Understanding: CVIU, 113(3): 425–434, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mclean:1993:CED

[Mcl93] G. F. Mclean. Codebook edge detection. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(1):48–57, January 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1003/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1003/production/
pdf.

McLean:1996:GCD

[McL96]  G. F. McLean.  Geo-

metric correction of digitized art. Graphical Models and Image Processing: GMIP, 58(2):142–154, March 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0012/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0012/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1996.
0013/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0013/production/
pdf.

Mignotte:1999:TCM

[MCPB99] M. Mignotte, C. Collet, P. Pérez, and P. Bouthemy. Three-class Markovian segmentation of high-resolution sonar images. Computer Vision and Image Understanding: CVIU, 76(3): 191–204, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0804/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0804/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0804/production/ref.

Mignotte:2000:MRF

[MCPB00] M. Mignotte, C. Collet, P. Pérez, and P. Bouthemy. Markov random field and fuzzy logic modeling in sonar imagery: Application to the classification of underwater floor. Computer Vision and Image Understanding: CVIU, 79(1):4–24, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0844;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0844/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0844/ref.

McDonnell:2005:DSB

[MCQ05] Kevin T. McDonnell, Yu-Sung Chang, and Hong Qin. DigitalSculpture: a subdivisionbased approach to interactive implicit surface modeling. Graphical Models, 67 (4):347–369, July 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Mavrinac:2010:ACM

[MCT10] Aaron Mavrinac, Xiang Chen, and Kemal Tepe. An automatic calibration method for stereo-based 3D distributed smart camera networks. Computer Vision and Image Understanding: CVIU, 114(8): 952–962, August 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mori:1982:STE

[MD82] Shunji Mori and Michihisa Doh. A sequential tracking extraction of shape features and its constructive description. Computer Graphics and Image Processing, 19(4): 349–366, August 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Manocha:1995:AIP

[MD95] Dinesh Manocha and James Demmel. Algorithms for intersecting parametric and algebraic curves II: multiple intersections. Graphical Models and Image Processing: GMIP, 57(2):81–100, March 1995. CODEN GMIPF4.

ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1010/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1010/production/
                                        pdf.

Maurin:2009:FAS

[MDdMG09] B. Maurin, C. Doignon, M. de Mathelin, and A. Gangi. A fast and automatic stereotactic registration with a single CT-slice. Computer Vision and Image Understanding: CVIU, 113(8):878–890, August 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Macrini:2011:BGM

[MDFS11a] Diego Macrini, Sven Dickinson, David Fleet, and Kaleem Siddiqi. Bone graphs: Medial shape parsing and abstraction. Computer Vision and Image Understanding: CVIU, 115(7):1044–1061, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000798.

Macrini:2011:OCU

[MDFS11b] Diego Macrini, Sven Dickinson, David Fleet, and Kaleem Siddiqi. Object categorization using bone graphs. Computer Vision and Image Understanding: CVIU, 115 (8):1187–1206, August 2011.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000907.

Monga:1991:EDU

[MDR91] O. Monga, R. Deriche, and J.-M. Rocchisani. 3D edge detection using recursive filtering: application to scanner images. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (1):76–87, January 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Moganti:1998:SPC

[ME98a] Madhav Moganti and Fikret Ercal. Segmentation of printed circuit board images into basic patterns. Computer Vision and Image Understanding: CVIU, 70(1):74–86, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0594/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0594/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0594/production/ref.

Moganti:1998:SLI

[ME98b] Madhav Moganti and Fikret Ercal. A subpattern level inspection system for printed circuit boards. Computer Vision and Image Understanding: CVIU, 70(1):51–62, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0600/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0600/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0600/production/ref.

Meagher:1982:GMU

[Mea82] Donald Meagher. Geometric modeling using octree encoding. Computer Graphics and Image Processing, 19(2): 129–147, June 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Medek:1984:CHE

[Med84] V. Medek. Culling hidden edges of rectangular parallelepipeds. Computer Vision, Graphics, and Image Processing, 28(2):263–268, November 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Moganti:1996:API

[MEDT96] Madhav Moganti, Fikret Ercal, Cihan H. Dagli, and Shou Tsunekawa. Automatic PCB inspection algorithms: a survey. Computer Vision and Image Understanding: CVIU, 63(2):287–313, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0020/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0020/production/
pdf.

Meer:1989:SIP

[Mee89] Peter Meer. Stochastic image pyramids. Computer Vision, Graphics, and Image Processing, 45(3):269– 294, March 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Meer:1994:CVG

[Mee94] Peter Meer. Computer vision: The goal and the means. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):257–259, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1053/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1053/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1058/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1058/production/
pdf.

Mero:1981:ASR

[Mer81] Laszlo Mero. Algorithm for scale-and rotation-invariant recognition of two-dimensional objects. Computer Graphics and Image Processing, 15(3): 279–287, March 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Merickel:1988:RRP

[Mer88] Michael Merickel. 3D reconstruction: the registration problem. Computer Vision, Graphics, and Image Processing, 42(2):206–219, May 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Meyer:1986:ASC

[Mey86] Fernand Meyer. Automatic screening of cytological specimens. Computer Vision, Graphics, and Image Processing, 35(3):356–369, September 1986. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Meyer:1988:WSS

[Mey88]  Gary W. Meyer.  Wave-

length selection for synthetic image generation. Computer Vision, Graphics, and Image Processing, 41(1):57– 79, January 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Matov:2011:OFM

[MEYD11] Alexandre Matov, Marcus M.

Edvall, Ge Yang, and Gaudenz Danuser. Optimalflow minimum-cost correspondence assignment in particle flow tracking. Computer Vision and Image Understanding: CVIU, 115(4):531–540, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Modestino:1977:EDN

[MF77] J. W. Modestino and R. W.

Fries. Edge detection in noisy images using recursive digital filtering. Computer Graphics and Image Processing, 6(5): 409–433, October 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Mitiche:1989:CSL

[MFA89] A. Mitiche, O. Faugeras, and J. K. Aggarwal. Counting straight lines. Computer Vision, Graphics, and Image Processing, 47(3):353– 360, September 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Mackay:2011:MCA

[MFB11] Matthew D. Mackay, Robert G.

Fenton, and Beno Benhabib. Multi-camera active surveillance of an articulated human form — an implementation strategy. Computer Vision and Image Understanding: CVIU, 115(10): 1395–1413, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001433.

Michalke:2010:BIV

[MFG10] Thomas Michalke, Jannik Fritsch, and Christian Goerick. A biologically-inspired vision architecture for resourceconstrained intelligent vehicles. Computer Vision and Image Understanding: CVIU, 114(5):548–563, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mao:1995:IMF

[MFJ95] Jianchang Mao, Patrick J. Flynn, and Anil K. Jain. Integration of multiple feature groups and multiple views into a 3D object recognition system. Computer Vision and Image Understanding: CVIU, 62(3):309–325, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1057/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1057/production/
pdf.

Medioni:2007:RRT

[MFS^{+}07] Gérard Medioni, Alexandre R. J. François, Matheen Siddiqui, Kwangsu Kim, and Hosub Yoon. Robust realtime vision for a personal service robot. Computer Vision and Image Understanding: CVIU, 108(1–2):196– 203, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Modestino:1980:SIM

[MFV80] J. W. Modestino, R. W. Fries, and A. L. Vickers. Stochastic image models generated by random tessellations of the plane. Computer Graphics and Image Processing, 12(1): 74–98, January 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Mason:1995:ASP

[MG95a] Scott O. Mason and Armin Grün. Automatic sensor placement for accurate dimensional inspection. Computer Vision and Image Understanding: CVIU, 61(3):454–467, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1034/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1034/production/
pdf.

Miller:1995:GAD

[MG95b] James R. Miller and Ronald N. Goldman. Geometric algorithms for detecting and calculating all conic sections in the intersection of any two natural quadratic surfaces. Graphical Models and Image Processing: GMIP, 57(1):55–66, January 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1006/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1006/production/
                                        pdf.

Moeslund:2001:SCV

[MG01] Thomas B. Moeslund and Erik Granum. A survey of computer visionbased human motion capture. Computer Vision and Image Understanding: CVIU, 81(3):231–268, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0897;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0897/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0897/ref.

Matas:2000:RDL

[MGK00] J. Matas, C. Galambos, and J. Kittler. Robust detection of lines using the progressive probabilistic Hough transform. Computer Vision and Image Understanding: CVIU, 78(1):119–137, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0831;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0831/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0831/ref.

Manzotti:2001:DEL

[MGMS01] R. Manzotti, A. Gasteratos, G. Metta, and G. Sandini. Disparity estimation on logpolar images and vergence control. Computer Vision and Image Understanding: CVIU, 83(2):97–117, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0924;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0924/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0924/ref.

Morillas:2008:IIN

[MGPF08] Samuel Morillas, Valentín Gregori, and Guillermo Peris-Fajarnés. Isolating impulsive noise pixels in color images by peer group techniques. Computer Vision and Image Understanding: CVIU, 110(1): 102–116, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Moreno:2011:EPC

[MGPJ11] Rodrigo Moreno, Miguel Angel Garcia, Domenec Puig, and Carme Julià. Edgepreserving color image denoising through tensor voting. Computer Vision and Image Understanding: CVIU, 115(11):1536–1551, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001706.

Melendez:2011:UTB

[MGPP11] Jaime Melendez, Miguel Angel Garcia, Domenec Puig, and Maria Petrou. Unsupervised texture-based image segmentation through pattern discovery. Computer Vision and Image Understanding: CVIU, 115(8):1121–1133, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000968.

Ma:2010:HSM

[MGW10] Yu Ma, Xiaodong Gu, and Yuanyuan Wang. Histogram similarity measure using variable bin size distance. Computer Vision and Image Understanding: CVIU, 114(8):

981–989, August 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Milgram:1979:CEV

[MH79] David L. Milgram and Martin Herman. Clustering edge values for threshold selection. Computer Graphics and Image Processing, 10(3): 272–280, July 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Miyaoku:1998:APC

[MH98] Kento Miyaoku and Koichi Harada. Approximating polygonal curves in two and three dimensions. Graphical Models and Image Processing: GMIP, 60(3):222–225, May 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0468/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0468/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0468/production/ref.

Moeslund:2006:SAV

[MHK06] Thomas B. Moeslund, Adrian Hilton, and Volker Krüger. A survey of advances in visionbased human motion capture and analysis. Computer Vision and Image Understanding: CVIU, 104(2–3):

90–126, November/December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Medrano:2009:MFA

[MHMO09] C. Medrano, J. E. Herrero, J. Martínez, and C. Orrite. Mean field approach for tracking similar objects. Computer Vision and Image Understanding: CVIU, 113(8): 907–920, August 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Matsuyama:1984:FOG

[MHN84] Takashi Matsuyama, Le Viet Hao, and Makoto Nagao. A file organization for geographic information systems based on spatial proximity. Computer Vision, Graphics, and Image Processing, 26(3):303–318, June 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Medeiros:2010:PHB

[MHSP10] Henry Medeiros, Germán Holguín, Paul J. Shin, and Johnny Park. A parallel histogram-based particle filter for object tracking on SIMDbased smart cameras. Computer Vision and Image Understanding: CVIU, 114(11): 1264–1272, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

McKeown:1989:AKA

[MHW89] D. M. McKeown, Jr., W. A.

Harvey, and L. E. Wixson. Automating knowledge acquisition for aerial image interpretation. Computer Vision, Graphics, and Image Processing, 46(1):37–81, April 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Midgley:1979:IFP

[Mid79] James E. Midgley. Isotropic four-point interpolation. Computer Graphics and Image Processing, 11(2):192– 196, October 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Mignotte:2012:MBS

[Mig12] Max Mignotte. MDS-based segmentation model for the fusion of contour and texture cues in natural images. Computer Vision and Image Understanding: CVIU, 116 (9):981–990, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421200077X.

Milgram:1979:REU

[Mil79a] D. L. Milgram. Region extraction using convergent evidence. Computer Graphics and Image Processing, 11(1): 1–12, September 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Milgram:1979:CTR

[Mil79b] David L. Milgram. Constructing trees for region description. Computer Graphics and Image Processing, 11(1):88– 99, September 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Miles:1980:SGP

[Mil80a] R. E. Miles. Survey of geometrical probability in the plane, with emphasis on stochastic image modeling. Computer Graphics and Image Processing, 12(1):1–24, January 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Mills:1980:ICP

[Mil80b] Harlan D. Mills. On information content in patterns. Computer Graphics and Image Processing, 14(3):183– 202, November 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Milios:1989:SMU

[Mil89] Evangelos E. Milios. Shape matching using curvature processes. Computer Vision, Graphics, and Image Processing, 47(2):203–226, August 1989. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Miller:1999:ATI

[Mil99] Erik G. Miller. Alternative tilings for improved surface area estimates by local counting algorithms. Computer Vision and Image Understanding: CVIU, 74(3):193–211, June 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0754/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0754/production/
pdf.

Mille:2009:NBR

[Mil09] Julien Mille. Narrow band region-based active contours and surfaces for 2D and 3D segmentation. Computer Vision and Image Understanding: CVIU, 113(9):946– 965, September 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Minerbo:1979:MME

[Min79] Gerald Minerbo. Ment: a maximum entropy algorithm for reconstructing a source from projection data. Computer Graphics and Image Processing, 10(1):48–68, May 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Mintz:1994:RCB

[Min94] Doron Mintz. Robust consensus based edge detection. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):137–153, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1009/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1009/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1011/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1011/production/
pdf.

Misra:1984:DAU

[Mis84] Rachita Misra. A deblurring algorithm using edge gradient for the Walsh-Hadamard transform image coding system. Computer Vision, Graphics, and Image Processing, 27(3):369–379, September 1984. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Mitiche:1988:TDS

[Mit88]  Amar Mitiche.  Three-

dimensional space from optical flow correspondence. Computer Vision, Graphics, and Image Processing, 42(3):306–317, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

McLean:1988:HED

[MJ88] G. F. McLean and M. E.

Jernigan. Hierarchical edge detection. Computer Vision, Graphics, and Image Processing, 44(3):350–366, December 1988. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Martinez:2011:TMG

[MJ11] Antonio Martínez and Juan José Jiménez. Tracking by means of geodesic region models applied to multidimensional and complex medical images. Computer Vision and Image Understanding: CVIU, 115 (8):1083–1098, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000932.

Meer:1988:RIP

[MJBR88] Peter Meer, Song-Nian Jiang, Ernest S. Baugher, and Azriel Rosenfeld. Robustness of image pyramids under structural perturbations. Computer Vision, Graphics, and Image Processing, 44(3):307– 331, December 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

McKenna:2000:TGP

[MJD^{+}00] Stephen J. McKenna, Sumer Jabri, Zoran Duric, Azriel Rosenfeld, and Harry Wechsler. Tracking groups of people. Computer Vision and Image Understanding: CVIU, 80(1):42–56, October 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0870;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0870/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0870/ref.

Mann:1997:CPS

[MJS97] Richard Mann, Allan Jepson, and Jeffrey Mark Siskind. The computational perception of scene dynamics. Computer Vision and Image Understanding: CVIU, 65 (2):113–128, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0576/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0576/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0576/production/ref.

Mohwinkel:1976:CPP

[MK76] C. Mohwinkel and L. Kurz. Computer picture processing and enhancement by localized operations. Computer Graphics and Image Processing, 5 (4):401–424, December 1976. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Milgram:1979:RRE

[MK79] David L. Milgram and Daryl J.

Kahl. Recursive region extraction. Computer Graphics and Image Processing, 9 (1):82–88, January 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Medasani:2001:CID

[MK01] Swarup Medasani and Raghu Krishnapuram. Categorization of image databases for efficient retrieval using robust mixture decomposition. Computer Vision and Image Understanding: CVIU, 83(3):216–235, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0926;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0926/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0926/ref.

Maekawa:2002:SCF

[MK02] T. Maekawa and K. H. Ko.

Surface construction by fitting unorganized curves. Graphical Models, 64(5):316–332, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Marinov:2005:OMS

[MK05] Martin Marinov and Leif Kobbelt. Optimization methods for scattered data approximation with subdivision surfaces. Graphical Models, 67(5):452–473, September 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000068.

Mori:1973:IPC

[MKA73] Ken-Ichi Mori, Masatsugu Kidode, and Haruo Asada. An iterative prediction and correction method for automatic stereocomparison. Computer Graphics and Image Processing, 2(3/4):393– 401, December 1973. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Matas:2002:MNS

[MKK02] J. Matas, D. Koubaroulis, and J. Kittler. The multimodal neighborhood signature for modeling object color appearance and applications in object recognition and image retrieval. Computer Vision and Image Understanding: CVIU, 88(1):1–23, October 2002. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Mihalef:2008:ITP

[MKS^{+}08] Viorel Mihalef, Samet Kadioglu, Mark Sussman, Dimitris Metaxas, and Vassilios Hurmusiadis. Interaction of two-phase flow with animated models. Graphical Models, 70(3):33–42, May 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070307000264.

Milanfar:1994:RBP

[MKW94] P. Milanfar, W. C. Karl, and A. S. Willsky. Reconstructing binary polygonal objects from projections: a statistical view. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(5):371–391, September 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1034/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1034/production/
pdf.

Mokhtarian:2001:CCF

[MKY01] F. Mokhtarian, N. Khalili, and P. Yuen. Curvature computation on freeform 3-D meshes at multiple scales. Computer Vision and Image Understanding: CVIU, 83(2):118–139, August 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0919;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0919/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0919/ref.

Milstein:1978:STI

[ML78] Laurence B. Milstein and Thomas Lazicky. Statistical tests for image tracking. Computer Graphics and Image Processing, 17(3): 413–424, June 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Malgouyres:2000:TPW

[ML00] Rémy Malgouyres and Alexandre Lenoir. Topology preservation within digital surfaces. Graphical Models, 62 (2):71–84, March 2000. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
1999.0517;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.1999.0517/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
1999.0517/ref.

Ma:2013:HSI

[ML13] Jun Ma and Le Lu. Hierarchical segmentation and identification of thoracic vertebra using learning-based edge detection and coarse-to-fine deformable model. Computer Vision and Image Understanding: CVIU, 117(9): 1072–1083, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000763.

Ma:2012:VNG

[MLF^{+}12] Teng Ma, Xiang Long, Lu Feng, Pei Luo, and Zhuangzhi Wu. Visible neighborhood graph of point clouds. Graphical Models, 74(4):184–196, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000276.

Mekonnen:2013:CPT

[MLH13] A. A. Mekonnen, F. Lerasle, and A. Herbulot. Cooperative passers-by tracking with a mobile robot and external cameras. Computer Vision and Image Understanding: CVIU, 117(10): 1229–1244, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212002044.

Marcelli:1997:SIC

[MLP97] Angelo Marcelli, Natasha Likhareva, and Theo Pavlidis. Structural indexing for character recognition. Computer Vision and Image Understanding: CVIU, 66(3): 330–346, June 1997. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/artid/
                                        cviu.1996.0518/production;
                                         http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0518/production/pdf;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0518/production/
                                        ref.

Marangelli:1980:PAR

[MM80] B. Marangelli and N. Mirizzi. A pointer for analog refresh memories. Computer Graphics and Image Processing, 14 (2):170–176, October 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Marangelli:1981:PEA

[MM81] B. Marangelli and N. Mirizzi. Pseudocolor encoder for analog and digital refresh memories. Computer Graphics and Image Processing, 17(1):60– 64, September 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

McIntosh:1988:MSL

[MM88] James H. McIntosh and Kathleen M. Mutch. Matching straight lines. Computer Vision, Graphics, and Image Processing, 43(3):386– 408, September 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Meisels:1990:SRO

[MM90] A. Meisels and D. Mintz.

Symbolic reasoning in object extraction. Computer Vision, Graphics, and Image Processing, 52(3):447– 459, December 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Metaxas:1992:RCI

[MM92] Dimitris Metaxas and Evangelos Milios. Reconstruction of a color image from nonuniformly distributed sparse and noisy data. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(2): 103–111, March 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Morimoto:2005:EGT

[MM05] Carlos H. Morimoto and Marcio R. M. Mimica. Eye gaze tracking techniques for interactive applications. Computer Vision and Image Understanding: CVIU, 98(1): 4–24, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mokhtarian:2006:PEC

[MM06] Farzin Mokhtarian and Farahnaz Mohanna. Performance evaluation of corner detectors using consistency and accuracy measures. Computer Vision and Image Understanding: CVIU, 102(1): 81–94, April 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Maschino:2006:JRA

[MMA06] Emeric Maschino, Yves Maurin, and Philippe Andrey. Joint registration and averaging of multiple 3D anatomical surface models. Computer Vision and Image Understanding: CVIU, 101(1): 16–30, January 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Martinez:2004:CGP

[MMK04] Aleix M. Martínez, Pradit Mittrapiyanuruk, and Avinash C. Kak. On combining graph-partitioning with non-parametric clustering for image segmentation. Computer Vision and Image Understanding: CVIU, 95(1): 72–85, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mansouri:1987:LDD

[MML87] Abdol-Reza R. Mansouri, Alfred S. Malowany, and Martin D. Levine. Line detection in digital pictures: a hypothesis prediction/verification paradigm. Computer Vision, Graphics, and Image Processing, 40(1):95–114, October 1987. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Matsuyama:1983:SAN

[MMN83] Takashi Matsuyama, Shu-Ichi Miura, and Makoto Nagao. Structural analysis of natural textures by Fourier transformation. Computer Vision, Graphics, and Image Processing, 24(3):347– 362, December 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Merelli:1985:ADD

[MMP85] D. Merelli, P. Mussio, and M. Padula. An approach to the definition, description, and extraction of structures in binary digital images. Computer Vision, Graphics, and Image Processing, 31(1):19–49, July 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Mittal:2009:SMC

[MMP09] Anurag Mittal, Antoine Monnet, and Nikos Paragios.

Scene modeling and change detection in dynamic scenes: a subspace approach. Computer Vision and Image Understanding: CVIU, 113(1): 63–79, January 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Marchand-Maillet:1997:DCS

[MMS97] Stéphane Marchand-Maillet and Yazid M. Sharaiha. Discrete convexity, straightness, and the 16-neighborhood. Computer Vision and Image Understanding: CVIU, 66(3):316–329, June 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0521/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0521/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0521/production/ref.

Marchand-Maillet:1999:EOC

[MMS99] Stéphane Marchand-Maillet and Yazid M. Sharaiha. Euclidean ordering via chamfer distance calculations. Computer Vision and Image Understanding: CVIU, 73(3):404–413, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.

0743/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0743/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0743/production/ref.

Moenning:2007:MGS

[MMS^{+}07] C. Moenning, F. Mémoli, G. Sapiro, N. Dyn, and N. A. Dodgson. Meshless geometric subdivision. Graphical Models, 69(3–4): 160–179, May/July 2007. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000956.

Mansouri:2006:MCL

[MMV06] Abdol-Reza Mansouri, Amar Mitiche, and Carlos Vázquez. Multiregion competition: a level set extension of region competition to multiple region image partitioning. Computer Vision and Image Understanding: CVIU, 101(3): 137–150, March 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Medioni:1985:SBS

[MN85] Gerard Medioni and Ramakant Nevatia. Segmentbased stereo matching. Computer Vision, Graphics, and Image Processing, 31(1):2– 18, July 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Mount:1994:CEA

[MN94] D. M. Mount and N. S. Netanyahu. Computationally efficient algorithms for high-dimensional robust estimators. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(4):289–303, July 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1026/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1026/production/
pdf.

Michel:1995:UMM

[MN95] Jonathan Michel and N. Nandhakumar. Unified 3D models for multisensor image synthesis. Graphical Models and Image Processing: GMIP, 57(4):283–302, July 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1026/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1026/production/
pdf.

Madjidi:2006:RLA

[MN06] Hossein Madjidi and Shahriar Negahdaripour. On robustness and localization accuracy of optical flow computation for underwater color images. Computer Vision and Image Understanding: CVIU, 104 (1):61–76, October 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

McCane:2001:BOF

[MNCG01] B. McCane, K. Novins, D. Crannitch, and B. Galvin. On benchmarking optical flow. Computer Vision and Image Understanding: CVIU, 84(1):126–143, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0930;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0930/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0930/ref.

Maki:2000:ASS

[MNE00] Atsuto Maki, Peter Nordlund, and Jan-Olof Eklundh. Attentional scene segmentation: Integrating depth and motion. Computer Vision and Image Understanding: CVIU, 78(3):351–373, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0840;   http://www.
idealibrary.com/links/doi/

                                        10.1006/cviu.2000.0840/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0840/ref.

Meribout:2000:HTA

[MNHO00] Mahmoud Meribout, Mamoru Nakanishi, Eiichi Hosoya, and Takeshi Ogura. Hough transform algorithm for threedimensional segment extraction and its parallel hardware implementation. Computer Vision and Image Understanding: CVIU, 78(2):177–205, May 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0834;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0834/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0834/ref.

Michel:1998:GAT

[MNSK98] J. D. Michel, N. Nandhakumar, Tushar Saxena, and Deepak Kapur. Geometric, algebraic, and thermophysical techniques for object recognition in IR imagery. Computer Vision and Image Understanding: CVIU, 72(1):84–97, October 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0669/production;   http:

//www.idealibrary.com/links/
artid/cviu.1997.0669/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0669/production/ref.

McGuinness:2011:TAE

[MO11] Kevin McGuinness and Noel E.

O’Connor. Toward automated evaluation of interactive segmentation. Computer Vision and Image Understanding: CVIU, 115(6):868–884, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mokrane:1992:NIC

[Mok92] A. Mokrane. A new image contrast enhancement technique based on a contrast discrimination model. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(2): 171–180, March 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Mokhtarian:1997:TMT

[Mok97] Farzin Mokhtarian. A theory of multiscale, torsionbased shape representation for space curves. Computer Vision and Image Understanding: CVIU, 68(1):1–17, October 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0544/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0544/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0544/production/ref.

Montani:1984:RRP

[Mon84] Claudio Montani. Region representation: Parallel connected stripes. Computer Vision, Graphics, and Image Processing, 28(2):139– 165, November 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Moonis:1977:MPG

[Moo77] A. Moonis. Mathematical picture grammar applied to script generation. Computer Graphics and Image Processing, 6(??):93–102, ???? 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Morrin:1976:CLC

[Mor76] T. H. Morrin, II. Chainlink compression of arbitrary black-white images. Computer Graphics and Image Processing, 5(2):172– 189, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Morgenthaler:1981:NHE

[Mor81] David G. Morgenthaler. New hybrid edge detector. Computer Graphics and Image Processing, 16(2):166– 176, June 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Moran:1990:MTS

[Mor90] C. J. Moran. A morphological transformation for sharpening edges of features before segmentation. Computer Vision, Graphics, and Image Processing, 49(1):85– 94, January 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Moshfeghi:1991:EMM

[Mos91] M. Moshfeghi. Elastic matching of multimodality medical images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(3): 271–282, May 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Muhlthaler:2003:CMS

[MP03] Heidrun Mühlthaler and Helmut Pottmann. Computing the Minkowski sum of ruled surfaces. Graphical Models, 65 (6):369–384, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Magee:2009:CVB

[MP09a] Derek Magee and Janez Pers.

Computer vision based analysis in sport environments.

Computer Vision and Image Understanding: CVIU, 113 (5):589, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Montoliu:2009:GLS

[MP09b] R. Montoliu and F. Pla. Generalized least squares-based parametric motion estimation. Computer Vision and Image Understanding: CVIU, 113(7):790–801, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Meer:1994:MAI

[MPC94] Peter Meer, Rae-Hong H. Park, and Kyujin J. Cho. Multiresolution adaptive image smoothing. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(2):140–148, March 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1013/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1013/production/
pdf.

MacLean:2007:SIS

[MPF07] W. James MacLean, Nikos Paragios, and David Fleet. Special issue on spatial coherence for visual motion analysis. Computer Vision and Image Understanding: CVIU, 108(3):205–206, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Martinez-Perez:1987:TAB

[MPJN87] M. Pilar Martinez-Perez, Javier Jimenez, and Jose L. Navalon. A thinning algorithm based on contours. Computer Vision, Graphics, and Image Processing, 39 (2):186–201, August 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Morse:1998:ZIV

[MPPG98] B. S. Morse, S. M. Pizer, D. T. Puff, and C. Gu. Zoom-invariant vision of figural shape. effects on cores of image disturbances. Computer Vision and Image Understanding: CVIU, 69(1): 72–??, ???? 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Marcon:2008:FPA

[MPST08] Marco Marcon, Luca Piccarreta, Augusto Sarti, and Stefano Tubaro. Fast PDE approach to surface reconstruction from large cloud of points. Computer Vision and Image Understanding: CVIU, 112(3):274–285, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Martinez:2013:SRO

[MPV13] Jonàs Martínez, Núria Pla, and Marc Vigo. Skeletal representations of orthogonal shapes. Graphical Models, 75(4):189–207, July 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000155.

Maximo:2011:RRI

[MPVF11] A. Maximo, R. Patro, A. Varshney, and R. Farias. A robust and rotationally invariant local surface descriptor with applications to non-local mesh processing. Graphical Models, 73(5):231–242, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070311000166.

Manohar:1989:CCL

[MR89] M. Manohar and H. K. Ramapriyan. Connected component labeling of binary images on a mesh connected massively parallel processor. Computer Vision, Graphics, and Image Processing, 45 (2):133–149, February 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Margalit:1990:UFP

[MR90a] Avraham Margalit and Azriel Rosenfeld. Using feature probabilities to reduce the expected computational cost of template matching. Computer Vision, Graphics, and Image Processing, 52(1):110– 123, October 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Margalit:1990:UPD

[MR90b] Avraham Margalit and Azriel Rosenfeld. Using probabilistic domain knowledge to reduce the expected computational cost of template matching. Computer Vision, Graphics, and Image Processing, 51 (3):219–234, September 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Myrheim:1992:NAM

[MR92] Jan Myrheim and Håvard Rue. New algorithms for maximum entropy image restoration. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(3): 223–238, May 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Moreau:1996:GSI

[MR96] Patrick Moreau and Christian Ronse. Generation of shading-off in images by extrapolation of Lipschitz functions. Graphical Models and Image Processing:

GMIP, 58(4):314–333, July 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0026/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0026/production/
pdf.

Mateer:2005:VBP

[MR05] J. W. Mateer and J. A. Robinson. A vision-based postproduction tool for footage logging, analysis, and annotation. Graphical Models, 67 (6):565–583, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000123.

Murino:1996:GSP

[MRF96] Vittorio Murino, Carlo S. Regazzoni, and Gian Luca Foresti. Grouping as a searching process for minimumenergy configurations of labelled random fields. Computer Vision and Image Understanding: CVIU, 64(1): 157–174, July 1996. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1996.0051/production;
 http://www.idealibrary.
com/links/artid/cviu.1996.
0051/production/pdf.

Mair:1997:SAD

[MRW^{+}97] Bernard A. Mair, Zoltán Réti, David C. Wilson, Edward A. Geiser, and Bryn David. A q-series approach to deblurring the discrete Gaussian. Computer Vision and Image Understanding: CVIU, 66(2):247–254, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0609/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0609/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0609/production/ref.

Mallgren:1978:GTH

[MS78] W. R. Mallgren and A. C.

Shaw. Graphical transformations and hierarchic picture structures. Computer Graphics and Image Processing, 8 (2):237–258, October 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Messner:1985:IPA

[MS85] Richard A. Messner and Harold H. Szu. An image processing architecture for real time generation of scale and rotation invariant patterns. Computer Vision, Graphics, and Image Processing, 31(1):50–66, July 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Montani:1994:UMC

[MS94] Claudio Montani and Roberto Scopigno. Using marching cubes on small machines. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(2):182–183, March 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1017/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1017/production/
                                        pdf.

Ma:1996:FPT

[MS96a] C. Min Ma and Milan Sonka. A fully parallel 3D thinning algorithm and its applications. Computer Vision and Image Understanding: CVIU, 64(3):420–433, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0069/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0069/production/
                                        pdf.

Malladi:1996:IPF

[MS96b] R. Malladi and J. A. Sethian. Image processing: Flows under min/max curvature and mean curvature. Graphical Models and Image Processing: GMIP, 58(2):127–141, March 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0011/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0011/production/
pdf.

Murray:1996:DUP

[MS96c] D. W. Murray and L. S.

Shapiro. Dynamic updating of planar structure and motion: The case of constant motion. Computer Vision and Image Understanding: CVIU, 63(1):169–181, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0012/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0012/production/
pdf.

Malassiotis:1997:MBJ

[MS97a] Sotiris Malassiotis and Michael G.

Strintzis. Model-based joint motion and structure estimation from stereo images. Computer Vision and Image Understanding: CVIU, 65(1):79–94, January 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.

                                        0481/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0481/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0481/production/ref.

Maxwell:1997:PBS

[MS97b] Bruce A. Maxwell and Steven A. Shafer. Physics-based segmentation of complex objects using multiple hypotheses of image formation. Computer Vision and Image Understanding: CVIU, 65(2):269–295, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0573/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0573/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0573/production/ref.

Maxwell:2000:SIM

[MS00] Bruce A. Maxwell and Steven A. Shafer. Segmentation and interpretation of multicolored objects with highlights. Computer Vision and Image Understanding: CVIU, 77(1):1–24, January 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0801/production;   http:
                                        //www.idealibrary.com/links/

artid/cviu.1999.0801/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0801/production/ref.

Mahmoudi:2009:TDP

[MS09] Mona Mahmoudi and Guillermo Sapiro. Three-dimensional point cloud recognition via distributions of geometric distances. Graphical Models, 71(1):22–31, January 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000313.

Michikawa:2010:SGD

[MS10] Takashi Michikawa and Hiromasa Suzuki. Sparse grid distance transforms. Graphical Models, 72(4):35–45, July 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031000010X.

Minetto:2012:IVS

[MSF^{+}12] R. Minetto, T. V. Spina, A. X.

Falcão, N. J. Leite, J. P. Papa, and J. Stolfi. IFTrace : Video segmentation of deformable objects using the Image Foresting Transform. Computer Vision and Image Understanding: CVIU, 116 (2):274–291, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002141.

Mallik:2010:CBI

[MSG10] Joyita Mallik, Ashok Samal, and Scott L. Gardner. A content based image retrieval system for a biological specimen collection. Computer Vision and Image Understanding: CVIU, 114(7):745–757, July 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mulgaonkar:1986:SPR

[MSH86] Prasanna G. Mulgaonkar, Linda G. Shapiro, and Robert M. Haralick. Shape from perspective: a rulebased approach. Computer Vision, Graphics, and Image Processing, 36(2/3):298–320, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

MacLean:2010:LCM

[MSI10] W. James MacLean, Siraj Sabihuddin, and Jamin Islam. Leveraging cost matrix structure for hardware implementation of stereo disparity computation using dynamic programming. Computer Vision and Image Understanding: CVIU, 114(11): 1126–1138, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Monne:1981:BTS

[MSM81] J. Monne, F. Schmitt, and D. Massaloux. Bidimensional texture synthesis by Markov chains. Computer Graphics and Image Processing, 17(1): 1–23, September 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Matsuyama:1982:SAR

[MSN82] Takashi Matsuyama, Kinjiro Saburi, and Makoto Nagao. A structural analyzer for regularly arranged textures. Computer Graphics and Image Processing, 18(3): 259–278, March 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). MATSUYAMA82.

Magid:2007:CGM

[MSR07] Evgeni Magid, Octavian Soldea, and Ehud Rivlin. A comparison of Gaussian and mean curvature estimation methods on triangular meshes of range image data. Computer Vision and Image Understanding: CVIU, 107(3):139– 159, September 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Manohar:1990:TQR

[MSS90] M. Manohar, P. Sudarsana Rao, and S. Sitarama Iyengar. Template quadtrees for representing region and line data present in binary images. Computer Vision, Graphics, and Image Processing, 51 (3):338–354, September 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Miura:2009:BRE

[MSSS09] Jun Miura, Takumi Shimawaki, Takuro Sakiyama, and Yoshiaki Shirai. Ball route estimation under heavy occlusion in broadcast soccer video. Computer Vision and Image Understanding: CVIU, 113(5):653–662, May 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Massone:1985:FIT

[MST85] L. Massone, G. Sandini, and V. Tagliasco. “forminvariant” topological mapping strategy for 2D shape recognition. Computer Vision, Graphics, and Image Processing, 30(2):169– 188, May 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Meer:2000:RCV

[MST00] Peter Meer, Charles V. Stewart, and David E. Tyler. Robust computer vision: An interdisciplinary challenge. Computer Vision and Image Understanding: CVIU, 78(1):1–7, April 2000. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0833;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0833/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0833/ref.

Mount:1996:AOT

[MSW96] David M. Mount, Ruth Silverman, and Angela Y. Wu. On the area of overlap of translated polygons. Computer Vision and Image Understanding: CVIU, 64(1):53–61, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0045/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0045/production/
pdf.

Melter:1984:MBD

[MT84] Robert A. Melter and Ioan Tomescu. Metric bases in digital geometry. Computer Vision, Graphics, and Image Processing, 25(1):113– 121, January 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Metaxas:1997:GEI

[MT97] Dimitris Metaxas and Demetri Terzopoulos. GUEST ED-ITORS’ INTRODUCTION.

Computer Vision and Image Understanding: CVIU, 65 (2):111–112, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0593/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0593/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0593/production/ref.

Murino:2000:UCV

[MT00] Vittorio Murino and Andrea Trucco. Underwater computer vision and pattern recognition. Computer Vision and Image Understanding: CVIU, 79(1):1–3, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0852;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0852/
pdf.

Morales:2011:RTA

[MTAA11] Néstor Morales, Jonay T. Toledo, Leopoldo Acosta, and Rafael Arnay. Real-time adaptive obstacle detection based on an image database. Computer Vision and Image Understanding: CVIU, 115(9):1273–1287, September 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001263.

Mancas-Thillou:2007:CTE

[MTG07] Céline Mancas-Thillou and Bernard Gosselin. Color text extraction with selective metric-based clustering. Computer Vision and Image Understanding: CVIU, 107(1– 2):97–107, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mindru:2004:MIR

[MTVM04] Florica Mindru, Tinne Tuytelaars, Luc Van Gool, and Theo Moons. Moment invariants for recognition under changing viewpoint and illumination. Computer Vision and Image Understanding: CVIU, 94(1–3):3–27, April/ June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mikolajczyk:2011:ARA

[MU11] K. Mikolajczyk and H. Uemura. Action recognition with appearance–motion features and fast search trees. Computer Vision and Image Understanding: CVIU, 115 (3):426–438, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mukundan:1992:EQP

[Muk92]  R. Mukundan.  Estima-

tion of quaternion parameters from two dimensional image moments. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54 (4):345–350, July 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Mukawa:1997:OMB

[Muk97]  Naoki Mukawa.  Optical-

model-based analysis of consecutive images. Computer Vision and Image Understanding: CVIU, 66(1):25–32, April 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0500/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0500/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0500/production/ref.

Mulder:1988:DV

[Mul88] Jan A. Mulder. Discrimination vision. Computer Vision, Graphics, and Image Processing, 43(3):313– 336, September 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Mullikin:1992:VDT

[Mul92] James C. Mullikin. The vector distance transform in two and three dimensions. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(6): 526–535, November 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Munkelt:1995:ATG

[Mun95]  Olaf Munkelt.  Aspect-

trees: Generation and interpretation. Computer Vision and Image Understanding: CVIU, 61(3):365–386, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1029/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1029/production/
pdf.

Murray:1987:MBR

[Mur87] D. W. Murray. Model-based recognition using 3D shape alone. Computer Vision, Graphics, and Image Processing, 40(2):250–266, November 1987. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Murray:1995:RRU

[Mur95] David W. Murray. Recovering range using virtual multicamera stereo. Computer Vision and Image Understanding: CVIU, 61(2):285–291, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1021/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1021/production/
pdf.

Madabhushi:2006:GST

[MUS06] Anant Madabhushi, Jayaram K. Udupa, and Andre Souza. Generalized scale: Theory, algorithms, and application to image inhomogeneity correction. Computer Vision and Image Understanding: CVIU, 101(2):100– 121, February 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mukherjee:1986:DCR

[MV86] A. Mukherjee and Y. V. Venkatesh. Digital color reproduction on color television monitors. Computer Vision, Graphics, and Image Processing, 36(1):114– 132, October 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Macaire:2006:CIS

[MVP06] Ludovic Macaire, Nicolas Vandenbroucke, and Jack-Gérard Postaire. Color image segmentation by analysis of subset connectedness and color homogeneity properties. Computer Vision and Image Understanding: CVIU, 102 (1):105–116, April 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Manseur:1991:DMC

[MW91] Z. Z. Manseur and D. C. Wilson. Decomposition methods for convolution operators. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):428– 434, September 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Ma:2000:PTA

[MW00] Cherng-Min Ma and Shu-Yen Wan. Parallel thinning algorithms on 3D (18, 6) binary images. Computer Vision and Image Understanding: CVIU, 80(3):364–378, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0879;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0879/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0879/ref.

Metternich:2013:TBR

[MW13] Michael J. Metternich and Marcel Worring. Track based relevance feedback for tracing persons in surveillance videos. Computer Vision and Image Understanding: CVIU, 117(3):229–237, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001889.

Ma:2007:FVP

[MWF07] Zhonghua Ma, Hong Ren Wu, and Dagan Feng. Fuzzy vector partition filtering technique for color image restoration. Computer Vision and Image Understanding: CVIU, 107(1–2):26–37, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ma:1986:FSD

[MWL86] Jun Ma, Cheng-Ke K. Wu, and Xin-Ru R. Lu. A fast shape descriptor. Computer Vision, Graphics, and Image Processing, 34(3):282– 291, June 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Milun:1999:GRB

[MWL99] Elyse H. Milun, Deborah K. W. Walters, and Yiming Li. General ribbonbased thinning algorithms for stylus-generated images. Computer Vision and Image Understanding: CVIU, 76 (3):267–277, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0807/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0807/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0807/production/ref.

Milun:1999:GRM

[MWLA99] Elyse H. Milun, Deborah K. W. Walters, Yiming Li, and Bemina Atanacio. General ribbons: a model for stylus-generated images. Computer Vision and Image Understanding: CVIU, 76(3):259–266, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0806/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0806/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0806/production/ref.

Matsuyama:2004:RTS

[MWTN04] T. Matsuyama, X. Wu, T. Takai, and S. Nobuhara. Real-time 3D shape reconstruction, dynamic 3D mesh deformation, and high fidelity visualization for 3D video. Computer Vision and Image Understanding: CVIU, 96(3):

393–434, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Medioni:1987:CDC

[MY87a] Gerard Medioni and Yoshio Yasumoto. Corner detection and curve representation using cubic B-splines. Computer Vision, Graphics, and Image Processing, 39 (3):267–278, September 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Mitra:1987:TAS

[MY87b] Sanjit K. Mitra and Tian-Hu H. Yu. Transform amplitude sharpening: a new method of image enhancement. Computer Vision, Graphics, and Image Processing, 40(2):205–218, November 1987. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Masuda:1995:RMR

[MY95] Takeshi Masuda and Naokazu Yokoya. A robust method for registration and segmentation of multiple range images. Computer Vision and Image Understanding: CVIU, 61(3):295–307, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1995.

1024/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1024/production/
pdf.

Ma:2009:TTI

[MYC09] Yunqian Ma, Qian Yu, and Isaac Cohen. Target tracking with incomplete detection. Computer Vision and Image Understanding: CVIU, 113 (4):580–587, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Martinez:2003:SIF

[MYK03] Aleix M. Martínez, Ming-Hsuan Yang, and David J. Kriegman. Special issue on face recognition. Computer Vision and Image Understanding: CVIU, 91(1–2): 1–5, July/August 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Mirhosseini:1998:HFI

[MYLP98] Ali Reza Mirhosseini, Hong Yan, Kin-Man Lam, and Tuan Pham. Human face image recognition: An evidence aggregation approach. Computer Vision and Image Understanding: CVIU, 71(2):213–230, August 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0710/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0710/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0710/production/ref.

Mehrotra:1996:CAZ

[MZ96] Rajiv Mehrotra and Shiming Zhan. A computational approach to zero-crossing-based two-dimensional edge detection. Graphical Models and Image Processing: GMIP, 58(1):1–17, January 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0001/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0001/production/
                                        pdf.

Marzotto:2010:RTV

[MZB^{+}10] Roberto Marzotto, Paul Zoratti, Daniele Bagni, Andrea Colombari, and Vittorio Murino. A real-time versatile roadway path extraction and tracking on an FPGA platform. Computer Vision and Image Understanding: CVIU, 114(11):1164–1179, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Martin:2005:HIO

[MZC^{+}05] José L. Martín, Aitzol Zuloaga, Carlos Cuadrado, Jesús Lázaro, and Unai Bidarte. Hardware implementation of optical flow constraint equation using FPGAs. Computer Vision and Image Understanding: CVIU, 98(3): 462–490, June 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nitzan:1979:FMF

[NA79] David Nitzan and Gerald J.

Agin. Fast methods for finding object outlines. Computer Graphics and Image Processing, 9(1):22–39, January 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Nakamura:1984:DC

[NA84] Akira Nakamura and Kunio Aizawa. Digital circles. Computer Vision, Graphics, and Image Processing, 26(2):242–255, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nakamura:1985:DIG

[NA85] Akira Nakamura and Kunio Aizawa. Digital images of geometric pictures. Computer Vision, Graphics, and Image Processing, 30(1):107–120, April 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nakamura:1990:DS

[NA90] Akira Nakamura and Kunio Aizawa. Digital squares.

Computer Vision, Graphics, and Image Processing, 49(3):357–368, March 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nackman:1982:CRT

[Nac82] L. R. Nackman. Curvature relation in three-dimensional symmetric axes. Computer Graphics and Image Processing, 20(1):43–57, September 1982. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Nadler:1984:DSC

[Nad84]  Morton Nadler.  Docu-

ment segmentation and coding techniques. Computer Vision, Graphics, and Image Processing, 28(2):240– 262, November 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nadler:1990:NCC

[Nad90] Morton Nadler. A note on the coefficients of compass mask convolutions. Computer Vision, Graphics, and Image Processing, 51(1):96– 101, July 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nagel:1978:FOC

[Nag78] H. H. Nagel. Formation of an object concept by analysis of systematic time variations in the optically perceptible environment. Computer Graphics and Image Processing, 7 (2):149–194, April 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Nagel:1983:DVD

[Nag83] Hans-Hellmut Nagel. Displacement vectors derived from second-order intensity variations in image sequences. Computer Vision, Graphics, and Image Processing, 21 (1):85–117, January 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nagao:1986:CPP

[Nag86] Makato Nagao. Comment on the position paper “Expert Vision Systems”. Computer Vision, Graphics, and Image Processing, 34(1):104– 105, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See [Ros86b, Kov86, Tho86, Uhr86].

Nair:1987:RPB

[Nai87] Hemraj Nair. Reconstruction of planar boundaries from incomplete information. Computer Vision, Graphics, and Image Processing, 39 (3):383–387, September 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nevatia:1980:LFE

[NB80] Ramakant Nevatia and K. Ramesh Babu. Linear feature extraction and description. Computer Graphics and Image Processing, 13(3):257–269, July 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Nichau:2010:PIN

[NB10] Marco Nichau and Volker Blanz. Pose-insensitive nose detection in TOF-scans. Computer Vision and Image Understanding: CVIU, 114(12): 1346–1352, December 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nunes:2004:MER

[NBDB04] J. C. Nunes, Y. Bouaoune, E. Delechelle, and Ph. Bunel. A multiscale elastic registration scheme for retinal angiograms. Computer Vision and Image Understanding: CVIU, 95(2):129–149, August 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Natali:2011:GBR

[NBPF11] Mattia Natali, Silvia Biasotti, Giuseppe Patanè, and Bianca Falcidieno. Graphbased representations of point clouds. Graphical Models, 73 (5):151–164, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000099.

Nguyen:1993:GRF

[NC93] Hong Hai Nguyen and Paul Cohen. Gibbs random fields, fuzzy clustering, and the unsupervised segmentation of textured images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(1):1–19, January 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1001/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1001/production/
pdf.

Natsupakpong:2010:DEP

[NÇ10] Suriya Natsupakpong and M. Cenk Çavusoğlu. Determination of elasticity parameters in lumped element (massspring) models of deformable objects. Graphical Models, 72(6):61–73, November 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070310000147.

Narayanan:1992:RDA

[ND92] P. J. Narayanan and Larry S.

Davis. Replicated data algorithms in image processing. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(3): 351–365, November 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Nilsson:1997:FMS

[ND97] Frederik Nilsson and Per-Erik Danielsson. Finding the minimal set of maximum disks for binary objects. Graphical Models and Image Processing: GMIP, 59(1):55–60, January 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0412/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0412/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1996.
0412/production/ref.

Nesi:1995:RAO

[NDBT95] P. Nesi, A. Del Bimbo, and D. Ben-Tzvi. A robust algorithm for optical flow estimation. Computer Vision and Image Understanding: CVIU, 62(1):59–68, July 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1041/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1041/production/
pdf.

Neveu:1986:TDO

[NDC86] Charles F. Neveu, Charles R.

Dyer, and Roland T. Chin. Two-dimensional object recognition using multiresolution models. Computer Vision, Graphics, and Image Processing, 34(1):52–65, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Niessen:1997:MAI

[NDN^{+}97] W. J. Niessen, J. S. Duncan, M. Nielsen, L. M. J. Florack, B. M. ter Haar Romeny, and M. A. Viergever. A multiscale approach to image sequence analysis. Computer Vision and Image Understanding: CVIU, 65(2):259–268, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0582/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0582/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0582/production/ref.

Noceti:2009:STC

[NDO09] Nicoletta Noceti, Elisabetta Delponte, and Francesca Odone. Spatio-temporal constraints for on-line 3D object recognition in videos. Computer Vision and Image Understanding: CVIU, 113(12): 1198–1209, December 2009.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Negahdaripour:1996:DCF

[Neg96] Shahriar Negahdaripour. Direct computation of the FOE with confidence measures. Computer Vision and Image Understanding: CVIU, 64(3):323–350, November 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0063/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0063/production/
                                        pdf.

Negahdaripour:2012:VMA

[Neg12] S. Negahdaripour. Visual motion ambiguities of a plane in 2-D FS sonar motion sequences. Computer Vision and Image Understanding: CVIU, 116(6):754–764, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000379.

Nuchter:2010:SPR

[NESP10] Andreas Nüchter, Jan Elseberg, Peter Schneider, and Dietrich Paulus. Study of parameterizations for the rigid body transformations of the scan registration problem. Computer Vision and Image Understanding: CVIU, 114(8):963–980, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nevatia:1976:DMM

[Nev76] Ramakant Nevatia. Depth measurement by motion stereo. Computer Graphics and Image Processing, 5(2): 203–214, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Neycenssac:1993:CEU

[Ney93] Franck Neycenssac. Contrast enhancement using the Laplacian-of-a-Gaussian filter. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(6):447– 463, November 1993. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1034/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1034/production/
pdf.

Neff:2006:MEE

[NF06] Michael Neff and Eugene Fiume. Methods for exploring expressive stance. Graphical Models, 68(2):133–157, March 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000329.

Neumann:2004:HCP

[NFA04] Jan Neumann, Cornelia Fermüller, and Yiannis Aloimonos. A hierarchy of cameras for 3D photography. Computer Vision and Image Understanding: CVIU, 96(3): 274–293, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Newman:1993:MBC

[NFJ93] Timothy S. Newman, Patrick J. Flynn, and Anil K. Jain. Model-based classification of quadric surfaces. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(2):235–249, September 1993. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1040/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1040/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1042/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1042/production/
pdf.

Nascimento:2008:IIP

[NFM08] J. Nascimento, M. Figueiredo, and J. Marques. Independent increment processes for human motion recognition. Computer Vision and Image Understanding: CVIU, 109 (2):126–138, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nguyen:2013:FDS

[NFSD13] Kien Nguyen, Clinton Fookes, Sridha Sridharan, and Simon Denman. Feature-domain super-resolution for iris recognition. Computer Vision and Image Understanding: CVIU, 117(10):1526–1535, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001306.

Neuenschwander:1997:VSF

[NFSK97] W. Neuenschwander, P. Fua, G. Székely, and O. Kübler. Velcro surfaces: Fast initialization of deformable models. Computer Vision and Image Understanding: CVIU, 65 (2):237–245, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0578/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0578/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0578/production/ref.

Nyul:2002:FCI

[NFU02] László G. Nyúl, Alexandre X. Falcão, and Jayaram K. Udupa. Fuzzy-connected 3D image segmentation at interactive speeds. Graphical Models, 64(5):259–281, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Nagao:1998:AMP

[NG98a] Kenji Nagao and W. E. L. Grimson. Affine matching of planar sets. Computer Vision and Image Understanding: CVIU, 70(1):1–22, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0623/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0623/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0623/production/ref.

Nagao:1998:UPI

[NG98b] Kenji Nagao and W. Eric. L. Grimson. Using photometric invariants for 3D object recognition. Computer Vision and Image Understanding: CVIU, 71(1):74–93, July 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0603/production;   http:
                                        //www.idealibrary.com/links/

artid/cviu.1997.0603/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0603/production/ref.

Niblack:1992:GSC

[NGC92] C. Wayne Niblack, Phillip B.

Gibbons, and David W. Capson. Generating skeletons and centerlines from the distance transform. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(5):420– 437, September 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Negahdaripour:1989:DML

[NH89] Shahriar Negahdaripour and Berthold K. P. Horn. A direct method for locating the focus of expansion. Computer Vision, Graphics, and Image Processing, 46(3):303– 326, June 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nurre:1992:EGG

[NH92] Joseph H. Nurre and Ernest L.

Hall. Encoded grid generation from a computer solid model. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (2):131–138, September 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Nam:2008:SBL

[NHK08] Yunyoung Nam, Eenjun Hwang, and Dongyoon Kim. A similarity-based leaf image retrieval scheme: Joining shape and venation features. Computer Vision and Image Understanding: CVIU, 110 (2):245–259, May 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nagin:1981:VRL

[NHR81] Paul A. Nagin, Allen R. Hanson, and Edward M. Riseman. Variations in relaxation labeling techniques. Computer Graphics and Image Processing, 17(1):33– 51, September 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ni:2009:UMS

[NHSC09] Kangyu Ni, Byung-Woo Hong, Stefano Soatto, and Tony Chan. Unsupervised multiphase segmentation: a recursive approach. Computer Vision and Image Understanding: CVIU, 113(4): 502–510, April 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nejhum:2010:OVT

[NHY10] S. M. Shahed Nejhum, Jeffrey Ho, and Ming-Hsuan Yang. Online visual tracking with histograms and articulating blocks. Computer Vision and Image Understanding: CVIU, 114(8):901–914, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nishihara:1982:FCR

[NI82] Seiichi Nishihara and Katsuo Ikeda. False-contour removal by random blurring. Computer Graphics and Image Processing, 20(4):391– 397, December 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Nicol:1995:SAR

[Nic95] C. J. Nicol. Systolic approach for real time connected component labeling. Computer Vision and Image Understanding: CVIU, 61(1):17–31, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1002/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1002/production/
pdf.

Nishida:1995:SFE

[Nis95] Hirobumi Nishida. Structural feature extraction using multiple bases. Computer Vision and Image Understanding: CVIU, 62(1):78–89, July 1995. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1043/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1043/production/
pdf.

Nishida:1996:SRI

[Nis96a] Hirobumi Nishida. Shape recognition by integrating structural descriptions and geometrical/statistical transforms. Computer Vision and Image Understanding: CVIU, 64(2):248–262, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0057/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0057/production/
pdf.

Nishida:1996:SMC

[Nis96b] Hirobumi Nishida. A structural model of curve deformation by discontinuous transformations. Graphical Models and Image Processing: GMIP, 58(2):164–179, March 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0014/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0014/production/
pdf.

Nishida:1997:ASD

[Nis97] Hirobumi Nishida. Analysis and synthesis of deformed patterns based on structural models. Computer Vision and Image Understanding: CVIU, 68(1):59–71, October 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0541/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0541/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0541/production/ref.

Nishida:1998:BEG

[Nis98] Hirobumi Nishida. Boundary extraction from grayscale document images based on surface data structures. Graphical Models and Image Processing: GMIP, 60 (1):035–045, January 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0452/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0452/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0452/production/ref.

Nishida:1999:SSI

[Nis99] Hirobumi Nishida. Structural shape indexing with feature generation models.

Computer Vision and Image Understanding: CVIU, 73(1):121–136, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0712/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0712/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0712/production/ref.

Newman:1995:SAV

[NJ95] Timothy S. Newman and Anil K. Jain. A survey of automated visual inspection. Computer Vision and Image Understanding: CVIU, 61(2):231–262, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1017/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1017/production/
pdf.

Negahdaripour:2000:MBC

[NK00] S. Negahdaripour and A. Khamene. Motion-based compression of underwater video imagery for the operations of unmanned submersible vehicles. Computer Vision and Image Understanding: CVIU, 79(1):162–183, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0845;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0845/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0845/ref.

Noris:2011:WGT

[NKB11] Basilio Noris, Jean-Baptiste Keller, and Aude Billard. A wearable gaze tracking system for children in unconstrained environments. Computer Vision and Image Understanding: CVIU, 115(4):476–486, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nemeth:2011:TCI

[NKP11] Gábor Németh, Péter Kardos, and Kálmán Palágyi. Thinning combined with iteration-by-iteration smoothing for 3D binary images. Graphical Models, 73(6):335– 345, November 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000063.

Ngo:2013:CSR

[NKPT13] Phuc Ngo, Yukiko Kenmochi, Nicolas Passat, and Hugues Talbot. Combinatorial structure of rigid transformations in 2D digital images. Computer Vision and Image Understanding: CVIU, 117(4):393–408, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001786.

Nagasamy:1990:EDP

[NL90] Vijay Nagasamy and Noshir A. Langrana. Engineering drawing processing and vectorization system. Computer Vision, Graphics, and Image Processing, 49(3):379– 397, March 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nguyen:1996:RDO

[NL96] Quang-Loc Nguyen and Martin D. Levine. Representing 3-D objects in range images using geons. Computer Vision and Image Understanding: CVIU, 63(1):158–168, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0011/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0011/production/
                                        pdf.

Noureddin:2005:NCD

[NLM05] B. Noureddin, P. D. Lawrence, and C. F. Man. A noncontact device for tracking gaze in a human computer interface. Computer Vision and Image Understanding: CVIU, 98(1):52–82, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lee:1991:ROR

[nLPR91] Chung nim Lee, T. Poston, and A. Rosenfeld. Representation of orthogonal regions by vertices. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2): 149–156, March 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Niessen:2013:SIS

[NLW13] Wiro Niessen, Shuo Li, and Song Wang. Special issue on Shape Modeling in Medical Image Analysis. Computer Vision and Image Understanding: CVIU, 117 (9):965, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001227.

Nagao:1979:EPS

[NM79] Makoto Nagao and Takashi Matsuyama. Edge preserving smoothing. Computer Graphics and Image Processing, 9 (4):394–407, April 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Nagao:1979:RES

[NMI79] Makoto Nagao, Takashi Matsuyama, and Yoshio Ikeda. Region extraction and shape analysis in aerial photographs. Computer Graphics and Image Processing, 10(3):195– 223, July 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). NAGAO79a.

Nastar:1997:FIM

[NMP97] Chahab Nastar, Baback Moghaddam, and Alex Pentland. Flexible images: Matching and recognition using learned deformations. Computer Vision and Image Understanding: CVIU, 65 (2):179–191, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0583/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0583/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0583/production/ref.

Nikolaev:2004:LCS

[NN04] Dmitry P. Nikolaev and Petr P. Nikolayev. Linear color segmentation and its implementation. Computer Vision and Image Understanding: CVIU, 94(1–3):115– 139, April/June 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Natarajan:2013:HMC

[NN13] Pradeep Natarajan and Ramakant Nevatia. Hierarchical multi-channel hidden semi Markov graphical models for activity recognition. Computer Vision and Image Understanding: CVIU, 117(10):1329–1344, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421200166X.

Nitschke:2011:DCC

[NNT11] Christian Nitschke, Atsushi Nakazawa, and Haruo Takemura. Display-camera calibration using eye reflections and geometry constraints. Computer Vision and Image Understanding: CVIU, 115(6): 835–853, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nordberg:2009:TT

[Nor09] Klas Nordberg. The triangulation tensor. Computer Vision and Image Understanding: CVIU, 113(9):935– 945, September 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nalwa:1987:EAE

[NP87] Vishvjit S. Nalwa and Eric Pauchon. Edgel aggregation and edge description. Computer Vision, Graphics, and Image Processing, 40(1):79–94, October 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nelson:1992:QRM

[NP92] Randal C. Nelson and Ramprasad Polana. Qualitative recognition of motion using temporal texture. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):78–89, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Naito:1988:SRP

[NR88a] Seiichiro Naito and Azriel Rosenfeld. Shape from random planar features. Computer Vision, Graphics, and Image Processing, 42(3):345– 370, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Nirmal:1988:PGD

[NR88b] Nalinakshi Nirmal and R. Roma. Picture generation and developmental matrix systems. Computer Vision, Graphics, and Image Processing, 43(1): 67–80, July 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Narappanawar:2011:GTB

[NRJ11] Nitin Narappanawar, B. Madhusudan Rao, and Maduri Joshi. Graph theory based segmentation of traced boundary into open and closed sub-sections. Computer Vision and Image Understanding: CVIU, 115(11): 1552–1558, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001676.

Nielsen:1991:PAI

[NS91] Lars Nielsen and Gunnar Sparr. Projective areainvariants as an extension of the cross ratio. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(1):145–159, July 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Nogly:1996:DTG

[NS96] Daniel Nogly and Markus Schladt. Digital topology on graphs. Computer Vision and Image Understanding: CVIU, 63(2):394–396, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0029/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0029/production/
                                        pdf.

Noordmans:1998:DCI

[NS98] H. J. Noordmans and A. W. M. Smeulders. Detection and characterization of isolated and overlapping spots. Computer Vision and Image Understanding: CVIU, 70(1): 23–35, April 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/artid/
                                        cviu.1998.0604/production;
                                         http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0604/production/pdf;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0604/production/
                                        ref.

Normand:2013:MDD

[NSEA13] Nicolas Normand, Robin Strand, Pierre Evenou, and Aurore Arlicot. Minimaldelay distance transform for neighborhood-sequence distances in 2D and 3D. Computer Vision and Image Understanding: CVIU, 117(4): 409–417, April 2013. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001798.

Naf:1997:VST

[NSK^{+}97] M. Näf, G. Székely, R. Kikinis, M. E. Shenton, and O. Kübler. 3D Voronoi skeletons and their usage for the characterization and recognition of 3D organ shape. Computer Vision and Image Understanding: CVIU, 66(2):147–161, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0610/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0610/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0610/production/ref.

Negahdaripour:2010:DME

[NT10] S. Negahdaripour and Ali Taatian. 3-D motion estimation by integrating visual cues in 2-D multi-modal optiacoustic stereo sequences. Computer Vision and Image Understanding: CVIU, 114(8):928–941, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Nurmi:1986:TSO

[Nur86]  Otto Nurmi.  On trans-

lating a set of objects in 2- and 3-dimensional space. Computer Vision, Graphics, and Image Processing, 36(1):42–52, October 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Niessen:1997:NMR

[NVWV97] Wiro J. Niessen, Koen L. Vincken, Joachim A. Weickert, and Max A. Viergever. Nonlinear multiscale representations for image segmentation. Computer Vision and Image Understanding: CVIU, 66(2):233–245, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0614/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0614/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0614/production/ref.

Noble:1997:CAG

[NWP97] Alison Noble, Dale Wilson, and Jean Ponce. On computing aspect graphs of smooth shapes from volumetric data. Computer Vision and Image Understanding: CVIU, 66(2):179–192, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0615/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0615/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0615/production/ref.

Orrite:2004:ESM

[OBH04] Carlos Orrite, Sergio Blecua, and J. Elías Herrero. Erratum to “Shape matching of partially occluded curves invariant under projective transformation” [Comp. Vision Image Understanding 93 (2004) 34– 64]. Computer Vision and Image Understanding: CVIU, 95(1):127, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [OH04].

Ohtake:2005:SDI

[OBS05] Yutaka Ohtake, Alexander Belyaev, and Hans-Peter Seidel. 3D scattered data interpolation and approximation with multilevel compactly supported RBFs. Graphical Models, 67(3):150–165, May 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ohtake:2006:CAM

[OBS06a] Yutaka Ohtake, Alexander Belyaev, and Hans-Peter Seidel. A composite approach to meshing scattered data. Graphical Models, 68 (3):255–267, May 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000221.

Ohtake:2006:SSR

[OBS06b] Yutaka Ohtake, Alexander Belyaev, and Hans-Peter Seidel. Sparse surface reconstruction with adaptive partition of unity and radial basis functions. Graphical Models, 68(1):15–24, January 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000548.

ORourke:1987:CDN

[OBW87] Joseph O’Rourke, Heather Booth, and Richard Washington. Connect-the-dots: a new heuristic. Computer Vision, Graphics, and Image Processing, 39(2):258– 266, August 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Oakley:1990:FSM

[OC90] J. P. Oakley and M. J. Cunningham. A function space model for digital image sampling and its application in image reconstruction. Computer Vision, Graphics, and Image Processing, 49 (2):171–197, February 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

ORourke:1982:NLA

[OCON82] Joseph O’Rourke, Chi Bin Chien, Thomas Olson, and David Naddor. New linear algorithm for intersecting convex polygons. Computer Graphics and Image Processing, 19(4):384– 391, August 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ozden:2004:RTI

[OCVV04] Kemal Egemen Ozden, Kurt Cornelis, Luc Van Eycken, and Luc Van Gool. Reconstructing 3D trajectories of independently moving objects using generic constraints. Computer Vision and Image Understanding: CVIU, 96(3): 453–471, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Okatani:1997:SRE

[OD97] Takayuki Okatani and Koichiro Deguchi. Shape reconstruction from an endoscope image by shape from shading technique for a point light source at the projection center. Computer Vision and Image Understanding: CVIU, 66(2):119–131, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0613/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0613/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0613/production/ref.

Okatani:1999:CSG

[OD99] Takayuki Okatani and Koichiro Deguchi. Computation of the sign of the Gaussian curvature of a surface from multiple unknown illumination images without knowledge of the reflectance property. Computer Vision and Image Understanding: CVIU, 76(2):125–134, November 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0792/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0792/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0792/production/ref.

Okatani:2001:UST

[OD01] Takayuki Okatani and Koichiro Deguchi. On uniqueness of solutions of the three-light-source photometric stereo: Conditions on illumination configuration and surface reflectance. Computer Vision and Image Understanding: CVIU, 81(2):211–226, February 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0887;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0887/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0887/ref.

Okatani:2002:MFR

[OD02] Ikuko Shimizu Okatani and Koichiro Deguchi. A method for fine registration of multiple view range images considering the measurement error properties. Computer Vision and Image Understanding: CVIU, 87(1–3):66–77, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Oberkampf:1996:IPE

[ODD96] Denis Oberkampf, Daniel F.

DeMenthon, and Larry S. Davis. Iterative pose estimation using coplanar feature points. Computer Vision and Image Understanding: CVIU, 63(3):495–511, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0037/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0037/production/
pdf.

Olsson:2008:ISR

[OEK08] Carl Olsson, Anders P. Eriksson, and Fredrik Kahl. Improved spectral relaxation methods for binary quadratic optimization problems. Computer Vision and Image Understanding: CVIU, 112(1): 3–13, October 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

OGorman:1988:NHE

[O’G88]  Lawrence O’Gorman.  A

note on histogram equalization for optimal intensity range utilization. Computer Vision, Graphics, and Image Processing, 41(2):229– 232, February 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

OGorman:1990:T

[O’G90] Lawrence O’Gorman. k × k thinning. Computer Vision, Graphics, and Image Processing, 51(2):195–215, August 1990. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

OGorman:1994:BMD

[O’G94]  L. O’Gorman.  Binariza-

tion and multithresholding of document images using connectivity. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(6):494–506, November 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1044/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1044/production/
                                        pdf.

Oblonsek:1998:FSB

[OG98] C. Oblonsek and N. Guid.

A fast surface-based procedure for object reconstruction from 3D scattered points. Computer Vision and Image Understanding: CVIU, 69(2):185–195, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0584/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0584/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0584/production/ref.

Oliver:2004:LRL

[OGH04] Nuria Oliver, Ashutosh Garg, and Eric Horvitz. Layered representations for learning and inferring office activity from multiple sensory channels. Computer Vision and Image Understanding: CVIU, 96(2):163–180, November 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

OConnor:1981:APU

[OH81] B. T. O’Connor and T. S.

Huang. Application of phase unwrapping to image restoration. Computer Graphics and Image Processing, 15(1): 25–44, January 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Orrite:2004:SMP

[OH04] Carlos Orrite and J. Elias Herrero. Shape matching of partially occluded curves invariant under projective transformation. Computer Vision and Image Understanding: CVIU, 93(1):34–64, January 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See erratum [OBH04].

Oliver:2005:SPP

[OH05] Nuria Oliver and Eric Horvitz. Selective perception policies for guiding sensing and computation in multimodal systems: a comparative analysis. Computer Vision and Image Understanding: CVIU, 100(1–2):198–224, October/ November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ong:2006:LIK

[OH06] Eng-Jon Ong and Adrian Hilton. Learnt inverse kinematics for animation synthesis. Graphical Models, 68(5–6):472–483, September/November 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070306000609.

Osadchy:2008:USC

[OJRT08] Margarita Osadchy, David Jacobs, Ravi Ramamoorthi, and David Tucker. Using specularities in comparing 3D models and 2D images. Computer Vision and Image Understanding: CVIU, 111(3):275– 294, September 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Osadchy:2004:EDU

[OK04] Margarita Osadchy and Daniel Keren. Efficient detection under varying illumination conditions and image plane rotations. Computer Vision and Image Understanding: CVIU, 93(3):245–259, March 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Oh:2007:DPF

[OK07] Seungtaik Oh and Bon Ki Koo. Data perturbation for fewer triangles in marching tetrahedra. Graphical Models, 69(3–4):211–218, May/July 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070307000148.

Okawa:1981:CMC

[Oka81]  Yoshikuni Okawa.  Com-

plexity measure for colored pictures in commercial design. Computer Graphics and Image Processing, 17(4):345– 361, December 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Okawa:1984:AIS

[Oka84]  Yoshikuni Okawa.  Auto-

matic inspection of the surface defects of cast metals. Computer Vision, Graphics, and Image Processing, 25(1):89–112, January 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Okawa:1988:SAV

[Oka88] Yoshikuni Okawa. Structural analysis of visual form on packaging graphics and its use in an automated design system. Computer Vision, Graphics, and Image Processing, 43(2):265– 278, August 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

OLeary:1988:SAA

[O’L88] Dianne P. O’Leary. Some algorithms for approximating convolutions. Computer Vision, Graphics, and Image Processing, 41(3):333– 345, March 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Oliensis:1991:SSP

[Oli91] J. Oliensis. Shape from (or and??) shading as a partially well-constrained problem. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (2):163–183, September 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Olivo:1994:ATS

[Oli94]  J. C. Olivo.  Automatic

threshold selection using the wavelet transform. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(3):205–218, May 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1019/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1019/production/
pdf.

Oliensis:2000:CSM

[Oli00] John Oliensis. A critique of structure-from-motion algorithms. Computer Vision and Image Understanding: CVIU, 80(2):172–214, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0869;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0869/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0869/ref. See corrigen-

dum [Oli01].

Oliensis:2001:CSM

[Oli01] John Oliensis. A critique of structure-from-motion algorithms: Volume 80, Number 2 (2000), pages 172–214. Computer Vision and Image Understanding: CVIU, 84(3): 407–408, December 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [Oli00].

Olkkonen:1995:DBS

[Olk95] Hannu Olkkonen. Discrete binomial splines. Graphical Models and Image Processing: GMIP, 57(2):101–106, March 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1011/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1011/production/
                                        pdf.

Olsen:1993:ENI

[Ols93]  S. I. Olsen.  Estimation

of noise in images: An evaluation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(4):319–323, July 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1993.
                                        1022/production;   http:
                                        //www.idealibrary.com/links/

artid/cgip.1993.1022/production/
pdf.

Olson:1999:CHT

[Ols99] Clark F. Olson. Constrained Hough transforms for curve detection. Computer Vision and Image Understanding: CVIU, 73(3):329–345, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0728/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0728/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0728/production/ref.

OCallaghan:1984:EDN

[OM84] John F. O’Callaghan and David M. Mark. The extraction of drainage networks from digital elevation data. Computer Vision, Graphics, and Image Processing, 28 (3):323–344, December 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ong:2006:VIE

[OMBH06] Eng-Jon Ong, Antonio S.

Micilotta, Richard Bowden, and Adrian Hilton. Viewpoint invariant exemplarbased 3D human tracking. Computer Vision and Image Understanding: CVIU, 104(2–3):178–189, November/December 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ogier:1998:MAD

[OMLL98] Jean Marc Ogier, Rémy Mullot, Jacques Labiche, and Yves Lecourtier. Multilevel approach and distributed consistency for technical map interpretation: Application to cadastral maps. Computer Vision and Image Understanding: CVIU, 70(3):438–451, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0690/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0690/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0690/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0708/production;
 http://www.idealibrary.
com/links/artid/cviu.1998.
0708/production/pdf.

Olson:2007:VTM

[OMW^{+}07] Clark F. Olson, Larry H. Matthies, John R. Wright, Rongxing Li, and Kaichang Di. Visual terrain mapping for Mars exploration. Computer Vision and Image Understanding: CVIU, 105(1): 73–85, January 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Olkkonen:1996:GPW

[OP96] H. Olkkonen and P. Pesola.

Gaussian pyramid wavelet transform for multiresolution analysis of images. Graphical Models and Image Processing: GMIP, 58(4):394–398, July 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0032/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0032/production/
pdf.

Ohlander:1978:PSU

[OPR78] Ron Ohlander, Keith Price, and D. Raj Reddy. Picture segmentation using a recursive region splitting method. Computer Graphics and Image Processing, 8(3):313– 333, December 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Otto:1981:CHP

[OR81] G. P. Otto and D. E.

Reynolds. Counting hardware for parallel processors. Computer Graphics and Image Processing, 17(2):185– 186, October 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

ORourke:1982:PDS

[O’R82] Joseph O’Rourke. Polygon decomposition and switching function minimization. Computer Graphics and Image Processing, 18(4):382– 391, April 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

ORourke:1985:CMC

[O’R85] Joseph O’Rourke. Counterexamples to a minimal circumscription algorithm. Computer Vision, Graphics, and Image Processing, 30(3):364– 366, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

ORourke:1994:SHR

[O’R94] J. O’Rourke. On the scaling heuristic for reconstruction from slices. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(5):420–423, September 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1038/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1038/production/
pdf.

Oliviero:1981:NAC

[OS81] A. Oliviero and G. Scarpetta. A new approach to contour coding. Computer Graphics and Image Processing, 15(1): 87–92, January 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

OGorman:1987:CMC

[OS87] Lawrence O’Gorman and Arthur C. Sanderson. A comparison of methods and computation for multi-resolution low- and band-pass transforms for image processing. Computer Vision, Graphics, and Image Processing, 37(3):386–401, March 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Oishi:1995:TOD

[OS95] Yasuaki Oishi and Kokichi Sugihara. Topology-oriented divide-and-conquer algorithm for Voronoi diagrams. Graphical Models and Image Processing: GMIP, 57(4):303–314, July 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1027/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1027/production/
pdf.

Olson:1996:PPI

[OTL96] Thomas J. Olson, John R.

Taylor, and Robert J. Lockwood. Programming a pipelined image processor. Computer Vision and Image Understanding: CVIU, 64 (3):351–367, November 1996.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0064/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0064/production/
                                        pdf.

Ogiela:2006:GIL

[OTO06] Marek R. Ogiela, Ryszard Tadeusiewicz, and Lidia Ogiela. Graph image language techniques supporting radiological, hand image interpretations. Computer Vision and Image Understanding: CVIU, 103(2):112–120, August 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ottmann:1983:TSL

[OW83] Thomas Ottmann and Peter Widmayer. On translating a set of line segments. Computer Vision, Graphics, and Image Processing, 24 (3):382–390, December 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ottmann:1984:DSP

[OW84] Thomas Ottmann and Derick Wood. Dynamical sets of points. Computer Vision, Graphics, and Image Processing, 27(2):157–166, August 1984. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ottmann:1986:SEP

[OW86] Thomas Ottmann and Derick Wood. Space-economical plane-sweep algorithms. Computer Vision, Graphics, and Image Processing, 34(1):35– 51, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ottmann:1985:FAB

[OWW85] Thomas Ottmann, Peter Widmayer, and Derick Wood. A fast algorithm for the Boolean masking problem. Computer Vision, Graphics, and Image Processing, 30(3):249–268, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ouksel:1992:IBB

[OY92] M. Aris Ouksel and Anan Yaagoub. The interpolationbased bintree and encoding of binary images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1): 75–81, January 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Onoe:1998:TRT

[OYTY98] Yoshio Onoe, Kazumasa Yamazawa, Haruo Takemura, and Naokazu Yokoya. Telepresence by real-time viewdependent image generation from omnidirectional video streams. Computer Vision and Image Understanding: CVIU, 71(2):154–165, August 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0705/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0705/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0705/production/ref.

Posdamer:1982:SMS

[PA82] J. L. Posdamer and M. D. Altschuler. Surface measurement by space-encoded projected beam systems. Computer Graphics and Image Processing, 18(1):1–17, January 1982. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Prager:1983:COF

[PA83] John M. Prager and Michael A. Arbib. Computing the optic flow: the match algorithm and prediction. Computer Vision, Graphics, and Image Processing, 24(3):271– 304, December 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pikaz:1997:ETC

[PA97] Arie Pikaz and Amir Averbuch. An efficient topological characterization of gray-levels textures, using a multiresolution representation. Graphical Models and Image Processing: GMIP, 59(1):1–17, January 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0410/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0410/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1996.
0410/production/ref.

Pikaz:1998:RBS

[PA98] Arie Pikaz and Amir Averbuch. On the relation between second-order statistics. connectivity analysis, and percolation models in digital textures. Graphical Models and Image Processing: GMIP, 60(3):226–232, May 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0462/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0462/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0462/production/ref.

Papamarkos:2000:GLR

[PA00] Nikos Papamarkos and Antonios Atsalakis. Gray-level reduction using local spatial features. Computer Vision and Image Understanding: CVIU, 78(3):336–350, June 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0838;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0838/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0838/ref.

Park:2006:STM

[PA06] Sangho Park and J. K. Aggarwal. Simultaneous tracking of multiple body parts of interacting persons. Computer Vision and Image Understanding: CVIU, 102(1):1–21, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Papadourakis:2010:MOT

[PA10a] Vasilis Papadourakis and Antonis Argyros. Multiple objects tracking in the presence of long-term occlusions. Computer Vision and Image Understanding: CVIU, 114(7): 835–846, July 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Phan:2010:CBR

[PA10b] Raymond Phan and Dimitrios Androutsos. Contentbased retrieval of logo and trademarks in unconstrained color image databases using Color Edge Gradient Cooccurrence Histograms. Computer Vision and Image Understanding: CVIU, 114(1): 66–84, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Perdigoto:2013:CMP

[PA13] Luis Perdigoto and Helder Araujo. Calibration of mirror position and extrinsic parameters in axial noncentral catadioptric systems. Computer Vision and Image Understanding: CVIU, 117(8):909–921, August 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000635.

Pizer:1987:AHE

[PAA^{+}87] Stephen M. Pizer, E. Philip Amburn, John D. Austin, Robert Cromartie, Ari Geselowitz, Trey Greer, Bart ter Haar Romeny, John B. Zimmerman, and Karel Zuiderveld. Adaptive histogram equalization and its variations. Computer Vision, Graphics, and Image Processing, 39 (3):355–368, September 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Paglieroni:1992:DTP

[Pag92] David W. Paglieroni. Distance transforms: Properties and machine vision applications. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1): 56–74, January 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Paglieroni:1997:DDT

[Pag97] David W. Paglieroni. Directional distance transforms and height field preprocessing for efficient ray tracing. Graphical Models and Image Processing: GMIP, 59(4):253–264, July 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0434/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0434/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0434/production/ref.

Paglieroni:1999:CAD

[Pag99] David W. Paglieroni. A complexity analysis for directional parametric height field ray tracing. Graphical Models and Image Processing: GMIP, 61(5):299–321, September 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0503/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0503/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0503/production/ref.

Panda:1978:NSP

[Pan78a] Durga P. Panda. Nonlinear smoothing of pictures. Computer Graphics and Image Processing, 8(2):259– 270, October 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Panda:1978:SPT

[Pan78b] Durga P. Panda. Statistical properties of thresholded images. Computer Graphics and Image Processing, 8(3): 334–354, December 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Pandzic:2003:FMC

[Pan03] Igor S. Pandzic. Facial motion cloning. Graphical Models, 65(6):385–404, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Pasko:2001:CHM

[PASS01] Alexander Pasko, Valery Adzhiev, Benjamin Schmitt, and Christophe Schlick. Constructive hypervolume modeling. Graphical Models, 63(6):

413–442, November 2001. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Paton:1979:LDL

[Pat79] Keith Paton. Line detection by local methods. Computer Graphics and Image Processing, 9(4):316–332, April 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Patane:2013:MRS

[Pat13]  Giuseppe Patanè.  Multi-

resolutive sparse approximations of d-dimensional data. Computer Vision and Image Understanding: CVIU, 117(4):418–428, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001816.

Pavlidis:1978:RAS

[Pav78]  Theodosios Pavlidis.  Re-

view of algorithms for shape analysis. Computer Graphics and Image Processing, 7 (2):243–258, April 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Pavlidis:1979:FAR

[Pav79] Theo Pavlidis. Filling algorithms for raster graphics. Computer Graphics and Image Processing, 10(2):

126–141, June 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Pavlidis:1980:TAD

[Pav80] Theo Pavlidis. Thinning algorithm for discrete binary images. Computer Graphics and Image Processing, 13(2): 142–157, June 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Pavlidis:1986:VFE

[Pav86] Theo Pavlidis. A vectorizer and feature extractor for document recognition. Computer Vision, Graphics, and Image Processing, 35(1):111– 127, July 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pnueli:1996:GHR

[PB96] Yachin Pnueli and Alfred M.

Bruckstein. Gridless halftoning: a reincarnation of the old method. Graphical Models and Image Processing: GMIP, 58(1):38–64, January 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0003/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0003/production/
pdf.

Puzicha:1999:MAG

[PB99] Jan Puzicha and Joachim M. Buhmann. Multiscale annealing for grouping and unsupervised texture segmentation. Computer Vision and Image Understanding: CVIU, 76(3):213–230, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0805/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0805/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0805/production/ref.

Papazov:2011:SGO

[PB11] Chavdar Papazov and Darius Burschka. Stochastic global optimization for robust point set registration. Computer Vision and Image Understanding: CVIU, 115(12):1598–1609, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001561.

Prasad:2004:RBI

[PBG04] B. G. Prasad, K. K. Biswas, and S. K. Gupta. Regionbased image retrieval using integrated color, shape, and location index. Computer Vision and Image Understanding: CVIU, 94(1–3):193– 233, April/June 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pereira:2011:SBW

[PBM^{+}11] Thiago Pereira, Emilio Vital Brazil, Ives Macêdo, Mario Costa Sousa, Luiz Henrique de Figueiredo, and Luiz Velho. Sketch-based warping of RGBN images. Graphical Models, 73(4):97–110, July 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031000041X.

Pradal:2009:PPB

[PBN^{+}09] C. Pradal, F. Boudon, C. Nouguier, J. Chopard, and C. Godin. PlantGL: a Pythonbased geometric library for 3D plant modelling at different scales. Graphical Models, 71(1):1–21, January 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000143.

Peng:1999:PFR

[PBQ99] Jing Peng, Bir Bhanu, and Shan Qing. Probabilistic feature relevance learning for content-based image retrieval. Computer Vision and Image Understanding: CVIU, 75 (1–2):150–164, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0770/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0770/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0770/production/ref.

Puig:2012:COC

[PBSG12] Luis Puig, J. Bermúdez, Peter Sturm, and J. J. Guerrero. Calibration of omnidirectional cameras in practice: a comparison of methods. Computer Vision and Image Understanding: CVIU, 116 (1):120–137, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001858.

Popescu:2006:M

[PBSM06] V. Popescu, G. Bahmutov, E. Sacks, and M. Mudure. The ModelCamera. Graphical Models, 68(5–6):385–401, September/November 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070306000397.

Park:1989:NIC

[PC89] Rae-Hong Park and Woo Young Choi. A new interpretation of the compass gradient edge operators. Computer Vision, Graphics, and Image Processing, 47(2):259– 265, August 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pearce:1999:IMH

[PC99] Adrian R. Pearce and Terry Caelli. Interactively matching hand-drawings using induction. Computer Vision and Image Understanding: CVIU, 73(3):391–403, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0742/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0742/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0742/production/ref.

Parameswaran:2005:HAR

[PC05] Vasu Parameswaran and Rama Chellappa. Human action-recognition using mutual invariants. Computer Vision and Image Understanding: CVIU, 98(2):294–324, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Perri:2013:ACT

[PCC13] Stefania Perri, Pasquale Corsonello, and Giuseppe Cocorullo. Adaptive Census Transform: a novel hardwareoriented stereovision algorithm. Computer Vision and Image Understanding: CVIU, 117(1):29–41, January 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001336.

Paparoditis:1998:BDR

[PCJC98] N. Paparoditis, M. Cord, M. Jordan, and J.-P. Cocquerez. Building detection and reconstruction from midand high-resolution aerial imagery. Computer Vision and Image Understanding: CVIU, 72(2):122–142, November 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0722/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0722/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0722/production/ref.

Plante:2002:CCH

[PCP02] Eric Plante, Marie-Paule Cani, and Pierre Poulin. Capturing the complexity of hair motion. Graphical Models, 64 (1):40–58, January 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Phillips:1986:DAT

[PCR86] Tsaiyun Phillips, Robert Cannon, and Azriel Rosenfeld. Decomposition and approximation of three-dimensional solids. Computer Vision, Graphics, and Image Processing, 33(3):307–317, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Provost:2004:HMS

[PCR^{+}04] J.-N. Provost, C. Collet, P. Rostaing, P. Pérez, and P. Bouthemy. Hierarchical Markovian segmentation of multispectral images for the reconstruction of water depth maps. Computer Vision and Image Understanding: CVIU, 93(2):155–174, February 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Poli:1994:RCS

[PCV94] Riccardo Poli, Giuseppe Coppini, and Guido Valli. Recovery of 3D closed surfaces from sparse data. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):1–25, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1028/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1028/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1033/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1033/production/
pdf.

Panda:1979:SAS

[PD79] Durga P. Panda and Tsvi Dubitzki. Statistical analysis of some edge operators. Computer Graphics and Image Processing, 11(4):313– 348, December 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Paquin:1983:STG

[PD83] R. Paquin and E. Dubois. Spatio-temporal gradient method for estimating the displacement field in time-varying imagery. Computer Vision, Graphics, and Image Processing, 21(2):205–221, February 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Paragios:2005:GAR

[PD05] Nikos Paragios and Rachid Deriche. Geodesic active regions and level set methods for motion estimation and tracking. Computer Vision and Image Understanding: CVIU, 97(3):259–282, March 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pehlivan:2011:NPB

[PD11] Selen Pehlivan and Pınar Duygulu. A new posebased representation for recognizing actions from multiple cameras. Computer Vision and Image Understanding: CVIU, 115(2):140– 151, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Picinbono:2003:NLA

[PDA03] Guillaume Picinbono, Hervé Delingette, and Nicholas Ayache. Non-linear anisotropic elasticity for real-time surgery simulation. Graphical Models, 65(5):305–321, September 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Palmer:1996:PMB

[PDK96] P. L. Palmer, H. Dabis, and J. Kittler. Performance measure for boundary detection algorithms. Computer Vision and Image Understanding: CVIU, 63(3):476–494, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0036/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0036/production/
pdf.

Pavlidis:2007:IHP

[PDS^{+}07] I. Pavlidis, J. Dowdall, N. Sun, C. Puri, J. Fei, and M. Garbey. Interacting with human physiology. Computer Vision and Image Understanding: CVIU, 108(1–2): 150–170, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Prasad:2006:EAP

[PDTE06] V. S. N. Prasad, Larry S. Davis, Son Dinh Tran, and Ahmed Elgammal. Edge affinity for pose-contour matching. Computer Vision and Image Understanding: CVIU, 104(1):36–47, October 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pahlavan:1992:HES

[PE92] Kourosh Pahlavan and Jan-Olof Eklundh. A head-eye system–analysis and design. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):41–56, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Paglieroni:2009:RAG

[PE09] David W. Paglieroni and Walter G. Eppler. Resolution analysis for Gradient Direction Matching of object model edges to overhead images. Computer Vision and Image Understanding: CVIU, 113 (2):235–248, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Peckinpaugh:1991:IMC

[Pec91] S. H. Peckinpaugh. An improved method for computing gray-level co-occurrence matrix based texture measures. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6):574– 580, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Pece:2007:CRG

[Pec07] Arthur E. C. Pece. On the computational rationale for generative models. Computer Vision and Image Understanding: CVIU, 106(1): 130–143, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Poggio:1992:LVM

[PEF92] Tomaso Poggio, Shimon Edelman, and Manfred Fahle. Learning of visual modules from examples: a framework for understanding adaptive visual performance. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):22–30, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Pizer:1998:ZIV

[PEFM98] Stephen M. Pizer, David Eberly, Daniel S. Fritsch, and Bryan S. Morse. Zoominvariant vision of figural shape: The mathematics of cores. Computer Vision and Image Understanding: CVIU, 69(1):055–071, January 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0563/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0563/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0563/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0564/production;
 http://www.idealibrary.
com/links/artid/cviu.1997.
0564/production/pdf;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0564/production/
ref.

Peleg:1979:ARH

[Pel79]  Shmuel Peleg.  Ambigu-

ity reduction in handwriting with ambiguous segmentation and uncertain interpretation. Computer Graphics and Image Processing, 10 (3):235–245, July 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Peleg:1984:CDO

[Pel84] Shmuel Peleg. Classification by discrete optimization. Computer Vision, Graphics, and Image Processing, 25 (1):122–130, January 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Penna:1989:LSL

[Pen89] Michael A. Penna. Local and semi-local shape from shading for a single perspective image of a smooth object. Computer Vision, Graphics, and Image Processing, 46(3):346–366, June 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Penna:1992:NRM

[Pen92] Michael A. Penna. Nonrigid motion analysis: Isometric motion. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (3):366–380, November 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Penna:1994:IAN

[Pen94] Michael A. Penna. The incremental approximation of nonrigid motion. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):141–156, September 1994. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1043/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1043/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1048/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1994.1048/production/
                                        pdf.

Penna:1999:MAN

[Pen99] Michael A. Penna. The motion analysis of nonrigid membranes. Computer Vision and Image Understanding: CVIU, 75(3):281–306, September 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0780/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0780/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0780/production/ref.

Peng:2003:MCR

[Pen03] Jing Peng. Multi-class relevance feedback content-based image retrieval. Computer Vision and Image Understanding: CVIU, 90(1):42–67, April 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Persoon:1976:NED

[Per76] Eric Persoon. New edge detection algorithm and its applications in picture processing. Computer Graphics and Image Processing, 5(4):425– 446, December 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Perkins:1981:UCS

[Per81] W. A. Perkins. Using circular symmetry and intensity profiles for computer vision inspection. Computer Graphics and Image Processing, 17(2): 161–172, October 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Peters:1985:ATP

[Pet85] Frans J. Peters. An algorithm for transformations of pictures represented by quadtrees. Computer Vision, Graphics, and Image Processing, 32 (3):397–403, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Peterfreund:1999:VSD

[Pet99]  Natan Peterfreund.  The

velocity snake: Deformable contour for tracking in spatio-velocity space. Computer Vision and Image Understanding: CVIU, 73 (3):346–356, March 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0732/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0732/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0732/production/ref.

Peternell:2000:GPB

[Pet00]  Martin Peternell.  Geo-

metric properties of bisector surfaces. Graphical Models, 62(3):202–236, May 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        1999.0521;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/gmod.1999.0521/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        1999.0521/ref.

Peuquet:1979:RMA

[Peu79] Donna J. Peuquet. A rastermode algorithm for interactive modification of line drawing data. Computer Graphics and Image Processing, 10 (2):142–158, June 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Peuquet:1983:HSS

[Peu83]  Donna J. Peuquet.  Hy-

brid structure for the storage and manipulation of very large spatial data sets. Computer Vision, Graphics, and Image Processing, 24(1):14–27, October 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Peyre:2009:MMS

[Pey09] Gabriel Peyré. Manifold models for signals and images.

Computer Vision and Image Understanding: CVIU, 113 (2):249–260, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ponce:1987:OCH

[PF87] Jean Ponce and Olivier Faugeras. An object centered hierarchical representation for 3D objects: the prism tree. Computer Vision, Graphics, and Image Processing, 38(1):1–28, April 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Plankers:2001:TMP

[PF01] Ralf Plänkers and Pascal Fua. Tracking and modeling people in video sequences. Computer Vision and Image Understanding: CVIU, 81(3):285–302, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0891;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0891/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0891/ref.

Philipp-Foliguet:2009:FIR

[PFGG09] Sylvie Philipp-Foliguet, Julien Gony, and Philippe-Henri Gosselin. FReBIR: an image retrieval system based on fuzzy region matching. Computer Vision and Image Understanding: CVIU, 113(6): 693–707, June 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pasko:2011:PFB

[PFV^{+}11] Alexander Pasko, Oleg Fryazinov, Turlif Vilbrandt, Pierre-Alain Fayolle, and Valery Adzhiev. Procedural functionbased modelling of volumetric microstructures. Graphical Models, 73(5):165–181, September 2011. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070311000087.

Papamarkos:1994:NAM

[PG94] N. Papamarkos and B. Gatos. A new approach for multilevel threshold selection. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(5):357–370, September 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1033/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1994.1033/production/
                                        pdf.

Picard:2013:EIS

[PG13] David Picard and Philippe-Henri Gosselin. Efficient image signatures and similarities using tensor products of local descriptors. Computer Vision and Image Understanding: CVIU, 117(6):680–687, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000337.

Pua:2004:RTR

[PGGM04] Kok Meng Pua, John M.

Gauch, Susan E. Gauch, and Jedrzej Z. Miadowicz. Real time repeated video sequence identification. Computer Vision and Image Understanding: CVIU, 93(3): 310–327, March 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Poppl:1982:BDS

[PH82] Siegfried J. Pöppl and Günter Herrmann. Boundary detection in scintigraphic images. Computer Graphics and Image Processing, 19(3): 281–290, July 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Pham:1986:DSS

[Pha86] Son Pham. Digital straight segments. Computer Vision, Graphics, and Image Processing, 36(1):10–30, October 1986. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pham:1989:CBS

[Pha89] Binh Pham. Conic B-splines for curve fitting: a unifying approach. Computer Vision, Graphics, and Image Processing, 45(1):117– 125, January 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pham:1991:EBS

[Pha91]  Binh Pham.  Expressive

brush strokes. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53 (1):1–6, January 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Pham:2001:SMF

[Pha01]  Dzung L. Pham.  Spa-

tial models for fuzzy clustering. Computer Vision and Image Understanding: CVIU, 84(2):285–297, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0951;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0951/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0951/ref.

Ponce:1992:UCM

[PHK92] J. Ponce, A. Hoogs, and D. J. Kriegman. On using CAD models to compute the pose of curved 3D objects. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (2):184–197, March 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Pang:2011:STD

[PHY^{+}11] Yanwei Pang, Qiang Hao, Yuan Yuan, Tanji Hu, Rui Cai, and Lei Zhang. Summarizing tourist destinations by mining user-generated travelogues and photos. Computer Vision and Image Understanding: CVIU, 115(3): 352–363, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pieroni:1979:MAD

[Pie79] Goffredo G. Pieroni. Method for analyzing dynamic processes represented by sequences of maps. Computer Graphics and Image Processing, 10(4):375– 387, August 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Piech:1988:CFT

[Pie88]  M. Ann Piech.  Com-

ments on fingerprints of two-dimensional edge models. Computer Vision, Graphics, and Image Processing, 42(3):381–386, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Princen:1990:HAL

[PIK90] John Princen, John Illingworth, and Josef Kittler. A hierarchical approach to line extraction based on the Hough transform. Computer Vision, Graphics, and Image Processing, 52(1):57– 77, October 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pizer:1981:IML

[Piz81] Stephen M. Pizer. Intensity mappings to linearize display devices. Computer Graphics and Image Processing, 17(3): 262–268, November 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Paglieroni:1988:CPT

[PJ88] David W. Paglieroni and Anil K. Jain. Control point transforms for shape representation and measurement. Computer Vision, Graphics, and Image Processing, 42(1):87–111, April 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pan:2011:MSP

[PJW11] Yongsheng Pan, Won-Ki Jeong, and Ross Whitaker. Markov surfaces: a probabilistic framework for userassisted three-dimensional image segmentation. Computer Vision and Image Understanding: CVIU, 115(10): 1375–1383, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001408.

Palagyi:1999:PST

[PK99] Kálmán Palágyi and Attila Kuba. A parallel 3D 12-subiteration thinning algorithm. Graphical Models and Image Processing: GMIP, 61(4):199–221, July 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0498/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0498/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0498/production/ref.

Pont:2005:RLG

[PK05] Sylvia C. Pont and Jan J.

Koenderink. Reflectance from locally glossy thoroughly pitted surfaces. Computer Vision and Image Understanding: CVIU, 98(2):211–222, May 2005. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Pitie:2007:ACG

[PKD07] François Pitié, Anil C. Kokaram, and Rozenn Dahyot. Automated colour grading using colour distribution transfer. Computer Vision and Image Understanding: CVIU, 107(1–2):123– 137, July/August 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Perse:2009:TBA

[PKK^{+}09] Matej Perše, Matej Kristan, Stanislav Kovačič, Goran Vučkovič, and Janez Perš. A trajectory-based analysis of coordinated team activity in a basketball game. Computer Vision and Image Understanding: CVIU, 113(5): 612–621, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Palmer:1997:OLF

[PKP97] P. L. Palmer, J. Kittler, and M. Petrou. An optimizing line finder using a Hough transform algorithm. Computer Vision and Image Understanding: CVIU, 67(1):1–23, July 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0491/production;   http:

//www.idealibrary.com/links/
artid/cviu.1996.0491/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0491/production/ref.

Pece:2007:GMB

[PL07] Arthur E. C. Pece and Rasmus Larsen. Generative model based vision. Computer Vision and Image Understanding: CVIU, 106(1):3–4, April 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Potetz:2008:EBP

[PL08] Brian Potetz and Tai Sing Lee. Efficient belief propagation for higher-order cliques using linear constraint nodes. Computer Vision and Image Understanding: CVIU, 112 (1):39–54, October 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Perko:2010:FVC

[PL10] Roland Perko and Aleš Leonardis. A framework for visual-context-aware object detection in still images. Computer Vision and Image Understanding: CVIU, 114(6): 700–711, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Platt:1992:GDC

[Pla92] John Platt. A generalization of dynamic constraints. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(6):516–525, November 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Pla:1996:RPC

[Pla96] Filiberto Pla. Recognition of partial circular shapes from segmented contours. Computer Vision and Image Understanding: CVIU, 63(2):334–343, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0023/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0023/production/
pdf.

Pottmann:2004:RI

[PLH04] Helmut Pottmann, Stefan Leopoldseder, and Michael Hofer. Registration without ICP. Computer Vision and Image Understanding: CVIU, 95(1):54–71, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Park:2000:LFM

[PLL00] Sang Ho Park, Kyoung Mu Lee, and Sang Uk Lee. A line feature matching technique based on an eigenvector approach. Computer Vision and Image Understanding:

CVIU, 77(3):263–283, March 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0808;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0808/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0808/ref.

Park:2012:MST

[PLL12] Min Ki Park, Seung Joo Lee, and Kwan H. Lee. Multiscale tensor voting for feature extraction from unstructured point clouds. Graphical Models, 74(4):197–208, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000288.

Park:2003:RPO

[PLLL03] Bo Gun Park, Kyoung Mu Lee, Sang Uk Lee, and Jin Hak Lee. Recognition of partially occluded objects using probabilistic ARG (attributed relational graph)-based matching. Computer Vision and Image Understanding: CVIU, 90(3): 217–241, June 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pottmann:1996:RRS

[PLR96] Helmut Pottmann, Wei Lü, and Bahram Ravani. Rational ruled surfaces and their offsets. Graphical Models and Image Processing: GMIP, 58 (6):544–552, November 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0045/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0045/production/
pdf.

Pei:1997:ECA

[PLS97] Soo-Chang Pei, Chin-Lun Lai, and Frank Y. Shih. An efficient class of alternating sequential filters in morphology. Graphical Models and Image Processing: GMIP, 59(2):109–116, March 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0416/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0416/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1996.
0416/production/ref.

Peli:1982:SED

[PM82] Tamar Peli and David Malah. A study of edge detection algorithms. Computer Graphics and Image Processing, 20(1): 1–21, September 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). PELI82.

Parvin:1989:AMF

[PM89] B. Parvin and G. Medioni.

Adaptive multiscale feature extraction from range data. Computer Vision, Graphics, and Image Processing, 45(3):346–356, March 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pla:1997:MFP

[PM97a] Filiberto Pla and John A.

Marchant. Matching feature points in image sequences through a regionbased method. Computer Vision and Image Understanding: CVIU, 66(3):271–285, June 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0512/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0512/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0512/production/ref.

Prescott:1997:LBC

[PM97b] B. Prescott and G. F.

McLean. Line-based correction of radial lens distortion. Graphical Models and Image Processing: GMIP, 59(1):39–47, January 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0407/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0407/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0407/production/ref.

Poiesi:2013:MTT

[PMC13] Fabio Poiesi, Riccardo Mazzon, and Andrea Cavallaro. Multi-target tracking on confidence maps: an application to people tracking. Computer Vision and Image Understanding: CVIU, 117(10): 1257–1272, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001634.

Pridmore:1990:EIP

[PMF90] Tony P. Pridmore, John E. W. Mayhew, and John P. Frisby. Exploiting imageplane data in the interpretation of edge-based binocular disparity. Computer Vision, Graphics, and Image Processing, 52(1):1–25, October 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Paranjape:1992:ANH

[PMR92] R. B. Paranjape, W. M. Morrow, and R. M. Rangayyan. Adaptive-neighborhood histogram equalization for image enhancement. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(3): 259–267, May 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Pluim:2000:IAM

[PMV00] Josien P. W. Pluim, J. B. Antoine Maintz, and Max A. Viergever. Interpolation artefacts in mutual informationbased image registration. Computer Vision and Image Understanding: CVIU, 77 (2):211–232, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0816;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0816/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0816/ref.

Planitz:2005:CFS

[PMW05] B. M. Planitz, A. J. Maeder, and J. A. Williams. The correspondence framework for 3D surface matching algorithms. Computer Vision and Image Understanding: CVIU, 97(3): 347–383, March 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Poggio:1985:EVC

[Pog85] Tomaso Poggio. Early vision: from computational structure to algorithms and parallel hardware. Computer Vision, Graphics, and Image Processing, 31(2):139– 155, August 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ponce:1990:CRF

[Pon90] J. Ponce. On characterizing ribbons and finding skewed symmetries. Computer Vision, Graphics, and Image Processing, 52(3):328– 340, December 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Poppe:2007:VBH

[Pop07] Ronald Poppe. Vision-based human motion analysis: an overview. Computer Vision and Image Understanding: CVIU, 108(1–2):4–18, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Portegys:2000:RHP

[Por00] Thomas E. Portegys. Recognizing hand-printed digits with a distance quasimetric. Computer Vision and Image Understanding: CVIU, 80(3):289–294, December 2000. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0876;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0876/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0876/ref.

Posdamer:1977:VDF

[Pos77] Jeffrey L. Posdamer. A vector development of the fundamentals of computational geometry. Computer Graphics and Image Processing, 6(4): 382–393, August 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Potter:1977:SSU

[Pot77] Jerry L. Potter. Scene segmentation using motion information. Computer Graphics and Image Processing, 6(6): 558–581, December 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Potmesil:1987:GOM

[Pot87] Michael Potmesil. Generating octree models of 3D objects from their silhouettes in a sequence of images. Computer Vision, Graphics, and Image Processing, 40(1):1–29, October 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Peng:1995:AME

[PP95] Anrong Peng and Wojciech Pieczynski. Adaptive mixture estimation and unsupervised local Bayesian image segmentation. Graphical Models and Image Processing: GMIP, 57(5):389–399, September 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1033/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1033/production/
pdf.

Palmer:1993:HTA

[PPK93] P. L. Palmer, M. Petrou, and J. Kittler. A Hough transform algorithm with a 2D hypothesis testing kernel. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(2):221–234, September 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1039/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1039/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1041/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1041/production/
pdf.

Petrou:2006:TRS

[PPT06] M. Petrou, R. Piroddi, and A. Talebpour. Texture recognition from sparsely and irregularly sampled data. Computer Vision and Image Understanding: CVIU, 102(1): 95–104, April 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Peng:2011:MFE

[PQML11] Bo Peng, Gang Qian, Yunqian Ma, and Baoxin Li. Multifactor feature extraction for human movement recognition. Computer Vision and Image Understanding: CVIU, 115 (3):375–389, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Proffitt:1979:MEC

[PR79] D. Proffitt and D. Rosen.

Metrication errors and coding efficiency of chain-encoding schemes for the representation of lines and edges. Computer Graphics and Image Processing, 10(4):318– 332, August 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Pizlo:1992:RPS

[PR92a] Zygmunt Pizlo and Azriel Rosenfeld. Recognition of planar shapes from perspective images using contour-based invariants. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (3):330–350, November 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Prokop:1992:SMB

[PR92b] Richard J. Prokop and Anthony P. Reeves. A survey of moment-based techniques for unoccluded object representation and recognition. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(5):438–460, September 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Pickering:2003:EKF

[PR03] Marcus J. Pickering and Stefan Rüger. Evaluation of key frame-based retrieval techniques for video. Computer Vision and Image Understanding: CVIU, 92(2–3):217– 235, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Prazdny:1981:DID

[Pra81] K. Prazdny. Determining the instantaneous direction of motion from optical flow generated by a curvilinearly moving observer. Computer Graphics and Image Processing, 17(3): 238–248, November 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Prazdny:1983:IOF

[Pra83a] K. Prazdny. On the information in optical flows. Computer Vision, Graphics, and Image Processing, 22(2):239–259, May 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Prazdny:1983:WSD

[Pra83b] K. Prazdny. Waveform segmentation and description using edge preserving smoothing. Computer Vision, Graphics, and Image Processing, 23 (3):327–333, September 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Price:1986:AYC

[Pri86a] Keith Price. Anything you can do, I can do better (no you can’t) …. Computer Vision, Graphics, and Image Processing, 36(2/3):387–391, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Price:1986:HMU

[Pri86b]  Keith E. Price.  Hierar-

chical matching using relaxation. Computer Vision, Graphics, and Image Processing, 34(1):66–75, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Paragios:2003:NRR

[PRR03] Nikos Paragios, Mikael Rousson, and Visvanathan Ramesh. Non-rigid registration using distance functions. Computer Vision and Image Understanding: CVIU, 89(2– 3):142–165, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pizlo:1997:GVS

[PRW97a] Zygmunt Pizlo, Azriel Rosenfeld, and Isaac Weiss. The geometry of visual space: About the incompatibility between science and mathematics. Computer Vision and Image Understanding: CVIU, 65(3):425–433, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0492/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0492/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0492/production/ref.

Pizlo:1997:VSM

[PRW97b] Zygmunt Pizlo, Azriel Rosenfeld, and Isaac Weiss. Visual space: Mathematics, engineering, and science. Computer Vision and Image Understanding: CVIU, 65(3):450–454, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0498/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0498/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0498/production/ref.

Park:1983:IRP

[PS83] Stephen K. Park and Robert A.

Schowengerdt. Image reconstruction by parametric cubic convolution. Computer Vision, Graphics, and Image Processing, 23(3):258– 272, September 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Palazzi:1994:CRR

[PS94] L. Palazzi and J. Snoeyink.

Counting and reporting red/ blue segment intersections. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(4):304–310, July 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1027/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1027/production/
pdf.

Pu:1995:TDG

[PS95] Christopher C. Pu and Frank Y. Shih. Threshold decomposition of gray-scale soft morphology into binary soft morphology. Graphical Models and Image Processing: GMIP, 57(6):522–526, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1042/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1042/production/
                                        pdf.

Philippou:1997:VFA

[PS97] Paul A. Philippou and Robin N. Strickland. Vector field analysis and synthesis using three-dimensional phase portraits. Graphical Models and Image Processing: GMIP, 59(6):446–462, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0445/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0445/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0445/production/ref.

Pulli:2000:SRD

[PS00] Kari Pulli and Linda G. Shapiro. Surface reconstruction and display from range and color data. Graphical Models, 62(3):165–201, May 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
1999.0519;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.1999.0519/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
1999.0519/ref.

Paris:2003:RAI

[PS03] Sylvain Paris and François Sillion. Robust acquisition of 3D informations from short image sequences. Graphical Models, 65(4):222–238, July 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Pasko:2005:SSI

[PS05a] Alexander Pasko and Michela Spagnuolo. SMI 2003 special issue. Graphical Models, 67(3):149, May 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Pham:2005:ORU

[PS05b] Thang V. Pham and Arnold W. M. Smeulders. Object recognition with uncertain geometry and uncertain part detection. Computer Vision and Image Understanding: CVIU, 99(2):241–258, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Peternell:2007:MSB

[PS07] Martin Peternell and Tibor Steiner. Minkowski sum boundary surfaces of 3Dobjects. Graphical Models, 69(3–4):180–190, May/July 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070307000021.

Proenca:2012:FCS

[PS12] Hugo Proença and Gil Santos. Fusing color and shape descriptors in the recognition of degraded iris images acquired at visible wavelengths. Computer Vision and Image Understanding: CVIU, 116 (2):167–178, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002359.

Papadopoulos:2011:CSO

[PSE^{+}11] G. Th. Papadopoulos, C. Saathoff, H. J. Escalante, V. Mezaris, I. Kompatsiaris, and M. G. Strintzis. A comparative study of object-level spatial context techniques for semantic image analysis. Computer Vision and Image Understanding: CVIU, 115(9): 1288–1307, September 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001275.

Patane:2007:FCG

[PSF07] G. Patanè, M. Spagnuolo, and B. Falcidieno. Families of cut-graphs for bordered meshes with arbitrary genus. Graphical Models, 69(2):119–138, March 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407030600066X.

Page:2002:NVV

[PSK^{+}02] D. L. Page, Y. Sun, A. F.

Koschan, J. Paik, and M. A. Abidi. Normal vector voting: Crease detection and curvature estimation on large, noisy meshes. Graphical Models, 64(3–4):199–229, May 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Potel:1980:GIT

[PSM80] Michael J. Potel, Richard E.

Sayre, and Steven A. Mackay. Graphics input tools for interactive motion analysis. Computer Graphics and Image Processing, 14(1):1– 23, September 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Pechuk:2008:LFB

[PSR08] Michael Pechuk, Octavian Soldea, and Ehud Rivlin. Learning function-based object classification from 3D imagery. Computer Vision and Image Understanding: CVIU, 110(2):173–191, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pong:1984:ESU

[PSWH84] Ting-Chuen Pong, Linda G. Shapiro, Layne T. Watson, and Robert M. Haralick. Experiments in segmentation using a facet model region grower. Computer Vision, Graphics, and Image Processing, 25(1):1–23, January 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Pei:2013:LPV

[PSYZ13] Mingtao Pei, Zhangzhang Si, Benjamin Z. Yao, and Song-Chun Zhu. Learning and parsing video events with goal and intent prediction. Computer Vision and Image Understanding: CVIU, 117(10):1369–1383, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421200197X.

Park:2008:UHI

[PT08] Sangho Park and Mohan M. Trivedi. Understanding human interactions with track and body synergies (TBS) captured from multiple views. Computer Vision and Image Understanding: CVIU, 111 (1):2–20, July 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pothos:2012:LSD

[PTE12] Vasileios K. Pothos, Christos Theoharatos, and George Economou. A local spectral distribution approach to face recognition. Computer Vision and Image Understanding: CVIU, 116(6):663–675, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000318.

Pudney:1998:DOH

[Pud98]  Chris Pudney.  Distance-

ordered homotopic thinning: a skeletonization algorithm for 3D digital images. Computer Vision and Image Understanding: CVIU, 72(3): 404–413, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0680/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0680/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0680/production/ref.

Pun:1981:ETN

[Pun81] T. Pun. Entropic thresholding, a new approach. Computer Graphics and Image Processing, 16(3):210– 239, July 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Pun:2003:RIT

[Pun03]  Chi-Man Pun.  Rotation-

invariant texture feature for image retrieval. Computer Vision and Image Understanding: CVIU, 89(1):24–43, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Poli:1997:SRD

[PV97] Riccardo Poli and Guido Valli. Shape from radiological density. Computer Vision and Image Understanding: CVIU, 65(3):361–381, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0489/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0489/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0489/production/ref.

Pauwels:2006:OIR

[PV06] Karl Pauwels and Marc M. Van Hulle. Optimal instantaneous rigid motion estimation insensitive to local minima. Computer Vision and Image Understanding: CVIU, 104(1):77–86, October 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pellegrini:2013:TMC

[PV13] Stefano Pellegrini and Luc Van Gool. Tracking with a mixed continuous-discrete Conditional Random Field. Computer Vision and Image Understanding: CVIU, 117(10):1215–1228, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001683.

Prenter:1986:TAD

[PW86] P. M. Prenter and E. R.

Westwater. Three adaptive discrete least squares cubic spline procedures for the compression of data. Computer Vision, Graphics, and Image Processing, 33(3):327– 345, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Prince:1991:CSR

[PW91] J. L. Prince and A. S. Willsky. Convex set reconstruction using prior shape information. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):413– 427, September 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Park:2006:SMC

[PW06] Eunkwang Park and Kwangyun Wohn. Stereo and motion correspondences using nonlinear optimization method. Computer Vision and Image Understanding: CVIU, 101(3): 194–203, March 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Palanivel:2008:MPA

[PY08a] S. Palanivel and B. Yegnanarayana. Multimodal person authentication using speech, face and visual speech. Computer Vision and Image Understanding: CVIU, 109(1): 44–55, January 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Park:2008:PAU

[PY08b] Youngsup Park and Kyunghyun Yoon. Painterly animation using motion maps. Graphical Models, 70(1–2):1–15, January/March 2008. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S152407030700015X.

Polat:2003:RTH

[PYS03] Ediz Polat, Mohammed Yeasin, and Rajeev Sharma. Robust tracking of human body parts for collaborative human computer interaction. Computer Vision and Image Understanding: CVIU, 89(1): 44–69, January 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Pavlidis:1992:PSC

[PZ92] Theo Pavlidis and Jiangying Zhou. Page segmentation and classification. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(6): 484–496, November 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Paragios:2008:DOC

[PZ08] Nikos Paragios and Ramin Zabih. Discrete optimization in computer vision. Computer Vision and Image Understanding: CVIU, 112(1): 1–2, October 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Paragios:2009:CDO

[PZ09] Nikos Paragios and Ramin Zabih. Corrigendum to “Discrete optimization in computer vision” [Comput. Vis. Image Understanding 112 (2008) 1–2]. Computer Vision and Image Understanding: CVIU, 113(4):588, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). See [PZ08].

Prankl:2013:IOM

[PZV13] Johann Prankl, Michael Zillich, and Markus Vincze. Interactive object modelling based on piecewise planar surface patches. Computer Vision and Image Understanding: CVIU, 117(6):718–731, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421300026X.

Pan:2013:EAF

[PZX13] Hong Pan, Yaping Zhu, and Liangzheng Xia. Efficient and accurate face detection using heterogeneous feature descriptors and feature selection. Computer Vision and Image Understanding: CVIU, 117(1):12–28, January 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001294.

Qazi:2011:PML

[QAB^{+}11] Imtnan-Ul-Haque Qazi, Olivier Alata, Jean-Christophe Burie, Mohamed Abadi, Ahmed Moussa, and Christine Fernandez-Maloigne. Parametric models of linear prediction error distribution for color texture and satellite image segmentation. Computer Vision and Image Understanding: CVIU, 115(8):1245–1262, August 2011. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000944.

Qian:2004:BSC

[QC04] Gang Qian and Rama Chellappa. Bayesian self-calibration of a moving camera. Computer Vision and Image Understanding: CVIU, 95(3): 287–316, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Quynh:2012:IAC

[QHXC12] Dao T. P. Quynh, Ying He, Shi-Qing Xin, and Zhonggui Chen. An intrinsic algorithm for computing geodesic distance fields on triangle meshes with holes. Graphical Models, 74(4):209–220, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031200029X.

Qureshi:2012:CAA

[QKH^{+}12] Rashid Jalal Qureshi, Laszlo Kovacs, Balazs Harangi, Brigitta Nagy, Tunde Peto, and Andras Hajdu. Combining algorithms for automatic detection of optic disc and macula in fundus images. Computer Vision and Image Understanding: CVIU, 116 (1):138–145, January 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001883.

Qu:1996:STA

[QL96] Xiaoqing Qu and Xiaobo Li. A 3D surface tracking algorithm. Computer Vision and Image Understanding: CVIU, 64(1):147–156, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0050/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0050/production/
                                        pdf.

Qian:2010:AIT

[QT10] Xiaoning Qian and Hemant D. Tagare. Adapting indexing trees to data distribution in feature spaces. Computer Vision and Image Understanding: CVIU, 114(1): 111–124, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Quan:1998:JIT

[QV98] Long Quan and Francoise Veillon. Joint invariants of a triplet of coplanar conics: Stability and discriminating power for object recognition. Computer Vision and Image Understanding: CVIU, 70(1):111–119, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0617/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0617/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0617/production/ref.

Qin:2002:EPP

[QY02] Xuejie Qin and Yee-Hong Yang. Estimating parameters for procedural texturing by genetic algorithms. Graphical Models, 64(1):19–39, January 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Riseman:1977:CTV

[RA77] Edward M. Riseman and Michael A. Arbib. Computational techniques in the visual segmentation of static scenes. Computer Graphics and Image Processing, 6(3): 221–276, June 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rabut:1992:EDB

[Rab92] Christophe Rabut. Even degree B-spline curves and surfaces. A note on the paper “B-Spline Curves and Surfaces Viewed as Digital Filters”, by A. Goshtasby, F. Cheng and B. Barsky. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4):

351–356, July 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). See [GCB90].

Rekik:2013:TGP

[RAC^{+}13] Islem Rekik, Stéphanie Allassonnière, Olivier Clatz, Ezequiel Geremia, Erin Stretton, Hervé Delingette, and Nicholas Ayache. Tumor growth parameters estimation and source localization from a unique time point: Application to low-grade gliomas. Computer Vision and Image Understanding: CVIU, 117(3):238–249, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001476.

Ragnemalm:1992:NDT

[Rag92] Ingemar Ragnemalm. Neighborhoods for distance transformations using ordered propagation. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(3):399–409, November 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Radeva:1997:DBS

[RAH97] Petia Radeva, Amir A. Amini, and Jiantao Huang. Deformable B-solids and implicit snakes for 3D localization and tracking of SPAMM MRI data. Computer Vision and Image Understanding: CVIU, 66(2):163–178, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0611/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0611/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0611/production/ref.

Raveaux:2011:LGP

[RAHT11] Romain Raveaux, Sébastien Adam, Pierre Héroux, and Éric Trupin. Learning graph prototypes for shape recognition. Computer Vision and Image Understanding: CVIU, 115(7):905–918, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421100083X.

Ramer:1972:IPP

[Ram72]  U. Ramer.  An iterative

procedure for the polygonal approximation of plane curves. Computer Graphics and Image Processing, 1(3): 244–256, November 1972. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ram:1976:AIS

[Ram76] Giora Ram. Analysis of images specified by graphlike descriptions. Computer Graphics and Image Processing, 5 (1):137–148, March 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ram:1984:ERI

[Ram84] Giora Ram. On the encoding and representing of images. Computer Vision, Graphics, and Image Processing, 26(2):224–232, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Radack:1982:JPM

[RB82] Gerald M. Radack and Norman I. Badler. Jigsaw puzzle matching using a boundary-centered polar encoding. Computer Graphics and Image Processing, 19 (1):1–17, May 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Radack:1989:LMS

[RB89] Gerald M. Radack and Norman I. Badler. Local matching of surfaces using a boundary-centered radial decomposition. Computer Vision, Graphics, and Image Processing, 45(3):380– 396, March 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1992:MLE

[RB92] Azriel Rosenfeld and Saibal Banerjee. Maximum-likelihood edge detection in digital signals. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (1):1–13, January 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Rao:1994:NIE

[RBA94] K. Raghunath Rao and Jezekiel Ben-Arie. Nonorthogonal image expansion related to optimal template matching in complex images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(2):149–160, March 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1014/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1014/production/
pdf.

Rajagopalan:1997:SVA

[RC97] A. N. Rajagopalan and S. Chaudhuri. Space-variant approaches to recovery of depth from defocused images. Computer Vision and Image Understanding: CVIU, 68(3):309–329, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0534/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0534/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0534/production/ref.

Richard:2003:NIR

[RC03] Frédéric J. P. Richard and Laurent D. Cohen. A new image registration technique with free boundary constraints: application to mammography. Computer Vision and Image Understanding: CVIU, 89(2–3):166–196, February/March 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rodgman:2006:RVG

[RC06] David Rodgman and Min Chen. Refraction in volume graphics. Graphical Models, 68(5–6):432–450, September/November 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070306000580.

Rouchdy:2013:GVA

[RC13] Youssef Rouchdy and Laurent D. Cohen. Geodesic voting for the automatic extraction of tree structures. Methods and applications. Computer Vision and Image Understanding: CVIU, 117(10):1453–1467, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001057.

Ripolles:2009:RCL

[RCG^{+}09] Oscar Ripolles, Miguel Chover, Jesus Gumbau, Francisco Ramos, and Anna Puig-Centelles. Rendering continuous level-of-detail meshes by Masking Strips. Graphical Models, 71(5):184–195, September 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407030900023X.

Ryoo:2013:PDD

[RCJ^{+}13] M. S. Ryoo, Sunglok Choi, Ji Hoon Joung, Jae-Yeong Lee, and Wonpil Yu. Personal driving diary: Automated recognition of driving events from first-person videos. Computer Vision and Image Understanding: CVIU, 117(10):1299–1312, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000118.

Rios-Cabrera:2012:EMC

[RCTV12] Reyes Rios-Cabrera, Tinne Tuytelaars, and Luc Van Gool. Efficient multi-camera vehicle detection, tracking, and identification in a tunnel surveillance application. Computer Vision and Image Understanding: CVIU, 116(6):742–753, June 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000380.

Rodriguez:2011:CCL

[RCVA11] Jorge Ernesto Rodríguez, Irving Cruz, Eduard Vergés, and Dolors Ayala. A connectedcomponent-labeling-based approach to virtual porosimetry. Graphical Models, 73 (5):296–310, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S152407031100018X.

Rajasekaran:1977:RPT

[RD77] S. N. S. Rajasekaran and B. L. Deekshatulu. Recognition of printed telugu characters. Computer Graphics and Image Processing, 6(4): 335–360, August 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Reed:1993:RRT

[RD93] Todd R. Reed and J. M. Hans Du Buf. A review of recent texture segmentation and feature extraction techniques. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):359–372, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1024/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1024/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1024/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1024/production/
pdf.

Radgui:2011:OFE

[RDM^{+}11] A. Radgui, C. Demonceaux, E. Mouaddib, M. Rziza, and D. Aboutajdine. Optical flow estimation from multichannel spherical image decomposition. Computer Vision and Image Understanding: CVIU, 115(9):1263–1272, September 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421100124X.

Rivlin:1995:RFP

[RDR95] Ehud Rivlin, Sven J. Dickinson, and Azriel Rosenfeld. Recognition by functional parts. Computer Vision and Image Understanding: CVIU, 62(2):164–176, September 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.

                                        com/links/artid/cviu.1995.
                                        1048/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1995.1048/production/
                                        pdf.

Rebordao:1989:LTL

[Reb89] J. M. Rebordao. Lookup table loadings for image processing with controlled knots. Computer Vision, Graphics, and Image Processing, 47 (2):189–202, August 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Reeves:1979:APS

[Ree79]  A. P. Reeves.  An ar-

ray processing system with a Fortran-based realization. Computer Graphics and Image Processing, 9(3):267– 281, March 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Reeves:1980:EGI

[Ree80] A. P. Reeves. On efficient global information extraction methods for parallel processors. Computer Graphics and Image Processing, 14(2):159– 169, October 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Reeves:1982:LMO

[Ree82] Anthony P. Reeves. Local median and other window operations on SIMD computers. Computer Graphics and Image Processing, 19(2): 165–178, June 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Reed:1984:CSC

[Ree84a] George M. Reed. On the characterization of simple closed surfaces in three-dimensional digital images. Computer Vision, Graphics, and Image Processing, 25(2):226– 235, February 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Reeves:1984:PCA

[Ree84b] Anthony P. Reeves. Parallel computer architectures for image processing. Computer Vision, Graphics, and Image Processing, 25(1):68– 88, January 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Reese:1992:IEI

[Ree92] Greg Reese. Image enhancement by intensity-dependent spread functions. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1): 45–55, January 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Reissell:1996:WMR

[Rei96] L.-M. Reissell. Wavelet multiresolution representation of curves and surfaces. Graphical Models and Image Processing: GMIP, 58(3):198–217, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0017/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0017/production/
                                        pdf.

Remondino:2004:DRS

[Rem04] Fabio Remondino. 3-D reconstruction of static human body shape from image sequence. Computer Vision and Image Understanding: CVIU, 93(1):65–85, January 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rewo:1984:EDC

[Rew84] Ludomir Rewo. Enhancement and detection of convex objects using regression models. Computer Vision, Graphics, and Image Processing, 25 (2):257–269, February 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Robertson:2002:PER

[RF02] Craig Robertson and Robert B. Fisher. Parallel evolutionary registration of range data.

Computer Vision and Image Understanding: CVIU, 87(1– 3):39–50, July 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rothwell:1997:CPR

[RFC97] Charlie Rothwell, Olivier Faugeras, and Gabriella Csurka. A comparison of projective reconstruction methods for pairs of views. Computer Vision and Image Understanding: CVIU, 68(1):37–58, October 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0525/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0525/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0525/production/ref.

Rodrigues:2002:SIR

[RFL02] Marcos Rodrigues, Robert Fisher, and Yonghuai Liu. Special issue on registration and fusion of range images. Computer Vision and Image Understanding: CVIU, 87 (1–3):1–7, July 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Richard:2011:DDR

[RFLSA11] A. Richard, L. Fuchs, G. Largeteau-Skapin, and E. Andres. Decomposition of nD-rotations: Classification, properties and algorithm. Graphical Models, 73(6):346–353, November 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070311000221.

Rohr:2003:SBE

[RFS03] K. Rohr, M. Fornefett, and H. S. Stiehl. Spline-based elastic image registration: integration of landmark errors and orientation attributes. Computer Vision and Image Understanding: CVIU, 90(2): 153–168, May 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rocha:2010:PRS

[RG10] Anderson Rocha and Siome Goldenstein. Progressive randomization: Seeing the unseen. Computer Vision and Image Understanding: CVIU, 114(3):349–362, March 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Raposo:2012:MAB

[RG12] Adriano N. Raposo and Abel J. P. Gomes. 3D molecular assembling of B-DNA sequences using nucleotides as building blocks. Graphical Models, 74(4):244–254, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000331.

Ryoo:2010:TDI

[RGA10] M. S. Ryoo, Kristen Grauman, and J. K. Aggarwal. A task-driven intelligent workspace system to provide guidance feedback. Computer Vision and Image Understanding: CVIU, 114(5): 520–534, May 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Reynolds:1987:DST

[RGC87] R. Anthony Reynolds, Dan Gordon, and Lih-Shyang S. Chen. A dynamic screen technique for shaded graphics display of slice-represented objects. Computer Vision, Graphics, and Image Processing, 38(3):275–298, June 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Richards:1985:CCC

[RH85] Whitman Richards and Donald D. Hoffman. Codon constraints on closed 2D shapes. Computer Vision, Graphics, and Image Processing, 31 (3):265–281, September 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ronse:1991:ABM

[RH91] C. Ronse and H. J. A. M. Heijmans. The algebraic basis of mathematical morphology. II Openings and closings. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(1):74–97, July 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Rigoutsos:1995:BAM

[RH95] Isidore Rigoutsos and Robert Hummel. A Bayesian approach to model matching with geometric hashing. Computer Vision and Image Understanding: CVIU, 62(1):11–26, July 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1038/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1038/production/
pdf.

Ragheb:2006:TNV

[RH06] Hossein Ragheb and Edwin R. Hancock. Testing new variants of the Beckmann– Kirchhoff model against radiance data. Computer Vision and Image Understanding: CVIU, 102(2):145–168, May 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rich:1984:ICC

[Ric84] Robert Rich. Image contrast, complexity, and stability. Computer Vision, Graphics, and Image Processing, 26(3):394–399, June 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Riesenfeld:1975:CA

[Rie75] R. Riesenfeld. On Chaikin’s algorithm. Computer Graphics and Image Processing, 4 (3):304–310, ???? 1975. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Risse:1989:HTL

[Ris89] Thomas Risse. Hough transform for line recognition: Complexity of evidence accumulation and cluster detection. Computer Vision, Graphics, and Image Processing, 46(3):327–345, June 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Raja:1994:OGP

[RJ94] Narayan Sriranga Raja and Anil K. Jain. Obtaining generic parts from range images using a multi-view representation. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):44–64, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1030/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1030/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1035/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1035/production/
pdf.

Raviv:2000:VLN

[RJ00] Daniel Raviv and Kunal Joarder. The visual looming navigation cue: a unified approach. Computer Vision and Image Understanding: CVIU, 79(3):331–363, September 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0862;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0862/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0862/ref.

Raftopoulos:2011:GLT

[RK11] Konstantinos A. Raftopoulos and Stefanos D. Kollias. The Global–Local transformation for noise resistant shape representation. Computer Vision and Image Understanding: CVIU, 115(8):1170–1186, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421100097X.

Rogelj:2003:PSM

[RKG03] Peter Rogelj, Stanislav Kovačič, and James C. Gee. Point similarity measures for non-rigid registration of multi-modal data. Computer Vision and Image Understanding: CVIU, 92(1):112–140, October 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Robles-Kelly:2005:ESR

[RKH05] Antonio Robles-Kelly and Edwin R. Hancock. Estimating the surface radiance function from single images. Graphical Models, 67 (6):518–548, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000135.

Rajagopalan:2000:LHF

[RKK^{+}00] A. N. Rajagopalan, K. Sunil Kumar, Jayashree Karlekar, R. Manivasakan, M. Milind Patil, U. B. Desai, P. G. Poonacha, and S. Chaudhuri. Locating human faces in a cluttered scene. Graphical Models, 62(5):323– 342, September 2000. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
1999.0511;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.1999.0511/

                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        1999.0511/ref.

Rosenfeld:1991:DS

[RKW91] A. Rosenfeld, T. Yung Kong, and A. Y. Wu. Digital surfaces. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53 (4):305–312, July 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Rao:1993:IHL

[RL93a] A. Ravishankar Rao and Gerald L. Lohse. Identifying high level features of texture perception. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(3):218–233, May 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1993.
                                        1016/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1993.1016/production/
                                        pdf.

Roth:1993:EGP

[RL93b] Gerhard Roth and Martin D. Levine. Extracting geometric primitives. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(1):1–22, July 1993. CODEN CIUNEJ.

ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1028/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1028/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1029/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1029/production/
pdf.

Roshtkhari:2013:LRT

[RL13a] Mehrsan Javan Roshtkhari and Martin D. Levine. An on-line, real-time learning method for detecting anomalies in videos using spatiotemporal compositions. Computer Vision and Image Understanding: CVIU, 117(10): 1436–1452, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001239.

Rosin:2013:AMR

[RL13b] Paul L. Rosin and Yu-Kun Lai. Artistic minimal rendering with lines and blocks. Graphical Models, 75(4):208–229, July 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070313000143.

Ryberg:2011:AEG

[RLC^{+}11] A. Ryberg, B. Lennartson, A.- K. Christiansson, M. Ericsson, and L. Asplund. Analysis and evaluation of a general camera model. Computer Vision and Image Understanding: CVIU, 115(11): 1503–1515, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001548.

Ramalingam:2006:GSM

[RLS06] Srikumar Ramalingam, Suresh K. Lodha, and Peter Sturm. A generic structure-from-motion framework. Computer Vision and Image Understanding: CVIU, 103(3):218– 228, September 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ranganathan:1991:VAD

[RM91] N. Ranganathan and R. Mehrotra. A VLSI architecture for dynamic scene analysis. Computer Vision, Graphics, and Image Processing. Image Understanding, 53(2): 189–197, March 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Robert:1998:FBI

[RM98] L. Robert and G. Malandain. Fast binary image processing using binary decision diagrams. Computer Vision and Image Understanding: CVIU, 72(1):1–9, October 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0655/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0655/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0655/production/ref.

Rivera:2002:ARC

[RM02] Mariano Rivera and Jose L.

Marroquin. Adaptive rest condition potentials: First and second order edgepreserving regularization. Computer Vision and Image Understanding: CVIU, 88(2):76– 93, November 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Raytchev:2003:URM

[RM03] Bisser Raytchev and Hiroshi Murase. Unsupervised recognition of multi-view face sequences based on pairwise clustering with attraction and repulsion. Computer Vision and Image Understanding: CVIU, 91(1–2):22–52, July/August 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rosin:2006:SCM

[RM06] Paul L. Rosin and Christine L. Mumford. A symmetric convexity measure. Computer Vision and Image Understanding: CVIU, 103(2): 101–111, August 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rahimi:2008:RDD

[RMD08] Ali Rahimi, Louis-Philippe Morency, and Trevor Darrell. Reducing drift in differential tracking. Computer Vision and Image Understanding: CVIU, 109(2): 97–111, February 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ruan:2002:FMS

[RMFB02] Su Ruan, Bruno Moretti, Jalal Fadili, and Daniel Bloyet. Fuzzy Markovian segmentation in application of magnetic resonance images. Computer Vision and Image Understanding: CVIU, 85(1): 54–69, January 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ramakrishna:1985:NIA

[RMR85] R. S. Ramakrishna, S. K. Mullick, and R. K. S. Rathore. A new iterative algorithm for image restoration. Computer Vision, Graphics, and Image Processing, 30(1):47– 55, April 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rao:1993:DSS

[RN93] Kashi Rao and Ram Nevatia. Describing and segmenting scenes from imperfect and incomplete data. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):1–23, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1001/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1001/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1001/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1001/production/
pdf.

Raudies:2012:REM

[RN12] Florian Raudies and Heiko Neumann. A review and evaluation of methods estimating ego-motion. Computer Vision and Image Understanding: CVIU, 116(5):606–633, May 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000021.

Rolland-Neviere:2013:RDB

[RNDA13] Xavier Rolland-Nevière, Gwenaël Doërr, and Pierre Alliez. Robust diameter-based thickness estimation of 3D objects. Graphical Models, 75 (6):279–296, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000192.

Robinson:1977:EDC

[Rob77] Guner S. Robinson. Edge detection by compass gradient masks. Computer Graphics and Image Processing, 6(5): 492–501, October 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Roberge:1985:DRA

[Rob85] James Robergé. A data reduction algorithm for planar curves. Computer Vision, Graphics, and Image Processing, 29(2):168– 195, February 1985. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Robert:1996:CCF

[Rob96a] Luc Robert. Camera calibration without feature extraction. Computer Vision and Image Understanding: CVIU, 63(2):314–325, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0021/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0021/production/
pdf.

Robinson:1996:LSC

[Rob96b] Julia Jean Robinson. Line symmetry of convex digital regions. Computer Vision and Image Understanding: CVIU, 64(2):263–285, September 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0058/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0058/production/
pdf.

Rosenfeld:1988:HTA

[ROH88] Azriel Rosenfeld, John Ornelas, Jr., and Yubin Hung. Hough transform algorithms for mesh-connected SIMD parallel processors. Computer Vision, Graphics, and Image Processing, 41(3):293– 305, March 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rohr:1994:TMB

[Roh94] K. Rohr. Towards modelbased recognition of human movements in image sequences. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):94–115, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1006/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1994.1006/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1006/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1006/production/
                                        pdf.

Ren:2009:TSB

[ROJX09] Jinchang Ren, James Orwell, Graeme A. Jones, and Ming Xu. Tracking the soccer ball using multiple fixed cameras. Computer Vision and Image Understanding: CVIU, 113 (5):633–642, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ronse:1986:SCP

[Ron86] Christian Ronse. A strong chord property for 4-connected convex digital sets. Computer Vision, Graphics, and Image Processing, 35(2):259– 269, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1976:SPP

[Ros76]  Azriel Rosenfeld.  Survey

picture processing: 1975. Computer Graphics and Image Processing, 5(2):215– 237, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rosenfeld:1977:PP

[Ros77] Azriel Rosenfeld. Picture processing: 1976. Computer Graphics and Image Processing, 6(2):157–183, April 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rosenfeld:1979:PP

[Ros79] Azriel Rosenfeld. Picture processing: 1978. Computer Graphics and Image Processing, 9(4):354–386, April 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rosen:1980:ABQ

[Ros80a] D. Rosen. On the areas and boundaries of quantized objects. Computer Graphics and Image Processing, 13 (1):94–98, May 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rosenfeld:1980:PP

[Ros80b] Azriel Rosenfeld. Picture processing: 1979. Computer Graphics and Image Processing, 13(1):46–79, May 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rosenfeld:1981:PP

[Ros81] Azriel Rosenfeld. Picture processing: 1980. Computer Graphics and Image Processing, 16(1):52–82, May 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rosenfeld:1982:PP

[Ros82] Azriel Rosenfeld. Picture processing: 1981. Computer Graphics and Image Processing, 19(1):35–75, May 1982. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rosenfeld:1983:PP

[Ros83] Azriel Rosenfeld. Picture processing: 1982. Computer Vision, Graphics, and Image Processing, 22(3):339– 387, June 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1984:PP

[Ros84] Azriel Rosenfeld. Picture processing: 1983. Computer Vision, Graphics, and Image Processing, 26(3):347– 393, June 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1985:PP

[Ros85] Azriel Rosenfeld. Picture processing: 1984. Computer Vision, Graphics, and Image Processing, 30(2):189– 242, May 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1986:ARS

[Ros86a] Azriel Rosenfeld. Axial representations of shape. Computer Vision, Graphics, and Image Processing, 33(2):156– 173, February 1986. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1986:DEV

[Ros86b] Azriel Rosenfeld. Dialog: “expert” vision systems: Some issues. Computer Vision, Graphics, and Image Processing, 34(1):99–102, 117, April 1986. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic). See also [Kov86, Nag86, Tho86, Uhr86].

Rosenfeld:1986:SPP

[Ros86c] Azriel Rosenfeld. Survey:

Picture processing — 1985. Computer Vision, Graphics, and Image Processing, 34(2):204–251, May 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1987:PP

[Ros87a] Azriel Rosenfeld. Picture processing: 1986. Computer Vision, Graphics, and Image Processing, 38(2):147– 225, May 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1987:TWH

[Ros87b] Azriel Rosenfeld, editor. Third Workshop on Human and Machine Vision, volume 37(1) of Computer Vision, Graphics, and Image Processing. Academic Press, New York, NY, USA, January 1987. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1988:IAC

[Ros88]  Azriel Rosenfeld.  Image

analysis and computer vision: 1987. Computer Vision, Graphics, and Image Processing, 42(2):234–293, May 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). ROSENFELD88.

Rosenfeld:1989:SIA

[Ros89] Azriel Rosenfeld. Survey: Image analysis and computer vision: 1988. Computer Vision, Graphics, and Image Processing, 46(2):196–264, May 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1990:IAC

[Ros90]  Azriel Rosenfeld.  Image

analysis and computer vision. 1989. Computer Vision, Graphics, and Image Processing, 50(2):188–240, May 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosenfeld:1991:IAC

[Ros91] A. Rosenfeld. Image analysis and computer vision: 1990. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (3):322–365, May 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Rosenfeld:1992:IAC

[Ros92] Azriel Rosenfeld. Image analysis and computer vision: 1991 (survey). Computer Vision, Graphics, and Image Processing. Image Understanding, 55(3):349–380, May 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Rosenfeld:1993:IAC

[Ros93a]  Azriel Rosenfeld.  Image

analysis and computer vision: 1992. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(1):85–135, July 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1033/production;   http:
//www.idealibrary.com/links/

artid/ciun.1993.1033/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1034/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1034/production/
pdf.

Rosin:1993:MRM

[Ros93b] Paul L. Rosin. Multiscale representation and matching of curves using codons. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(4):286–310, July 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1020/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1020/production/
pdf.

Rosenfeld:1994:IAC

[Ros94]  Azriel Rosenfeld.  Image

analysis and computer vision: 1993. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):367–404, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1026/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1026/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.

1030/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1030/production/
pdf.

Rosenfeld:1995:IAC

[Ros95]  Azriel Rosenfeld.  Image

analysis and computer vision: 1994. Computer Vision and Image Understanding: CVIU, 62(1):90–143, July 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1044/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1044/production/
pdf.

Rosenfeld:1996:IAC

[Ros96a]  Azriel Rosenfeld.  Image

analysis and computer vision: 1995. Computer Vision and Image Understanding: CVIU, 63(3):568–612, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0041/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0041/production/
pdf.

Rosin:1996:AEF

[Ros96b] Paul L. Rosin. Assessing error of fit functions for ellipses. Graphical Models and Image Processing: GMIP, 58(5):494–502, September 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0041/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0041/production/
pdf.

Rosin:1996:ACD

[Ros96c] Paul L. Rosin. Augmenting corner descriptors. Graphical Models and Image Processing: GMIP, 58(3):286–294, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0023/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0023/production/
pdf.

Rosenfeld:1997:IAC

[Ros97]  Azriel Rosenfeld.  Image

analysis and computer vision: 1996. Computer Vision and Image Understanding: CVIU, 66(1):33–93, April 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0602/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0602/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0602/production/ref.

Rosenfeld:1998:IAC

[Ros98a]  Azriel Rosenfeld.  Image

analysis and computer vision: 1997. Computer Vision and Image Understanding: CVIU, 70(2):239–284, May 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0697/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0697/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0697/production/ref.

Rosin:1998:EFU

[Ros98b] Paul L. Rosin. Ellipse fitting using orthogonal hyperbolae and Stirling’s oval. Graphical Models and Image Processing: GMIP, 60(3):209–213, May 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0471/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0471/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0471/production/ref.

Rosenfeld:1999:IAC

[Ros99a]  Azriel Rosenfeld.  Image

analysis and computer vision: 1998. Computer Vision and Image Understanding: CVIU, 74(1):36–95, April 1999. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0746/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0746/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0746/production/ref.

Rosin:1999:FFP

[Ros99b]  Paul L. Rosin.  Further

five-point fit ellipse fitting. Graphical Models and Image Processing: GMIP, 61 (5):245–259, September 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0500/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0500/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0500/production/ref.

Rosin:1999:MCP

[Ros99c]  Paul L. Rosin.  Measur-

ing corner properties. Computer Vision and Image Understanding: CVIU, 73 (2):291–307, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0719/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0719/production/

pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0719/production/ref.

Rosenfeld:2000:CLR

[Ros00a] Azriel Rosenfeld. Classifying the literature related to computer vision and image analysis. Computer Vision and Image Understanding: CVIU, 79(2):308–323, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0851;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0851/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0851/ref.

Rosenfeld:2000:IAC

[Ros00b]  Azriel Rosenfeld.  Image

Analysis and Computer Vision: 1999. Computer Vision and Image Understanding: CVIU, 78(2):222–302, May 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0835;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0835/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0835/ref.

Rosenfeld:2001:IAC

[Ros01] Azriel Rosenfeld. From image analysis to computer vision:

An annotated bibliography, 1955–1979. Computer Vision and Image Understanding: CVIU, 84(2):298–324, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0953;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0953/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0953/ref.

Rosin:2002:TCD

[Ros02] Paul L. Rosin. Thresholding for change detection. Computer Vision and Image Understanding: CVIU, 86(2): 79–95, May 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rosin:2008:TCR

[Ros08]  Paul L. Rosin.  A two-

component rectilinearity measure. Computer Vision and Image Understanding: CVIU, 109(2):176–185, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rosin:2010:IPU

[Ros10a] Paul L. Rosin. Image processing using 3-state cellular automata. Computer Vision and Image Understanding: CVIU, 114(7):790–802, July 2010. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Rossignac:2010:GCP

[Ros10b] Jarek Rossignac. GMOD:

Creation, processing, animation, visualization, and dissemination of graphical models. Graphical Models, 72 (1):iii–v, January 2010. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070310000032.

Roth:1982:RCM

[Rot82] Scott D. Roth. Ray casting for modeling solids. Computer Graphics and Image Processing, 18(2):109– 144, February 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). Also in SIGGRAPH ’87, ’88, ’89 Introduction to Ray Tracing course notes. The other classic ray tracing paper.

Reddy:1988:RMT

[RP88] B. R. K. Reddy and A. L. Pai.

Reed–Muller transform image coding. Computer Vision, Graphics, and Image Processing, 42(1):48–61, April 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rahbar:2008:ILC

[RP08] Kambiz Rahbar and Hamid Reza Pourreza. Inside looking out camera pose estimation for virtual studio. Graphical Models, 70(4):57–75, July 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000027.

Rashwan:2012:IRV

[RPG12] Hatem A. Rashwan, Domenec Puig, and Miguel Angel Garcia. Improving the robustness of variational optical flow through tensor voting. Computer Vision and Image Understanding: CVIU, 116 (9):953–966, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000756.

Rubner:2001:EED

[RPTB01] Yossi Rubner, Jan Puzicha, Carlo Tomasi, and Joachim M. Buhmann. Empirical evaluation of dissimilarity measures for color and texture. Computer Vision and Image Understanding: CVIU, 84(1):25–43, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0934;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0934/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/cviu.
2001.0934/ref.

Rivlin:1995:NF

[RR95] Ehud Rivlin and Azriel Rosenfeld. Navigational functionalities. Computer Vision and Image Understanding: CVIU, 62(2): 232–244, September 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1052/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1052/production/
pdf.

Robertson:2006:GMH

[RR06] Neil Robertson and Ian Reid.

A general method for human activity recognition in video. Computer Vision and Image Understanding: CVIU, 104(2–3):232–248, November/December 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Raskin:2011:DRU

[RRR11] Leonid Raskin, Michael Rudzsky, and Ehud Rivlin. Dimensionality reduction using a Gaussian Process Annealed Particle Filter for tracking and classification of articulated body motions. Computer Vision and Image Understanding: CVIU, 115(4):503–519, April 2011. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Rajala:1983:AOD

[RRS83] Sarah A. Rajala, Alfy N. Riddle, and Wesley E. Snyder. Application of the onedimensional Fourier transform for tracking moving objects in noisy environments. Computer Vision, Graphics, and Image Processing, 21 (2):280–293, February 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ranganathan:1988:VAC

[RS88] N. Ranganathan and Mubarak Shah. A VLSI architecture for computing scale space. Computer Vision, Graphics, and Image Processing, 43 (2):178–204, August 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rangarajan:1991:EMC

[RS91a] Krishnan Rangarajan and Mubarak Shah. Establishing motion correspondence. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(1):56–73, July 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Rao:1991:COT

[RS91b] A. Ravishankar Rao and B. G. Schunck. Computing oriented texture fields. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2): 157–185, March 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Rosales:2003:FHG

[RS03] Rómer Rosales and Stan Sclaroff. A framework for heading-guided recognition of human activity. Computer Vision and Image Understanding: CVIU, 91(3):335– 367, September 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Raman:1993:HSE

[RSB93] S. V. Raman, S. Sarkar, and K. L. Boyer. Hypothesizing structures in edge-focused cerebral magnetic resonance images using graph-theoretic cycle enumeration. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):81–98, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1005/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1005/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1005/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1993.1005/production/
                                        pdf.

Rueda:2004:RCP

[RSFdM04] A. J. Rueda, R. J. Segura, F. R. Feito, and J. Ruiz de Miras. Rasterizing complex polygons without tessellations. Graphical Models, 66 (3):127–132, May 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Ramalingam:2010:GSC

[RSL10] Srikumar Ramalingam, Peter Sturm, and Suresh K. Lodha. Generic self-calibration of central cameras. Computer Vision and Image Understanding: CVIU, 114(2):210– 219, February 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ribnick:2012:RAP

[RSPD12] Evan Ribnick, Ravishankar Sivalingam, Nikolaos Papanikolopoulos, and Kostas Daniilidis. Reconstructing and analyzing periodic human motion from stationary monocular views. Computer Vision and Image Understanding: CVIU, 116(7):815–826, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000513.

Rivlin:2007:ESV

[RSS07] Guy Froimovich Ehud Rivlin, Ilan Shimshoni, and Octavian Soldea. Efficient search and verification for function based classification from real range images. Computer Vision and Image Understanding: CVIU, 105(3):200–217, March 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Rangarajan:1989:OCD

[RSv89] Krishnan Rangarajan, Mubarak Shah, and David van Brackle. Optimal corner detector. Computer Vision, Graphics, and Image Processing, 48 (2):230–245, November 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rubin:1980:NSR

[Rub80] Steven M. Rubin. Natural scene recognition using locus search. Computer Graphics and Image Processing, 13(4): 298–333, August 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rubin:1982:NRD

[Rub82] Steven M. Rubin. Note: The representation and display of scenes with a wide range of detail. Computer Graphics and Image Processing, 19(3): 291–298, July 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rutkowski:1979:SC

[Rut79] Wallace S. Rutkowski. Shape completion. Computer Graphics and Image Processing, 9(1):89–101, January 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rutkowski:1981:SSU

[Rut81] W. S. Rutkowski. Shape segmentation using arc/chord properties. Computer Graphics and Image Processing, 17 (2):114–129, October 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rutkowski:1982:ROS

[Rut82] Wallace S. Rutkowski. Recognition of occluded shapes using relaxation. Computer Graphics and Image Processing, 19(2):111–128, June 1982. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rothstein:1976:PSS

[RW76] Jerome Rothstein and Carl Weiman. Parallel and sequential specification of a context sensitive language for straight lines on grids. Computer Graphics and Image Processing, 5(1):106– 124, March 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Rothstein:1979:PSS

[RW79] Jerry Rothstein and Carl F. R. Weiman. Parallel and sequential specification of a context sensitive language for straight lines on grids. Computer Graphics and Image Processing, 5(??): 106–124, ???? 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Raafat:1988:TID

[RW88] Hazem M. Raafat and Andrew K. C. Wong. A texture information-directed region growing algorithm for image segmentation and region classification. Computer Vision, Graphics, and Image Processing, 43(1):1–21, July 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Rosin:1995:SDT

[RW95] Paul L. Rosin and Geoff A. W.

West. Salience distance transforms. Graphical Models and Image Processing: GMIP, 57 (6):483–521, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1041/production;   http:
//www.idealibrary.com/links/

artid/gmip.1995.1041/production/
pdf.

Rivlin:1997:DIO

[RW97] Ehud Rivlin and Isaac Weiss. Deformation invariants in object recognition. Computer Vision and Image Understanding: CVIU, 65 (1):95–108, January 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0478/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0478/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0478/production/ref.

Ritter:1990:IAO

[RWD90] G. X. Ritter, J. N. Wilson, and J. L. Davidson. Image algebra. an overview. Computer Vision, Graphics, and Image Processing, 49(3):297–331, March 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Robey:1995:IUP

[RWV95] M. Robey, G. West, and S. Venkatesh. An investigation into the use of physical modeling for the prediction of various feature types visible from different viewpoints. Computer Vision and Image Understanding: CVIU, 61(3):417–429, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1031/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1031/production/
pdf.

Reinhardt:2000:CBS

[RWWH00] Joseph M. Reinhardt, Andrien J. Wang, Thomas P. Weldon, and William E. Higgins. Cue-based segmentation of 4D cardiac image sequences. Computer Vision and Image Understanding: CVIU, 77(2):251–262, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0818;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0818/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0818/ref.

Rodriguez:1995:HRH

[RY95] Jeffrey J. Rodriguez and Christopher C. Yang. Highresolution histogram modification of color images. Graphical Models and Image Processing: GMIP, 57 (5):432–440, September 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.

                                        1037/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1037/production/
                                        pdf.

Reisfeld:1998:PFI

[RY98] Daniel Reisfeld and Yehezkel Yeshurun. Preprocessing of face images: Detection of features and pose normalization. Computer Vision and Image Understanding: CVIU, 71(3):413–430, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0640/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0640/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0640/production/ref.

Rosenfeld:1998:TPD

[RYN98] Azriel Rosenfeld, T. Yung Kong, and Akira Nakamura. Topology-preserving deformations of two-valued digital pictures. Graphical Models and Image Processing: GMIP, 60(1):024–034, January 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0459/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0459/production/
                                        pdf;   http://www.idealibrary.

com/links/artid/gmip.1997.
0459/production/ref.

Rosin:2005:MR

[RŽ05] Paul L. Rosin and Joviša Žunić. Measuring rectilinearity. Computer Vision and Image Understanding: CVIU, 99(2):175–188, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schachter:1979:RPG

[SA79] Bruce Schachter and Narendra Ahuja. Random pattern generation processes. Computer Graphics and Image Processing, 10(2):95– 114, June 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shlien:1981:FFA

[SA81] Seymour Shlien and Paul Allard. A fir filtering approach for the generation of smooth curves on a graphics terminal. Computer Graphics and Image Processing, 17(3):269– 280, November 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Suzuki:1985:TSA

[SA85] Satoshi Suzuki and Keiichi Abe. Topological structural analysis of digitized binary images by border following. Computer Vision, Graphics, and Image Processing, 30(1):32–46, April 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Srivastava:1990:OGO

[SA90] Sanjay K. Srivastava and Narendra Ahuja. Octree generation from object silhouettes in perspective views. Computer Vision, Graphics, and Image Processing, 49(1):68–84, January 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sabata:1991:EMP

[SA91] Bikash Sabata and J. K. Aggarwal. Estimation of motion from a pair of range images: a review. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (3):309–324, November 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Singh:1992:IFC

[SA92] Ajit Singh and Peter Allen. Image-flow computation: An estimation-theoretic framework and a unified perspective. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (2):152–177, September 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Sull:1995:IMS

[SA95] Sanghoon Sull and Narendra Ahuja. Integrated matching and segmentation of multiple features in two views. Computer Vision and Image Understanding: CVIU, 62(3):279–297, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1055/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1055/production/
pdf.

Sabata:1996:SCM

[SA96] Bikash Sabata and J. K. Aggarwal. Surface correspondence and motion computation from a pair of range images. Computer Vision and Image Understanding: CVIU, 63(2):232–250, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0017/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0017/production/
pdf.

Stamos:2002:GTR

[SA02] Ioannis Stamos and Peter K.

Allen. Geometry and texture recovery of scenes of large scale. Computer Vision and Image Understanding: CVIU, 88(2):94–118, November 2002.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sato:2004:TSV

[SA04] Koichi Sato and J. K. Aggarwal. Temporal spatio-velocity transform and its application to tracking and interaction. Computer Vision and Image Understanding: CVIU, 96(2): 100–128, November 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sabata:1993:SRI

[SAA93] Bikash Sabata, Farshid Arman, and J. K. Aggarwal. Segmentation of 3D range images using pyramidal data structures. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):373–387, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1993.
                                        1025/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1993.1025/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1993.
                                        1025/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1993.1025/production/
                                        pdf.

Sabah:1976:NCO

[Sab76] G. Sabah. Notion of command operator in pattern recognition. Computer Graphics and Image Processing, 5(1): 125–136, March 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Skaff:2009:SBA

[SAC09] Sandra Skaff, Tal Arbel, and James J. Clark. A sequential Bayesian approach to color constancy using non-uniform filters. Computer Vision and Image Understanding: CVIU, 113(9):993–1004, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sederberg:1984:IRP

[SAG84] T. W. Sederberg, D. C. Anderson, and R. N. Goldman. Implicit representation of parametric curves and surfaces. Computer Vision, Graphics, and Image Processing, 28(1):72–84, October 1984. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sederberg:1985:III

[SAG85] Thomas W. Sederberg, David C.

Anderson, and Ronald N. Goldman. Implicitization, inversion, and intersection of planar rational cubic curves. Computer Vision, Graphics, and Image Processing, 31(1): 89–102, July 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Saha:2005:TSL

[Sah05] Punam Kumar Saha. Tensor scale: a local morphometric parameter with applications to computer vision and image processing. Computer Vision and Image Understanding: CVIU, 99 (3):384–413, September 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Salzman:1990:MGM

[Sal90] David B. Salzman. A method of general moments for orienting 2D projections of unknown 3D objects. Computer Vision, Graphics, and Image Processing, 50(2):129– 156, May 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Samet:1980:RRQ

[Sam80] H. Samet. Region representation: quadtrees from binary arrays. Computer Graphics and Image Processing, 13 (1):88–93, May 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Samet:1982:NFT

[Sam82a] Hanan Samet. Neighbor finding techniques for images represented by quadtrees.

Computer Graphics and Image Processing, 18(1):37– 57, January 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). SAMET82.

Sampson:1982:FCS

[Sam82b] Paul D. Sampson. Fitting conic sections to “very scattered” data: an iterative refinement of the Bookstein algorithm. Computer Graphics and Image Processing, 18(1): 97–108, January 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Samet:1984:ACQ

[Sam84] Hanan Samet. Algorithms for the conversion of quadtrees to rasters. Computer Vision, Graphics, and Image Processing, 26(1):1–16, April 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Samet:1985:RQQ

[Sam85] Hanan Samet. Reconstruction of quadtrees from quadtree medial axis transforms. Computer Vision, Graphics, and Image Processing, 29(3):311– 328, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Samet:1989:NFI

[Sam89] Hanan Samet. Neighbor finding in images represented by octrees. Computer Vision, Graphics, and Image Processing, 46(3):367–386, June 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sankar:1977:VCS

[San77]  P. V. Sankar.  A ver-

tex coding scheme for interpreting ambiguous trihedral solids. Computer Graphics and Image Processing, 6(1): 61–89, February 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sankar:1978:GIQ

[San78] P. V. Sankar. Grid intersect quantization schemes for solid object digitization. Computer Graphics and Image Processing, 8(1):25–42, August 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

SannitidiBaja:1990:CPM

[San90] Gabriella Sanniti di Baja. O(N) computation of projections and moments from the labeled skeleton. Computer Vision, Graphics, and Image Processing, 49(3):369– 378, March 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sapiro:1997:CS

[Sap97]  Guillermo Sapiro.  Color

snakes. Computer Vision and Image Understanding: CVIU, 68(2):247–253, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0562/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0562/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0562/production/ref.

Sarraga:1983:AMI

[Sar83] Ramon F. Sarraga. Algebraic methods for intersections of quadric surfaces in GM-SOLID. Computer Vision, Graphics, and Image Processing, 22(2):222–238, May 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sanroma:2012:NGM

[SAS12] Gerard Sanromà, René Alquézar, and Francesc Serratosa. A new graph matching method for point-set correspondence using the EM algorithm and softassign. Computer Vision and Image Understanding: CVIU, 116(2): 292–304, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002360.

Sauer:1991:ELB

[Sau91] K. Sauer. Enhancement of low bit-rate coded images using edge detection and estimation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1):52– 62, January 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Saund:1993:ISC

[Sau93]  Eric Saund.  Identify-

ing salient circular arcs on curves. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3):327–337, November 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1045/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1045/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1047/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1047/production/
pdf.

Saund:1999:POO

[Sau99] Eric Saund. Perceptual organization of occluding contours of opaque surfaces. Computer Vision and Image Understanding: CVIU, 76(1):70–82, October 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0789/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0789/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0789/production/ref.

Savoy:1987:CAI

[Sav87]  Robert L. Savoy.  Con-

tingent aftereffects and isoluminance: Psychophysical evidence for separation of color, orientation, and motion. Computer Vision, Graphics, and Image Processing, 37(1):3–19, January 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sloan:1979:CMT

[SB79] Kenneth R. Sloan, Jr. and Christopher M. Brown. Color map techniques. Computer Graphics and Image Processing, 10(4):297–317, August 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shani:1984:SEG

[SB84] Uri Shani and Dana H. Ballard. Splines as embeddings for generalized cylinders. Computer Vision, Graphics, and Image Processing, 27 (2):129–156, August 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Savoji:1985:DMB

[SB85] M. H. Savoji and R. E. Burge. On different methods based on the Karhunen-Loeve expansion and used in image analysis. Computer Vision, Graphics, and Image Processing, 29(2):259– 269, February 1985. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Stevens:1987:DSS

[SB87] Kent A. Stevens and Allen Brookes. Detecting structure by symbolic constructions on tokens. Computer Vision, Graphics, and Image Processing, 37(2):238– 260, February 1987. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sotak:1989:LGK

[SB89] G. E. Sotak, Jr. and K. L. Boyer. The Laplacian-of-Gaussian kernel: a formal analysis and design procedure for fast, accurate convolution and full-frame output. Computer Vision, Graphics, and Image Processing, 48 (2):147–189, November 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Stewman:1990:DCP

[SB90] John H. Stewman and Kevin W. Bowyer. Direct construction of the perspective projection aspect graph of convex polyhedra. Computer Vision, Graphics, and Image Processing, 51(1):20–37, July 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sarkar:1991:OII

[SB91] S. Sarkar and K. L. Boyer.

Optimal infinite impulse response zero crossing based edge detectors. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(2):224–243, September 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Sato:1993:FRS

[SB93] Hiroaki Sato and Thomas O.

Binford. Finding and recovering SHGC objects in an edge image. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):346–358, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1023/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1023/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1023/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1993.1023/production/
                                        pdf.

Stark:1994:FBG

[SB94] Louise Stark and Kevin Bowyer. Function-based generic recognition for multiple object categories. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):1–21, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1001/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1994.1001/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1001/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1994.1001/production/
                                        pdf.

Sarkar:1995:UPI

[SB95] Sudeep Sarkar and Kim L. Boyer. Using perceptual inference networks to manage vision processes. Computer Vision and Image Understanding: CVIU, 62(1):27–46, July 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1995.
                                        1039/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1995.1039/production/
                                        pdf.

Sahabi:1996:AED

[SB96a] Hossein Sahabi and Anup Basu. Analysis of error in depth perception with vergence and spatially varying sensing. Computer Vision and Image Understanding: CVIU, 63(3):447–461, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0034/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0034/production/
pdf.

Shaked:1996:CA

[SB96b] Doron Shaked and Alfred M.

Bruckstein. The curve axis. Computer Vision and Image Understanding: CVIU, 63(2):367–379, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0026/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0026/production/
pdf.

Sarkar:1998:QMC

[SB98a] Sudeep Sarkar and Kim L.

Boyer. Quantitative measures of change based on feature organization: Eigenvalues and eigenvectors. Computer Vision and Image Understanding: CVIU, 71(1):110–136, July 1998. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0637/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0637/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0637/production/ref.

Sengupta:1998:MPU

[SB98b] Kuntal Sengupta and Kim L. Boyer. Modelbase partitioning using property matrix spectra. Computer Vision and Image Understanding: CVIU, 70(2):177–196, May 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0631/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0631/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0631/production/ref.

Shaked:1998:PMA

[SB98c] Doron Shaked and Alfred M. Bruckstein. Pruning medial axes. Computer Vision and Image Understanding: CVIU, 69(2):156–169, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0598/production;   http:
//www.idealibrary.com/links/

artid/cviu.1997.0598/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0598/production/ref.

Stark:19xx:IFB

[SBxx] L. Stark and K. Bowyer. Indexing function-based categories for generic object recognition. Computer Vision, Graphics, and Image Processing. Image Understanding, ?? (??):??, ???? 19xx. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Stevens:2000:LSI

[SB00] Mark R. Stevens and J. Ross Beveridge. Localized scene interpretation from 3D models, range, and optical data. Computer Vision and Image Understanding: CVIU, 80 (2):111–129, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0821;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0821/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0821/ref.

Svensson:2002:DDT

[SB02] Stina Svensson and Gunilla Borgefors. Digital distance transforms in 3D images using information from neighbourhoods up to 5 × 5 × 5.

Computer Vision and Image Understanding: CVIU, 88(1): 24–53, October 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Strand:2005:DTT

[SB05] Robin Strand and Gunilla Borgefors. Distance transforms for three-dimensional grids with non-cubic voxels. Computer Vision and Image Understanding: CVIU, 100 (3):294–311, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shang:2013:FRF

[SB13] Changjing Shang and Dave Barnes. Fuzzy-rough feature selection aided support vector machines for Mars image classification. Computer Vision and Image Understanding: CVIU, 117(3):202–213, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001968.

Shellshear:2013:PPD

[SBA13] Evan Shellshear, Fadi Bitar, and Ulf Assarsson. PDQ: Parallel Distance Queries for deformable meshes. Graphical Models, 75(2):69–78, March 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070313000027.

Saha:2010:DIE

[SBB10] Sankalita Saha, Neal K.

Bambha, and Shuvra S. Bhattacharyya. Design and implementation of embedded computer vision systems based on particle filters. Computer Vision and Image Understanding: CVIU, 114(11):1203– 1214, November 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sbert:2000:OAP

[Sbe00] Mateu Sbert. Optimal absorption probabilities for random walk radiosity. Graphical Models, 62(1):56–70, January 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/artid/gmod.1999.
0513/production;   http:
//www.idealibrary.com/links/
artid/gmod.1999.0513/production/
pdf;   http://www.idealibrary.
com/links/artid/gmod.1999.
0513/production/ref.

Shyu:1999:APL

[SBK^{+}99] Chi-Ren Shyu, Carla E. Brodley, Avinash C. Kak, Akio Kosaka, Alex M. Aisen, and Lynn S. Broderick. AS-SERT: a physician-in-the-loop content-based retrieval system for HRCT image databases. Computer Vision and Image Understanding: CVIU, 75(1–2): 111–132, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0768/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0768/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0768/production/ref.

Shokoufandeh:2006:RMC

[SBM^{+}06] Ali Shokoufandeh, Lars Bretzner, Diego Macrini, M. Fatih Demirci, Clas Jönsson, and Sven Dickinson. The representation and matching of categorical shape. Computer Vision and Image Understanding: CVIU, 103(2): 139–154, August 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sohn:2004:VVC

[SBS04] Bong-Soo Sohn, Chandrajit Bajaj, and Vinay Siddavanahalli. Volumetric video compression for interactive playback. Computer Vision and Image Understanding: CVIU, 96(3):435–452, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shoucri:1985:NDD

[SBT85] R. Shoucri, R. Benesch, and S. Thomas. Note on the determination of a digital straight line from chain codes. Computer Vision, Graphics, and Image Processing, 29 (1):133–139, January 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sudhir:1997:FPC

[SBZ97] G. Sudhir, Subhashis Banerjee, and Andrew Zisserman. Finding point correspondences in motion sequences preserving affine structure. Computer Vision and Image Understanding: CVIU, 68(2):237–246, November 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0545/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0545/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0545/production/ref.

Shen:1992:OLO

[SC92] Jun Shen and Serge Castan.

An optimal linear operator for step edge detection. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(2):112–133, March 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Sharaiha:1993:OAS

[SC93] Y. M. Sharaiha and N. Christofides. An optimal algorithm for the straight segment approximation of digital arcs. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(5):397–407, September 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1993.
                                        1030/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1993.1030/production/
                                        pdf.

Saha:1996:DTU

[SC96] P. K. Saha and B. B. Chaudhuri. 3D digital topology under binary transformation with applications. Computer Vision and Image Understanding: CVIU, 63(3):418–429, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0032/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0032/production/
                                        pdf.

Sapiro:1997:CEI

[SC97a] Guillermo Sapiro and Vicent Caselles. Contrast enhancement via image evolution flows. Graphical Models and Image Processing: GMIP, 59 (6):407–416, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0446/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0446/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0446/production/ref.

Schreiber:1997:UNN

[SC97b] F. A. Schreiber and R. Calvo Wolfler. Use of neural networks to estimate the number of nodes of an edge quadtree. Graphical Models and Image Processing: GMIP, 59(2):61–72, March 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0417/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0417/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1996.
0417/production/ref.

Sziranyi:1998:TCS

[SC98] Tamás Szirányi and Márton Csapodi. Texture classification and segmentation by cellular neural networks using genetic learning. Computer Vision and Image Understanding: CVIU, 71(3): 255–270, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0646/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0646/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0646/production/ref.

Sakkalis:1999:ACA

[SC99a] Takis Sakkalis and Ch. Charitos. Approximating curves via alpha shapes. Graphical Models and Image Processing: GMIP, 61(3):165–176, May 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0496/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1999.0496/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0496/production/ref.

Sato:1999:EGT

[SC99b] Jun Sato and Roberto Cipolla. Extracting group transformations from image moments. Computer Vision and Image Understanding: CVIU, 73(1):29–42, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0702/production;   http:
                                        //www.idealibrary.com/links/

artid/cviu.1998.0702/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0702/production/ref.

Sarkar:2000:MPS

[SC00a] Sudeep Sarkar and Srikanth Chavali. Modeling parameter space behavior of vision systems using Bayesian networks. Computer Vision and Image Understanding: CVIU, 79(2):185–223, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0854;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0854/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0854/ref.

Squire:2000:ISC

[SC00b] David McG. Squire and Terry M. Caelli. Invariance signatures: Characterizing contours by their departures from invariance. Computer Vision and Image Understanding: CVIU, 77(3):284–316, March 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0809;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0809/
pdf;   http://www.idealibrary.

                                        com/links/doi/10.1006/cviu.
                                        2000.0809/ref.

Sturm:2005:FLC

[SCCP05] P. Sturm, Z. L. Cheng, P. C. Y. Chen, and A. N. Poo. Focal length calibration from two views: method and analysis of singular cases. Computer Vision and Image Understanding: CVIU, 99 (1):58–95, July 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Santamaria:2011:CSS

[SCD11] J. Santamaría, O. Cordón, and S. Damas. A comparative study of state-of-the-art evolutionary image registration methods for 3D modeling. Computer Vision and Image Understanding: CVIU, 115(9):1340–1354, September 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001366.

Salvador:2004:CSS

[SCE04] Elena Salvador, Andrea Cavallaro, and Touradj Ebrahimi. Cast shadow segmentation using invariant color features. Computer Vision and Image Understanding: CVIU, 95(2): 238–259, August 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schachter:1976:INM

[Sch76]  Bruce Schachter.  Im-

proved nonlinear mapping algorithm and its application to picture prototype selection. Computer Graphics and Image Processing, 5(2): 271–279, June 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Schachter:1978:NMA

[Sch78] Bruce Schachter. A nonlinear mapping algorithm for large data sets. Computer Graphics and Image Processing, 8(2): 271–276, October 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Schachter:1980:LCW

[Sch80a] Bruce J. Schachter. Long crested wave models. Computer Graphics and Image Processing, 12(2):187– 201, February 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Schrack:1980:CGK

[Sch80b]  G. F. Schrack.  Com-

puter graphics: a keywordindexed bibliography for the years 1976, 1977, and 1978. Computer Graphics and Image Processing, 14(1):24– 79, September 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Schrack:1981:CGK

[Sch81] G. F. Schrack. Computer graphics: a keyword-indexed bibliography for the year 1979. Computer Graphics and Image Processing, 15(1): 45–78, January 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Schrack:1982:CGK

[Sch82]  G. F. Schrack.  Com-

puter graphics — a keywordindexed bibliography for the year 1980. Computer Graphics and Image Processing, 18(2): 145–187, February 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Schunck:1986:IFC

[Sch86]  Brian G. Schunck.  The

image flow constraint equation. Computer Vision, Graphics, and Image Processing, 35(1):20–46, July 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Schrack:1992:FNE

[Sch92] Gunther Schrack. Finding neighbors of equal size in linear quadtrees and octrees in constant time. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(3):221–230, May 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Schaum:1993:TDL

[Sch93] A. Schaum. Theory and design of local interpolators. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55 (6):464–481, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1035/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1035/production/
pdf.

Schindler:2006:GCS

[Sch06] Konrad Schindler. Geometry and construction of straight lines in log-polar images. Computer Vision and Image Understanding: CVIU, 103 (3):196–207, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shin:2013:MOR

[SCL13] Young Min Shin, Minsu Cho, and Kyoung Mu Lee. Multiobject reconstruction from dynamic scenes: an objectcentered approach. Computer Vision and Image Understanding: CVIU, 117(11): 1575–1588, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001276.

Strand:2013:MBD

[SCMS13] Robin Strand, Krzysztof Chris Ciesielski, Filip Malmberg, and Punam K. Saha. The minimum barrier distance. Computer Vision and Image Understanding: CVIU, 117(4):429–437, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001750.

Spagnuolo:2009:SSI

[SCOG09] Michela Spagnuolo, Daniel Cohen-Or, and Xianfeng David Gu. SMI 2008 special issue. Graphical Models, 71 (2):33, March 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070309000137.

Shao:1991:RDM

[SCS91] M. Shao, R. Chellappa, and T. Simchony. Reconstructing a 3D depth map from one or more images. Computer Vision, Graphics, and Image Processing. Image Understanding, 53(2): 219–226, March 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Ser:1999:GAE

[SCS99] P. K. Ser, Clifford S. T. Choy, and W. C. Siu. Genetic algorithm for the extraction of nonanalytic objects from multiple dimensional parameter space. Computer Vision and Image Understanding: CVIU, 73(1):1–13, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0695/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0695/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0695/production/ref.

Sobieranski:2011:LND

[SCvW11] Antonio Carlos Sobieranski, Eros Comunello, and Aldo von Wangenheim. Learning a nonlinear distance metric for supervised region-merging image segmentation. Computer Vision and Image Understanding: CVIU, 115(2): 127–139, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Song:1990:AMF

[SD90] Jisheng Song and Edward J.

Delp. The analysis of morphological filters with multiple structuring elements. Computer Vision, Graphics, and Image Processing, 50(3):308–328, June 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Seales:1992:VOC

[SD92] W. B. Seales and C. R. Dyer. Viewpoint from occluding contour. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (2):198–211, March 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Svensson:2003:SCS

[SdB03] Stina Svensson and Gabriella Sanniti di Baja. Simplifying curve skeletons in volume images. Computer Vision and Image Understanding: CVIU, 90(3): 242–257, June 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sivignon:2004:DIM

[SDC04] Isabelle Sivignon, Florent Dupont, and Jean-Marc Chassery. Digital intersections: minimal carrier, connectivity, and periodicity properties. Graphical Models, 66 (4):226–244, July 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Suetens:1981:SDM

[SDPO81] P. Suetens, P. Dierckx, R. Piessens, and A. Oosterlinck. A semiautomatic digitization method and the use of spline functions in processing line drawings. Computer Graphics and Image Processing, 15(4):390–400, April 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sprague:1991:MAI

[SDR91] A. P. Sprague, M. J. Donahue, and S. I. Rokhlin. A method for automatic inspection of printed circuit boards. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (3):401–415, November 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Saez:2011:EMS

[SE11] Juan M. Sáez and Francisco Escolano. 6DOF entropy minimization SLAM for stereobased wearable devices. Computer Vision and Image Understanding: CVIU, 115(2): 270–285, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Seemuller:1989:EOV

[See89] William W. Seemuller. The extraction of ordered vector drainage networks from elevation data. Computer Vision, Graphics, and Image Processing, 47(1):45–58, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Selesnick:1981:LIT

[Sel81] S. A. Selesnick. Local invariants and twist vectors in computer-aided geometric design. Computer Graphics and Image Processing, 17(2):145– 160, October 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Selfridge:1986:LNB

[Sel86] Peter G. Selfridge. Locating neuron boundaries in electron micrograph images using “primal sketch” primitives. Computer Vision, Graphics, and Image Processing, 34(2):156–165, May 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Serra:1980:BMR

[Ser80] J. Serra. Boolean model and random sets. Computer Graphics and Image Processing, 12(2):99–126, February 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Serra:1986:IMM

[Ser86]  Jean Serra.  Introduction

to mathematical morphology. Computer Vision, Graphics, and Image Processing, 35 (3):283–305, September 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sezan:1990:PDA

[Sez90] M. Ibrahim Sezan. A peak detection algorithm and its application to histogram-based image data reduction. Computer Vision, Graphics, and Image Processing, 49(1):36– 51, January 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Seales:1995:BTD

[SF95] W. Brent Seales and Olivier D.

Faugeras. Building threedimensional object models from image sequences. Computer Vision and Image Understanding: CVIU, 61(3):308–324, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1025/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1025/production/
pdf.

Stark:1996:AAA

[SF96] J. A. Stark and W. J. Fitzgerald. An alternative algorithm for adaptive histogram equalization. Graphical Models and Image Processing: GMIP, 58(2):180–185, March 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0015/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0015/production/
pdf.

Soucy:1997:SRR

[SF97] G. Soucy and F. P. Ferrie. Surface recovery from range images using curvature and motion consistency. Computer Vision and Image Understanding: CVIU, 65(1):1–18, January 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0485/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0485/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0485/production/ref.

Shi:2007:QCT

[SF07] Lilong Shi and Brian Funt. Quaternion color texture segmentation. Computer Vision and Image Understanding: CVIU, 107(1–2):88–96, July/August 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sun:2008:CVM

[SFWG08] Yaoru Sun, Robert Fisher, Fang Wang, and Herman Martins Gomes. A computer vision model for visual-object-based attention and eye movements. Computer Vision and Image Understanding: CVIU, 112(2):126–142, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Slump:1982:NFA

[SG82] Cornelis H. Slump and Jan J.

Gerbrands. A network flow approach to reconstruction of the left ventricle from two projections. Computer Graphics and Image Processing, 18(1): 18–36, January 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sandini:1994:WPV

[SG94] Giulio Sandini and Enrico Grosso. Why purposive vision? Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):109–112, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1040/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1040/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1045/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1045/production/
pdf.

Sengur:2011:CTI

[SG11] Abdulkadir Sengur and Yanhui Guo. Color texture image segmentation based on neutrosophic set and wavelet transformation. Computer Vision and Image Understanding: CVIU, 115(8): 1134–1144, August 2011.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000993.

Shu:2012:AAM

[SGA12] Xianbiao Shu, Chunyu Gao, and Narendra Ahuja. Aperture access and manipulation for computational imaging. Computer Vision and Image Understanding: CVIU, 116 (2):222–237, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002384.

Shin:2001:CED

[SGB01] Min C. Shin, Dmitry B. Goldgof, and Kevin W. Bowyer. Comparison of edge detector performance through use in an object recognition task. Computer Vision and Image Understanding: CVIU, 84(1):160–178, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0932;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0932/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0932/ref.

Song:2001:MPB

[SGDP01] Yang Song, Luis Goncalves, Enrico Di Bernardo, and Pietro Perona. Monocular perception of biological motion in Johansson displays. Computer Vision and Image Understanding: CVIU, 81(3):303–327, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0890;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0890/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0890/ref.

Stefanov:2007:RTH

[SGH07] Nikolay Stefanov, Aphrodite Galata, and Roger Hubbold. A real-time hand tracker using variable-length Markov models of behaviour. Computer Vision and Image Understanding: CVIU, 108(1–2): 98–115, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Senasli:2000:RVL

[SGHM00] M. Senasli, L. Garnero, A. Herment, and E. Mousseaux. 3D reconstruction of vessel lumen from very few angiograms by dynamic contours using a stochastic approach. Graphical Models, 62(2):105–127, March 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        1999.0520;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/gmod.1999.0520/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        1999.0520/ref.

Stokman:2000:CMI

[SGK00] H. M. G. Stokman, Th. Gevers, and J. J. Koenderink. Color measurement by imaging spectrometry. Computer Vision and Image Understanding: CVIU, 79(2):236–249, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0860;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0860/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0860/ref.

Sarris:2001:BTD

[SGS01] Nikos Sarris, Nikos Grammalidis, and Michael G. Strintzis. Building three dimensional head models. Graphical Models, 63(5):333–368, September 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Sabatini:2010:CHC

[SGS^{+}10] Silvio P. Sabatini, Giulia Gastaldi, Fabio Solari, Karl Pauwels, Marc M. Van Hulle, Javier Diaz, Eduardo Ros, Nicolas Pugeault, and Norbert Krüger. A compact harmonic code for early vision based on anisotropic frequency channels. Computer Vision and Image Understanding: CVIU, 114(6): 681–699, June 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sewall:2009:VSS

[SGTL09] Jason Sewall, Nico Galoppo, Georgi Tsankov, and Ming Lin. Visual simulation of shockwaves. Graphical Models, 71(4):126–138, July 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000150.

Suto:1977:SPF

[SH77] Yasuzo Suto and Yoshio Hayashi. Study on picture feature extraction and picture recognition. Computer Graphics and Image Processing, 6 (5):502–510, October 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Suk:1984:EET

[SH84] Minsoo Suk and Soonho Hong. An edge extraction technique for noisy images. Computer Vision, Graphics, and Image Processing, 25(1):24–45, January 1984.

CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Starck:2005:VVS

[SH05] J. Starck and A. Hilton. Virtual view synthesis of people from multiple view video sequences. Graphical Models, 67(6):600–620, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000159.

Starck:2008:MBH

[SH08] Jonathan Starck and Adrian Hilton. Model-based human shape reconstruction from multiple views. Computer Vision and Image Understanding: CVIU, 111(2): 179–194, August 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schmitt:2009:MMD

[SH09] Oliver Schmitt and Maria Hasse. Morphological multiscale decomposition of connected regions with emphasis on cell clusters. Computer Vision and Image Understanding: CVIU, 113(2): 188–201, February 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shapiro:1975:TCD

[Sha75]  S. Shapiro.  Transforma-

tions for the computer detection of curves in noisy pictures. Computer Graphics and Image Processing, 4 (??):??, ???? 1975. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shapiro:1978:PTD

[Sha78] Stephen D. Shapiro. Properties of transforms for the detection of curves in noisy pictures. Computer Graphics and Image Processing, 8(2): 219–236, October 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shapiro:1979:DSP

[Sha79a] Linda G. Shapiro. Data structures for picture processing: a survey. Computer Graphics and Image Processing, 11(2): 162–184, October 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shaw:1979:LRE

[Sha79b]  Gilbert B. Shaw.  Lo-

cal and regional edge detectors: Some comparisons. Computer Graphics and Image Processing, 9(2):135– 149, February 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shanbhag:1994:UIM

[Sha94] A. G. Shanbhag. Utilization of information measure as a means of image thresholding. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(5):414–419, September 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1037/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1037/production/
pdf.

Shah:2005:GSS

[Sha05] Jayant Shah. Gray skeletons and segmentation of shapes. Computer Vision and Image Understanding: CVIU, 99(1): 96–109, July 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shapiro:2006:ASL

[Sha06] Vladimir Shapiro. Accuracy of the straight line Hough Transform: The non-voting approach. Computer Vision and Image Understanding: CVIU, 103(1):1–21, July 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shah:2011:CTS

[Sha11] Mili Shah. Comparing two sets of corresponding six degree of freedom data. Computer Vision and Image Understanding: CVIU, 115(10): 1355–1362, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001378.

Sharma:2012:RPI

[SHC^{+}12] Abhishek Sharma, Murad Al Haj, Jonghyun Choi, Larry S. Davis, and David W. Jacobs. Robust pose invariant face recognition using coupled latent space discriminant analysis. Computer Vision and Image Understanding: CVIU, 116(11):1095–1110, November 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001051.

Silberberg:1986:ORU

[SHD86] Teresa M. Silberberg, David A.

Harwood, and Larry S. Davis. Object recognition using oriented model points. Computer Vision, Graphics, and Image Processing, 35(1):47– 71, July 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sher:1986:ESV

[She86] David Sher. Expert systems for vision based on Bayes’ rule. Computer Vision, Graphics, and Image Processing, 34(1):114– 116, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shen:1996:MED

[She96] Jun Shen. On multi-edge detection. Graphical Models and Image Processing: GMIP, 58(2):101–114, March 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0009/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0009/production/
                                        pdf.

Sheffer:2003:SM

[She03] Alla Sheffer. Skinning 3D meshes. Graphical Models, 65 (5):274–285, September 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Sevigny:1988:TFT

[SHG^{+}88] L. Sevigny, C. Hedegaard, J. P. Gambotto, M. Bohner, S. Grinaker, D. E. Lloyd, L. E. Garn, and J. A. Knecht. A tape format for transferral of image data and source programs. Computer Vision, Graphics, and Image Processing, 41(1):107– 113, January 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shimizu:1981:AGD

[Shi81]  K. Shimizu.  Algorithm

for generating a digital circle on a triangular grid. Computer Graphics and Image Processing, 15(4):401– 402, April 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shiozaki:1983:IED

[Shi83] Akira Shiozaki. Image enhancement in a dithered picture. Computer Vision, Graphics, and Image Processing, 24(1):107–113, October 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shiozaki:1986:EEU

[Shi86] Akira Shiozaki. Edge extraction using entropy operator. Computer Vision, Graphics, and Image Processing, 36(1):1–9, October 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shirai:1994:PCC

[Shi94] Yoshiaki Shirai. Performance characterization in computer vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):260–261, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1054/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1054/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1059/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1059/production/

pdf. See also [DB94, Har94a, Har94b].

Shimshoni:1999:EUL

[Shi99] Ilan Shimshoni. On estimating the uncertainty in the location of image points in 3D recognition from match sets of different sizes. Computer Vision and Image Understanding: CVIU, 74(3):163–173, June 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0755/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0755/production/
pdf.

Sevigny:1983:DCV

[SHJB^{+}83] L. Sevigny, G. Hvedstrup-Jensen, M. Bohner, E. Ostevold, S. Grinaker, and J. Dehne. Discrimination and classification of vehicles in natural scenes from thermal imagery. Computer Vision, Graphics, and Image Processing, 21(3):229– 243, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sibbing:2011:MRS

[SHK11] Dominik Sibbing, Martin Habbecke, and Leif Kobbelt. Markerless reconstruction and synthesis of dynamic facial expressions. Computer Vision and Image Understanding: CVIU, 115(5):668–680, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sim:1998:AMK

[SHKP98] Dong-Gyu Sim, Young Kug Ham, In Kwon Kim, and Rae-Hong Park. Analysis of mixed Korean documents using the branch and bound algorithm based on DP matching. Computer Vision and Image Understanding: CVIU, 71(3):373–384, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0651/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0651/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0651/production/ref.

Shlien:1983:SDC

[Shl83] Seymour Shlien. Segmentation of digital curves using linguistic techniques. Computer Vision, Graphics, and Image Processing, 22(2):277– 286, May 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shneier:1981:CGP

[Shn81a] Michael Shneier. Calculations of geometric properties using quadtrees. Computer Graphics and Image Processing, 16 (3):296–302, July 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shneier:1981:THL

[Shn81b] Michael Shneier. Two hierarchical linear feature representations: Edge pyramids and edge quadtrees. Computer Graphics and Image Processing, 17(3):211– 224, November 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shepp:1979:TFA

[SHS79] L. A. Shepp, S. K. Hilal, and R. A. Schulz. Tuning fork artifact in computerized tomography. Computer Graphics and Image Processing, 10 (3):246–255, July 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Suzuki:2003:LTC

[SHS03] Kenji Suzuki, Isao Horiba, and Noboru Sugie. Lineartime connected-component labeling based on sequential local operations. Computer Vision and Image Understanding: CVIU, 89(1):1–23, January 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shufelt:1997:TAE

[Shu97] Jefferey A. Shufelt. Texture analysis for enhanced color image quantization. Graphical Models and Image Processing: GMIP, 59(3):149–163, May 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0428/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0428/production/
pdf.

Sato:1996:RAC

[SI96] Yoichi Sato and Katsushi Ikeuchi. Reflectance analysis for 3D computer graphics model generation. Graphical Models and Image Processing: GMIP, 58(5):437– 451, September 1996. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0036/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0036/production/
pdf.

Sclaroff:2003:ABR

[SI03] Stan Sclaroff and John Isidoro. Active blobs: regionbased, deformable appearance models. Computer Vision and Image Understanding: CVIU, 89(2–3):197–225, February/March 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sato:1992:MBR

[SIK92] K. Sato, K. Ikeuchi, and T. Kanade. Model based recognition of specular objects using sensor models. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(2): 155–169, March 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Sinha:1987:WIA

[Sin87] R. M. K. Sinha. A widthindependent algorithm for character skeleton estimation. Computer Vision, Graphics, and Image Processing, 40 (3):388–397, December 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sumi:2007:RVA

[SIT07] Yasushi Sumi, Yutaka Ishiyama, and Fumiaki Tomita. Robotvision architecture for realtime 6-DOF object localization. Computer Vision and Image Understanding: CVIU, 105(3):218–230, March 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shah:1984:DTV

[SJ84] Mubarak A. Shah and Ramesh Jain. Detecting time-varying corners. Computer Vision, Graphics, and Image Processing, 28(3):345–355, December 1984. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Skifstad:1989:IIC

[SJ89] Kurt Skifstad and Ramesh Jain. Illumination independent change detection for real world image sequences. Computer Vision, Graphics, and Image Processing, 46(3):387–399, June 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sauer:1993:BBW

[SJ93a] Ken Sauer and Coleen Jones.

Bayesian block-wise segmentation of interframe differences in video sequences. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(2):129–139, March 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1009/production;   http:
//www.idealibrary.com/links/

artid/cgip.1993.1009/production/
pdf.

Shu:1993:DEE

[SJ93b] Chiao-Fe F. Shu and Ramesh C. Jain. Direct estimation and error analysis for oriented patterns. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3):383–398, November 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1049/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1049/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1051/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1051/production/
pdf.

Satherley:2001:VCV

[SJ01] Richard Satherley and Mark W. Jones. Vector-city vector distance transform. Computer Vision and Image Understanding: CVIU, 82(3): 238–254, June 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0915;
 http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0915/pdf;   http://
www.idealibrary.com/links/

doi/10.1006/cviu.2001.0915/
ref.

Shen:2012:DTA

[SJ12] Jingjing Shen and Xiaogang Jin. Detailed traffic animation for urban road networks. Graphical Models, 74 (5):265–282, September 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000227.

Spies:2002:RFE

[SJB02] Hagen Spies, Bernd Jähne, and John L. Barron. Range flow estimation. Computer Vision and Image Understanding: CVIU, 85(3):209– 231, March 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schmugge:2007:OEA

[SJST07] Stephen J. Schmugge, Sriram Jayaram, Min C. Shin, and Leonid V. Tsap. Objective evaluation of approaches of skin detection using ROC analysis. Computer Vision and Image Understanding: CVIU, 108(1–2):41–51, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sahasrabudhe:1979:SVI

[SK79] S. C. Sahasrabudhe and A. D.

Kulkarni. Shift variant image degradation and restoration using svd. Computer Graphics and Image Processing, 9 (3):203–212, March 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Shafer:1983:USF

[SK83a] Steven A. Shafer and Takeo Kanade. Using shadows in finding surface orientations. Computer Vision, Graphics, and Image Processing, 22(1):145–176, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Stuller:1983:KFF

[SK83b] John Stuller and Goplan Krishnamurthy. Kalman filter formulation of low-level television image motion estimation. Computer Vision, Graphics, and Image Processing, 21 (2):169–204, February 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Suk:1984:NMS

[SK84] Minsoo Suk and Hwanil Kang. New measures of similarity between two contours based on optimal bivariate transforms. Computer Vision, Graphics, and Image Processing, 26(2):168–182, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Smith:1985:ASD

[SK85] David R. Smith and Takeo Kanade. Autonomous scene description with range imagery. Computer Vision, Graphics, and Image Processing, 31(3):322–334, September 1985. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shapiro:1986:CIT

[SK86] Linda Shapiro and Avi Kak, editors. Current Issues and Trends in Computer Vision, volume 36(2/3) of Computer Vision, Graphics, and Image Processing. Academic Press, New York, NY, USA, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sinha:1988:PBS

[SK88] R. M. K. Sinha and H. C.

Karnick. Plang based specification of patterns with variations for pictorial data bases. Computer Vision, Graphics, and Image Processing, 43(1):98–110, July 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Soffer:1998:GCH

[SK98] Menashe Soffer and Nahum Kiryati. Guaranteed convergence of the Hough transform. Computer Vision and Image Understanding: CVIU, 69(2):119–134, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0557/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0557/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0557/production/ref.

Stolte:2001:NTR

[SK01] Nilo Stolte and Arie Kaufman. Novel techniques for robust voxelization and visualization of implicit surfaces. Graphical Models, 63(6):387– 412, November 2001. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Sablatnig:2002:MBR

[SK02] Robert Sablatnig and Martin Kampel. Model-based registration of front- and backviews of rotationally symmetric objects. Computer Vision and Image Understanding: CVIU, 87(1–3):90–103, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shimshoni:1996:GSS

[SKB96] Ilan Shimshoni, Ron Kimmel, and Alfred M. Bruckstein. Global shape from shading. Computer Vision and Image Understanding:

CVIU, 64(1):188–189, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0053/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0053/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0054/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0054/production/
pdf.

Steiner:1998:PSE

[SKB98] Ami Steiner, Ron Kimmel, and Alfred M. Bruckstein. Planar shape enhancement and exaggeration. Graphical Models and Image Processing: GMIP, 60(2):112–124, March 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0461/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0461/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0461/production/ref;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0465/production;
 http://www.idealibrary.
com/links/artid/gmip.1998.
0465/production/pdf;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0465/production/
ref.

Sun:2013:ODS

[SKBS13] Min Sun, Shyam Sunder Kumar, Gary Bradski, and Silvio Savarese. Object detection, shape recovery, and 3D modelling by depth-encoded Hough voting. Computer Vision and Image Understanding: CVIU, 117(9): 1190–1202, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000969.

Shekhovtsov:2008:EMD

[SKH08] Alexander Shekhovtsov, Ivan Kovtun, and Václav Hlaváč. Efficient MRF deformation model for non-rigid image matching. Computer Vision and Image Understanding: CVIU, 112(1):91–99, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shafer:1983:GSU

[SKK83] Steven A. Shafer, Takeo Kanade, and John R. Kender. Gradient space under orthography and perspective. Computer Vision, Graphics, and Image Processing, 21(3):182–199, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sminchisescu:2006:CMC

[SKM06] Cristian Sminchisescu, Atul Kanaujia, and Dimitris Metaxas.

Conditional models for contextual human motion recognition. Computer Vision and Image Understanding: CVIU, 104(2–3):210–220, November/December 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Skolnick:1986:AMT

[Sko86] Michael M. Skolnick. Application of morphological transformations to the analysis of two-dimensional electrophoretic gels of biological materials. Computer Vision, Graphics, and Image Processing, 35(3):306– 332, September 1986. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sakane:1995:PFA

[SKOS95] Shigeyuki Sakane, Toshiji Kuruma, Toru Omata, and Tomomasa Sato. Planning focus of attention for multifingered hand with consideration of time-varying aspects. Computer Vision and Image Understanding: CVIU, 61(3):445–453, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1033/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1033/production/
pdf.

Siddiqi:1997:GSC

[SKS97] Kaleem Siddiqi, Benjamin B. Kimia, and Chi-Wang Shu. Geometric shock-capturing ENO schemes for subpixel interpolation, computation and curve evolution. Graphical Models and Image Processing: GMIP, 59(5):278– 301, September 1997. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0438/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0438/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0438/production/ref.

Strandmark:2011:PDV

[SKS11] Petter Strandmark, Fredrik Kahl, and Thomas Schoenemann. Parallel and distributed vision algorithms using dual decomposition. Computer Vision and Image Understanding: CVIU, 115(12): 1721–1732, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001652.

Smith:2008:RCR

[SKSR08] Eric R. Smith, Bradford J. King, Charles V. Stewart, and Richard J. Radke. Registration of combined range– intensity scans: Initialization through verification. Computer Vision and Image Understanding: CVIU, 110(2): 226–244, May 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shinozaki:2009:CCI

[SKU^{+}09] Megumi Shinozaki, Masato Kusanagi, Kazunori Umeda, Guy Godin, and Marc Rioux. Correction of color information of a 3D model using a range intensity image. Computer Vision and Image Understanding: CVIU, 113(11): 1170–1179, November 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Smeets:2013:MLS

[SKVS13] Dirk Smeets, Johannes Keustermans, Dirk Vandermeulen, and Paul Suetens. mesh-SIFT: Local surface features for 3D face recognition under expression variations and partial data. Computer Vision and Image Understanding: CVIU, 117 (2):158–169, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001324.

Sedgwick:1985:ECV

[SL85] H. A. Sedgwick and S. Levy.

Environment-centered and viewer-centered perception of surface orientation. Computer Vision, Graphics, and Image Processing, 31(2):248– 260, August 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Soucy:1996:MSM

[SL96] Marc Soucy and Denis Laurendeau. Multiresolution surface modeling based on hierarchical triangulation. Computer Vision and Image Understanding: CVIU, 63(1):1–14, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0001/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0001/production/
pdf.

Smith:1999:ICQ

[SL99] John R. Smith and Chung-Sheng Li. Image classification and querying using composite region templates. Computer Vision and Image Understanding: CVIU, 75 (1–2):165–174, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0771/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0771/production/
pdf;   http://www.idealibrary.

com/links/artid/cviu.1999.
0771/production/ref.

Shu:2000:EMC

[SLB^{+}00] Huazhong Shu, Limin Luo, Xudong Bao, Wenxue Yu, and Guoniu Han. An efficient method for computation of Legendre moments. Graphical Models, 62(4):237–262, July 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0523;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0523/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0523/ref.

Sperling:1985:IEA

[SLCP85] George Sperling, Michael Landy, Yoav Cohen, and M. Pavel. Intelligible encoding of ASL image sequences at extremely low information rates. Computer Vision, Graphics, and Image Processing, 31 (3):335–391, September 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shneier:1986:MBS

[SLK86] Michael O. Shneier, Ronald Lumia, and Ernest W. Kent. Model-based strategies for high-level robot vision. Computer Vision, Graphics, and Image Processing, 33(3):293– 306, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Son:2011:SGD

[SLKL11] Minjung Son, Yunjin Lee, Henry Kang, and Seungyong Lee. Structure grid for directional stippling. Graphical Models, 73(3):74–87, May 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000433.

Song:2001:MBO

[SLL01] Bong Seop Song, Kyoung Mu Lee, and Sang Uk Lee. Modelbased object recognition using geometric invariants of points and lines. Computer Vision and Image Understanding: CVIU, 84(3): 361–383, December 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Stoksik:1995:PSA

[SLN95] Marc A. Stoksik, R. G. Lane, and D. T. Nguyen. Practical synthesis of accurate fractal images. Graphical Models and Image Processing: GMIP, 57(3):206–219, May 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1020/production;   http:
                                        //www.idealibrary.com/links/

artid/gmip.1995.1020/production/
pdf.

Sole:2001:CED

[SLS01] Andres Fco. Solé, Antonio López, and Guillermo Sapiro. Crease enhancement diffusion. Computer Vision and Image Understanding: CVIU, 84(2):241–248, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0945;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0945/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0945/ref.

Sebe:2003:VRS

[SLS03] Nicu Sebe, Michael S. Lew, and Arnold W. M. Smeulders. Video retrieval and summarization. Computer Vision and Image Understanding: CVIU, 92(2–3):141–146, November/December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sclaroff:1999:UTV

[SLST99] Stan Sclaroff, Marco La Cascia, Saratendu Sethi, and Leonid Taycher. Unifying textual and visual cues for content-based image retrieval on the world wide web. Computer Vision and Image Understanding: CVIU, 75(1–2):86–98, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0765/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0765/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0765/production/ref.

Shih:1989:SAE

[SLY89] Zen-Chung C. Shih, R. C. T. Lee, and S. N. Yang. A systolic algorithm for extracting regions from a planar graph. Computer Vision, Graphics, and Image Processing, 47 (2):227–242, August 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Simard:1990:VFR

[SM90] P. Y. Simard and G. E. Mailloux. Vector field restoration by the method of convex projections. Computer Vision, Graphics, and Image Processing, 52(3):360– 385, December 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shufelt:1993:FMC

[SM93] Jefferey A. Shufelt and David M. McKeown, Jr. Fusion of monocular cues to detect man-made structures in aerial imagery. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):307–330, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1021/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1021/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1021/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1021/production/
pdf.

Strickland:1994:CME

[SM94] Robin N. Strickland and Zuhua H. Mao. Contour motion estimation using relaxation matching with a smoothness constraint on the velocity field. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):157–167, September 1994. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1044/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1044/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1049/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1049/production/
pdf.

Syeda-Mahmood:1997:DMD

[SM97] Tanveer Fathima Syeda-Mahmood. Data- and modeldriven selection using parallel line groups. Computer Vision and Image Understanding: CVIU, 67(3): 205–222, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0542/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0542/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0542/production/ref.

Syeda-Mahmood:1999:DPS

[SM99] T. F. Syeda-Mahmood. Detecting perceptually salient texture regions in images. Computer Vision and Image Understanding: CVIU, 76(1):93–108, October 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0784/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0784/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0784/production/ref.

Sekkati:2006:JOF

[SM06] H. Sekkati and A. Mitiche. Joint optical flow estimation, segmentation, and 3D interpretation with level sets. Computer Vision and Image Understanding: CVIU, 103 (2):89–100, August 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sugano:2010:HOI

[SM10] Hiroki Sugano and Ryusuke Miyamoto. Highly optimized implementation of OpenCV for the Cell Broadband Engine. Computer Vision and Image Understanding: CVIU, 114(11):1273–1281, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

SanMiguel:2012:SBP

[SM12] Juan C. SanMiguel and José M. Martínez. A semantic-based probabilistic approach for real-time video event recognition. Computer Vision and Image Understanding: CVIU, 116(9): 937–952, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000665.

Sparks:2013:SSM

[SM13a] Rachel Sparks and Anant Madabhushi. Statistical shape model for manifold regularization: Gleason grading of prostate histology. Computer Vision and Image Understanding: CVIU, 117(9): 1138–1146, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000672.

Subburaman:2013:AST

[SM13b] Venkatesh Bala Subburaman and Sébastien Marcel. Alternative search techniques for face detection using location estimation and binary features. Computer Vision and Image Understanding: CVIU, 117(5):551–570, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000039.

Starck:1995:MSA

[SMB95] Jean-Luc Starck, Fionn Murtagii, and Albert Bijaoui. Multiresolution support applied to image filtering and restoration. Graphical Models and Image Processing: GMIP, 57(5):420–431, September 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1036/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1036/production/
                                        pdf.

Singh:2008:LVC

[SMD^{+}08] Vikas Singh, Lopamudra Mukherjee, Petru M. Dinu, Jinhui Xu, and Kenneth R. Hoffmann. Limited view CT reconstruction and segmentation via constrained metric labeling. Computer Vision and Image Understanding: CVIU, 112(1):67–80, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Syeda-Mahmood:2004:CSI

[SMHH04] Tanveer Syeda-Mahmood, Ismail Haritaoglu, and Thomas Huang. CVIU special issue on event detection in video. Computer Vision and Image Understanding: CVIU, 96(2): 97–99, November 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Smith:1985:POR

[Smi85] Beverly J. Smith. Perception of organization in a random stimulus. Computer Vision, Graphics, and Image Processing, 31(2):242– 247, August 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sarkar:2002:POB

[SMK02] Sudeep Sarkar, Daniel Majchrzak, and Kishore Korimilli. Perceptual organization based computational model for robust segmentation of moving objects. Computer Vision and Image Understanding: CVIU, 86(3):141–170, June 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Saha:1998:LTP

[SMR98] P. K. Saha, D. Dutta Majumder, and Azriel Rosenfeld. Local topological parameters in a tetrahedral representation. Graphical Models and Image Processing: GMIP, 60 (6):423–436, November 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0481/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1998.0481/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0481/production/ref.

Seo:2004:EBA

[SMT04] Hyewon Seo and Nadia Magnenat-Thalmann. An example-based approach to human body manipulation. Graphical Models, 66(1):1– 23, January 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Selinger:1999:PGH

[SN99] Andrea Selinger and Randal C. Nelson. A perceptual grouping hierarchy for appearance-based 3D object recognition. Computer Vision and Image Understanding: CVIU, 76(1):83–92, October 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0788/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0788/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0788/production/ref.

Semwal:2001:SFU

[SO01] Sudhanshu K. Semwal and Jun Ohya. Spatial filtering using the active-space indexing method. Graphical Models, 63(3):135–150, May 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0540;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0540/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0540/ref.

Shahid:2007:ILS

[SO07] Kamran Shahid and Galina Okouneva. Intelligent LI-DAR scanning region selection for satellite pose estimation. Computer Vision and Image Understanding: CVIU, 107(3):203–209, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sobel:1978:NCB

[Sob78] Irwin Sobel. Neighborhood coding of binary images for fast contour following and general binary array processing. Computer Graphics and Image Processing, 8(1):127– 135, August 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Smeaton:2010:VSB

[SOD10] Alan F. Smeaton, Paul Over, and Aiden R. Doherty. Video shot boundary detection: Seven years of TRECVid activity. Computer Vision and Image Understanding: CVIU, 114(4):411–418, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Steinemann:2009:SMD

[SOG09] Denis Steinemann, Miguel A. Otaduy, and Markus Gross. Splitting meshless deforming objects with explicit surface tracking. Graphical Models, 71(6):209–220, November 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070309000034.

Sobh:1995:IIR

[SOJ^{+}95] Tarek M. Sobh, J. Owen, C. Jaynes, M. Dekhil, and T. C. Henderson. Industrial inspection and reverse engineering. Computer Vision and Image Understanding: CVIU, 61(3):468–474, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1035/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1035/production/
pdf.

Soroka:1981:GCS

[Sor81] Barry I. Soroka. Generalized cones from serial sections. Computer Graphics and Image Processing, 15(2):154– 166, February 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Selfridge:1981:ODA

[SP81] Peter G. Selfridge and Judith M. S. Prewitt. Organ detection in abdominal computerized tomography scans

— application to the kidney. Computer Graphics and Image Processing, 15(3): 265–278, March 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Scarlatos:1992:HTU

[SP92] Lori Scarlatos and Theo Pavlidis. Hierarchical triangulation using cartographic coherence. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(2): 147–161, March 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Salzenstein:1997:PEH

[SP97a] Fabien Salzenstein and Wojciech Pieczynski. Parameter estimation in hidden fuzzy Markov random fields and image segmentation. Graphical Models and Image Processing: GMIP, 59(4):205–220, July 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0431/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0431/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0431/production/ref.

Shimshoni:1997:RSP

[SP97b] Ilan Shimshoni and Jean Ponce. Recovering the shape of polyhedra using line-drawing analysis and complex reflectance models. Computer Vision and Image Understanding: CVIU, 65 (2):296–310, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0569/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0569/production/

pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0569/production/ref.

Shnaider:1997:ICT

[SP97c] Mikhail Shnaider and Andrew P. Papliński. Image coding through D lattice quantization of wavelet coefficients. Graphical Models and Image Processing: GMIP, 59(4):193–204, July 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0429/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0429/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0429/production/ref.

Sim:1997:TSA

[SP97d] Dong-Gyu Sim and Rae-Hong Park. A two-stage algorithm for motion discontinuitypreserving optical flow estimation. Computer Vision and Image Understanding: CVIU, 65(1):19–37, January 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0483/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0483/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0483/production/ref.

Sinha:2006:PTZ

[SP06] Sudipta N. Sinha and Marc Pollefeys. Pan-tilt-zoom camera calibration and highresolution mosaic generation. Computer Vision and Image Understanding: CVIU, 103 (3):170–183, September 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Spetsakis:1992:LAP

[Spe92]  Minas E. Spetsakis.  A

linear algorithm for point and line-based structure from motion. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (2):230–241, September 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Spetsakis:1994:MSV

[Spe94] Minas Spetsakis. Models of statistical visual motion estimation. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3):300–312, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1994.
                                        1059/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1994.1059/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1994.
                                        1065/production;   http:
                                        //www.idealibrary.com/links/

artid/cviu.1994.1065/production/
pdf.

Spetsakis:1997:OFE

[Spe97] Minas E. Spetsakis. Optical flow estimation using discontinuity conforming filters. Computer Vision and Image Understanding: CVIU, 68(3):276–289, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0555/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0555/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0555/production/ref.

Spitz:1998:ACD

[Spi98] A. Lawrence Spitz. Analysis of compressed document images for dominant skew, multiple skew, and logotype detection. Computer Vision and Image Understanding: CVIU, 70(3):321–334, June 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0686/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0686/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0686/production/ref.

Shyu:2002:UHP

[SPK^{+}02] Chi-Ren Shyu, Christina Pavlopoulou, Avinash C. Kak, Carla E. Brodley, and Lynn S. Broderick. Using human perceptual categories for content-based retrieval from a medical image database. Computer Vision and Image Understanding: CVIU, 88(3): 119–151, December 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shapiro:1981:SGB

[SPS81] Bruce Shapiro, Jim Pisa, and Jack Sklansky. Skeleton generation from X, Y boundary sequences. Computer Graphics and Image Processing, 15 (2):136–153, February 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sherbrooke:1996:DTP

[SPW96] Evan C. Sherbrooke, Nicholas M. Patrikalakis, and Franz-Erich Wolter. Differential and topological properties of medial axis transforms. Graphical Models and Image Processing: GMIP, 58(6):574– 592, November 1996. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0047/production;   http:
                                        //www.idealibrary.com/links/

artid/gmip.1996.0047/production/
pdf.

Saha:2000:DTS

[SR00] Punam K. Saha and Azriel Rosenfeld. The digital topology of sets of convex voxels. Graphical Models, 62(5):343–352, September 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0527;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0527/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0527/ref.

Segvic:2009:MLF

[ŠRDC09] Siniša Šegvić, Anthony Remazeilles, Albert Diosi, and François Chaumette. A mapping and localization framework for scalable appearancebased navigation. Computer Vision and Image Understanding: CVIU, 113(2): 172–187, February 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Suau:2013:DEE

[SRHC13] Xavier Suau, Javier Ruiz-Hidalgo, and Josep R. Casas. Detecting end-effectors on 2.5D data using geometric deformable models: Application to human pose estimation. Computer Vision and Image Understanding: CVIU, 117(3):281–288, March 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001907.

Suzuki:2002:SIN

[SRK02a] Hiromasa Suzuki, Alyn Rockwood, and Leif Kobbelt. Special issue on the Ninth Pacific Graphics Conference (PG 2001). Graphical Models, 64 (2):59–60, March 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Szymczak:2002:PRM

[SRK02b] Andrzej Szymczak, Jarek Rossignac, and Davis King. Piecewise regular meshes: Construction and compression. Graphical Models, 64(3– 4):183–198, May 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Siero:1982:CDP

[SRL82] P. L. J. Siero, G. Rozenberg, and A. Lindenmayer. Cell division patterns: Syntactical description and implementation. Computer Graphics and Image Processing, 18(4): 329–346, April 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sun:2009:NAS

[SRML09] Xianfang Sun, Paul L.

Rosin, Ralph R. Martin, and Frank C. Langbein. Noise analysis and synthesis for 3D laser depth scanners. Graphical Models, 71 (2):34–48, March 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000337.

Subashini:2010:AAB

[SRP10] T. S. Subashini, V. Ramalingam, and S. Palanivel. Automated assessment of breast tissue density in digital mammograms. Computer Vision and Image Understanding: CVIU, 114(1):33–43, January 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sole-Ribalta:2011:MAC

[SRS11] Albert Solé-Ribalta and Francesc Serratosa. Models and algorithms for computing the common labelling of a set of attributed graphs. Computer Vision and Image Understanding: CVIU, 115(7):929–945, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000750.

Sun:2001:ECF

[SRT01] Zhaohui Sun, Visvanathan Ramesh, and A. Murat Tekalp. Error characterization of the factorization method. Computer Vision and Image Understanding: CVIU, 82(2):110–137, May 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0910;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0910/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0910/ref.

Safaee-Rad:1991:APE

[SRTBS91] R. Safaee-Rad, I. Tchoukanov, B. Benhabib, and K. C. Smith. Accurate parameter estimation of quadratic curves from Grey-Level Images. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (2):259–274, September 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Siromoney:1976:HAR

[SS76] Gift Siromoney and Rani Siromoney. Hexagonal arrays and rectangular blocks. Computer Graphics and Image Processing, 5(3):353– 381, September 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sankar:1978:PPD

[SS78] P. V. Sankar and C. U.

Sharma. A parallel procedure for the detection of dominant points on a digital curve. Computer Graphics and Image Processing, 17 (3):403–412, June 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Suresh:1979:SSR

[SS79] B. R. Suresh and B. A.

Shenoi. State-space realization of a certain class of twodimensional systems with applications to image processing. Computer Graphics and Image Processing, 11(2):101– 110, October 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sugihara:1984:RRS

[SS84a] Kokichi Sugihara and Noboru Sugie. Recovery of rigid structure from orthographically projected optical flow. Computer Vision, Graphics, and Image Processing, 27 (3):309–320, September 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Suk:1984:CFE

[SS84b] Minsoo Suk and Ohyoung Song. Curvilinear feature extraction using minimum spanning trees. Computer Vision, Graphics, and Image Processing, 26(3):400–411, June 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shaffer:1987:OQC

[SS87] Clifford A. Shaffer and Hanan Samet. Optimal quadtree construction algorithms. Computer Vision, Graphics, and Image Processing, 37(3):402– 419, March 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shaffer:1990:SOU

[SS90a] Clifford A. Shaffer and Hanan Samet. Set operations for unaligned linear quadtrees. Computer Vision, Graphics, and Image Processing, 50(1): 29–49, April 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Singh:1990:GLC

[SS90b] Ajit Singh and Michael Shneier. Grey level corner detection. A generalization and a robust real time implementation. Computer Vision, Graphics, and Image Processing, 51(1):54–69, July 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shaffer:1991:LTD

[SS91] Clifford A. Shaffer and Quentin F. Stout. Linear time distance transforms for Quadtrees. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (2):215–223, September 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Schmitz:1995:CPR

[SS95a] Barbara E. Schmitz and Robert L. Stevenson. Color palette restoration. Graphical Models and Image Processing: GMIP, 57(5):409– 419, September 1995. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1035/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1035/production/
pdf.

Sederberg:1995:RRS

[SS95b] Thomas W. Sederberg and Takafumi Saito. Rationalruled surfaces: Implicitization and section curves. Graphical Models and Image Processing: GMIP, 57(4):334–342, July 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1029/production;   http:
//www.idealibrary.com/links/

                                        artid/gmip.1995.1029/production/
                                        pdf.

Shamir:2006:SBS

[SS06] Ariel Shamir and Amir Shaham. Skeleton based solid representation with topology preservation. Graphical Models, 68(3):307–321, May 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070305000779.

Shirazi:2009:OED

[SS09] Sayed Kamaledin Ghiasi Shirazi and Reza Safabakhsh. Omnidirectional edge detection. Computer Vision and Image Understanding: CVIU, 113(4):556–564, April 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sadeghi:2011:SRL

[SS11a] Javad Sadeghi and Faramarz F. Samavati. Smooth reverse loop and Catmull–Clark subdivision. Graphical Models, 73(5):202–217, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070311000117.

Sundareswara:2011:BDC

[SS11b] Rashmi Sundareswara and Paul R. Schrater. Bayesian discounting of camera parameter uncertainty for optimal 3D reconstruction from images. Computer Vision and Image Understanding: CVIU, 115(1):117–126, January 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Smith:2006:IML

[SSdVL06] Paul Smith, Mubarak Shah, and Niels da Vitoria Lobo. Integrating multiple levels of zoom to enable activity analysis. Computer Vision and Image Understanding: CVIU, 103(1):33–51, July 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sankar:1994:CSG

[SSF94] P. V. Sankar, M. J. Silbermann, and L. A. Ferrari. Curve and surface generation and refinement based on a high speed derivative algorithm. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):94–101, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1008/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1008/production/
pdf.

Shah:1986:PSE

[SSJ86] Mubarak Shah, Arun Sood, and Ramesh Jain. Pulse and staircase edge models. Computer Vision, Graphics, and Image Processing, 34(3):321–343, June 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Saracchini:2012:RMS

[SSL^{+}12] Rafael F. V. Saracchini, Jorge Stolfi, Helena C. G. Leitão, Gary A. Atkinson, and Melvyn L. Smith. A robust multi-scale integration method to obtain the depth from gradient maps. Computer Vision and Image Understanding: CVIU, 116(8):882–895, August 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000537.

Shi:2006:HEL

[SSM06] J. Shi, A. Samal, and D. Marx. How effective are landmarks and their geometry for face recognition? Computer Vision and Image Understanding: CVIU, 102(2): 117–133, May 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sankar:1978:COS

[SSN78] P. V. Sankar, C. U. Sharma, and R. Narasimhan. Computing the organizations and shapes of two-dimensional dot patterns. Computer Graphics and Image Processing, 8(2): 203–218, October 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Socolinsky:2003:FRV

[SSN03] Diego A. Socolinsky, Andrea Selinger, and Joshua D. Neuheisel. Face recognition with visible and thermal infrared imagery. Computer Vision and Image Understanding: CVIU, 91(1–2): 72–114, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sakkalis:2001:TGP

[SSP01a] T. Sakkalis, G. Shen, and N. M. Patrikalakis. Topological and geometric properties of interval solid models. Graphical Models, 63 (3):163–175, May 2001. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0539;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2001.0539/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2001.0539/ref.

Shen:2001:BRM

[SSP01b] G. Shen, T. Sakkalis, and N. M. Patrikalakis. Boundary representation model rectification. Graphical Models, 63(3):177–195, May 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2001.0543;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/gmod.2001.0543/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2001.0543/ref.

Siromoney:1982:STA

[SSS82] Gift Siromoney, Rani Siromoney, and K. G. Subramanian. Stochastic table arrays. Computer Graphics and Image Processing, 18(2):202– 211, February 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Soladie:2013:IRF

[SSS13] Catherine Soladié, Nicolas Stoiber, and Renaud Séguier. Invariant representation of facial expressions for blended expression recognition on unknown subjects. Computer Vision and Image Understanding: CVIU, 117(11): 1598–1609, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213001355.

Sturm:2006:SIO

[SST06] Peter Sturm, Tomas Svoboda, and Seth Teller. Special issue: Omnidirectional vision and camera networks. Computer Vision and Image Understanding: CVIU, 103(3): 155, September 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Sahoo:1988:STT

[SSWC88] P. K. Sahoo, S. Soltani, A. K. C. Wong, and Y. C. Chen. A survey of thresholding techniques. Computer Vision, Graphics, and Image Processing, 41(2):233– 260, February 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sandini:1980:ARL

[ST80] Giulio Sandini and Vincenzo Tagliasco. Anthropomorpic retina-like structure for scene analysis. Computer Graphics and Image Processing, 14(4): 365–372, December 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sato:1996:DPC

[ST96] Yoshinobu Sato and Shinichi Tamura. Detecting planar and curved symmetries of 3D shapes from a range image. Computer Vision and Image Understanding: CVIU, 64(1):175–187, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0052/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0052/production/
                                        pdf.

Shao:2007:AP

[ST07] Wei Shao and Demetri Terzopoulos. Autonomous pedestrians. Graphical Models, 69(5–6):246–274, September/November 2007. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070307000252.

Shubina:2010:VSO

[ST10] Ksenia Shubina and John K. Tsotsos. Visual search for an object in a 3D environment using a mobile robot. Computer Vision and Image Understanding: CVIU, 114(5): 535–547, May 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Stark:1995:FOR

[Sta95] L. Stark. Functionality in object recognition. Computer Vision and Image Understanding: CVIU, 62(2): 145–??, ???? 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Stahlhut:2005:ENT

[Sta05] O. Stahlhut. Extending natural textures with multi-scale synthesis. Graphical Models, 67(6):496–517, November 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000147.

Sternberg:1986:GM

[Ste86] Stanley R. Sternberg. Grayscale morphology. Computer Vision, Graphics, and Image Processing, 35(3):333–355, September 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Stevens:2001:EIC

[Ste01] Mark R. Stevens. Evaluating 2D image comparison metrics for 3D scene interpretation. Computer Vision and Image Understanding: CVIU, 84(1):179–197, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0940;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0940/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0940/ref.

Steger:2013:UEL

[Ste13]  Carsten Steger.  Unbi-

ased extraction of lines with parabolic and Gaussian profiles. Computer Vision and Image Understanding: CVIU, 117(2):97–112, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421200118X.

Saber:1996:AIA

[STEK96] Eli Saber, A. Murat Tekalp, Reiner Eschbach, and Keith Knox. Automatic image annotation using adaptive color classification. Graphical Models and Image Processing: GMIP, 58(2):115–126, March 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0010/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0010/production/
pdf.

Sebe:2008:SMC

[STLH08] Nicu Sebe, Qi Tian, Michael S. Lew, and Thomas S. Huang. Similarity Matching in Computer Vision and Multimedia. Computer Vision and Image Understanding: CVIU, 110 (3):309–311, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Stockman:1987:ORL

[Sto87] George C. Stockman. Object recognition and localization via pose clustering. Computer Vision, Graphics, and Image Processing, 40 (3):361–387, December 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Stuller:1976:LRE

[Stu76] J. A. Stuller. Linear resolution enhancement. Computer Graphics and Image Processing, 5(3):291–318, September 1976. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Savarese:2009:SIR

[STV09] Silvio Savarese, Tinne Tuytelaars, and Luc Van Gool. Special issue on 3D representation for object and scene recognition. Computer Vision and Image Understanding: CVIU, 113(12):1181–1182, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Saha:2001:FCO

[SU01a] Punam K. Saha and Jayaram K. Udupa. Fuzzy connected object delineation: Axiomatic path strength definition and the case of multiple seeds. Computer Vision and Image Understanding: CVIU, 83(3):275–295, September 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.

2001.0927;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0927/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0927/ref.

Saha:2001:RFC

[SU01b] Punam K. Saha and Jayaram K. Udupa. Relative fuzzy connectedness among multiple objects: Theory, algorithms, and applications in image segmentation. Computer Vision and Image Understanding: CVIU, 82(1):42–56, April 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0902;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0902/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0902/ref.

Subramanian:1979:HAG

[Sub79] K. G. Subramanian. Hexagonal array grammars. Computer Graphics and Image Processing, 10(4):388– 394, August 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Subbarao:1990:BTC

[Sub90] Muralidhara Subbarao. Bounds on time-to-collision and rotational component from firstorder derivatives of image flow. Computer Vision, Graphics, and Image Processing, 50(3):329–341, June 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sugihara:1978:PLS

[Sug78] Kokichi Sugihara. Picture language for skeletal polyhedra. Computer Graphics and Image Processing, 8(3):382– 405, December 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Sugihara:1988:SLP

[Sug88] Kokichi Sugihara. Some location problems for robot navigation using a single camera. Computer Vision, Graphics, and Image Processing, 42(1):112–129, April 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sugihara:1993:AGV

[Sug93]  Kokichi Sugihara.  Ap-

proximation of generalized Voronoi diagrams by ordinary Voronoi diagrams. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(6):522–531, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1039/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/cgip.1993.1039/production/
                                        pdf.

Saha:2000:SBF

[SUO00] Punam K. Saha, Jayaram K. Udupa, and Dewey Odhner. Scale-based fuzzy connected image segmentation: Theory, algorithms, and validation. Computer Vision and Image Understanding: CVIU, 77(2):145–174, February 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0813;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.1999.0813/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0813/ref.

Supanekar:1979:CSK

[Sup79]  S. D. Supanekar.  Com-

puter study of knots. Computer Graphics and Image Processing, 11(2):150– 161, October 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Super:2002:FRI

[Sup02] Boaz J. Super. Fast retrieval of isolated visual shapes. Computer Vision and Image Understanding: CVIU, 85(1): 1–21, January 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Suri:1986:LTA

[Sur86]  Subhash Suri.  A linear

time algorithm for minimum link paths inside a simple polygon. Computer Vision, Graphics, and Image Processing, 35(1):99–110, July 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Santos-Victor:1997:VBD

[SVS97] José Santos-Victor and Giulio Sandini. Visual behaviors for docking. Computer Vision and Image Understanding: CVIU, 67(3): 223–238, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0528/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0528/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0528/production/ref.

Shen:1983:GTR

[SW83a] Helen C. Shen and Andrew K. C. Wong. Generalized texture representation and metric. Computer Vision, Graphics, and Image Processing, 23(2):187–206, August 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Sun:1983:NGL

[SW83b] Chengjun Sun and William G. Wee. Neighboring gray level dependence matrix for texture classification. Computer Vision, Graphics, and Image Processing, 23(3):341– 352, September 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Subbarao:1986:CFS

[SW86] Muralidhara Subbarao and Allen M. Waxman. Closed form solutions to image flow equations for planar surfaces in motion. Computer Vision, Graphics, and Image Processing, 36(2/3):208–228, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Shih:1994:IFA

[SW94] F. Y. Shih and Wai-Tak T. Wong. An improved fast algorithm for the restoration of images based on chain codes description. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(4):348–351, July 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1994.
                                        1031/production;   http:
                                        //www.idealibrary.com/links/

artid/cgip.1994.1031/production/
pdf.

Shih:2004:FED

[SW04] Frank Y. Shih and Yi-Ta Wu.

Fast Euclidean distance transformation in two scans using a 3 × 3 neighborhood. Computer Vision and Image Understanding: CVIU, 93(2): 195–205, February 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shih:2005:DBM

[SW05] Frank Y. Shih and Yi-Ta Wu. Decomposition of binary morphological structuring elements based on genetic algorithms. Computer Vision and Image Understanding: CVIU, 99(2):291–302, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Schnieders:2013:CLC

[SW13] Dirk Schnieders and Kwan-Yee K. Wong. Camera and light calibration from reflections on a sphere. Computer Vision and Image Understanding: CVIU, 117(10): 1536–1547, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001082.

Saha:2002:FDT

[SWG02] Punam K. Saha, Felix W. Wehrli, and Bryon R. Gomberg. Fuzzy distance transform: Theory, algorithms, and applications. Computer Vision and Image Understanding: CVIU, 86(3):171–190, June 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Scollar:1984:IEU

[SWH84] Irwin Scollar, Bernd Weidner, and Thomas S. Huang. Image enhancement using the median and the interquartile distance. Computer Vision, Graphics, and Image Processing, 25(2):236– 251, February 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Stolpner:2011:SML

[SWS11] Svetlana Stolpner, Sue Whitesides, and Kaleem Siddiqi. Sampled medial loci for 3D shape representation. Computer Vision and Image Understanding: CVIU, 115(5): 695–706, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Singh:2000:MMU

[SWYP00] Hanumant Singh, Louis Whitcomb, Dana Yoerger, and Oscar Pizarro. Microbathymetric mapping from underwater vehicles in the deep ocean. Computer Vision and Image Understanding: CVIU, 79(1):143–161, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0850;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0850/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0850/ref.

Shen:1998:DOM

[SY98] Jie Shen and Yee-Hong Yang.

Deformable object modeling using the time-dependent finite element method. Graphical Models and Image Processing: GMIP, 60(6):461– 487, November 1998. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0484/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0484/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0484/production/ref.

Sahillioglu:2010:CFS

[SY10] Y. Sahillioğlu and Y. Yemez.

Coarse-to-fine surface reconstruction from silhouettes and range data using mesh deformation. Computer Vision and Image Understanding: CVIU, 114(3):334–348, March 2010.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Srestasathiern:2011:PSR

[SY11] P. Srestasathiern and A. Yilmaz. Planar shape representation and matching under projective transformation. Computer Vision and Image Understanding: CVIU, 115(11):1525–1535, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421100169X.

Sakalli:1999:RBS

[SYF99] Mustafa Sakalli, Hong Yan, and Alan Fu. A regionbased scheme using RKLT and predictive classified vector quantization. Computer Vision and Image Understanding: CVIU, 75(3): 269–280, September 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0776/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0776/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0776/production/ref.

Shaked:1996:DSR

[SYK96] D. Shaked, O. Yaron, and N. Kiryati. Deriving stopping rules for the probabilistic Hough transform by sequential analysis. Computer Vision and Image Understanding: CVIU, 63(3):512–526, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0038/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0038/production/
pdf.

Srivastava:2013:UOG

[SYPK13] Gaurav Srivastava, Josiah A.

Yoder, Johnny Park, and Avinash C. Kak. Using objective ground-truth labels created by multiple annotators for improved video classification: a comparative study. Computer Vision and Image Understanding: CVIU, 117(10):1384–1399, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421300129X.

Sederberg:1996:PBS

[SZ96a] Thomas W. Sederberg and Alan K. Zundel. Pyramids that bound surface patches. Graphical Models and Image Processing: GMIP, 58(1):75–81, January 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.

                                        0005/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0005/production/
                                        pdf.

Shao:1996:CFB

[SZ96b] Lejun Shao and Hao Zhou. Curve fitting with Bézier cubics. Graphical Models and Image Processing: GMIP, 58(3):223–232, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0019/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0019/production/
                                        pdf.

Schaffalitzky:2003:ALM

[SZ03] F. Schaffalitzky and A. Zisserman. Automated location matching in movies. Computer Vision and Image Understanding: CVIU, 92(2–3):236–264, November/ December 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Shih:2007:LOC

[SZ07] Frank Y. Shih and Kai Zhang. Locating object contours in complex background using improved snakes. Computer Vision and Image Understanding: CVIU, 105(2):93– 98, February 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Szeliski:1991:FSS

[Sze91] R. Szeliski. Fast shape from shading. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (2):129–153, March 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Szeliski:1993:ROC

[Sze93] Richard Szeliski. Rapid octree construction from image sequences. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(1):23–32, July 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1029/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1029/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1030/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1030/production/
pdf.

Sederberg:1999:AIU

[SZKD99] Thomas W. Sederberg, Jianmin Zheng, Kris Klimaszewski, and Tor Dokken. Approximate implicitization using monoid curves and surfaces. Graphical Models and Image Processing:

GMIP, 61(4):177–198, July 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0497/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1999.0497/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0497/production/ref.

Tsukune:1988:AOP

[TA88] H. Tsukune and J. K. Aggarwal. Analyzing orthographic projection of multiple 3D velocity vector fields in optical flow. Computer Vision, Graphics, and Image Processing, 42(2):157–191, May 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tzevanidis:2011:ULB

[TA11] Konstantinos Tzevanidis and Antonis Argyros. Unsupervised learning of background modeling parameters in multicamera systems. Computer Vision and Image Understanding: CVIU, 115(1): 105–116, January 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tavakoli:2013:SSB

[TA13] Vahid Tavakoli and Amir A. Amini. A survey of shapedbased registration and segmentation techniques for cardiac images. Computer Vision and Image Understanding: CVIU, 117(9): 966–989, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000775.

Tajima:1983:UCS

[Taj83] Johji Tajima. Uniform color scale applications to computer graphics. Computer Vision, Graphics, and Image Processing, 21(3):305– 325, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tamaki:2009:RIR

[TAK09] Toru Tamaki, Toshiyuki Amano, and Kazufumi Kaneda. Representing images of a rotating object with cyclic permutation for view-based pose estimation. Computer Vision and Image Understanding: CVIU, 113(12):1210– 1221, December 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tamminen:1983:PAC

[Tam83] Markku Tamminen. Performance analysis of cell based geometric file organizations. Computer Vision, Graphics, and Image Processing, 21(3):160–181, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tamminen:1984:EPT

[Tam84]  Markku Tamminen.  En-

coding pixel trees. Computer Vision, Graphics, and Image Processing, 28(1):44– 57, October 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tanimoto:1976:PFD

[Tan76] Steven L. Tanimoto. Pictorial feature distortion in a pyramid. Computer Graphics and Image Processing, 5(3):333– 352, September 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Tanimoto:1979:ITG

[Tan79] Steven L. Tanimoto. Image transmission with gross information first. Computer Graphics and Image Processing, 9(1):72–76, January 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Tang:1981:MSI

[Tan81a] Gregory Y. Tang. Management system for an integrated database of pictures and alphanumerical data. Computer Graphics and Image Processing, 16(3):270–286, July 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Tanimoto:1981:TMP

[Tan81b] Steven L. Tanimoto. Template matching in pyramids. Computer Graphics and Image Processing, 16(4):356– 369, August 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Tanimoto:1986:CAA

[Tan86] S. Tanimoto. The case for appropriate architecture. Computer Vision, Graphics, and Image Processing, 34(1):116– ??, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tang:1989:ISD

[Tan89] Kai Tang. On the intersection of a set of direction cones. Computer Vision, Graphics, and Image Processing, 45(3):357–361, March 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tanaka:1995:ABS

[Tan95] Hiromi T. Tanaka. Accuracybased sampling and reconstruction with adaptive meshes for parallel hierarchical triangulation. Computer Vision and Image Understanding: CVIU, 61(3): 335–350, May 1995. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.

idealibrary.com/links/artid/
cviu.1995.1027/production;
 http://www.idealibrary.
com/links/artid/cviu.1995.
1027/production/pdf.

Tang:1999:CCC

[Tan99] Kai Tang. On computing contact configurations of a curved chain. Graphical Models and Image Processing: GMIP, 61 (6):341–361, November 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0507/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0507/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0507/production/ref.

Tang:2011:ARF

[Tan11] Min Tang. Automatic registration and fast volume reconstruction from serial histology sections. Computer Vision and Image Understanding: CVIU, 115(8):1112–1120, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000865.

Tasto:1977:RRO

[Tas77] M. Tasto. Reconstruction of random objects from noisy projections. Computer Graphics and Image Processing, 6 (2):103–122, April 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Taubin:2002:DMR

[Tau02a] Gabriel Taubin. Dual mesh resampling. Graphical Models, 64(2):94–113, March 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Taubin:2002:SIP

[Tau02b] Gabriel Taubin. Special issue on processing of large polygonal meshes. Graphical Models, 64(3–4):145–146, May 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Taylor:2000:RAO

[Tay00] Camillo J. Taylor. Reconstruction of articulated objects from point correspondences in a single uncalibrated image. Computer Vision and Image Understanding: CVIU, 80(3):349–363, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0878;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0878/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0878/ref.

Tarr:1994:CEP

[TB94a] Michael J. Tarr and Michael J. Black. A computational and evolutionary perspective on the role of representation in vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):65–73, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1031/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1031/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1036/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1036/production/

pdf. See also [CM94b].

Tarr:1994:RP

[TB94b] Michael J. Tarr and Michael J. Black. Reconstruction and purpose. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):113–118, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1041/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1041/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1046/production;   http:
//www.idealibrary.com/links/

artid/cviu.1994.1046/production/
pdf.

Tarel:1999:CFR

[TB99] Jean-Philippe Tarel and Nozha Boujemaa. A coarse to fine 3D registration method based on robust fuzzy clustering. Computer Vision and Image Understanding: CVIU, 73(1):14–28, January 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0673/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0673/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0673/production/ref.

Thurfjell:1995:BAF

[TBN95] Lennart Thurfjell, Ewert Bengtsson, and Bo Nordin. Boundary approach for fast neighborhood operations on three-dimensional binary data. Graphical Models and Image Processing: GMIP, 57 (1):13–19, January 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1002/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1002/production/
pdf.

Teh:1986:DAM

[TC86] Cho-Huak H. Teh and Roland T. Chin. On digital approximation of moment invariants. Computer Vision, Graphics, and Image Processing, 33(3):318–326, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tan:1987:BQB

[TC87] S. T. Tan and K. C. Chan. Bi-quadratic B-spline surfaces generated from arbitrary polyhedral meshes: a constructive approach. Computer Vision, Graphics, and Image Processing, 39(2):144– 166, August 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Thomas:1989:NSA

[TC89] Samuel M. Thomas and Y. T. Chan. Note: a simple approach for the estimation of circular arc center and its radius. Computer Vision, Graphics, and Image Processing, 45(3):362– 370, March 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Thomas:1995:CRL

[TC95] Samuel M. Thomas and Y. T. Chan. Cramer-Rao lower bounds for estimation of a circular arc center and its radius. Graphical Models and Image Processing: GMIP, 57 (6):527–532, November 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1043/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1043/production/
pdf.

Tsechpenakis:2011:DPM

[TC11] Gavriil Tsechpenakis and Sotirios P. Chatzis. Deformable probability maps: Probabilistic shape and appearancebased object segmentation. Computer Vision and Image Understanding: CVIU, 115 (8):1157–1169, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000853.

Tortora:1990:PAI

[TCAC90] Genoveffa Tortora, Gennaro Costagliola, Timothy Arndt, and Shi-Kuo Chang. Pyramidal algorithms for iconic indexing. Computer Vision, Graphics, and Image Processing, 52(1):26–56, October 1990. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Thacker:2008:PCC

[TCB^{+}08] Neil A. Thacker, Adrian F.

Clark, John L. Barron, J. Ross Beveridge, Patrick Courtney, William R. Crum, Visvanathan Ramesh, and Christine Clark. Performance characterization in computer vision: a guide to best practices. Computer Vision and Image Understanding: CVIU, 109(3):305–334, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Trivedi:1990:ODS

[TCC90] Mohan M. Trivedi, ChuXin Chen, and Daniel H. Cress. Object detection by stepwise analysis of spectral, spatial, and topographic features. Computer Vision, Graphics, and Image Processing, 51 (3):235–255, September 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tsui:1990:OES

[TCCK90] Hung-Tat Tsui, Ming-Hong Chan, Kin-Cheong Chu, and Shao-Hua Kong. Orientation estimation of 3D surface patches. Computer Vision, Graphics, and Image Processing, 50(1):112– 124, April 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Teng:2007:CTM

[TCH07] Chin-Hung Teng, Yung-Sheng Chen, and Wen-Hsing Hsu. Constructing a 3D trunk model from two images. Graphical Models, 69(1):33– 56, January 2007. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000427.

Theobalt:2004:CFF

[TCMS04] Christian Theobalt, Joel Carranza, Marcus A. Magnor, and Hans-Peter Seidel. Combining 3D flow fields with silhouette-based human motion capture for immersive video. Graphical Models, 66 (6):333–351, November 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Thi:2012:ILA

[TCZ^{+}12] Tuan Hue Thi, Li Cheng, Jian Zhang, Li Wang, and Shinichi Satoh. Integrating local action elements for action analysis. Computer Vision and Image Understanding: CVIU, 116(3):378–395, March 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002062.

Thrift:1983:APS

[TD83] Philip R. Thrift and Stanley M. Dunn. Approximating point-set images by line segments using a variation of the Hough transform. Computer Vision, Graphics, and Image Processing, 21(3):383–394, March 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tagare:1992:SES

[Td92] Hemant D. Tagare and Rui J. P. de Figueiredo. Simultaneous estimation of shape and reflectance map from photometric stereo. Computer Vision, Graphics, and Image Processing. Image Understanding, 55(3):275–286, May 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Tagare:1993:FCR

[Td93] Hemant D. Tagare and Rui J. P. de Figueiredo. A framework for the construction of reflectance maps for machine vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):265–282, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1019/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1019/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1019/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1019/production/
pdf.

Tan:2004:DVC

[TD04] Robin Tan and James W.

Davis. Differential video coding of face and gesture events in presentation videos. Computer Vision and Image Understanding: CVIU, 96(2): 200–215, November 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tek:2010:PDI

[TDK10] F. Boray Tek, Andrew G.

Dempster, and İzzet Kale. Parasite detection and identification for automated thin blood film malaria diagnosis. Computer Vision and Image Understanding: CVIU, 114 (1):21–32, January 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Thalmann:1985:LRD

[TDMT85] Daniel Thalmann, Louis-Philippe Demers, and Nadia Magnenat-Thalmann. Locating, replacing, and deleting patterns in graphics editing of line drawings. Computer Vision, Graphics, and Image Processing, 29(1):37– 46, January 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tran:2012:MPD

[TDT12] Cuong Tran, Anup Doshi, and Mohan Manubhai Trivedi.

Modeling and prediction of driver behavior by foot gesture analysis. Computer Vision and Image Understanding: CVIU, 116(3):435– 445, March 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002086.

Tisse:2007:ONS

[TDWH07] Christel-Loic Tisse, Hugh Durrant-Whyte, and R. Andrew Hicks. An optical navigation sensor for micro aerial vehicles. Computer Vision and Image Understanding: CVIU, 105(1):21–29, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Terzopoulos:1983:MCP

[Ter83] Demetri Terzopoulos. Multilevel computational processes for visual surface reconstruction. Computer Vision, Graphics, and Image Processing, 24(1):52–96, October 1983. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Thorstensen:2011:DMF

[TÉSK11] Nicolas Thorstensen, Patrick Étyngier, Florent Ségonne, and Renaud Keriven. Diffusion maps as a framework for shape modeling. Computer Vision and Image Understanding: CVIU, 115(4): 520–530, April 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tsao:1981:PTA

[TF81] Y. F. Tsao and K. S.

Fu. Parallel thinning algorithm for 3D pictures. Computer Graphics and Image Processing, 17(4):315– 331, December 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Tsad:1984:SSM

[TF84a] Y. Tsad and K. Fu. Stochastic skeleton modeling of objects. Computer Vision, Graphics, and Image Processing, 26(3):348–370, March 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tsao:1984:SSM

[TF84b] Yea-Fu F. Tsao and King-Sun S. Fu. Stochastic skeleton modeling of objects. Computer Vision, Graphics, and Image Processing, 25(3):348– 370, March 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tenenbaum:1980:SMS

[TFB80] Jay M. Tenenbaum, Martin A.

Fischler, and Harry G. Barrow. Scene modeling: a structural basis for image description. Computer Graphics and Image Processing, 12 (4):407–425, April 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Thomas:2009:SRR

[TFL^{+}09] Alexander Thomas, Vittorio Ferrari, Bastian Leibe, Tinne Tuytelaars, and Luc Van Gool. Shape-from-recognition: Recognition enables meta-data transfer. Computer Vision and Image Understanding: CVIU, 113(12):1222–1234, December 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tarbox:1995:IIV

[TG95a] G. H. Tarbox and S. N. Gottschlich. IVIS: An Integrated Volumetric Inspection System. Computer Vision and Image Understanding: CVIU, 61(3):430–444, May 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1032/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1032/production/
pdf.

Tarbox:1995:PCS

[TG95b] G. H. Tarbox and S. N. Gottschlich. Planning for complete sensor coverage in inspection. Computer Vision and Image Understanding: CVIU, 61(1):84–111, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1007/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1007/production/
pdf.

Thirion:1995:CDC

[TG95c] Jean-Philippe Thirion and Alexis Gourdon. Computing the differential characteristics of isointensity surfaces. Computer Vision and Image Understanding: CVIU, 61(2):190–202, March 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1015/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1015/production/
pdf.

Thirion:1996:MLA

[TG96] Jean-Philippe Thirion and Alexis Gourdon. The 3D marching lines algorithm. Graphical Models and Image Processing: GMIP, 58 (6):503–509, November 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.

                                        0042/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0042/production/
                                        pdf.

Taati:2011:LSD

[TG11] Babak Taati and Michael Greenspan. Local shape descriptor selection for object recognition in range data. Computer Vision and Image Understanding: CVIU, 115 (5):681–694, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tamstorf:2013:DBF

[TG13] Rasmus Tamstorf and Eitan Grinspun. Discrete bending forces and their Jacobians. Graphical Models, 75 (6):362–370, November 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000209.

Tolani:2000:RTI

[TGB00] Deepak Tolani, Ambarish Goswami, and Norman I. Badler. Real-time inverse kinematics techniques for anthropomorphic limbs. Graphical Models, 62(5):353–388, September 2000. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/gmod.
                                        2000.0528;   http://www.

idealibrary.com/links/doi/
10.1006/gmod.2000.0528/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0528/ref.

Tsap:1998:ENF

[TGSH98] Leonid V. Tsap, Dmitry B.

Goldgof, Sudeep Sarkar, and Wen-Chen Huang. Efficient nonlinear finite element modeling of nonrigid objects via optimization of mesh models. Computer Vision and Image Understanding: CVIU, 69(3):330–350, March 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0663/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0663/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0663/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0698/production;
 http://www.idealibrary.
com/links/artid/cviu.1998.
0698/production/pdf.

Tian:1986:ASR

[TH86] Qi Tian and Michael N.

Huhns. Algorithms for subpixel registration. Computer Vision, Graphics, and Image Processing, 35(2):220– 233, August 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Torsello:2004:SMS

[TH04] Andrea Torsello and Edwin R. Hancock. A skeletal measure of 2D shape similarity. Computer Vision and Image Understanding: CVIU, 95(1):1–29, July 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tang:2006:CSA

[TH06] W. K. Tang and Y. S. Hung. A column-space approach to projective reconstruction. Computer Vision and Image Understanding: CVIU, 101 (3):166–176, March 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tsuchie:2012:SMD

[TH12] Shoichi Tsuchie and Masatake Higashi. Surface mesh denoising with normal tensor framework. Graphical Models, 74(4):130–139, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000173.

Trivedi:1984:ODB

[THCG84] Mohan M. Trivedi, Charles A. Harlow, Richard W. Conners, and Semoon Goh. Object detection based on gray level co-occurrence. Computer Vision, Graphics, and Image Processing, 28(2):199– 219, November 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Therrien:1983:ETA

[The83] Charles W. Therrien. An estimation-theoretic approach to terrain image segmentation. Computer Vision, Graphics, and Image Processing, 22(3):313–326, June 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). THERRIEN83.

Thirion:1992:RSS

[Thi92] Jean-Phillipe Thirion. Realistic 3D simulation of shapes and shadows for image processing. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1):82– 90, January 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Tubic:2003:VAI

[THL03] Dragan Tubic, Patrick Hébert, and Denis Laurendeau. A volumetric approach for interactive 3D modeling. Computer Vision and Image Understanding: CVIU, 92(1): 56–77, October 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Thomee:2013:SIV

[THL13] Bart Thomee, Mark Huiskes, and Michael S. Lew. Special issue on visual concept detection in the MIR-FLICKR/ImageCLEF benchmark. Computer Vision and Image Understanding: CVIU, 117(5):451–452, May 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000350.

Thurfjell:1992:NTD

[THN92] Lennart Thurfjell, Ewert Hengtsson, and Bo Nordin. A new three-dimensional connected components labeling algorithm with simultaneous object feature extraction capability. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54 (4):357–364, July 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Thompson:1986:CEV

[Tho86] William B. Thompson. Comments on ““Expert” Vision Systems: Some Issues”. Computer Vision, Graphics, and Image Processing, 34(1):109– 110, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See [Ros86b, Kov86, Nag86, Uhr86].

Thomas:1994:RRT

[THO94] J. Inigo Thomas, Allen Hanson, and John Oliensis. Refining 3D reconstructions: a theoretical and experimental study of the effect of cross-correlations. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3): 359–370, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1062/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1062/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1068/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1068/production/
pdf.

Thonnat:2010:SII

[Tho10] Monique Thonnat. Special issue on intelligent vision systems. Computer Vision and Image Understanding: CVIU, 114(5):501–502, May 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tsai:1998:OSP

[THT^{+}98] Horng-Ren Tsai, Shi-Jinn Horng, Shun-Shan Tsai, Shung-Shing Lee, Tzong-Wann Kao, and Chia-Ho Chen. Optimal speed-up parallel image template matching algorithms on processor arrays with a reconfigurable bus system. Computer Vision and Image Understanding: CVIU, 71(3):393–412, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0638/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0638/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0638/production/ref.

Thurmer:2003:CCD

[Thü03] Grit Thürmer. Closed curves in n-dimensional discrete space. Graphical Models, 65 (1–3):43–60, May 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

terHaar:2009:FMF

[tHV09] Frank B. ter Haar and Remco C. Veltkamp. A 3D face matching framework for facial curves. Graphical Models, 71(2):77–91, March 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070308000325.

Tan:2001:DSR

[TI01] Joo Kooi Tan and Seiji Ishikawa. Deformable shape recovery by factorization based on a spatiotemporal measurement matrix. Computer Vision and Image Understanding: CVIU, 82(2):101–109, May 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0904;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0904/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0904/ref.

Taycher:2007:COF

[TID07] Leonid Taycher, John W. Fisher III, and Trevor Darrell. Combining object and feature dynamics in probabilistic tracking. Computer Vision and Image Understanding: CVIU, 108(3):243–260, December 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Takacs:2012:RPS

[TJ12] T. Takacs and B. Jüttler. H^{2} regularity properties of singular parameterizations in isogeometric analysis. Graphical Models, 74(6):361–372, November 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000380.

Tek:1997:VSM

[TK97] Hüseyin Tek and Benjamin B.

Kimia. Volumetric segmentation of medical images by three-dimensional bubbles. Computer Vision and Image Understanding: CVIU, 65(2):246–258, February 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0579/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0579/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0579/production/ref.

Truc:2009:VEF

[TKL^{+}09] Phan T. H. Truc, Md. A. U. Khan, Young-Koo Lee, Sungyoung Lee, and Tae-Seong Kim. Vessel enhancement filter using directional filter bank. Computer Vision and Image Understanding: CVIU, 113(1):101–112, January 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Thurey:2009:DPF

[TKPR09] N. Thürey, R. Keiser, M. Pauly, and U. Rüde. Detailpreserving fluid control. Graphical Models, 71(6):221–228, November 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000022.

Troxel:1978:ENP

[TL78] Donald E. Troxel and Charles Lynn, Jr. Enhancement of news photos. Computer Graphics and Image Processing, 7(2):266–281, April 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Tobler:1979:IIH

[TL79] W. Tobler and J. Lau. Interpolation of images via histosplines. Computer Graphics and Image Processing, 9 (1):77–81, January 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Tang:1988:RFU

[TL88] Gregory Y. Tang and Brian Lien. Region filling with the use of the discrete Green theorem. Computer Vision, Graphics, and Image Processing, 42(3):297–305, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tang:2005:OAF

[TL05] Kai Tang and Yong-Jin Liu.

An optimization algorithm for free-form surface partitioning based on weighted Gaussian image. Graphical Models, 67 (1):17–42, January 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Teng:2005:AOF

[TLCH05] Chin-Hung Teng, Shang-Hong Lai, Yung-Sheng Chen, and Wen-Hsing Hsu. Accurate optical flow computation under non-uniform brightness variations. Computer Vision and Image Understanding: CVIU, 97(3):315–346, March 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Theoharatos:2006:CSO

[TLEF06] Christos Theoharatos, Nikolaos Laskaris, George Economou, and Spiros Fotopoulos. Combining self-organizing neural nets with multivariate statistics for efficient color image retrieval. Computer Vision and Image Understanding: CVIU, 102(3):250–258, June 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tarini:2005:AMO

[TLGS05] Marco Tarini, Hendrik P. A. Lensch, Michael Goesele, and Hans-Peter Seidel. 3D acquisition of mirroring objects using striped patterns. Graphical Models, 67(4):233–259, July 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Tsotsos:2005:AVM

[TLMT^{+}05] John K. Tsotsos, Yueju Liu, Julio C. Martinez-Trujillo, Marc Pomplun, Evgueni Simine, and Kunhao Zhou. Attending to visual motion. Computer Vision and Image Understanding: CVIU, 100(1–2):3–40, October/November 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tan:1991:RRW

[TLT91a] Kah-Chye Tan, Hock Lim, and B. T. G. Tan. Restoration of real-world motionblurred images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(3): 291–299, May 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Tan:1991:WTI

[TLT91b] Kah-Chye Tan, Hock Lim, and B. T. G. Tan. Windowing techniques for image restoration. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):491– 500, September 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Tong:2012:SSF

[TLWT12] Yan Tong, Xiaoming Liu, Frederick W. Wheeler, and Peter H. Tu. Semi-supervised facial landmark annotation. Computer Vision and Image Understanding: CVIU, 116(8):922–935, August 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000616.

Tan:1986:DSA

[TM86] Chew L. Tan and W. N. Martin. A distributed system for analyzing timevarying multiresolution imagery. Computer Vision, Graphics, and Image Processing, 36(2/3):162–174, November/December 1986. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Theimer:1994:PBB

[TM94] Wolfgang M. Theimer and Hanspeter A. Mallot. Phasebased binocular vergence control and depth reconstruction using active vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3):343–358, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1061/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1061/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1067/production;   http:
//www.idealibrary.com/links/

artid/cviu.1994.1067/production/
pdf.

Tordoff:2004:IRD

[TM04] Ben Tordoff and David W.

Murray. The impact of radial distortion on the selfcalibration of rotating cameras. Computer Vision and Image Understanding: CVIU, 96(1):17–34, October 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Terra:2007:PBT

[TM07a] Sílvio César Lizana Terra and Ronald Anthony Metoyer. A performance-based technique for timing keyframe animations. Graphical Models, 69(2):89–105, March 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000646.

Tordoff:2007:MRZ

[TM07b] B. J. Tordoff and D. W.

Murray. A method of reactive zoom control from uncertainty in tracking. Computer Vision and Image Understanding: CVIU, 105(2): 131–144, February 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Torabi:2012:IIF

[TMB12] Atousa Torabi, Guillaume Massé, and Guillaume-Alexandre Bilodeau. An iterative integrated framework for thermalvisible image registration, sensor fusion, and people tracking for video surveillance applications. Computer Vision and Image Understanding: CVIU, 116 (2):210–221, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002335.

Tillett:2000:EDF

[TML00] Robin Tillett, Nigel McFarlane, and Jeff Lines. Estimating dimensions of freeswimming fish using 3D point distribution models. Computer Vision and Image Understanding: CVIU, 79(1):123–141, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0847;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0847/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0847/ref.

Tsechpenakis:2006:LBD

[TMN06] Gabriel Tsechpenakis, Dimitris Metaxas, and Carol Neidle. Learning-based dynamic coupling of discrete and continuous trackers. Computer Vision and Image Understanding: CVIU, 104 (2–3):140–156, November/ December 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Takai:2009:DSA

[TMNM09] Takeshi Takai, Atsuto Maki, Koichiro Niinuma, and Takashi Matsuyama. Difference sphere: an approach to near light source estimation. Computer Vision and Image Understanding: CVIU, 113(9): 966–978, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tao:2013:RNR

[TMQM13] Lili Tao, Stephen J. Mein, Wei Quan, and Bogdan J. Matuszewski. Recursive nonrigid structure from motion with online learned shape prior. Computer Vision and Image Understanding: CVIU, 117(10):1287–1298, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000726.

Tang:2010:MMC

[TMT10] Min Tang, Dinesh Manocha, and Ruofeng Tong. MCCD: Multi-core collision detection between deformable models using front-based decomposition. Graphical Models, 72(2):7–23, March 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000020.

Thurley:2005:IVC

[TN05] Matthew J. Thurley and Kim C. Ng. Identifying, visualizing, and comparing regions in irregularly spaced 3D surface data. Computer Vision and Image Understanding: CVIU, 98(2):239–270, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Todorovic:2007:ICS

[TN07] Sinisa Todorovic and Michael C. Nechyba. Interpretation of complex scenes using dynamic tree-structure Bayesian networks. Computer Vision and Image Understanding: CVIU, 106(1):71–84, April 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Thurley:2008:ISE

[TN08] Matthew J. Thurley and Kim C. Ng. Identification and sizing of the entirely visible rocks from a 3D surface data segmentation of laboratory rock piles. Computer Vision and Image Understanding: CVIU, 111(2): 170–178, August 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Thomas:1999:DNM

[TO99] J. Inigo Thomas and John Oliensis. Dealing with noise in multiframe structure from motion. Computer Vision and Image Understanding: CVIU, 76(2):109–124, November 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0779/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0779/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0779/production/ref.

Tou:1980:PFE

[Tou80] Julius T. Tou. Pictorial feature extraction and recognition via image modeling. Computer Graphics and Image Processing, 12(4):376– 406, April 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Tanimoto:1975:HDS

[TP75] S. L. Tanimoto and Theo Pavlidis. A hierarchical data structure for picture processing. Computer Graphics and Image Processing, 4(2): 104–119, June 1975. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Thompson:1992:QCS

[TP92] William B. Thompson and James S. Painter. Qualitative constraints for structurefrom-motion. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):69–77, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Traver:2005:SME

[TP05] V. Javier Traver and Filiberto Pla. Similarity motion estimation and active tracking through spatial-domain projections on log-polar images. Computer Vision and Image Understanding: CVIU, 97(2): 209–241, February 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Trucco:2000:FTV

[TPR^{+}00] E. Trucco, Y. R. Petillot, I. Tena Ruiz, K. Plakas, and D. M. Lane. Feature tracking in video and sonar subsea sequences with applications. Computer Vision and Image Understanding: CVIU, 79(1):92–122, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0846;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0846/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/cviu.
2000.0846/ref.

Toffoli:1997:TDR

[TQ97] Tommaso Toffoli and Jason Quick. Three-dimensional rotations by three shears. Graphical Models and Image Processing: GMIP, 59 (2):89–95, March 1997. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0420/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0420/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0420/production/ref.

Tresadern:2009:VSH

[TR09] Philip A. Tresadern and Ian D. Reid. Video synchronization from human motion using rank constraints. Computer Vision and Image Understanding: CVIU, 113(8): 891–906, August 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Treisman:1985:PPV

[Tre85] Anne Treisman. Preattentive processing in vision. Computer Vision, Graphics, and Image Processing, 31 (2):156–177, August 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Toth:2013:SSP

[TRG^{+}13] Robert Toth, Justin Ribault, John Gentile, Dan Sperling, and Anant Madabhushi. Simultaneous segmentation of prostatic zones using Active Appearance Models with multiple coupled levelsets. Computer Vision and Image Understanding: CVIU, 117(9): 1051–1060, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000702.

Trivedi:1990:SAM

[Tri90] Harit P. Trivedi. A semianalytic method of determining stereo camera geometry from matched points in a pair of images: coincident meridional planes, exact or noisy data. Computer Vision, Graphics, and Image Processing, 51(3):299– 312, September 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tobor:2006:RMS

[TRS06] Ireneusz Tobor, Patrick Reuter, and Christophe Schlick. Reconstructing multi-scale variational partition of unity implicit surfaces with attributes. Graphical Models, 68(1):25–41, January 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000573.

Topa:1986:AAD

[TS86] Leonardo C. Topa and Robert J.

Schalkoff. An analytical approach to the determination of planar surface orientation using active-passive image pairs. Computer Vision, Graphics, and Image Processing, 35 (3):404–418, September 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tistarelli:1992:DAA

[TS92] Massimo Tistarelli and Giulio Sandini. Dynamic aspects in active vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(1):108–129, July 1992. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Tang:2000:OSI

[TS00a] Xiaoou Tang and W. Kenneth Stewart. Optical and sonar image classification: Wavelet packet transform vs Fourier transform. Computer Vision and Image Understanding: CVIU, 79(1):25–46, July 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0843;   http://www.
idealibrary.com/links/doi/

10.1006/cviu.2000.0843/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0843/ref.

Tari:2000:NLS

[TS00b] Sibel Tari and Jayant Shah. Nested local symmetry set. Computer Vision and Image Understanding: CVIU, 79(2):267–280, August 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0856;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0856/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0856/ref.

Tissainayagam:2001:PPA

[TS01] P. Tissainayagam and D. Suter. Performance prediction analysis of a point feature tracker based on different motion models. Computer Vision and Image Understanding: CVIU, 84(1):104–125, October 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0939;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0939/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0939/ref.

Thomas:2011:RRR

[TS11] Diego Thomas and Akihiro Sugimoto. Robustly registering range images using local distribution of albedo. Computer Vision and Image Understanding: CVIU, 115(5): 649–667, May 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tsai:1985:MPT

[Tsa85] Wen-Hsiang H. Tsai. Momentpreserving thresholding: a new approach. Computer Vision, Graphics, and Image Processing, 29(3):377– 393, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tsai:1996:PAG

[Tsa96] Frank C. D. Tsai. A probabilistic approach to geometric hashing using line features. Computer Vision and Image Understanding: CVIU, 63(1):182–195, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0013/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0013/production/
pdf.

Topkar:1994:PAD

[TSK94] V. Topkar, A. K. Sood, and B. Kjell. Performance analysis of 1-D scalespace algorithms for pulse detection in noisy image scans. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):191–209, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1047/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1047/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1052/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1052/production/
pdf.

Tsotsos:1994:TNO

[Tso94] John K. Tsotsos. There is no one way to look at vision. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(1):95–97, July 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1036/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1036/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1041/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1041/production/
pdf.

Tari:1997:ESS

[TSP97] Z. Sibel Göktepe Tari, Jayant Shah, and Homer Pien. Extraction of shape skeletons from grayscale images. Computer Vision and Image Understanding: CVIU, 66(2):133–146, May 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0612/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0612/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0612/production/ref.

Taylor:1989:FSR

[TSR89] Russell W. Taylor, Massimo Savini, and Anthony P. Reeves. Fast segmentation of range imagery into planar regions. Computer Vision, Graphics, and Image Processing, 45(1):42–60, January 1989. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tamura:1983:DIS

[TST^{+}83] Hideyuki Tamura, Shigeyuki Sakane, Fumiaki Tomita, Naokazu Yokoya, Masahide Kaneko, and Katsuhiko Sakaue. Design and implementation of Spider — a transportable image processing software package. Computer Vision, Graphics, and Image Processing, 23 (3):273–294, September 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Tekalp:1991:CSS

[TT91] A. M. Tekalp and H. J. Trussell. Comparative study of some statistical and settheoretic methods for image restoration. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(2): 108–120, March 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Tarabanis:1994:ACF

[TTA94] Konstantinos Tarabanis, Roger Y. Tsai, and Peter K. Allen. Analytical characterization of the feature detectability constraints of resolution, focus, and field-of-view for vision sensor planning. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(3):340–358, May 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1024/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1024/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1028/production;   http:
//www.idealibrary.com/links/

artid/cviu.1994.1028/production/
pdf.

Takahashi:2004:TVS

[TTF04] Shigeo Takahashi, Yuriko Takeshima, and Issei Fujishiro. Topological volume skeletonization and its application to transfer function design. Graphical Models, 66 (1):24–49, January 2004. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Tarabanis:1994:CCC

[TTG94] Konstantinos Tarabanis, Roger Y.

Tsai, and Douglas S. Goodman. Calibration of a computer controlled robotic vision sensor with a zoom lens. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(2):226–241, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1015/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1015/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1017/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1017/production/
pdf.

Tu:2007:FMT

[TTH07] Jilin Tu, Hai Tao, and Thomas Huang. Face as mouse through visual face tracking. Computer Vision and Image Understanding: CVIU, 108(1–2):35–40, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Toklu:1996:TMI

[TTIM96] Candemir Toklu, A. Tanju Erdem, M. Ibrahim Sezan, and A. Murat Tekalp. Tracking motion and intensity variations using hierarchical 2-D mesh modeling for synthetic object transfiguration. Graphical Models and Image Processing: GMIP, 58 (6):553–573, November 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0046/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0046/production/
pdf.

Takala:1999:DTA

[TV99] Jarmo H. Takala and Jouko O. Viitanen. Distance transform algorithm for bit-serial SIMD architectures. Computer Vision and Image Understanding: CVIU, 74(2):150–161, May 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0756/production;   http:

//www.idealibrary.com/links/
artid/cviu.1999.0756/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0756/production/ref;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0757/production;
http://www.idealibrary.
com/links/artid/cviu.1999.
0757/production/pdf.

Turaga:2009:UVR

[TVC09] Pavan Turaga, Ashok Veeraraghavan, and Rama Chellappa. Unsupervised view and rate invariant clustering of video sequences. Computer Vision and Image Understanding: CVIU, 113(3): 353–371, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tahoces:2008:ICM

[TVLS08] Pablo G. Tahoces, J. Ramón Varela, María J. Lado, and Miguel Souto. Image compression: Maxshift ROI encoding options in JPEG2000. Computer Vision and Image Understanding: CVIU, 109(2): 139–145, February 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Takacs:1998:DMM

[TW98] Barnabás Takács and Harry Wechsler. A dynamic and multiresolution model of visual attention and its application to facial landmark detection. Computer Vision and Image Understanding: CVIU, 70(1):63–73, April 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0619/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0619/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0619/production/ref.

Tabbone:2006:NSD

[TWS06] S. Tabbone, L. Wendling, and J.-P. Salmon. A new shape descriptor defined on the Radon transform. Computer Vision and Image Understanding: CVIU, 102(1): 42–51, April 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Tankus:2001:CBV

[TY01] Ariel Tankus and Yehezkel Yeshurun. Convexity-based visual camouflage breaking. Computer Vision and Image Understanding: CVIU, 82(3):208–237, June 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0912;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0912/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/cviu.
2001.0912/ref.

Tankus:2005:SCD

[TY05] Ariel Tankus and Yehezkel Yeshurun. Scene-consistent detection of feature points in video sequences. Computer Vision and Image Understanding: CVIU, 97(1): 1–29, January 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Terzopoulos:1982:DOI

[TZ82] Demetri Terzopoulos and Steven W. Zucker. Detection of osteogenesis imperfecta by automated texture analysis. Computer Graphics and Image Processing, 20(3): 229–243, November 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Torr:2000:MNR

[TZ00] P. H. S. Torr and A. Zisserman. MLESAC: a new robust estimator with application to estimating image geometry. Computer Vision and Image Understanding: CVIU, 78(1):138–156, April 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
1999.0832;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.1999.0832/

                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        1999.0832/ref.

Torr:1998:RDD

[TZM98] P. H. S. Torr, A. Zisserman, and S. J. Maybank. Robust detection of degenerate configurations while estimating the fundamental matrix. Computer Vision and Image Understanding: CVIU, 71(3):312–333, September 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0559/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0559/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0559/production/ref.

Tu:2008:SMR

[TZY08] Zhuowen Tu, Songfeng Zheng, and Alan Yuille. Shape matching and registration by data-driven EM. Computer Vision and Image Understanding: CVIU, 109(3): 290–304, March 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Underwood:1977:ICA

[UA77] S. A. Underwood and J. K. Aggarwal. Interactive computer analysis of aerial color infrared photographs. Computer Graphics and Image Processing, 6(1):1–24, February 1977. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Udupa:1990:BOL

[UA90] Jayaram K. Udupa and Venkatramana G. Ajjanagadde. Boundary and object labelling in three-dimensional images. Computer Vision, Graphics, and Image Processing, 51(3):355–369, September 1990. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Unsalan:2005:SDH

[ÜB05] Cem Ünsalan and Kim L.

Boyer. A system to detect houses and residential street networks in multispectral satellite images. Computer Vision and Image Understanding: CVIU, 98(3): 423–461, June 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Uz:2009:MBT

[UBEP09] Tamer Uz, George Bebis, Ali Erol, and Salil Prabhakar. Minutiae-based template synthesis and matching for fingerprint authentication. Computer Vision and Image Understanding: CVIU, 113 (9):979–992, September 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Untereiner:2013:DMR

[UCB13] Lionel Untereiner, David Cazier, and Dominique Bechmann. n-dimensional multiresolution representation of subdivision meshes with arbitrary topology. Graphical Models, 75(5):231–246, September 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000131.

Udupa:1981:DSP

[Udu81] Jayaram K. Udupa. Determination of 3D shape parameters from boundary information. Computer Graphics and Image Processing, 17(1):52– 59, September 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Udupa:1982:ISB

[Udu82]  Jayaram K. Udupa.  In-

teractive segmentation and boundary surface formation for 3D digital images. Computer Graphics and Image Processing, 18(3):213– 235, March 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Udupa:1994:MDB

[Udu94]  J. K. Udupa.  Multidi-

mensional digital boundaries. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(4):311–323, July 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1028/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1028/production/
pdf.

Urtasun:2006:TMM

[UFF06] Raquel Urtasun, David J.

Fleet, and Pascal Fua. Temporal motion models for monocular and multiview 3D human body tracking. Computer Vision and Image Understanding: CVIU, 104(2–3):157–177, November/December 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ur:1992:IRS

[UG92] Hanoch Ur and Daniel Gross.

Improved resolution from subpixel shifted pictures. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(2):181–186, March 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). See note [Ano92b].

Uhr:1986:WGD

[Uhr86]  Leonard Uhr.  Workshop

on goal-directed “expert” vision systems: My positions and comments. Computer Vision, Graphics, and Image Processing, 34(1):105– 109, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See [Ros86b, Kov86, Nag86, Tho86].

Ukita:2012:GPM

[UK12a] Norimichi Ukita and Takeo Kanade. Gaussian process motion graph models for smooth transitions among multiple actions. Computer Vision and Image Understanding: CVIU, 116(4): 500–509, April 2012. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211002608.

Ukita:2012:RCR

[UK12b] Norimichi Ukita and Takeo Kanade. Reference consistent reconstruction of 3D cloth surface. Computer Vision and Image Understanding: CVIU, 116(8):869–881, August 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212000628.

Umeyama:1988:RPT

[UKH88] Shinji Umeyama, T. Kasvand, and M. Hospital. Recognition and positioning of threedimensional objects by combining matchings of primitive local patterns. Computer Vision, Graphics, and Image Processing, 44(1):58– 76, October 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Unsalan:2001:CBP

[Ül01] Cem Ünsalan and Aytül Erç l. Conversions between parametric and implicit forms using polar/spherical coordinate representations. Computer Vision and Image Understanding: CVIU, 81(1): 1–25, January 2001. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0881;
 http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0881/pdf;   http://
www.idealibrary.com/links/
doi/10.1006/cviu.2000.0881/
ref.

Ullman:1979:RCO

[Ull79] S. Ullman. Relaxation and constrained optimization by local processes. Computer Graphics and Image Processing, 10(2):115–125, June 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ullman:1981:IOD

[Ull81] Shimon Ullman. Interfacing the one-dimensional scanning of an image with the application of two-dimensional operators. Computer Graphics and Image Processing, 16(2): 150–157, June 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Ullmann:1983:IOO

[Ull83] J. R. Ullmann. Investigation of occlusion in one dimension. Computer Vision, Graphics, and Image Processing, 22(1):194–203, April 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ulupinar:1990:RED

[UM90] Faith Ulupinar and Gerard Medioni. Refining edges detected by a LoG operator. Computer Vision, Graphics, and Image Processing, 51 (3):275–298, September 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Ukita:2005:RTC

[UM05] Norimichi Ukita and Takashi Matsuyama. Real-time cooperative multi-target tracking by communicating active vision agents. Computer Vision and Image Understanding: CVIU, 97(2): 137–179, February 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Ulupinar:1991:CIL

[UN91] F. Ulupinar and R. Nevatia.

Constraints for interpretation of line drawings under perspective projection. Computer Vision, Graphics, and Image Processing. Image Understanding, 53(1):88–96 (or 674– 676??), January 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Unzueta:2008:FBP

[UPBS08] Luis Unzueta, Manuel Peinado, Ronan Boulic, and Ángel Suescun. Full-body performance animation with Sequential Inverse Kinematics. Graphical Models, 70 (5):87–104, September 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000040.

Udupa:1996:FCO

[US96] Jayaram K. Udupa and Supun Samarasekera. Fuzzy connectedness and object definition: Theory, algorithms, and applications in image segmentation. Graphical Models and Image Processing: GMIP, 58(3):246–261, May 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0021/production;   http:
//www.idealibrary.com/links/

                                        artid/gmip.1996.0021/production/
                                        pdf.

Ulges:2010:LAC

[USKB10] Adrian Ulges, Christian Schulze, Markus Koch, and Thomas M. Breuel. Learning automatic concept detectors from online video. Computer Vision and Image Understanding: CVIU, 114(4): 429–438, April 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Urfalioglu:2011:AEA

[UTB^{+}11] Onay Urfalioglu, Thorsten Thormählen, Hellward Broszio, Patrick Mikulastik, and A. Enis Cetin. Algebraic error analysis of collinear feature points for camera parameter estimation. Computer Vision and Image Understanding: CVIU, 115(4):467–475, April 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Udomkesmalee:1997:IPP

[UZC97] Suraphol Udomkesmalee, David Q. Zhu, and Cheng-Chih Chu. Image processing for planetary limb/terminator extraction. Computer Vision and Image Understanding: CVIU, 67(3):274–284, September 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.

0543/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0543/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0543/production/ref.

VanderBrug:1977:EIE

[Van77] Gordon J. VanderBrug. Experiments in iterative enhancement of linear features. Computer Graphics and Image Processing, 6(1): 25–42, February 1977. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

VanWyk:1984:CBC

[Van84] Christopher J. Van Wyk.

Clipping to the boundary of a circular-arc polygon. Computer Vision, Graphics, and Image Processing, 25(3):383–392, March 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

vanLierop:1986:GTP

[van86] Marloes L. P. van Lierop.

Geometrical transformations on pictures represented by leafcodes. Computer Vision, Graphics, and Image Processing, 33(1):81–98, January 1986. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Vasa:2011:OMT

[Vás11] Libor Vása. Optimised mesh traversal for dynamic mesh compression. Graphical Models, 73(5):218–230, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000129.

Valentine:2010:ECC

[VAWW10] Brian Valentine, Senyo Apewokin, Linda Wills, and Scott Wills. An efficient, chromatic clustering-based background model for embedded vision platforms. Computer Vision and Image Understanding: CVIU, 114(11): 1152–1163, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Vaidya:1998:DPS

[VB98] Nitin M. Vaidya and Kim L. Boyer. Discontinuity-preserving surface reconstruction using stochastic differential equations. Computer Vision and Image Understanding: CVIU, 72(3):257–270, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0700/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0700/production/
pdf;   http://www.idealibrary.

com/links/artid/cviu.1998.
0700/production/ref.

Voigtmann:1997:HMM

[VBH97] Andreas Voigtmann, Ludger Becker, and Klaus Hinrichs. A hierarchical model for multiresolution surface reconstruction. Graphical Models and Image Processing: GMIP, 59(5):333–348, September 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0436/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0436/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0436/production/ref.

Varga:2011:DDB

[VBN11] László Varga, Péter Balázs, and Antal Nagy. Directiondependency of binary tomographic reconstruction algorithms. Graphical Models, 73 (6):365–375, November 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000245.

Vanrell:2004:IOC

[VBS^{+}04] M. Vanrell, R. Baldrich, A. Salvatella, R. Benavente, and F. Tous. Induction operators for a computational colour–texture representation. Computer Vision and Image Understanding: CVIU, 94(1–3):92–114, April/ June 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Vasudevan:1988:HIL

[VCBC88] Sridhar Vasudevan, Robert L. Cannon, James C. Bezdek, and William L. Cameron. Heuristics for intermediate level road finding algorithms. Computer Vision, Graphics, and Image Processing, 44 (2):175–190, November 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Vacavant:2009:FDI

[VCT09] Antoine Vacavant, David Coeurjolly, and Laure Tougne. A framework for dynamic implicit curve approximation by an irregular discrete approach. Graphical Models, 71(3):113–124, May 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070309000149.

Vemuri:1998:EAC

[VCVQ^{+}98] B. C. Vemuri, L. Chen, L. Vu-Quoc, X. Zhang, and O. Walton. Efficient and accurate collision detection for granular flow simulation. Graphical Models and Image Processing: GMIP, 60 (6):403–422, November 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0479/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0479/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0479/production/ref.

Veatch:1990:EAO

[VD90] Phillip A. Veatch and Larry S.

Davis. Efficient algorithms for obstacle detection using range data. Computer Vision, Graphics, and Image Processing, 50(1):50–74, April 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Venetianer:2010:PEI

[VD10] Péter L. Venetianer and Hongli Deng. Performance evaluation of an intelligent video surveillance system – A case study. Computer Vision and Image Understanding: CVIU, 114(11):1292– 1302, November 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

VanGool:1985:STA

[VDO85] L. Van Gool, P. Dewaele, and A. Oosterlinck. Survey: Texture analysis anno 1983. Computer Vision, Graphics, and Image Processing, 29(3):336–357, March 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

vandeWetering:1996:CCT

[vdWvO96] Huub van de Wetering and Kees van Overveld. Chain codes and their application in curve design. Graphical Models and Image Processing: GMIP, 58(5):464– 470, September 1996. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1996.
                                        0038/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1996.0038/production/
                                        pdf.

Veelaert:1997:CFE

[Vee97] Peter Veelaert. Constructive fitting and extraction of geometric primitives. Graphical Models and Image Processing: GMIP, 59(4):233–251, July 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0433/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1997.0433/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1997.
                                        0433/production/ref.

Veltkamp:1995:BTS

[Vel95] Remco C. Veltkamp. Boundaries through scattered points of unknown density. Graphical Models and Image Processing: GMIP, 57(6):441– 452, November 1995. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1038/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1038/production/
pdf.

Verly:1981:HRI

[Ver81] Jacques G. Verly. Highresolution imaging in 3D reconstructive tomography. Computer Graphics and Image Processing, 16(1):1–19, May 1981. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Verri:1997:TMV

[Ver97] Alessandro Verri. There is more to vision than geometry

— reply to Pizlo, Rosenfeld, and Weiss. Computer Vision and Image Understanding: CVIU, 65(3):439–441, March 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0495/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0495/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0495/production/ref.

Vieville:1996:FOE

[VF96] T. Viéville and O. D. Faugeras. The first order expansion of motion equations in the uncalibrated case. Computer Vision and Image Understanding: CVIU, 64(1):128–146, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0049/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0049/production/
                                        pdf.

VandePanne:1993:PBM

[VFV93] Michiel Van de Panne, Eugene Fiume, and Zvonko Vranesic. Physically based modeling and control of turning. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(6):507–521, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1993.
                                        1038/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1993.1038/production/
                                        pdf.

vanGemert:2010:CCC

[vGSV^{+}10] Jan C. van Gemert, Cees G. M. Snoek, Cor J. Veenman, Arnold W. M. Smeulders, and Jan-Mark Geusebroek. Comparing compact codebooks for visual categorization. Computer Vision and Image Understanding: CVIU, 114(4): 450–462, April 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Vayda:1991:RVS

[VK91] A. J. Vayda and A. C. Kak.

A robot vision system for recognition of generic shaped objects. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (1):1–46, July 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Venkatesh:1992:EEU

[VK92] Svetha Venkatesh and Leslie John Kitchen. Edge evaluation using necessary components. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(1):23–30, January 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Vijayakumar:1998:IBR

[VKP98] B. Vijayakumar, David Kriegman, and Jean Ponce. Invariantbased recognition of complex curved 3D objects from image contours. Computer Vision and Image Understanding: CVIU, 72(3):287–303, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0701/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0701/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0701/production/ref.

Venot:1984:NCS

[VLR84] A. Venot, J. F. Lebruchec, and J. C. Roucayrol. A new class of similarity measures for robust image registration. Computer Vision, Graphics, and Image Processing, 28 (2):176–184, November 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Vogler:2001:FRS

[VM01] Christian Vogler and Dimitris Metaxas. A framework for recognizing the simultaneous aspects of American Sign Language. Computer Vision and Image Understanding: CVIU, 81(3):358–384, March 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0895;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2000.0895/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2000.0895/ref.

Varadhan:2006:AMS

[VM06] Gokul Varadhan and Dinesh Manocha. Accurate Minkowski sum approximation of polyhedral models. Graphical Models, 68 (4):343–355, July 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000191.

Vandenbroucke:2003:CIS

[VMP03] Nicolas Vandenbroucke, Ludovic Macaire, and Jack-Gérard Postaire. Color image segmentation by pixel classification in an adapted hybrid color space. application to soccer image analysis. Computer Vision and Image Understanding: CVIU, 90 (2):190–216, May 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

VanGool:1995:CDS

[VMUO95] Luc Van Gool, Theo Moons, Dorin Ungureanu, and Andre Oosterlinck. The characterization and detection of skewed symmetry. Computer Vision and Image Understanding: CVIU, 61(1):138–150, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1010/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1010/production/
pdf.

Vossepoel:1988:NND

[Vos88] Albert M. Vossepoel. Note: a note on “Distance Transformations in Digital Images”. Computer Vision, Graphics, and Image Processing, 43(1):88–97, July 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic). See [Bor86, BS89].

Vigo:2012:EAB

[VPAM12] Marc Vigo, Núria Pla, Dolors Ayala, and Jonàs Martínez. Efficient algorithms for boundary extraction of 2D and 3D orthogonal pseudomanifolds. Graphical Models, 74(3):61–74, May 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000112.

Venkatesh:1995:DTD

[VR95] Svetha Venkatesh and Paul L. Rosin. Dynamic threshold determination by local and global edge evaluation. Graphical Models and Image Processing: GMIP, 57(2):146–160, March 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1995.
                                        1015/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1995.1015/production/
                                        pdf.

Vacavant:2013:CMS

[VRKL13] Antoine Vacavant, Tristan Roussillon, Bertrand Kerautret, and Jacques-Olivier Lachaud. A combined multiscale/irregular algorithm for the vectorization of noisy digital contours. Computer Vision and Image Understanding: CVIU, 117(4):438–450, April 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001841.

Vossepoel:1982:VCP

[VS82] A. M. Vossepoel and A. W. M.

Smeulders. Vector code probability and metrication error in the representation of straight lines of finite length. Computer Graphics and Image Processing, 20(4):347– 364, December 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Vanderhyde:2008:TSI

[VS08] James Vanderhyde and Andrzej Szymczak. Topological simplification of isosurfaces in volumetric data using octrees. Graphical Models, 70 (1–2):16–31, January/March 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070307000161.

Veeraraghavan:2006:RTD

[VSP06] Harini Veeraraghavan, Paul Schrater, and Nikos Papanikolopoulos. Robust target detection and tracking through integration of motion, color, and geometry. Computer Vision and Image Understanding: CVIU, 103(2): 121–138, August 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Varady:2012:TSI

[VSR12] Tamás Várady, Péter Salvi, and Alyn Rockwood. Transfinite surface interpolation with interior control. Graphical Models, 74(6):311–320, November 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000100.

Velho:1995:CIS

[VTG95] Luiz Velho, Demetri Terzopoulos, and Jonas Gomes. Constructing implicit shape models from boundary data. Graphical Models and Image Processing: GMIP, 57 (3):220–234, May 1995. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1021/production;   http:
//www.idealibrary.com/links/

artid/gmip.1995.1021/production/
pdf.

vandenBoomgaard:1992:MFM

[vv92a] Rein van den Boomgaard and Richard van Balen. Methods for fast morphological image transforms using bitmapped binary images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(3): 252–258, May 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Villasenor:1992:ASR

[VV92b] John Villasenor and Alain Vincent. An algorithm for space recognition and time tracking of vorticity tubes in turbulence. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (1):27–35, January 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

VandenWyngaerd:2002:ACP

[VV02] Joris Vanden Wyngaerd and Luc Van Gool. Automatic crude patch registration: Toward automatic 3D model building. Computer Vision and Image Understanding: CVIU, 87(1–3):8–26, July 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

vanLierop:1988:LRA

[vvv88] M. L. P. van Lierop, C. W. A. M. van Overveld, and H. M. M. van de Wetering. Line rasterization algorithms that satisfy the subset line property. Computer Vision, Graphics, and Image Processing, 41(2):210– 228, February 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Vaishnavi:1980:DSR

[VW80] Vijay Vaishnavi and Derick Wood. Data structures for the rectangle containment and enclosure problems. Computer Graphics and Image Processing, 13(4): 372–384, August 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Vega:1994:DST

[VY94] Omar E. Vega and Yee-Hong H. Yang. Default shape theory: With application to the computation of the direction of the light source. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3):285–299, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1058/production;   http:
//www.idealibrary.com/links/

artid/ciun.1994.1058/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1064/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1064/production/
pdf.

vanVliet:1989:NLO

[vYB89] Lucas J. van Vliet, Ian T.

Young, and Guus L. Beckers. A nonlinear Laplace operator as edge detector in noisy images. Computer Vision, Graphics, and Image Processing, 45(2):167– 195, February 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Vincze:2009:IVS

[VZP^{+}09] Markus Vincze, Michael Zillich, Wolfgang Ponweiser, Vaclav Hlavac, Jiri Matas, Stepan Obdrzalek, Hilary Buxton, Jonathan Howell, Kingsley Sage, Antonis Argyros, Christoph Eberst, and Gerald Umgeher. Integrated vision system for the semantic interpretation of activities where a person handles objects. Computer Vision and Image Understanding: CVIU, 113(6):682–692, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Weng:1987:OOA

[WA87] Juyang Weng and Narendra Ahuja. Octrees of objects in arbitrary motion: Representation and efficiency. Computer Vision, Graphics, and Image Processing, 39 (2):167–185, August 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wagner:1976:ROL

[Wag76] W. Wagner. Reconstruction of object layers from their Xray projections: a simulation study. Computer Graphics and Image Processing, 5(4): 470–483, December 1976. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wahl:1983:NDM

[Wah83] Friedrich M. Wahl. New distance mapping and its use for shape measurement on binary patterns. Computer Vision, Graphics, and Image Processing, 23(2):218– 226, August 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Walters:1987:SIP

[Wal87] Deborah Walters. Selection of image primitives for generalpurpose visual processing. Computer Vision, Graphics, and Image Processing, 37 (2):261–298, February 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Walsh:1988:EAT

[Wal88] T. R. Walsh. Efficient axistranslation of binary digital pictures by blocks in linear quadtree representation. Computer Vision, Graphics, and Image Processing, 41(3):282–292, March 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Walton:1989:NGE

[Wal89] D. J. Walton. A note on graphics editing of curved line drawings. Computer Vision, Graphics, and Image Processing, 45(1):61–67, January 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wildes:2000:REF

[WALL00] Richard P. Wildes, Michael J.

Amabile, Ann-Marie Lanzillotto, and Tzong-Shyng Leu. Recovering estimates of fluid flow from image sequence data. Computer Vision and Image Understanding: CVIU, 80(2):246–266, November 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0874;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0874/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0874/ref.

Wampler:1985:ERT

[Wam85]  John E. Wampler.  En-

hancing real-time perception of quantum limited images from a doubly intensified SIT camera system. Computer Vision, Graphics, and Image Processing, 32 (2):208–220, November 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Watson:1987:CTR

[Wat87] Andrew B. Watson. The cortex transform: Rapid computation of simulated neural images. Computer Vision, Graphics, and Image Processing, 39(3):311–327, September 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wu:1982:ABA

[WB82] J. K. Wu and R. E. Burge. Adaptive bit allocation for image compression. Computer Graphics and Image Processing, 19(4):392– 400, August 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

White:1990:TMI

[WB90] B. White and D. Brzakovic. Two methods of image extension. Computer Vision, Graphics, and Image Processing, 50(3):342–352, June 1990.

CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wang:1997:DCE

[WB97] Yang Wang and Prabir Bhattacharya. Digital connectivity and extended wellcomposed sets for gray images. Computer Vision and Image Understanding: CVIU, 68(3):330–345, December 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0551/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0551/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0551/production/ref.

Williams:2001:SRM

[WB01] John Williams and Mohammed Bennamoun. Simultaneous registration of multiple corresponding point sets. Computer Vision and Image Understanding: CVIU, 81(1):117–142, January 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0884;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0884/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0884/ref.

Wu:2011:MRF

[WB11] Dijia Wu and Kim L. Boyer. Markov random field based phase demodulation of interferometric images. Computer Vision and Image Understanding: CVIU, 115(6): 759–770, June 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2012:AGS

[WB12] Quan Wang and Kim L. Boyer. The active geometric shape model: a new robust deformable shape model and its applications. Computer Vision and Image Understanding: CVIU, 116(12): 1178–1194, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001154.

Wiejak:1985:CSM

[WBB85] J. S. Wiejak, H. Buxton, and B. F. Buxton. Convolution with separable masks for early image processing. Computer Vision, Graphics, and Image Processing, 32 (3):279–290, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wendt:2007:FVF

[WBOL07] Jeremy D. Wendt, William Baxter, Ipek Oguz, and Ming C. Lin. Finite volume flow simulations on arbitrary domains. Graphical Models, 69(1):19–32, January 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000415.

Wu:1986:CGP

[WBR86] Angela Y. Wu, S. K. Bhaskar, and Azriel Rosenfeld. Computation of geometric properties from the medial axis transform in O(n log n) time. Computer Vision, Graphics, and Image Processing, 34(1): 76–92, April 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wu:1988:PCG

[WBR88] Angela Wu, S. K. Bhaskar, and Azriel Rosenfeld. Parallel computation of geometric properties from the medial axis transform. Computer Vision, Graphics, and Image Processing, 41(3):323– 332, March 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Weiman:1979:LSG

[WC79] Carl F. R. Weiman and George Chaikin. Logarithmic spiral grids for image processing and display. Computer Graphics and Image Processing, 11(3):197– 226, November 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wu:1992:SFM

[WC92] Chung-Ming Wu and Yung-Chang Chen. Statistical feature matrix for texture analysis. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(5):407– 419, September 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Wang:1999:PMR

[WC99] Ranxiao Frances Wang and James E. Cutting. A probabilistic model for recovering camera translation. Computer Vision and Image Understanding: CVIU, 76(3): 205–212, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0798/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1999.0798/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1999.
                                        0798/production/ref.

Wu:2010:CCG

[WCF10] Lin Wu, Xiaochun Cao, and Hassan Foroosh. Camera calibration and geo-location estimation from two shadow trajectories. Computer Vision and Image Understanding: CVIU, 114(8):915–927, August 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wilson:1998:SMA

[WCH98] Richard C. Wilson, Andrew D. J. Cross, and Edwin R. Hancock. Structural matching with active triangulations. Computer Vision and Image Understanding: CVIU, 72(1):21–38, October 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0656/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0656/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0656/production/ref.

Wang:2002:ITM

[WCZ02] Qing Wang, Zheru Chi, and Rongchun Zhao. Image thresholding by maximizing the index of nonfuzziness of the 2-D grayscale histogram. Computer Vision and Image Understanding: CVIU, 85(2): 100–116, February 2002. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2007:VIS

[WCZ^{+}07] Qi Wang, Xilin Chen, Liang-Guo Zhang, Chunli Wang, and Wen Gao. Viewpoint invariant sign language recognition. Computer Vision and Image Understanding: CVIU, 108(1–2):87–97, October/November 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wall:1984:FSM

[WD84] Karin Wall and Per-Erik E. Danielsson. A fast sequential method for polygonal approximation of digitized curves. Computer Vision, Graphics, and Image Processing, 28 (2):220–227, November 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Won:1992:USN

[WD92] Chee Sun Won and Haluk Derin. Unsupervised segmentation of noisy and textured images using Markov random fields. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(4): 308–328, July 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Wang:1996:RTD

[WD96] Wendong Wang and James H. Duncan. Recovering the three-dimensional motion and structure of multiple moving objects from binocular image flows. Computer Vision and Image Understanding: CVIU, 63(3):430–446, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0033/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0033/production/
pdf.

Weinhaus:1999:PTM

[WD99] Frederick M. Weinhaus and Robert N. Devich. Photogrammetric texture mapping onto planar polygons. Graphical Models and Image Processing: GMIP, 61 (2):63–83, March 1999. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0491/production;   http:
//www.idealibrary.com/links/
artid/gmip.1999.0491/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1999.
0491/production/ref.

Widynski:2012:FSC

[WDB12] Nicolas Widynski, Séverine Dubuisson, and Isabelle Bloch. Fuzzy spatial constraints and ranked partitioned sampling approach for multiple object tracking. Computer Vision and Image Understanding: CVIU, 116(10):1076–1094, October 2012. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001038.

Wang:2012:DMB

[WDN^{+}12] Tao Wang, Guojun Dai, Bingbing Ni, De Xu, and Francois Siewe. A distance measure between labeled combinatorial maps. Computer Vision and Image Understanding: CVIU, 116(12):1168–1177, December 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001130.

Webb:1983:SC

[Web83] Jon A. Webb. Shape and correspondence. Computer Vision, Graphics, and Image Processing, 21(1):145– 160, January 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wechsler:1978:NLL

[Wec78] Harry Wechsler. New lowlevel procedure for image segmentation. Computer Graphics and Image Processing, 7 (1):120–129, February 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wechsler:1979:SAS

[Wec79]  Harry Wechsler.  Struc-

tural approach to shape analysis using mirroring axes.

Computer Graphics and Image Processing, 9(3):246– 266, March 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wechsler:1981:NFA

[Wec81] Harry Wechsler. New and fast algorithm for estimating the perimeter of objects for industrial vision tasks. Computer Graphics and Image Processing, 17(4):375– 385, December 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Weiss:1988:SRC

[Wei88]  Isaac Weiss.  3D shape

representation by contours. Computer Vision, Graphics, and Image Processing, 41 (1):80–100, January 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Weinshall:1990:QDS

[Wei90] Daphna Weinshall. Qualitative depth from stereo, with applications. Computer Vision, Graphics, and Image Processing, 49(2):222– 241, February 1990. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Weinshall:1992:SSC

[Wei92] Daphna Weinshall. Shortcuts in shape classification from two images. Computer Vision, Graphics, and Image Processing. Image Understanding, 56 (1):57–68, July 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Weszka:1978:STS

[Wes78] Joan S. Weszka. Survey of threshold selection techniques. Computer Graphics and Image Processing, 7(2): 259–265, April 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Welling:2006:RVF

[WEY06] Joel S. Welling, William F. Eddy, and Terence K. Young. Rotation of 3D volumes by fourier-interpolated shears. Graphical Models, 68(4):356– 370, July 2006. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000949.

Wechsler:1978:IPA

[WF78] H. Wechsler and K. S. Fu. Image processing algorithms applied to rib boundary detection in chest radiographs. Computer Graphics and Image Processing, 17(3): 375–390, June 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wanderley:2002:SFP

[WF02] Juliana F. Camapum Wanderley and Mark H. Fisher. Spatial-feature parametric clustering applied to motionbased segmentation in camouflage. Computer Vision and Image Understanding: CVIU, 85(2):144–157, February 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Woodard:2005:FSB

[WF05] Damon L. Woodard and Patrick J. Flynn. Finger surface as a biometric identifier. Computer Vision and Image Understanding: CVIU, 100 (3):357–384, December 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wong:1978:SMI

[WH78] Robert Y. Wong and Ernest L.

Hall. Scene matching with invariant moments. Computer Graphics and Image Processing, 8(1):16–24, August 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wang:1984:AMS

[WH84] Shyuan Wang and Robert M.

Haralick. Automatic multithreshold selection. Computer Vision, Graphics, and Image Processing, 25(1):46– 67, January 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Weng:1994:PCV

[WH94] Juyang Y. Weng and T. S. Huang. Performance of computer vision algorithms. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2): 253–256, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1052/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1052/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1057/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1057/production/
pdf.

Williams:1996:PCO

[WH96] Lance R. Williams and Allen R. Hanson. Perceptual completion of occluded surfaces. Computer Vision and Image Understanding: CVIU, 64(1):1–20, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0043/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0043/production/
pdf.

Wilson:2000:BVA

[WH00] Richard C. Wilson and Edwin R. Hancock. Biasvariance analysis for controlling adaptive surface meshes. Computer Vision and Image Understanding: CVIU, 77(1):25–47, January 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0810/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0810/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0810/production/ref.

Wang:2001:STM

[WH01] Xiaoguang Wang and Allen R.

Hanson. Surface texture and microstructure extraction from multiple aerial images. Computer Vision and Image Understanding: CVIU, 83(1):1–37, July 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0916;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0916/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0916/ref.

Whatmough:1991:ATS

[Wha91] R. J. Whatmough. Automatic threshold selection from a histogram using the ‘exponential hull’. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(6): 592–600, November 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Weber:2012:SFP

[WHHB12] Christopher Weber, Stefanie Hahmann, Hans Hagen, and Georges-Pierre Bonneau. Sharp feature preserving MLS surface reconstruction based on local feature line approximations. Graphical Models, 74(6):335–345, November 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031200032X.

Whitaker:1993:GLD

[Whi93] Ross T. Whitaker. Geometrylimited diffusion in the characterization of geometric patches in images. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):111–120, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1007/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1007/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.

1007/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1007/production/
pdf.

Wermser:1984:SBS

[WHL84] D. Wermser, G. Haussmann, and C. E. Liedtke. Segmentation of blood smears by hierarchical thresholding. Computer Vision, Graphics, and Image Processing, 25 (2):151–168, February 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wang:2009:SSK

[WHM^{+}09] Meng Wang, Xian-Sheng Hua, Tao Mei, Richang Hong, Guojun Qi, Yan Song, and Li-Rong Dai. Semi-supervised kernel density estimation for video annotation. Computer Vision and Image Understanding: CVIU, 113(3): 384–396, March 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wu:2008:MNR

[WHN08] Xiao Wu, Alexander G.

Hauptmann, and Chong-Wah Ngo. Measuring novelty and redundancy with multiple modalities in crosslingual broadcast news. Computer Vision and Image Understanding: CVIU, 110(3): 418–431, June 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Williams:1978:EAP

[Wil78] Charles M. Williams. An efficient algorithm for the piecewise linear approximation of planar curves. Computer Graphics and Image Processing, 8(2):286–293, October 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Williams:1979:IPC

[Wil79] Robin Williams. Image processing and computer graphics. Computer Graphics and Image Processing, 10(2): 183–193, June 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Williams:1981:BSL

[Wil81] Charles M. Williams. Bounded straight-line approximation of digitized planar curves and lines. Computer Graphics and Image Processing, 16(4): 370–381, August 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Williams:1984:TAD

[Wil84] Charles M. Williams. The trapezoidal approximation of digitized images. Computer Vision, Graphics, and Image Processing, 27(1):64– 77, July 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Willson:1989:CIM

[Wil89] Stephen J. Willson. Convergence of iterated median rules. Computer Vision, Graphics, and Image Processing, 47(1):105–110, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wilkinson:1998:OED

[Wil98] Michael H. F. Wilkinson.

Optimizing edge detectors for robust automatic threshold selection: Coping with edge curvature and noise. Graphical Models and Image Processing: GMIP, 60 (5):385–401, September 1998. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1998.
0478/production;   http:
//www.idealibrary.com/links/
artid/gmip.1998.0478/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1998.
0478/production/ref.

Wang:2007:MVF

[WJ07] Peng Wang and Qiang Ji.

Multi-view face and eye detection using discriminant features. Computer Vision and Image Understanding: CVIU, 105(2):99–111, February 2007. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2002:CQS

[WJG02] Wenping Wang, Barry Joe, and Ronald Goldman. Computing quadric surface intersections based on an analysis of plane cubic curves. Graphical Models, 64(6):335– 367, November 2002. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wang:1994:DMD

[WJW94] Wenping P. Wang, Barry Joe, and C. Y. Wang. On the difference method for drawing conic arcs. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(1):8–18, January 1994. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1994.
1002/production;   http:
//www.idealibrary.com/links/
artid/cgip.1994.1002/production/
pdf.

Wang:2013:MRF

[WKP13] Chaohui Wang, Nikos Komodakis, and Nikos Paragios. Markov Random Field modeling, inference & learning in computer vision & image understanding: a survey. Computer Vision and Image Understanding: CVIU, 117(11):1610–1627, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001343.

Woo:1985:TCC

[WL85] Tony C. Woo and Hyun-Chan C. Lee. On the time complexity for circumscribing a convex polygon. Computer Vision, Graphics, and Image Processing, 30(3):362– 363, June 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wu:1988:LIB

[WL88] Zhongquan Wu and Lingxiao Li. A line-integration based method for depth recovery from surface normals. Computer Vision, Graphics, and Image Processing, 43(1):53–66, July 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wang:1989:MIR

[WL89] Yuan Mei Wang and Wei Xue Lu. Multicriterion image reconstruction and implementation. Computer Vision, Graphics, and Image Processing, 46(1):131–135, April 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wagner:1999:RME

[WLD99] Robert Wagner, Feiyu Liu, and Klaus Donner. Robust motion estimation for calibrated cameras from monocular image sequences. Computer Vision and Image Understanding: CVIU, 73 (2):258–268, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0739/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0739/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0739/production/ref.

Watson:1985:TCD

[WLH85] Layne T. Watson, Thomas J. Laffey, and Robert M. Haralick. Topographic classification of digital image intensity surfaces using generalized splines and the discrete cosine transformation. Computer Vision, Graphics, and Image Processing, 29 (2):143–167, February 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wong:2008:MBA

[WLI08] Sam T. S. Wong, Howard Leung, and Horace H. S. Ip. Model-based analysis of Chinese calligraphy images. Computer Vision and Image Understanding: CVIU, 109(1):

69–85, January 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wu:2008:TSO

[WLMG08] Yiming Wu, Xiuwen Liu, Washington Mio, and K. A. Gallivan. Two-stage optimal component analysis. Computer Vision and Image Understanding: CVIU, 110(1): 91–101, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2006:STI

[WLW06] Wencheng Wang, Kuiyu Li, and Enhua Wu. Stick textures for image-based rendering. Graphical Models, 68(3):294–306, May 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000154.

Wang:2004:DCS

[WLZW04] Ze Wang, Chi-Sing Leung, Yi-Sheng Zhu, and Tien-Tsin Wong. Data compression with spherical wavelets and wavelets for the image-based relighting. Computer Vision and Image Understanding: CVIU, 96(3):327–344, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Williams:1992:FAA

[WM92] Donna J. Williams and Shah Mubarak. A fast algorithm for active contours and curvature estimation. Computer Vision, Graphics, and Image Processing. Image Understanding, 55 (1):14–26, January 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Williams:1993:ECU

[WM93] Donna J. Williams and Shah Mubarak. Edge characterization using normalized edge detector. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(4):311–318, July 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1021/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1021/production/
pdf.

Wiliem:2012:SBD

[WMBY12] Arnold Wiliem, Vamsi Madasu, Wageeh Boles, and Prasad Yarlagadda. A suspicious behaviour detection using a context space model for smart surveillance systems. Computer Vision and Image Understanding: CVIU, 116 (2):194–209, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211002001.

Westphal:1986:TDT

[WN86] H. Westphal and H. H.

Nagel. Toward the derivation of three-dimensional descriptions from image sequences for nonconvex moving objects. Computer Vision, Graphics, and Image Processing, 34(3):302–320, June 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Westphal:1987:ERP

[WN87] H. Westphal and H. H. Nagel.

Exploiting reflectance properties to analyze images of moving objects needs local constraints. Computer Vision, Graphics, and Image Processing, 38(1):90–98, April 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wachter:1999:TPM

[WN99] S. Wachter and H.-H. Nagel.

Tracking persons in monocular image sequences. Computer Vision and Image Understanding: CVIU, 74(3): 174–192, June 1999. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.
idealibrary.com/links/artid/
cviu.1999.0758/production;
 http://www.idealibrary.

                                        com/links/artid/cviu.1999.
                                        0758/production/pdf.

Wong:2005:FAD

[WNH05] Andrew K. C. Wong, Peiyi Niu, and Xiang He. Fast acquisition of dense depth data by a new structured light scheme. Computer Vision and Image Understanding: CVIU, 98(3):398–422, June 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2010:GED

[WO10] Hongzhi Wang and John Oliensis. Generalizing edge detection to contour detection for image segmentation. Computer Vision and Image Understanding: CVIU, 114 (7):731–744, July 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wojcik:1984:ARC

[Woj84] Zbigniew M. Wojcik. An approach to the recognition of contours and lineshaped objects. Computer Vision, Graphics, and Image Processing, 25(2):184– 204, February 1984. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wojcik:1987:RAS

[Woj87]  Zbigniew Wojcik.  Rough

approximation of shapes in pattern recognition. Computer Vision, Graphics, and Image Processing, 40(2):228– 249, November 1987. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Worthington:2005:RDS

[Wor05] Philip L. Worthington. Reilluminationdriven shape from shading. Computer Vision and Image Understanding: CVIU, 98(2): 325–343, May 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Watson:1984:SST

[WP84] D. F. Watson and G. M.

Philip. Survey: Systematic triangulations. Computer Vision, Graphics, and Image Processing, 26(2):217– 223, May 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Werman:1988:GLR

[WP88] Michael Werman and Shmuel Peleg. Gray level requantization. Computer Vision, Graphics, and Image Processing, 43(1):81–87, July 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wang:1993:DCS

[WP93a] Li Wang and Theo Pavlidis.

Detection of curved and straight segments from gray scale topography. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3):352– 365, November 1993. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1047/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1047/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1049/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1049/production/
pdf.

Whitaker:1993:MSA

[WP93b] Ross T. Whitaker and Stephen M. Pizer. A multi-scale approach to nonuniform diffusion. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(1):99–110, January 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1006/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1006/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1006/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1006/production/
pdf.

Wolter:2000:SIC

[WP00] Franz-Erich Wolter and Nicholas M.

Patrikalakis. Special issue for CGI ’98. Graphical Models, 62(1):1, January 2000. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/artid/gmod.1999.
0516/production;   http:
//www.idealibrary.com/links/
artid/gmod.1999.0516/production/
pdf.

Withey:2009:DET

[WPK09] D. J. Withey, W. Pedrycz, and Z. J. Koles. Dynamic edge tracing: Boundary identification in medical images. Computer Vision and Image Understanding: CVIU, 113(10):1039–1052, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Werman:1985:DMM

[WPR85] Michael Werman, Shmuel Peleg, and Azriel Rosenfeld. A distance metric for multidimensional histograms. Computer Vision, Graphics, and Image Processing, 32 (3):328–336, December 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wang:2003:BSM

[WPS03] Yongmei Wang, Bradley S.

Peterson, and Lawrence H.

Staib. 3D brain surface matching based on geodesics and local geometry. Computer Vision and Image Understanding: CVIU, 89(2– 3):252–271, February/March 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wu:1987:DSI

[WR87] Xiaolin Wu and Jon G. Rokne. Double-step incremental generation of lines and circles. Computer Vision, Graphics, and Image Processing, 37(3):331– 344, March 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Waksman:1993:SOT

[WR93] Adlai Waksman and Azriel Rosenfeld. Sparse, opaque three-dimensional texture. I. arborescent patterns. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(3):388–399, May 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1026/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1026/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1026/production;   http:
//www.idealibrary.com/links/

artid/cviu.1993.1026/production/
pdf.

Waksman:1996:SOT

[WR96] Adlai Waksman and Azriel Rosenfeld. Sparse, opaque three-dimensional texture. 2a. visibility. Graphical Models and Image Processing: GMIP, 58(2):155–163, March 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic).

Williams:2005:MMS

[WR05] Jason Williams and Jarek Rossignac. Mason: morphological simplification. Graphical Models, 67(4):285–303, July 2005. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Worz:2008:PBE

[WR08] Stefan Wörz and Karl Rohr.

Physics-based elastic registration using non-radial basis functions and including landmark localization uncertainties. Computer Vision and Image Understanding: CVIU, 111(3):263–274, September 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Weinland:2006:FVA

[WRB06] Daniel Weinland, Remi Ronfard, and Edmond Boyer. Free viewpoint action recognition using motion history volumes.

Computer Vision and Image Understanding: CVIU, 104(2–3):249–257, November/December 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Weinland:2011:SVB

[WRB11] Daniel Weinland, Remi Ronfard, and Edmond Boyer. A survey of vision-based methods for action representation, segmentation and recognition. Computer Vision and Image Understanding: CVIU, 115 (2):224–241, February 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Williams:1997:IDM

[WRH97] Christopher K. I. Williams, Michael Revow, and Geoffrey E. Hinton. Instantiating deformable models with a neural net. Computer Vision and Image Understanding: CVIU, 68(1):120–126, October 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0540/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0540/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0540/production/ref.

Walther:2005:SVA

[WRKP05] Dirk Walther, Ueli Rutishauser, Christof Koch, and Pietro Perona. Selective visual attention enables learning and recognition of multiple objects in cluttered scenes. Computer Vision and Image Understanding: CVIU, 100(1– 2):41–63, October/November 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:1989:CNI

[WS89] Dacheng Wang and Sargur N. Srihari. Classification of newspaper image blocks using texture analysis. Computer Vision, Graphics, and Image Processing, 47 (3):327–352, September 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Williams:1990:ECU

[WS90] Donna J. Williams and Mubarak Shah. Edge contours using multiple scales. Computer Vision, Graphics, and Image Processing, 51 (3):256–274, September 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wuthrich:1991:ACB

[WS91] C. A. Wuthrich and P. Stucki.

An algorithmic comparison between square- and hexagonal-based grids. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):324–339, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Worring:1993:DCE

[WS93] Marcel Worring and Arnold W. M. Smeulders. Digital curvature estimation. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(3): 366–382, November 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1993.
                                        1048/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1993.1048/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1993.
                                        1050/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1993.1050/production/
                                        pdf.

Wang:2003:EMD

[WS03] Yang Wang and Dimitris Samaras. Estimation of multiple directional light sources for synthesis of augmented reality images. Graphical Models, 65(4):185–205, July 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wong:2006:TBM

[WS06] King Yuen Wong and Minas E. Spetsakis. Tracking based motion segmentation under relaxed statistical assumptions. Computer Vision and Image Understanding: CVIU, 101(1): 45–64, January 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2008:VLR

[WS08] Liang Wang and David Suter.

Visual learning and recognition of sequential data manifolds with applications to human movement analysis. Computer Vision and Image Understanding: CVIU, 110 (2):153–172, May 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2012:EMD

[WSC^{+}12] Hui Wang, Zhixun Su, Junjie Cao, Ye Wang, and Hao Zhang. Empirical mode decomposition on surfaces. Graphical Models, 74 (4):173–183, July 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000252.

Wan:2012:SUP

[WSCO^{+}12] Guowei Wan, Noah Snavely, Daniel Cohen-Or, Qian Zheng, Baoquan Chen, and Sikun Li. Sorting unorganized photo sets for urban reconstruction. Graphical Models, 74(1):14–28, January 2012.

CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000658.

Woo:2013:MSU

[WSKH13] Jonghye Woo, Piotr J. Slomka, C.-C. Jay Kuo, and Byung-Woo Hong. Multiphase segmentation using an implicit dual shape prior: Application to detection of left ventricle in cardiac MRI. Computer Vision and Image Understanding: CVIU, 117(9):1084–1094, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000684.

Worring:1996:PFB

[WSSD96] Marcel Worring, Arnold W. M. Smeulders, Lawrence H. Staib, and James S. Duncan. Parameterized feasible boundaries in gradient vector fields. Computer Vision and Image Understanding: CVIU, 63(1):135–144, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0009/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0009/production/
pdf.

Wang:2013:ASI

[WSSS13] Zhimin Wang, Qing Song, Yeng Chai Soh, and Kang Sim. An adaptive spatial information-theoretic fuzzy clustering algorithm for image segmentation. Computer Vision and Image Understanding: CVIU, 117(10): 1412–1420, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000957.

Walker:1991:EDL

[WSV91] Michael W. Walker, Lejun Shao, and Richard A. Volz. Estimating 3-D location parameters using dual number quaternions. Computer Vision, Graphics, and Image Processing. Image Understanding, 54(3):358–367, November 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Wang:2005:EEG

[WSV05] Jian-Gang Wang, Eric Sung, and Ronda Venkateswarlu. Estimating the eye gaze from one eye. Computer Vision and Image Understanding: CVIU, 98(1):83–103, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wu:2013:MSG

[WSZL13] Jinliang Wu, Xiaoyong Shen, Wei Zhu, and Ligang Liu. Mesh saliency with global rarity. Graphical Models, 75 (5):255–264, September 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000180.

Wu:1993:HTU

[Wu93] Z. Y. Wu. Homogeneity testing for unlabeled data: a performance evaluation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(5):370–380, September 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cgip.1993.
                                        1028/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cgip.1993.1028/production/
                                        pdf.

Wu:2002:DMD

[Wu02] Wen-Yen Wu. A dynamic method for dominant point detection. Graphical Models, 64(5):304–315, September 2002. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Winkler:1978:MC

[WV78] G. Winkler and K. Vattrodt. Measures for conspicuousness. Computer Graphics and Image Processing, 8(3): 355–368, December 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wan:1997:SPC

[WV97] Wenhua Wan and José A.

Ventura. Segmentation of planar curves into straightline segments and elliptical arcs. Graphical Models and Image Processing: GMIP, 59 (6):484–494, November 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0450/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0450/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0450/production/ref.

Wang:1981:GIW

[WVL81] David C. C. Wang, Anthony H. Vagnucci, and C. C. Li. Gradient inverse weighted smoothing scheme and the evaluation of its performance. Computer Graphics and Image Processing, 15(2):167– 181, February 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wang:1983:DIE

[WVL83] David C. C. Wang, Anthony H. Vagnucci, and C. C. Li. Digital image enhancement: a survey. Computer Vision, Graphics, and Image Processing, 24(3):363– 381, December 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wallace:1980:ETD

[WW80] Timothy P. Wallace and Paul A. Wintz. An efficient three-dimensional aircraft recognition algorithm using normalized Fourier descriptors. Computer Graphics and Image Processing, 13 (2):99–126, June 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Widmayer:1988:TSO

[WW88] Peter Widmayer and Derick Wood. A time- and space-optimal algorithm for Boolean mask operations for orthogonal polygons. Computer Vision, Graphics, and Image Processing, 41(1):14– 27, January 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wohn:1990:ASI

[WW90] Kwangyoen Wohn and Allen M. Waxman. The analytic structure of image flows: deformation and segmentation. Computer Vision, Graphics, and Image Processing, 49 (2):127–151, February 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wu:1991:DII

[WW91] Jian Wu and K. Wohn.

On the deformation of image intensity and zerocrossing contours under motion. Computer Vision, Graphics, and Image Processing. Image Understanding, 53 (1):66–75, January 1991. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Wu:1993:DDP

[WW93] Wen-Yen Y. Wu and Mao-Jiun J. J. Wang. Detecting the dominant points by the curvature-based polygonal approximation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(2):79–88, March 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1006/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1006/production/
pdf.

Wang:1994:MCF

[WW94] Y. F. Wang and Jih-Fang F.

Wang. On 3D model construction by fusing heterogeneous sensor data. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):210–229, September 1994. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1048/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1048/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1053/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1053/production/
pdf.

Wang:1995:HOD

[WW95] Guo-Zhao Wang and Guo-Jin Wang. Higher order derivatives of a rational Bézier curve. Graphical Models and Image Processing: GMIP, 57(3):246–253, May 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1023/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1023/production/
pdf.

Williams:1997:APV

[WW97] James Williams and Lawrence Wolff. Analysis of the pulmonary vascular tree using differential geometry based vector fields. Computer Vision and Image Understanding: CVIU, 65 (2):226–236, February 1997.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0571/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0571/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1996.
0571/production/ref.

Wu:1984:ADS

[WWB84] C. K. Wu, D. Q. Wang, and R. K. Bajcsy. Acquiring 3-D spatial data of a real object. Computer Vision, Graphics, and Image Processing, 28 (1):126–133, October 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wahl:1982:BST

[WWC82] Friedrich M. Wahl, Kwan Y.

Wong, and Richard G. Casey. Block segmentation and text extraction in mixed text/ image documents. Computer Graphics and Image Processing, 20(4):375– 390, December 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Wachsmuth:2007:CIV

[WWH07] Sven Wachsmuth, Sebastian Wrede, and Marc Hanheide. Coordinating interactive vision behaviors for cognitive assistance. Computer Vision and Image Understanding: CVIU, 108(1–2):135– 149, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Waldman:1988:NCM

[WWHL88] G. Waldman, J. Wootton, G. Hobson, and K. Luetkemeyer. A normalized clutter measure for images. Computer Vision, Graphics, and Image Processing, 42(2):137– 156, May 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wu:2013:ADF

[WWJ13a] Yuwei Wu, Yuanquan Wang, and Yunde Jia. Adaptive diffusion flow active contours for image segmentation. Computer Vision and Image Understanding: CVIU, 117(10): 1421–1435, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213001033.

Wu:2013:SLV

[WWJ13b] Yuwei Wu, Yuanquan Wang, and Yunde Jia. Segmentation of the left ventricle in cardiac cine MRI using a shape-constrained snake model. Computer Vision and Image Understanding: CVIU, 117(9):990–1003, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000787.

Wu:1992:PES

[WWL92] Wen-Yen Wu, Mao-Jiun J. J.

Wang, and Chih-Ming Liu. Performance evaluation of some noise reduction methods. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(2):134– 146, March 1992. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Weise:2011:OLC

[WWLV11] Thibaut Weise, Thomas Wismer, Bastian Leibe, and Luc Van Gool. Online loop closure for real-time interactive 3D scanning. Computer Vision and Image Understanding: CVIU, 115(5):635–648, May 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wallace:1989:NAI

[WWW89a] Richard S. Wallace, Jon A.

Webb, and I-Chen C. Wu. Note: An architectureindependent programming language for low-level vision. Computer Vision, Graphics, and Image Processing, 48 (2):246–264, November 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wallace:1989:NMI

[WWW89b] Richard S. Wallace, Jon A. Webb, and I-Chen C. Wu. Note: Machine-independent image processing: Performance of apply on diverse architectures. Computer Vision, Graphics, and Image Processing, 48(2):265– 276, November 1989. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Wu:1995:RAF

[WWW95] Zhongquan Wu, Lide Wu, and Aicheng Wu. The robust algorithms for finding the center of an arc. Computer Vision and Image Understanding: CVIU, 62(3):269–278, November 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1054/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1054/production/
pdf.

Wang:2012:RCS

[WWWM12] Sen Wang, Jianhuang Wu, Mingqiang Wei, and Xin Ma. Robust curve skeleton extraction for vascular structures. Graphical Models, 74(4):109–120, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S152407031200015X.

Wang:1991:NTC

[WX91] Guojin Wang and Wei Xu.

Note: The termination criterion for subdivision of the rational Bézier curves. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1):93–96, January 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Wang:2007:VNF

[WXRA07] Hongcheng Wang, Ning Xu, Ramesh Raskar, and Narendra Ahuja. Videoshop: a new framework for spatiotemporal video editing in gradient domain. Graphical Models, 69(1):57–70, January 2007. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000439.

Willick:1991:EEM

[WY91] Darryl Willick and Yee-Hong Yang. Experimental evaluation of motion constraint equations. Computer Vision, Graphics, and Image Processing. Image Understanding, 54 (2):206–214, September 1991. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Wang:2007:STM

[WY07] Jun Wang and Lijun Yin.

Static topographic modeling for facial expression recognition and analysis. Computer Vision and Image Understanding: CVIU, 108(1– 2):19–34, October/November 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2011:QMS

[WY11] Jun Wang and Zeyun Yu. Quality mesh smoothing via local surface fitting and optimum projection. Graphical Models, 73(4):127–139, July 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070311000051.

Wyvill:2003:SII

[Wyv03] Brian Wyvill. Special issue on the International Conference of Shape Modeling (SMI) 2002. Graphical Models, 65 (5):259, September 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Wang:1997:BMC

[WZ97] Guo-Zhao Wang and Jian-Min Zheng. Bounds on the moving control points of hybrid curves. Graphical Models and Image Processing: GMIP, 59(1):19–25, January 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.

0411/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0411/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1996.
0411/production/ref.

Wang:2004:DIO

[WZ04] Yongmei Michelle Wang and Hongjiang Zhang. Detecting image orientation based on low-level visual content. Computer Vision and Image Understanding: CVIU, 93(3): 328–346, March 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wan:2008:SVU

[WZ08] Dingrui Wan and Jie Zhou.

Stereo vision using two PTZ cameras. Computer Vision and Image Understanding: CVIU, 112(2):184–194, November 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Wang:2003:LKB

[WZL^{+}03] Tian-Shu Wang, Nan-Ning Zheng, Yan Li, Ying-Qing Xu, and Heung-Yung Shum. Learning kernel-based HMMs for dynamic sequence synthesis. Graphical Models, 65 (4):206–221, July 2003. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Williams:1999:CSC

[WZWT99] Lance Williams, John Zweck, Tairan Wang, and Karvel Thornber. Computing stochastic completion fields in lineartime using a resolution pyramid. Computer Vision and Image Understanding: CVIU, 76(3):289–297, December 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0800/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0800/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0800/production/ref.

Xu:2007:OSU

[XAB07] Ning Xu, Narendra Ahuja, and Ravi Bansal. Object segmentation using graph cuts based active contours. Computer Vision and Image Understanding: CVIU, 107(3): 210–224, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xie:2013:RMI

[XFSC13] Jing Xie, Evan Fletcher, Baljeet Singh, and Owen Carmichael. Robust measurement of individual localized changes to the aging hippocampus. Computer Vision and Image Understanding: CVIU, 117(9): 1128–1137, September 2013.

CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000696.

Xiang:2008:IAA

[XG08a] Tao Xiang and Shaogang Gong. Incremental and adaptive abnormal behaviour detection. Computer Vision and Image Understanding: CVIU, 111(1):59–73, July 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xiang:2008:ODG

[XG08b] Tao Xiang and Shaogang Gong. Optimising dynamic graphical models for video content analysis. Computer Vision and Image Understanding: CVIU, 112(3):310– 323, December 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2012:SST

[XHJF12] Yong Xu, Sibin Huang, Hui Ji, and Cornelia Fermüller. Scale-space texture description on SIFT-like textons. Computer Vision and Image Understanding: CVIU, 116(9):999–1013, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000781.

Xiao:2009:GMG

[XHW09] Bai Xiao, Edwin R. Hancock, and Richard C. Wilson. A generative model for graph matching and embedding. Computer Vision and Image Understanding: CVIU, 113(7):777–789, July 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2012:ISP

[XJK12] Li Xu, Jiaya Jia, and Sing Bing Kang. Improving sub-pixel correspondence through upsampling. Computer Vision and Image Understanding: CVIU, 116 (2):250–261, February 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S107731421100258X.

Xu:1988:SRO

[XL88] S. B. Xu and W. X. Lu. Surface reconstruction of 3D objects in computerized tomography. Computer Vision, Graphics, and Image Processing, 44(3):270–278, December 1988. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Xiong:1998:ESC

[XL98] Wei Xiong and John Chung-Mong Lee. Efficient scene change detection and camera motion annotation for video classification. Computer Vision and Image Understanding: CVIU, 71 (2):166–181, August 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0711/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0711/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0711/production/ref.

Xu:1993:RHT

[XO93] Lei Xu and Erkki Oja.

Randomized Hough transform (RHT): basic mechanisms, algorithms, and computational complexities. Computer Vision, Graphics, and Image Processing. Image Understanding, 57 (2):131–154, March 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1009/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1009/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1009/production;   http:
//www.idealibrary.com/links/
artid/cviu.1993.1009/production/
pdf.

Xiao:2005:EPS

[XOF05] G. Xiao, S. H. Ong, and K. W. C. Foong. Efficient partial-surface registration for 3D objects. Computer Vision and Image Understanding: CVIU, 98(2):271–293, May 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2011:SIC

[XP11] Mai Xu and Maria Petrou. 3D Scene interpretation by combining probability theory and logic: The tower of knowledge. Computer Vision and Image Understanding: CVIU, 115(11):1581–1596, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211001834.

Xiong:1998:DSD

[XS98] Yalin Xiong and Steven A. Shafer. Dense structure from a dense optical flow sequence. Computer Vision and Image Understanding: CVIU, 69(2):222–245, February 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0601/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0601/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0601/production/ref.

Xiao:2004:TVM

[XS04] Jiangjian Xiao and Mubarak Shah. Tri-view morphing. Computer Vision and Image Understanding: CVIU, 96(3): 345–366, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2012:TSA

[XSD12] Ziyue Xu, Punam K. Saha, and Soura Dasgupta. Tensor scale: an analytic approach with efficient computation and applications. Computer Vision and Image Understanding: CVIU, 116(10):1060–1075, October 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000896.

Xu:2004:DLM

[XST04] Yaowu Xu, Eli Saber, and A. Murat Tekalp. Dynamic learning from multiple examples for semantic object segmentation and search. Computer Vision and Image Understanding: CVIU, 95(3): 334–353, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2004:VHB

[XTLP04] Songhua Xu, Min Tang, Francis C. M. Lau, and Yunhe Pan. Virtual hairy brush for painterly rendering. Graphical Models, 66(5):263–302, September 2004. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Xumin:2010:HPU

[XWYY10] Liu Xumin, Xu Weixiang, Guan Yong, and Shang Yuanyuan. Hyperbolic polynomial uniform B-spline curves and surfaces with shape parameter. Graphical Models, 72(1):1–6, January 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000265.

Xu:2008:ASD

[XYW^{+}08] Shuchang Xu, Xiuzi Ye, Yin Wu, Franck Giron, Jean-Luc Leveque, and Bernard Querleux. Automatic skin decomposition based on single image. Computer Vision and Image Understanding: CVIU, 110(1):1–6, April 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xu:2011:LST

[XYW11] Jiang Xu, Junsong Yuan, and Ying Wu. Learning spatiotemporal dependency of local patches for complex motion segmentation. Computer Vision and Image Understanding: CVIU, 115(3): 334–351, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Xiao:2011:LIS

[XYZH11] Bai Xiao, Song Yi-Zhe, and Peter Hall. Learning invariant structure for object identification by using graph methods. Computer Vision and Image Understanding: CVIU, 115(7):1023–1031, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211000841.

Xu:2006:PSI

[XZWB06] Dong Xu, Hongxin Zhang, Qing Wang, and Hujun Bao. Poisson shape interpolation. Graphical Models, 68(3):268–281, May 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070306000208.

Yang:2012:SRN

[YA12] Qingxiong Yang and Narendra Ahuja. Surface reflectance and normal estimation from photometric stereo. Computer Vision and Image Understanding: CVIU, 116(7):793–802, July 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000483.

Yachida:1983:DVM

[Yac83] Masahiko Yachida. Determining velocity maps by spatiotemporal neighborhoods from image sequences. Computer Vision, Graphics, and Image Processing, 21(2):262– 279, February 1983. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yao:2008:ILR

[YAK^{+}08] Yi Yao, Besma R. Abidi, Nathan D. Kalka, Natalia A. Schmid, and Mongi A. Abidi. Improving long range and high magnification face recognition: Database acquisition, evaluation, and enhancement. Computer Vision and Image Understanding: CVIU, 111(2):111–125, August 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yamaguchi:1978:NCF

[Yam78] Fujio Yamaguchi. A new curve fitting method using a CRT computer display. Computer Graphics and Image Processing, 17(3): 425–437, June 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yamamoto:1979:MDC

[Yam79]  Hiromichi Yamamoto.  A

method of deriving compatibility coefficients for relaxation operators. Computer Graphics and Image Processing, 10(3):256–271, July 1979. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). YA-MAMOTO79.

Yamamoto:1980:SEL

[Yam80] Hiromichi Yamamoto. Some experiments on Landsat pixel classification using relaxation operators. Computer Graphics and Image Processing, 13 (1):31–45, May 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yan:1993:SCD

[Yan93a] Hong Yan. Skew correction of document images using interline cross- correlation. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(6):538–543, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1041/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1041/production/
pdf.

Yang:1993:ESA

[Yan93b]  Chengda D. Yang.  Effi-

cient stochastic algorithms on locally bounded image space. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55 (6):494–506, November 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1037/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1037/production/
pdf.

Yang:1997:FCT

[YAT97] Luren Yang, Fritz Albregtsen, and Torfinn Taxt. Fast computation of three-dimensional geometric moments using a discrete divergence theorem and a generalization to higher dimensions. Graphical Models and Image Processing: GMIP, 59(2):97–108, March 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0418/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0418/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0418/production/ref.

Yau:1984:GQC

[Yau84] Mann-May M. Yau. Generating quadtrees of cross sections from octrees. Computer Vision, Graphics, and Image Processing, 27(2):211– 238, August 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yanowitz:1989:NMI

[YB89] S. D. Yanowitz and A. M.

Bruckstein. A new method for image segmentation. Computer Vision, Graphics, and Image Processing, 46(1):82– 95, April 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yu:1995:GSA

[YB95] Shan Yu and Marc Berthod.

A game strategy approach for image labeling. Computer Vision and Image Understanding: CVIU, 61(1):32–37, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1003/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1003/production/
pdf.

Yacoob:1999:PMR

[YB99] Yaser Yacoob and Michael J.

Black. Parameterized modeling and recognition of activities. Computer Vision and Image Understanding: CVIU, 73(2):232–247, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0726/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/cviu.1998.0726/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1998.
                                        0726/production/ref.

Yin:2001:GRF

[YB01] Lijun Yin and Anup Basu. Generating realistic facial expressions with wrinkles for model-based coding. Computer Vision and Image Understanding: CVIU, 84(2): 201–240, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0949;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0949/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0949/ref.

Yan:2007:FAI

[YB07] Ping Yan and Kevin W. Bowyer. A fast algorithm for ICP-based 3D shape biometrics. Computer Vision and Image Understanding: CVIU, 107(3):195–202, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yoo:1993:NPD

[YBDC93] J. Yoo, C. A. Bouman, E. J. Delp, and E. J. Coyle. The nonlinear prefiltering and difference of estimates approaches to edge detection:

Applications of stack filters. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 55(2):140–159, March 1993. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). URL

http://www.idealibrary.
com/links/artid/cgip.1993.
1010/production;   http:
//www.idealibrary.com/links/
artid/cgip.1993.1010/production/
pdf.

Yakimovsky:1978:SED

[YC78a] Y. Yakimovsky and R. Cunningham. A system for extracting 3-D measurements from a stereo pair of TV cameras. Computer Graphics and Image Processing, 7(??): 195–209, ???? 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yakimovsky:1978:SET

[YC78b] Y. Yakimovsky and R. Cunningham. System for extracting three-dimensional measurements from a stereo pair of TV cameras. Computer Graphics and Image Processing, 7(2):195–210, April 1978. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yi:1998:MBO

[YC98] June Ho Yi and David M.

Chelberg. Model-based 3D object recognition using Bayesian indexing. Computer Vision and Image Understanding: CVIU, 69 (1):087–105, January 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0597/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1997.0597/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1997.
                                        0597/production/ref.

Yoo:2005:NNI

[YC05] Dong Hyun Yoo and Myung Jin Chung. A novel non-intrusive eye gaze estimation using cross-ratio under large head motion. Computer Vision and Image Understanding: CVIU, 98(1):25–51, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2010:ABF

[YCA^{+}10] Ming Yang, James Crenshaw, Bruce Augustine, Russell Mareachen, and Ying Wu. AdaBoost-based face detection for embedded systems. Computer Vision and Image Understanding: CVIU, 114(11):1116–1125, November 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yalcin:2007:BEU

[YCH07] Hulya Yalcin, Robert Collins, and Martial Hebert. Background estimation under rapid gain change in thermal imagery. Computer Vision and Image Understanding: CVIU, 106(2–3):148– 161, May/June 2007. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yao:2010:AOC

[YCKA10] Yi Yao, Chung-Hao Chen, Andreas Koschan, and Mongi Abidi. Adaptive online camera coordination for multicamera multi-target surveillance. Computer Vision and Image Understanding: CVIU, 114(4):463–474, April 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yu:2007:NSM

[YCL07] Yuan-Hui Yu, Chin-Chen Chang, and Iuon-Chang Lin. A new steganographic method for color and grayscale image hiding. Computer Vision and Image Understanding: CVIU, 107(3):183–194, September 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yacoob:1994:LHF

[YD94] Yaser Yacoob and Larry S.

Davis. Labeling of human face components from range data. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(2):168–178, September 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1045/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1045/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1050/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1050/production/
pdf.

You:1980:DSR

[YF80] K. C. You and K. S. Fu. Distorted shape recognition using attributed grammars and error-correcting techniques. Computer Graphics and Image Processing, 13 (1):1–16, May 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yuille:1998:IWS

[YFZ98] Alan L. Yuille, Mario Ferraro, and Tony Zhang. Image warping for shape recovery and recognition. Computer Vision and Image Understanding: CVIU, 72(3):351–359, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0676/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0676/production/

pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0676/production/ref.

Yilmaz:2007:COC

[YG07] Türker Yilmaz and Ugur Güdükbay. Conservative occlusion culling for urban visualization using a slicewise data structure. Graphical Models, 69(3–4):191– 210, May/July 2007. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070307000033.

Yang:2013:TLA

[YGC13] Wanqi Yang, Yang Gao, and Longbing Cao. TRASMIL: a local anomaly detection framework based on trajectory segmentation and multiinstance learning. Computer Vision and Image Understanding: CVIU, 117(10): 1273–1286, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001658.

Yang:2011:RLD

[YGH11] Kun Yang, Shuzhi Sam Ge, and Hongsheng He. Robust line detection using twoorthogonal direction image scanning. Computer Vision and Image Understanding: CVIU, 115(8):1207–1222, August 2011. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000981.

Yokoyama:1978:TSU

[YH78] Ryuzo Yokoyama and Robert M. Haralick. Texture synthesis using a growth model. Computer Graphics and Image Processing, 8(3):369– 381, December 1978. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yang:1981:EMF

[YH81] G. J. Yang and T. S. Huang. Effect of median filtering on edge location estimation. Computer Graphics and Image Processing, 15(3):224– 245, March 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yen:1983:DDM

[YH83] B. L. Yen and T. S. Huang. Determining 3-D motion and structure of a rigid body using the spherical projection. Computer Vision, Graphics, and Image Processing, 21(1):21–32, January 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yang:2011:SOD

[YHN11] Bo Yang, Chang Huang, and Ram Nevatia. Segmentation of objects in a detection window by Nonparametric Inhomogeneous CRFs. Computer Vision and Image Understanding: CVIU, 115(11): 1473–1482, November 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001779.

Yang:2005:SBD

[YHR^{+}05] Allen Y. Yang, Kun Huang, Shankar Rao, Wei Hong, and Yi Ma. Symmetry-based 3-D reconstruction from perspective images. Computer Vision and Image Understanding: CVIU, 99(2):210–240, August 2005. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Young:1989:ESE

[YHS89] A. A. Young, P. J. Hunter, and B. H. Smaill. Epicardial surface estimation from coronary angiograms. Computer Vision, Graphics, and Image Processing, 47(1):111– 127, July 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yi:1995:OSL

[YHS95] Seungku K. Yi, Robert M.

Haralick, and Linda G. Shapiro. Optimal sensor and light source positioning for machine vision. Computer Vision and Image Understanding: CVIU, 61(1):122–137, January 1995. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1995.
1009/production;   http:
//www.idealibrary.com/links/
artid/cviu.1995.1009/production/
pdf.

You:1984:PES

[YJ84] Zhisheng You and Anil K. Jain. Performance evaluation of shape matching via chord length distribution. Computer Vision, Graphics, and Image Processing, 28 (2):185–198, November 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yla-Jaaski:1996:GSS

[YJA96] Antti Ylä-Jääski and Frank Ade. Grouping symmetrical structures for object segmentation and description. Computer Vision and Image Understanding: CVIU, 63(3):399–417, May 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0031/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0031/production/
pdf.

Yu:2009:ACC

[YJC^{+}09] Xinguo Yu, Nianjuan Jiang, Loong-Fah Cheong, Hon Wai Leong, and Xin Yan. Automatic camera calibration of broadcast tennis video with applications to 3D virtual content insertion and ball detection and tracking. Computer Vision and Image Understanding: CVIU, 113(5): 643–652, May 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yla-Jaaski:1991:FDD

[YJKK91] J. Yla-Jaaski, F. Klein, and O. Kubler. Fast direct display of volume data for medical diagnosis. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(1): 7–18, January 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Yokokawa:1980:DNR

[YK80] Kanji Yokokawa and Tosiyasu L.

Kunii. Definition of neighborhood of a region for picture processing. Computer Graphics and Image Processing, 14 (2):112–144, October 1980. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yang:1986:DIP

[YK86] H. S. Yang and A. C. Kak.

Determination of the identity, position, and orientation of the topmost object in a pile. Computer Vision, Graphics, and Image Processing, 36(2/3):229–255, November/December 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yamada:1987:TOE

[YK87] Hiromitsu Yamada and Tony Kasvand. Transparent object extraction from regular textured backgrounds by using binary parallel operations. Computer Vision, Graphics, and Image Processing, 40(1):41–53, October 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yen:1995:DGS

[YK95] Chinching Yen and Shyh-Shiaw Kuo. Degraded grayscale text recognition using pseudo-2D hidden Markov models and N-best hypotheses. Graphical Models and Image Processing: GMIP, 57(2):131–145, March 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1014/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1014/production/
pdf.

Yitzhaky:1997:IBP

[YK97] Y. Yitzhaky and N. S.

Kopeika. Identification of blur parameters from motion blurred images. Graphical Models and Image Processing: GMIP, 59(5):310– 320, September 1997. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1997.
0435/production;   http:
//www.idealibrary.com/links/
artid/gmip.1997.0435/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1997.
0435/production/ref.

Yoon:2008:DSM

[YK08] Kuk-Jin Yoon and In So Kweon. Distinctive Similarity Measure for stereo matching under point ambiguity. Computer Vision and Image Understanding: CVIU, 112(2): 173–183, November 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2001:FDU

[YKA01] Ming-Hsuan Yang, David Kriegman, and Narendra Ahuja. Face detection using multimodal density models. Computer Vision and Image Understanding: CVIU, 84 (2):264–284, November 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0937;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/cviu.2001.0937/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/cviu.
                                        2001.0937/ref.

Yang:1986:ACF

[YKC^{+}86] Mark C. K. Yang, Chong-Kyo Kim, Kuo-Young Cheng, Chung-Chun Yang, and S. S. Liu. Automatic curve fitting with quadratic B-spline functions and its applications to computer-assisted animation. Computer Vision, Graphics, and Image Processing, 33(3):346–363, March 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Kwon:2011:ABF

[yKL11] Ji yong Kwon and In-Kwon Lee. An animation bilateral filter for slow-in and slowout effects. Graphical Models, 73(5):141–150, September 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070311000075.

Yuille:1990:SCD

[YL90] A. Yuille and M. Leyton. 3D symmetry-curvature duality theorems. Computer Vision, Graphics, and Image Processing, 52(1):124– 140, October 1990. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yokoya:1994:VSS

[YL94] Naokazu Yokoya and Martin D. Levine. Volumetric shapes of solids of revolution from a single-view range image. Computer Vision, Graphics, and Image Processing. Image Understanding, 59(1):43–52, January 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1003/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1003/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1003/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1003/production/
pdf.

Yoon:2008:SCN

[YL08] Jong-Chul Yoon and In-Kwon Lee. Stable and controllable noise. Graphical Models, 70(5):105–115, September 2008. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070308000052.

Yu:2009:SSS

[YLA09] Tianli Yu, Jiebo Luo, and Narendra Ahuja. Search strategies for shape regularized active contour. Computer Vision and Image Understanding: CVIU, 113(10): 1053–1063, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yu:2012:FSM

[YLL12] Wei Yu, Maoqing Li, and Xin Li. Fragmented skull modeling using heat kernels. Graphical Models, 74 (4):140–151, July 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000185.

Yang:2011:DSE

[YLM11] Peng Yang, Qingshan Liu, and Dimitris Metaxas. Dynamic soft encoded patterns for facial event analysis. Computer Vision and Image Understanding: CVIU, 115(3): 456–465, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yu:1992:IPR

[YLWY92] Bin Yu, Xinggang Lin, Youshou Wu, and Baozong Yuan. Isothetic polygon representation for contours. Computer Vision, Graphics, and Image Processing. Image Understanding, 56(2):264–268, September 1992. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic).

Yalamanchili:1982:EMO

[YMA82] S. Yalamanchili, W. N. Martin, and J. K. Aggarwal. Extraction of moving object descriptions via differencing. Computer Graphics and Image Processing, 18(2):188– 201, February 1982. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yin:2011:HKP

[YNCO11] Shimin Yin, Jin Hee Na, Jin Young Choi, and Songhwai Oh. Hierarchical Kalmanparticle filter with adaptation to motion changes for object tracking. Computer Vision and Image Understanding: CVIU, 115(6):885–900, June 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yao:2011:FHD

[YO11] Jian Yao and Jean-Marc Odobez. Fast human detection from joint appearance and foreground feature subset covariances. Computer Vision and Image Understanding: CVIU, 115(10): 1414–1426, October 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001391.

Youssef:1986:NAO

[You86] Saul Youssef. A new algorithm for object oriented ray tracing. Computer Vision, Graphics, and Image Processing, 34(2):125–137, May 1986. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Young:1981:NIB

[YPVv81] Ian T. Young, Ricardo L. Peverini, Piet W. Verbeek, and Peter J. van Otterloo. New implementation for the binary and Minkowski operators. Computer Graphics and Image Processing, 17(3):189– 210, November 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Yue:2006:VMP

[YR06] Shigang Yue and F. Claire Rind. Visual motion pattern extraction and fusion for collision detection in complex dynamic scenes. Computer Vision and Image Understanding: CVIU, 104(1):48–60, October 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yilmaz:2006:MAP

[YS06] Alper Yilmaz and Mubarak Shah. Matching actions in presence of camera motion. Computer Vision and Image Understanding: CVIU, 104(2–3):221–231, November/December 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yilmaz:2008:DGA

[YS08] Alper Yilmaz and Mubarak Shah. A differential geometric approach to representing the human actions. Computer Vision and Image Understanding: CVIU, 109(3): 335–351, March 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2009:CGM

[YS09] Ruiduo Yang and Sudeep Sarkar. Coupled grouping and matching for sign and gesture recognition. Computer Vision and Image Understanding: CVIU, 113(6):663–681, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yu:2003:MMA

[YSD03] Weichuan Yu, Gerald Sommer, and Kostas Daniilidis. Multiple motion analysis: in spatial or in spectral domain? Computer Vision and Image Understanding: CVIU, 90(2): 129–152, May 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yu:2011:GOE

[YSL11] Chanki Yu, Yongduek Seo, and Sang Wook Lee. Global optimization for estimating a multiple-lobe analytical BRDF. Computer Vision and Image Understanding: CVIU, 115(12):1679–1688, December 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314211001615.

Ye:1999:SPO

[YT99] Yiming Ye and John K. Tsotsos. Sensor planning for 3D object search. Computer Vision and Image Understanding: CVIU, 73 (2):145–168, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0736/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0736/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0736/production/ref.

Yi:2013:TES

[YT13] Chucai Yi and Yingli Tian. Text extraction from scene images by character appearance and structure modeling. Computer Vision and Image Understanding: CVIU, 117 (2):182–194, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001701.

Yanagihara:1983:GRE

[YTTT83] Yoshio Yanagihara, Minoru Tanaka, Shinichi Tamura, and Kokichi Tanaka. Generation of radar echo images from a contour map. Computer Vision, Graphics, and Image Processing, 24 (1):114–128, October 1983. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yuille:1989:ZCL

[Yui89] A. L. Yuille. Zero crossings on lines of curvature. Computer Vision, Graphics, and Image Processing, 45(1):68–87, January 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yu:1999:HMV

[YW99] Hong Heather Yu and Wayne Wolf. A hierarchical multiresolution video shot transition detection scheme. Computer Vision and Image Understanding: CVIU, 75 (1–2):196–213, July/August 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1999.
0773/production;   http:
//www.idealibrary.com/links/
artid/cviu.1999.0773/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1999.
0773/production/ref.

Yemez:2007:VFT

[YW07] Y. Yemez and C. J. Wetherilt. A volumetric fusion technique for surface reconstruction from silhouettes and range data. Computer Vision and Image Understanding: CVIU, 105(1):30–41, January 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2008:USN

[YWMS08] Allen Y. Yang, John Wright, Yi Ma, and S. Shankar Sastry. Unsupervised segmentation of natural images via lossy data compression. Computer Vision and Image Understanding: CVIU, 110(2):212–225, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yan:2011:RRT

[YWZ11] Chao Yan, Yuangqing Wang, and Zhaoyang Zhang. Robust real-time multi-user pupil detection and tracking under various illumination and large-scale head motion. Computer Vision and Image Understanding: CVIU, 115 (8):1223–1238, August 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000890.

Yang:2000:CHM

[YXYW00] Xue Dong Yang, Zhan Xu, Jun Yang, and Tao Wang.

The cluster hair model. Graphical Models, 62(2):85– 103, March 2000. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
1999.0518;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.1999.0518/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
1999.0518/ref.

Yang:1984:AIP

[YY84a] Mark C. K. Yang and Chung-Chun C. Yang. Assessment of the intersection probabilities of random line segments and squares. Computer Vision, Graphics, and Image Processing, 26(3):319–330, June 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yuang:1984:AIP

[YY84b] M. C. K. Yuang and C.- C. Yang. Assessment of the intersection probabilities of random line segments and squares. Computer Vision, Graphics, and Image Processing, 27(??):319– 330, ???? 1984. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yokoya:1989:FBA

[YYF89] Naokazu Yokoya, Kazuhiko Yamamoto, and Noboru Funakubo. Fractal-based analysis and interpolation of 3D natural surface shapes and their application to terrain modeling. Computer Vision, Graphics, and Image Processing, 46(3):284–302, June 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Yamamoto:1996:AFV

[YYL96] Hiroyuki Yamamoto, Yehezkel Yeshurun, and Martin D. Levine. An active foveated vision system: Attentional mechanisms and scan path convergence measures. Computer Vision and Image Understanding: CVIU, 63 (1):50–65, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0004/production;   http:
//www.idealibrary.com/links/
artid/cviu.1996.0004/production/
pdf.

Yeung:1998:SVC

[YYL98] Minerva Yeung, Boon-Lock Yeo, and Bede Liu. Segmentation of video by clustering and graph analysis. Computer Vision and Image Understanding: CVIU, 71(1):94–109, July 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.

0628/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0628/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0628/production/ref.

Yao:2006:HSD

[YZ06] Jian Yao and Zhonefei (Mark)

Zhang. Hierarchical shadow detection for color aerial images. Computer Vision and Image Understanding: CVIU, 102(1):60–69, April 2006. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yan:2013:GSC

[YZT^{+}13] Pingkun Yan, Wuxia Zhang, Baris Turkbey, Peter L. Choyke, and Xuelong Li. Global structure constrained local shape prior estimation for medical image segmentation. Computer Vision and Image Understanding: CVIU, 117(9):1017–1026, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S107731421300074X.

Yang:2011:UCA

[YZY11] Lei Yang, Nanning Zheng, and Jie Yang. A unified context assessing model for object categorization. Computer Vision and Image Understanding: CVIU, 115(3): 310–322, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yang:2010:PCB

[YZZ^{+}10] Yi-Jun Yang, Wei Zeng, Hui Zhang, Jun-Hai Yong, and Jean-Claude Paul. Projection of curves on B-spline surfaces using quadratic reparameterization. Graphical Models, 72 (5):47–59, September 2010. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000123.

Zunic:1998:GCS

[ŽA98] Joviša Žunić and Dragan M. Acketa. A general coding scheme for families of digital curve segments. Graphical Models and Image Processing: GMIP, 60(6):437– 460, November 1998. CO-DEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0482/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1998.0482/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1998.
                                        0482/production/ref.

Zhao:2005:AVM

[ZB05] Liwei Zhao and Norman I. Badler. Acquiring and validating motion qualities from live limb gestures. Graphical Models, 67(1):1–16, January 2005. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Zappella:2013:JES

[ZBLS13] Luca Zappella, Alessio Del Bue, Xavier Lladó, and Joaquim Salvi. Joint estimation of segmentation and structure from motion. Computer Vision and Image Understanding: CVIU, 117 (2):113–129, February 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001312.

Zielke:1993:IEB

[ZBV93] Thomas Zielke, Michael Brauckmann, and Werner Von Seelen. Intensity and edge-based symmetry detection with an application to car-following. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(2):177–190, September 1993. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.
1037/production;   http:
//www.idealibrary.com/links/
artid/ciun.1993.1037/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1993.
1039/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1993.1039/production/
                                        pdf.

Zetzsche:1989:IPR

[ZC89] Christoph Zetzsche and Terry Caelli. Invariant pattern recognition using multiple filter image representations. Computer Vision, Graphics, and Image Processing, 45 (2):251–262, February 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Zlotnick:1993:FRS

[ZC93] A. Zlotnick and P. D. Carnine, Jr. Finding road seeds in aerial images. Computer Vision, Graphics, and Image Processing. Image Understanding, 57(2):243– 260, March 1993. CO-DEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/ciun.1993.
                                        1016/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1993.1016/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1993.
                                        1016/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1993.1016/production/
                                        pdf.

Zordan:2006:BEM

[ZCCD06] Victor B. Zordan, Bhrigu Celly, Bill Chiu, and Paul C. DiLorenzo. Breathe easy: Model and control of human respiration for computer animation. Graphical Models, 68(2):113–132, March 2006. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070305000287.

Zhou:2013:NBS

[ZCF13] Hong Zhou, Yiru Chen, and Rong Feng. A novel background subtraction method based on color invariants. Computer Vision and Image Understanding: CVIU, 117(11):1589–1597, November 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213001380.

Zivkovic:2009:ABM

[ZCK09] Zoran Zivkovic, Ali Taylan Cemgil, and Ben Kröse. Approximate Bayesian methods for kernel-based object tracking. Computer Vision and Image Understanding: CVIU, 113(6):743–749, June 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhu:1999:ICE

[ZCL99] Hui Zhu, Francis H. Y.

Chan, and F. K. Lam. Image contrast enhancement by constrained local histogram equalization. Computer Vision and Image Understanding: CVIU, 73 (2):281–290, February 1999. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.
0723/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0723/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0723/production/ref.

Ziou:2001:DDE

[ZD01] Djemel Ziou and Francois Deschenes. Depth from defocus estimation in spatial domain. Computer Vision and Image Understanding: CVIU, 81(2):143–165, February 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0899;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0899/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0899/ref.

Zhu:2010:KSR

[ZDF10] Youding Zhu, Behzad Dariush, and Kikuo Fujimura. Kinematic self retargeting: a framework for human pose estimation. Computer Vision and Image Understanding: CVIU, 114(12):1362– 1375, December 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2011:SGS

[ZDL^{+}11] Guo-Xin Zhang, Song-Pei Du, Yu-Kun Lai, Tianyun Ni, and Shi-Min Hu. Sketch guided solid texturing. Graphical Models, 73(3):59–73, May 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070310000226.

Zhang:1994:FPC

[ZF94] Zhengyou Y. Zhang and Olivier D. Faugeras. Finding planes and clusters of objects from 3D line segments with application to 3D motion determination. Computer Vision, Graphics, and Image Processing. Image Understanding, 60(3): 267–284, November 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1057/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1057/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1063/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1063/production/
pdf.

Zhang:2008:ISE

[ZFG08] Hui Zhang, Jason E. Fritts, and Sally A. Goldman. Image segmentation evaluation: a survey of unsupervised methods. Computer Vision and Image Understanding: CVIU, 110(2):260–280, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhou:1991:GNB

[ZG91] Xiaohua Zhou and R. Gordon. Generation of noise in binary images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(5):476– 478, September 1991. CO-DEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic). See comment [Liu93].

Zagorchev:2006:PLR

[ZG06] Lyubomir Zagorchev and Ardeshir Goshtasby. A paintbrush laser range scanner. Computer Vision and Image Understanding: CVIU, 101(2):65–86, February 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2010:ACS

[ZG10] Jianguo Zhang and Shaogang Gong. Action categorization by structural probabilistic latent semantic analysis. Computer Vision and Image Understanding: CVIU, 114(8): 857–864, August 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhu:1995:RMR

[ZGK95] Wenwu W. Zhu, Nikolas P.

Galatsanos, and Aggelos K. Katsaggelos. Regularized multichannel restoration using cross-validation. Graphical Models and Image Processing: GMIP, 57(1):38–54, January 1995. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1995.
1005/production;   http:
//www.idealibrary.com/links/
artid/gmip.1995.1005/production/
pdf.

Zeng:2012:DHK

[ZGLG12] Wei Zeng, Ren Guo, Feng Luo, and Xianfeng Gu. Discrete heat kernel determines discrete Riemannian metric. Graphical Models, 74 (4):121–129, July 2012. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000161.

Zhu:2012:ATR

[ZGLP12] Hua Zhu, Shuming Gao, Ming Li, and Wanbin Pan. Adaptive tetrahedral remeshing for modified solid models. Graphical Models, 74(4):76–86, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000124.

Zucker:1979:TLL

[ZH79] Steven W. Zucker and Robert A. Hummel. Toward a lowlevel description of dot clusters: Labeling edge, interior, and noise points. Computer Graphics and Image Processing, 9(3):213– 233, March 1979. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Zhuang:1986:MSE

[ZH86] Xinhua Zhuang and Robert M. Haralick. Morphological structuring element decomposition. Computer Vision, Graphics, and Image Processing, 35(3):370–382, September 1986. CODEN CVG-PDB. ISSN 0734-189X (print), 1557-895X (electronic).

Zhu:2004:LLA

[ZH04] Zhigang Zhu and Allen R. Hanson. LAMP: 3D layered, adaptive-resolution, and multi-perspective panorama— a new scene representation. Computer Vision and Image Understanding: CVIU, 96(3): 294–326, December 2004. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:1997:AOD

[Zha97a] Yuefeng Zhang. Adaptive ordered dither. Graphical Models and Image Processing: GMIP, 59(1):49–53, January 1997. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0414/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0414/production/
pdf;   http://www.idealibrary.
com/links/artid/gmip.1996.
0414/production/ref.

Zhang:1997:TLB

[Zha97b] Zhengyou Zhang. A tighter lower bound on the Spetsakis-Aloimonos trilinear constraints. Computer Vision and Image Understanding: CVIU, 67(2):202–204, August 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1997.
0526/production;   http:
//www.idealibrary.com/links/
artid/cviu.1997.0526/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1997.
0526/production/ref.

Zhang:1999:CBC

[Zha99] Jiwen Zhang. C-Bézier curves and surfaces. Graphical Models and Image Processing: GMIP, 61(1):2–15, January 1999. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1999.
0490/production;   http:

                                        //www.idealibrary.com/links/
                                        artid/gmip.1999.0490/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0490/production/ref;   http:
                                        //www.idealibrary.com/links/
                                        artid/gmip.1999.0492/production;
                                         http://www.idealibrary.
                                        com/links/artid/gmip.1999.
                                        0492/production/pdf.

Zhuang:1988:SLO

[ZHAH88] Xinhua Zhuang, Thomas S. Huang, Narendra Ahuja, and Robert M. Haralick. A simplified linear optic flow-motion algorithm. Computer Vision, Graphics, and Image Processing, 42(3):334–344, June 1988. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Zhang:2011:RME

[ZHM11] Shaoting Zhang, Junzhou Huang, and Dimitris N. Metaxas. Robust mesh editing using Laplacian coordinates. Graphical Models, 73(1):10–19, January 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000160.

Zhuang:1989:SLT

[Zhu89] Xinhua Zhuang. A simplification to linear two-view motion algorithms. Computer Vision, Graphics, and Image Processing, 46(2):175–178, May 1989. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Zucker:1987:OSO

[ZI87] Steven W. Zucker and Lee Iverson. From orientation selection to optical flow. Computer Vision, Graphics, and Image Processing, 37 (2):196–220, February 1987. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Zito:1988:SSH

[Zit88] Richard R. Zito. The shape of SAR histograms (synthetic aperture radar). Computer Vision, Graphics, and Image Processing, 43(3):281– 293, September 1988. CO-DEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Zivkovic:2010:WSC

[Ziv10]  Zoran Zivkovic.  Wireless

smart camera network for real-time human 3D pose reconstruction. Computer Vision and Image Understanding: CVIU, 114(11):1215– 1222, November 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhu:2005:RRT

[ZJ05] Zhiwei Zhu and Qiang Ji. Robust real-time eye detection and tracking under variable lighting conditions and various face orientations. Computer Vision and Image Understanding: CVIU, 98(1): 124–154, April 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zabele:1981:ILD

[ZK81] G. Stephen Zabele and Jack Koplowitz. On improving line detection in noisy images. Computer Graphics and Image Processing, 15(2):130– 135, February 1981. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Zhou:2001:ESE

[ZK01] Lin Zhou and Chandra Kambhamettu. Extending superquadrics with exponent functions: Modeling and reconstruction. Graphical Models, 63(1):1–20, January 2001. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0529;   http://www.
idealibrary.com/links/doi/
10.1006/gmod.2000.0529/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/gmod.
2000.0529/ref.

Zhang:2005:ECU

[ZK05] Jiwen Zhang and Frank L. Krause. Extending cubic uniform B-splines by unified trigonometric and hyperbolic basis. Graphical Models, 67 (2):20, March 2005. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Zhang:2008:ETS

[ZK08] Xinyu Zhang and Young J.

Kim. Efficient texture synthesis using strict Wang Tiles. Graphical Models, 70 (3):43–56, May 2008. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070307000276.

Zhou:2003:PRH

[ZKC03] Shaohua Zhou, Volker Krueger, and Rama Chellappa. Probabilistic recognition of human faces from video. Computer Vision and Image Understanding: CVIU, 91(1–2): 214–245, July/August 2003. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhu:2004:DMC

[ZKRH04] Zhigang Zhu, Deepak R.

Karuppiah, Edward M. Riseman, and Allen R. Hanson. Dynamic mutual calibration and view planning for cooperative mobile robots with panoramic virtual stereo vision. Computer Vision and Image Understanding: CVIU, 95(3):261–286, September 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2001:EFM

[ZL01] Zhengyou Zhang and Charles Loop. Estimating the fundamental matrix by transforming image points in projective space. Computer Vision and Image Understanding: CVIU, 82(2):174–180, May 2001. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0909;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2001.0909/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2001.0909/ref.

Zhang:2013:ESG

[ZLH13] Guo-Xin Zhang, Yu-Kun Lai, and Shi-Min Hu. Efficient synthesis of gradient solid textures. Graphical Models, 75(3):104–117, May 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000744.

Zhou:2013:MSB

[ZLS^{+}13] Huiyu Zhou, Xuelong Li, Gerald Schaefer, M. Emre Celebi, and Paul Miller. Mean shift based gradient vector flow for image segmentation. Computer Vision and Image Understanding: CVIU, 117(9): 1004–1016, September 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314213000751.

Zhan:1994:ZCB

[ZM94] Shiming M. Zhan and Rajiv Mehrotra. A zerocrossing-based optimal threedimensional edge detector. Computer Vision, Graphics, and Image Processing. Image Understanding, 59 (2):242–253, March 1994. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1994.
1016/production;   http:
//www.idealibrary.com/links/
artid/ciun.1994.1016/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1994.
1018/production;   http:
//www.idealibrary.com/links/
artid/cviu.1994.1018/production/
pdf.

Zhao:1996:GTD

[ZM96] ChangSheng Zhao and Roger Mohr. Global three-dimensional surface reconstruction from occluding contours. Computer Vision and Image Understanding: CVIU, 64(1):62–96, July 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1996.
0046/production;   http:
//www.idealibrary.com/links/

                                        artid/cviu.1996.0046/production/
                                        pdf.

Zapater:2005:GAS

[ZMCA05] V. Zapater, L. Martínez-Costa, and G. Ayala. A granulometric analysis of specular microscopy images of human corneal endothelia. Computer Vision and Image Understanding: CVIU, 97(3): 297–314, March 2005. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2008:EPT

[ZN08] Hongsheng Zhang and Shahriar Negahdaripour. Epiflow — a paradigm for tracking stereo correspondences. Computer Vision and Image Understanding: CVIU, 111(3):307– 328, September 2008. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhou:2013:CRU

[ZN13] Qian-Yi Zhou and Ulrich Neumann. Complete residential urban area reconstruction from dense aerial Li-DAR point clouds. Graphical Models, 75(3):118–125, May 2013. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000689.

Zhu:2013:MMO

[ZNG^{+}13] Y. Zhu, N. Nayak, U. Gaur, B. Song, and A. Roy-Chowdhury. Modeling multiobject interactions using “string of feature graphs”. Computer Vision and Image Understanding: CVIU, 117(10):1313–1328, October 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212001646.

Zhao:2000:INS

[ZOMK00] Hong-Kai Zhao, Stanley Osher, Barry Merriman, and Myungjoo Kang. Implicit and nonparametric shape reconstruction from unorganized data using a variational level set method. Computer Vision and Image Understanding: CVIU, 80(3):295–314, December 2000. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0875;   http://www.
idealibrary.com/links/doi/
10.1006/cviu.2000.0875/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/cviu.
2000.0875/ref.

Zafeiriou:2011:EGM

[ZP11] Stefanos Zafeiriou and Maria Petrou. 2.5D Elastic graph matching. Computer Vision and Image Understanding: CVIU, 115(7):1062–1072, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000762.

Zhou:2011:BMP

[ZQ11] Pei Zhou and Wen-Han Qian. Blending multiple parametric normal ringed surfaces using implicit functional splines and auxiliary spheres. Graphical Models, 73(4):87–96, July 2011. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070310000457.

Zhao:2011:GAE

[ZRKZ^{+}11] Haifeng Zhao, Antonio Robles-Kelly, Jun Zhou, Jianfeng Lu, and Jing-Yu Yang. Graph attribute embedding via Riemannian submersion learning. Computer Vision and Image Understanding: CVIU, 115(7):962–975, July 2011. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314211000737.

Zhang:2011:KFS

[ZRL^{+}11] Nan Zhang, Su Ruan, Stéphane Lebonvallet, Qingmin Liao, and Yuemin Zhu. Kernel feature selection to fuse multispectral MRI images for brain tumor segmentation. Computer Vision and Image Understanding: CVIU, 115(2):

256–269, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhao:2009:AGC

[ZS09] Liming Zhao and Alla Safonova. Achieving good connectivity in motion graphs. Graphical Models, 71(4):139– 152, July 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000174.

Zheng:2011:ART

[ZS11] Liying Zheng and Daming Shi. Advanced Radon transform using generalized interpolated Fourier method for straight line detection. Computer Vision and Image Understanding: CVIU, 115(2): 152–160, February 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zeng:2008:TCN

[ZSCP08] Yun Zeng, Dimitris Samaras, Wei Chen, and Qunsheng Peng. Topology cuts: a novel min-cut/max-flow algorithm for topology preserving segmentation in N–D images. Computer Vision and Image Understanding: CVIU, 112(1):81–90, October 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:1996:BRD

[ZSN96] Bing Zhang, Mehdi N. Shirazi, and Hideki Noda. Blind restoration of degraded binary Markov random field images. Graphical Models and Image Processing: GMIP, 58(1):90–98, January 1996. CODEN GMIPF4. ISSN 1077-3169 (print), 1090-2481 (electronic). URL

http://www.idealibrary.
com/links/artid/gmip.1996.
0007/production;   http:
//www.idealibrary.com/links/
artid/gmip.1996.0007/production/
pdf.

Zucker:1980:FSC

[ZT80] Steven W. Zucker and Demetri Terzopoulos. Finding structure in co-occurrence matrices for texture analysis. Computer Graphics and Image Processing, 12(3):286– 308, March 1980. CO-DEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic).

Zheng:1998:GDP

[ZT98] Jiang Yu Zheng and Saburo Tsuji. Generating dynamic projection images for scene representation and understanding. Computer Vision and Image Understanding: CVIU, 72(3):237–256, December 1998. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

http://www.idealibrary.
com/links/artid/cviu.1998.

0678/production;   http:
//www.idealibrary.com/links/
artid/cviu.1998.0678/production/
pdf;   http://www.idealibrary.
com/links/artid/cviu.1998.
0678/production/ref.

Zhao:2009:MOR

[ZT09] Qi Zhao and Hai Tao. A motion observable representation using color correlogram and its applications to tracking. Computer Vision and Image Understanding: CVIU, 113 (2):273–290, February 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2011:MSS

[ZTH^{+}11] Shiliang Zhang, Qi Tian, Gang Hua, Wengang Zhou, Qingming Huang, Houqiang Li, and Wen Gao. Modeling spatial and semantic cues for large-scale near-duplicated image retrieval. Computer Vision and Image Understanding: CVIU, 115(3):403– 414, March 2011. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:1996:P

[ZTS96] Ruo Zhang, Ping-Sing Tsai, and Mubarak Shah. Photomotion. Computer Vision and Image Understanding: CVIU, 63(2):221–231, March 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0016/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0016/production/
                                        pdf.

Zhuge:2009:ISS

[ZU09] Ying Zhuge and Jayaram K. Udupa. Intensity standardization simplifies brain MR image segmentation. Computer Vision and Image Understanding: CVIU, 113(10): 1095–1103, October 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zucker:1976:RGC

[Zuc76a] Steven W. Zucker. Region growing: Childhood and adolescence. Computer Graphics and Image Processing, 5 (3):382–399, September 1976. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). ZUCKER76b.

Zucker:1976:TMT

[Zuc76b] Steven W. Zucker. Toward a model of texture. Computer Graphics and Image Processing, 5(2):190–202, June 1976. CODEN CGIPBG. ISSN 0146-664X (print), 1557-9697 (electronic). ZUCKER76a.

Zucker:1985:EOS

[Zuc85] Steven W. Zucker. Early orientation selection: Tangent fields and the dimensionality of their support.

Computer Vision, Graphics, and Image Processing, 32 (1):74–103, October 1985. CODEN CVGPDB. ISSN 0734-189X (print), 1557-895X (electronic).

Zunic:2003:DTC

[Žun03] Joviša Žunić. On discrete triangles characterization. Computer Vision and Image Understanding: CVIU, 90(2): 169–189, May 2003. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhuge:2006:VSB

[ZUS06] Ying Zhuge, Jayaram K.

Udupa, and Punam K. Saha. Vectorial scale-based fuzzyconnected image segmentation. Computer Vision and Image Understanding: CVIU, 101(3):177–193, March 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:1993:PMC

[ZW93] Guanghua H. Zhang and Andrew Wallace. Physical modeling and combination of range and intensity edge data. Computer Vision, Graphics, and Image Processing. Image Understanding, 58(2):191–220, September 1993. CODEN CIUNEJ. ISSN 1049-9660 (print), 1557-7635 (electronic). URL

http://www.idealibrary.
com/links/artid/ciun.1993.

                                        1038/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/ciun.1993.1038/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1993.
                                        1040/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1993.1040/production/
                                        pdf.

Zabrodsky:1997:UBS

[ZW97] Hagit Zabrodsky and Daphna Weinshall. Using bilateral symmetry to improve 3D reconstruction from image sequences. Computer Vision and Image Understanding: CVIU, 67(1):48–57, July 1997. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL

                                        http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0506/production;   http:
                                        //www.idealibrary.com/links/
                                        artid/cviu.1996.0506/production/
                                        pdf;   http://www.idealibrary.
                                        com/links/artid/cviu.1996.
                                        0506/production/ref.

Zheng:2003:PBC

[ZW03] Jianmin Zheng and Guozhao Wang. Perturbing Bézier coefficients for best constrained degree reduction in the L_{2}- norm. Graphical Models, 65 (6):351–368, November 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Zhu:2002:RTA

[ZXK02] Yuanxin Zhu, Guangyou Xu, and David J. Kriegman. A real-time approach to the spotting, representation, and recognition of hand gestures for human–computer interaction. Computer Vision and Image Understanding: CVIU, 85(3):189–208, March 2002. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2012:ERL

[ZXY^{+}12] Long Zhang, Jiazhi Xia, Xiang Ying, Ying He, Wolfgang Mueller-Wittig, and Hock-Soon Seah. Efficient and robust 3D line drawings using difference-of-Gaussian. Graphical Models, 74(4):87–98, July 2012. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000136.

Zheng:2009:SB

[ZYP09] Wen Zheng, Jun-Hai Yong, and Jean-Claude Paul. Simulation of bubbles. Graphical Models, 71(6):229–239, November 2009. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070309000253.

Zhou:2009:OTU

[ZYS09] Huiyu Zhou, Yuan Yuan, and Chunmei Shi. Object tracking using SIFT features and mean shift. Computer Vision and Image Understanding: CVIU, 113(3):345–352, March 2009. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zheng:2010:DOB

[ZYT10] Songfeng Zheng, Alan Yuille, and Zhuowen Tu. Detecting object boundaries using low-, mid-, and high-level information. Computer Vision and Image Understanding: CVIU, 114(10):1055– 1067, October 2010. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2013:RIR

[ZYXZ13] Jun Zhang, Lei Ye, Yang Xiang, and Wanlei Zhou. Robust image retrieval with hidden classes. Computer Vision and Image Understanding: CVIU, 117(6):670–679, June 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000520.

Zhang:2007:JDZ

[ZZ07] Lei Zhang and David Zhang. A joint demosaicking–zooming scheme for single chip digital color cameras. Computer Vision and Image Understanding: CVIU, 107(1– 2):14–25, July/August 2007. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhao:2010:SAL

[ZZ10] Keke Zhao and Zhenyue Zhang. Successively alternate least square for lowrank matrix factorization with bounded missing data. Computer Vision and Image Understanding: CVIU, 114(10): 1084–1096, October 2010. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhao:1992:PER

[ZZAA92] Yunxin Zhao, Xinhua Zhuang, Les Atlas, and Lars Anderson. Parameter estimation and restoration of noisy images using Gibbs distributions in hidden Markov models. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 54(3):187–197, May 1992. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Zhang:2013:ASA

[ZZC^{+}13] Shaoting Zhang, Yiqiang Zhan, Xinyi Cui, Mingchen Gao, Junzhou Huang, and Dimitris Metaxas. 3D anatomical shape atlas construction using mesh quality preserved deformable models. Computer Vision and Image Understanding: CVIU, 117(9):1061–1071, September 2013. CODEN CVIUF4.

ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314213000799.

Zhang:2013:CRI

[ZZL13] Jun Zhang, Heng Zhao, and Jimin Liang. Continuous rotation invariant local descriptors for texton dictionarybased texture classification. Computer Vision and Image Understanding: CVIU, 117(1):56–75, January 2013. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1077314212001348.

Zhang:2013:LBS

[ZZLZ13] Yu-Wei Zhang, Yi-Qi Zhou, Xue-Lin Li, and Li-Li Zhang. Line-based sunken relief generation from a 3D mesh. Graphical Models, 75(6):297– 304, November 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070313000222.

Zhou:2012:HOS

[ZZP12] Bingyin Zhou, Fan Zhang, and Lizhong Peng. Higherorder SVD analysis for crowd density estimation. Computer Vision and Image Understanding: CVIU, 116(9): 1014–1021, September 2012. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1077314212000884.

Zhang:2006:AMI

[ZZZ06] Yongjun Zhang, Zuxun Zhang, and Jianqing Zhang. Automatic measurement of industrial sheetmetal parts with CAD data and non-metric image sequence. Computer Vision and Image Understanding: CVIU, 102(1):52–59, April 2006. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zeng:2013:OBF

[ZZZL13] Ming Zeng, Fukai Zhao, Jiaxiang Zheng, and Xinguo Liu. Octree-based fusion for realtime 3D reconstruction. Graphical Models, 75 (3):126–136, May 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1524070312000768.

Zhao:2009:OSP

[ZZZP09] Y. Zhao, L. Zhang, D. Zhang, and Q. Pan. Object separation by polarimetric and spectral imagery fusion. Computer Vision and Image Understanding: CVIU, 113(8): 855–866, August 2009. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Zhang:2013:RTB

[ZZZY13] Yu-Wei Zhang, Yi-Qi Zhou, Xiao-Feng Zhao, and Gang Yu. Real-time bas-relief generation from a 3D mesh. Graphical Models, 75(1):2– 9, January 2013. CO-DEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070312000719.